A Complete Bibliography of Publications in Journal of Computational Physics: 2015–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

18 January 2025 Version 1.86

Title word cross-reference

(2 + 1) [PS17]. (d^{2}/dx^{2} − h^{2}) [Kas15]. (r, v_{r}, v_{θ}) [VSC18]. (S_{N}) [OWKE16]. 0 [TCS16a]. 1 [ALTR17, BK17c, CGMH18, CSK^{+}16, EL18, JKE^{+}17, LLA19, LMG19, NMM19, Nor15, PCMC19, VLP^{+}16, XQ17, YC17, ZGJ16]. 13 [SGP17b, WT19]. 2 [AMXJ19, BK17c, BT16, BC16c, CGK17, CDL19, CZ17, CSK^{+}16, FL18, FNGV18, FNGDMNR18, GFL17, GMWC19, GHL^{+}16, Hu17, IG15, KQB18, LMPS15, LM15a, LGB17, LZT^{+}15, LY16a, LGD17, LZL^{+}19, Mue18, NMM18, NMM19, PG17, PKJ^{+}18, PMGW16, PAFT19, QDH15, RLP16, SCS16, TCS16a, WY17, WHRL19, XDSX17, YFJ18, ZND16, ZJ18, ZZW^{+}16]. 3 [AG16, AHHC18, AMXJ19, ACS16, BHZ16, BGV17, BDK^{+}17, BS15a, BSM16, BC16c, BCRS19, CSS17, CDLR19, CBC^{+}18, CDL17, CGL18, CCZC16, CZ16, CX16, Cho19, CSK^{+}16, DBD^{+}17, DS15a, DMRB19, DWGW16, DF16, Dod17, DD16b, EDW19, FDS^{+}15, FGLB16, FC19a, FYC^{+}18, GBM16, GWC18, HWH^{+}16, HLCL19, JBLO15, KE15, KES18, KC17c, KFWK17, LFRH17, LML^{+}16, LHMB16, LZ17a, LLJJ18, LZHM19, LFAR19, LMC19, MKYZ17, MG15b, MC15, MBS19, MF16a, MW17b, MFF^{+}19, MSF^{+}19, Noe15, PGCG18, PK17, PR16a, PAFT19, PTT18, RBY19, SNSG16, SFT16, ST18b, SA15, ST18c, STV19, Sto16, SSL^{+}16b, TCD17, TB19, TRL15, TCL15, VLP^{+}16, WXW15, WSH^{+}17, YDLC19, Yam19, YSWS16, YFJ18, YPC19, YTW15, YXD^{+}16, YPK16, YT19, ZBH^{+}18, ZZZ17, ZHS18, ZSL^{+}19, ZVO15, ZYCK15, dBIM16, dJRP^{+}15]. 4 [CSC19, MCHL16]. ^{2} [RKO^{+}17b]. A [SLH18]. α [DW19]. β [CS16a, OLHD17]. C [KD17b]. C^{0} [XJG18]. C^{1} [KS16a]. D [RRS19b, RRS19a]. δf [SP16c, SPCH16, BPL19]. ℓ_{1} [JES15]. f [Ido16, KYPK15]. h [BCB17, JKE^{+}17, TXKvdV16]. hp [MSP15, HEPG15, HZ15, SL19c]. K [ZZT^{+}16, HY16, KL16, KFL17, LHS^{+}18, LZB^{+}17, LHMB18, PBC^{+}17]. L(α) [PKK18]. l^{1} [LT17c]. L^{2} [FSB16]. l_{1} [GNZ18, PHD16]. LU [RTV17]. R [dlHC16]. R^{3} [GHV19]. H [CDC17]. μ(I) [FNGDMNR18, FC19a]. N [CBZ18, Don15b, Don17, Don18, SLL19, YD18, PF15]. n^{th} [FYZ^{+}15]. O(N) [CV16b, XL17b]. O(N log N) [CC19b]. ω [LZB^{+}17, LHMB18]. p [BST^{+}18, BG19b, FSWW17, GHS19, HZ15, HSF19, LL19a, NdlLPCC19, RRMF^{+}19, SS17b, TABR17]. P^{2} [WNW^{+}19]. P_{0} [WYZZ18]. P_{1} [WYZZ18]. P_{N} [Her16, LMH16, ZM16b]. P_{n} − P_{n−1}^{DG} [Mel18]. P_{N}P_{M} [BB19]. Π [HAPK15]. P N [HE15]. r [BRW15, DTA^{+}15, VW16]. R^{2} [CHCC18]. S [PKLS17]. S_{N} [HR18a, SWG^{+}17]. t [GWE^{+}15]. τ [Ani16]. θ [GCVCHH18, KTK15]. φ [DTA^{+}15]. X [WHEK18]. y′ = F (y, t)y [JFS17]. z [DTA^{+}15, VW16].

-adaptation [RN18a]. -Adaptive [HEPG15, BST^{+}18, BRW15, NdlLPCC19, TXKvdV16, TABR17]. -adaptivity [JKE^{+}17]. -boxes [SS17b]. -continuous [KS16a]. -coupled [CBZ18]. -criterion [KTK15]. -curl [LL19a]. -D [AMXJ19, MSF^{+}19, Sto16, YDLC19, CCZC16, CSC19, FGLB16, JKE^{+}17, JBLO15, LM15a, LZT^{+}15, LMG19, Nor15, PCMC19, WSH^{+}17, YC17, ZSL^{+}19]. -eigenvalue [KL16, KFL17]. -exact [HY16, PBC^{+}17]. -FEM [BG19b]. -forward [SL19c]. -grids [KD17b]. -gyrokinetic [BPL19]. -isothermal [TXKvdV15]. -Laplacian [FSWW17]. -matrix [CDC17]. -method [GCVCHH18]. -micron [LLA19]. -minimization [GNZ18, JES15, PHD16]. -moment [WT19]. -multigrid [BCB17, HSF19, RRMF^{+}19]. -order [FYZ^{+}15]. -parameters [PKLS17]. -periodic [ZHS18]. -point [WHEK18]. -projection [FSB16]. -rheology [FC19a]. -second [LLA19]. -stable [PKK18, SLH18]. -stage [CSS17]. -symplectic [ZZT^{+}16]. -system [GHS19]. -T [AMXJ19, DS15a]. -th [LHS^{+}18]. -version [HZ15]. -version/ [HZ15].

0012 [FW17].

1 [Ano19a, Ano19-29, MRRRF18, RRS19a]. 1- [RRS19a]. 1-772 [Ano19-28]. 1-904 [Ano19-29]. 148 [CNG17]. 15 [Ano19b, Ano19c, Ano19-28]. 199 [MN17].

2 [FNGV18, RRS19b]. 2- [RRS19b]. 2.0 [LLD19]. 2019 [Ano19a, Ano19b, Ano19c, Ano19-29, Ano19-28]. 229 [Dav15]. 230 [Gho17, HSK^{+}15]. 231 [TK15b]. 259 [PS15a]. 262 [EH15, XS15]. 265 [SYV17]. 270 [ZJS15]. 294 [GBCF16]. 299 [BR16]. 2nd [Cac15a, Cac15b]. 2nd-ASAM [Cac15a, Cac15b].

307 [KYW^{+}18]. 343 [STEK22]. 344 [NG18, SWMD17a]. 348 [HGN17a]. 361 [DK18a]. 367 [DvW19]. 372 [GRT21]. 374 [ABG^{+}19b, Hag25]. 375 [KTK19]. 376 [BLL20]. 381 [SZN20]. 388 [EFO20]. 390 [SYOS21]. 395 [Pan20].

4U [HAPK15].

A-SLEIPNNIR [PC19]. ABCD [PD15]. ABCs [TT19]. ability [KKZ15]. able [BDJP19]. above [GP16b, YS18b]. Absorbing [Pin15, SK15a, AMP16, BG19a, GGT15, KCHW19, LH16, LMC19, MNW19, SJH^{+}15, VAD17]. absorption [DZR18, DCA^{+}16, DJV^{+}18, WWRS17]. abyss [OY19]. accelerate [BT17a, Lan19, PKW17]. Accelerated [CMR^{+}16, AC17, BGHK19, CWM^{+}16, CG18a, CRZ17, GKE15, HPY18, HPC19, JTD16, KCSW19, KH18, LO19, PCMC19, PTMF18, RGW16, SMAG17, VWV17, WWGW18, WL16, YZW^{+}18, ZMCC18]. Accelerating [Gen15, LZSG19, MN18a, SLR^{+}16, WNW^{+}19, XZZ15]. Acceleration [BPF^{+}16, HLQ16, KJHA19, PT18, SAOW17, WB16, AJW17, BBBG15, BM19b, CLGA17, DMAM15, Fon16, Mas18, PSP16, SB18, SWG^{+}17, TWM18, YXD^{+}16, ZNS19, MAM16]. accelerator [VMN^{+}18, TC15a, TKC15]. accelerators [SR18]. acceptance [PDS15]. account [LYDB17, SSL^{+}16a]. accounting [EDW19, Noe15]. accumulation [HMBH15]. Accuracy [CNG99, CSB15, GDS^{+}16, GR15, Nis15, NL17, Pei16, BK17a, CBB16, CNG17, COdLL18, CSK^{+}16, DBZ17, Fal17, FFBB16, FYO^{+}15, GH17a, GO16, KDF15, KGS17, LS16a, MDMS18, MP19, MA16, MSH^{+}15, NJL19, OMLdL16, SM16, Sla16, TLR16, WZ15, WKPS18, ZS18, ZS19a]. Accuracy-preserving [Nis15, NL17]. Accurate [CCZC16, CLvS17, EMZ16, FYC^{+}18, GLMC16, HM16b, IM15, JG19, RCRF16, SR19, SWLW19, WCL15, AD15, AASRT17, ALO18, ABH^{+}19, ABG18c, ANL^{+}16, BLM18, BZ16a, BXY17, BOA17, Bat17, BH18, BST15, BDZ15, CSW^{+}19, CKT17, CM18b, CYS17, CD17, Cho19, CLV19, CG16, CLMZ17, CC19c, DS15b, DVP^{+}16, DY17, DL18b, DL18c, DL16, DvWZ18, EMM^{+}18, FFM19, FS18, GPS17a, GPS17b, HK19a, JL16, KTN15, KB19, LS19a, lLLNS16, LY16c, LTWZ18, LDHJ15, LPR19, MA17, MH19, MBHS17, MPFL16, MDP18, NMM19, OLDN17, OS15a, OSKN18, OT15, OV17, PXLL16, PLB18, PMB18, RT16, RJ19b, RSD17, STHW17, Say17a, Say17b, SUR18, STG17, SCS16, ST18c, TD18, TST^{+}15, WR15, WT15, WA18, XX17, Yan17, ZMF15, ZL15b, ZGD^{+}16]. accurately [LYL19]. ACDS [CCDL19]. achieving [MN16c]. Acoustic [AN15, BBN18, Gib18, APKP16, BHJ18, BG19b, BLS16, CCFC19, CGL18, CDDL19, CCZ15, Cho19, CLQ17, DBD^{+}17, DSS18, GFC18, GHH^{+}16, HLL^{+}16, IML15, KZR15, LWWY18, LB16, LZS^{+}19, LMM17, MBJ16, MBNJ16, MKYZ17, MDW18, MSS16, MBD19, MNW19, PA19, PGM17, PWC18b, RZ17, RRD16, SZW^{+}16, TP17, TRLK18, TLB^{+}18, VAD17, WLW^{+}18, WAZ19, Yam19, ZGD^{+}16, bWAW15, dFGS^{+}17, tEDKT17]. acoustic-convective [tEDKT17]. acoustic-structure [LZS^{+}19]. acoustic-transport [PGM17]. acoustic-wave [GFC18]. acoustical [LMC19]. acoustically [DXvW18, DvW19]. acoustically-conservative [DXvW18, DvW19]. acoustics [BDJP19, SK15a, ZR17, dCMR19]. across [KF17, LMM17]. action [WY17]. activation [VLP^{+}16]. Active [ZLC^{+}18, BDMZ19, CDX18b, CELI15, DCP15, DKPC15, KBG^{+}15, RBJS15, TS19, RC18]. active-strain [LHY^{+}19]. actuated [BBMN18]. actuator [GPAO^{+}18]. Adams [ZM16a]. Adaptation [KRFV16, ALO18, BOA17, BAD19, BD16, CGL18, DLK17, FBG15, GCVMK15, GSN17, HK15b, KLA17, PES19, RN18a, RRMF^{+}19, SW15, TST17, TVB^{+}16, TG17, THS^{+}19, WBBC16, WCT18]. adaptations [VLAB18]. adapted [BWR15, BOD19, DBMB15, Rag15, RHvR^{+}15, SL17, dCMR19]. Adapting [BHdD18]. Adaptive [ABP^{+}16, ASWvD19, BV15, BDM17, BHP19, BS15b, CWB^{+}19, CEH16, FW18, FPASS16, GBvZB16, GK19b, HEPG15, HXB15, Lan19, LW17a, LH17b, LNM15, MG17, MBBKTH17, PS16, PR17b, QB16, RKO^{+}17b, SZN19, SZN20, SS15a, SPD19, SL18, TCA16, TWF19, TWH15, VN15, YZ19, YYJ^{+}19, ZAK15, ZLH^{+}17, ZH15, AMS17, AWS16, APLK19, AC17, BGS16, BHL15, BST^{+}18, BHS^{+}18, BST15, BRW15, CPV16, CPdS19, Cen19, CC17a, CQ15, CS16c, CTG16, CYYL18, CJC19, CYWL17, DS16, DMS17, DS15d, EL17, FOF15, FGLW18, FBM16, FLHA17, FHA18, FC16, GB15a, GRT18, GRT21, GTG15, HD18, HS17b, HS18a, HIN^{+}16, HHLY17, HY16, HYL17, Hu17, HXX18, HW16c, HC18b, IGQ15, JW15a, JJ18a, KCW17, KG15, KDPK15, KC18, LS16a, LL15, LZB^{+}17, LHMB18, LHLL19, LZT17, LL19b, LH18, LM19d, LDSM19, LWJV19, MBSS15, MNG15a]. adaptive [Mar19, MBM^{+}18, MMSS15, MGBG16, MW17b, MDAB18, MSB^{+}16, MNW19, MM16d, MM18, NdlLPCC19, NVBDV15, NBH18, PR19, Pop15, PVB17, PC19, QLF16, RDG17, RDM15, RS15b, SRBÓ17, SwS16, SWHV16, SC16, SLdTV18, SD18, TCS16a, TMWF18, TDC^{+}19, TXKvdV16, TL15, TSB^{+}18, TABR17, URL16, WDG^{+}17, WDS15, WBM15a, XDLX19, XTYL18, XWZ^{+}18, YFC19, dlAC17, BB19, BDV17, COdLL18, PSB^{+}18, WSJY16]. Adaptive-Mesh-Refinement [SL18]. Adaptive-Order [BB19]. adaptively [HC18a, TR17]. adaptivity [APP^{+}16, AKM^{+}19, BHZ16, BG19b, CPP19, DBSS^{+}19, GBD^{+}15, JKE^{+}17, OKWE17, WKOE17]. added [BHKS16, SBHS19]. added-mass [BHKS16, SBHS19]. adding [DL18a]. additional [Abg18a]. additive [ADE^{+}17, CHZ16]. address [AM17b]. addressed [CSCM16]. ADER [BK16a, BLD15, BDLM18, BTVC16, CTM^{+}16, DPRZ16, DPRZ17, Jac17b, JC17, NMM15, NMM16, NMM18, NMJFM19, Nor15]. ADER-MOOD [BDLM18]. ADER-type [BTVC16]. adherence [RBK19]. adhesion [ISST18]. adhesive [FRL15]. ADI [BC16c, FBF15]. Adiabatic [DRZ^{+}19, BLVC16]. adjacent [GMP16, ZYK18]. Adjoint [AMJ17, Blo17, Mag19, NT19, RPC^{+}18, SW15, Cac15a, Cac15b, CYYL18, DK18a, DK18b, HL15a, JW15a, KPKGH19, KPKG15, LYPP17, Loz17, MMMS15, SSC^{+}16, Stü15, VBF15, WWZ19, XRMM15, ZP16]. Adjoint-based [AMJ17, RPC^{+}18, SW15, CYYL18, JW15a, Loz17, MMMS15, SSC^{+}16]. adjoints [Fid17]. adjustable [HWA19]. adjusted [CW19]. adjustment [APT17, OSP17]. ADM [CvKH16, HCVH18, Ani16]. ADM- [Ani16]. admissibility [BT16]. ADOO [FS19a]. adsorption [ZQCT15]. Advanced [TK16, TM17, WB17, KH15, KPJ18, SSL^{+}16b, Zoh17]. advances [PC16]. advancing [AW16, CCdL15, ZJ18]. advected [HM17]. advecting [PR16b]. Advection [KL19, AAL15, APP^{+}16, BCS19, BFT17, BTVC16, CKK18b, CSH15, DJL^{+}19, EHXM15, FNNW19, GS15a, IM17b, JZ16, LN17, LLP^{+}16, LE16, LPB17, LZ17b, LLH19, LLLN18, MD18, MS18b, MK15, MN16a, MMvR18, MBD19, MSP16, MN16c, NL18a, O’S19, PCF15, PPCK17, QDRB15, RZOZ19, SF18b, SP15b, TSH17, TAR17, Vab18, Wil19, YWHP15, ZJL16]. advection-diffusion [BFT17, CKK18b, GS15a, LE16, LZ17b, LLLN18, MS18b, MK15, MN16a, MMvR18, MSP16, NL18a, TSH17]. advection-diffusion-reaction [BTVC16, DJL^{+}19, JZ16]. advection-dispersion [PCF15, PPCK17]. advective [AJVH17, BHdD18]. advective-diffusive [AJVH17]. Adversarial [SHTY19, YP19]. aeroacoustic [BBK19, CFSZ19, ZZH16]. aeroacoustics [BCD^{+}15, JC17, PCBG18, SWS17]. aerodynamic [CF19, GGW17, Loz17, TZ16]. aerodynamics [SPP^{+}16a, TVB^{+}16]. aeroelastic [LHY17, MM17, SPP^{+}16a]. aerofoil [KH15]. aerosol [CMR^{+}16, FSK^{+}16, FL16, SNB^{+}15]. aerosols [SNB^{+}15]. aerothermal [ED16]. AETHER [TC15d]. affect [VW18]. affine [JST17]. affine-particle-in-cell [JST17]. Affordable [sCYxL^{+}18, WG16b]. against [sCYxL^{+}18]. age [LDWZ15]. agglomeration [BCB17]. aggregate [GPG17, LMY^{+}19, MGPG19]. aggregation [XR17]. aging [SAH17]. air [CHE^{+}17, DBD^{+}17]. aircraft [KYUO15]. airflows [WAF^{+}19]. airfoil [FW17]. airfoils [CPS17]. al. [CFO18, YM15]. ALE [BB19, BQCG17, BMC^{+}18b, CGP16, DG16a, DLM18, FRW16, Liu16, LMZ19, MCL19, OMLdL16, RDG17, RXS16, ZS16, ZFL^{+}19, ZCY^{+}19]. Algebraic [CvKH16, CFvKH18, HCVH18, TAH16, ANL^{+}16, BST^{+}18, HHR15, RWG18, TWH15, WS16]. Algorithm [CDOY19, MDH19, ABN15, AM19, ALM15, AA15, BSK15, BHKS16, BHST17a, BHST17b, BHST18, BDBEE15, BCM15a, BLK19, Bre18, BZ16b, BKG15, CPV16, CM16a, CC16a, CS16a, CC19a, CC17a, Cha16, CQ15, CZJ17, CM19, CRMP16, CY19b, CWJ18, DFM17, DSX19, DXvW18, DvW19, Don15b, Don18, DJL^{+}19, DHC16, EAAM15, ETAG15, EBQ15, FHY^{+}19, FL18, GHV19, GLZ16, Gen11, Gho17, GJ15, Gro18, HSK^{+}15, HTZG17, HPV16, HLS19, HYL17, HHK15, KBK15a, KC17b, KKLS17, KF17, KK16, KL18b, KJ17b, KJ18, LM15b, LL18, LL15, LMC16, LHB^{+}16, LY17, MS15a, MBHS17, MBD19, MA16, MKV^{+}17, NOM^{+}17, NKN^{+}17, NSB15, NLL^{+}15, PVFN15, PSB^{+}18, PKLS17, PN17, PLWJ16, Ram17, RYZ18, RC18, RZZ19, RLP16, RL18, RL17, ST16, SLL19, DD17c, SPD19, SWS^{+}18, SKS17, SBHS19, SP19a, SS18b, SPM^{+}15]. algorithm [SW17b, SCC19, SO15, SR18, TCSM15, TPTT18, TH16, VYP15, VBG^{+}17a, WK18, WK19, WS16, WKOE17, WS15a, Wu16, XL17a, XDSX17, XZZ15, XL17b, YJ17, YD18, YM19, YLLH19, Zad11, ZMF15, Zau16, ZZDB15]. Algorithms [SK19b, ZzSK15, AN15, BSN19, BK18, BSP18, CJWS19, CWF16, CCK^{+}17, CQL^{+}17, CM18b, DNBH15, DPW^{+}15, DLC15, DFS16, Dom18, EG18a, FAY19, HBC^{+}16, HGR16, HS17b, HSLQ15, IPSG15, JHT^{+}18, LL16a, LSLA16, LLL16, LS19a, LKSM17, Loh17, LT17b, MGCW18, NSK^{+}16, PT17a, PWP15, PT17b, RKO17a, RKL18, RRL19, RS15b, RO19, SSDN15, SBT17, Sla16, TASA19, WZ18b, ZPE^{+}16, ZRE16]. aligned [KKLS17]. all-hex [RGW16]. all-Mach [BQRX19, FP18]. all-regime [CGK17]. all-scale [SDH^{+}16, SSX16, SKW19]. all-speed [AIP17, WNDB19]. Allen [JJ18a, KJYC17, WX17, YLLH19]. allowing [CSCM16]. alloy [AZK16, BGJ^{+}15, MTL^{+}17, RTO15]. alloys [DMS17, OTS17]. almost [VK15, BPTA16]. Alternating [LP16b, LZT^{+}15, SS16b, SZ17]. Alternative [BVG^{+}16, MG15b, Pei16, PSP16, SG19b, WLGD18]. AMP [SBHS19]. Ampére [DL17, CCZ15, TC15a, TKC15, WBBC16]. Amplitude [GHL15, GHL^{+}16]. AMR [DWGW16, PSB^{+}18, DD17c, ZFL^{+}19]. AMS [TAH16]. analogue [BN17]. Analyses [YM15]. Analysis [AJP15, ADOP18, BG16b, CYYL18, COdLL18, Cif19, CL19b, GK18, HTMP17, JL16, LZL^{+}17, UWH17, YCBC15, ADG19, AMJ17, AW18, AM19, AA15, ACJ17, Ani19, ADP^{+}17, ARTG^{+}19, BK17b, BK19a, BHST17a, BHS^{+}18, BM16, Blo17, BW18a, BGM16, BTVC16, Cac15a, Cac15b, CKK18a, CBC^{+}18, CW16, CNW17, CNOS15, CC17c, CSLL15, CV18, CwYjS16, DC18a, DDM^{+}19, DDV18, DK18a, DK18b, DMSC16, DMM19, EL18, EL19, GB15b, GMS19, HD15, HW16b, JRPPS18, JPLL15, JSB^{+}19, KS16a, KMdB16, KSV^{+}15, KFL17, KD17b, Kri17, Lap16, LSM19, LT15, LDL^{+}16, LLJJ18, LWZ19, Lia16, LZS^{+}19, DV17, LBB^{+}17, MDH19, MBJ16, MBNJ16, MH19, Mel18, MDMS18, MHGM^{+}15, MDDM17, MMP18, MTL^{+}17, MF16b, MSP15, MSP16, NJL19, NW17, NT19, NWFT19, NF17, OWKE16, PXXZ15, PRXC19, Par17, Pei16, PZF16, QS16]. analysis [RWKW16, SZM19b, SVG18, SUR18, SP19a, SSM^{+}17, SW17b, SPB16, TCA16, TST^{+}15, URGT18, VPM15, VPV^{+}17, WHCN17, WSB19, XB18, YJB18, ZMF15, ZOG19, ZZW^{+}16, ZBZ^{+}18, dLDG^{+}18]. analytic [KCHW19, LGB17]. Analytical [AHHC18, QWXZ17, SWML17, AB17, ALTR17, CZ16, DF16, DH18a, EAAM15, FKF17, GT19, LC18, LC17a, LHGF19, MD15, MHT^{+}19, MTD15, TM15b]. analytical-stochastic [DH18a]. analytically [RRM^{+}16]. analyze [EDW19, UG16]. Analyzing [RJ19a, VMK^{+}19]. Anatomically [ANL^{+}16]. anchored [MS16b]. anchored-ANOVA [MS16b]. anchoring [AHHC18]. Anderson [AJW17, PSP16]. anelastic [SHLG15]. aneurysms [YPK16]. angiogenesis [BCC^{+}18]. angle [CDL19, CHE^{+}17, Don17, Gan15, Hig17, KL15, TSR15]. angles [BFP18, HKS^{+}16]. Angular [DBSS^{+}19, DL15, ABP^{+}16, BCG^{+}15, GBD^{+}15, JST17, KL15, MFG15, OWKE16, ZM16b]. animals [PBP18]. Anisotropic [BG19b, BD16, BDV17, CNK19, CS18a, DPK17, ALO18, Azi19, BJWZ17, BOA17, BAD19, CPV16, CPdS19, CGL18, CLS^{+}18, CSG17, CHL^{+}19, Chu17, EH18, FBG15, GMP16, GFG^{+}15, GH17a, HHA16, MDT16, PS15b, PC19, RMA17, RN18a, RRMF^{+}19, SAEF17, SS17c, SDW18, TW17, TMT17, TTN^{+}16, VLAB18, WHY18, YDN19, YL19, ZSW17, vEKdB16]. anisotropy [CGG18, YC16]. annealing [ZWYW18]. annular [MBJ16, MBNJ16]. annuli [MF16b]. anomalies [BM19b, NMM17, PKLS17]. anomalous [ADHN15, GBU15, MP15a, YZZ19]. anomalously [MDH19, ZHWQ18]. Anomaly [AKK^{+}19, KS15b]. ANOVA [CC16b, LL16c, LL19b, MS16b, TCA16]. Antarctic [IPSG15]. antenna [SFDE15]. anti [KSSL18, ZZW^{+}16]. anti-hourglass [KSSL18]. anti-plane [ZZW^{+}16]. any [RCRF16]. AP [WSJY16]. AP-Cloud [WSJY16]. aperture [LS19b, SG18]. apertures [SL16a]. applicable [QYJ19]. Application [APP^{+}16, AS17, AP16, Bal15, BLK19, CC17a, CY19b, CGM15, DWZ19, EG17, EFO19, EFO20, GPS17b, GCVCHH18, GN19b, HHCG15, KSV^{+}15, LSLA16, MMNI16, MG15b, MB15, NOM^{+}17, NMM15, OS16, Pis18, RC18, SRBÓ17, SWS^{+}18, SZ15a, SiI16, SiI17, TCD17, Tav16, TWM18, VALT16, Zau16, ZWL^{+}19, ASB^{+}15, Ama18, AJW17, BCSK17, BAD19, BLG^{+}16, BTA17, BZ16b, Cac15b, CGSS18, CKK18b, CP16, CLG^{+}19, CVM^{+}19, Cot16, CWJ18, DS16, DL17, DAO17, DS15d, DMM19, DYL19, ECC18, EJMI18, FBL17, FFJ^{+}19, FPT17, GWC17, HKH^{+}16, HTMP17, HM19c, IPSG15, JL18c, KG15, KFWK17, LLW19, Liu16, LEB^{+}17, MRA16, MKYZ17, MP16, MSP15, NMM16, NMM19, NBH18, NBMB19, PTT19, PKK18, RXS16, Ren19, SDMS17, SWS17, Sir19, SW17b, TMWF18, TSB^{+}18, TD16b, TRLK18, Vog17, WY17]. application [WSS^{+}15, WWZ19, WHR19, WZLS19, WB17, WKSS15, XYF^{+}17, YR15, Yas17, ZM16a, Abg18a, BD15b, HTBG15, NMM17]. Applications [AZ19b, Chu17, KKL15, KHP15, MM16c, NFG15, NYD19, PSB^{+}18, PQR17, RSBS19, TBG16, TCS^{+}16b, ABDN19, ACCCDA16, ALKZ16, AAD16, AdSŚ^{+}15, BT19, BMY19, BHGK18, BDPM18, BW18b, CCK^{+}17, CCB^{+}19, CBN^{+}16, DDJ17, DLC15, DDV^{+}15, DY17, DY19b, DZC16, FK17, GBR15, GFO18, GK19b, HWH^{+}16, pHzSrC15, JL15, Jou15, KADE15, KADE17, LB17, LSD18, MWD16, MW16b, MS17, MS18d, NLFM16, PLPR19, PC19, Ram17, RG15, RO19, Say17a, Say17b, SA15, SK19b, Spe15, SCLG15, TP17, TMH18, WW19, YNW17, YL16, ZzSK15, ZS19b, ZPE^{+}16]. Applied [SL19c, AGRB18, BC16a, CG19, DCP15, DZ16, DGL^{+}15, GMS19, GBD^{+}15, GFvR18, HR18b, JdR^{+}18, LML^{+}16, NRZS17, PBA^{+}15, PA15, SWPS17, WS16, ZCHS15, dFVJ15]. Approach [TK12, TK15b, ADFG17, AMJ17, AS17, APLK19, AR16b, AMM^{+}15, AM17b, BVM^{+}17a, BB17, BHS^{+}18, BCM19, BSM16, BDPM18, CGS18, CMP19, CKK18b, CNK19, CE18, CNOS15, CJL16, CFPB17, CN16, CGJ16, DG18, DvB17, Dom18, DLWY19, EO15, EZG16, Eva18, EE16, FQZNZ18, FG16, FKR16, FFJT16, FG18, FYC^{+}18, GTL18, GR18, GLS15, GWE^{+}15, GPG17, HFND18, HB16, HF19, ISST18, JH17, KD17a, KKS15, KM16b, KP15b, KES18, KL15, KW16, KV16, KBD19, LZ18, LSWF16, LZ15a, LSP^{+}18, LO16, LMSK17, LH18, LGZ^{+}19, LFAR19, MVKD15, MRP^{+}15, MCN18, MKSP19, MNR19, Mar19, MNC19, MN16a, MRN16, MPR^{+}18, MD15, MGPG19, MWB^{+}15a, MWB^{+}15b, MTBT18, NJPB17, NDH19, NNV19, OB19, OS16, OB17, PMS15, PHÖ^{+}16, PHRA16, PE15, PJB^{+}19, RFGSV15, RT16, RO16, RS16a, Ric15]. approach [RJ19b, RSD17, RRD19, SP18, STEK17, STEK22, SKS17, SZ15a, SXY18, SLB^{+}16, SDFA17, SW18b, SWHV16, SMOM^{+}17, SP15b, SV17, SIX16, TBHG18, TFGK18, Tav15, TDC^{+}19, TT17b, TAJ^{+}17, TND18, TABR17, Vos17, WLL16, WTL19, WT16, XYPT16, XWW^{+}16, XWZ^{+}18, YS18a, YZ19, YL16, ZL15b, ZC18, ZZPH18b, ZZPH18a, ZC19a, ZCL17, ZS19b, dJRP^{+}15, tEDKT17]. approaches [LL17, CFG16, CPSF17, GHF19, HAX19, KWHB19, LYB18, MBA19, MMPS17, SGA^{+}15, YGEM17]. appropriate [FG19]. Approximate [BST19, EAAM15, KEJ18, KKLS17, MKYZ17, PP18b, Ama15, AB17, BSWG15, CLY^{+}15, DYL19, LZT^{+}15, MM16a, NJL19, SPW18, WHL17, WSN^{+}15, WLK^{+}16, XM18, YKMM19, ZSO^{+}19]. approximated [LDGH16]. Approximating [CFO18, GH19, JS19, WX19]. Approximation [ABM16, BC16b, CT15, KK17a, LB15, OS16, WC19, ALKZ16, AEL^{+}15a, AEL^{+}15b, AA19, VMN^{+}18, BDKK17, BKKY19, BA15, BZ15, BKKRB16, CL18, CSY19, CCZ18, CQ15, CVG19, CSLL15, CLX19, CLP16a, Cot16, DMRB19, DDM^{+}19, DH18b, DCBK15, DZC16, EMZ16, GNZ18, GLZ19, pHzSrC15, HKLZ18, HB16, HJS19, Ike18, Jou15, KVKS19, KZR15, LYM19, LTKA15, LLY15, LZ17a, LHLL19, lLLNS16, LYD19, LLVF^{+}15, LYA16, LY17, MML17, MP15b, MP16, MB15, MN18c, PCX17, PYK19, QWX19, ST18a, SS18a, SWX18, SLdTV18, SAOW17, VSDW18, WTL19, WX18, YYL16, YY16, ZCL17, ZNX15, ZV18, Zil15]. approximation-based [LLY15]. approximations [AEL^{+}17, BGN15, BFFB17, BFF19, CLC16, CMW16, DY17, FFM19, FFBB16, FPV18, GMS16, Hig15, JW15a, Kay15, KS15a, KS16b, LLS15, LZW19, LN15, LHA15a, MN04, MN17, MSP16, PUA^{+}15, PPM^{+}19, VLAB18, WYZZ18, WGJS19, WG15, WF17, YZW17, YY17, ZGJ16, ZzSK15]. APR [COdLL18]. aquarium [AA19]. Arakawa [SLN15]. Arakawa-like [SLN15]. Arbitrarily [LW17d, GLS15, GBS15, LBTK18, LIW18, OLV16, PN18, SUR18, TCS16a, TSN16, TYROG19, WX18, ZLGS18]. Arbitrary [BMR^{+}16, BLD15, BD17, CNG99, FFM19, TD18, WW15, ADGN17, ATF16, BtTBI18, BS15b, CNG17, CDL19, CJL16, CYL^{+}16, CE17, CR18, CCGH17, DL15, DY16, DF16, GWC18, GAIN19, GTG15, GL17, HR18a, HL16a, HHR^{+}19, JH17, KTN15, LL16b, LTR16, LK16a, LHW^{+}17, LYPP17, LVL18, LMN18, MN15, MW16b, MTK17, MWB^{+}15b, NSB15, OLD^{+}16, OKE17, PWC18b, Rag15, SWMD17a, SWMD17b, Spe15, SGT16, SGT17, TCS17, TKF17, WHL17, XX16, XX17, ZSW17, ZCL19, ZYD^{+}19a, dTP16, ABM16, BB19]. Arbitrary-Lagrangian [BLD15, BD17, BB19]. Arbitrary-order [FFM19, CCGH17, JH17]. arc [Par15]. arc-like [Par15]. Archimedean [CY19b]. Architect [MAM16]. architecture [TCS^{+}16b]. architectures [AAB^{+}16, RGPS17, RHvR^{+}15, ZAK15]. area [JB15]. Aris [GR15]. arising [BKRB15, GSMR18, HLTC18, lLNS17, PGH15, SG19a, YDN19, ZNS19]. arrangements [BDAA^{+}18]. array [BDMZ19, GD19, SCE^{+}19]. arrays [LB16, SFDE15]. Arrow [DFM17]. Arrow-Hurwicz [DFM17]. arterial [DHC16]. arteries [GDFL17, GFL17, PTT19, YPK16]. artery [BFI^{+}16, LSCC19]. articulated [BGRC19]. artifacts [MSG18a, MSG18b]. Artificial [Rod17, Rod18, WBM^{+}15b, CM18a, CSD19, CJD^{+}17, DRM15, EMM^{+}18, FRRV16, HK19a, HIN^{+}16, HP17, MTT19, Mue18, RRS19b, RRS19a, RH18, Str17, TLB^{+}18, UY19, WHR19, Wil18, YM17c, FBC^{+}16]. artificial-dissipation [EMM^{+}18]. ASAM [Cac15a, Cac15b]. aspect [Sti16]. aspects [PM16, TC15a, TKC15, WX19]. asphaltenes [ELH^{+}16]. Asselin [MRY19]. assemblies [LL18]. assembly [MNC19]. Assessment [BD18, CFSZ19, XWL^{+}16, HBC^{+}16, JD19, Shi19]. assimilation [ADP^{+}17, CZ19a, FDS^{+}15, FG18, GS15c, GM16, KYUO15, MP17, MCGS16, MWZ19, NP16, RS16a, RVMR17, SD17, SWHV16, SSN15, WWZ19, YNW17, YQNW19, AMPG19]. associated [AÁPB17, Bre18, Don18, OvdHVH16]. astrophysical [CCBF19, KB18]. astrophysics [KFF^{+}17, Teu16]. Asymptotic [BLMY17, CKK18a, CwYjS16, DD17b, DMTB15, JXZ15, JS19, MC15, SSM15, BLS16, BT16, CX15, CDN17, CDV17, DLM18, GJL19, Hiv18, HW15b, JLQX15, JL17c, JS17, LCK19, Liu19b, LP17a, SJX15, SJXL15, TW17, WY16, XJLQ15, ZLJ16, DDD17]. Asymptotic-Preserving [DD17b, JXZ15, BT16, CDN17, GJL19, HW15b, JL17c, JS17, Liu19b, WY16, DDD17]. Asymptotically [CCDL19, NMM16]. Asymptotics [LLS15]. Asynchronous [LPBR15, SGL17, AM17a]. asynchrony [AD17]. asynchrony-tolerant [AD17]. athermal [RKL18]. atmosphere [WSH19]. atmospheric [AZ16, AZ19a, CGSS18, FL16, KMS^{+}18, KG15, KS17, Mel18, MM16c, SZY16, SWMD17a, SWMD17b, SDH^{+}16, SSX16, SKG17, SKW19, SZS15, TLLF15, ZCHS15, ZA15a, GPAO^{+}18]. Atom [LLEK17]. Atom-partitioned [LLEK17]. atomic [TY17]. atomically [FGLB16]. atomically/kinetically [FGLB16]. atomistic [CDX^{+}18a, FOF15, FHS17, FS15, VKE^{+}18]. atomistically [FRL15]. atomization [GHR17, LSYF15, SLC^{+}18]. attractive [Rua18]. attractive-repulsive [Rua18]. Auction [JME18]. augmented [AGRB18, AT18, LXC^{+}15, NMM16, NMM18, NMJFM19, RTG18, XLY15, NMM15, Vog17]. Augmenting [SNK18]. AUSMD [Niu16]. auto [ZKS^{+}15]. auto-covariance [ZKS^{+}15]. autofocusing [SG18]. automated [LBB^{+}17, TST^{+}15]. Automatic [FS19a, KJ18, LZS^{+}19, TO15, ZJ18]. Automating [BHNS19]. automaton [DMS17]. autoregressive [HHR15]. auxiliary [HAX19, LYD19, SXY18, YD19]. avalanche [VBG^{+}15]. avalanches [FNGV18]. Avazzadeh [Pan20]. Average [ZSL^{+}19]. Averaged [TUJ19, BTVB15, CLW18, GZ19, LZB^{+}17, LHMB18, NMJFM19, XWW^{+}16]. averages [RL17]. averaging [BB15, DWGW17, MSG18b, SPCH16, ZSL^{+}19]. averaging-reconstruction [ZSL^{+}19]. avoid [MSG18a]. avoiding [NWKC16, Wal16]. avoids [SYM15]. Aware [TS18, BKS18, DS15a, DS15b, GK19c, LRZ17, NBH18]. axi [RZ17]. axi-symmetric [RZ17]. axial [ZCL17]. axis [ZSL^{+}19]. Axisymmetric [NOM^{+}17, BGN19, FH17, GBD17, HR17, LO19, LLFX18, LB16, LWB^{+}16, LMB18, VSC18, Xie15, ZCHS15, ZCSZ19, TBLJ15]. axon [MW16a].

B [CZBC^{+}18, FGLB16, SLVE18, YZT^{+}18]. B-spline [FGLB16, SLVE18, YZT^{+}18]. B-splines [CZBC^{+}18]. Babich [LQB16]. back [BFP18, Ols15]. back-scattered [BFP18]. backflow [BC16a]. background [BJK17, ION^{+}17]. backward [PBBK15, PKK18]. Baer [CHS17, DG16a, FRRV16, LDGH16, TT16]. Baer-Nunziato [DG16a]. Balance [PMF^{+}18, BAK19, CTM^{+}16, LPWK15, LM16, MRXI17, MN16c, NLFM16, RPC^{+}18, TM15b, VK18, WYA^{+}17b, XZZ15]. balance-Monte [XZZ15]. balanced [AASPT18, ABT16, CCK^{+}18, FNGDMNR18, FGLB16, GLK19, LX18, LYKW19, LMKS15, LAEK18, MDBCF17, NMM15, NMM16, NMM17, NMM18, PN17, PME^{+}15, XCX17]. balanced-force [LYKW19]. balancing [CV17, GFA^{+}16, JBLO15, KJ18]. ball [CWJ18]. ballistic [TP16b]. ballooning [WSH^{+}17]. band [AAB^{+}16, KH18, MHJ15, WHZ18]. band-Krylov [AAB^{+}16]. banded [JH17]. bandgap [DBD^{+}17]. bands [BVMW16]. Bandwidth [WCT18, HXB15]. Bandwidth-based [WCT18]. Baroclinic [OLHD17]. Barotropic [CK16a, CLV19, DDM^{+}19, XWB15, YR15]. barrier [AW16]. BASE [HD18]. BASE-PC [HD18]. Based [ABM16, DJV^{+}18, AGC19, AAE17, APV^{+}18, AMJ17, AS15, AS16, AB16b, AMXJ19, AA15, ABT17, ABdC^{+}18, ARTG^{+}19, BJO18, BSK15, BTD16, BFI^{+}16, BD15a, BK16a, BVG^{+}16, BLM18, BCO^{+}15, BBF^{+}17, BIR18, BC18c, BM16, BZ15, BDBEE15, BTVB15, BCM15a, BS15b, BGG16, BCB17, BPD19, BC16d, BKL17, CCFC19, CKKD19, CDM^{+}16, CCHL15, CSW^{+}19, CARdN19, CGS18, CDC17, CGL18, CC16b, CJD^{+}17, CQ15, CJL16, CZ16, CYYL18, CNQ^{+}19, CLL17, CHL^{+}19, CLX15, CGJ16, CGJ19, CGJY19, CLQ17, CELZ18, CMH15, CV16b, CYWL17, CC19c, DRP^{+}16, DCA^{+}16, DRM15, DC18b, DXvW18, DvW19, DPW^{+}15, DF16, DLK17, DL18b, EH14, EH15, ES18, Ein19, EMZ16, EE16, FRL15, FBY19, FHY^{+}19, FWK17, FG16, Fid17, FB15, FPDT17, FK17, FSK^{+}16, FC16, GTL18, GJL19, GSS15a, GZM^{+}17, GMLD18, GHH15, GOR17, GMS19]. based [GO15, GCVMK15, GFA^{+}16, GAIN19, GBD^{+}15, GN16, GZLH19, GAJ15, GJ18, GSN17, GLG^{+}19, GFW16, GYZ19, GSL^{+}19, HGR16, HEPG15, HZL^{+}15, HTZG17, HP17, HDA^{+}18, HMFJ18, HCLT19, HHL19, HW15c, HW16c, HLL^{+}18, HKS^{+}16, iI15, ISST18, JW15a, JL18a, JKE^{+}17, JZSX18, JL17b, JT18, JL19, JLKF17, JTD16, KM17, KFF^{+}17, KKJB16, KPJ18, KRFV16, KC17c, KG15, KP15c, KK17b, KGP^{+}17, KSSL18, KBF17, LH17a, LY15a, LR19a, LKK17a, LC18, LLY15, LJZ15, LLL16, LDL^{+}16, LLSJ19, LHLL19, LXL17, LC17a, LLY18, LGB16, LYD19, LLW19, LSTkM15, LJ16, LW17d, LHY17, LZL^{+}19, LLH19, LKSM17, LYA16, LLLN18, Loz17, LZW^{+}17, LP17a, MN18a, MMB18, MNG15b, MKS18, MW16b, MP17, MGBG16, MMMS15, MVZ16, MGPG19, MCHL16, Moh15, MZ15, MSF^{+}19, NBT19, NPB19, NPP15, NBZ^{+}19, NPRC15, NYD19, NL17, Niu16, OC18, OS16, OSP17, ÖPHA15]. based [OV17, PXXZ15, PCX17, PDdG^{+}17, PES19, PD17, PPLC16, PUA^{+}15, PBP18, PR16a, PHRA16, PLWJ16, PSMPG17, PR16c, PMB18, PPM^{+}19, RO16, RYZ18, RRM^{+}16, RMLvR18, RRL19, RXSG15, RXS16, Ric15, RN18a, RM16, RPC^{+}18, RRMF^{+}19, SNSG16, STR15, SRBÓ17, SNB^{+}15, SBT17, SPD19, SPB17, SLH18, SL18, SGC^{+}17, SSC^{+}16, SP16a, SF18b, SW15, SLL16, SKO17, SL16a, SGS^{+}19, SBH19, SLA^{+}19, Sto16, SWZ17, TC15a, TKC15, TBHG18, TW17, TWF19, TLQ16, TE19, TD17, UY19, VLAB18, VCEK19, VMK^{+}19, VV19, VBG^{+}17b, Vos17, WG16a, WW15, WR15, WDS15, WHL17, WC19, WHRL19, WTL19, WSN^{+}15, WNW^{+}19, WCT18, WCCB16, WH16b, WHZ18, XB18, XYPT16, XDvW17, XDSX17, XS15, XTYL18, XWZ^{+}18, XGZ19, XM18, XZT18, YJ17, YSW15, YY16, YZW17, YZT^{+}18, YBSTT19, YD19, YXX^{+}16, YLLH19, YLH^{+}19, YB17, YZZ15, YCS^{+}17, ZL15a]. based [ZHA17a, ZC18, ZC19a, ZKG19, ZCSZ19, ZVO15, ZSX17, ZYCK15, ZGD^{+}16, ZM16b, ZCL17, ZL15c, ZLGS18, ZWG17, ZPE^{+}16, dFGS^{+}17, tEDKT17]. bases [AAE17, LMBZ15, MJ17, RSB16]. Bashforth [ZM16a]. Bashforth/Moulton [ZM16a]. Basic [DC18b, WRL16a]. Basis [HD18, Mue18, SNK18, TST17, AH15, Alm19, BVS18, CMP19, CQ15, CS16b, CS18a, CPP19, FHY^{+}19, FGLB16, FBW16, GBvZB16, HXB15, JES15, JL17b, JWH16, KKL15, KMGR16, LB15, LL16c, LL19b, LHY17, Lot18, MVKD15, MF17, MMBP19, ML16, MR16b, OS16, PDdG^{+}17, RRD19, SKS17, SMT^{+}16, Sha17b, SF18b, SW18a, SP15b, TG17, WQZ15, WF17, XYPT16, XL17a, YYL18, ZLH^{+}17]. Basset [CFO18]. Bassi [MRRRF18, QN19]. Bateman [BP18, JFS17]. bathymetry [WWGK17]. Bayesian [AÁPB17, AZ19b, BST19, CZB15, CS16b, CN16, CMW16, EZG16, FOF15, FK17, GZ19, GWE^{+}15, HAPK15, HYL17, KKL15, KL17a, Kou16, LBTCG16, LPU18, LZ18, LL15, LL19b, LMTC15, MPP15, NJL19, NS16, PPCK17, SPP^{+}16a, SCE^{+}19, SPB16, UHKT19, WLL16, XWW^{+}16, YZ19, YGEM17, ZZ18, dFGS^{+}17]. BBM [BNS17]. BCR [FBY19]. BCR-Net [FBY19]. BDDC [KCW17]. BDF [HEPG15]. BDF2 [Nis19b]. be [CSCM16, ZBZ^{+}18]. Bead [DCP15]. beads [RLH19]. beam [MS15c, MMW15, MSF^{+}19, YZW17]. beams [LHB^{+}16]. bed [LMKS15, NMM18, RZ15]. bedding [ST16]. beds [BVM17b]. behavior [HLCL19, KGT15, KYKS19, LWY17, MTL^{+}17, SDJU15, YG18, BFF19]. behaviors [ZW15]. behaviour [LIW18, RS15b]. Beltrami [LYPP17]. BEM [BLJ17, Dod17, FH17, ZGD^{+}16]. BEM/FEM [Dod17]. Bénard [BGM16]. Benchmark [LP16a, ZR17]. benchmarking [PJE^{+}16]. bending [WJD16]. Bernoulli [RSSSE18]. Bernstein [LKSM17]. Bessel [Ike18]. best [LN15]. Beta [SSM^{+}17]. Bethe [BDKK17]. between [APP^{+}16, ACCCD^{+}17, Buk16, HHZZ19, KR17, KLC19, LZHM19, MTK17, Moc17, NG17, NG18, PR16a, QYF15, RZ15, TFGK18, WSK19, ZYK18, ZN16]. BEUT [SCS16]. beyond [FKF17, LYA16]. BGK [BIR18, CKT17, CLL19, Eva18, HHY15, KCHW19, LP17a, XJLQ15, XQ17, YZZ15, ZLGS18]. BGKW [JL16]. BGT [MTT19]. Bi [MHJ15, ZYK18, BCSK17, HFND18, MSH^{+}15, SS16b]. bi-diagonal [SS16b]. Bi-directional [MHJ15, ZYK18, MSH^{+}15]. bi-fidelity [HFND18]. bi-orthogonal [BCSK17]. bi-orthogonal/dynamically-orthogonal [BCSK17]. bianisotropic [DKTH15]. bifidelity [CPX19]. bifurcation [MH19]. bifurcations [EEG^{+}15, MC15, PQR17]. Big [KZG16]. biharmonic [MLMM17]. bilayer [BBMN18]. bilayers [SDMS17]. bimodular [RYZ18]. bimolecular [RFGSV15]. Binarized [HS18a]. Binarized-octree [HS18a]. binary [BGJ^{+}15, DMS17, ES18, Lau17, OTS17, RTO15, ZW19]. binding [GH17b, PD16b]. biochemical [LYB18, MPT16]. biofilm [SHP^{+}16]. biofluids [RO19]. biogeochemical [PP18a]. bioheat [HHRA19, Pan20]. Bioinspired [BI16]. biology [GH17b, SCLG15]. biomass [SNB^{+}15]. biomechanics [FK17]. biomembranes [LSMS17]. biomolecular [FHS17]. biomolecule [LYL19, YX15]. biomolecules [EG18a, XJ16]. biopolymer [MMW15]. biopolymers [LLA19]. biorthogonal [ZFPB16]. biosensors [SCE^{+}19]. biperiodic [LWZ16]. bipolar [HB15a, Liu19b]. Birkhoff [JWH16, LIW18]. Birkhoffian [SL16c]. BKZ [TBO^{+}16]. black [HS17b, WLC15, YSLY19]. blame [MSG18b]. blast [CSY15]. blasting [XYF^{+}17]. blended [SS16c]. blends [Yan16b]. Bloch [BTWY15, GYZ19, KH18, WZLS19, WH16b, ZBZT17]. Bloch-periodic [WZLS19]. Bloch-Torrey [BTWY15]. Block [LWLC17, AAD16, AA15, BMMP19, BST15, CTG16, DFGQ16, FGLW18, FLHA17, JCNK19, KNS15, LL16b, LM19c, MSD^{+}17, PLWJ16, RJLW19]. block-adaptive [BST15, CTG16]. Block-diagonalization [LWLC17]. block-preconditioners [BMMP19]. block-structured [FGLW18, FLHA17]. blocking [LH15]. blocky [SSL17]. blood [APR^{+}15, BB17, GZM^{+}17, GDFL17, GFL17, KLC19, LSCC19, MB15, MLB16, PTT19, ZZDB15]. blow [GY15]. blow-up [GY15]. blue [YZW^{+}18]. Board [Ano18y, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18s, Ano18r, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z, Ano19-27]. bodies [BHST17a, BHST17b, BHST18, BGRC19, CFSN18, CGRV17, EGO19, HM19b, JBM19, LTB16a, LC17a, MP19, MM16d, NJPB17, PN18, PR16a, QYF15, RW15a, RXS16, SGMS16, TOR^{+}15, ZLGS18, dTP16]. body [ABT17, BOA17, CZBC^{+}18, LC15, LSP^{+}18, MNC19, NBT19, PLWJ16, Say17a, Say17b, SD15, STKH15, TRM16, WG16a, WE15, YXF^{+}16, YDCK16, ZJ18]. body-fitted [BOA17, ZJ18]. body-force [WG16a]. body-forces [YDCK16]. body-of-revolution [NBT19]. Bohm [MP15b]. boiling [JS16, SN15, VALT16]. Boltzmann [GBCF16, GSS15b, Hag25, ARF18, AS16, APT17, AJVH17, BP18, BWR15, BTVB15, BAR15, Bou19, CT15, CG18a, CVG18, CLM15, CSB15, CYWL17, DBSS^{+}19, DLNR18, DCBK15, EG18b, Eva18, FGL16, FB17, FBL17, FKF17, FBJS19, GR18, GPS17a, GPS17b, GR15, GBCF15, GN19b, GW16, HK15a, HLML17, HW15b, HLU15, HJ16, HM19c, HJS19, HY15, HHY15, HW15c, HHY16, HW16c, HW16b, HW18, HWA19, Hwa16, JAH19, JSY15, KCHW19, KGT15, KP15a, KL15, KS15b, KS16d, LMPS15, LFDP16, LL16b, Li17, LLSJ19, LDWZ15, LWB^{+}16, LXSC16, Liu19b, LM15d, MG17, MK15, MHGM^{+}15, MFF^{+}19, MKV^{+}17, NSL16, Ols15, PL16b, PMGW16, PGGW18, PF16, Poë19, RS15a, RTO15, ST18a, SYI^{+}19, STW16, Shi17, STG17, SA19, SWLW19, TS17, VMM19, WSY15, WSHT15, WSY16, WC19, WGME17, WSB19, WZRZ15, WZL^{+}17, Xie15, XJ16, XTYL18, YFKS15, YYY^{+}16]. Boltzmann [YC16, ZLJ16, ZYW16, ZCSZ19, ZZL19, ZY17, ZQCT15, ZWG17]. Boltzmann-BGK [Eva18, HHY15]. Boltzmann/Finite [GSS15b]. Bond [TRM16]. bookkeeping [HB15b]. boosted [YXD^{+}16]. bootstrap [CY19b]. Boris [EBQ15, WSR15]. Born [OLV16]. Bose [ALT17, Rua18, ZS19b]. both [CFF18]. bottleneck [OZ17]. bottom [AB19]. bounce [Ols15]. bounce-back [Ols15]. bouncing [SGP17b]. Bound [EHXM15, HS18b, QSY16, CXY19, XYG19]. Bound-preserving [EHXM15, HS18b, QSY16, CXY19, XYG19]. boundaries [AB17, BFF19, BLS16, EG18b, FB17, GSN16, GT19, HF18, JSY15, LH16, LSLA16, MAK15, MM18, RF18, ST18a, YM17b, YTW15]. Boundary [BCD^{+}15, BCO^{+}15, BDG^{+}17, BAR15, CV18, DKK15, GPAO^{+}18, GZ17, GD19, GBS15, HY15, KSM19, KZR15, MAvdW18, Pan15, PF16, RVZB15, RMF^{+}18, SGT17, TSN16, TBLJ15, WSY16, AR16a, ABN15, AB16a, AMS17, AB18, AZ19a, AMP16, AHHC18, ACS16, ABG18c, AR16b, AKM^{+}19, AEvW19, Azi19, BG19a, BC18a, BBKS16, BKP16, BXY17, BRK^{+}18, BNM15, BBF^{+}17, BDB18, BC18b, BNK18, BBN18, BNS17, BHP19, BPTA16, BSP18, BG19c, BG16a, BHF15, Bre17, BBK19, BHFB19, BCRS19, CJWS19, CC19a, CDL17, CGL18, Cha16, CG18a, CYWL17, DDJ18, DRZ^{+}19, DGHP17, De18, DSH^{+}16, DC18b, DSX19, Dod17, Don15a, DS15c, Don17, DSSP18, Du18, DL18c, FR18, Fal15, FH17, FG16, FPDT17, FG19, FN17, GP17, GGT15, GLMC16, GC17, GVTQ16, GSL^{+}19, HL15a, HTFL18, HGW18, HP17, HR17, HKH^{+}16, HCLT19, HLY15]. boundary [HLSY16, HHY15, HHY16, HDF18, Hue15, IKI15, JSP16, JL17a, JW15b, JL19, JSY15, KCHW19, KDF15, KLSF15, KADE15, KLC18, KHHN16, LTB16a, LC15, LLEK17, LM18, LXC^{+}15, LFDP16, LBZA16, LCK16, LCLY19, LZ17b, LC17b, LD15, LTWZ18, LZS^{+}19, LHW^{+}17, LYPP17, Loz17, LFT^{+}16, LHA16a, LWTF19, MS18a, MS18b, MCIGO19, MK15, Mar19, MAP17, MA17, MKS18, MTT19, MP15b, MMP18, Mue18, MNW19, MN18c, NPB19, NBZ^{+}19, NYD19, Nis15, NL18a, NW15, NRS19, OB19, Ols15, OM19, ÖPHA15, PLL^{+}15a, PHHR17, PNZ18, PPLC16, PKJ^{+}18, PCN15b, PN18, PLL15b, Pes15, PTT18, PE16b, PMF15, PDRB17, PG18, PGH15, PVB17, QSB18, QM18, RS16b, RS18, RDG17, RZ17, SYOS19, SYOS21, SS17a, SWS17, SL17, SKF15, SKF16, SHKL16, STKL19, SF18a, SK15a, SMA^{+}16, STG17, SLVE18, Smi18, SMLB15, SMSR18, ST18c]. boundary [STV19, SMOM^{+}17, SLdTV18, SGT16, SHP^{+}16, SCLG15, Stü15, SWL19, SJH^{+}15, TCD17, TZGW18, TP17, TNB21, TTN^{+}16, Tsa15, Tsa16, TKF17, Vai15, VAD17, VMC^{+}19, WN18, WG16a, WZ15, WE15, WCH^{+}17, WL18, WVB19, WZLS19, WNW^{+}19, WS15a, WBM^{+}15b, WGME17, XY17, XTYL18, YK15, YS15, YD18, YM17c, YZZ15, ZL15a, ZB15, Zha16, ZG18b, ZZL19, ZY17, ZSX17, ZLL^{+}17b, ZZH16, ZRT18, ZSM19, dDPG19, dTP16, SCS16, SIX16]. boundary-constraint [XY17]. boundary-finite [GSL^{+}19]. boundary-integral [QM18]. Boundary-Lattice [PF16, LFDP16, WSY16, XTYL18]. boundary-layer [BHFB19, NL18a]. boundary-value [WZ15]. Bounded [MSP19, AG18, BLS16, Don17, IM17b, JHPAT17, KBR17, LDB19, LI15, MS18c, NGY^{+}17, PBCR19, YG19, YLA15]. Boundedness [HDA^{+}18, SMD18a, SKC17]. Boundedness-preserving [HDA^{+}18]. bounds [BMC^{+}18b, HFND18, MSK18, MM15, Tso18, WK18]. Boussinesq [UL16, ZA15a]. boxes [SS17b]. Bracket [Suz18]. Braginskii [MP16]. brain [TT17a]. branches [XL17b]. break [GWYS18]. breaking [AW16, FKR16]. breathing [MCHL16]. Breit [JdR^{+}18]. Brenner [UY19]. brick [WR16]. brick-tetrahedron [WR16]. Bridging [KLC19, DPW^{+}15, SDJU15]. brief [Shu16]. Brinkman [GX15, HKLW15, LPB17, STG17, SHW17]. brittle [ZHLZ18]. broad [JB15]. broad-area [JB15]. broadband [ZZH16]. broadening [DJD^{+}17, JDFS16]. Brownian [BT17a, BRK^{+}18, DH18a, MMW15, SPRW15]. bubble [FP18, JSVD17, LZHM19, ZL15a]. bubbles [HTBG15, KZR15, NBMB19, SKF16, WB17]. bubbly [MLL18]. building [ARG^{+}17, CC17a]. built [BC18a, TBG16, dLKK19]. built-in [TBG16, dLKK19]. bulk [CM18a, COV18, PK17, ZV16]. bulk-surface [COV18]. buoyancy [KA18, LT15]. Burgers [EAAM15, MK17, dlHC16]. burning [SNB^{+}15]. butterfly [Yan19]. BVD [SIX16]. bypass [BFI^{+}16]. bypassing [CPT16].

C [SRBÓ17]. CAD [MBS19]. CAF [GBR15]. Cahn [HTMP17, BKR19, CS16c, CLS^{+}18, DD16a, DJLQ18, GX15, HW15a, JJ18a, KS16a, KMdB16, KJYC17, LJZ15, LCK16, MGCW18, SL19b, Tav16, WX17, XGZ19, YLD19, YLLH19, ZSX17, ZYCK15]. calculate [LSP^{+}18, SLL19, WT16]. calculating [DB16a, SWZ17]. calculation [AAL15, CSW^{+}19, CLY^{+}15, CHE^{+}17, For16, GZ18, HS17a, HM16b, KH18, LZSG19, Mac15, MH19, MDP18, QS16, SY17, SFP16, WNW^{+}19, Yan17]. calculations [ADFG17, CSN18, EH14, EH15, GLZ16, HED^{+}16, HLTC18, KK16, LHS^{+}18, LC19, LKN17, LLVF^{+}15, LYZ19, LY16d, LY17, MJ16, Mas18, MDP^{+}15, PDdG^{+}17, PUA^{+}15, PD16b, RO16, WKSS15, XS15, ZJLC15, ZLH^{+}17]. calculus [CC17c, MHS16, NBT19, SMC15, SP18, VBL^{+}16]. Calderon [GPRA18, AAE19, DDV^{+}15]. calibrate [LSWF16]. calibration [FOF15, KL17a, NHM17]. Can [WDG^{+}17]. Canny [FKK19]. Canny-Edge-Detection [FKK19]. Canny-Edge-Detection/Rankine [FKK19]. Canonical [CQL^{+}17, LBZ16, KS16b, RBD17, ZZH16, ZZT^{+}16]. capabilities [AKZ16, BBK19, PJE^{+}16, SSC^{+}16, SP16b]. capability [MMPS17]. capacity [BHP19]. capillarity [Sun19a]. capillary [DvW15a, HM16a, LSMS17, LW18, LT15, MC17, SPD^{+}17, TBLJ15, ZZ17b]. capsule [BLJ17, ISST18]. Capturing [Sid18, ABIR19, Bar19, BJ15, CLG^{+}19, DSX19, GHR17, JBM19, JSS15, JLC15, KYW^{+}16, KYW^{+}18, KLWQ17, LTWZ18, OSKN18, PSS17, QWX18, RLGT19, SP15b, Wil19, WL17, XX17, ZSMP19]. Caputo [DYL19, DZC16]. carbon [GGL^{+}17]. carbuncle [sCYxL^{+}18, Rod17, Rod18]. cardiac [CGG18, MSV^{+}16, VLP^{+}16]. cardiology [PQR17]. Carlo [BC16b, Gho17, Mac16, YZZ19, AR16a, BP18, BTA17, CSS15, Cha16, CL17, CSN18, CG15, CW18, CHE^{+}17, Cos16, DPW^{+}15, DG16c, EARA15, EN17, FDKI17, GB15b, GMS16, Gen11, GDS^{+}16, GAJ15, GBU15, Hig17, HC17, HMRG16, ION^{+}17, KM17, KMS^{+}18, KL16, KC17b, KES18, KK17b, KLGO18, LS15a, LBTCG16, Lan19, LPU18, LYCC17, LYB18, LB17, LXL17, LWL18, MNO^{+}17, MZTS16, MSS16, NHA18, PJE^{+}16, PUA^{+}15, PDS15, Poë19, PvL19, QPK19, RFPSSA18, RRL19, RKH15, SY17, Swe18, TSR15, WBC^{+}16, WL16, XZZ15, XR17, YC15, Yas17, ZLJ16, Zil15, vdKK16]. carrier [vdKK16]. carriers [SU15]. cartesian [ADOP18, FGLW18, ACS16, BNK18, Cai16, CARdN19, CXL16, DDJ18, DM16, GP17, GNK18a, GNK18b, HS17b, HS18a, HLL^{+}18, LPW15, LGB17, Mar19, MM16d, MM18, QDRB15, QLF16, RBI18, STK^{+}16, SLY16, Sti16, TDC^{+}19, XTYL18, dBIM16]. cascade [SFT16]. cascades [FBL17]. cascadic [PHHR17]. case [BHZ16, CGS18, FNGV18, MRRRF18, PP19, RRS19b, RRS19a, Rod18, VSM16a, VSM16b, WY19, WCWY19, WLE17, ZR17]. CASL [TK16]. casting [Swe18]. cathodic [PYAG19]. Cauchy [LY16a, MST15, PZF16]. Causality [UHKT19]. cavitating [ESHA16]. cavitation [MC18, ÖPHA15, PS14, PS15a]. cavities [VMN^{+}18, GFvR18, HK16b, LGO17, NMJFM19, PLL15b, UWH17, ZZ19]. cavity [EN17, GKE15, MH19]. CCH [BMCK15]. CCS [SFT16]. CCS-RG [SFT16]. CDG [LXC19]. CDG-FE [LXC19]. CE [WMS18, ZWL^{+}19]. CE/SE [WMS18, ZWL^{+}19]. Cell [CLMZ17, DFS16, LAL18, TMT17, AR16b, BTGM17, BGTM18, Bat17, BNK18, BMRA^{+}15, BDZ15, BDLM18, BLC^{+}17, Bra16a, BMCK15, CC19a, CWB^{+}19, CHJT17, CGP16, DM16, DJV^{+}18, DL15, DL16, FGLW18, FLW16, FS17b, GBM16, GFA^{+}16, GNK18a, GNK18b, GZLH19, GH17b, GPG17, HWH^{+}16, HXLL15, HLS19, ISST18, JST17, KKH18, KHTZA16, KBF17, Lap17, LPW15, LYZ15, LY15b, LSD^{+}17, LSTkM15, MMNI16, MHZ^{+}15, MGPG19, MDD^{+}19, MM16d, MM18, NRZS17, PxRS17, PE16a, PHÖ^{+}16, PMF15, RH18, RLH19, SGMS16, SSM15, SCLG15, SPCH16, dCPDC^{+}17, TM15a, VSM16a, VSM16b, WHY18, WCCB16, YXD^{+}16, ZOG19, ZXDL17, AG18, CT19, DDD17, MNO^{+}17, MSD^{+}17, TUJ19]. cell-based [KBF17]. Cell-centered [LAL18, TMT17, BDZ15, BDLM18, BMCK15, CHJT17, CGP16, DL15, FGLW18, FLW16, GBM16, LYZ15, LY15b, LSTkM15, VSM16a, VSM16b, ZOG19, ZXDL17]. cell-centred [Bat17]. cells [DF16, HXLL15, HGR16, KLC19, LMB19, PG18, RH19, SPD19, XL16]. cellular [BB17, DMS17]. cellular-scale [BB17]. Cellwise [CSH15]. center [PKK18]. centered [BDZ15, BDLM18, BMCK15, CHJT17, CGP16, DL15, FGLW18, FLW16, GBM16, HHK15, LAL18, LYZ15, LY15b, LSTkM15, MWB^{+}15a, MWB^{+}15b, SP18, TMT17, TLB^{+}18, VSM16a, VSM16b, ZOG19, ZSW17, ZXDL17]. central [HC18a, HLA19, IDSG15, LN15, LZZS15, LMKS15, LAEK18, SY18b, TLQ15, TLQ16, TK12, TK15b, WLW^{+}18, XL16]. central-upwind [HC18a, LMKS15, LAEK18]. centred [AGBL15, Bat17]. centroidal [YGJ18, FHA17a]. cerebral [YPK16]. certain [GSS15a]. Certified [SFDE15]. CESE [SP19b, YFJ17, YFJ18]. CFD [EH15, XS15, ALO18, AdSŚ^{+}15, BLG^{+}16, CJK^{+}19, EH14, JMM19, LKK17a, LKK17b, MS16b, MMSS15, MH18b, PJB^{+}19, TR19, VLTPS16, WTL19, YCPD15]. CFD-DEM [BLG^{+}16]. CFOSLS [VLN^{+}18]. CGR [BMCK15]. Chain [Lan19, KBK15a, RKL18, YZZ19]. chains [MWD16]. Challenges [PEVG18, CSCM16, TK16]. Chance [CSS15]. Change [KYKS19, AT18, DD15, HW15c, HW16c, LRA17, LSD^{+}17, NLW^{+}16, VW18, ZN18, ZKG19, ZCY^{+}19]. changes [FB15]. changing [DCCC16, Liu16]. changing-connectivity [Liu16]. channel [BKG15, DG16c, KCS^{+}17, KP15c, KFWK17, SHLG15, ZV16]. channels [WBM^{+}15b, ZMF15]. Chaos [ABM16, ARG^{+}17, ATM^{+}18, AM18, BSN19, GMS19, GGW17, GNZ18, GLZ19, HD15, HD18, JES15, KSV^{+}15, KS16b, LMTC15, NDH19, OB17, PHD16, SS17b, SG17, TE19, TMES19, THS^{+}19, YZ19, JL18a, TG17]. chaotic [Blo17, BW18a, CNW17, LSM19, Lia16, NW17, NT19, NWFT19, SP19a]. Characteristic [HTZG17, NF17, FL16, HP17, Hue15, JSY15, LLP^{+}16, Mag19, SWZ15, SW16, WGME17]. Characteristic-based [HTZG17]. Characteristics [FSK^{+}16, APR^{+}15, BR15b, BR16, HL16b, LM15b, WPB15, ZCL17, ZWG17, ZG19]. Characteristics-based [FSK^{+}16]. characterization [AABD15, AÁPB17, DKTH15, SNB^{+}15]. Charge [TC15a, TKC15, AP16, BVS18, GZ17, MXL16, NOM^{+}17, RMC15, SU15, vdKK16]. Charge-and-energy [TC15a, TKC15]. charge-conservative [NOM^{+}17]. charged [BMR19, CFPB17, HSLQ15, HSLQ16, SZCL18, Tao16]. Charney [HK15a]. Chebyshev [BW18b, Fal17, HB15a, JB15, Kas15, KKJB16, LB15, MJ17, Moo17, O’S15b, PBKK17, PKA^{+}16, VK16]. Chebyshev-like [Fal17, LB15]. chemical [BHdD18, BWR15, DEZ16, EEG^{+}15, FRW16, MTK^{+}16, SPB17, VBG^{+}17a, WHCN17]. chemistry [MA16, PD15, SZY16, SXBB15, XXR18]. chemotaxis [MMNI16, Yas17, ZM16a]. Cherenkov [NT16]. chiral [HLCL19]. choice [FYZ^{+}15]. choices [KBF17]. Choleskey [FHY^{+}19]. chosen [DJD^{+}17]. CIP [FYO^{+}15]. circuit [SDFA17]. circular [SHW17]. class [BGS16, BDB18, DLS15, EE16, FHA18, GSC19, GHV19, GSS15a, LM15a, LZZS15, Lot18, MD17, MW15, NPC15, O’S15b, RBL16, SSO^{+}15, ZS17]. classical [LYA16, Spe15, TSR15]. classification [ACC^{+}15]. CLBM [ZZPH18b, ZZPH18a]. cleaning [PMF15, TPB16, YJ17]. clinical [CVM^{+}19]. cloak [Chu17]. cloaks [WHL17]. close [CKK18a]. Closed [Shi19]. Closed-loop [Shi19]. closely [CB19]. closely-spaced [CB19]. Closest [Vog17, KR17, PR16c, PLR18, PLPR19]. closure [BKS18, BFM19, GHH15, MKC17, SGC^{+}18a, SGS^{+}19]. closures [AS15, Bre17, BV18, DC18b, SBT17, Sch16a, Sch16b, SWL19, ZM16b]. cloud [MC18, SKG17, TUJ19, WSJY16]. Cloud-In-Cell [TUJ19]. cloud-resolving [SKG17]. Cluster [CELZ18, TAJ^{+}17, AZK16, HMBH15, KMD^{+}18]. Cluster-based [CELZ18]. clustering [FS16]. clusters [BEJ15]. CMFD [JPLL15, KL16]. co [Eng18, Kla15, GMS16]. co-located [Kla15]. co-volume [Eng18]. coagulation [KK17b, LPWK15, MZTS16, XZZ15]. Coalescence [SRS19, FBL17, MOR18]. Coanda [PQR17]. Coarray [GBR15]. Coarse [KGT15, MLL19, dlCGCA17, CWB^{+}19, CSCM16, FOF15, GK19c, HKKP16, KKP15, KCW17, KC17c, LKN17, MVKD15, SZK17, TWN19, YFC19, Mas18]. Coarse- [KGT15]. coarse-grained [FOF15, HKKP16, KKP15, TWN19]. Coarse-graining [MLL19, dlCGCA17, GK19c, MVKD15, SZK17]. Coarse-Mesh [Mas18]. coarsening [DD16a]. coastal [CK16a, RBY19]. coaxial [SR18]. code [BPL19, CDX18b, DTA^{+}15, EKV^{+}16, GFA^{+}16, HRJ^{+}16, HED^{+}16, HdBH^{+}16, JL18a, JdR^{+}18, KB18, KFF^{+}17, KYPK15, MAM16, MNO^{+}17, MSD^{+}17, MHZ^{+}15, PJE^{+}16, TBC^{+}16, WLM15, WSH^{+}17, XWL^{+}16, YXX^{+}16, GFA^{+}16]. codes [GDS^{+}16, JH15, LSWF16, MH18b, PMF15, PD15, RTG15, TSN16, WS15b]. coefficient [Cif19, DCA^{+}16, KBK15b, LMMS16, NKN^{+}17, RBI18, SWLW19, WZ15, WW17, YD19]. coefficients [BST19, BSWG15, CLX19, CELZ18, CR18, Cui15, DJV^{+}18, DCCC16, DKTH15, ED16, FCL19, GT18, HHLY17, KCW17, MS16a, MNR17, MH18b, Nis18b, OZ17, Ran18, RL18, TST17, Wu19]. coherent [BGD19, CWS18, FBG15, NJ15, PEVG18]. CoKriging [YBSTT19]. cold [BJK17]. ColDICE [SC16]. Cole [LP17a]. collapse [LEB^{+}17]. collection [JBM19, TBHG18]. Collective [DG16c]. Collision [Mac16, MNO^{+}17, AWS16, BVM17b, BTVB15, CT15, HYK^{+}16, HM19c, JAH19, JdR^{+}18, RRS19b, RRS19a, RSSSE18, RKH15, SK19a, WC19, YSWW16]. Collisional [TKC15, CBB16, GJL19, LLD^{+}16, ZG17]. Collisionless [TC15a, AG18, HK15a, KHTZA16]. collisions [HMRG16, LYCC17, Mac16, SWHK15, TSR15, YXW19, YC15]. Collocated [DPO16, BDAA^{+}18, Bra16b, TLH15, WWR16, ZKG19]. Collocation [BDV17, Mue18, RTV17, BHS^{+}18, BA15, EDC16, FBF15, JWH16, JTD16, Kas15, LWL17, LZT17, NVBDV15, PBKK17, PGH15, SGN16, SMLB15, SP15b, Sub15, WW17, YC17, YDLC19, ZK15, ZTBW19, ZZW^{+}16, ZLX17]. colloidal [FKY15, NWZ18]. colloids [Han19]. color [KLWQ17]. colored [MGT18]. column [CK16a]. columns [DLY17]. Combination [DC18a, HF18, LTB16b, Lot18, Zil15, DJD^{+}17, HB16, OMLdL16]. Combined [YJM19, AAL15, CZ19a, CH17, DLK17, DS15d, GS15a, KE15, DV17, MBD19, RPNP18, SDW16, TZGW18, YWHP15, ZB15, ZD17]. combines [MBA19]. Combining [DZ18, GK19a, PKW17, CGG18, HLML17, KSV^{+}15, LSMS17, YYL16]. combustion [CPV16, CPdS19, SLC^{+}18, TMWF18, WHR19, WMYG16]. combustors [MBJ16, MBNJ16]. Comment [EBQ15, Pan20, QHZ^{+}15, XS15, ZJS15, CC17b, EH15]. Comments [Gho17, HSK^{+}15]. common [LLJJ18]. common-refinement [LLJJ18]. Communication [NBH18, LH15, NWKC16]. communication-avoiding [NWKC16]. Communication-aware [NBH18]. Comp [GBCF16]. Compact [Cui15, GT18, GGT15, TMH16, WRL16a, WRL16b, WRPL17, YLA15, ZRW19, AW18, BZ19a, Bre17, CZL18, Fan16, GS15a, HF18, HJW19, JPSX18, JJ19, JS19, Ler16, LZZS15, MSP19, OVP15, PX16, RSH^{+}17, SS16b, SWL19, WLW^{+}18, WL17, YWHP15, YT19, SG19b]. Compact-WENO [SG19b]. compaction [MDP18]. Comparative [ED16, KS16a, KGS17, CX15, MVZ16, RS15a, TK15a, WMM^{+}18, ZFL^{+}19, ZED15]. Comparing [GBR15]. Comparison [BZ19b, BKR19, EMM^{+}18, GWB^{+}15, PUA^{+}15, Pas16, SS15b, YM17c, BD18, CFSZ19, Ein19, FKF17, JZSX18, KS16b, LGB16, NdlLPCC19, PBL^{+}19, RMC15, VWV17, WG15]. Comparisons [MAM16]. Compatible [BMC^{+}18b, MO18b, SY18a, BC18b, BCS19, EL18, EL19, GBM16, KSVB18, KSSL18, SGC18b, SLVE18, SO15, TMH18, YSC^{+}17]. Compatible-strain [SY18a]. complete [MG15a, SD16]. complex [AMS17, ALM^{+}17, AZ19a, AC17, AEvW19, BB17, BN19, BHP19, CGSS18, CZL^{+}15, CM19, CD17, CRZ17, De18, DOO17, DD16b, FLT18, GLS15, GS18, GEZK16, HAPK15, KWHB19, KJ17b, LCK16, LBTK18, Mar19, MRK15, MHT^{+}19, MR16b, Noe15, RS16b, SMLB15, TK15a, TDC^{+}19, TP16b, VBG^{+}17a, VD16, WXW15, WWRS17, XDvW17, YDCK16, ZYW16]. complexes [KSVB18]. complexity [CCDL19, LYCC17, OZ17]. complexity-bottleneck [OZ17]. complexly [GN16]. compliant [HJY19]. complicated [ABFR16, SYOS19, SYOS21, TNB21]. component [ATC19, Did17, FB15, GZ17, HHM17, KS16c, LFDP16, LCK16, STW16, Tav16, Vos17, ZS19b]. components [TBB^{+}19]. componentwise [CLP16a]. Composable [JCNK19]. Composite [SGP17a, BCM15b, JHPAT17, JW15c, LJZ15, LSS16, ZWYW18, RZ15]. composition [KL19]. compositional [BMT18, CFvKH18, GV18, MTZ16, MTJ17, MTJ18, MF16a, WKSS15, XML17]. compound [MT17, PZNG15]. compounding [KL19]. Comprehensive [RLV16]. compressibility [GZM^{+}17, HP17]. compressible [AIP17, AD15, AMS17, AZ16, ALA16, BTD16, BHKS16, BMR^{+}16, Bar19, BJ15, BAD19, BHF15, BC16c, BCJ19, Cai16, CM18a, CBS18, CFSN18, CYL^{+}16, CYYL18, CZL18, CL19a, CSN17, CXY19, CC16c, CPS17, CCPdL19, DDJ18, DDJ19, DRZ^{+}19, DG18, DWR18, DLM18, DIX^{+}18, DXvW18, DvW19, FMRZ17, FST15, FBJS19, FHA16, GOR17, GHR17, GWK16, Ger17, GN19a, GNK18b, GMT19, HL15a, HZL^{+}15, HTZG17, HTBG15, IGQ15, JSP16, JYY18, JL18c, JLC18, KD17a, KKH18, KS17, KTK18, KTK19, Lap16, Ler15, Ler16, LW17b, LPR18, LSD^{+}17, LSR16, LH17b, LSZ18, LMZ19, LNM15, LZW^{+}17, MA19, MDAB18, MM16d, MM18, MPMB19, NDCB17, NF17, OB19, OSKN18, ÖPHA15, PX16, PHHA18, PL18, PSS17, PSB^{+}18, PM16, PWC18a, PCN15a, PCN15b, PHÖ^{+}16, PS16, PT18, PBC^{+}17, QLF16, QSB18, QSBY19, RLV16, RMF^{+}18, RJ19b]. compressible [SWC18, SWS17, SPD^{+}17, SPD19, SP15a, SGMS16, SHA16, SWPS17, SP19b, SG19b, SY18b, SKC17, SWL19, SST^{+}15, Svä15, TD17, TWH15, TGY18, TT16, TABR17, VM15, VGZ18, VSM16a, VSM16b, VBF15, WW15, WLM15, WCH^{+}17, WS15a, WDGW17, WL17, XYF^{+}17, XDLX19, YSW15, YSWS16, YWS^{+}16, YYJ^{+}19, Zha17c, ZHA17a, ZMCC18, ZCY^{+}19, ZXW^{+}19, ZYD^{+}19a, ZW19, dBIM16, dFVJ15, dLDG^{+}18, dPSS16, vOMB17]. compressible-fluid [FHA16]. Compression [LY15c, CC19b]. Compressive [HD15, THS^{+}19, LSD18]. Comput [ABG^{+}19b, ASS17, BLL20, CNG17, Dav15, DK18a, DvW19, EFO20, GRT21, Gho17, Hag25, HGN17a, KYW^{+}18, KTK19, MN17, NG18, PS15a, SZN20, SYOS21, STEK22, SWMD17a, SYV17, TK15b, Vre21, ZJS15, ZCQ20]. Computation [BDMC15, GGL^{+}17, HKLZ18, MHL17, MTD15, NL15, Pru18, TBB^{+}19, AQ19, ALT17, AL19a, BJRF18, BLL16, BHP19, CGTH18, CC17b, CPS17, CG16, DG16a, Dod17, EMZ16, FFW17, FCL17, FBG15, FYC^{+}18, GCI19, GHH15, GFvR18, GLMC16, GT19, ION^{+}17, KH15, KSVB18, KB19, LVTR15, LS19a, LZ19, LO16, LDGH16, LDHJ15, MBSS15, MT18, NCP^{+}17, PSB^{+}18, PK17, SCQP16, SWLW19, Tre16, ZZH16, ZLX17, dMRHJ17]. Computational [AK17, BTGM17, BGTM18, BR16, Cac15a, DD16a, EH15, FKF17, Fon16, FAC^{+}19, Gam15, HSK^{+}15, Kat16, MSV^{+}16, PQR17, TS19, VS17, WHCN17, XS15, YG18, Zoh17, ATF16, BB17, BBK19, LL17, BZ16b, BKL17, CCBdL15, CV16b, HHCG15, HHRA19, JC17, KS16a, KSV^{+}15, KP15b, KLC19, KZG16, KPP^{+}19, KBF17, LFR17, LGZ^{+}19, MMNI16, MNC19, Moh15, NPC15, NJHL18, NGS16, PVFN15, Pan20, PBP18, SBG^{+}17, SLC^{+}18, XTS^{+}16, ZR17]. Computationally [HMBH15, Tav15, FK19, PMS15, SXBB15]. Computationally-efficient [HMBH15, PMS15, SXBB15]. Computations [Niu16, Cif19, EN17, Fal17, FH17, FSB16, ISP^{+}15, KD17a, KH17, MS18a, MS18b, MC15, MMSS15, PKA^{+}16, RDG17, RXSG15, SGC^{+}17, Sha17a, SMSR18, WMM^{+}18, WF17, ZS16]. compute [FDS^{+}15, PWC18b, RG15, SKF15]. computed [HR19]. Computer [Fed17, KL17a, TT19]. computers [GP18, WLC15, YM17a]. Computing [BJTZ15, VMN^{+}18, BKKY19, CAA18, GN16, HLTC18, QLS^{+}19, VCNP18, XP15, YPC19, ABR16, ARTG^{+}19, Cac15a, Cac15b, DLN15, FHY^{+}19, GH17b, GYZ19, GP16c, HAPK15, HXB15, NJPB17, OD15, RGPS17, RLP16, Roy15, Rua18, VYP15, VCNGP15, VCNOP18, XJ16, XZZ15, YXW19, YX15, ZAK15, ZRT18]. concave [AKM^{+}19, WT16]. concentrated [ZVO15]. concentration [BHdD18, Han16, LSS16, LDWZ15, SG16]. concept [AB15, SKO17]. concepts [KK17b]. concurrent [TKB^{+}15]. Condensates [ZS19b, ALT17, Rua18]. condensation [FSK^{+}16, KKLS17, ZSO^{+}19]. condensed [MN16b]. condition [BSP18, BG16a, Don15a, Don17, GSK18, GSN17, HGW18, HHY15, KLSF15, LM18, LHA16a, MK15, Mue18, Ols15, PLL^{+}15a, PZNG15, PKJ^{+}18, SF18a, SL16b, SJH^{+}15, Vai15, Vel19, WSY16, YD18]. condition-enforced [WSY16]. Conditional [FLV18, LDHJ15]. conditioned [CCB^{+}19, Cot16, JWH16, PPLC16, SO17, WBC^{+}16]. conditions [AMN18, AR16a, AMP16, AHHC18, ABG18c, BG19a, BC18b, BHdD18, BJ15, BBN18, BNS17, BPTA16, BG19c, BAR15, BHMS18, CJWS19, CC19a, Cha16, DRZ^{+}19, DGL^{+}15, DS15c, DSSP18, DL18c, DKK15, EG18b, FN17, FKK19, GGT15, GVTQ16, GSL^{+}19, HL15a, HTFL18, HP17, HKH^{+}16, HY15, Hue15, JSP16, JW15b, JSY15, KSM19, KZR15, KHHN16, LLEK17, LXC^{+}15, LCK16, LZ17b, LFT^{+}16, MNR19, MTT19, MP15b, MMP18, MNW19, MN18c, NYD19, NW15, PGCG18, PCN15b, PE16b, PMF15, PDRB17, RZ15, SYOS19, SYOS21, SS17a, SK15a, STG17, Stü15, SC18b, TSN16, TNB21, TTN^{+}16, VAD17, VMC^{+}19, WN18, WSY16, WZLS19, WNW^{+}19, WGME17, XP15, ZSX17, ZZH16, dDPG19, Pan15]. conductance [DWZ19]. conducting [DPRZ16, MML17, Par15, Par17, Par18b, SF18a]. conduction [CP16, HC17]. conductivity [BMPS18, HS17a, KK17a, LYDB17]. conductor [CC19a]. cone [PS17]. configuration [MP16]. configurations [BPL19, RG15]. confined [GBCF15, GBCF16, GSS15b, HHZZ19, TS19]. Confinement [Ram17, Sid18, RKL18]. conformal [ADGN17, BC16d, Dom18, Fuj19, iI15, MC17, RMBN18]. conformation [MOAA15]. Conforming [FKS19, ÅN19, CZBC^{+}18, FNNB19, GM19, RRD16, ZFL^{+}19]. conjugate [ALT17, EMS^{+}19, MBHS17, NSK^{+}16, PLC18, STK^{+}16, VYP15, VBG16, YK15, ZVO15]. connected [LDL^{+}16]. connectivity [HM19a, Liu16]. Conservation [Sla16, Abg18a, AW18, BD15b, Bal15, BT16, BK16b, BLD15, Bra16a, CCRdL17, CHOR17, CS17a, CH19, Cho15, CGJY19, Del15, DC18b, DL18a, DL18b, EFT15, FPASS16, FS15, FS17b, FHA17b, FHA18, GJL19, GNK18a, HLS15, HAH16, IBML16, IC17, IDSG15, JL18c, KGS17, KG15, KC18, LMS17, LPG18, LLSJ19, LMBZ15, LYZ15, LY15b, LHGF16, LHGF19, LHQ19, LSI16, MDVM16, MDHC15, MRXI17, MB15, MFG15, NT15, NMM16, NR17, NG17, NG18, Nor15, PxRS17, SW17a, SL18, SWZ15, SWLZ15, SW16, SWPS17, SPP16b, SKC17, TLQ15, TM15a, TKP16, VNA15, WLGD18, ZP16, ZPW18, ZQ16b]. conservation-moment-based [LLSJ19]. Conservative [ARF18, ADGN17, ADN19, CCS18, CNG99, CCK^{+}18, DRZ^{+}19, Hag25, HLJ^{+}19, IM17b, MCL19, PF15, TPT16, VSC18, ZSL^{+}19, AHNF15, AMH^{+}18, Abg18a, AASPT18, APP^{+}16, AM17b, BN17, BNGI19, BTVB15, CQQ16, CNG17, CC16a, CC17b, Cha18, CD17, CSH15, DGMT17, DXvW18, DvW19, DSS18, Du18, DY19b, DB16b, EHXM15, FGLW18, FL16, GSK18, GWWC17, HHA15, HSK^{+}15, HHY15, JKM19, JW16, JH17, JJ18a, JJ18b, KJYC17, KL18a, KL18b, LGH^{+}18, LM19a, LHA15a, LHA15b, MS18d, NOM^{+}17, NN17, NF17, OvdHVH16, PNZ18, PHHA18, PA15, QWX18, QWZ19b, SGMS16, SA16, SFT16, SWLZ15, STV18, SOS19, SLY16, SY18b, SMAG17, SCC19, SK18, TCS16a, TND18, Wac15, WWR16, WH15, WZ15, WKOE17, Wil19, WRL18, XS19, Zad11, ZA15b, ZG17, ZKG19, dLKK19, FRO17]. conserved [Sto17, WSS^{+}15]. conserving [BC18b, BMR19, BMC^{+}18b, CC19a, CCZ15, CLX19, FGL16, FS19b, GK19a, HJZC17, JST17, Lap17, LYKW19, LSYF15, OD17, PG17, SLN15, SD16, TC15a, TKC15, TCSM15, WGJS19, WG16b, CT19]. considerations [COdLL18]. considering [MKV^{+}17]. consistence [LHA15a]. Consistency [Don17, AWJ17, DDJ19, NG17, NG18, Stü15, Stü17]. Consistent [MLB16, ADFG17, BAGK16, BKR19, Bre17, Cen19, DK18a, DK18b, DWG^{+}18, Don18, HHR15, HL15a, JSP16, KS18b, KRK^{+}18, LZW19, MD18, NN19, OMLdL16, OLD^{+}16, OLB^{+}17, PKP^{+}17, PNZ18, PN17, Pei16, PS14, PS15a, PMGW16, RMC15, SV19, STK^{+}16, SK18, TFGK18, TTN^{+}16, TKP16, TSR15, Wac15, WY17, WXSJ19]. consolidation [AGRB18]. Consortium [TM17, TK16]. Constant [WY19, BMPS18, Cif19, LTKA15, MNR17, OKE17, WG15, ZC18, ZC19a]. constant-coefficient [Cif19]. Constant-density [WY19]. constant-property [WY19]. constants [OKE17]. constitutive [CWB^{+}19, TBO^{+}16, ZLC^{+}18]. Constrained [BKS18, Cot16, CLNH15, TPB16, VCEK19, BGHK19, CF19, CVG19, FS18, FMPT18, JME18, MAP17, Moc17, Tav15, TD16b, VLN^{+}18, XX16, ZZKP19]. Constraint [SV19, BTGM17, BGTM18, BK19a, CEL18a, FG18, HCB19, PBP18, RS16a, SD17, XY17, YSLY19]. constraint-based [PBP18]. Constraint-consistent [SV19]. constraint-preservation [BTGM17, BGTM18]. constraint-preserving [BK19a, HCB19, YSLY19]. constraints [ADE^{+}17, CG19, DRP^{+}16, EST17, HX16, LDT19, RKB19, SHTY19, WT15]. construct [Abg18a, SGC^{+}18a]. constructed [SGC^{+}17]. Constructing [AEAM15, FN17, LTR16, DB18, EG18a, HHR15, KV16, RT16, XY17]. Construction [HY17, RSB16, AG16, AA19, CPP19, MW16b, OS15a]. consumption [FYO^{+}15]. Contact [BM19a, LRZ17, ABG^{+}18b, ABG^{+}19b, BKR19, DL17, DHH^{+}18, Don17, FB17, FRL15, FFJ^{+}19, FPV18, Gan15, HW18, HKS^{+}16, LPGT16, LD15, Liu16, LDGH16, LHA16a, MAK15, MHGL19, PR16b, SYY15, SSA17, TP16a, Wil19, XZT18, YY17, ZDGW16, ZVO15]. contact- [FFJ^{+}19]. contact-angle [Don17]. Contact-aware [LRZ17]. contact-capturing [Wil19]. containing [LKB15]. contaminant [Har18]. Contents [Ano15a, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano15-35, Ano15-36, Ano15-37, Ano15-38, Ano15-39, Ano15-40, Ano15-41, Ano15-42, Ano15-43, Ano15-44, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano16-48, Ano16-49, Ano16-50, Ano16-51, Ano16-44, Ano16-45, Ano16-52, Ano16a, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano16-37, Ano16-38, Ano16-39, Ano16-40, Ano16-41, Ano16-42, Ano16-43, Ano16-46, Ano16-47, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t]. Contents [Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano16z, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, Ano17-53, Ano17a, Ano17-27, Ano17-28, Ano17-29, Ano17-30, Ano17-31, Ano17-32, Ano17-33, Ano17-34, Ano17-35, Ano17-36, Ano17-37, Ano17-52, Ano17-38, Ano17-39, Ano17-40, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano17u, Ano17v, Ano17w, Ano17x, Ano17y, Ano17z, Ano18a, Ano18b, Ano18c, Ano18d]. context [KGS17]. continental [CS18a]. continua [CEL^{+}18b]. Continuation [BVM^{+}17a, BZ16b, JT18]. continued [Ano16-48, Ano16-49, Ano16-50, Ano16-51, Ano16-52, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, Ano17-53, Ano17-52]. Continuity [MAP17, CRW16, YJ17]. continuity-preserving [YJ17]. continuous [AG16, BKP16, BST15, DGMT17, DKK^{+}18, Fid17, HR18b, HY17, KS16a, KPKGH19, KPKG15, KLSF15, KG15, LKSM17, MSK18, MKS18, MSP16, PL16a, SS16c, VMC^{+}19]. continuous-discontinuous [SS16c]. continuous-in-time [Fid17]. Continuously [Bar18]. Continuum [ISST18, CCP19, CX15, CDX^{+}18a, DKC15, DPRZ16, DPRZ17, GSL18, HS17a, Har18, HKS^{+}16, Jac17a, KLC19, KGP^{+}17, LXSC16, LWX19, MSH^{+}15, NS19a, SSDN15, YSWW16, YXX^{+}16, ZWG17]. continuum-kinetic [Har18]. contour [ZGD^{+}16]. contoured [DKC15]. contraction [EFO19, EFO20]. contrast [FAC^{+}19, KCW17, ML16, RVZB15]. contrasts [BDPM18]. Control [AEL^{+}15a, ABG^{+}15, APP^{+}16, AEL^{+}15b, BMRA^{+}15, CF19, CVG19, DP19, FW18, FDKI17, GM16, KMD^{+}18, KYW^{+}16, KYW^{+}18, KSSL18, LC17a, LYKW19, LM19d, Lot18, NJPB17, Pea15, SPX^{+}18, SWHK15, SPM16, TASA19, VLAB18, WBM15a, YK15, ZILZ15]. Control-volume [AEL^{+}15a, APP^{+}16, AEL^{+}15b]. controllable [ZZH16]. controlled [EMSS17, MRP^{+}15, MCL19, PD15]. Controlling [ZV16]. convection [BLC^{+}17, BGM16, Cai16, CHY16, CY19a, CB18b, Cui15, DY19a, HY15, HHY16, JJ17, Kay15, KS15a, Lap16, LWZ19, LP16b, Liu16, LFT^{+}16, PKLC16, PKLC17, RTO15, SGN16, SL17, Shu16, Sir19, SPZ18, WLM15, WB17, WC18, WSF17, YL19, ZZL19]. convection-diffusion [Cui15, DY19a, HY15, HHY16, LP16b, LFT^{+}16, SPZ18, YL19, ZZL19]. convection-diffusion-reaction [JJ17, KS15a, LWZ19]. convection-dominated [LWZ19, Shu16, WB17]. convective [CARdN19, CCPdL19, Don15a, KTK18, KTK19, MS15b, MBD19, STK^{+}16, tEDKT17]. convective-like [Don15a]. Convergence [FHE15, HD15, HZ15, JPLL15, SAEF17, ADG19, AWJ17, Ata15, CBC^{+}18, GB15b, GSS15a, GPAO^{+}18, GDS^{+}16, GDA16, HR19, KMdB16, KW15b, KDL15, LHA15a, MZAF17, NNW17, Nis19a, PA19, PWP15, SHA16, SVG18, SDJU15, SWZW19, SN19, WTL17, YBSTT19, YJB18, ZNS19, ZHLZ18, ZS17]. convergent [BCC^{+}18, IZ18, NKN^{+}17, OS15a, OLV16, SLH18, Svä15]. conversion [IG15, JLKF17]. Convex [GZ18, CFF18, DF16, EEG^{+}15, IM15, JW16, LM15b, LHGF16, LHGF19, SLL16, SGD18, SLL17]. convexity [DRP^{+}16, GO16]. Convolution [SS17a, VGF16]. Convolutional [WRL19, SZF15, ZZ18, DZR18]. convolutions [Han16, RB18]. ConvPDE [WRL19]. ConvPDE-UQ [WRL19]. Cook [ZYCK15]. coordinate [BDV17, MSA19, PX15, Pru18, SMS16]. coordinates [BtTBI18, CX16, DCCH19, EHXM15, EEG^{+}15, FRW16, HB15a, LBZ16, LMB18, OvdHVH16, PS15b, QSBY19, TLH15, TVB^{+}16, VBL^{+}16, VMM19, VSC18, WWZ19, YFJ18]. coplanar [KW15b]. copolymer [CYS17]. core [CPSF17, Cos16, HBC^{+}16]. Coriolis [ADOP18, LCK19, SD16]. Corner [DBZ17, BMCK15, ZFZL15]. Corner-corrected [DBZ17]. corner-free [ZFZL15]. corners [AKM^{+}19, DCCC16, HK18b, SR16, Tsa16]. corona [VBG^{+}17a]. coronary [BFI^{+}16]. Corrected [CW18, DBZ17, FNNW19, HR18b, LKSM17, Loh17, RMF^{+}18, RSD17, SFDE15]. Correcting [BH16b]. Correction [Kat16, AMN18, ALL18, BLL19, BLL20, BG16a, BDJP19, CWS18, CWB^{+}19, CLX15, CCGH17, DRP^{+}16, DvW15b, DS15c, EH18, GLTB18, GXX17, GSL^{+}19, HX16, HDA^{+}18, HLQ16, HXX18, JLC15, JLKF17, KW15a, KS16d, LLB19, MBD19, PK16, PBC^{+}17, RÖS16, RS17, SMS16, SM16, SW15, Sir19, Vil19, WMYG16, BK17a, MNR19]. correction-based [GSL^{+}19]. correction/finite [KW15a]. corrections [HSM19, WWR16]. corrector [BK16a, NS19b, PHRA16]. correlated [Zau16]. correlation [AKZ16, LT17b, TMWF18]. correlators [BPF^{+}16]. Correspondence [Moc17]. corresponding [STR15]. Corrigendum [ABG^{+}19b, ASS17, BLL20, BR16, CNG17, Dav15, DK18a, DvW19, EFO20, GRT21, GBCF16, Hag25, HGN17a, KYW^{+}18, KTK19, MN17, NG18, PS15a, SZN20, SYOS21, STEK22, SWMD17a, SYV17, TK15b, Vre21, ZCQ20]. corrosion [JS16]. CORS [ZD15a]. cosmic [BPF^{+}16, CL19b]. cosmological [SPM^{+}15]. Cosserat [AMM^{+}15]. cost [CCBdL15, LHMB18, LY15c]. Couette [JL16, SWLW19]. Coulomb [HLL^{+}18, RKH15, TSR15, YC15]. count [HSF17]. counter [ZW15]. counter-intuitive [ZW15]. counterpart [SPRW15]. Counting [GP18]. couple [BMT18]. Coupled [BLS16, CMDL18, FKDL17, GAS^{+}18, QWX18, RTG15, AEL^{+}15a, AEL^{+}15b, AEL^{+}17, BT19, BP18, BK16b, Buk16, BKRB15, CBZ18, CSW^{+}19, CGS18, CWF16, CYS17, COV18, CFPB17, Chi19, CGM15, DGW18, DMAM15, DLM18, DPRZ17, EH18, FFJ^{+}19, GDS^{+}16, GC17, GQC^{+}19, GN19b, HGN17a, HGN17b, HM16b, HCLT19, HHL19, HWA19, JTR16, JGS16, LGH^{+}18, LMKS15, LY16c, LHW^{+}17, LRGO18, MMNI16, MRP^{+}15, MG15b, MNO^{+}17, MMMS15, MDD^{+}19, MKV^{+}17, MPMB19, PF16, QYF15, RSBS19, RRD16, SDM^{+}17, SMOM^{+}17, SF16, SM19b, TH18, TMWF18, TPT16, TT17b, TUJ19, TV19, TPTT18, TC15d, VCEK19, VLP^{+}16, WE15, WED15, XDvW17, YS15, ZL15b, ZZPH18b, ZZPH18a, ZSX17, ZBZ^{+}18, MHL17, SGD18]. Coupling [CFSN18, CFG16, JH15, LB17, MNG15a, MDL16, MTZ16, Wic16, ALKZ16, BCD^{+}15, BKO18, BRK17, CDM18, CLL19, DKPC15, ED16, FH17, FHE15, GN19a, HBC^{+}16, HG17, HLSY16, HF19, ISST18, ID17, KLC18, LPB17, LMC16, LZHM19, LPBR15, LMN18, PCN15a, PHÖ^{+}16, PAL^{+}16, PWP15, PME^{+}15, TKB^{+}15, TAJ^{+}17, VKE^{+}18, WWR16, WPB15, WED15, XYF^{+}17, YG18, ZYK18, ZRE16, dSPDH15]. couplings [FKS19, PJB^{+}19]. Courant [GSK18]. covariance [BCSK17, ZKS^{+}15]. covariances [FDS^{+}15, ZH15]. covered [ELH^{+}16, MDW18, STV19]. covering [PLWJ16]. CPR [CLNH15, ZLFW18]. CPR-MS [CLNH15]. CPU [COdLL18, FJLC18]. cracks [Par15, Par17, Par18b]. Crank [FBF15, HYL17]. Creating [RRM^{+}16, Zau16]. creeping [PZNG15, PGCG18]. crisis [GPS17b]. criteria [GKRB17]. criterion [KTK15, RMP18, TZ16]. Critique [TNB21]. CRKSPH [FRO17]. cross [ABT16, CCFC19, CV16b, Dod17, DJD^{+}17, JDFS16, KFL17, LMGG17, MSF^{+}19]. cross-beam [MSF^{+}19]. cross-section [ABT16]. cross-sections [LMGG17]. crossed [HN18]. crystal [DBD^{+}17, GHL^{+}16, LSL15, LZW19, SLL16, YH17, YC16]. crystals [CSG17, CHL^{+}19, HHZZ19, KLWQ17, NWZ18, PD16b, WZLS19, ZYSW16, ZZW^{+}16]. Cubature [PR17a, LTXB17, vdBKD17]. cubed [IDSG15, KC17a, LP18, YP17]. cubed-sphere [IDSG15, KC17a]. Cubic [LFR17, LT17b, LY15c, LY17, PK17, SP18, SL15, ZYW16]. cubic-quintic [ZYW16]. CUDA [JL18a, PTMF18]. cumulant [GPS17a, GPS17b, SA19]. cumulative [CNQ^{+}19, Hig17]. cure [Rod17, Rod18, SM19a]. curl [DGL^{+}15, LL19a, LYZ18]. curl-curl [DGL^{+}15]. Current [MSV^{+}16, BGV17, BCB15, CCZ15, KE15, MTD15, PYAG19, RBGV15, WMY18, dSPDH15]. current-driven [CCZ15]. currents [AAL15, PK17]. curse [CDOY19]. Curvature [LHA16a, AZ17, BDPM18, CRMP16, CG16, EDvW17, IM15, LAA16, OD15, OCSC18, QLS^{+}19, Vog17]. Curvature-Augmented [Vog17]. curvature-inducing [LAA16]. curve [WTL19]. curved [BD18, Bou19, CE17, Chu17, CHD^{+}18, FB17, FP16, GSN16, GA18, HHY16, LD15, NN16, Pas16, RRD16, WXSJ19, YK19, ZL15b, Zha16]. curves [Wal16]. curvilinear [ÅN19, AB17, BDV17, BC16c, CSW^{+}19, CC16a, CC19a, CW19, CTG16, CX16, DWR18, DCCH19, EHXM15, JG15, LM19a, PS15b, QSBY19, SSVL18, TLR16, VMM19, WR15, WWZ19, WWGK17, WWGW18, YFJ18, ZH19]. Cut [ÖPHA15, BNK18, CWJ18, DM16, FZ19, GNK18a, GNK18b, GEZK16, LSD^{+}17, MM16d, MM18, PHÖ^{+}16, SGMS16]. cut-cell [BNK18, LSD^{+}17, MM16d, MM18, PHÖ^{+}16, SGMS16]. Cut-element [ÖPHA15]. cut-stencil [GEZK16]. CVD [AEL^{+}17]. CVFEM [Dom18]. CVFEM/DG [Dom18]. cycle [SPP^{+}16a]. cyclic [TBHG18]. cycling [SPCH16]. cylinder [AA19]. cylinders [BPGS16, HK19b]. cylindrical [CJH^{+}19, KS15b, OVP15, OvdHVH16, SCQP16, TLH15, VSC18]. cylindrically [MTD15, MDT16].

D [AMXJ19, CZ17, CSK^{+}16, DWGW16, DSSP18, MSF^{+}19, PS17, Sto16, TCS16a, TRL15, VLP^{+}16, YDLC19, ZJS15, AG16, AHHC18, ACS16, ALTR17, BHZ16, BK17c, BGV17, BDK^{+}17, BS15a, BT16, BSM16, BC16c, BCRS19, CDLR19, CBC^{+}18, CGMH18, CDL17, CGL18, CGK17, CDL19, CCZC16, CZ16, CX16, CSC19, Cho19, CSK^{+}16, DBD^{+}17, DMRB19, DF16, Dod17, DD16b, EL18, EDW19, FDS^{+}15, FL18, FNGV18, FNGDMNR18, FGLB16, FC19a, FYC^{+}18, GBM16, GFL17, GWC18, GMWC19, GHL^{+}16, HWH^{+}16, Hu17, HLCL19, IG15, JKE^{+}17, JBLO15, KQB18, KE15, KES18, KC17c, KFWK17, LMPS15, LM15a, LFRH17, LML^{+}16, LGB17, LHMB16, LZ17a, LLJJ18, LZHM19, LZT^{+}15, LY16a, LGD17, LZL^{+}19, LFAR19, LMG19, LMC19, MKYZ17, MG15b, MC15, MBS19, MF16a, MW17b, MFF^{+}19, Mue18, NMM18, NMM19, Noe15, Nor15, PG17, PGCG18]. D [PK17, PKJ^{+}18, PCMC19, PR16a, PMGW16, PAFT19, PTT18, QDH15, RBY19, RLP16, SNSG16, SFT16, ST18b, SCS16, SA15, ST18c, STV19, SSL^{+}16b, TCD17, TB19, TCL15, VLP^{+}16, WY17, WXW15, WHRL19, WSH^{+}17, WSU^{+}15, XDSX17, XQ17, YC17, Yam19, YSWS16, YFJ18, YPC19, YTW15, YXD^{+}16, YPK16, YT19, ZBH^{+}18, ZND16, ZGJ16, ZZZ17, ZJ18, ZSL^{+}19, ZVO15, ZYCK15, ZZW^{+}16, dBIM16, dJRP^{+}15]. D- [TCS16a]. D-TDIBC [DSSP18]. D-VAR [FDS^{+}15]. D/ [CSK^{+}16]. dam [GWYS18]. dam-break [GWYS18]. Damage [CF15, BHJ15, HMBH15]. damped [CZW17, YJ17, DFM17]. damping [CGTH18, HSC16, NNW17, ZC19a]. Darcy [AEL^{+}17, BMT18, GMT19, KLGO18, LTW18, MTZ16, Noe15, STG17, Sun19a, TV19, WSN^{+}18, XML17]. Darcy-flux [AEL^{+}17]. Darcy-scale [Sun19a]. d’Arolla [AS17]. Darwin [CC19a]. Data [AMPG19, BN19, BV18, GK19c, GLZ19, LK17, LSD18, LS19b, QWX19, SG16, TMES19, YQNW19, Yeo19, ZYK18, ATM^{+}18, AKK^{+}19, ACC^{+}15, AM19, ADP^{+}17, AÁPB17, AZ19b, BFP18, BFM19, BGG16, BDMZ19, BHJ15, CR17, CJK^{+}19, CGM18, CZ19a, CCM15, FG18, GS15c, GM16, GBvZB16, GLG^{+}19, IPSG15, KL17a, KYUO15, KZG16, LSP19, LW15a, LZB^{+}17, LHMB18, LLD19, Lor19, LBB^{+}17, MM15, MP17, MCGS16, MWZ19, MBA19, NKN^{+}17, OY19, PPCK17, PD16a, PLB18, PND16, PF15, RPK17a, RS16a, RRM^{+}16, RVMR17, STHW17, SG18, SR19, SWS17, SD17, SWX18, Sla16, SWHV16, SSN15, TBLM15, WWZ19, WX19, XWW^{+}16, XY18, YNW17, YBSTT19, YM19, ZZ17a, ZWB^{+}18, ZZKP19]. data-assimilation [CZ19a]. Data-domain [ZYK18]. Data-driven [BN19, BV18, GLZ19, LK17, LSD18, SG16, TMES19, Yeo19, AZ19b, BFM19, CZ19a, LSP19, LZB^{+}17, LHMB18, MBA19, PD16a, XWW^{+}16, YM19]. data-model [YBSTT19]. data-sets [STHW17]. databases [ATF16, SG17]. datasets [BN19]. DCT [MCHL16]. deal [BDJP19]. dealiased [RB18]. Dealiasing [MDM^{+}15, WMM^{+}18]. Debye [GHJ15, MG15a]. decay [LWY17]. decaying [Bra16c]. December [Ano19a, Ano19b]. decentered [Fal15]. decoder [ZZ18]. Decomposition [APLK19, JHPAT17, AH15, AABD15, AA15, BBB15, BLK15, BMT18, DA17, DGL^{+}15, ETAG15, FHY^{+}19, FHA17a, GM19, GJL19, GFW16, HLS15, Jer19, JX17, LH15, LLS15, LWZ19, LYA16, LMGG17, MN18a, Mag19, MBST17, MCS16, PLL15b, RO16, RTV17, RBD17, RMBN18, RKB19, SLB^{+}19, SW18b, SL16b, SC18b, SWZ17, TCA16, TWF19, Tav16, TT17b, TST17, TSST16, VSM17, WH16b, ZFPB16, ZOY^{+}19, ZYCK15, AM19, CJC19]. decomposition-based [WH16b, ZYCK15]. decomposition-synthesis [MCS16]. Deconvolution [VV19, WSN^{+}15]. Deconvolution-based [VV19]. Decoupled [CBZ18, CJYZ15, FLV15, OD15, PKLC16, PKLC17, SDW18, ZYSW16]. decoupling [CFG16, QWXZ17]. Deep [BFM19, RKB19, YM19, AZ16, CJK^{+}19, DGW18, FS17a, GZ19, GHF19, HZE19, LZSG19, LLD19, PLL15b, QWX19, RPK19, SS18b, TB18, ZZ18, ZZKP19, TB18]. deep-water [FS17a]. deexcitation [YCBC15]. defect [HHZZ19, RSD17, SM16, SWZ17, WMYG16]. defect-correction [WMYG16]. defects [GP18, GBS15, PD16b, TZSS17]. deferred [CLX15, CCGH17, GXX17, HSM19]. defined [WDGW17]. definition [CP16, MBS19]. definitions [TMdO19]. Deflation [JTD16, SLR^{+}16, vdLJLV16]. Deflation-accelerated [JTD16]. deflection [DHH^{+}18]. deformable [LRGO18, PME^{+}15, SMA^{+}16, SMOM^{+}17, YM17b]. deformation [ANL^{+}16, FHY^{+}19, FRW16, GBvZB16, KAR17, LY16c, MTL^{+}17, SKS17, WQZ15, WY19, WCWY19, WTS^{+}17, YG18, ZL15c]. deformations [AQ19, GSL18, GLS15]. deformed [QYJ19]. Deforming [SYV17, LHB^{+}16, LY15b, MMSS15, NN15a, NN17, NN19, RB15, SYV14, ZP16]. degenerate [AHZ19, BTA17]. degree [Bre18]. DEIM [SSN15]. delaminations [GD19]. Delaunay [WQZ15]. Delayed [DSSP18, AC16]. Delayed-time [DSSP18]. delta [EG17, HNS16, OM19, XM18]. Delves [HB15b]. DEM-IB-CLBM [ZZPH18b]. Demonstration [WSK19]. dendrite [RTO15]. dendritic [CY19a, DMS17]. denoising [CWL^{+}16]. denominators [HPV16]. dense [AAD16, BVM17b, DAO17, Han19, KBK15a, RO19, SG19a, SYM15, WZR15]. dense-to-dilute [DAO17]. densities [BCST17, YD19, YY17]. Density [GS16, AKZ16, BVM^{+}17a, BEJ15, BKKY19, BC16b, Cai16, CZL^{+}15, CNQ^{+}19, CVK16, Cif19, CDV17, DLY17, EJMI18, FB17, GZLH19, HK19b, HW18, KP15a, KLC18, LMH16, LTB16a, LL16a, LT17b, NPB19, NGPB19, NGY^{+}17, NSK^{+}16, OM19, PAFT19, PLB18, RFGSV15, SP15a, SHP^{+}16, SK18, TASA19, TKF17, WSY15, WSS^{+}15, WSHT15, WDT^{+}19, WY19, WCWY19, WSF17, XDSX17, ZLH^{+}17, Zil15]. density-stratified [Cai16]. density/viscosity [WDT^{+}19]. dependence [FW17]. dependent [AL19a, AL19b, AR16b, AWJ17, BHL15, BST19, BOA17, BCB15, BSWG15, CR17, CHY16, CX16, CDOY19, CGJ19, Chu17, CLP16a, DD16a, DKPC15, DBMB15, FM15, GSN16, Gan15, Gen11, Gho17, GKNA17, HPC19, HL16b, IKS19, KBK15b, KFL17, LZ15a, lLNS17, LR19b, MCN18, MMMS15, OKWE17, PLC18, PCMC19, PHRA16, RDQ19, RRD16, STEK17, STEK22, SS15b, Shu16, SP15b, Sub15, SJXL15, ZTBW19, Zha16]. depletable [SN15]. depletion [GLTB18]. deposited [AASRT17]. deposition [MZ15, TT17a, TP16b, Zoh17]. deposits [JS16]. Depth [NMJFM19, DDM^{+}19, DV17]. Depth-averaged [NMJFM19]. depth-independent [DDM^{+}19]. Derivation [GPS17a, RSSSE18, Sch16a]. derivative [AHKT17, CF15, CGM18, Cha16, DBZ17, DLC15, DYL19, DZC16, FC16, HL16b, JW15b, Kat16, MBSS15, NDCB17, OM15, OLB^{+}17, Par17, RCRF16, Roy15, TWF19, TVB^{+}16]. derivative-based [TWF19]. derivative-free [FC16]. derivatives [ARTG^{+}19, BKP16, CDLR19, CZ16, CZ17, DMM19, GMS19, GGT15, pHzSrC15, Mac15, MD17, MN04, MN17, SZM19a, SZM19b, TMdO19, ZzSK15]. derived [JL16]. deriving [DC18b, KBD19]. descending [XL17b]. descent [FSWW17, MH18b, TP16b]. described [CF15]. describing [AMM^{+}15, WT19]. Description [ALKZ16, DTA^{+}15, RZOZ19, SG19a]. Design [BTVC16, Dom18, DMM19, FBC^{+}16, GM16, GSS^{+}19, TCS^{+}16b, BDB^{+}17, BHNS19, CC16b, GGW17, KL17a, Kou16, NP16, NW15, RPC^{+}18, WLL16, WHL17, vLtTBI17]. Design-order [Dom18]. Detailed [Did17, MHGM^{+}15, JD19, LMY^{+}19, LSYF15, MA16, VLP^{+}16, XXR18]. Detecting [RH19]. Detection [ACC^{+}15, ABT17, EEG^{+}15, LR19a, AKK^{+}19, CW16, CW17, Gno17, JdR^{+}18, KLA17, LLW19, PDN19, PQR17, WLL16]. Detection/Rankine [FKK19]. detector [LSI16]. deteriorating [PT17a]. Determination [HK16b, EZG16]. determine [LGZ^{+}19]. deterministic [Dav10, Dav15, HPC19, PCMC19, RMC15, SS15b, TAJ^{+}17]. deterministic/stochastic [TAJ^{+}17]. detonation [Hu17, RA17, RA19, WDS15]. detonations [DY19b, ZWL^{+}19]. develop [LSP19]. Development [AKZ16, BV15, BLG^{+}16, BBK19, CBC^{+}18, DDJ17, KYPK15, RBY19, TM17, YSWS16, ZLGS18, Ani16, CYYL18, LPR19, Shi19]. Developments [IC17, PMF^{+}18, Shu16]. deviational [Yan16a]. device [BPD19, FKF17]. devices [BLL16, LDO^{+}19, NOM^{+}17, RKH15, WPB15]. dewetting [ABG^{+}18b, ABG^{+}19b, BJWZ17]. DFFD [CH17]. DFM [BHTT17]. DFT [BW18b]. DG [AW18, BBF^{+}17, CFSZ19, CEHM19, Dom18, FG17, HCB19, JLQX15, JL18c, NJ15, RXS16, Shu16, TABR17, XQ17, YL19, SL17]. DG-IMEX [CEHM19]. DG-schemes [FG17]. DGFEM [KL15]. DGM [SS18b]. DGSEM [Ren19].

DGTD [TRL15, BK19a, SSL^{+}16a, SSVL18]. Diagonal [Mat17, DBZ17, KNS15, MAvdW18, MO18b, SS16b, WLK^{+}16]. Diagonal-norm [Mat17, DBZ17, MAvdW18, MO18b]. diagonalization [LWLC17, PKA^{+}16]. diagonally [BZ18, CCGH19, HK18a, NMC15]. diagonally-implicit [BZ18]. diamond [AA15, JKE^{+}17]. diamond-difference [AA15]. diatomic [WYLX17]. diblock [CYS17]. dielectric [GWB^{+}15, HK19b, LSP^{+}18, PKLS17]. dielectrically [RMLvR18]. dielectrics [MG15a]. diffeomorphisms [CRW16]. Difference [NWFT19, SYV17, ÅN19, AD17, AW18, Ali15, AA15, BBKS16, BH16a, BS19a, BH18, Bra16c, Bre17, BTT18, BTWY15, CBS18, CLC16, CTG16, Che18, Cho15, CR18, CYWL17, DWZ19, DLK17, DvWZ18, EN18, Fan16, FGLW18, GSS15a, hGwSzS15, GS15a, GH17a, GK19a, GS16, GT18, GQC^{+}19, GHL^{+}16, GL17, HZL^{+}15, HF18, HAH16, JKE^{+}17, JW15b, JSB^{+}19, KW15a, Kay15, KS15a, KJYC17, KL17b, KPJ18, KWHB19, LH16, LYC16, LHMB16, Li17, LMBZ15, LWWY18, LYZ15, LY15b, LN15, LMMS16, MN04, MN17, NN15a, NF17, OLDN17, OS15a, OV17, PR19, PS15b, PS17, QWZ19b, RBI18, RLGT19, RWN18, Sha17b, SF18b, SG19b, SYV14, SZ17, SLA^{+}19, SK18, SN19, TLH15, TBO^{+}16, TKP16, WLM15, WH15, WDS15, WH16a, WLGD18, WLW^{+}18, WT15, WA18, YYL16, YHQ15, YLA15, YX15, YM15, YWHP15, ZZK16, ZL15b, ZG18a, ZSQ17, ZQ16b, dFJN16]. Difference [GSS15b, Mas18]. difference-boundary [BBKS16]. difference/embedded [Cho15]. difference/finite [BTWY15, ZG18a]. difference/spectral [CLC16]. Differences [BHJ18, ABFR16, CZL^{+}15, CPS17, FBW16, KBD19, LTB16b, MF17, TRLK18]. differencing [DvW15b, FAZ16, LMG19, PSG19, TK12, TK15b, WJD16, WBM15a]. different [LCK16, OTS17, TBHG18, YD19]. differentiable [Bar18]. differential [ABDN19, AD17, ADH^{+}16, AEAM15, Beg15, BZ15, BSWG15, BR17, BOD19, BT15, CGS18, CSD19, CAA18, CXH15, CC19c, DLL^{+}17, hGwSzS15, Gno17, GN16, GXX17, HO15, HBR15, HZ15, JW15c, JX15, JX17, KNS15, KR17, KB19, LYC16, LL16c, LLH19, Lor19, MS16a, MR16a, MPR^{+}18, MTK^{+}16, MSA19, MTBT18, NYNYM15, NBH18, Opp17, Pis18, PF15, RPK17a, RPK17b, RK18, RPK19, RMP18, RDQ19, SR16, SS18b, SLN15, Sub15, Sub18, Sun19b, TY17, TST17, TO15, VCNGP15, WRL19, WZ18b, XY18, XHC15, YHKPF17, YDN19, YJB18, ZTBW19, ZHWQ18]. differentiation [CWL^{+}16, FS19a, LAK^{+}16, YCPD15]. differentiator [SZF15]. diffraction [CDL19, CDDL19, HN17b, ZED15]. Diffuse [FB17, PN18, ZDGW16, CSN17, De18, KS16c, KS18b, LD15, MA19, NFG15, TDC^{+}19, WSS^{+}15, ZFL^{+}19]. diffuse-interface [KS18b, LD15, MA19]. Diffusion [BSWG15, CNK19, LLS15, Ali15, ADHN15, ACJ17, AHZ19, AHKT17, BJO18, BL18, BBW16, BDBEE15, BGHK19, BFT17, BTVC16, Cac15b, CG19, CKK18b, CNOS15, CLC16, CHY16, CLZ18, CLZZ19, CLR15, CG15, CCM15, Cui15, CwYjS16, DS15a, DS15b, DD16a, DMSC16, DJL^{+}19, DY17, DYL19, DY19a, DB18, Fal16, FNNW19, FBF15, FHE15, FLW19, GSS15a, GS15a, GPS17a, GPS17b, GBU15, GLW18, GL17, HG17, HSC16, HY15, HHY16, IZ18, JPLL15, JW15b, JW16, JZ16, JLLZ15, JJ19, JJ17, Kay15, KS15a, KSM19, KKLS17, KBK15b, LE16, LAL18, LP16a, LPB17, LW17c, LS19a, LWZ19, lLLNS16, LZ17b, lLNS17, LMMS16, LP16b, LCCZ19, LLH19, LM19d, LM15c, LLLN18, LFT^{+}16, Luc15, MBSS15, MMNI16, MD18, MS18b, MIM^{+}19, MK15, MN16a, MM15, MP15a, MMvR18, MDDM17, MSP15, MSP16, MW15, MN16c, NJHL18, NN18, NL18a, Nis18a]. diffusion [Nis18b, O’S19, OADN19, PD15, QDH15, QN19, Rag15, RRL19, RB15, RZ18, SAEF17, SWG^{+}17, SF18b, SY16, SYM15, SYM17, Sir19, SMD18b, SSM15, SX15, SGA^{+}15, SPZ18, SPRW15, SDW18, SLZ^{+}17, TWN15, TW17, TK15a, TSH17, TMT17, WZ15, WY16, WW17, WHY17, WHY18, WCL15, WZ17, YZZ19, YHQ15, YYN^{+}17, YM17b, YLA15, YL19, ZSP15, ZSW17, ZG18a, ZZL19, ZJL16, ZLL^{+}17b, ZC19b, ZSO^{+}19, vEKdB16]. diffusion-controlled [PD15]. diffusion-limited [BL18]. diffusion-reaction [FNNW19, MN16c]. diffusion-wave [BJO18, BDBEE15, HSC16, YLA15]. diffusive [AJVH17, BHdD18, BR15b, BR16, BLC^{+}17, CCDL19, JLQX15, JXZ15, JL17c, Liu19b, MDH19, MP15a, MFF^{+}19, SAOW17, VDPP15, ZHWQ18]. diffusivity [HK19a, YL19]. DIII [WSU^{+}15]. DIII-D [WSU^{+}15]. dilatancy [MDP18]. Dilute [LWX19, DAO17, SGP17b, Yan17]. Dimension [CLM16, TLQ16, AS15, CQ15, Lan19, YM17b]. dimension-adaptive [CQ15]. Dimension-by-dimension [TLQ16]. Dimension-independent [CLM16]. Dimensional [CJC19, NN18, AR16a, APR^{+}15, AEL^{+}15a, AEL^{+}15b, AB16b, APT17, An17, AZ19b, ADOP18, Bal15, BVG^{+}16, BOA17, BH16b, BGL^{+}17, BA15, BH18, BLS16, BVT18, BGG16, BGJ^{+}15, BHMS18, BTWY15, CJWS19, CB15, CCZ18, CQ15, CP16, CZJ17, CLZ18, CM19, CHJT17, CVK16, CLX19, CHL^{+}19, CGJY19, CGP16, CM18c, Cot18, CLMZ17, CYWL17, DCA^{+}16, Del15, DvW15b, DZ16, DHH^{+}18, DvB17, EDvW17, FDKI17, FS17a, FST15, FNNW19, FPDT17, FK17, FLW19, GSC19, GMD19, GIF18, GS18, GGL^{+}17, GN16, GT19, GK19c, Gri19, GQC^{+}19, HPC19, HTFL18, Hiv18, HHL19, HYL17, Hue15, IGQ15, IDSG15, IM15, JL18c, JGS16, JJ17, JSY15, KF15, KA15, KCW17, KHTZ19, Kou16, KS15b, LGO17, Lan19, LLL16, LPR18, LXC19, LL16c, LL19b, lLLNS16, lLNS17, LDT19, LD15, LSTkM15, LK16a, LW17d, LW17e, LMSK17, LCK19, LZS^{+}19]. dimensional [LCCZ19, LLH19, LEB^{+}17, MN18a, Mag19, MIM^{+}19, Mar19, MHL17, MDDM17, MTM19, MBM^{+}15, MB15, MLB16, PST19, PxRS17, PHHR17, PK16, PCN15a, PCN15b, PR16b, PF15, QWZ19b, QSBY19, QPK19, RNO19, Ram17, RG15, RS16a, RH19, RDG17, RKRGW17, RXSG15, RXS16, Rod18, RLH19, SV19, SG18, SSVL18, SD17, SWJG19, SSA17, SX15, SSN15, SF16, SWZ17, SK15b, SLZ^{+}17, TCSM15, TCS17, TCA16, TWF19, TD16a, TSH17, TZSS17, TBO^{+}16, TSB^{+}18, Tre16, TEP19, TBG16, TB18, VCNOP18, VNA15, VSM16a, VSM16b, WSY15, WDS15, WCN15, WRL16a, WRL16b, WTGC16, WHY17, WLE17, WHE17, WWGK17, XML17, YSW15, YSYW19, YK18, ZMF15, ZZK16, ZTBW19, ZL15a, ZLL16a, ZYW16, ZBZT17, ZCL17, ZL15c, ZWB^{+}18, ZZKP19]. dimensionality [AZ19b, BGG16, CDOY19, TBG16]. dimensionally [GNK18a, GNK18b]. dimensions [BHJ18, BHST18, BXY17, CC16a, CM18b, CB18b, CGRV17, DS15a, DS15b, DL17, Ein19, ECC18, FR18, FS16, GR19, HN17a, KSVB18, PES19, RVZB15, SHKL16, Vee16, WK19, WCT18, ZYD19b]. Diminishing [SIX16, CCDL19, DSX19, DWG^{+}18, DLMDV18]. diodes [DS15d, JB15]. dioxide [GGL^{+}17]. dipolar [WNW^{+}19]. Dipole [MML17]. Dirac [ASS17, ASS13, Alm19, AL19a, EG17, FGLB16, HNS16, KML18, Pin15, PS17]. Direct [BLD15, CR17, FKY15, KLNH17, LRA17, MTT19, OMYvdP^{+}15, Par18b, PGGW18, RW15a, SAK18, YK19, ZN18, AA19, ABT19, BS15a, BCM19, BB19, CDC17, CHY16, CYL^{+}16, CYYL18, CC16c, CGP16, CWJ18, DY16, EGO19, Eva18, GB15b, IM17a, JL19, KNS15, LO19, LCLY19, LWTF19, PPLC16, PG18, PVB17, RS16b, RLV16, STK^{+}16, TFGK18, YS15, ZG18b, ZCL19, ZMZC19, Mac16]. direct-forcing [LWTF19, PVB17]. direction [BCG^{+}15, GGT15, LZT^{+}15, LK16a, SX15, SZ17]. directional [BNK18, FYO^{+}15, HKA19, MHJ15, MSH^{+}15, ZYK18]. directional-splitting [FYO^{+}15]. Directly [ZQ16a]. Dirichlet [ABN15, ED16, GBD17, KHHN16, VMC^{+}19, WZ15, YK15]. Dirichlet-to-Neumann [GBD17]. disc [SHW17]. discharge [DBMB15, VBG^{+}17b, ZCHS15]. disconnected [GT19]. discontinuities [GLTG15, HZL^{+}15, PMS19, WS15b]. discontinuity [DS15a, DS15b, DIX^{+}18, PE16a, PDN19, RSBS19]. discontinuity-aware [DS15a, DS15b]. discontinuity-resolving [DIX^{+}18]. Discontinuous [BHGK18, BD17, BKRB15, BKKRB16, CG19, FNP17, HGN17a, JHT^{+}18, KCHW19, NLW^{+}16, OWKE16, Rag15, TSC17, TRL15, ZX19, ZK18, ZN16, AG16, AM17a, AGC19, AB19, AS15, APKP16, ADK^{+}17, BMMP19, BST^{+}18, BDM17, BGD19, BCJL17, BFT17, BCB17, BD18, Bou19, BT15, CGQ18, CPX19, CGMH18, CWM^{+}16, Cha18, CW19, CJD^{+}17, CHY16, CS17a, CYL^{+}16, CYYL18, CZL18, CL19a, CCKQ15, CLG^{+}19, CXY19, CR18, CK16a, CK16b, CCGH19, DM17b, DKK^{+}18, DLL^{+}17, DY19a, DL16, Ein19, EHXM15, EDC19, FWK17, FWK18, Fer17, FX18, FBM16, FSB16, FS17b, FS19b, GWK16, GCVMK15, GBC^{+}18, GSN17, GX15, GY15, HR18a, HKA19, HL16a, Hig15, HS18b, Ism15, JAH19, JSB^{+}19, JH17, JJ19, JTD16, KCSW19, KDF15, KM16b, KFF^{+}17, KRFV16, KG15, KFWK17, LMH16, LLP^{+}16, LP16a, LPR18, LX18, LC19, LHLL19]. discontinuous [LSR16, LM19a, LM19b, LTB16b, LP16b, LY16b, LW17e, LSZ18, LMB18, LMB19, LFAR19, LCCZ19, LLLN18, LMG19, LHL15, LHQ19, LI15, LSI16, MSG18a, MLM18, MRRRF18, MK17, MN16a, MSP19, MKC17, MF16a, MLB18, MSP15, MSB^{+}16, MMPS17, MH17, NdlLPCC19, NMM17, NJ15, NPC15, NPRC15, NDCB17, Nis18b, OADN19, OKWE17, OKE17, PL16a, PA19, PE16a, PP19, PCN15a, PP17, PP18b, PMB18, QWZ^{+}19a, QSY16, QDH15, RXSG15, RdM19, RDM15, RRMF^{+}19, RBL16, SPX^{+}18, Say17a, Say17b, Sch16b, SWG^{+}17, SMP16, SL19b, SZ15b, SS16c, SPZ18, Sti16, SWZW19, SCS18, TH18, TD16a, TD17, TD18, Teu16, TM15a, TXKvdV15, TXKvdV16, TLB^{+}18, UL16, VPV^{+}17, VCNOP18, Vil19, WW15, WTGC16, WLE17, WWGK17, WWGW18, WG15, WMM^{+}18, WBM^{+}15b, Xia15, XOX19, XJLQ15, XL16, XYG19, YY16, YK19, Zha16, ZLH^{+}17, Zha17c, ZF18]. discontinuous [ZCL19, ZY19, ZCQ19, ZCY^{+}19, ZCQ20, ZT17, ZYD19b, dFVJ15, vOMB17, HGN17b, OLHD17, PSB^{+}18, DDM18, RHS18]. discontinuous-Galerkin [NJ15, Sch16b]. Discovering [PPCK17]. discovery [BN19]. Discrete [ACGR15, BNS17, LMPS15, LPG18, MHS16, SP18, WYZZ18, WWZ19, AEL^{+}17, ADHN15, BCST17, BBB^{+}16, BPS16, BC18c, BSP18, BHTT17, CFKK19, CC17c, Che19, CVG18, CwYjS16, Del15, DWGW17, EFHZ17, EDW19, FNNB19, FKS19, HLML17, HCVH18, HHY15, Hwa16, JLQX15, JKE^{+}17, LFRH17, LC15, Loz17, MWD16, MRM16, Mas18, MZ15, NBT19, NMA15, NHA18, NN15a, NN17, NN15b, OM19, OWKE16, OKWE17, PL16b, RW19, SSDN15, SVG18, SMG19, SWK18, SGL17, SDW18, SLZ^{+}17, TZGW18, TAH16, VLTPS16, VBF15, Xia15, XRMM15, YSWS16, YSYW19, ZTBW19, ZW19, ZNX15, SMAG17, dPSS16]. discrete-adjoint [VBF15]. discrete-forcing [LC15]. discrete-ordinates [Mas18, RW19]. discrete-time [BSP18, MWD16]. discrete-velocity [HLML17, JLQX15]. discrete/continuum [SSDN15]. discretely [Cha18, CW19]. discretisation [ABP^{+}16, AGC19, DXvW18, DvW19, GBD^{+}15, OLHD17, OWKE16, SSM15, Smi18, TFGK18, DDM18]. discretisations [BMMP19, MRRRF18, OKE17]. Discretization [BS19a, CF19, Dav10, Dav15, FPDT17, AD15, AVT17, BHE^{+}17, BFNGDNR18, BKRB15, CDM^{+}16, CGS18, CP17, CM15, CHD^{+}18, DvB17, DS15d, DCD^{+}18, DL18c, EG17, EDK19, FNGDMNR18, FW17, FKS19, GZLH19, GDA16, GSMR18, HR18a, Her16, HLML17, HK15b, JCWX19, KL18a, KML18, LMMS16, MSK18, MMvR18, MCS^{+}19, MHS16, MMP18, Nis15, NL17, Nor15, OvdHVH16, PG17, PG18, DM18, QLF16, RBL16, STK^{+}16, SKF15, SUR18, TCS17, Vel19, VDPP15, VBG^{+}15, VK16, YP17, ZP16, ZZKF15]. Discretizations [SYV17, BGGM15, BCB17, BSM16, CJK^{+}19, CHOR17, DK19, DCCH19, FKF17, FWK17, KCHW19, KD17b, MXL16, PE16a, RN18a, RN18b, SF18b, SLVE18, SYV14, TMH16, VLN^{+}18, WX17, ZNS19, ZSX17]. discretize [DBMB15]. discretized [HR18b, JW15c, SWG^{+}17]. Discretizing [POSB16, SP18]. discs [GPAO^{+}18]. disk [ZG19]. disk-like [ZG19]. dislocation [BC18c]. disordered [SU15]. disparate [TCS16a]. dispersal [Har18]. disperse [JS17, LWX19]. Dispersion [BGGM15, EL18, EL19, JSB^{+}19, Mel18, SL15, An17, CSD19, CHLZ17, GZY16, GR15, HK18a, JLC18, KMS^{+}18, KD17b, LKN17, MRN16, MT17, MHZ^{+}15, MSP15, NMC15, PCF15, PPCK17, Ram18, SSL^{+}16a, Sto16, URGT18, WA18, YWHP15]. dispersion-diffusion [MSP15]. dispersion-relation-preserving [YWHP15]. Dispersive [SU15, ABH^{+}19, AEAM15, EDC19, Iwa15, LM15a, LXC19, DM18, SSVL18, ZWUR16]. Displacement [RVMR17, BST^{+}18, LW17a, RSBS19, SWML17]. displacements [BQCG17, CXY19, RDG17]. dissipating [CG18b, QWZ19b]. Dissipation [CZW17, JT18, BR15a, BMCK15, CSD19, DLLV17, DWGW17, EMM^{+}18, HK18a, HWA15, JLC18, KCS^{+}17, KYW^{+}16, KYW^{+}18, KV16, LHO^{+}19, MGCW18, NMC15, SMD18a, SL16c, TWN19, UY19, WDGW17, WL17, ZHA17a]. Dissipation-based [JT18]. Dissipation-preserving [CZW17, SL16c]. Dissipative [YG19, AMH^{+}18, Abg18a, AF18, DPK17, DJL^{+}19, JKM19, KP15c, LS15b, LS16a, LBTK18, MD17, MHT^{+}19, MBM^{+}15, MFG15, PLL^{+}15a, Sto17, YDCK16]. dissociating [WMS18]. Distance [XL17b]. distorted [Nis19a]. distortion [TAR17]. distributed [AEL^{+}15a, AEL^{+}15b, CPT16, CLC16, FG16, hGwSzS15, GMS19, LAA16, MR16a, WLC15, WX18, YLA15]. distributed-order [hGwSzS15, YLA15]. distribution [AD15, AB16a, EG17, FL18, GMLD18, GMD19, HNS16, iI15, iI17, IC17, LN17, MN15, STR15]. distributions [BC18a, GWE^{+}15, LL15, MTM19]. div [LYZ18]. Divergence [Ama15, BD15a, CZBC^{+}18, BK17b, BDG^{+}17, DWG^{+}18, KBR17, PMF15, RRM^{+}16, TPB16, XL16, YJ17, YFJ17]. divergence-cleaning [YJ17]. Divergence-conforming [CZBC^{+}18]. Divergence-free [Ama15, BD15a, BK17b, BDG^{+}17, RRM^{+}16, XL16, YFJ17]. divertor [MP16, TTN^{+}16]. divertors [BDB^{+}17]. DLM [NPB19, PZNG15, PGCG18]. DLM/FD [PGCG18]. DLM/FD/IB [PZNG15]. DNS [HW19, KCS^{+}17, KP15c, KFWK17, MA16, MMPS17, RL17, SLC^{+}18]. Do [RFGSV15]. docking [PLWJ16]. Domain [IBML16, JHPAT17, JX17, TRL15, AM17a, And16, AA15, BFF19, BLK15, BMT18, BGD19, BG16a, BCJ19, CXH15, CLC16, CC17c, Che18, CLQ17, DZ16, DZ18, DvB17, DDV^{+}15, DGL^{+}15, DSSP18, ETAG15, FHA17a, GM19, GFC18, GBD17, GHJ15, GHH^{+}16, HXLL15, HGW18, IML15, JSP16, JCWX19, KPJ18, LH16, LS15a, LH15, LLS15, LZ16, LHMB16, LHLL19, LC17a, LHY^{+}19, LZT^{+}15, LK16a, LMM17, LYA16, MS18a, MS18b, MMSS15, MJ17, MMP18, MH17, NBT19, PR16a, PR19, PLL15b, PT17a, PBA^{+}15, QDH15, RW19, RZ17, STFK19, SZW^{+}16, SW18b, SZ17, SMSR18, SL16b, SC18b, SWZ17, TT17b, TP17, TST17, WR16, ZP16, ZLY15, ZD17, ZYK18, ZYCK15, ZBZ^{+}18, ZZH16]. domain-decomposition [TT17b]. domain/active [LHY^{+}19]. domain/active-strain [LHY^{+}19]. domains [AB16b, ABFR16, And16, ABG18c, AEvW19, BLS16, BTT18, BC16c, CLZZ19, CFF18, DGHP17, DH18b, ECC18, FH17, FYZ^{+}15, FBF15, FLT18, GSN16, GLS15, GN16, GT19, GLTG15, HK18b, JW16, JGS16, JTD16, KADE15, KADE17, KBR17, KJ17b, LPGT16, LB15, LCK16, LC16, MMNI16, MTZ16, MS18c, MSA19, NN15a, NN19, NGY^{+}17, NSK^{+}16, NN16, OLD^{+}16, PKN17, RB15, ST17, SHW18, SGT16, SGT17, Tow18, Tsa16, WRL19, YYN^{+}17, YG19, YDCK16, YLA15, ZL15b, ZG19]. dominated [LWZ19, RZOZ19, Shu16, WB17]. Doppler [DJD^{+}17, JDFS16]. dosimetry [KSV^{+}15]. Double [LH16, BLC^{+}17, CSC19, EG16, JCNK19, TB19]. double-diffusive [BLC^{+}17]. double-slit [CSC19]. double-sweeping [EG16]. Doubly [YYL18, BLS16, Cho19, GD19, HTFL18, HN18, LB16, NL15]. doubly-asymptotic [BLS16]. doubly-periodic [Cho19, HTFL18, LB16, NL15]. down [CLL17]. DP [KCW17]. DPD [GZM^{+}17]. DPD-based [GZM^{+}17]. DPG [FKDL17]. drag [BLL19, BLL20, Eva18, GPS17b, HM16b, ID17, SGC^{+}18a]. drift [DDH^{+}18, HK15a, LC17a, RRL19]. drift-diffusion [RRL19]. driven [AZ17, AZ19b, BFM19, BN19, BV18, BC16d, CPT16, CZ19a, CCZ15, CEL15, CV16b, DM17a, DS16, DVP^{+}16, EN17, GLZ19, KA18, LK17, LSP19, LZB^{+}17, LHMB18, LSD18, LAA16, MBA19, NS19b, PD16a, QWX19, SG16, Str17, TMES19, XWW^{+}16, YDCK16, Yeo19, YM19, Zoh17, dLGT^{+}17]. driving [BHZ16]. drop [BLJ17, JJS15]. droplet [BKG15, GLTB18, JRPPS18, LZ15b, LWC17, MOR18, OM19, WY19, WCWY19]. droplet-droplet [MOR18]. droplet-laden [BKG15]. droplets [Did17, Gan15, PKB15]. drops [Fed17, PST19, SRS19, ST18c, STV19]. DRP [Bra16c, CFSZ19]. Drucker [LEB^{+}17]. drum [Ant17]. dry [LAEK18, PP19, WWGW18]. drying [ABT16, FKY15]. DSA [OLD^{+}18]. DSA-lithography [OLD^{+}18]. DSMC [Mac16, GJ15, GRS15, JL18a, KJ17b, KJ18, MC16, RMC15, RSSSE18, WPB15]. DT [Nor15]. Dual [HB15b, MN18c, Stü17, WSN^{+}18, AAE17, CLP16b, DZ16, DZ18, Eng18, GCVCHH18, HHRA19, JS19, LWY18, NN19, NG17, NG18, OKWE17, Pan20, Par18a, PJB^{+}19, SPD19, SFDE15, TC15b, Yi18, ZD17]. Dual-consistency [Stü17]. dual-corrected [SFDE15]. dual-grid [PJB^{+}19]. dual-porosity [GCVCHH18]. Dual-scale [WSN^{+}18, JS19]. dualism [Luc15]. duality [SDW18]. duals [DPO16]. duct [BBKS16, TRLK18]. ducts [CV16a]. due [LM16, MCS16, SZY16]. DUGKS [ZWG17]. during [TYD16]. DVM [YSWW16]. Dynamic [DD16b, GSN17, LWY18, NLK^{+}15, VKE^{+}18, APP^{+}16, AF18, CL16, CvKH16, CFvKH18, EST17, FGL16, Gan15, GQC^{+}19, HCVH18, HKS^{+}16, IGQ15, KSVB18, LMC16, LGD17, MRA16, MG15b, MNG15b, MS15c, MHGL19, MDD^{+}19, MM17, NFG15, OCSC18, PD17, RPC^{+}18, TYD16, WY17, WS15b, YDLC19, YLH^{+}19, ZZ17b, ZXDL17, AM19]. dynamic-solver-consistent [WY17]. dynamical [Blo17, BW18a, BV18, CL18, CM19, EL17, GS15b, Lia16, NW17, NT19, NWFT19, OSP17, OB17, YM19]. Dynamically [ALKZ16, MN18c, CYS17, DBMB15, KG15, MD18, PS16]. dynamically-orthogonal [BCSK17]. Dynamics [BL18, WB16, ABG^{+}18b, ABG^{+}19b, AGBL15, AGKD19, AWS16, AF18, ATZ16, ABR16, BJTZ15, BKS18, BHdD18, BC18c, BBW16, BLS15, BLK19, BLJ17, BZ16b, BMR19, CJK^{+}19, CZBC^{+}18, CVM^{+}19, DMAM15, Dav10, Dav15, DPK17, DZ18, DG16c, DLR15, DFS16, DJL^{+}19, DPRZ17, EJZ17, FB17, FP18, GS15c, Gen15, GBM16, GK19b, HSLQ15, HSLQ16, HK15a, HM16a, HMBH15, HM17, iI15, iI17, JME18, JRPPS18, JB15, JLKF17, KM17, KBK15b, KP15b, Kor17, KS17, LSMS17, LFR17, LS15b, LS16a, LBTK18, LK16b, MLL19, MDH19, MNC19, MT18, MHL17, MD15, MHT^{+}19, MGB^{+}18, MMW15, MFG15, NPC15, NLL^{+}15, NLW^{+}16, PLL^{+}15a, Par18a, PQR17, PCBG18, RS17, RHS18, RKB19, SWC18, Say17a, Say17b, SVG18, SHKL16, ST15, SY18b, SSX16, SKW19, SMAG17, Sto17, SiI17, SAOW17, SZCL18, SZS15, TY17]. dynamics [TP16a, TAJ^{+}17, TPTT18, TR17, WE15, WWX19, WGJS19, WTS^{+}17, WH16b, WYA^{+}17b, XZZ15, XWW17, YZW^{+}18, Yeo19, ZL15a, ZLH^{+}17, ZD17, ZFL^{+}19, ZLC^{+}18, ZHWQ18, ZK18, HSB16, YG19, YDCK16].

Eady [YSC^{+}17]. EAM [YZW^{+}18]. EAM/FS [YZW^{+}18]. EAM/FS-type [YZW^{+}18]. Earth [DMM19]. earthquake [CCWY18, DD16b]. Eca [EH15, XS15]. ECGs [NCP^{+}17]. echo [BG19a]. Eddington [LMG19]. Eddy [FNP17, PD17, TABR17, BGV17, BR15a, BPM18, BJ16, CWS18, CLB^{+}16, CC16c, DLLV17, Fer17, FG17, KH15, LDB19, MD16, MMPS17, NYNYM15, PK17, RS16b, RWG18, RBGV15, SMD18a, VV16, CL16, CWS18, LLM17]. eddy-current [BGV17]. eddy-resolving [MMPS17]. eddy-viscosity [CWS18]. Edge [FCL19, FKK19, BGGM15, Dod17, DCD^{+}18, GDS^{+}16, GBC^{+}18, GYZ19, KHC^{+}16, MP15b, MP16, NL17, PF15, TBC^{+}16, WSH^{+}17]. edge-based [GZLH19, NL17]. edges [HK16b, Tsa16]. Editorial [Abg16, Abg19a, Ano18y, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18s, Ano18r, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z, Ano19-27]. effect [CFKK19, CM18a, EDW19, GR15, LYDB17, PQR17, SAH17, VALT16, WX17, XR17]. Effective [DGL^{+}15, GVTQ16, XLY15, BPS16, CPT16, CBC^{+}18, Cot16, HS17a, LK17, PVFN15, RLH19, VS17]. Effectivity [CGTH18]. Effects [NNW17, AAL15, GZM^{+}17, Gen11, Gho17, HCW15, HW15c, KD17a, KCS^{+}17, LW17b, LLW19, MAH16, MLB16, NWZ18, ST16, SPD^{+}17, SSL^{+}16a, SP16c, VCNOP18, WAF^{+}19, WTL17, YT17]. Efficiency [RW19, BHZ16, BT17b, CGTH18, Die15, HLL^{+}18, KK16, LWY18, TT19, WBC^{+}16]. Efficient [AG16, ALT17, BL18, BGV17, Cen19, CS16c, CLS^{+}18, CM18b, CY19a, CYS17, CLGA17, DNBH15, ESHA16, FNGV18, HE15, HHM17, Jer19, JYY18, KAR17, LZ16, Lia16, LB16, LLA19, LHA16b, MBSS15, MS16a, MPT16, MN16a, MMBP19, MFF^{+}19, MGCW18, MSA19, NMA15, NCP^{+}17, Nis19a, PLC18, SBT17, SMG19, SYY15, SDM^{+}17, SPRW15, TRM16, VSM17, WJD16, XL17a, YM17a, ZS15, ZS19b, bWAW15, ARG^{+}17, ADGN17, ALM^{+}17, APKP16, BGS16, BLM18, BCM15a, BAK19, BST15, BHFB19, CCdL15, CC17a, CE18, CCZC16, CGC17, CZJ17, CSC19, CM19, CNQ^{+}19, CPS17, DZR18, DY16, DLN15, DLNR18, DOO17, DB16b, EDC19, EMZ16, FHY^{+}19, FWK18, FBG15, FG19, FLW19, FYC^{+}18, GWB^{+}15, GHV19, GS15b, GLZ16, GP16a, GLTB18, GWC17, GX15, HD18, HTFL18, HHCG15, HMBH15, HF18, HWA15, HH19, HC17, IPSG15, JBM19, JCWX19]. efficient [KC17a, KH17, KHTZ19, LM15a, LKK17b, Ler16, LWY17, LHY17, LM19c, LPBR15, LPR19, LWC17, OSKN18, PXLL16, PL18, PES19, PSB^{+}18, PMS15, PKJ^{+}18, PSP16, RT16, SXBB15, SR19, SGMS16, SCL19, SO15, SSN15, SF16, SWLW19, Tav15, TRL15, TASA19, VBG^{+}17a, VD16, WLWW17, WSOW16, WAF^{+}19, WS15a, XX17, XWW17, ZZDB15, ZL15b, ZGD^{+}16, ZSM19, dlCGCA17]. efficiently [Cac15a, Cac15b, SLL19, ZWUR16]. eigenfields [HK16b]. eigenfrequency [ZC18]. Eigenmode [GFvR18]. eigenmodes [ABT17]. eigenpair [CG18b]. eigenpairs [VYP15]. eigenproblems [MBJ16, MBNJ16]. Eigensolution [MSP16, MDMS18]. eigensolver [AAB^{+}16, CHL^{+}19, RNO19, ZGD^{+}16]. eigenvalue [ABN15, Alm19, VMN^{+}18, BDKK17, CXX16, GFvR18, HLTC18, JPLL15, KL16, KFL17, LHS^{+}18, Loh17, PKA^{+}16, PGH15, XZ15, YM17c, ZCZ19]. eigenvalues [ABFR16, ABT17, HXB15, HSSZ16, Jac17b, XJG18]. eikonal [LP17b, NCP^{+}17, TH16, YS17, bWAW15]. Einstein [ALT17, Rua18, ZS19b]. elastic [AGKD19, AHHC18, ABT17, BHJ18, BXY17, Buk16, CHT17, CDL19, CHJT17, DL17, DHH^{+}18, DWW15, DPRZ16, DKK15, DD16b, GTL18, GFG^{+}15, GH17a, GK19a, GSL18, GFL17, GD19, GC17, GBS15, Heu17, Heu19, HJY19, KTK15, KDL15, KLRT15, KH18, LC15, LWZ16, LZL^{+}19, MKS18, PS15b, RM16, RRD16, SBHS19, SZW^{+}16, SCQP16, SiI17, SZF15, TDC^{+}19, VSDW18, VK15, WJD16, WTL17, XJG18, ZZZ17, ZZW^{+}16, ZBZ^{+}18, dTP16]. elastic-acoustic [RRD16]. elastic-electrostatic [DL17, DHH^{+}18]. elastic-plastic [CHJT17, GSL18, Heu17, Heu19, KTK15]. elastic-viscous-plastic [KDL15, WTL17]. elastic-wave [GH17a]. elasticity [FKDL17, RJ19b, SY18a, TD18, WXW15, ZZYC19]. elastodynamic [AB16b, CDC17, SGL18]. elastodynamics [CDL17, LM19c]. elastomers [SAH17]. elastoplastic [MN18b]. elastoplasticity [PBL^{+}19, RSB15]. elastostatics [GBD17]. electric [AAE17, AAE19, BGGM15, CCHL15, CJH^{+}19, DvB17, HK16b, KBR17, LDL^{+}16, LYDB17, NWZ18, ZRT18]. electrical [MS15a, VLP^{+}16, YG18]. electrically [Zoh17]. electrically-driven [Zoh17]. electro [DPRZ17, HGN17a, HGN17b]. electro-dynamics [DPRZ17]. electro-thermal [HGN17a, HGN17b]. electrocardiography [CGM15]. electrocardiology [PVFN15]. electroconvective [GN19b]. electrodes [MTD15]. electrodynamic [BAGK16, DPO16]. electrodynamics [BTGM17, BGTM18, PT17a]. electroencephalography [GPRA18, RMA17]. electrograms [NCP^{+}17]. electrohydrodynamic [HHL19, HLY15, HLSY16, JGS16, TND18]. electrohydrodynamics [STV19, Vee16]. electrokinetic [MXL16, PKP^{+}17]. electrolytes [GWB^{+}15]. Electromagnetic [HLCL19, AQ19, AJP15, ACC^{+}15, BAGK16, BGV17, CC16a, CC17c, CLFL17, DZR18, DC18a, DK18a, DK18b, DDV^{+}15, EGO19, FCL17, GHJ15, GKE15, HN18, Ism15, JG19, KS18a, KPJ18, LGO17, LO19, MHZ^{+}15, NOM^{+}17, PLL15b, ST16, SUR18, SF18a, SCS16, SLVE18, SCC19, SSL^{+}16b, Tao16, TSN16, TRL15, TBLM15, UWH17, VCNOP18, XB18, ZZ19, ZWUR16]. electromagnetics [AM17a, LH16, QWZ^{+}19a]. electromagnetism [BAGK16]. electromechanics [ANL^{+}16]. electron [ALM15, BTA17, CHE^{+}17, HZE19, HMRG16, Ido16, JL18a, KKS15, KKS16, KB18, LLVF^{+}15, LY15c, MP16, SCC19, VBG^{+}15, WSH19, YCBC15]. electron-electron [BTA17, HMRG16]. Electronic [CSCM16, LHS^{+}18, LC19, MRZG16, NOM^{+}17, PDdG^{+}17, PD16b, RO16]. electrons [CKK18b, KM16a]. electropermeabilization [GPG17, LPW15, MGPG19]. electrophysiology [CGG18]. electrostatic [AG18, BHP19, DL17, DHH^{+}18, HK15a, LLEK17, LYL19, LSP^{+}18, MSD^{+}17, PMF15, dCPDC^{+}17]. electrostatics [BCO^{+}15, DS16, XJ16, YX15]. Element [ADG19, CEH16, GFG^{+}15, GBS15, SCS16, SMAG17, TLB^{+}18, TBLJ15, AM17a, ABG^{+}15, AVT17, ADFG17, AB18, ASS13, ASS17, Alm19, AAD16, ADK^{+}17, ARTG^{+}19, BJRF18, BCD^{+}15, BCO^{+}15, BBKS16, BHL15, BXY17, BJWZ17, BGN15, BGN19, BBF^{+}17, BK17c, BC18b, BKO18, BCS19, BBB^{+}16, BS19b, BHP19, BSM16, BTWY15, BOD19, BKRB15, BFTVC18, CZW17, CCHL15, CWF16, CHT17, CDL17, CGL18, CL16, CJD^{+}17, CH17, CWW17, COV18, CHL^{+}19, CEL15, CELZ18, CEL18a, CPP19, CMH15, CLFL17, DSH^{+}16, DGMT17, Did17, DM19, EKEB16, FNNW19, FBM16, FZ19, FC19b, GM19, GFC18, GK19a, GG15, GBD17, GDA16, GY17, GSMR18, HR18a, HWH^{+}16, HHR^{+}19, HS17a, HTFL18, HdBH^{+}16, HLL^{+}16, HR17, HMFJ18, HHLY17, HXX18, HSF17, JTR16, JL15, JLLZ15, JTD16, Jou15, KC17a, KDF15, KVKS19, KE15, KG15, LP18]. element [LHO^{+}19, LTKA15, LMC16, LZ17a, LPR18, LGH^{+}18, LYZ18, LCLY19, LJ19, LYKW19, LTXB17, LTW18, LZS^{+}19, LYPP17, LMG19, LWC17, MML17, MIM^{+}19, MR17, Mel18, MDM^{+}15, MP16, MM16c, MF16a, MWYZ16, MN16c, MZ15, MMW15, NBT19, NNV19, NH17, NJHL18, NS19a, NS19b, ÖPHA15, PKF16, PG17, PCX17, PL16a, PHÖ^{+}16, PR17a, Rag15, RG15, RZ17, RMBN18, RAMB15, RRD16, RBGV15, RSD17, RBL16, SNSG16, SPX^{+}18, SDMS17, SC18a, SWZ15, SW16, SWPS17, SZW^{+}16, SGC18b, SY18a, SLVE18, SW18b, SA15, SFDE15, SSO^{+}15, SZ15b, SDW16, SS16c, Sov16, TCD17, TH18, TD18, TC15b, Tre16, URL16, URGT18, VSDW18, VKE^{+}18, WYZZ18, WZLS19, WXSJ19, WT19, WG15, WSF17, WHZ18, XWL^{+}16, XZ15, XJ16, XGZ19, YSC^{+}17, YYN^{+}17, YX15, ZS16, ZS15, ZL15a, ZGJ16, ZHLZ18, ZCY^{+}19, ZZYC19, ZBZT17, dCMR19, DJV^{+}18]. element-based [HMFJ18, JTD16, KG15]. element-integral [BKO18]. element-wise [MN16c]. Elementary [KD17b]. elements [Bou19, CV15, CHD^{+}18, Dod17, HR18b, JG15, LMH16, LPG18, LKSM17, MG15b, MSP19, MT17, MMW15, OKE17, Pas16, QN19, RGW16, RSB16, SWG^{+}17, SM16, SFP16, YP17, ZS16, ZILZ15, ZG19]. elevation [NMM18]. ELF [Chu17]. elimination [LSCC19]. ellipsoidal [SK19a]. ellipsoids [PGCG18]. elliptic [AGC19, AR16b, BFFB17, LL17, CWW17, CELZ18, CR18, CFF18, DMM19, EJMI18, FSWW17, FPDT17, FCL19, GH19, GLTG15, GY17, GY18, HL15b, HHLY17, HSF17, HSF19, KKL15, KCW17, LJ19, LMMS16, MWYZ16, NRS19, OKE17, PHHR17, SR16, SDW16, Vab15, VCNGP15, VMC^{+}19, WTGC16, WHE17, WRL19, ZILZ15, ZHW18, ZG19]. Embedded [CK16b, SMSR18, vLtTBI17, AMS17, BKO18, Cho15, DD16b, HCVH18, HDF18, KKJB16, KP15b, MS18a, MS18b, Mar19, MA17, NPC15, NRS19, PBKK17, PvL19, RS18, TAH16, WBM15a, BM15]. embedding [KYKS19]. emergent [BWR15]. emission [AP16]. emphasis [KS16a].

Empirical [dPSS16, ABdC^{+}18, NMA15, YZW^{+}18]. Emulation [LBTCG16, MRA16, XTS^{+}16]. emulator [ZKS^{+}15]. enabled [KMD^{+}18]. Enabling [YXD^{+}16]. Encapsulated [ÅN19]. encoder [ZZ18]. encoder-decoder [ZZ18]. endocytosis [LAA16]. energetic [Ama18, CSY15, DLWY19]. energetic-particle-magnetohydrodynamics [Ama18]. Energetically [MXL16]. energies [BJWZ17]. Energy [BC18b, BCJL17, CCBdL15, CGS18, CC19a, CCZ15, CLX19, CG18b, LZW19, LCF16, MRXI17, MMP18, NMM15, NYD19, OLDN17, OKWE17, RKH15, SL16c, WGJS19, AK17, AJW17, AMXJ19, BC18a, BGRC19, BKR19, Bra16a, BMR19, BMC^{+}18b, CBZ18, CCdL15, CCRdL17, CJYZ15, CS16c, CLS^{+}18, CLL17, CEL18a, CVG18, CL19b, Don15a, DS15c, FPASS16, FG17, FS19b, GK19a, GN19a, GZ18, GHL^{+}16, GGT18, GX15, HPV16, HJZC17, HLS15, HW15b, HJY19, JLKF17, KTK18, KTK19, LMH16, Lap17, LM18, LW15b, LYD19, LSS16, LLVF^{+}15, LW17e, MKS18, MHH19, MDMS18, MGCW18, MSF^{+}19, NMM16, NMM17, NN15a, PG17, PS14, PS15a, PME^{+}15, QWZ19b, QN19, RSB16, SYY15, SLN15, SD16, Sto17, Suz18, SN19, TC15a, TKC15, TCSM15, Tav15, Tav16, TT17a, TKP16, VW18, VCEK19, Vel19, VV16, WH15, WJD16, WW18, WCL15]. energy [Yan16b, YH17, YZW17, YLD19, YD19, YCS^{+}17, ZYSW16, ZW19, CT19, ZN16]. energy- [Suz18]. energy-balanced [PME^{+}15]. Energy-based [CGS18, MKS18, YCS^{+}17]. energy-conservation [CCRdL17]. Energy-conserving [CC19a, CCZ15, WGJS19, BMR19, FS19b, GK19a, HJZC17, CT19]. energy-preserving [CBZ18, CCdL15, LW15b, WW18]. Energy-stable [NYD19, Don15a, DS15c, LM18, MDMS18, SN19, YLD19, YD19]. energy-transport [HW15b]. Energy/dissipation [SL16c]. Energy/dissipation-preserving [SL16c]. enforced [WSY16]. enforcement [LHGF16, LHGF19]. Enforcing [GSK18, SHTY19, MN16c]. engulfment [TYD16]. enhanced [BHMS18, GNZ18, iI17, LXC19, MZAF17, MMB18, MH18a, MWZ19, PHD16, SW18b, XR17, XM18]. Enhancement [EST17, FL18, TT19, BTD16]. Enhancements [SP16b]. enhances [CSCM16]. Enhancing [CSN18, FBW16, JW15a, JES15, YLBL16]. EnKF [WHRL19]. ENO [Sid18, CH19, FHA16, FHA17b, FHA18, IDSG15, LJ16]. ENO/WENO [Sid18]. Enriched [LW18, VSDW18, LW17a, NRS19, SA15]. enrichment [KW16]. Ensemble [RMK15, BJO18, MWZ19, ZH15]. Ensemble-type [RMK15]. Enskog [SG19a, WZR15]. enstrophy [BC18b, PG17, SLN15, SD16]. enthalpy [HW15c, HW16c]. enthalpy-based [HW15c, HW16c]. entries [CC19b]. entropic [DCBK15]. Entropy [AKM^{+}19, CS17a, CHD^{+}18, DRM15, LSZ18, LI15, PCN15a, PCN15b, Ren19, ST18a, SGS^{+}19, UY19, YDLC19, Abg18a, AS15, Bra16b, Cha18, CW19, CJD^{+}17, CHS17, DRZ^{+}19, DCCH19, DWGW16, DWGW17, DWG^{+}18, DB18, GMLD18, GHH15, IC17, KTK18, KTK19, LW17a, LZW19, LCF16, LSI16, MLI17, MCL19, Opp17, SBT17, SW17a, SY18b, WWGK17, WWGW18, WG15, WG16b, WDGW17, YC17, BC16b]. Entropy-based [DRM15, SGS^{+}19, AS15, GHH15, SBT17]. Entropy-bounded [LI15]. entropy-residual [LSI16]. entropy-satisfying [CHS17]. Entropy-stable [CHD^{+}18, DCCH19, DWGW17, IC17, MLI17]. entropy-variables-based [GMLD18]. Environment [TCS^{+}16b]. environments [LCLY19, MWZ19]. EOS [FSB16, DLP19, RVK^{+}18]. EoS-independent [DLP19]. EPIRK [RT16]. epitaxial [YZW17]. epitaxy [Xia15]. epsilon [Lot18]. Equation [ACGR15, AMN18, AAE17, AAE19, Ali15, ASS13, ASS17, AMP16, ABFR16, An17, And16, ADK^{+}17, ABH18, AL19a, AHKT17, Ata15, ALTR17, ADOP18, BJRF18, BM15, BK17b, BJTZ15, BHL15, BZ19b, BLA^{+}15, BNM15, BBF^{+}17, BK19b, BKO18, BDK^{+}17, BIR18, BP18, BNS17, BH18, BWR15, BBKS18, BCM15a, BGGM15, Bou19, BR17, BTT18, CQQ16, CJWS19, CSD19, CP17, CNK19, CCZ18, Cha16, CCZC16, CLC16, CM18b, CSC19, CZ19b, CD17, CHCC18, Cif19, CMH15, CVG18, CV16b, CGRV17, CV18, CLMZ17, Cui15, CHLZ17, DD16a, DvB17, DLNR18, DS15d, DLL^{+}17, DJLQ18, DYL19, DLWY19, DBMB15, DKK15, EG17, EO15, EAAM15, EG16, EMSS17, FS16, FAY19, FQZNZ18, FGLB16, FLT17, FLW19, FYC^{+}18, FSM16, GR18, GMP16, GMP15, GMS16, GBD^{+}15, GN16, GD19, GWWC17, GHL15, GL17, HW15a, HZE19, HR18b, HB16]. equation [HW16a, HK19b, Her16, Hiv18, HLML17, HSC16, HHRA19, HF18, HW15b, HJ16, HXX18, HJS19, HXB15, HHY15, HWA19, HMRG16, IZ18, IKS19, JAH19, JW15c, JL16, JLLZ15, JCWX19, JJ17, KS16a, KMdB16, KTN15, KKS15, KKS16, KJYC17, KL17b, KDL15, KL15, KS18b, LFRH17, LSL15, LAL18, LTKA15, LJZ15, LWWY18, lLLNS16, LZ17b, LZ19, LSP^{+}18, LDWZ15, LY16b, LK16a, LTW18, LFAR19, LLH19, LYA16, LM15d, LQB16, LP17a, LP17b, MKYZ17, MK17, MS15c, MST15, MZTS16, MR17, MS17, MS18d, MLMM17, MFF^{+}19, MGCW18, NH17, NPRC15, OC18, OT15, OLV16, OWKE16, PKF16, PA19, Pan20, PKJ^{+}18, PL16b, PS14, PS15a, PLL15b, PDRB17, Pin15, Poë19, PS17, PCA19, PSV18, PC19, QWZ19b, RBI18, Rag15, RJ19a, RM16, ST18a, SYI^{+}19, SS17a, STEK17, STEK22, SM16, SLR^{+}16, SwS16, SL15, SK15a]. equation [SL19b, SLL16, SL16a, SWK18, Smi18, ST18c, SHW18, SPRW15, Sto16, SWLW19, SWZW19, SV17, SK15b, SLZ^{+}17, TCSM15, TK15a, Tav16, TSH17, Ter18, TGY18, TBO^{+}16, TCL15, Tou18, TH16, Tsa15, Tsa16, Vab15, Vab18, VSM17, Vee16, VMM19, Wac15, WSJY16, WY16, WH16a, WLW^{+}18, WL18, WKOE17, WBBC16, WSH19, WZRZ15, WZR15, WZL^{+}17, WA18, Wu19, WAZ19, XWB15, XJ16, XJLQ15, XQ17, XGZ19, XZJK19, YYL16, YZZ19, YS17, Yan17, YJB18, YLA15, YM15, YM17c, YC16, YL17, ZND16, ZLJ16, ZYW16, ZG18a, ZBZT17, ZSX17, ZHWQ18, ZYCK15, ZLL^{+}17b, ZV18, aKT16, bWAW15, dlHC16, MSG18a, MSG18b]. equation-based [OC18]. Equations [HO15, NMM18, AG16, AD15, AR16a, ABDN19, AD17, ALKZ16, AB19, AS15, AJW17, AMXJ19, ABH18, ABH^{+}19, ADH^{+}16, ATZ16, ABR16, ALL18, ABG18c, AEAM15, AB17, ABIR19, BJO18, BTD16, BK19a, BHJ18, BV15, BGN19, BK17c, BC18b, Beg15, BCB15, BP18, BDM17, BA15, BZ15, BDBEE15, BLMY17, BTVB15, BG19c, BAK19, BCJL17, BSWG15, BHGK18, BPD19, BHF15, BC16c, BTWY15, BT15, BTVC16, CZW17, CBZ18, Cap18, CGS18, CRW16, CG19, CW16, CAA18, CXH15, CCZC16, CTG16, CHZ16, CS16b, CHY16, CLZ18, CLL19, CYL^{+}16, CYYL18, CDN17, CCK^{+}18, CCKQ15, CVK16, CLX19, CDOY19, CFST16, CGJ19, CE17, Chu17, CRZ17, CCM17, CLP16b, CHD^{+}18, CCDL19, CEF15, CC19c, DRZ^{+}19, DK19, DA17, Del15, DWG^{+}18, DG16b, DP19, DWZ19, DM19, DKK^{+}18, DGL^{+}15, DMSC16, DMM19, Du18, DY19a]. equations [DMTB15, ETAG15, EDK19, EFHZ17, Fal16, FKF17, FWK17, FSWW17, FNNW19, FX18, FBW16, FP16, FLV18, FRRV16, GSN16, GJL19, GS15b, GSS15a, hGwSzS15, GS15a, GZY19, GCI19, GWK16, GP16a, GBM16, GCVMK15, GN19a, GO16, Gno17, GNK18b, GT18, GLK19, GHL^{+}16, GP16b, GPTK19, GTG15, GY15, GXX17, GLW18, HPY18, HE15, HSM19, HKLZ18, HdBH^{+}16, HCB19, HBR15, HHCG15, HLJ^{+}19, HJZC17, HTMP17, HY16, HY15, HZ15, HHY16, HS18b, HLCL19, HJW19, HSF17, Ism15, JLQX15, JG19, JPSX18, JW15b, JW16, JX15, JZ16, JX17, JL18c, JSB^{+}19, JXZ15, JL17c, JFS17, KCHW19, Kay15, KNS15, KA15, KÁGR18, KR17, KHTZ19, KL18a, LPWK15, LM18, LP18, Ler15, Ler16, LLS15, LZ15a, LXC^{+}15, LYC16, LX16, LDL^{+}16, LT17a, LZ17a, LGH^{+}18, LX18, LHLL19, LHS^{+}19, LL16c, lLNS17, LYD19, LR19b, LP16b, LXSC16, LW17d, LIW18, LCK19, LCCZ19]. equations [LY19, Lor19, LLLN18, LMG19, LHL15, LFT^{+}16, LHQ16, LI15, MMNI16, MD17, MD18, MM16b, MS18b, MLM18, MGT18, MS15b, MS16a, MG15b, MR16a, MA17, MKC17, MH18a, MPFL16, MDBCF17, MS18c, MBBKTH17, MDDM17, MHS16, Moh15, Mon19, MMP18, MFB18, MTK^{+}16, MDAB18, MBM^{+}15, MN16c, MSA19, MTBT18, MN18c, NMM15, NLFM16, NBH18, NN19, NW15, NN16, OS15a, Opp17, OvdHVH16, PG17, PXLL16, PCF15, PPCK17, PP19, PJC16, PCN15a, PCN15b, PS16, PTT18, PE16b, PND16, PE15, PDRB17, PBBK15, PMB18, Pop15, QHZ^{+}15, QWX19, QDH15, RMA17, RPK17a, RPK17b, RK18, RPK19, RMP18, RDQ19, RDM15, RJ19b, RBK19, SKP^{+}15, SP18, SS18a, SP15a, Sch16a, Sch16b, SWG^{+}17, SZ15a, SR16, SMS16, SF18b, SP19b, SY16, SG19b, SLB^{+}16, SGC18b, SLY16, SYM15, SYM17]. equations [SS18b, SO17, SPP16b, SKW19, SX15, SE16, SPZ18, SD16, Stü15, SL16b, SC18b, SL16c, Sub15, SJX15, SJXL15, SJH^{+}15, SJX17, SCS18, Sun19b, Svä15, TW17, TY17, TD16a, TD17, TDC^{+}19, TP17, TC15b, TXKvdV15, TXKvdV16, TMS^{+}19, TST17, TT16, TCL15, TC15c, TO15, TMH18, TYROG19, UL16, UY19, VST16, Vel19, VGZ18, Ver19, VS17, VCNGP15, VV19, WN18, WY17, WW15, WR15, WH15, WZ15, WXW15, WRL16b, WTGC16, WHY17, WHRL19, WT19, WCL15, WR16, WRL19, WWGK17, WWGW18, WG15, WG16b, WBM^{+}15b, WZ17, WZ18b, WX19, XDSX17, XY18, XHC15, XL16, YC17, YDLC19, YJ17, YHQ15, YYN^{+}17, YHKPF17, YLD19, YDN19, Yi18, YLLH19, YTW15, YWHP15, YT19, ZNS19, ZZK16, ZTBW19, ZA15a, ZSP15, Zha17c, ZHS18, ZLFW18, ZCL19, ZX19, ZCQ19, ZZL19, ZCQ20, ZED15, ZJL16, ZYD^{+}19a, ZQ16a]. equations [ZSQ17, ZYD19b, vOMB17, NMM17, PMF^{+}18]. equatorially [iI17]. equidistant [WWRS17]. equilibrated [GHP15]. equilibria [HR17]. equilibrium [AHHC18, BJ15, BWR15, CwYjS16, DRM15, FH17, GMLD18, HFM17, HKKP16, IKI15, MPP15, RG15, STR15, Sun19a, TCS17, TZ16, WG16a, WMS18, YHQ15]. equilibrium-preserving [TCS17]. equilibriums [NF17]. equipped [ZYK18]. Equivalence [ZN16, ACCCD^{+}17]. Equivalent [TUJ19, BM19b, KE15]. ERENA [MTK^{+}16]. Ericksen [NWZ18]. ERKN [MW16b]. erodible [LMKS15]. erosion [MS17, QM18]. erroneous [NN16]. Error [Kri17, LDB19, PDRB17, RS17, SM19b, AMK17, AGRB18, AM19, AR16b, BAD19, BH16b, CP17, CNOS15, CF19, DZC16, FC16, GWE^{+}15, HFND18, Hwa16, JW15a, KKJB16, LKN17, MM15, Nis19b, OKWE17, RL17, SZM19b, SD17, SW15, TS17, TR19, TYROG19, VLAB18, WK18, WA18, YY16, ZH15]. errors [AÁPB17, CSD19, Dav10, Dav15, HDA^{+}18, Iwa15, KL19, LM16, LZL^{+}17, RRMF^{+}19, SZY16, TPA19]. esophageal [KBG^{+}15, KGP^{+}17]. essentially [HWA15, LJ16, MSP19, MWB^{+}15a, ZPW18, ZQ17]. estimate [BAD19]. estimates [DZC16, IM15, JW15a, JES15, TYROG19, YY16]. Estimating [SZY16, TR17, WLK^{+}16, RRD19, STR15]. Estimation [CLZZ19, EDvW17, AMK17, AGRB18, ABT19, BLL19, BLL20, CSD19, Cha16, CM19, CN16, EH14, EH15, FAZ16, GP16b, ISP^{+}15, KM17, KRFV16, KMP^{+}19, LFAR19, PKW17, RZOZ19, RFGSV15, RBK19, RL17, SW15, SCE^{+}19, SM19b, TT17a, TS17, TR19, WN17, XS15, YJM19, ZH15, dFGS^{+}17]. estimator [Opp17, Swe18]. estimators [KMP^{+}19, LB17, OKWE17, QPK19]. Euler [BLL20, AG16, ABIR19, BLL19, Bal15, BLMY17, BQRX19, CG19, CBC^{+}18, CLL19, CCK^{+}18, CGS15, CHD^{+}18, DLMDV18, DKK^{+}18, GWK16, GLK19, GP16b, HY16, ID17, JPSX18, JL18c, Ler15, Ler16, LX16, LX18, LLB19, LI15, MS15b, MH18a, MMP18, MDAB18, MMPS17, PXLL16, PS16, PDRB17, PZF16, SP19b, SG19b, TCL15, VGZ18, WW15, WR15, WRL16b, WDGW17, XDSX17, ZLFW18, ZYD^{+}19a, vOMB17]. Eulerian [BB19, AB16a, ALA16, BMR^{+}16, BS15b, BLD15, BD17, CBB16, DL15, DB16a, FRW16, GSL18, ISST18, JKM19, LS16c, MC18, MWB^{+}15b, PR16a, PBL^{+}19, RW15a, RSB15, SDM^{+}17, WG19, ZWL^{+}19, ZYD^{+}19a]. eutectic [DMS17]. Evaluation [BAVC17, DB16a, SDJU15, SGC^{+}17, SGC^{+}18a, AW18, BGHK19, Bre18, CKK18a, CDM19, KKL15, MDT16, OCSC18, PAFT19, RCRF16, RSD17, RSSSE18, SY17, ST18b, ZSMP19, ZWG17, ZPE^{+}16]. evaporating [DM17a, Did17]. evaporation [IM17a, PS14, PS15a, SLC^{+}18, VALT16, WAF^{+}19]. even [WKOE17]. even-parity [WKOE17]. event [Hig17, KBK15a, SMG19, SGL17]. events [CL17, GH17b, MCS16, SPB16]. evolution [BGN19, BMRA^{+}15, GSN17, HJW19, LP16b, Lor19, OB17, Pis18, UG16, ZXL17]. Evolutionary [STHW17, MHJ15, SWS^{+}18, WS16]. evolving [CRW16, FBF15, MMNI16, MW17b, SS16a, SK19c]. Ewald [GKE15, NPP15, ST17, WB16]. Exact [SFP16, AB19, BJ16, Del15, HY16, KAR17, MDVM16, MCS^{+}19, MMW15, PBC^{+}17, PA15, SY17, SL16b, BPTA16]. Exactly [Lap17, Cac15a, Cac15b, NMM16]. Examples [LL17]. exceeding [BPF^{+}16]. Exchange [Han19, AKZ16, VW18]. exchange-correlation [AKZ16]. excitation [LYCC17, MML17, YCBC15]. excitation/deexcitation [YCBC15]. excitations [ZC19a]. excited [LY17, XZJK19]. exciton [CLZZ19]. Excluded [EG18a]. exclusion [DG16c, RMC15]. excursions [MD16]. exhibiting [MDP18]. Existence [PJC16, HBR15]. existing [vLtTBI17]. exit [LM15c]. expanding [And16]. expansion [AÁPB17, CSY15, GRMK15, GP16c, LQB16, PHD16, RMLvR18, ST18b, TMES19, YYL16, ZM16b, aKT16, RKO17a, WK18, WK19]. expansions [ARG^{+}17, ATM^{+}18, AZK16, AM18, BSN19, CN16, For16, GNZ18, GLZ19, HD15, JES15, KKL15, KSV^{+}15, KS16b, LLEK17, LMTC15, NS16, RKO17a, SS17b, THS^{+}19, VAD17, YLBL16]. expensive [PKW17]. experiment [BHNS19, MRP^{+}15]. Experimental [SHP^{+}16, CC16b, KL17a, NP16, NKN^{+}17, RL17, WLL16]. Experiments [FBW16]. Explicit [BBBG15, BJ16, CCRdL17, CJD^{+}17, SST^{+}15, Tao16, Tie18, ZRW19, BPL19, BCJ19, CB15, Che18, CKQT15, CGJ19, CDV17, DDM^{+}19, DLMDV18, EMSS17, GZY19, GPTK19, HLJ^{+}19, JMM19, KA15, KFWK17, LH17a, MVK16, MAM16, MPMB19, NYNYM15, NWKC16, O’S19, PGM17, RS15a, Ric15, SLH18, SC18a, SL19b, Tie16, Ver19, YSW15, dSPDH15]. explicitly [GSK18]. Exploiting [CELI15]. Exploring [BMT16]. explosive [AA19]. explosives [MN16b, RA17, RA19]. Exponential [GDA16, MHH19, PSG19, WBM15a, AMP16, CGJY19, Cui15, ETL17, GP16a, GRT18, GRT21, HS18b, Ike18, LLWJ18, LTR17, MW17a, SL19a, SWZ17, WJD16, WW19, ZSP15]. exponents [MSB^{+}16]. expression [AHHC18, LB17]. expressions [FS15]. Extended [BQCG17, RÖS17, Tso18, AMB17, BMC^{+}18b, CPT16, CS16a, Fan19, GG15, Guo15, HZ17, JH15, LYZ19, NBMB19, PXXZ15, PCX17, PRXC19, RWG18, SL16a, WSB19, PMF^{+}18]. Extending [GZY19, LYZ15, LY15b, MLMM17]. Extension [GHJ15, PxRS17, SGT17, WRL16b, WMS18, WYA^{+}17a, ZLL16a, ZLL17a, ATC17, ATC19, ABT16, FG16, FLT18, LDSM19, MN15, NMM18, RA19, SW16, SO16, SGT16, SiI16, SiI17, VNA15]. extensions [LYPP17]. extent [CK16a]. exterior [BXY17, CC17c, MHS16, NBT19, SP18, dCMR19]. external [ESGS17, LC17b, WH16b]. Extra [CSS15]. extracting [KKP15]. extraction [LLM17, ZJ18]. Extrapolated [MVK16, EMSS17, FBF15]. Extrapolation [LH17a, SLH18, ABFR16, HK18a, LWY18, PHHR17, SHTY19]. Extrapolation-based [LH17a, SLH18]. extrema [Kri17]. extreme [RKL18, VYP15].

F [HCVH18, SMT^{+}16, TAH16]. F-ADM [HCVH18]. F-AMS [TAH16]. F-MsRSB [SMT^{+}16]. Faber [ZV18]. face [uKHGK19, MMB18, MDD^{+}19, PF15, SP18]. face-centered [SP18]. face-offsetting [uKHGK19]. face-smoothed [MDD^{+}19]. face-vertex [MMB18]. faces [SPD19, WHY18]. faceted [SMAG17]. FaCSI [DFGQ16]. factor [JZ16, LCCZ19, TWN15, WCN15, LMG19]. factored [TH16]. factorization [HSF17, MM16a, VMK^{+}19, Yan19]. Factorizing [HSF17]. factors [BPF^{+}16]. failure [CL17, PKW17, RC18, WLL16]. family [FHA16, HDF18, HJW19, JZSX18, YC17, YLD19]. far [ZZ17a]. far-field [ZZ17a]. farfield [VAD17]. farm [GPAO^{+}18]. FAS [FGLW18]. Fast [BFI^{+}16, BDKK17, CDL17, CKQT15, CEL18a, DY17, EG18a, EFT15, FAY19, For16, GWC18, GP16b, HS17a, JW15b, LS19a, LZL^{+}19, NPP15, PLB18, RKO17a, RKN19, RO19, SPM16, VGF16, WZR15, WZ17, YGJ18, Yan17, YLLH19, ZZK16, ZSP15, AAB^{+}16, AAD16, AC17, BSK15, CHL^{+}19, Cif19, CV16b, CC19c, DG16b, DLR15, DWW15, DHC16, FHS17, GRT18, GRT21, Gno17, HM19c, HLL^{+}18, HJW19, Jac17a, JAH19, JW15c, JW16, JLC15, KNS15, KLWQ17, LKB15, LGH^{+}18, LCLY19, lLLNS16, LC16, LC17b, LQB16, MCIGO19, MST15, Moo17, MTK^{+}16, NRZS17, NLL^{+}15, PA19, PWC18b, PR17b, RDQ19, RZZ19, RSBS19, RLP16, RL17, SZM19a, SO17, SLdTV18, Sti16, SL16b, SWZW19, SLZ^{+}17, TCD17, TH16, WK18, WK19, WQZ15, WW17, WZRZ15, WZL^{+}17, Wu19, XB18, YS17, ZG18b, ZVO15, ZZ19, ZGD^{+}16, aKT16, dCMR19, GKE15, GN19b, Str18]. fast-marching [NLL^{+}15]. faster [Mac16]. fatigue [CF15]. fault [FFJT16, LKK17a]. fault-resilient [LKK17a]. faults [DD16b]. FC [AB16b]. FC-based [AB16b]. FCIF [MHL17]. FCT [BK17c]. FD [MF17, NWFT19, PZNG15, Sha17b, SF18b, AW18, BFFB17, BFF19, DA17, FFBB16, FL16, KÁGR18, PGCG18, PLR18, PLPR19, Sha17b, TLLF15, ZNS19]. FD-NILSS [NWFT19]. FDTD [DPO16, FYZ^{+}15, Fuj19, DV17, MBM^{+}15, Ram18, SP18, SL15, SSM^{+}17, Tie16, Tie18, WSOW16, ZFZL15]. FE [LXC19, SSC^{+}16, YLLH19, ZFL^{+}19, ZBZ^{+}18]. feature [Gno17, KLA17]. feedback [KL16]. Fekete [PR17a]. FEM [HLL^{+}16, BG19b, BLJ17, Dod17, FH17, GS15b, HW19, HL15b, LHW^{+}17, SL17, vOMB17]. fermion [CEHM19]. fermions [SD15]. ferrofluid [GP16c]. FETI [KCW17]. FETI-DP [KCW17]. FFT [BJTZ15, BC18c, BM19b, GO15, LO19, MVZ16]. FFT-accelerated [LO19]. FFT-based [BC18c, GO15, MVZ16]. fiber [NRZS17, WAF^{+}19]. fibers [CPSF17, DCP15]. Fictitious [IML15, ZLY15, HXLL15, HGW18, LHY^{+}19, PR16a]. fidelity [AAI16, DIX^{+}18, HFND18, LKK17b, LSWF16, MS16b, MS15c, MW16a, NS19a, PPCK17, PVPK17, PKW17, RPK17a, RKN19, RS17, UG16, VBF15, YZ19, ZYK18, ZLX17]. Field [SRS19, ARF18, ATM^{+}18, ADFG17, AAE17, AAE19, BPL19, BJK17, BGJ^{+}15, BDPM18, BG16b, CLW18, CWF16, CC15, Cen19, CS16a, CJYZ15, CS16c, CWB^{+}19, CY19a, CKQT15, CYS17, Chi19, CLMZ17, DWG^{+}18, DMM19, DW19, ESGS17, Fed17, FCL17, GHL^{+}16, GGT18, Guo15, GFW16, Hag25, HLCL19, JTR16, JCNK19, JJ18a, JJ18b, KPJ18, LSL15, LW15a, LBZ16, LWZ16, LDL^{+}16, LWY17, LYL19, LY16c, Liu19a, LBB^{+}17, MAA18, NS19a, NWB19, OTS17, OLD^{+}16, OLB^{+}17, PPCK17, PMS15, PD16a, PKB15, RTO15, SYY15, SLL16, Shi19, SAOW17, TW17, TK15a, TS19, TSST16, TBB^{+}19, VSM17, VS17, WJD16, WDT^{+}19, WXSJ19, Wic16, WCCB16, WHZ18, XCX17, XL16, XZT18, Yan16b, YH17, YY17, Zau16, ZW16, ZZ17a, ZHLZ18, ZYSW16, ZYCK15]. fields [AQ19, BAGK16, BMC^{+}18b, JG19, KBR17, LE16, LX16, MLMM17, PVPK17, RRM^{+}16, RSD17, SAF^{+}19, Tao16, TG17, TBB^{+}19, XTS^{+}16, XY17, YG18, ZFPB16]. fifth [CTG16, DSX19, SG19b, WLGD18, ZQ16b]. fifth-order [CTG16, DSX19]. Filament [JSS15]. filamentation [VK18]. filaments [DCP15, Mar19]. filled [DSH^{+}16, SAH17]. film [JS16, KHP17, Pes15, Xia15]. films [AASRT17, HJY19, JTR16, LVB^{+}15]. filter [ABT19, BJO18, BJ16, KC17a, MRP^{+}15, NYNYM15, PE16a, PKA^{+}16, PvL19]. Filtered [OS15a, LMH16, MM16b, SA19, ZN16]. Filtering [GO15, ALMJ15, BW18b, BJ16, EMM^{+}18, Fal15, FBM16, GZY19, MG17, MP17, NYNYM15, SD18]. filters [Fal15, Fal17, HHR15, LAK^{+}16, MRY19, SKO17, VLTPS16, ZH15]. filtration [VK18]. finding [BSWG15, SPM^{+}15]. fine [CGS15, KGT15, NLFM16, NS19a]. fine-grid [KGT15]. fine-scale [NS19a]. fingering [BST^{+}18]. Finite [ADG19, AGBL15, Alm19, AHZ19, AMM^{+}15, BGN19, BS19b, BTWY15, CLC16, CEH16, DG18, DSH^{+}16, DJV^{+}18, EN18, FPT17, GFG^{+}15, GBS15, Gri19, GSS15b, GSMR18, HWK19, IGQ15, Kay15, KS15a, LTKA15, LYC16, LYPP17, MML17, MDL16, Mas18, MHZ^{+}15, NBT19, RBGV15, SKO17, SYV17, SP16c, TV19, TVB^{+}16, TRLK18, WT19, YYN^{+}17, ZSQ17, vEKdB16, AM17a, ABG^{+}15, AVT17, ADFG17, AD17, ASS13, ASS17, AAD16, ABFR16, AM19, ADK^{+}17, AA19, ABT16, ARTG^{+}19, AM17b, BJRF18, BCD^{+}15, BD15a, BBKS16, BHL15, BJWZ17, BGN15, BK17c, Bat17, BC18b, BKO18, BGV17, BS19a, BCS19, BLVC17, BH18, BLMY17, BLD15, BDZ15, BD17, BDLM18, BHTT17, Bou19, Bra16c, Bre17, BOD19, BKRB15, BFTVC18, CCHL15, CBS18, CCS18, CWF16, CHT17, CCZ18, CTG16, Che18, COV18, Cho15]. finite [CDX18b, CHL^{+}19, CEL15, CELZ18, CEL18a, CPP19, CMH15, CGP16, CR18, CPS17, CHS17, CCM17, CYWL17, DK19, DGMT17, Did17, DWZ19, DM19, DLK17, DMS17, DDH^{+}18, DVP^{+}16, DMM19, DL16, DvWZ18, FAZ16, FS18, FGLW18, FNNW19, FBW16, FZ19, FLW19, FC19b, GM19, GSS15a, GH17a, GFC18, GK19a, GOR17, GDS^{+}16, GS16, GG15, GBD17, GHL15, GLK19, GQC^{+}19, GHL^{+}16, GDA16, GY17, GSL^{+}19, GL17, HR18a, HWH^{+}16, HR18b, HHR^{+}19, HS17a, HdBH^{+}16, HZL^{+}15, HLL^{+}16, HR17, Heu17, HMFJ18, HHLY17, HY16, Hu17, HXX18, HAH16, Ism15, IDSG15, JTR16, JMM19, JW15b, JW15c, JW16, JSB^{+}19, JLLZ15, JG15, Jou15, KKH18, KDF15, KW15a, KVKS19, KW15b, KE15, KJYC17, KL17b, KPJ18, Kla15, KS17, KBD19, LH16, LMH16, LLD^{+}16, LY15a, LN17, LAL18, LHO^{+}19, LX16, LL16b, LHMB16, LMC16, LZ17a, LGH^{+}18, LYZ18, LLSJ19, LWWY18]. finite [LYZ15, LY15b, LZ17b, LYKW19, LDT19, LMMS16, LTB16b, LJ16, LTW18, LZS^{+}19, LKSM17, LMG19, LWC17, MIM^{+}19, MF17, MN04, MN17, MDHC15, MR17, MT17, Mel18, MH18a, MMvR18, MRK15, MH18b, MSS16, MM16c, MF16a, MLB18, MWYZ16, MN16c, MMW15, NNV19, NH17, NJHL18, NS19a, NS19b, NN15a, Nis15, Nis19a, NF17, Nor15, OLDN17, OV17, PxRS17, PCX17, PL16a, PL18, PHÖ^{+}16, PS16, PR19, Pei16, PS15b, PWP15, PS17, QWZ19b, RBI18, Rag15, RG15, RGW16, RLGT19, RMBN18, RAMB15, RRD16, RSD17, RWN18, RBL16, SNSG16, SPX^{+}18, SDMS17, SGL18, SAEF17, SWG^{+}17, SM16, Sha17b, SF18b, SY16, SG19b, SGC18b, SY18a, SLY16, SYM17, SW18b, SYV14, SDH^{+}16, SKG17, SA15, SBH19, SFDE15, SSO^{+}15, SZ15b, SDW16, SZ17, SS16c, SLA^{+}19, SP15b, SDW18, SK18, Sub18]. finite [SN19, TLH15, TD18, TMT17, TC15b, TBO^{+}16, TKP16, Tso18, TR19, URL16, VSDW18, VKE^{+}18, Vil19, VSC18, WR15, WDS15, WRL16a, WRL16b, WRPL17, WYZZ18, WLGD18, WLW^{+}18, WZLS19, WXSJ19, WT15, WSF17, WA18, WHZ18, XWL^{+}16, XDvW17, XZ15, XX16, XJ16, XDSX17, XDLX19, XGZ19, XM18, YG18, YSC^{+}17, YYL16, YHQ15, YP17, YX15, YM15, ZCHS15, ZOG19, ZS16, ZS15, ZILZ15, ZGJ16, ZZZ17, ZG18a, ZHLZ18, ZSL^{+}19, ZRW19, ZCY^{+}19, ZBZT17, ZQ16b, ZQ17, ZG19, ZXDL17, dCMR19, dFJN16, ADN19, CJ17, NWFT19]. Finite-Difference [SYV17, Bra16c, Bre17, CR18, CYWL17, GH17a, GS16, KPJ18, LH16, LHMB16, LWWY18, NF17, PR19, RBI18, SYV14, SN19, TLH15, WA18, YYL16]. Finite-Element [GFG^{+}15, AVT17, AAD16, BOD19, CHT17, DM19, GFC18, JTR16, LMG19, SFDE15]. finite-element-based [CMH15]. finite-elements [SM16]. Finite-Volume [DG18, IGQ15, TVB^{+}16, vEKdB16, CCS18, CDX18b, DK19, DDH^{+}18, GOR17, GDS^{+}16, IDSG15, KS17, LLD^{+}16, LN17, LL16b, LZ17b, MDHC15, Nis15, Nis19a, Nor15, PS16, SDH^{+}16, SKG17, TMT17, Tso18, TR19, VSC18, XDvW17, ZSL^{+}19, CJ17]. finite-volume-based [SBH19]. Finite-volume-concept-based [SKO17]. finite-volume/Monte [GDS^{+}16]. fins [LGZ^{+}19]. First [CC15, CNK19, LSL15, SLL16, ALKZ16, AZK16, ARTG^{+}19, Cac15a, Cac15b, DBZ17, DPRZ16, DPRZ17, DKK15, Hiv18, LWWY18, LM15c, LLLN18, MA17, MN16a, MRN16, OWKE16, PTMF18, Roy15, SM16, VSM16a, VSM16b, VLN^{+}18, WTX17, Yan16b, YH17]. first- [Cac15a, Cac15b, YH17]. first-order [Hiv18, LWWY18, LLLN18, MN16a, MRN16, OWKE16, SM16, VSM16a, VSM16b, VLN^{+}18, WTX17]. first-principles [AZK16]. fish [LGZ^{+}19]. fitted [BOA17, CZBC^{+}18, CWW17, DSH^{+}16, RA17, RA19, WW18, ZJ18, ZSX17]. fitting [LT17b, XDLX19, ZXDL17]. FitzHugh [LZT^{+}15]. five [RJ19b, TGY18]. five-equation [TGY18]. five-equations [RJ19b]. FIVER [MZAF17]. fixed [DGW18, IKI15, RZ15, SY17]. flame [KP15b, LZL^{+}17]. flames [SWS^{+}18]. flash [LZSG19, WKSS15]. flat [KJYC17, KMGR16, WF17]. flexibility [iI15]. flexible [BSK15, BPGS16, DG18, DCP15, FKR16, GLS15, JSP16, Moo17, NRZS17, SWG^{+}17]. flexible-wing [Moo17]. Flexibly [YS18a]. flexural [MDW18]. flexural-gravity [MDW18]. floating [CGSS18, LC17a]. flocking [ZK18]. flooding [DD17a]. Flow [BPS17, KLA17, YDCK16, ABI17, AASRT17, ABG^{+}15, APR^{+}15, APP^{+}16, AAG16, APV^{+}18, AS17, AMS17, AZ19a, ASWvD19, BCSK17, BZ19a, BCST17, BB17, BBKS16, BHST17a, BHST17b, BHST18, BGN15, BAD19, BLVC16, BLVC17, BS19b, BPS16, BNGI19, BLK15, BAVC17, BLG^{+}16, BLJ17, Bon17, BCB17, BHMS18, BC16d, BB15, BKG15, BKRB15, BKKRB16, CB18a, CCRdL17, CC15, CX15, CKT17, CV16a, CS17b, CGRV17, CV18, CM18d, CLNH15, CvKH16, CFvKH18, DM17a, DGW18, DM16, DWR18, DDV18, DGMT17, DB16a, DL18c, ESGS17, EST17, EN17, Fal17, Fan16, FMRZ17, FST15, FW17, FITcD19, FSB16, FS19a, GZM^{+}17, GPAO^{+}18, GPS17b, GSL18, GDFL17, GFL17, GAIN19, GGL^{+}17, GO16, GCVCHH18, GN19b, GMT19, HXLL15, HTFL18, HSK^{+}15, HG17, HGW18, HKH^{+}16, HW16b, HDF18, HY17, IPSG15, ION^{+}17, JSP16, JSVD17, JL16]. flow [JT18, JD19, KD17a, KPKGH19, KCSW19, KHP17, KA18, KJ17a, KEJ18, KH17, Kla15, KF17, KLC19, KS16c, KS18b, KW16, KFWK17, KRK^{+}18, KJ17b, KL19, KS16d, Kwo19, LVB^{+}15, LE16, LRA17, LW18, LH15, LPGT16, LHB^{+}16, LZB^{+}17, LLFX18, LLY18, LZT17, LKN17, LGZ^{+}19, LNM15, LAA16, LRGO18, LSCC19, MZAF17, MNG15a, MCN18, MTZ16, MT18, MT19, MH19, MP19, MHL17, MN18b, MS17, MDP^{+}15, MTJ17, MF16a, MB15, MLB16, MM16d, NL15, Noe15, NSK^{+}16, NSL16, OT15, OSP17, PST19, PZNG15, PGCG18, PM19, PHÖ^{+}16, PT18, PLWJ16, PGGW18, PTT19, PZF16, PSN^{+}19, PME^{+}15, QYF15, QLF16, RVZB15, RWG18, RW15a, RXSG15, RdM19, RJLW19, Rua18, RPC^{+}18, SK19a, SPX^{+}18, SV19, SYI^{+}19, Say17a, Say17b, SBHS19, Sha17a, SRBB18, SLL17, SCJ^{+}18, STG17, SPW18, SHW17, SPN^{+}19, Str17, SK18]. flow [SWLW19, SKC17, SCS18, SZCL18, TH18, TE19, TP16a, TLH15, TWH15, TAH16, TV19, TLLF15, TB19, TT16, TD16b, TSST16, VCEK19, Vel19, VWV17, Vos17, Vre17, Vre21, WWR16, WYLX17, WSN^{+}18, WDT^{+}19, WPB15, WC18, WNDB19, WCVF16, WKSS15, XCX17, XXR18, XDLX19, YYY^{+}16, YSY17, YNW17, YFC19, YR15, YM17b, YTW15, YZZ15, Zad11, ZP16, ZLY15, ZW15, ZV16, ZW16, ZZS^{+}17, ZZ17b, ZOY^{+}19, ZYD^{+}19a, ZZX16, ZZX19, ZRE16, aKT16, dFJN16, dMRHJ17, dPSS16, tEDKT17]. flow-field [TSST16]. flow-induced [KPKGH19]. flow-structure [LGZ^{+}19]. flow-transport [BKRB15]. flowing [ZZDB15]. flows [ACGR15, AMB17, ALO18, AB18, ACS16, ATC19, AB15, AEvW19, Bal15, BMR^{+}16, BJ15, BT19, BFI^{+}18, BC16a, BM19a, BS15a, BVM17b, BMT18, BDPM18, BFTVC18, Cai16, CV17, CBS18, CGS18, CFSN18, CGK17, CL16, CJD^{+}17, CZL^{+}15, CX16, CHJT17, CZL18, CL19a, CS18b, CLG^{+}19, CD17, Chi19, CGS15, CEL^{+}18b, Cif19, CC16c, CLGA17, CG18b, CPS17, CLV19, CCPdL19, CG16, CM16b, DG18, DLM18, DIX^{+}18, DXvW18, DvW19, DY16, DAO17, Don15a, DS15c, Don17, Don18, DVP^{+}16, ESHA16, EJMI18, EFO19, EFO20, FGL16, FBL17, FWK18, FBJS19, FNGDMNR18, FNP17, FMPT18, Fid17, FBM16, FZ19, FPDT17, FC19a, FG19, FKK19, GMLD18, GOR17, GHR17, GIF18, Ger17, GHF19, GWC17, GG15, GBCF15, GBCF16, GZLH19, GJ15, GRS15, GEZK16, GA18, GWYS18, GAS^{+}18, GSL^{+}19, GSS15b, HHA15, HFM17]. flows [HL15a, HEPG15, HZL^{+}15, HTZG17, HP17, HSB16, HM19b, HM16b, HTMP17, HAX19, HCLT19, HW18, HWA19, HLA19, HTBG15, IGQ15, JL19, JS17, JLC18, JGS16, JJ18b, KKH18, KCHW19, KYUO15, KTN15, KP15a, KLNH17, KCS^{+}17, KF17, KP15c, KYW^{+}16, KYW^{+}18, KL18b, KV16, KS15b, KTK18, KTK19, LMPS15, LS15a, LVTR15, LSP19, LPG18, LHO^{+}19, LPB17, LFDP16, LL16b, LW17b, LPR18, LLWJ18, LLSJ19, LSD^{+}17, LSR16, LC16, LC17b, LD15, LMKS15, LWB^{+}16, LXSC16, LH17b, LH18, LMZ19, LWX19, LDHJ15, LZW^{+}17, LHA15b, LHA16b, LWTF19, LEB^{+}17, MLI17, MM16a, MNG15a, MDL16, MOAA15, MTZ16, MA19, MC15, MP17, MRK15, MRXI17, MCGS16, MDP18, MF16b, MLL18, MA16, MSB^{+}16, MR16b, MM18, MPMB19, NdlLPCC19, NPB19, NGPB19, NMJFM19, NYD19, NDCB17, OB19, OVP15, OSKN18, ÖPHA15, OD17, PD19, PKP^{+}17, PNZ18, PHHA18, PL16a]. flows [PL18, PSS17, PSB^{+}18, PM16, PPLC16, PWC18a, PN17, PN18, PGM17, PLWJ16, PF16, PEVG18, PBCR19, DM18, PCBG18, PWP15, QSB18, QSBY19, RZOZ19, RS16b, RDG17, Ren19, RV16, RMF^{+}18, Ric15, RZ15, SXBB15, SWS17, SPD^{+}17, SPD19, SP15a, SGMS16, SHA16, SL17, SKF15, SWML17, SVG18, SAK18, SWMD17a, SWMD17b, SXY18, STW16, SDM^{+}17, SA19, SY18b, SDH^{+}16, SKG17, SSA17, SMSR18, SLA^{+}19, SGT17, SWL19, SST^{+}15, SGP17b, Suz18, SSZC19, TZGW18, TUJ19, TK12, TK15b, TGS19, TND18, TASA19, TBO^{+}16, Tou18, TMH18, TKP16, TABR17, TO19, UG16, VPM15, VSM16a, VSM16b, VALT16, WDG^{+}17, WSY15, WSS^{+}15, WSHT15, WSY16, WSP17, WCH^{+}17, WZ18a, WHR19, WTL19, WKPS18, WSN^{+}15, WMS18, WG19, WT19, WGME17, WL17, XWL^{+}16, XDvW17, XX16, XML17, XSL18, XWZ^{+}18, XS19, YXW19]. flows [YS18b, YLZ^{+}19, YSW15, YSWS16, YXF^{+}16, YSWW16, YGEM17, YD18, YSYW19, YG19, YD19, YL16, YYJ^{+}19, YCS^{+}17, ZBH^{+}18, ZMF15, ZMCC18, ZKG19, ZCSZ19, ZLC^{+}18, ZLGS18, ZWG17, ZSM19, ZS19b, dFVJ15, dLGT^{+}17, OY19]. fluctuating [BRK^{+}18, DSH^{+}16, HM17, LLA19, SC18a, dSPDH15]. Fluctuation [BNGI19]. fluence [Swe18]. Fluid [AAL15, CGSS18, FB15, HWK19, HM17, JBLO15, LVTR15, LGD17, MHGL19, RW15b, SAK18, SDM^{+}17, VALT16, ABI17, ABG^{+}15, AAI16, Ama15, Ama18, AB15, BAGK16, BHKS16, Bar19, BQCG17, BZ16b, BCM15b, Buk16, CSW^{+}19, CGS18, CZBC^{+}18, CM16a, CDM18, CH17, CV16a, CLGA17, CVM^{+}19, CSH15, CLM15, CM16b, CYWL17, DG16a, DG18, DFGQ16, Dom18, EST17, EKSS15, ED16, ELH^{+}16, FW18, FLV15, FG19, FRW16, FHA16, GLTB18, GSL18, GLS15, GCVCHH18, HXLL15, Har18, HSK^{+}15, HM16a, HW19, HJS19, HDF18, HTvdH^{+}19, HF19, ISST18, IM17b, JSVD17, JBM19, JH15, JS17, KB18, KLC18, KF17, KC17c, KRK^{+}18, KM15, Kwo19, LTB16a, LLD^{+}16, LC15, LFR17, LGB17, LFDP16, LW17b, LLJJ18, LSD^{+}17, LHY^{+}19, LY16c, LDGH16, MLI17, MAK15, Mar19, MNC19, MOAA15, MAM16, MPR^{+}18, MC15, MT18, MCS^{+}19, MMMS15]. fluid [MTK17, MRXI17, MAA18, Mon19, MM17, MKV^{+}17, NFG15, NBZ^{+}19, NPC15, NBMB19, Niu16, NF17, NSK^{+}16, NLW^{+}16, PD19, PSS17, PHÖ^{+}16, PN18, PR16b, PAL^{+}16, PP17, PQR17, PCBG18, PME^{+}15, QLS^{+}19, RG15, RKRGW17, RV16, RTG15, RVK^{+}18, RPC^{+}18, SSL17, SWC18, Say17a, Say17b, STKL19, SRBB18, SMP16, STW16, SLB^{+}19, SMA^{+}16, SS16c, SJH^{+}15, TFGK18, TBC^{+}16, TCA16, TZGW18, TND18, WSP17, WCH^{+}17, WB17, Wic16, WS15a, WS15b, XDvW17, XYF^{+}17, XTYL18, YYY^{+}16, YK15, YS15, YXF^{+}16, ZAK15, Zad11, ZZPH18b, ZZPH18a, ZKG19, ZW19, ZBZ^{+}18, ZPE^{+}16, dJRP^{+}15, dTP16, BAVC17, JSS15, SGD18]. fluid-capsule [ISST18]. fluid-composite [BCM15b]. fluid-dynamic [EST17, MM17, WS15b]. fluid-dynamics [PQR17]. fluid-electron [SCC19]. fluid-fluid [LSD^{+}17]. Fluid-particle [HM17, Ama18, Har18, HW19, KF17]. fluid-porous [NSK^{+}16]. fluid-saturated [SSL17]. fluid-solid [MPR^{+}18, YK15, YS15]. Fluid-structure [CGSS18, LGD17, BHKS16, BQCG17, CZBC^{+}18, CM16a, CM16b, DG16a, DG18, DFGQ16, EKSS15, ED16, FW18, FRW16, HDF18, KLC18, KC17c, LC15, LLJJ18, LHY^{+}19, MMMS15, MAA18, MKV^{+}17, NBZ^{+}19, PHÖ^{+}16, Say17a, Say17b, SMP16, SLB^{+}19, SMA^{+}16, WCH^{+}17, Wic16, XTYL18, YXF^{+}16, dTP16]. fluid-structure-interaction [LTB16a]. fluid/fluid [MAK15]. Fluid/Level [VALT16]. fluid/thin [FLV15]. fluid/thin-walled [FLV15]. fluidic [MKV^{+}17]. fluids [ARF18, AJVH17, BHKS16, CCBF19, CFPB17, CSN17, DSH^{+}16, Don15b, Don17, Don18, DPRZ16, ES18, Hag25, JN19, JKM19, KKH18, KKS15, KKS16, KBK15b, LYM19, Liu16, LLA19, MS18d, PR16a, PS14, PS15a, RJ19b, SG19a, SK15b, TS19, TGY18, TOR^{+}15, TL17, WE15, YD18, ZDGW16, ZYSW16]. Flux [ALMJ15, AWJ17, HR18b, LKSM17, Loh17, NMM15, VK18, ZN16, AHNF15, AMH^{+}18, APV^{+}18, AEL^{+}15a, AEL^{+}15b, AEL^{+}17, BMT16, BNK18, BND16, CSY19, CJL16, CLP16b, FKF17, FNNW19, FS15, GHL15, GSL^{+}19, HK19a, HWH^{+}16, HHR^{+}19, HZL^{+}15, KW15a, KKS16, KFL17, Kri17, LBZ16, LYL19, LK16b, MS15b, MMB18, MDMS18, NMM16, Nis15, PPM^{+}19, QN19, RdM19, SYOS19, SYOS21, STW16, Stü15, SST^{+}15, TNB21, TT16, TCL15, VST16, VV16, VDPP15, Vil19, WSY15, WSHT15, WSY16, WG16b, YSW15, ZJLC15, ZCSZ19, ZXL17, BK17a, LWJV19]. flux-ADER [NMM16]. Flux-balance [VK18]. flux-based [MMB18]. Flux-corrected [HR18b, LKSM17, Loh17, FNNW19]. flux-dependent [KFL17]. flux-jump [LYL19]. flux-limiter [ZJLC15]. flux-reconstruction [AHNF15, AMH^{+}18, HK19a]. flux-split [HZL^{+}15]. flux-splitting [KKS16]. fluxes [CCK^{+}18, DH18a, QN19]. fly [EZG16, Mas18]. FMM [CHCC18, YS18a]. foams [SS16a]. focused [TSN16]. focusing [KLWQ17]. Fokas [CFF18]. Fokker [FLT17, JG19, TC15a, TKC15, TCSM15, TCS16a, TCS17, CM18b, CCL16, GJ15, GAJ15, HYK^{+}16, KJ17b, KJ18, MS18d, QWZ19b, SV17, SK15b]. folded [CLR15]. Force [HLU15, TP16a, ZLH^{+}17, AAL15, BDG^{+}17, CFO18, DKPC15, DKC15, KK16, LYKW19, LBB^{+}17, MBA19, SD16, VSM17, WG16a, YZT^{+}18, YCS^{+}17, Zau16]. force-coupling [DKPC15]. force-field [LBB^{+}17]. forced [ABG^{+}18b, ABG^{+}19b, CM18a, GTL18]. forces [CG16, GLTB18, GLMC16, HWK19, LT15, LM16, LCK19, NJPB17, YDCK16]. forcing [BCRS19, CK16a, Hig15, JL19, KLSF15, LC15, LWTF19, PPLC16, PBCR19, PG18, PVB17, YS15]. form [ABH18, Del15, DS15c, DKK15, FBY19, GWK16, GZ19, JFS17, KML18, NYD19, OWKE16, RÖS17, Ren19, RSSSE18, RWN18, RN18b, SPP16b, WKOE17, WMM^{+}18, XWW^{+}16, ZZH16]. formalism [PD17, SD15, SSL^{+}16b, TZSS17]. formalisms [OMLdL16]. format [GKMS17, GJ18, LY15c]. formation [AZ17, GP17, MDH19, SG18, SPM16]. forming [CLFL17, PR16b]. forms [AMH^{+}18, BOD19, KTK18, KTK19, PF15]. formula [DF16, LDOK17, PBKK17, RPL^{+}18]. formulas [DC18b, Loz17]. Formulation [Kor17, KSVB18, Teu16, TSB^{+}18, BVG^{+}16, BHST17b, BDAA^{+}18, BBF^{+}17, BC18c, BS15a, CZBC^{+}18, CCK^{+}18, CMH15, CGRV17, DG16a, DSH^{+}16, DCP15, Don15b, Don18, DPRZ16, DB16b, DPRZ17, EFO19, EFO20, FRL15, FNNW19, GPRA18, GMLD18, GHS19, GS16, GC17, GMT19, HTFL18, HL16b, JKM19, Jou15, Kim15, Lap17, LL19a, Ler15, LYKW19, LHA16b, MRRRF18, MT19, MN16b, MTJ18, MTD15, MR16b, NN17, NN19, NF17, NRS19, PBP18, PND16, QSB18, RG15, RWN18, SDMS17, Sel15, SM16, SL16a, TUJ19, Vil19, WZ18a, YLH^{+}19, YTW15, YT19, ZHA17b, ZCY^{+}19, dCMR19]. formulations [AG16, CCPdL19, FAY19, FKDL17, JHPAT17, LGO17, PBL^{+}19, RB15, SST^{+}15, Suz18, VS17, WRL16a]. Fortran [GBR15]. forward [CG19, FYC^{+}18, GPRA18, GHV19, RMA17, RPK19, SH19, SL19c, SSZC19, ZLGK19]. forward-in-time [SSZC19]. forward-peaked [FYC^{+}18]. four [Ein19, JCNK19, RS16a, SD17, SSN15]. four-dimensional [RS16a, SD17]. four-field [JCNK19]. Fourier [GKE15, GN19b, Str18, AW18, ALMJ15, DY17, Fer17, GSN16, GWWC17, HB15a, KFL17, MDVM16, MP16, MMBP19, MH17, ST15, SGT16]. Fourier-spectral [ALMJ15, MP16]. fourth [BGN19, CG16, DL17, DLL^{+}17, DL18b, DL18c, FS18, GH17a, GPS17a, GPS17b, GLW18, pHzSrC15, JPSX18, LHMB16, PXLL16, VSC18, YC17]. fourth-order [CG16, DLL^{+}17, FS18, GH17a, GLW18, pHzSrC15, JPSX18, LHMB16, PXLL16, VSC18, YC17]. FPDEs [SZM19a, SZM19b, ZK15]. FPGA [LWL18]. FR [HCB19]. FR/DG [HCB19]. FraC [FNNB19]. fraction [DB16a, SR19]. Fractional [ECC18, KHP15, KADE15, KADE17, MK17, SK19b, YPK16, ZK15, ZM16a, ABDN19, ASB^{+}15, Ali15, ADH^{+}16, ATZ16, AL19b, AEAM15, AHKT17, Ata15, BJO18, Beg15, BA15, BZ15, BDBEE15, BK18, BSWG15, BTWY15, CF15, CC15, CG19, CXH15, CNOS15, CLC16, CP16, CWL^{+}16, CLZ18, CV16a, Cui15, CGG18, Die15, DMSC16, DLL^{+}17, DYL19, DZC16, DvWZ18, DW19, EAAM15, EE16, FLW19, GSS15a, GS15a, GMP15, GLW18, HPY18, pHzSrC15, HO15, HB16, HSC16, HHRA19, HZ15, JW15b, JW15c, JW16, JX15, JX17, JWH16, JLLZ15, Kat16, KNS15, KSM19, LZ16, LYC16, LW17c, LWY17, LGH^{+}18, lLLNS16, lLNS17, LZT^{+}15, LLH19, Lot18, Luc15, MBSS15, Mac15, MD17, MD18, MS16a, MR16a, MM15, MP15a, MDDM17, OM15, PCF15, PPCK17, Pan20, QDH15, RRD19, RZ18, SMC15, SYM15, SYM17, SX15, SPRW15]. fractional [SLZ^{+}17, TY17, TSH17, TMdO19, Vab15, WH15, WZ15, WY16, WH16a, WWZ19, Wu16, WZ17, WZ18b, XY18, XHC15, YZZ19, YYN^{+}17, YJB18, YLA15, YLLH19, ZZK16, ZSP15, ZC15, ZG18a, ZzSK15, ZJL16, ZBZT17, ZHWQ18, ZLL^{+}17b]. fractional-order [ZC15]. fractional-step [HPY18, WWZ19]. fractions [EDvW17, KB19]. fracture [AEL^{+}15a, AEL^{+}15b, AEL^{+}17, BBB^{+}16, BPS16, BHTT17, BHMS18, FNNB19, FFJT16, FKS19, HCVH18, NFG15, NS19a, NWB19, Noe15, Wic16, ZHLZ18]. fracture-matrix [AEL^{+}17]. fractured [ABG^{+}15, AEVW18, BHMS18, CEL^{+}18b, Noe15, SMT^{+}16, VSDW18, XYF^{+}17, XML17, ZOY^{+}19]. fractures [BPS17, CB19, TAH16]. fracturing [RSBS19]. fragmentation [LPWK15, NFG15, WYA^{+}17a]. frame [ALA16, DCA^{+}16, DvB17, NMM18, SL16a, YZT^{+}18, YXD^{+}16, ZLGS18]. frame-based [ZLGS18]. frames [CE17, Chu17]. framework [Abg18a, BKS18, BGV17, BT15, BKL17, CLY^{+}15, Chi19, CLX15, CG16, CJ17, DEZ16, ES18, FOF15, FFJ^{+}19, GM16, GK19c, GGW17, HAPK15, HHA16, HL15b, ISP^{+}15, JTR16, JSVD17, JC17, KKP15, Lap16, LSP19, LKK17a, LKK17b, LLM17, LS16b, LMN18, MLM18, MR17, MW15, MZ15, OD17, PD19, PWC18b, PC19, QM18, RPK19, RW15a, RN18a, SLC^{+}18, SK19c, TKB^{+}15, TMH16, TABR17, TR19, WAF^{+}19, XDvW17, Yan19, ZLH^{+}17, ZWL^{+}19, ZSMP19]. frameworks [AAL15]. Fredholm [XZ15]. Free [MK15, ZBZ^{+}18, ZH19, AAE19, APV^{+}18, AS17, Ama15, ALM15, BD15a, BK17b, BDG^{+}17, BMMP19, BFI^{+}18, BMT18, BAR15, CDL19, CDDL19, CMH15, DDM^{+}19, DDV18, FH17, FPDT17, FKY15, FPV18, FC16, GP17, GG15, HHA15, HR17, KLSF15, KO17, LXC^{+}15, LLW19, LS16c, LW17e, LTWZ18, LEB^{+}17, MNG15a, MG15b, MNG15b, MDL16, MTZ16, MTK^{+}16, NWKC16, PSS17, Pes15, PBBK15, DM18, QWZ19b, RRM^{+}16, RDG17, RZ15, RRD19, Say17a, Say17b, SW18a, SP19b, SLB^{+}16, SLB^{+}19, Sla16, TBO^{+}16, VGF16, XLY15, XYPT16, XL16, XS19, YSWW16, YFJ17, YKMM19, YCS^{+}17, ZFZL15, ZD15a, ZZS^{+}17, GS16]. free-boundary [FH17, HR17]. free-flow [MTZ16]. free-slip [KLSF15]. free-space [VGF16]. Free-stream [ZH19]. free-surface [FPV18, HHA15, LS16c, MNG15a, MDL16, NWKC16, RZ15, XS19]. Freeman [HB15b]. freestream [AHNF15]. frequencies [ALM^{+}17, LQB16]. Frequency [KÁGR18, LS15a, WT16, BZ16a, BNM15, CGTH18, CDL17, CC17c, CLQ17, FQZNZ18, GM19, HPC19, HBR15, LHMB16, NKN^{+}17, NNW17, Par15, Par17, RW19, SZW^{+}16, SJXL15, Tre16, ZZ17a, ZF18, dCMR19]. frequency-dependent [SJXL15]. Frequency-domain [LS15a, LHMB16]. Frequency-independent [WT16]. Fresnel [DKTH15]. fretting [CLB^{+}16]. friction [DLM18, FNGDMNR18, MDBCF17, MHGL19]. frictional [LL18, ZVO15]. Friedrichs [GSK18]. friendly [YZW^{+}18]. frond [ASB^{+}15]. Front [MP15a, CCM15, Fan19, FL18, GSN17, Gro18, Hiv18, IM17a, Kim15, LTWZ18, dJRP^{+}15]. front-tracking [FL18, dJRP^{+}15]. fronts [Kim15, MW17b, SP15b, WWGW18]. Froude [CDV17]. Frozen [LYA16, LLY15]. FS-type [YZW^{+}18]. FSHL [WHL17]. FSHL-based [WHL17]. FSI [BHST17a, BHST17b, BHST18, LHB^{+}16, Liu16, LHW^{+}17]. FT [GR19]. FTLE [NJ15]. fuel [CLB^{+}16, MTL^{+}17, PBA^{+}15]. Full [LXSC16, ST16, ZKS^{+}15, AEL^{+}17, BFP18, CXX16, DBD^{+}17, HdBH^{+}16, Hig17, HLTC18, Ido16, JAH19, KYPK15, LC19, LZHM19, LS19b, MKYZ17, MAM16, MDP^{+}15, NS19a, PKN17, Par15, YLH^{+}19]. full- [Ido16, KYPK15, Par15]. full-angle [Hig17]. full-aperture [LS19b]. full-potential [LC19]. Full-wave [ST16]. full-waveform [BFP18, MKYZ17, PKN17]. Fully [AVT17, FLV15, KSI17, LSMS17, NN15a, NLW^{+}16, PKLC17, WSP17, Wil19, XDvW17, BA15, CZBC^{+}18, CC16a, CS16a, CCGH17, CLNH15, CvKH16, Del15, EKEB16, FRW16, GS15b, HYK^{+}16, HW19, JL19, JHT^{+}18, KLC19, KBG^{+}15, KL18b, LLD^{+}16, LM15a, LMKS15, MJ16, MNO^{+}17, MTJ17, MTJ18, NN17, OvdHVH16, PR16a, PP17, DM18, PBC^{+}17, QWXZ17, RSBS19, SMOM^{+}17, SCC19, SK19c, TCSM15, TH18, WMY18, Xia15, YSLY19, ZLY15, ZW19, MHL17]. Fully-conservative [Wil19]. Fully-coupled [XDvW17, TH18, MHL17]. fully-discrete [ZW19]. Fully-implicit [NLW^{+}16, CZBC^{+}18, Del15, HW19, LLD^{+}16, MNO^{+}17]. Fully-resolved [WSP17]. Function [BL18, MNR19, AMN18, BR17, Cha16, CS18a, CNQ^{+}19, CVK16, GBvZB16, GKE15, HXB15, HLTC18, Ike18, KMGR16, KW16, LB15, LC16, MG15a, MF17, MJ17, OD15, PYK19, PD15, RdM19, Sha17b, SF18b, SWX18, SP15b, SWLW19, TZSS17, WQZ15, WX18, XYPT16, YSW15, YYL18, YLH^{+}19, YC16, ZXL17, Mue18]. function-based [CNQ^{+}19, YSW15]. function-generated [MF17]. Functional [GS16, OLB^{+}17, AJP15, AKZ16, AAB^{+}16, BHZ16, BEJ15, GJ18, GZ18, NP16, NGY^{+}17, RS17, Tav15, TVB^{+}16, ZLH^{+}17]. Functionally [WW18]. Functionally-fitted [WW18]. Functions [SNK18, ABDN19, Alm19, BVM^{+}17a, Bar18, BC16b, Bre18, CDDL19, CLP16a, FHY^{+}19, FBW16, FLT18, FC16, GT19, Gri15, Gri19, HBR15, KDF15, KMGR16, LC18, LC17b, LHY17, LVL18, MVKD15, MR16a, MDT16, MFB18, PLHA18, RRD19, STR15, SKS17, SW18a, VGF16, WG16b, WF17, XL17a, XM18, Yam19, ZKS^{+}15]. Further [LPR19]. fusion [FBC^{+}16, GDS^{+}16, HYK^{+}16, HLS19, JH15, KMP^{+}19, LKK17a, LKK17b, RKH15, Ram17]. future [MSV^{+}16]. fuzzy [ASB^{+}15]. FV [CMDL18]. FVM [HW19, LH18]. FVTD [BTGM17, BGTM18].

G [MBM^{+}15]. G-FDTD [MBM^{+}15]. Gabor [DvB17]. Galerkin [HGN17a, RHS18, TRL15, ZCQ20, ZN16, AG16, AM17a, AGC19, AS15, ADK^{+}17, BFI^{+}16, BH16a, BHJ18, BK19b, BMMP19, BST^{+}18, BFI^{+}18, BDM17, BGD19, BCJL17, BHGK18, BFT17, BD17, BCB17, BD18, BT15, CGQ18, CBA17, CPdS19, CG19, CGMH18, CWM^{+}16, Cha18, CW19, CHOR17, CHY16, CS17a, CYL^{+}16, CYYL18, CZL18, CL19a, CCKQ15, CLG^{+}19, CDM19, CLX19, CSN18, CXY19, CK16a, CK16b, CCGH19, DGMT17, DM17b, DKK^{+}18, DLL^{+}17, DY19a, DL16, Ein19, EL18, EL19, EHXM15, EG18b, EDC19, FWK17, FWK18, FNP17, Fer17, FX18, FGLB16, FBM16, FSB16, FS17b, FS19b, GR18, GS15b, GWK16, GCVMK15, GHS19, GBC^{+}18, GSN17, GX15, GY15, HKA19, HL16a, HGR16, Hig15, HGN17b, HJ16, HJS19, HB15b, HS18b, Jac17b, JAH19, JSB^{+}19, JH17, JLLZ15, JL17c, JJ19, JHT^{+}18, KCHW19, KCSW19, KFF^{+}17]. Galerkin [KRFV16, KG15, KWHB19, KFWK17, LMH16, LLP^{+}16, LW17a, LW18, LPR18, LX18, LYZ18, LC19, LHLL19, LM19a, LM19b, LTB16b, LP16b, LY16b, LW17e, LSZ18, LMB18, LTW18, LMB19, LFAR19, LCCZ19, LKSM17, LLLN18, LMG19, LHL15, LHQ19, LI15, LSI16, MSK18, MLM18, MRRRF18, MS16a, MK17, MNG15b, MN16a, MSP19, MKC17, MF16a, MLB18, MSP15, MSP16, MSB^{+}16, MMPS17, MWYZ16, MH17, NdlLPCC19, NJ15, NPC15, NPRC15, NDCB17, NLW^{+}16, OLHD17, OKE17, PL16a, PE16a, PSB^{+}18, PP17, PP18b, PND16, PMB18, QWZ^{+}19a, QSY16, QDH15, RXSG15, RDM15, RRMF^{+}19, Say17a, Say17b, SS18a, Sch16b, SMP16, SL19b, SLB^{+}16, SLB^{+}19, SZ15b, SDW16, SE16, SPZ18, Sti16, SWZW19, SCS18, TH18, TSC17, TD16a, TD17, TD18, Teu16, TXKvdV15, TXKvdV16, TLB^{+}18, UL16, URL16, VSDW18, VCNOP18, Vil19, VMC^{+}19, WW15, WZ15, WTGC16]. Galerkin [WYZZ18, WSN^{+}18, WZLS19, WLE17, WGJS19, WWGK17, WWGW18, WG15, WMM^{+}18, WBM^{+}15b, WH16b, WTX17, Xia15, XOX19, XJLQ15, XL16, XYG19, YY16, YK19, Zha16, ZLH^{+}17, Zha17c, ZCL19, ZX19, ZY19, ZCQ19, ZBZT17, ZT17, ZK18, ZYD19b, dFVJ15, vOMB17]. Galerkin-Fourier [Fer17]. Galerkin-free [BFI^{+}18, SLB^{+}16, SLB^{+}19]. Galerkin-mixed [GS15b]. Galilei [GBU15]. Gamblets [OZ17]. game [LL18]. gap [KLC19, MHJ15]. gaps [QYF15]. gas [AAI16, AEVW18, ATC19, BLVC17, BS19b, BMT18, BTA17, CBS18, CX15, CXL16, CCL16, CLM15, DMAM15, DY16, DLR15, FSB16, GBM16, GK19b, GJ15, JPSX18, JZSX18, KJ17b, LS15a, LVB^{+}15, LLFX18, LLY18, LXSC16, LYZ19, LWX19, LK16b, MTZ16, OCSC18, PD19, PX15, PXLL16, PX16, Par18a, PLWJ16, PCBG18, RXSG15, RXS16, RL18, SYI^{+}19, SVG18, STKH15, SWJG19, SY18b, SWZW19, SJX15, SJXL15, SST^{+}15, SJX17, TZGW18, TK12, TK15b, TO19, WYLX17, WT19, WZRZ15, Wu19, XCX17, YSW15, YSWS16, YG19, Yan17, YZZ15, ZCHS15, ZLY15, ZLFW18, ZWL^{+}19, ZQCT15, ZXL17, ZLGS18, ZZX16, ZZX19]. gas-kinetic [CX15, JPSX18, JZSX18, LLFX18, LXSC16, LYZ19, LWX19, PX15, PXLL16, PX16, PLWJ16, RXS16, SST^{+}15, SJX17, WYLX17, XCX17, YSW15, ZLFW18, ZXL17, ZLGS18, ZZX16, ZZX19]. gas-liquid [OCSC18, TO19]. gas-liquid-like [TK12, TK15b]. gas-particle [LWX19, ZWL^{+}19]. gas-solid [ZLY15, ZQCT15]. gas-solids [AAI16]. gas-surface [LLY18]. gases [AIP17, ATC17, BLVC16, HM19c, Lap16, PL16b, SBT17, WZR15]. Gauge [WSF17, GHL15, Say17a, Say17b]. gauss [GS15c, AP16, AS16, CDLR19, CT19, Kas15, MRRRF18, Nis18a, PKN17, PR17a, RW19, Spe15, Stü17, Teu15, Tsa16]. Gauss-type [Spe15]. gauss-verifiability [GS15c]. Gaussian [AZ19b, BKP16, BVS18, CZB15, EMZ16, FITcD19, GLG^{+}19, LLY15, LYA16, MKC17, NDH19, NP16, PYK19, PVPK17, PSMPG17, RZOZ19, RPK17b, SL16a, TBG16, WLL16, XX17, YBSTT19, ZFPB16, ZKS^{+}15]. Gaussian-like [BKP16]. Gaussian-process-regression-based [YBSTT19]. Gaussian-sum [EMZ16]. Gaussian-windowed [SL16a]. GBS [HRJ^{+}16]. GCL [SYV17, GBM16, SYV14]. Gegenbauer [O’S19, SP16a]. Gegenbauer-based [SP16a]. gels [RBJS15]. gene [LB17]. General [BHST17b, LW15b, Abg18a, AVT17, AB16b, BPL19, BDV17, CSW^{+}19, CZ19b, CV15, CHD^{+}18, DCP15, Don15b, DB16b, FFJ^{+}19, Her16, HJW19, JAH19, KYPK15, LMG15, LKK17a, LHGF16, LMN18, PA15, RBL16, SR19, SAEF17, SMS16, SGD18, SDW18, Tao16, TCL15, Tou18, Vel19, WHY18, WXSJ19, YFJ18, ZLL16b, ZZL19, ZSMP19, ZWG17]. Generalised [Eng18, Ran18, CC16b]. generalization [Sha17b]. Generalized [BPGS16, BLA^{+}15, CSS15, DKTH15, HL15a, KH18, PX15, VGZ18, Ama18, ABH^{+}19, ABdC^{+}18, BVM^{+}17a, BLM18, BLS15, CEL15, CELZ18, CEL18a, CPP19, DS15c, DWW15, EARA15, Fal17, GFvR18, GQC^{+}19, IG15, KKJB16, LY16b, MS17, MSA19, RSSSE18, SNSG16, TM15b, WZR15, Yam19, ZHA17b, ADG19, ALA16, CEH16, CTM^{+}16, GFG^{+}15, MSG18a, MSG18b]. generate [WG16a]. generated [BRW15, iI17, MF17, dlAC17]. generating [TMH16]. generation [CJD^{+}17, FP16, GPAO^{+}18, GG15, HS18a, HK15b, iI15, MKSP19, PVFN15, SAF^{+}19, Spe15, TST^{+}15, WHE17, WHEK18, YGJ18, ZJ18]. generative [BGG16, SHTY19]. generator [CWW17, DW19]. generators [Cot16, LWL18]. generic [CSD19]. Gennes [BHZ16]. Gentile [Gho17]. genuine [CKT17]. genuinely [FKR16, MS15b, QSBY19]. Geodesic [AZ17, Pei16]. geological [VD16]. geomechanics [CHT17]. Geometric [BGL^{+}17, EG17, SL19a, BGN19, BMRA^{+}15, CQL^{+}17, DK19, EJMI18, HB16, HCW15, HK15b, KML18, Lan19, LYZ15, LY15b, MM16c, NMM16]. Geometrical [JG15, GDFL17, LHGF19, NLL^{+}15, SCE^{+}19, TLR16]. geometrically [DD16b, MMW15]. geometries [AC17, BOA17, CXL16, CRZ17, CGG18, GWC18, GEZK16, GTG15, GSS15b, HS18a, HR17, KWHB19, LBTK18, Mar19, MR16b, NSB15, OC18, OVP15, ÖPHA15, RMBN18, TBC^{+}16, TK15a, XDvW17, ZG18b]. geometry [AMS17, BB17, BBKS18, BCM19, BC16d, BRW15, Cai16, DCD^{+}18, GBC^{+}18, GFvR18, KB19, KYPK15, MBS19, PWC18b, RS16b, SYOS19, SYOS21, Sch16a, Sch16b, TCS17, TNB21, TSST16, WKOE17, WHEK18, XTYL18, YDCK16, YXD^{+}16]. geometry-adaptive [XTYL18]. geometry-specific [YDCK16]. geophysical [MRP^{+}15, ST16]. George [ABT16]. geostatistics [Shi19]. geothermal [WHT18]. GFM [LH17b]. GFRD [BL18]. Ghost [LVTR15, CR18, GLTB18, KTK15, LHW^{+}17, PG18, HWK19, VALT16]. ghost-point [CR18]. Gibbs [LBZA16, LAK^{+}16]. Gibou [YM15]. Gilbert [EMSS17]. Ginzburg [GS15b, LZ15a, SKP^{+}15, SSL^{+}16b, Tav15, WH16a, ZYW16]. Girsanov [KM17]. given [CCHL15]. Glacier [AS17]. glaciology [ALKZ16]. GLM [DWG^{+}18]. global [BCRS19, CBZ18, CCK^{+}18, DC18b, FC19b, GMS19, GBC^{+}18, Kor17, SZY16, STEK17, STEK22, SDH^{+}16, SKG17, SW17b, Swe18, TCA16, TMWF18, TL15, TO15, YP17]. globalization [BDB^{+}17]. Globally [HCB19, RKO17a, BK17b, BK19a, NKN^{+}17, XL16]. GMDH [SW17b]. GMDH-NN [SW17b]. GMRES [SL16b]. GMsFE [JL17b]. GMsFEM [GFG^{+}15]. GN [ZED15]. GNAT [JD19]. Goal [GBD^{+}15, JKE^{+}17, VLAB18, AMK17, BAD19, CPP19, TVB^{+}16]. Goal-based [GBD^{+}15, JKE^{+}17]. Goal-oriented [VLAB18, AMK17, BAD19, CPP19, TVB^{+}16]. Godunov [AAG16, ADOP18, BV15, Bar19, BT15, sCYxL^{+}18, Jac17a, MDP^{+}15, MWB^{+}15a, Rod17, SiI16, SiI17, SIX16, XLL^{+}17]. Godunov-like [MWB^{+}15a]. Godunov-type [AAG16, sCYxL^{+}18, MDP^{+}15, Rod17, XLL^{+}17]. golf [CWJ18]. Good [BRK17]. Gordon [BZ16a, AMP16, BZ19b, CJWS19, LW17d, LIW18]. governed [MP15a]. governing [Beg15, QWX19, RBK19, WX19, XZJK19]. GP [RLGT19]. gPC [KKL15, Poë19]. gPC-intrusive [Poë19]. GPR [JN19]. GPU [BGHK19, BBBG15, CWM^{+}16, HPY18, KCSW19, LXL17, MN18a, RGW16, SB18, SMAG17, SSL^{+}16b, TRM16, VWV17, XZZ15, YZW^{+}18, ZBH^{+}18, ZMCC18, dlAC17]. GPU-accelerated [BGHK19, CWM^{+}16, HPY18, YZW^{+}18, ZMCC18]. GPU-advanced [SSL^{+}16b]. GPU-based [LXL17, MN18a]. GPUs [RTG15, Swe18, WWGW18]. Grad [DCBK15, PKF16, RCRF16]. graded [Bat17, KVKS19, SLA^{+}19]. Gradient [AT18, GJ18, GY17, GY18, DDM18, ALT17, BMCK15, CPT16, DH18a, ES18, GO16, GGT15, GMA18, GNZ18, GYZ19, HAX19, JJ19, LR19a, LYL19, Nis18a, Nis19a, PHD16, RYZ18, Rua18, SXY18, SLL17, SZ15b, Stü17, SCS18, TAR17, VYP15, ZVO15, ZS19b, dMRHJ17]. Gradient-based [GJ18, ES18]. gradient-direction [GGT15]. gradient-driven [CPT16]. gradient-enhanced [PHD16]. Gradients [WN17, ABG18c, Bat17, BHdD18, Loz17, MIM^{+}19, Nis18a, NRS19]. grafts [BFI^{+}16]. grain [BVM17b, JTR16]. grain-resolving [BVM17b]. grained [FOF15, HKKP16, KKP15, TWN19]. graining [CSCM16, GK19c, MLL19, MVKD15, SZK17, dlCGCA17]. Granular [RLH19, AAG16, BVM17b, FNGDMNR18, FC19a, HM19c, IML15, LEB^{+}17, MDP18]. graph [WQZ15, ZSM19]. graph-partitioned [ZSM19]. graphene [BTA17, GYZ19, KM16a, LYDB17, Ram18, RMC15]. graphene-reinforced [LYDB17]. graphical [LZ18]. graphics [AAB^{+}16, GK19b, GP18]. Grassmann [GS18]. grating [FM15, ZS15]. gratings [HN17a, HN18]. gravitation [BLMY17, CCK^{+}18, GLK19, LX18]. gravitational [LX16, XCX17]. gravity [DMM19, GR19, MDW18, Sun19a, vOMB17]. gray [CG15, SJX15]. greedy [BSN19, FHY^{+}19, SKS17]. Green [BR15b, BR16, Pop15, BL18, Cha16, GT19, GKE15, GPTK19, HLTC18, LM15a, LXC19, LC16, LC17b, MDT16, Nis18a, PD15, Stü17, TZSS17, VGF16]. Green-Naghdi [LXC19]. Grey [DRM15, LMG19, MRM16, TWM18]. Grid [RO16, AZ16, ACJ17, Ani16, BGG16, CCFC19, CQ15, CXL16, CS16b, CWB^{+}19, CLB^{+}16, DCCH19, EH14, EH15, FGL16, FAZ16, FPDT17, GH17a, GCVMK15, Gro18, HK16a, HCLT19, HHL19, HF19, iI15, iI17, JW15a, KLA17, KPKG15, KS15a, KGT15, KLNH17, Kor17, KS16d, LML^{+}16, LLM17, LHMB16, LWWY18, LKN17, MNG15a, MNC19, MPFL16, MP19, MHZ^{+}15, MAH16, PMGW16, PLWJ16, PR16c, PJB^{+}19, RRM^{+}16, SPB18, SFT16, SP16c, Sub18, SZF15, Teu15, TBB^{+}19, VPM15, Vre17, Vre21, WDG^{+}17, WTGC16, WG19, WHE17, WHEK18, Wil18, XL17a, XS15, XTYL18, YYL16, YFC19, ZZKF15, dLDG^{+}18]. Grid-based [RO16, RRM^{+}16]. grid-independent [WDG^{+}17]. grid-refinement [KS16d]. grid-to-rod [CLB^{+}16]. gridding [PLB18]. gridfree [CB18b]. gridless [DTA^{+}15]. Grids [SYV17, ÅN19, APLK19, ABH18, AB17, AG18, BNK18, BST15, BHTT17, CSW^{+}19, CBC^{+}18, CTG16, CYL^{+}16, CZL18, CL19a, CLP16b, DDJ18, DWR18, DPO16, DL15, DBMB15, Eng18, EH18, FGLW18, GZLH19, HR18a, HL16a, Hu17, IGQ15, IDSG15, IM17b, JL17a, KF15, KW15a, KG15, KD17b, KL18a, KSI17, LAL18, LGB17, LB15, LPR18, LYZ15, LY15b, LZT17, LAEK18, LCCZ19, LHGF16, MM16b, MN15, MDHC15, MDM^{+}15, MGBG16, MGB^{+}18, MHGM^{+}15, MF16a, ML16, NOM^{+}17, NYNYM15, Nis15, Nis18a, Nis19a, OLDN17, PxRS17, PL16a, PN17, PS16, Pei16, PF15, PBC^{+}17, QDRB15, QLF16, Rag15, RH19, RDG17, RSD17, RKO^{+}17b, RHvR^{+}15, SP18, STK^{+}16, SS16b, SwS16, SLY16, SYV14, SGD18, SLA^{+}19, Sti16, TDC^{+}19, TRLK18, WR15, WCN15, WRL16a, WRL16b, WRPL17, WWRS17, WKSS15, WRL18, XX16, XDSX17, XX17, XL16, XWZ^{+}18]. grids [YDLC19, ZA15b, ZSW17, ZRW19, ZKG19, ZCL19, ZH19]. Gross [ATZ16, ABR16, MBM^{+}15]. ground [ATZ16, BJTZ15, Rua18]. group [JPLL15, KA15, LWLC17, MW16b, SWJG19]. growing [Bra16c]. growth [CB19, DMS17, FCW^{+}18, JTR16, LZW19, LTWZ18, RW15b, RTO15, YZW17, YC16, dlCGCA17]. GRP [DL18b, WW15]. Grünwald [MBSS15]. guaranteed [DWGW16]. Guermond [Sir19]. guided [GBS15, LGZ^{+}19]. guiding [PKK18]. GW [LLVF^{+}15]. gyrokinetic [BPL19, CBB16, CB18a, DCD^{+}18, Ido16, KB18, KHC^{+}16, KYPK15, YXX^{+}16]. gyrotropic [CJH^{+}19].

h [CYYL18, CC17a]. h-adaptive [CYYL18]. H-PCFE [CC17a]. Haar [ABP^{+}16, AAE17]. haemodynamics [BFI^{+}16, Gam15]. Hagstrom [AMP16]. Hagstrom-Warburton [AMP16]. half [AS16, GMP16]. half-range [AS16]. half-spaces [GMP16]. Hall [MAH16, SS17c]. Hamilton [CDOY19, CGJ19, DG16b, OS15a, ZQ16a, ZSQ17]. Hamiltonian [QHZ^{+}15, CEF15, EDK19, GAN^{+}16, KML18, LW15b, LHS^{+}19, LY16b, MW16b, MW17a, OLB^{+}17, RFPSSA18, SCN^{+}17, Sun19b, TSC17, TRLK18, ZZT^{+}16, vOMB17]. Hamiltonians [RNO19]. handling [ADGN17, KWHB19, MPR^{+}18]. hard [CT15, Cos16, KBK15a, SAH17]. hard-core [Cos16]. hard-sphere [CT15]. HARM [RKO^{+}17b]. Harmonic [PAFT19, BG19b, CJH^{+}19, DGL^{+}15, ETAG15, LY19, MSG18b, RM16, RPC^{+}18, SC18b, ZC19a]. harmonics [DBSS^{+}19]. harvesting [BGRC19]. Hasegawa [HK15a]. Haut [AS17]. HDG [MTBT18, SCN^{+}17]. HDMR [JL15]. heart [ANL^{+}16, KDPK15, NCP^{+}17, SBG^{+}17]. Heat [FS15, CMP19, CP16, CLG^{+}19, DRZ^{+}19, DPRZ16, EMS^{+}19, GSL^{+}19, HG17, HDA^{+}18, HC17, JL17c, LLB19, MBHS17, PLC18, STK^{+}16, ST15, VBG16, WSP17, WL18, WED15, YK15]. heat-conducting [DPRZ16]. heated [HM19c, KHP17]. height [OD15]. helices [XR17]. helicity [Suz18]. helicity-preserving [Suz18]. Helmholtz [ABFR16, Azi19, BBF^{+}17, BDK^{+}17, CDC17, Cha16, CHCC18, CMH15, DLS15, EFHZ17, EG16, FQZNZ18, GM19, HK18b, JHPAT17, JS19, LGB16, LQB16, NPRC15, OLV16, RSB16, SLR^{+}16, SwS16, Sto16, TCD17, WA18, YL17, ZND16]. hemodynamic [ISP^{+}15]. hemodynamics [DFGQ16, MSV^{+}16]. Hermite [AS16, DL18b, ECC18, HXB15, LIW18, Nor15, ST18a, TLQ15, TLQ16, WC19, YLBL16, ZQ16a, ZSQ17]. Hermitian [VYP15, ZD15a]. Hessian [XGZ19]. heterogeneity [BRK17, CGG18]. heterogeneous [ABG^{+}15, BST19, BC18c, BM16, BSWG15, BKKRB16, CGMH18, CHCC18, CEL^{+}18b, CFvKH18, DDV^{+}15, DD16b, FQZNZ18, FCL19, FC19b, GFG^{+}15, GVTQ16, HL15b, JT18, KYKS19, LKK17b, LJ19, LZT17, LMC19, MGKG17, MSS16, NWB19, SNSG16, SPX^{+}18, SAEF17, SMG19, SHP^{+}16, TKB^{+}15, TWH15, TAH16, TMT17, WRL19, YGEM17, YFC19, ZAK15, ZHWQ18, dMRHJ17]. hex [RGW16]. hexagonal [GHL^{+}16, RKRGW17]. hexahedral [WHY18]. Heydari [Pan20]. HFVS [CJL16]. Hidden [RK18]. HIE [Ram18]. Hierarchical [BABD16, DH18b, PK16, TS17, AAE17, CC19b, CCB^{+}19, LMBZ15, OS16, RBI18, SA15, XQ17]. High [AD17, ABFR16, And16, ADK^{+}17, ABH18, ABR16, BJRF18, BNM15, BKO18, BST^{+}18, BK16b, BDZ15, BB19, BD18, BPD19, BMR19, CKK18b, CLX15, CLTX15, CXY19, CC16c, DDJ19, DIX^{+}18, DS16, DLN15, DLC15, DDH^{+}18, DCD^{+}18, DPRZ16, DPRZ17, FP16, GLK19, GMA18, IDSG15, JLQX15, JFS17, KW15a, uKHGK19, LX16, Li17, LZ17b, LS16c, MRM16, MNR17, MS15c, MW16a, MDHC15, MC15, MP15b, MM16c, MSH^{+}15, NMC15, NRS19, OV17, PxRS17, PLHA18, PKA^{+}16, QPK19, RA17, Sch16b, Shu16, SYV17, SY18b, SC18b, TLQ15, TLQ16, TLH15, VAD17, WL17, WT15, XDLX19, XJLQ15, XYG19, YK18, ZCQ19, ZCQ20, ÅN19, AHNF15, AD15, AGC19, ALM^{+}17, AMP16, APKP16, AA19, ABH^{+}19, ALMJ15, AZ19b, ANL^{+}16, BZ19a, BD15a, BAGK16, BZ16a, BMR^{+}16, BVT18, BGG16, BZ18, BFT17, BQRX19, BSM16, BDPM18]. high [BTT18, BFTVC18, CGQ18, CBB16, CBS18, CJK^{+}19, CB15, CGMH18, CCK^{+}17, CDL17, CQ15, CJL16, CS17a, CZ17, CSC19, CM19, CZL18, CVK16, CFST16, CGJ19, CLQ17, Cif19, CL19b, Cot18, CLP16b, CWJ18, DC18b, DWGW16, DM19, DYL19, EGO19, EDC19, ECC18, FDKI17, Fal17, FQZNZ18, FAZ16, FWK17, FBJS19, FBM16, FK17, FHA16, FHA18, FAC^{+}19, FYO^{+}15, GM19, hGwSzS15, GFC18, Ger17, GSL18, GS18, GGL^{+}17, GZLH19, GGT15, GEZK16, GK19c, GY15, HAPK15, HK19a, HTZG17, HBR15, HW16a, HN17a, HN17b, HN18, HF18, HLQ16, HH19, Jer19, JZSX18, JTD16, KC17a, KCW17, KH17, KRFV16, KYW^{+}16, KYW^{+}18, Kou16, KFWK17, LMH16, LMS17, Lap16, LL19a, LLP^{+}16, LAL18, Ler15, Ler16, LSWF16, LLL16, LW17b, LL16c, LL19b, LGB16, LM19a, LM19b, LWB^{+}16, LW17d, LSZ18, LIW18, LMB19, LCCZ19, LKSM17, LWJV19, LQB16]. high [LP17b, LSI16, MCIGO19, MLM18, MKSP19, MS16b, MNG15b, MA17, MN16a, MTT19, MKC17, MDM^{+}15, MTM19, MA16, MSB^{+}16, MMPS17, ML16, MB15, MSA19, MM16d, NYNYM15, NPB19, NGPB19, NMM15, NJ15, NN15a, NNW17, NL18a, OLDN17, O’S15b, OB19, OSKN18, PE16a, PES19, PP18b, PKW17, PAFT19, PE16b, PMB18, PBC^{+}17, QWX18, QSB18, RNO19, RXSG15, RXS16, RLGT19, RGPS17, RN18a, RSB16, RJ19b, RA19, RRMF^{+}19, RWN18, SZN19, SZN20, STHW17, Say17a, Say17b, SSVL18, SL18, SC18a, SWLZ15, STG17, SYV14, SLN15, SPZ18, SGT16, SGT17, Sti16, SK18, SWZW19, SWL19, SK15b, Tao16, TD18, TK12, TK15b, Ter18, TMH16, TMH18, Tre16, TBG16, TB18, TKP16, Tso18, TYROG19, UG16, VPV^{+}17, VN15, VWV17, VLV19, VSM16a, VSM16b, Vil19, VBF15, WW15]. high [WLM15, WZ15, WSY15, WCN15, WRL16a, WRL16b, WTGC16, WRPL17, WLW^{+}18, WDT^{+}19, WY19, WCWY19, WSR15, WDGW17, WGME17, XQ17, XTYL18, YCPD15, YFJ17, YDN19, YT19, ZP16, ZZK16, ZL15b, ZZZ17, Zha17c, ZRW19, ZY19, ZED15, ZZ19, ZXL17, ZS17, ZZKP19, dLDG^{+}18, DL18c]. high-accuracy [CBB16, Fal17, WZ15]. high-aspect [Sti16]. High-dimensional [QPK19, AZ19b, BVT18, BGG16, CB15, CQ15, CM19, CVK16, Cot18, FDKI17, FK17, GK19c, Kou16, LLL16, LL16c, LL19b, LW17d, MTM19, RNO19, TBG16, WCN15, WTGC16, ZZK16, ZZKP19]. high-energy [LMH16]. High-fidelity [MS15c, MW16a, LSWF16, MS16b, PKW17, UG16, VBF15]. high-frequency [CDL17, CLQ17, FQZNZ18, HBR15, NNW17, Tre16]. high-level [ZED15]. High-Order [BD18, BPD19, DS16, AD17, ADK^{+}17, ABH18, ABR16, BNM15, BKO18, BMR19, CKK18b, CXY19, CC16c, DDJ19, DLC15, DDH^{+}18, DCD^{+}18, FP16, GLK19, IDSG15, KW15a, LZ17b, LS16c, MRM16, MDHC15, NMC15, NRS19, PLHA18, RA17, Sch16b, SC18b, TLQ15, TLQ16, TLH15, WL17, WT15, XDLX19, XYG19, YK18, AHNF15, AGC19, AMP16, APKP16, ABH^{+}19, ALMJ15, BZ19a, BAGK16, BZ18, BFT17, BSM16, BTT18, CBS18, CJK^{+}19, CGMH18, CCK^{+}17, CFST16, CLP16b, CWJ18, DC18b, DWGW16, DM19, DYL19, EGO19, FAZ16, FWK17, FHA16, FHA18, hGwSzS15, GFC18, GSL18, GGL^{+}17, GEZK16, GY15, HK19a, HTZG17, HN17a, HN17b, HN18, HF18, HLQ16, HH19, JZSX18, JTD16, KC17a, KFWK17, LMS17, LAL18, Ler15, Ler16, LGB16, LM19a, LM19b, LW17d, LIW18, LMB19, LWJV19, LSI16, MCIGO19, MLM18, MNG15b, MA17, MN16a, MKC17]. high-order [MDM^{+}15, MA16, MMPS17, MM16d, NJ15, O’S15b, OSKN18, PAFT19, PE16b, PMB18, PBC^{+}17, QWX18, RXSG15, RXS16, RLGT19, RN18a, RSB16, RJ19b, RA19, RRMF^{+}19, Say17a, Say17b, SC18a, SWLZ15, STG17, SGT16, SGT17, Sti16, SWZW19, SWL19, Tao16, TK12, TK15b, Ter18, TMH16, TMH18, TKP16, Tso18, TYROG19, VPV^{+}17, VN15, VWV17, VLV19, Vil19, WLM15, WSR15, YCPD15, YFJ17, YT19, ZP16, ZL15b, ZZZ17, ZY19, ZXL17, dLDG^{+}18]. high-order-accurate [OLDN17]. high-order/low-order [CCK^{+}17]. high-orders [VSM16a, VSM16b]. High-performance [PKA^{+}16, RGPS17]. High-performance-computing [DLN15]. high-plasma-frequency [BZ16a]. high-rank [Jer19]. High-resolution [GMA18]. high-Reynolds-number [NL18a]. high-speed [BMR^{+}16, GEZK16, MSB^{+}16, QSB18]. high-temperature [LL19a]. Higher [APP^{+}16, BH18, BC16c, BCJ19, GS16, HSLQ16, JC17, LTXB17, LAK^{+}16, MRY19, MS18c, PYAG19, SGC18b, Sub18, Tie16, WR16, BGTM18, BMMP19, BPF^{+}16, DPO16, DNBH15, DM17b, FRRV16, HB16, JYY18, KVKS19, LBTCG16, LW17c, QPK19, Rua18, TR19, WSOW16, XY18, YK19, ZLL16b, ZLL17a, ZSO^{+}19, ZS18, ZS19a]. Higher-order [APP^{+}16, BH18, BC16c, BCJ19, GS16, JC17, LTXB17, LAK^{+}16, MRY19, MS18c, PYAG19, SGC18b, Sub18, Tie16, WR16, BPF^{+}16, DPO16, DM17b, FRRV16, HB16, LBTCG16, QPK19, Rua18, TR19, WSOW16, YK19]. Highly [KB19, ABG^{+}15, FYC^{+}18, GCI19, GXX17, IKS19, JCWX19, JS19, Mon19, RKO^{+}17b, ST18c, TT17b, WA18, YS17, YFC19]. Hilbert [WTL19]. Hilliard [BKR19, HW15a, ZYCK15, CS16c, CLS^{+}18, DD16a, DJLQ18, GX15, HTMP17, KS16a, KMdB16, LJZ15, LCK16, MGCW18, SL19b, WX17, XGZ19, YLD19, ZSX17]. Hilliard-Brinkman [GX15]. Hinsberg [CFO18]. HLL [Bal15, DG16a, FLW16, SW17a, VNA15]. HLL- [DG16a]. HLL-type [SW17a]. HLLC [DG16a, GOR17, Guo15, LDGH16, SYY16, SM19a]. HLLC-based [GOR17]. HLLC-type [DG16a, LDGH16, SYY16]. HLLD [GFW16]. HLLEM [DB16b]. HLLI [BN17, BLM18]. Hodge [DPO16]. hodograph [ZX19]. Hoekstra [XS15]. hole [KP15b]. holes [ABT17]. HOLO [CCK^{+}17]. holographic [CDLR19]. Homogeneous [Wu19, Azi19, CGK17, DGHP17, DSS18, HWH^{+}16, MPT16, SUR18]. homogenization [ASWvD19, AR16b, GO15, MVZ16, NGS16]. homogenized [LMM17]. Homogenizing [TY17]. homology [WW16]. homopolymer [Yan16b]. homotopy [BZ16b, CC19c]. HOOMD [YZW^{+}18]. HOOMD-blue [YZW^{+}18]. Hopf [EEG^{+}15, LDOK17, LP17a, RPL^{+}18]. horizontal [FDS^{+}15, SHLG15]. Hosseininia [Pan20]. Hot [HED^{+}16]. hourglass [KSSL18]. hp [CC17a, CK16a, MBM^{+}18, MSP16, RN18a]. hp-adaptive [CC17a]. HPC [BLA^{+}15]. HRSSA [MPT16]. Hugoniot [FKK19]. Hugoniot-conditions [FKK19]. hull [LM15b]. human [ANL^{+}16, FK19, LSCC19, NCP^{+}17]. Hurwicz [DFM17]. Huygens [KLWQ17, LQB16]. HWENO [CQQ16, LHQ16]. Hybrid [BD18, BPD19, BHMS18, CSS15, Cho15, CCGH19, DG16b, DEZ16, FBJS19, Fuj19, HLML17, LZZS15, MJ17, MH17, SSDN15, SW17a, SGA^{+}15, AVT17, ALM^{+}17, Ama18, AdSŚ^{+}15, BT17a, BBKS16, BBB^{+}16, BAK19, BFTVC18, CARdN19, CWM^{+}16, Che18, CZL18, COV18, CFPB17, CBN^{+}16, CG15, CCGH17, CYWL17, DD17a, DZR18, DTA^{+}15, DJV^{+}18, Dod17, Dom18, DJL^{+}19, EARA15, FQZNZ18, FX18, FLW16, GBC^{+}18, HXLL15, HWA15, HLY15, Ido16, KD17a, KF15, KB18, KHC^{+}16, LSLA16, LML^{+}16, LTW18, LM19d, LPBR15, LLD19, LMN18, LHQ16, MLI17, MPT16, MS16b, MR16a, MAM16, MRY19, MN16b, MBA19, MF16a, MDAB18, Niu16, OB19, PL16a, PBC^{+}17, PWP15, RBJS15, SWLZ15, SCJ^{+}18, SCS16, SCC19, TAJ^{+}17, Tie16, WMY16, WPB15, WR16, XWL^{+}16, XDSX17, XX17, XML17, XWZ^{+}18, Yan16a, YWS^{+}16, YX15, YB17, ZZKF15, dlCGCA17]. Hybrid [AB15, GSS15b, RKO^{+}17b]. Hybrid-dimensional [BHMS18, XML17]. hybrid-Lagrangian [KHC^{+}16]. Hybridizable [UL16, NPRC15, SMP16, SWZW19, VCNOP18]. hybridized [BT15, FNP17, MLM18]. hydraulic [CB19, NMM19, RSBS19]. hydraulics [CBN^{+}16, SSC^{+}16, TSB^{+}18]. hydro [CYS17, MRP^{+}15]. hydro-dynamically [CYS17]. hydro-geophysical [MRP^{+}15]. hydrocode [VGZ18]. Hydrodynamic [GA18, BMC^{+}18b, HWA19, KV16, LMB18, LMB19, LCF16, MWB^{+}15a, MWB^{+}15b, MLB18, NJPB17, Ram17, WVB19, WRL18, ZYSW16, ZK18]. hydrodynamically [PBP18, PMGW16]. hydrodynamically-consistent [PMGW16]. Hydrodynamics [AWS16, DRM15, FRO17, KRK^{+}18, MDL16, YLZ^{+}19, AMXJ19, BKS18, BLK15, BLK19, BHE^{+}17, BDLM18, BMCK15, Cap18, CGP16, CG15, CM18c, DDJ18, DDJ19, DD15, DSH^{+}16, Guo15, GFW16, KSSL18, Kwo19, Li17, LLW19, LS16c, LSR16, LDT19, MPR^{+}18, NT15, OB19, PKP^{+}17, PLB18, QSY16, RKO^{+}17b, SKO18, SK19b, TOR^{+}15, TL17, WT15, ZS16, ZHA17b, ZC19b, TP16a]. hydrogels [LJZ15]. hydrology [MRA16]. hydrophobic [Fed17]. hydrostatic [AZ16, EDC19, GDFL17, LX18, YP17]. hyper [LM19c, MG15b, Tsa16]. hyper- [Tsa16]. hyper-elastodynamics [LM19c]. hyper-viscosity [MG15b]. Hyperbolic [GHS19, NL18a, NN16, PMF15, BD15b, BN17, BK16b, BLD15, CNK19, CTG16, CS17a, CH19, CTM^{+}16, DLP19, DL18a, DL18b, DPRZ16, DB16b, DPRZ17, EFT15, EDC19, FLV18, FS19a, FS17b, FHA17b, FHA18, FS19b, GNK18a, HS18b, IDSG15, JL18c, JXZ15, KKS15, KKS16, KA15, KC18, LMS17, LAL18, LMBZ15, LLLN18, LP17a, LHQ19, MSK18, MDH19, MA17, MN15, MN16a, MRN16, MDHC15, MB15, NMM16, NBH18, NN15a, NN18, Nis18a, Nis18b, NW15, OZ17, PxRS17, POSB16, PTT19, PTT18, Ren19, SS18a, SW17a, SL18, SPB18, SPP16b, SMSR18, TLQ15, TM15a, TM15b, TC15c, TPB16, VNA15, WLGD18, WTX17, YJ17, ZPW18, ZQ16b, ZG19]. hyperbolic-elliptic [ZG19]. hyperbolic-equation [KKS15, KKS16]. hyperbolicity [SS18a]. hyperbolicity-preserving [SS18a]. Hyperbolization [VST16]. hypercontraction [LY15c]. hyperdiffusion [URGT18]. hyperelastic [BM16, CWWZ17, PBL^{+}19]. hyperelasticity [HFM17]. hyperinterpolation [LB15]. hypersingular [Tsa15]. hypersonic [BHFB19, CLG^{+}19, GRS15, XWZ^{+}18]. Hyperviscosity [SF18b]. Hyperviscosity-based [SF18b]. hypoelastic [PBL^{+}19]. HyShot [CELI15]. hysteresis [MCHL16].

I2D [RHvR^{+}15]. IB [ZZPH18b, ZZPH18a, PZNG15]. IBM [SHP^{+}16]. IBSE [SGT17]. ice [ALKZ16, AS17, ALTR17, CCB^{+}19, CLvS17, CS18a, IPSG15, KDL15, MDW18, MR17, RW15b, SRBÓ17, WTL17]. ice-covered [MDW18]. icosahedral [Sub18]. Ideal [DWG^{+}18, BK17c, BND16, CFST16, DWGW17, DLK17, FS18, KW15b, LSZ18, PL16b, WSH^{+}17, WG16b, WDGW17, XL16, ZYD19b]. idealized [FCW^{+}18]. identifiability [WSK19]. Identification [CZ19a, CGM15, KM16b, TBLM15, BCB15, EFHZ17, MWZ19, RTG18, RYZ18, ST15, TPA19, ZFPB16]. identify [SPM^{+}15]. Identifying [LVL18, WTS^{+}17]. IGN [ZED15]. II [DLC15, BD15b, BGTM18, BHST17b, BFFB17, CELI15, DL18c, EFT15, GPS17b, GY17, LZZS15, MBNJ16, MS18b, PxRS17, PSN^{+}19, QN19, SZM19b, Say17b, Sch16b, SW16, SHP^{+}16, SiI17, TKC15, VSM16b, WRL16b, WCWY19, ZLH^{+}17]. II. [Cac15b]. III [BN17, BFF19, DLN15, GY18, LB15, WRPL17]. IIM [LXC^{+}15]. ill [CCB^{+}19]. ill-conditioned [CCB^{+}19]. illustrated [NS19a]. Illustrative [Cac15b]. Imada [Zil15]. image [LGZ^{+}19, NL15, SG18, WLWW17, WCVF16, YS18b, ZC15]. image-guided [LGZ^{+}19]. imagery [EFHZ17]. images [CV15]. imaginary [ZS19b]. imaging [AJP15, BCM19, CR17, CDLR19, Gib18, HZ17, LWZ16, Liu19a, MCHL16, Par15, PKLS17, SGP17a, WSU^{+}15, ZF18]. imaging-based [MCHL16]. Imbalance [KS16d]. Imbalance-correction [KS16d]. IMEX [ABR16, BLMY17, BQRX19, CEHM19, JLQX15, MCS^{+}19, VN15, XJLQ15, XQ17, SSZ19]. IMEX-spectral [ABR16]. IMEXP [LTR17]. Immersed [AB16a, BDG^{+}17, MCW16, MHL17, OM19, SGT16, ZILZ15, AZ19a, ACS16, AB17, AEvW19, BKP16, BPS17, BHF15, BBK19, CCHL15, CLM15, CYWL17, De18, FR18, FG16, FG19, FKY15, GWC18, GMWC19, GLMC16, GC17, GSL^{+}19, HWH^{+}16, HS18a, HF18, HHL19, HLY15, HLSY16, JL17a, JBM19, JL19, KLSF15, KLC18, LTB16a, LC15, LFDP16, LBZA16, LYL19, LC17b, LD15, LLA19, LWTF19, MAP17, MKS18, MM16d, NJPB17, NPB19, NBZ^{+}19, ÖPHA15, PNZ18, PPLC16, PN18, PG18, PVB17, QSB18, RS16b, SKF15, SKF16, SHKL16, STKL19, SMA^{+}16, SMOM^{+}17, SLdTV18, SHP^{+}16, SCLG15, TZGW18, TKF17, WE15, WSY16, WCH^{+}17, WVB19, WS15a, XLY15, XP15, XTYL18, YS15, YXF^{+}16, YZZ15, ZB15, ZSM19, dTP16, GY17, PF16, RMF^{+}18, SGT17]. immersed-body [YXF^{+}16]. immersed-boundary [BKP16, GC17, LD15, PVB17, YZZ15]. immersion [SWS17]. immiscible [AASRT17, CM18d, Don15b, Don17, Don18, FGL16, HTZG17, KKH18, WG19, YD18]. Impact [CSLL15, HCW15, KBF17, TT17a, URGT18, LBZA16, RHS18, YCBC15]. impacts [Heu17]. impedance [BG16a, DSSP18, JSP16, LR19a, MS15a, MMP18, SS17a, Yam19, ZZH16, dFGS^{+}17]. imperfectly [SPB17]. imperfectly-mixed [SPB17]. impinging [Gan15]. Implementation [ALTR17, BT17b, DTA^{+}15, HdBH^{+}16, JD19, PJE^{+}16, SA15, YZW^{+}18, BVS18, BG16a, CDC17, DY17, FLW19, HWK19, HK15b, JSP16, KC17a, KBF17, LZ16, MTT19, MM16c, PM16, PKA^{+}16, SRBÓ17, SE15, TRM16, ZZH16]. implementations [SBT17]. Implication [DDM^{+}19]. Implicit [CT19, Du18, GSL^{+}19, LMH16, Nis18a, PLWJ16, Say17a, Say17b, SD18, TM15b, ZZX16, AVT17, AGKD19, AZ19a, AB17, ABIR19, BPL19, BR15a, BFNGDNR18, BZ18, BQRX19, BCJ19, Cai16, CZBC^{+}18, CB15, CBC^{+}18, CC16a, CS16a, CC19a, Che18, CC16c, CG15, CW18, CTM^{+}16, CM16b, CLP16b, CCGH17, CCGH19, CLNH15, CvKH16, Del15, DLMDV18, DJLQ18, ESHA16, Fer17, FS19a, Gam15, GS15b, GZ17, Gen11, Gho17, GXX17, HPY18, HK18a, HDA^{+}18, HW19, HCLT19, JRPPS18, JZ16, KC17b, KL18b, LSMS17, LLD^{+}16, LH17a, Lap17, LYM19, LXC^{+}15, LL16b, LT17a, LLSJ19, LCCZ19, LTR17, MM16a, MNO^{+}17, MH18b, MTJ17, MTJ18, MMPS17, MPMB19, NMC15, NWKC16, NNW17, Nis19a, NLW^{+}16, OB19, OZ17, PKP^{+}17, PPLC16, PP17, PGM17, PLPR19, PME^{+}15, QWXZ17, QDRB15, Ram17, RMBN18, RAMB15, SXBB15, SLH18, SL19b, SKW19]. implicit [SZ17, SCC19, SPCH16, TCSM15, TD17, Tie16, VV16, WCN15, WH16a, WMYG16, WHT18, WKHT19, XXR18, YSLY19, YSYW19, YD19, ZBH^{+}18, ZMCC18, ZYCK15, ZRT18, ZZX19, FNP17]. implicit-explicit [BPL19, BCJ19, DLMDV18, MPMB19, PGM17, SLH18, SL19b]. implicit-exponential [LTR17]. implicit/explicit [CB15, Tie16]. implicitization [DV17]. implicitly [RB18]. implosion [FCW^{+}18]. Importance [CCL16, OCSC18, RC18, LL15, URGT18]. imposed [ADE^{+}17]. Imposing [KLSF15, MNR19, YS18a]. impressed [PYAG19]. Improve [HLL^{+}18, AB15, GO15, LWY18, NJL19, SIX16, vLtTBI17]. Improved [MN15, NMM19, NSK^{+}16, PA19, TZSS17, ZL15a, AdRBC16, AJVH17, BLL19, BLL20, BKP16, BSP18, CCRdL17, CZ16, CZ17, CH19, EDC16, GH17a, GRS15, HZ17, KMS^{+}18, KC17b, LS16a, LWTF19, MO18a, MRK15, MTBT18, STG17, SK18, TND18, WSY15, WLGD18, WWX19, YSYW19]. Improvement [SY17, BK17a, CSY19, OMLdL16, PSMPG17, RW19, WBC^{+}16, ZOG19]. Improvements [ACCCDA16, Ani16, HPV16, COdLL18]. Improving [GO16, ID17, PDS15, SM16, JW15a, KK16]. impulse [FHS17]. impulsively [GWC17, SHW17]. in-cell [KKH18]. in-depth [DV17]. in-plane [DBD^{+}17]. in-space [Har18]. incident [BFP18, GD19]. include [HMRG16]. Including [Gen11, Gho17, BM19b, BKG15, Guo15, HHM17, Hig17, LB15, SGP17b]. Inclusion [dDPG19, TSR15]. inclusions [DCA^{+}16, LKB15, WL18]. incommensurate [ZCZ19]. incomplete [GSC19]. incompletely [WN18]. incompressibility [ZZKF15]. Incompressible [LSR16, RV16, ZZS^{+}17, ACS16, ALL18, AEvW19, BHST17a, BHST17b, BHST18, BFI^{+}18, BM19a, BCB17, BPD19, BFTVC18, CCRdL17, CS16c, CX16, CCKQ15, Chi19, CS17b, CCM17, CLP16b, Don15a, DS15c, Don15b, Don17, Don18, Fan16, FWK17, FWK18, FLV15, Fer17, FZ19, FG19, GN19a, GTG15, HPY18, HGW18, HP17, HTMP17, HLA19, KCHW19, KCSW19, Kla15, KW16, KFWK17, LVTR15, LE16, LRA17, LM18, LHB^{+}16, Li17, LZ17a, LLSJ19, LYD19, LS16c, LC16, LC17b, LH18, LFAR19, LZW^{+}17, LHA15b, MHHX16, MS18b, MLM18, MC15, MH19, MPFL16, MP19, MS18c, MHS16, MR16b, MN18c, NT15, NGPB19, NYD19, NN19, OVP15, PG17, PKP^{+}17, PL16a, PM19, PPLC16, PF16, PND16, PBBK15, PMB18, PQR17, QYF15, RBJS15, RDM15, SV19, SL17, SBHS19, SMS16, SY18a, SLY16, SSO^{+}15, SLA^{+}19, SGT17, SST^{+}15, Suz18, SSZC19, TLH15]. incompressible [TD16a, TGS19, TOR^{+}15, Tou18, VV19, VL15, VK15, Vre17, Vre21, WDG^{+}17, WSS^{+}15, WSHT15, WZ18a, WWZ19, WSF17, XWL^{+}16, XX16, YSWS16, YD18, YD19, YT19, YZZ15, ZN18, ZCL19, dFJN16]. Incompressible-compressible [LSR16]. incorporated [LHW^{+}17]. increased [DBZ17]. Increasing [Die15]. increasingly [KMGR16, ZS18, ZS19a]. Incremental [SKS17, CBN^{+}16]. independent [Bre18, CLM16, DLP19, DDM^{+}19, HDF18, WDG^{+}17, WT16, YS18a]. index [LTKA15]. indicator [FS17b, HC18b, KC18, RRS19b, RRS19a, RH18]. indicators [NdlLPCC19]. indistinguishable [SD15]. induced [BLL19, BLL20, BPGS16, FRL15, HDA^{+}18, KPKGH19, LR19b, LLB19, PYK19, YR15]. inducing [LAA16]. induction [ACC^{+}15, BK17b]. inductively [MNO^{+}17, TC15d]. industry [VWV17]. industry-standard [VWV17]. inelastic [CKKD19, HM19c]. inequalities [KRBW17]. inequality [MCN18, OKE17, YSY17]. inertia [MDP^{+}15]. Inertial [Ram17]. inexact [LSCC19]. inextensibility [Vog17]. inextensible [RV16]. inference [CZB15, HKKP16, HYL17, IPSG15, LZ18, LYLK17, MPP15, NS16, SPP^{+}16a, YZ19]. Inferring [RPK17a]. infiltration [MRP^{+}15]. infinite [And16, BGL^{+}17, CZB15, GBD17, HYL17, Lan19, MJ17, SHLG15, STFK19, ZBZ^{+}18]. infinite-dimensional [BGL^{+}17, Lan19]. infinitesimal [DW19]. Inflow [KHHN16, CSLL15, ST18a]. Inflow/outflow [KHHN16]. Influence [SC18a, SWLW19, MDMS18]. Information [GKRB17, Fan19, KRBW17, LKK17a, LKK17b, LSWF16, RMK15]. informed [CLM16, CMW16, RPK19, XWW^{+}16, YP19, YBSTT19, ZLGK19]. inhomogeneities [Liu19a]. inhomogeneity [AJP15, LPWK15]. inhomogeneous [APKP16, CLQ17, DG16c, LM15d, OLV16, PA19, RMA17, SYOS19, SYOS21, SC18b, TNB21, TT19, Yan16a, YJ17]. initial [DZC16, MM15]. initial-value [DZC16]. initialization [LHGF19, Wac15]. initio [Gen15, LLVF^{+}15]. injection [KS18a, dCPDC^{+}17]. inputs [LL17, JXZ15, JL17c, LL16c, Liu19b]. Insights [MSP16, KS16b]. Insilico [HED^{+}16]. INSIM [GR19]. insoluble [HCLT19, SA16, dJRP^{+}15]. instabilities [MCS16, XLL^{+}17]. instability [CCZ15, DNOP15, MHZ^{+}15, RLV16, SM19a]. instantaneous [DLP19]. instationary [AMM^{+}15]. instrumented [GLG^{+}19]. Integral [BMR19, Vee16, AAE17, AAE19, ABN15, BC18a, BRK^{+}18, BNM15, BKO18, CCZC16, CG18a, CLX15, CGRV17, CRZ17, CV18, CCGH17, DvB17, Dod17, FAY19, GCI19, Gen15, GD19, HK19b, HHCG15, HCLT19, HSSZ16, JL16, KHP17, LGO17, LDL^{+}16, LZ19, LSP^{+}18, LY15c, MCIGO19, MKYZ17, MS17, Moh15, OC18, OT15, PA19, PLL15b, QM18, RVZB15, RMA17, RW19, SL16a, SO17, Smi18, ST18c, STV19, SV17, TP17, Tsa15, Tsa16, WZ18a, XZ15, XZJK19, Yan19, ZMZC19, ZGD^{+}16, ZRT18, Zil15, aKT16]. integrals [BPF^{+}16, LO16, TB19, Tsa15, Tsa16]. integrated [SSC^{+}16, ZHWQ18]. integration [BCM15a, BBBG15, CL19b, CCGH19, EBQ15, EMSS17, FCL17, GZY16, GP16a, HEPG15, JG19, JZ16, JFS17, LMS17, LLVF^{+}15, LCCZ19, LPR19, MM16b, MKB19, MS18c, Mon19, MTK^{+}16, MPMB19, NDCB17, PKK18, SXBB15, SGL18, SAOW17, TWN15, TW17, TC15c, WCN15, Web14, WBC^{+}16, WHE17, WHEK18, Yi18, YK19, ZJS15]. integrations [DY19b, RMK15, SKW19]. integrator [BZ16a, HJW19, LLWJ18, SS18c, WSR15]. integrators [CSS17, ETL17, FPASS16, GAN^{+}16, GRT18, GRT21, KTG16, LW15b, LWL17, LIW18, LTR17, MHH19, RFPSSA18, SL19a, SSZC19, Tao16, WW18, WW19, WZ18b]. integrodifferential [BCC^{+}18]. Intel [SGL18]. intense [Vai15]. inter [LLB19]. inter-phase [LLB19]. interacting [CGSS18, GBS15, JBM19, MFB18, MM18, SGMS16]. Interaction [CLM15, TFGK18, AQ19, AMB17, BQCG17, BCM15b, Buk16, CGSS18, CZBC^{+}18, CM16a, CDM18, CH17, CM16b, CYWL17, DG18, DFGQ16, EKSS15, FW18, FLV15, GKMS17, GLS15, Har18, HW19, HDF18, HLL^{+}18, HTBG15, KH15, KLC18, KC17c, LTB16a, LC15, LZHM19, LLY18, LHY^{+}19, LGD17, LZS^{+}19, MPR^{+}18, MTK17, MAA18, MOR18, NPB19, NBZ^{+}19, PHHA18, PHÖ^{+}16, PR16a, Rua18, Say17a, Say17b, SSL^{+}16a, SSVL18, SA16, SGC^{+}17, SMP16, SLB^{+}19, SMOM^{+}17, SCS18, TT19, Vai15, WCH^{+}17, Wic16, YXF^{+}16, dTP16]. interactions [ATZ16, BHKS16, BJTZ15, BTA17, BPM18, Cos16, DG16a, DWZ19, FRW16, HWH^{+}16, LML^{+}16, LXL17, LSP^{+}18, LGZ^{+}19, MKV^{+}17, PSV18, SG19a, SGC^{+}18a, SMA^{+}16, WMY18, WVB19, WNW^{+}19, WSB19, XTYL18, YS15]. interactive [CLFL17]. interatomic [TST^{+}15]. intercellular [CFG16]. interchange [Sov16]. Interface [ABG^{+}15, CNG99, DS15a, GZ17, GWYS18, GLTG15, GPG17, VK15, AAL15, AMN18, APKP16, ACS16, BKS18, Bar19, BJ15, BT19, CCHL15, CNG17, CTJ^{+}17, CWW17, CZ19b, CR18, DS15b, De18, DMRB19, DXvW18, DvW19, DF16, FB17, FMRZ17, GHR17, GLTB18, GHF19, GWC18, GMWC19, GD19, GY17, GY18, HHA15, HWH^{+}16, HTZG17, HG17, HHL19, HLY15, HLSY16, HW15c, IM15, JBM19, JLC15, KTK15, uKHGK19, KS16c, KS18b, KSVB18, LSD^{+}17, LSR16, LD15, LHA15b, MNG15a, MCW16, MNR17, MTZ16, MA19, MWYZ16, NFG15, OD15, PD19, PHHA18, PA19, PN18, PR17b, QWX18, RW15b, RV16, Say17a, Say17b, SA16, SHA16, SLC^{+}18, SCJ^{+}18, SGD18, SA15, SR18, TDC^{+}19, VPM15, WSS^{+}15, WZLS19, WL17, XLY15, XX17, XP15, ZL15b, ZD15b, ZDGW16, ZFL^{+}19, ZY19, ZSM19, TKB^{+}15]. Interface- [DS15a, DS15b]. interface-aware [BKS18]. interface-capturing [Bar19, WL17]. interface-compatible [KSVB18]. interface-conforming [ZFL^{+}19]. interface-correction [GLTB18]. interface-enriched [SA15]. interface-fitted [CWW17]. interface-interaction [PHHA18, SA16]. Interface-preserving [GWYS18, ZY19]. interface-resolved [PD19, SLC^{+}18]. interface-sharpening [HTZG17]. interfaces [ADGN17, ADN19, AB18, BG19c, BAR15, CZL^{+}15, CSN17, CLM15, DIX^{+}18, EN18, ELH^{+}16, HKLZ18, HGR16, KKLS17, LSMS17, MAK15, MNR19, MF17, NN17, OS16, OCSC18, PCN15a, PR16b, PS14, PS15a, PC19, QDRB15, SMOM^{+}17, WXW15, WB17, ZILZ15, ZMZC19, dFVJ15]. interfacial [DXvW18, DvW19, GOR17, GZLH19, HWK19, KRK^{+}18, LHA16b, NWB19, RW19, Say17a, Say17b, SZCL18, ZFL^{+}19]. interferences [CSC19]. Interior [MRRRF18, Fer17, LFAR19, OKE17, PKA^{+}16, DM18, QN19, SL17]. interiors [BLC^{+}17]. interlayers [SSL17]. intermediate [PDS15]. intermetallic [ZWYW18]. intermolecular [SWLW19]. internal [BD15b, BN17, Guo15, MCS16, SVG18, ZSM19, vOMB17]. interphase [HG17]. interpolating [WLK^{+}16]. Interpolation [SNK18, dPSS16, APP^{+}16, ADN19, ABdC^{+}18, BDG^{+}17, BDAA^{+}18, BST15, CGM18, CDM19, DJD^{+}17, ECC18, FYO^{+}15, HSC16, JYY18, JWH16, KAR17, KMGR16, LB15, MCW16, MAP17, MBD15, MBD19, NMA15, PJC16, PAFT19, PF15, RDG17, SLA^{+}19, SHTY19, Sun19b, WR15, WKOE17, XYPT16, ZWB^{+}18, FFBB16]. interpolation-free [RDG17]. interpolative [BBB15, LT17b]. intersecting [BPS17]. intersection [CZJ17]. interstitialcy [BBW16]. introducing [TTN^{+}16]. Intrusive [NWFT19, Blo17, CPX19, CMP19, HFND18, HU18, Poë19, TO19, WHR19, XYF^{+}17, YYJ^{+}19, vdBKD17, NW17, NT19]. intuitive [ZW15]. invadopodia [GP17]. invariance [BKP16, GHL15, LJT16]. Invariant [JL18c, YZW17]. Invariant-region-preserving [JL18c]. invariants [Hue15, LDHJ15]. Inverse [DDJ18, LBTCG16, LLL16, LFT^{+}16, AJP15, BJO18, BCSK17, BSK15, BtTBI18, BGL^{+}17, BDMZ19, BGHK19, BKL17, CT15, CGM15, CMW16, EZG16, FK17, GHV19, GZY16, GRMK15, GWE^{+}15, GHH^{+}16, KE15, LPU18, LW15a, LL19b, LY16a, LS19b, LMTC15, NKN^{+}17, OLD^{+}18, Par15, RPK19, RYZ18, SL19c, TCD17, WL18, WLK^{+}16, WSK19, YZ19, ZF18, ZLGK19]. inverse-power-law [CT15]. inversely [LGZ^{+}19]. inverses [For16]. inversion [AZ19b, BST19, BFP18, CS16b, LL15, LLY15, MKYZ17, MRP^{+}15, PKN17, PD16a, YLH^{+}19, dFGS^{+}17]. inverting [XL17b]. investigate [MZ15, WPB15]. Investigation [BR15a, CV16a, PKB15, VS17]. investigations [ZZ17b]. inviscid [BR15b, BR16, FKK19, LLWJ18, Loz17, PL18, RDG17, YSW15]. invisible [Chu17]. involving [Don15b, FSWW17, FS15, LGD17, RPK19, TBHG18, YD19]. inward [LWTF19]. Ion [dCPDC^{+}17, CCZ15, KB18, MP16, TTN^{+}16]. ion-acoustic [CCZ15]. ion-electron [MP16]. ion/fluid [SCC19]. ionic [YX15]. ionization [CV16b, LYCC17, YCBC15]. ionization/recombination [YCBC15]. ionized [PMS15, Zoh17]. ionizer [Fon16]. ions [SPCH16]. IP [XJG18, XWZ^{+}18]. IPDG [CLQ17]. irradiated [HMBH15, MSF^{+}19, TT17a]. irregular [ABG18c, CXL16, ECC18, GZLH19, GLTG15, LC18, LPGT16, LZT^{+}15, MDM^{+}15, Tow18, YYN^{+}17]. irrotational [LM16, OM19]. ISCAL [ALKZ16]. isentropic [BQRX19, PCBG18]. Ising [PT17b]. Ising-like [PT17b]. island [MGB^{+}18]. Isogeometric [BLJ17, BG16b, HTMP17, KMdB16, PXXZ15, PCX17, PRXC19, CRMP16, KC17c, LDL^{+}16, LWZ19, OWKE16, OKWE17, SLVE18, WKOE17, ZSX17]. isoparametric [Pas16]. isothermal [BLVC16, BS19b, KB18, OTS17, RJLW19, TUJ19, TXKvdV15, XML17, ZCSZ19]. isotropic [An17, CLS^{+}18, GK19a, RZZ19, SS17c]. ISPH [HKH^{+}16, KGS17, KHHN16]. Issue [KHP15, KZ17, KZG16, Kat16, TM17]. issues [NT15, WSB19]. Itô [AÁPB17, HHCG15, Moh15]. Itô-SDE [AÁPB17]. item [sCYxL^{+}18]. Iterated [ALL18]. iteration [BW18b, HB15b, KFL17, PBBK15, ZHLZ18]. iterations [CDLR19, TWM18, WZ17]. Iterative [AA15, GLZ16, HKLW15, KPP^{+}19, SHW17, AP16, AC16, BDK^{+}17, BDKK17, BSWG15, CCHL15, CDL17, CDC17, DGW18, DDV18, EAAM15, GM19, GWC17, KA15, Lau17, LR19a, LM19c, MM16a, MRP^{+}15, MVZ16, MBBKTH17, MTBT18, NDH19, NS19b, NNW17, Pea15, PE16b, PSP16, RMP18, SXBB15, SZ15a, SGD18, SWK18, TKF17, WLWW17, WZL^{+}17, XXR18, YS15, YLBL16, YJB18, YL17, ZF18, ZJL16, ZZ19]. iteratively [HHLY17]. itself [MSG18b]. IV [LXSC16, ZRW19].

J [ABG^{+}19b, ASS17, BLL20, CNG17, Dav15, DK18a, DvW19, EFO20, GRT21, Gho17, GBCF16, Hag25, HGN17a, KYW^{+}18, KTK19, MN17, NG18, PS15a, SZN20, SYOS21, STEK22, SWMD17a, SYV17, TK15b, Vre21, ZJS15, ZCQ20]. Jacobi [ACCCDA16, ACCCD^{+}17, AA15, BZ15, BDBEE15, CDOY19, CGJ19, DG16b, KJHA19, OS15a, PSP16, Tsa16, VBL^{+}16, YM17a, YM17c, ZQ16a, ZSQ17]. Jacobian [AB17, ALTR17, MTK^{+}16, SPW18, SLN15, YKMM19]. Jacobian-free [MTK^{+}16, YKMM19]. Jacobians [OKE17, RÖS17]. jet [GP16c]. jets [AMM^{+}15, RVK^{+}18]. Joint [SG18, AKK^{+}19, SGP17a, dFGS^{+}17]. Journal [BR16, EH15, HSK^{+}15, Kat16, Pan20, XS15]. July [Kat16]. jump [AMN18, HWH^{+}16, LYL19, MNR19, XZZ15, XP15]. jumps [MNR17, NMM19]. Junction [CTM^{+}16]. Junction-Generalized [CTM^{+}16]. junctions [BLVC17, MLB16].

K-BKZ [TBO^{+}16]. Kadomtsev [EO15]. Kalman [BJO18, MP17, ZH15]. Kansa [PCF15, SLZ^{+}17]. Kapila [tEDKT17]. Karhunen [CN16, LC18]. KDV [BNS17, EAAM15, LHQ16, ZHS18, dlHC16]. KdV-type [LHQ16, ZHS18]. Keller [ZM16a]. Kernel [DJD^{+}17, FRO17, MOAA15, SPB17, BKP16, CC19b, CGJ19, GH19, PvL19, RFGSV15, VD16, WTL19, YS18a, OY19]. Kernel-based [SPB17]. Kernel-conformation [MOAA15]. Kernels [EJZ17, CDC17, JAH19, OY19]. Kershaw [Sch16a, Sch16b]. key [GH17b]. kinematic [CNQ^{+}19]. kinematics [OSP17, PE15]. Kinetic [ESGS17, KTK18, KTK19, ASB^{+}15, AS15, BCC^{+}18, BBBG15, CX15, CXL16, CCDL19, CCGH17, DDJ17, DLN15, DP19, DAO17, FX18, FG17, FLV18, GHM15, GJL19, GMS16, Har18, Hiv18, Ido16, JL18a, JLQX15, JPSX18, JZSX18, JS17, JHT^{+}18, KM16a, LLFX18, LXSC16, LYZ19, LWX19, LP17a, LP17b, MAM16, PX15, PXLL16, PX16, PLWJ16, RXSG15, RXS16, RKH15, SWJG19, SWZW19, SJX15, SJXL15, SST^{+}15, SJX17, TZGW18, TKP16, Vel19, WY16, WYLX17, WCL15, XCX17, YSW15, YSWS16, Yas17, YZZ15, ZAK15, ZLFW18, ZCSZ19, ZXL17, ZLGS18, ZZX16, ZZX19, vdKK16]. kinetic-energy [Vel19]. kinetic-fluid [JS17, MAM16, ZAK15]. kinetically [FGLB16]. kinetics [DEZ16, LYB18, MTK^{+}16]. KIOPS [GRT21, GRT18]. Klein [AMP16, BZ16a, BZ19b, LW17d, LIW18]. KMC [GKRB17]. Knights [SGL18]. Knudsen [CLL19, KJ17b, LLFX18, SWLW19, YSYW19, ZZX19]. Kohn [BEJ15, BHL15, CDM^{+}16, HXX18, ZLH^{+}17]. Kolmogorov [FLT17]. Korteweg [CDN17, LY16b, TXKvdV15, TXKvdV16]. Kriging [CJC19, MWZ19, KSV^{+}15, SGC^{+}17, MS16b]. Kriging-enhanced [MWZ19]. Kriging-sparse [CJC19]. Kronecker [CLZ18]. Krylov [GRT21, ABDN19, AAB^{+}16, AdSŚ^{+}15, AB17, BSK15, GMP15, GRT18, GWC17, JZ16, LCCZ19, PSP16, TWM18, WNDB19, YKMM19]. Kuramoto [CCP19]. Kuroshio [YR15]. Kutta [BR16, O’S15b, O’S19, BR15b, BZ18, CCRdL17, CB15, CCGH19, HK18a, HS18b, JH17, LDSM19, MVK16, MW17a, MH18b, NMC15, PP17, SV19, SLL17, VLV19, Ver19, WJD16, WBM15a, ZT17, ZYD19b].

L [EH15, XS15]. labeled [ZZKP19]. laden [AMB17, BKG15, JL19, LWTF19, PST19, SRS19, ST18c, TUJ19, TASA19]. lag [HHRA19, Pan20]. Lagrange [BLL20, BLL19, Bra16b, BMCK15, CPdS19, CGK17, CGS15, DDJ18, FG16, ID17, LLB19]. Lagrange-Projection [CGK17]. Lagrange-remap [DDJ18]. Lagrangian [YLZ^{+}19, AGBL15, AB16a, AZ19a, BMR^{+}16, BKS18, BDM17, BLK19, BS15b, BLD15, BDZ15, BD17, BDLM18, BB19, BKKJ17, CQQ16, CGQ18, Cap18, CBM19, CM18c, Cot18, DL15, DB16a, DLR15, DAO17, Ein19, FFM19, FL18, FBG15, FFJT16, FLW16, GBM16, Ger17, HAH16, ISST18, KHC^{+}16, KSSL18, KYPK15, LS16c, LDT19, LSTkM15, LMB18, LMB19, LCF16, LWTF19, LHQ19, MC18, MWB^{+}15a, MWB^{+}15b, MLB18, NJ15, OMLdL16, OD17, PP18a, PLB18, PBBK15, PKK18, PZF16, PVB17, Ram17, SRBÓ17, SWC18, SPB17, SW18a, SFT16, SH19, SLdTV18, SK19c, TL17, VW16, VSM16a, VSM16b, WSN^{+}15, WRL18, ZWL^{+}19, ZYD^{+}19a, ZYD19b]. Lagrangian-based [SRBÓ17]. Lagrangian-type [BDZ15]. Laguerre [Ter18]. laminar [PBCR19]. laminar-turbulent [PBCR19]. laminate [ZWYW18]. Lanczos [ZWUR16]. Landau [BHZ16, EMSS17, FJLC18, GS15b, HYK^{+}16, KL17b, LZ15a, SKP^{+}15, SSL^{+}16b, Tav15, WH16a, ZYW16, ZG17]. Landau-de [BHZ16]. Landing [SGL18]. landslides [dlAC17]. Langevin [ALA16, MLL19, MGT18, TR17, VS17]. Laplace [ABN15, BSK15, HPV16, LYPP17, PAFT19, SLR^{+}16]. Laplacian [BGM16, CP16, DvWZ18, FSWW17, GT19, RM16, RRD19]. Laplacians [BDB18, SYM15]. Large [CLB^{+}16, DKPC15, DL18a, FNP17, AQ19, AM19, AG18, BQCG17, BR15a, BPM18, BBBG15, BJ16, Cac15a, Cac15b, CGSS18, CZL^{+}15, CGC17, CM18b, CC16c, CS17b, CHE^{+}17, CMW16, DLLV17, EG18a, ELH^{+}16, FB17, Fer17, FG17, FRW16, GHH15, GDFL17, GFL17, GLS15, HXB15, HLTC18, IPSG15, JdR^{+}18, JBM19, KP15a, KJHA19, KDPK15, LHS^{+}18, LLM17, LXL17, LWL18, Liu16, LDB19, MD16, NYNYM15, Nis19b, OM19, OLV16, PLL15b, PSP16, RFGSV15, RS16b, RWG18, RDG17, RO19, SKP^{+}15, SLL19, SMD18a, SSA17, SLdTV18, TRL15, TSR15, VV16, VKE^{+}18, WSY15, WSS^{+}15, WSHT15, WMY16, WDT^{+}19, WY19, WCWY19, WC18, XB18, CL16, CWS18, LLM17, PD17, TABR17]. large-angle [TSR15]. Large-Eddy [FNP17, CLB^{+}16, BR15a, LDB19, MD16, RWG18, SMD18a, CL16, CWS18, LLM17]. Large-scale [DKPC15, AG18, Cac15a, Cac15b, CGSS18, CGC17, CMW16, IPSG15, KDPK15, LHS^{+}18, LLM17, LXL17, LWL18, SKP^{+}15, SSA17, VKE^{+}18, XB18]. Larmor [CCZ18]. laser [ALM15, GCI19, JB15, MSF^{+}19, TSN16, Vai15, WTS^{+}17, YXD^{+}16]. laser-irradiated [MSF^{+}19]. laser-molecule [Vai15]. last [GG15]. latency [AW16]. latent [AZ19b]. lateral [NMJFM19]. latitude [SFT16]. latitude-longitude [SFT16]. Lattice [AS16, CSB15, GBU15, GW16, GSS15b, HK15a, MKV^{+}17, PF16, ZYW16, ZZL19, ZQCT15, ARF18, APT17, AJVH17, BTVB15, BAR15, CLM15, CYWL17, DCBK15, FGL16, FB17, FBL17, FST15, FBJS19, GPS17a, GPS17b, GR15, GBCF15, GBCF16, GN19b, Hag25, HLU15, HY15, HW15c, HHY16, HW16c, HW16b, HW18, JSY15, KGT15, KP15a, KS15b, KS16d, LFDP16, LL16b, Li17, LLSJ19, LC16, LC17b, LWB^{+}16, MWD16, MG17, MK15, MHGM^{+}15, NSL16, Ols15, PMGW16, PGGW18, RKL18, RS15a, RTO15, STW16, Shi17, STG17, SA19, WSY15, WSHT15, WSY16, WSB19, Xie15, XR17, XTYL18, YFKS15, YYY^{+}16, ZCSZ19, ZY17, ZWG17, HWA19, LDWZ15, VMM19, WGME17, YC16]. lattice-Boltzmann [GBCF16, Hag25, ARF18, GBCF15, KGT15, LL16b, WGME17]. lattices [CY19b, FST15, LS19a]. Laurent [For16, GRMK15]. lava [ZBH^{+}18]. law [AP16, CT15, JL18c, LYZ15, LY15b, CT19]. laws [AW18, BD15b, Bal15, BT16, BMRA^{+}15, BK16b, BLD15, CHOR17, CS17a, CH19, Cho15, CGJY19, CTM^{+}16, Del15, DL18a, DL18b, EFT15, FS15, FS17b, FHA17b, FHA18, GNK18a, HLS15, HAH16, IC17, IDSG15, KC18, LMS17, LMBZ15, LHQ19, LSI16, MDVM16, MDHC15, MB15, NMM16, Nor15, PxRS17, SW17a, SL18, SGC^{+}18a, SWZ15, SWLZ15, SW16, TLQ15, TM15a, TM15b, VNA15, WLGD18, ZP16, ZPW18, ZQ16b]. Lax [DDJ18, DL18c, FLW16, Heu17, Heu19, JCWX19, LDOK17, LFT^{+}16, RPL^{+}18]. layer [AKM^{+}19, BZ19a, BG16a, BHFB19, BCRS19, CKK18a, CLL19, Che19, CS18b, CMH15, DGHP17, DKK15, HRJ^{+}16, KM16a, KHP17, LCLY19, NL18a, PM16, PAFT19, SN15, SHLG15, STFK19, ST18b, SD16, SWLW19, SJH^{+}15, TB19, WG16a, ZC19a, dCMR19, DZR18, GPAO^{+}18]. layered [CHCC18, DvB17, Gib18, Hig15, HN17a, HN17b, HN18, LKB15, MSS16, NWB19, RZ17]. Layers [AL19b, BJK17, BBN18, BFNGDNR18, BLC^{+}17, CHCC18, GTL18, KCHW19, MDT16, Pin15, PD16b, SL17, WBM^{+}15b, AL19a, CJH^{+}19, DCCC16]. LB [ZFL^{+}19]. LBFGS [YGJ18]. LBM [STR15]. LDG [LHQ16]. leaky [Tre16]. leapfrog [MM16b]. Learning [LLD19, RC18, TB18, CJK^{+}19, CE18, GSS^{+}19, GHF19, GK19c, HKBR^{+}19, LKK17b, LZSG19, LJT16, OY19, PD16a, PT17b, QLS^{+}19, RPK17b, RK18, RPK19, RDQ19, RKB19, SS18b, SGS^{+}19, TMES19, VMK^{+}19, XTS^{+}16, YM19, ZLC^{+}18, ZZKP19]. Least [CCFC19, CNW17, NW17, NT19, NWFT19, SNK18, BVG^{+}16, BtTBI18, Blo17, CBA17, CZL18, GLZ19, HGW18, JL15, KP15c, LYZ18, LJ16, MNR19, MAP17, PLPR19, SX16, SLdTV18, TMWF18, TMH16, VLN^{+}18, ZTBW19, ZNX15, dTP16]. least-square [GLZ19]. Least-squares [CCFC19, BtTBI18, CBA17, CZL18, HGW18, LYZ18, MAP17, PLPR19, TMWF18, ZNX15]. least-squares/fictitious [HGW18]. left [CVM^{+}19, Mac15]. Legendre [BR17, Bre18, MDVM16, RW19]. Leidenfrost [VALT16]. length [MMW15, OSP17, RRL19]. lengths [CLZZ19]. LES/under [MMPS17]. LES/under-resolved [MMPS17]. Less [SL19c]. Letnikov [MBSS15]. Level [AAL15, AB15, HIN^{+}16, OLD^{+}18, PDN19, SGD18, VALT16, dLGT^{+}17, AASRT17, AGC19, AT18, BT19, CSW^{+}19, CWWZ17, CD17, CG16, CM16b, GLTB18, GHP15, GFO18, GWYS18, HSM19, HKJ17, JLC15, JH17, JGS16, KC18, LSMS17, LDO^{+}19, LYB18, LLY15, LYKW19, LM19b, LVL18, LSYF15, MGBG16, MSS16, MM16c, MW17b, MLMM17, NLK^{+}15, OLB^{+}17, PLHA18, PC19, QWX18, STV18, SCJ^{+}18, SSA17, TAR17, Vab18, Wac15, XSL18, YJM19, YCS^{+}17, ZC18, ZC19a, ZY19, ZED15, ZQCT15, ZHW18, dLKK19, BAVC17, IBML16, LVTR15]. Level-by-level [HIN^{+}16]. Level-set [OLD^{+}18, dLGT^{+}17, AASRT17, BT19, GFO18, HKJ17, MGBG16, OLB^{+}17, PLHA18, SSA17, Wac15, XSL18, YJM19, YCS^{+}17, ZY19]. levels [MBD19, RKN19]. levelset [vdLJLV16]. Leveraging [NJL19]. Lévy [DW19, XZJK19]. Lewy [GSK18]. lid [EN17]. lid-driven [EN17]. lidar [SNB^{+}15]. Lie [YXX^{+}16]. Lifschitz [EMSS17]. Lifshitz [KL17b]. lifting [STR15]. Light [TK16, TM17, CSLL15, SSL^{+}16a, SSVL18, SKF15]. like [BZ19a, BKP16, CGK17, Don15a, Fal17, GSMR18, KD17a, KNS15, LB15, LZZS15, MWB^{+}15a, NN15b, NLL^{+}15, Par15, PT17b, SPP16b, SLN15, TK12, TK15b, WWRS17, ZG19]. likelihood [CLM16, CMW16, NS16]. likelihood-informed [CLM16, CMW16]. limit [BZ16a, BZ19b, CX15, DDD17, Fal16, GZY19, HW15b, KMGR16, LB17, LTWZ18, LZL^{+}17, LP17a, PMF15, SPP^{+}16a, SSM15, WY16]. Limitations [LMTC15, AÁPB17]. limited [AJP15, AP16, BFP18, BL18, KKZ15, LS19b, Par15]. limited-aperture [LS19b]. limited-view [AJP15, Par15]. limiter [DL16, KH17, Kri17, Tso18, ZJLC15, KH17]. limiters [CK16b, GK18, HKA19, MSP19]. limiting [BD17, DKK^{+}18, LY16c, LK16b, Nor15, PK16, YK18, dFVJ15]. limits [MW15]. Line [BMR19, ABG^{+}18b, ABG^{+}19b, BKR19, FB17, FPV18, HB16, HW18, LPGT16, LBZ16, LHA16a, SYY15, SCS16, TW17, TP16a, YY17]. Linear [AD15, MS18b, MSP15, Yan16b, ALKZ16, AS15, Alm19, ATF16, ALMJ15, ADOP18, BP18, BVT18, BHFB19, Cac15a, Cac15b, CHT17, CW17, CCGH17, DJLQ18, DJD^{+}17, FAC^{+}19, FS19b, FKDL17, GH19, GN16, GDA16, HYK^{+}16, HN17a, HSF17, JSB^{+}19, KZR15, KJHA19, LS15a, LL16a, LLS15, LC17a, LWWY18, lLNS17, LYA16, Lor19, LLLN18, MHH19, PK17, PBL^{+}19, Pis18, Poë19, PSP16, QN19, RPK17b, RZ17, RSH^{+}17, SLL19, Sch16a, Sch16b, SMSR18, SCC19, TD18, TDC^{+}19, TWM18, URGT18, VDPP15, WY16, WT19, WSH^{+}17, YHKPF17, ZK15, ZD15a, ZW19, ZCL17, ZH19, ZR17, dSPDH15, vOMB17]. linear-scaling [LL16a]. linear-upwind [ZH19]. linearised [ST18a]. linearization [GMD19, Vos17]. linearized [CT15, CLL19, GS15b, JL16, LGH^{+}18, MMP18, NN19, SP15a, WZL^{+}17]. Linearly [ABIR19, YH17, BNM15, HSF19]. lined [RMLvR18]. Lines [CGJ16, HKS^{+}16, LD15, MG15b, MHGL19, PR16b, SWMD17a, SWMD17b, SSA17]. link [Ols15]. linking [Pan15]. Liouville [FLW19, KADE15, RJ19a, VSM17]. lipid [RAMB15, SDMS17]. Liquid [BG16b, BMT18, BLG^{+}16, BLC^{+}17, CTJ^{+}17, CGS15, CLM15, DD15, FMRZ17, HHZZ19, HW15c, HW16c, KLWQ17, LVB^{+}15, LSYF15, MTZ16, NBMB19, NWZ18, ÖPHA15, OCSC18, PD19, SDMS17, TK12, TK15b, TO19, VALT16, WY19, WCWY19, YG19, ZYSW16]. liquid-gas [PD19]. liquid-gas-particle [CLM15]. liquid-liquid [CTJ^{+}17]. Liquid-vapor [BG16b, DD15, FMRZ17]. List [Mac16, DFS16]. lithofacies [dFGS^{+}17]. lithography [OLD^{+}18]. Liu [GMD19]. Lloyd [YGJ18]. Lloyd-preconditioned [YGJ18]. load [GFA^{+}16, JBLO15, KJ18]. load-balancing [GFA^{+}16]. loading [LSS16]. loads [LC17a]. Lobatto [Kas15, MRRRF18, Teu15]. Local [CG19, DLL^{+}17, HSC16, KLRT15, LDSM19, MSK18, SL19b, TL15, ADK^{+}17, BBF^{+}17, BK19b, BDZ15, CBZ18, CPV16, CEL^{+}18b, DKTH15, DY19a, EDW19, FB15, FC19b, GH19, GSK18, GX15, GY15, HLJ^{+}19, JMM19, KL15, LW15b, DV17, MNG15a, MK17, MTM19, MGCW18, NS19b, OSP17, PR19, RPNP18, RGPS17, SSL^{+}16a, ST18b, TXKvdV15, TXKvdV16, VAD17, WZ18b, YS18a, YTW15, ZLH^{+}17, ZZW^{+}16, dlHC16]. Local-global [TL15, FC19b]. local/global [CBZ18]. Localised [RZOZ19, CK16b]. locality [MK17, NSB15]. Localization [BFP18, MTM19, NS19a]. Localized [DLY17, LL16a, AH15, CLR15, HK19a, WL17]. Locally [LWJV19, APLK19, BFNGDNR18, BHF15, CC17a, DGMT17, FGLW18, JW15c, KHP17, Rag15, SLY16, TABR17, ZG18b, ZCY^{+}19]. locally-cartesian [FGLW18]. locally-heated [KHP17]. located [Kla15]. location [CCWY18, Fan19, PKLS17]. locking [LR19b]. Loeve [LC18, CN16]. LOI [SNK18, SNK18]. Long [FRW16, LMC19, BPGS16, BLK19, CLMZ17, GZY16, GBS15, GQC^{+}19, JTD16, LXL17, LIW18, OB17, SG19a, XL17b]. long-range [LXL17, SG19a].

Long-term [FRW16, GZY16]. long-time [BLK19, GQC^{+}19, LIW18, OB17]. longitude [SFT16]. Loop [PRXC19, HWK19, PCX17, Shi19, PXXZ15]. loosely [Buk16]. loosely-coupled [Buk16]. Lorentz [JSB^{+}19, SPCH16, YXD^{+}16]. Lorenz [GHJ15]. loss [GKRB17]. lossy [KHTZ19]. Low [BKG15, CB15, GDFL17, KQB18, RNO19, STG17, AAI16, AMJ17, AAD16, BDKK17, BKKY19, BH16b, BLMY17, Bon17, BDJP19, CL18, CPV16, CS16a, CVG18, CDV17, DWR18, DCP15, DLMDV18, Dom18, Eva18, FDKI17, Fal16, FG18, FYO^{+}15, HKLZ18, HR19, HK18a, HLS15, HWA15, JLC18, KLC18, KS16b, KP15c, KYW^{+}16, KYW^{+}18, KV16, LTB16a, Lau17, LHO^{+}19, LSWF16, LT17a, LO16, MM16a, Mar19, MA19, MVZ16, MBD15, MA16, MDAB18, NMC15, OLHD17, RC18, SP15a, SK18, WAZ19, ZHA17a, ZWG17, ZWB^{+}18]. Low- [STG17, CS16a]. low-density [LTB16a]. low-dimensional [BH16b]. low-dispersion [HK18a, JLC18, NMC15]. low-dissipation [HK18a, HWA15, JLC18, KV16, LHO^{+}19, NMC15, ZHA17a]. low-energy [HLS15]. low-fidelity [AAI16, LSWF16]. low-Mach [CPV16, Dom18, LT17a, MDAB18, SP15a]. low-Mach-number [MA16]. low-memory [MVZ16]. low-order [AMJ17, CCK^{+}17, OLHD17]. Low-rank [RNO19, AAD16, BDKK17, BKKY19, CL18, FDKI17, FG18, HKLZ18, KS16b, Lau17, LO16]. Low-resolution [KQB18]. Low-Shapiro [GDFL17]. Low-storage [CB15]. lower [AEL^{+}15a, AEL^{+}15b, HLQ16, ZC19b]. lower-dimensional [AEL^{+}15a, AEL^{+}15b]. lower-order [HLQ16]. lowest [BCS19, LTW18]. lowest-order [BCS19, LTW18]. LPSE} [MSF^{+}19]. LU [LL16b]. LU-SGS [LL16b]. Lubricated [FR18]. Luenberger [CCM15]. lumped [BK16b, SM16]. lumping [Sir19]. lunar [HWH^{+}16]. lung [MCHL16]. LUPOD [RTV17]. Lyapunov [BVM^{+}17a, FW17, HR19, MSB^{+}16].

M [EH15, Pan20, XS15, YLLH19]. M.H [Pan20]. MAC [ZZKF15]. MacCormack [ZB15]. Mach [ABIR19, BLMY17, Bon17, BQRX19, BDJP19, BKG15, CPV16, DLMDV18, Dom18, FP18, HR19, LT17a, MM16a, MA19, MDP^{+}15, MBD15, MBD19, MA16, MDAB18, SP15a, SK18, TD17, WDGW17, XDSX17]. Mach-number [Bon17]. Machado [Kat16]. Machine [LJT16, RPK17b, RK18, RDQ19, CE18, GK19c, PD16a, QLS^{+}19, TMES19, VMK^{+}19]. macro [GJL19]. macromolecular [AAB^{+}16, LJZ15]. macromolecules [Han19, XL17b, ZRT18]. Macroscopic [Hwa16, KK17a, STR15, VS17, ZLC^{+}18]. magnetic [BBKS16, BPL19, BJK17, CLMZ17, DWG^{+}18, ESGS17, EBQ15, Guo15, GFW16, LBZ16, PMS15, RBGV15, RSD17, DD17c, SAF^{+}19, TBC^{+}16, TPTT18, TBB^{+}19, Web14, XL16, ZJS15]. magnetized [CKK18b, Iwa15, KKS15, KKS16, KS18a, LLD^{+}16, dSPDH15]. magneto [Guo15, GFW16, SK19b]. magneto-hydrodynamics [Guo15, GFW16, SK19b]. magnetohydrodynamic [Bal15, BMT16, CFST16, HdBH^{+}16, HHA16, JH15, MJ16, MAH16, SR18, XL16, YWS^{+}16]. magnetohydrodynamics [Ama18, BK16a, BND16, DWG^{+}18, DLK17, ETL17, HL16a, Iwa15, KW15b, KTG16, Moc17, SSO^{+}15, Sov16, SE15, SS17c, TPB16, ZN18, ZT17, ZYD19b]. magnetoquasistatic [NGS16]. magnetospheric [LML^{+}16]. magnets [FBC^{+}16]. Maintaining [BJ15, NF17]. MAN [ZR17]. management [MC16]. Manifold [HKBR^{+}19, XTS^{+}16, BTD16, BM16, GH19, GS18, GA18, KP15b, LBTCG16, SG16, ZWB^{+}18]. manifold-based [BM16]. manifolds [GSS^{+}19, KR17, LTR16, LYPP17, MMNI16, SG17, SGS^{+}19]. manipulated [BLL16]. manner [GK19a]. Manning [MDBCF17]. manufactured [VBG16]. manufacturing [ADE^{+}17, Zoh17]. many [BH18, HZE19, LSP^{+}18, LLVF^{+}15, RMBN18, SD15, TRM16, VYP15]. many-body [LSP^{+}18, SD15, TRM16]. many-dimensional [BH18]. many-electron [HZE19, LLVF^{+}15]. many-material [RMBN18]. map [BT19, KR17, KL19]. mapped [MDHC15, ZG19]. mapping [BCST17, CLFL17, iI15, MC17, PvL19, SZN19, SZN20, STFK19, SPRW15]. mappings [Pas16, PvL19]. Marangoni [Str17]. marching [FLV15, NLL^{+}15, PLHA18, SL19b, TH16, YS17]. markers [AB16a, FL18, ISST18]. Markov [Lan19, MWD16, XZZ15, YZZ19]. Mass [KG15, LHL15, SNB^{+}15, WWR16, BHKS16, CMDL18, FGL16, FGLW18, FB15, GLTB18, HSK^{+}15, HLL^{+}16, HG17, HDA^{+}18, LY15a, LSYF15, MTK15, MRXI17, OD17, PG17, SBHS19, SM16, Sir19, TCSM15, TCS16a, Teu15, WSS^{+}15, WY19, WCWY19, WB17, WG19, XL17b, Zad11]. mass-conservative [FGLW18, HSK^{+}15, Zad11]. mass-conserved [WSS^{+}15]. mass-conserving [FGL16]. Mass-corrections [WWR16]. mass-lumped [SM16]. mass-preserving [GLTB18]. mass-redistributed [HLL^{+}16]. Massively [TPTT18, vdKK16, DG16b, HM19a, HM19b, NN15b, PJE^{+}16, YS17]. massively-parallel [HM19a, HM19b]. master [GMS16, IZ18, MFB18]. Matched [AL19a, AL19b, CJH^{+}19, DZR18, DCCC16, BJK17, Che19, CMH15, DKK15, GTL18, PD16b, SJH^{+}15, dCMR19, GZ17]. matching [FKS19, LLJJ18, YZT^{+}18]. Material [DZ16, RHS18, APKP16, ABH^{+}19, BTGM17, BGTM18, Buk16, BHJ15, DZ18, GO15, HTZG17, KKLS17, LGZ^{+}19, MZAF17, MCL19, NDH19, PHHA18, PA19, PR16b, RMBN18, SC18a, VSM16a, VSM16b, Wil19, YG18, ZZS^{+}17, ZD17, ZYD^{+}19a, ZSO^{+}19]. materials [AIP17, BM16, CSY15, GHP15, HMBH15, KZ17, LK17, MGKG17, PS15b, PSN^{+}19, SNSG16, SU15, SPN^{+}19, TAJ^{+}17, VK15, YT17, KYKS19]. Mathematical [NLFM16, QS16, Lap16]. matrices [ABDN19, AAD16, CC19b, GRMK15, RO19, SYM15, YD19]. Matrix [BMMP19, AEL^{+}17, BDBEE15, BHMS18, CDC17, CWB^{+}19, CC19b, CLP16a, DLY17, DH18b, EE16, For16, GFvR18, LL16a, LWLC17, LM15d, NMA15, Noe15, SWZ17, Teu15, VYP15, WDGW17, WLK^{+}16, XLY15, XL17b, ZLL17a]. matrix-exponential [SWZ17]. Matrix-free [BMMP19, XLY15]. matrix-valued [LM15d]. matter [WTS^{+}17]. maximizing [ES18, ZC18].

Maximum [BC16b, DY19a, SWPS17, ADK^{+}17, CHY16, CLTX15, GP16b, LSS16, MN16c, WYZZ18, YL19]. Maximum-principle-preserving [DY19a, CLTX15]. Maximum-principle-satisfying [SWPS17, CHY16, YL19]. Maxwell [QHZ^{+}15, ABH18, ABH^{+}19, BK19a, BV15, BCB15, BCJL17, CW16, CCZC16, CHZ16, CQL^{+}17, Chu17, CEF15, DDD17, Del15, DGL^{+}15, ETAG15, Fal16, GSN16, GHV19, HKLZ18, HCB19, HJZC17, HHY15, HLCL19, Ism15, JSB^{+}19, KHTZ19, LHLL19, LHS^{+}19, LY19, MM16b, NBT19, PT17a, SP18, SZ15a, SOS19, SL16b, SC18b, SL16c, WR16, YJ17]. MBAR [XR17]. MBAR-enhanced [XR17]. MBO [JME18]. MC-IP [XWZ^{+}18]. MCMC [AÁPB17, BGL^{+}17, CLM16, HYL17, MTM19]. MD [WPB15]. MD-DSMC [WPB15]. MDF [Ger17]. mean [BDPM18, CRMP16, LC17a, SAOW17, TS19]. mean-field [TS19]. Meaningful [Cos16]. means [WNW^{+}19, ZBZ^{+}18]. measurements [EST17, SNB^{+}15]. measures [Opp17]. mechanical [GDFL17, KBG^{+}15, KGP^{+}17, LMC16, PD16b, YG18]. mechanically [ZSX17]. mechanics [BT17b, CGC17, DPRZ16, DPRZ17, FRL15, FFJ^{+}19, FFJT16, Jac17a, KGP^{+}17, MSH^{+}15, NDH19, NRZS17, Sel15, VCEK19, YT17]. mechanics-based [KGP^{+}17]. mechanisms [WTS^{+}17]. mechano [FRW16]. mechano-chemical [FRW16]. mechanochemical [DMRB19]. media [ABI17, AEVW18, An17, APKP16, Azi19, BTGM17, BGTM18, BDMC15, BPS17, BC18c, BCJL17, BSWG15, BHMS18, BKKRB16, CHCC18, Cho19, CXY19, CLQ17, CEL^{+}18b, CS17b, CJH^{+}19, CM18d, CLNH15, CvKH16, CFvKH18, DSS18, FQZNZ18, FPT17, FC19b, FYC^{+}18, GFG^{+}15, GH17a, Gib18, GAS^{+}18, HSK^{+}15, HN17b, HLCL19, JT18, JSB^{+}19, KJ17a, KYKS19, KLRT15, LW18, LP16a, LH15, LT15, LJ19, LZT17, LZL^{+}19, LFAR19, LNM15, MCN18, MP15a, MVZ16, MTD15, ML16, OLV16, PA19, PF16, RZZ19, RdM19, RJLW19, SSL17, SPX^{+}18, SWML17, SMT^{+}16, SMG19, SC18b, SiI16, TWH15, TAH16, UHKT19, VSDW18, VS17, Vos17, WC18, XOX19, XML17, YJ17, YGEM17, YSY17, YFC19, YB17, Zad11, ZZ17b, ZOY^{+}19, ZWUR16, dMRHJ17]. Medium [MSG18a, MSG18b, BNM15, BKL17, CGMH18, DvB17, DLWY19, GCVCHH18, Han19, HM17, Iwa15, KHTZ19, LTKA15, LH17b, LRGO18, LMC19, MSS16, NH17, QM18, TT19, ZF18]. MEEVC [PG17, dDPG19]. melt [CY19a, RTO15, WNDB19]. melts [SB18]. membrane [CJYZ15, GGT18, MTK15, TFGK18, XR17, YM17b]. membranes [LAA16, MTK17, RG15, RAMB15, SMA^{+}16, SMOM^{+}17]. Memory [SZ17, AMK17, DOO17, FYO^{+}15, MBSS15, MVZ16, TP17, WLC15, XOX19]. memory-efficient [DOO17]. Memory-optimized [SZ17]. Mercer [AABD15]. merging [KK17b]. Mesh [BV15, GPAO^{+}18, LS15c, PSB^{+}18, PWC18a, SPD19, SL18, SW18a, Sla16, WBBC16, APP^{+}16, AB16a, AMS17, ALO18, BHZ16, BOA17, BAD19, BGRC19, BHS^{+}18, BSM16, BD16, CGL18, CTJ^{+}17, CWW17, CHJT17, CGJY19, CBM19, DRP^{+}16, DvW15b, DHH^{+}18, DMS17, FHY^{+}19, Fid17, FBG15, FP16, FNNB19, GBR15, GBvZB16, GK19b, GSN17, HS17b, HS18a, HIN^{+}16, HLL^{+}16, HDA^{+}18, HM19a, HK15b, HW16c, IZ18, JW15c, KF15, KVKS19, KAR17, KLRT15, KJ18, KS17, LSLA16, LS16b, LMB19, LNM15, Loz17, LHQ19, MLM18, MKSP19, MCW16, MW17b, MLB18, MSB^{+}16, NH17, NSB15, OKWE17, OD15, PR19, QYJ19, RBJS15, RPNP18, RRD19, SRBÓ17, Say17a, Say17b, SKS17, SW15, SLdTV18, SFP16, Sub15, SJX17, SZS15, TVB^{+}16, WQZ15, WDS15, WKOE17, Wil18, WCT18, WHZ18, XL17a, YHQ15, YGJ18, ZL15b, ZZ17b, ZJ18]. mesh [ZHLZ18, ZL15c, dlAC17, GMWC19, Mas18]. mesh-decoupled [OD15]. Mesh-free [SW18a, Sla16, RRD19]. mesh-induced [HDA^{+}18]. mesh-to-mesh [WKOE17]. meshes [APP^{+}16, AAE17, ATF16, AHZ19, AM17b, ADOP18, BCST17, BD15a, BD15b, Bal15, Bar18, BKO18, BT16, BDZ15, BD17, BDLM18, BB19, BD18, Bou19, BRW15, BMC18a, CFSN18, CGK17, CWM^{+}16, CW19, CSY19, CHY16, CSN17, CLTX15, CXY19, CCM17, DK19, DSH^{+}16, DC18a, DvW15b, DY19a, DL16, DMTB15, EDvW17, FLHA17, GM19, GAIN19, GK18, HHR^{+}19, Her16, HR17, Ism15, IM15, JBLO15, KKLS17, KDPK15, LLD^{+}16, LSLA16, LMG15, LLP^{+}16, LYC16, LL16b, LLJJ18, LYKW19, LM19a, LM19b, LJ16, LSZ18, LTW18, LMN18, MSD^{+}17, MMvR18, MKB19, MHS16, MWB^{+}15b, MW17b, MDD^{+}19, MM17, PX16, PM16, PR17a, PL16b, DDM18, QDH15, RBI18, Rag15, RGW16, SR19, SAEF17, SL17, SWMD17a, SWMD17b, SWLZ15, SP19b, SYM17, SSX16, SDW18, TLQ15, TD16a, TD17, TD18, TC15b, TLR16, TLB^{+}18, Tso18]. meshes [VST16, WWR16, WHY17, WHY18, WWGK17, WWGW18, XP15, XYG19, YSYW19, YK18, YK19, ZOG19, ZZZ17, ZLFW18, ZCQ19, ZCQ20, ZPW18, ZYD^{+}19a, ZQ17, ZS19a, ZXDL17]. meshfree [AS17, SMLB15, SK19c, ZZW^{+}16, FPT17]. meshing [MMSS15]. Meshless [BK19b, IKI15, BDB18, CLR15, DA17, LMZ19, MBS19, SLA^{+}19, TSH17, TMH16, TMH18, XYPT16, YHKPF17, YTW15, ZNS19, ZMCC18]. mesoscale [CFPB17, DOO17, MGPG19, SGC^{+}17]. mesoscopic [FHE15, KLC19, LYLK17, ZLC^{+}18]. meta [KS16b]. meta-models [KS16b]. metafilms [DKTH15]. metal [CLFL17, YZW^{+}18, ZWYW18]. metal-friendly [YZW^{+}18]. metal-intermetallic [ZWYW18]. metallic [LZ19, SSL^{+}16a, VCNOP18]. metals [SSVL18]. metamaterials [Fuj19]. Metamodel [RC18]. metamodeling [SDJU15]. meteorological [LPR19]. Method [ACGR15, BQCG17, CE17, Chu17, GFG^{+}15, LFR17, MC15, PMF^{+}18, RKO^{+}17b, SMAG17, WZ18a, ABI17, AM17a, ARG^{+}17, AMN18, AASRT17, ABG^{+}15, AR16a, APR^{+}15, AGC19, ACCCDA16, ACCCD^{+}17, ALKZ16, ASB^{+}15, AB16a, AMS17, AB18, AZ19a, ASS13, ASS17, Alm19, AP16, AGRB18, ADK^{+}17, ACS16, ACJ17, Ani19, AL19a, AT18, AB15, ABT19, ABdC^{+}18, AÁPB17, AB17, ALTR17, ANL^{+}16, AJVH17, AG18, AEvW19, BCSK17, BJRF18, BC18a, BK17a, BM15, BFI^{+}16, BZ16a, BXY17, BDG^{+}17, BJWZ17, BRK^{+}18, BV15, BLA^{+}15, Bar19, BBF^{+}17, Bat17, BST^{+}18, BT19, BtTBI18, BBB^{+}16, BNK18, BC16a, BZ15, BC16b, BMT18, BGD19, BBKS18, BS15b, BVS18, BAR15, BGG16, BFT17, BTA17, BKKJ17, BHTT17, BPD19, BLC^{+}17, Bou19, BPM18, BHF15, BHFB19, BTWY15, BC16d, BFTVC18, CQQ16, Cai16, CZW17, CGQ18, CFKK19, CDM^{+}16]. method [CCHL15, CL18, CSW^{+}19, Cap18, CARdN19, CPdS19, CFO18, CHT17, CG19, CDM18, CGMH18, CTJ^{+}17, CW17, CJD^{+}17, CDDL19, CXH15, CCZC16, CXL16, CXX16, CZ16, CX16, CH17, CZ17, CL17, CWWZ17, Che18, CWB^{+}19, CT19, CSC19, CH19, CC19b, CLL19, CZ19b, CYL^{+}16, CYYL18, CZL18, CL19a, COV18, CSG17, CDN17, CLR15, CMDL18, Cho19, Cho15, CLL17, CFST16, CGJY19, CBN^{+}16, CEHM19, CLQ17, CELZ18, CEL18a, CPP19, Cif19, CVG18, CFF18, CPS17, CSK^{+}16, CCL16, CSH15, CLM15, CBM19, CV16b, CM18d, CM16b, CLP16b, CLMZ17, CYWL17, CC19c, CLNH15, CvKH16, CFvKH18, DM17a, DD17a, DMAM15, De18, DM16, DCA^{+}16, DJV^{+}18, DLM18, DC18a, DKPC15, Del15, DDV18, DGMT17, DG16b, DZ16, DZ18, DS16, Did17, DLR15, DLNR18, DF16, DH18a, DMS17, Dod17, DAO17, DVP^{+}16, DWW15, DLL^{+}17, DY19a]. method [DL16, DvWZ18, EDC16, ESHA16, EL17, EKSS15, EKEB16, EDW19, ELH^{+}16, FR18, FGL16, FBL17, Fal16, FS16, Fan19, FHS17, FMRZ17, FG16, FS18, FCL17, FBF15, FNP17, FGLB16, FBM16, FB15, FITcD19, FNNB19, FLV18, FZ19, FG19, FLHA17, FHA17a, FLW19, FC19b, FSM16, FP18, GSN16, GM19, GB15b, GP17, GHM15, Gam15, GZY16, GH17a, GFC18, GP16a, GLTB18, GPS17a, GPS17b, GSL18, GHP15, GWC17, GMWC19, GG15, GAIN19, GBC^{+}18, GBD17, GNK18a, GNK18b, GN16, GD19, GWWC17, GCVCHH18, GEZK16, GHJ15, GZ18, GWYS18, GQC^{+}19, GN19b, GTG15, GY15, GFW16, GHH^{+}16, GY18, GYZ19, GSL^{+}19, GP16c, GL17, HPY18, HK19a, HL16a, HPC19, HHA15, HWH^{+}16, Han19, HXLL15, HS17a, HHR15, HB16, HZL^{+}15, HLL^{+}16, HG17, HGW18, HKJ17, HW16a, HP17, HM16a, HR17, HHCG15, HMBH15, HL16b]. method [HGN17a, HGN17b, HJZC17, HN17a, HN17b, HN18, HMFJ18, HM19a, HM19b, HSC16, HCVH18, HHRA19, HF18, HHLY17, HZ17, HCLT19, HHL19, HLU15, HLY15, HJ16, HLSY16, HXX18, HM19c, HC17, HY15, HHY15, HZ15, HSSZ16, HHY16, HW16c, HLL^{+}18, HJW19, HMRG16, Hwa16, iI15, iI17, IKI15, IML15, IM17a, ION^{+}17, JAH19, JL17a, JBM19, JKE^{+}17, JSS15, JPLL15, JW16, JLC15, JST17, JT18, JL18b, JL19, JLLZ15, JL17c, JLKF17, JCWX19, JGS16, JJ19, JS19, JTD16, JJ17, KKH18, KPKGH19, KTN15, KKS16, KNS15, KC17b, KLSF15, KJ17a, KH15, KP15a, KKJB16, KJYC17, KL17b, KPJ18, KLC18, uKHGK19, KDL15, KR17, KO17, KLNH17, KCS^{+}17, KJHA19, KP15c, KK17b, KLGO18, KS16d, KLWQ17, KM15, LTB16a, LS15a, LMS17, LLD^{+}16, LY15a, LM15b, LFRH17, LMG15, LDO^{+}19, LML^{+}16, LC15, LAL18, LM18, LPW15, LH15, LLS15]. method [LFDP16, LBZA16, LW17c, Li17, LLJJ18, LBTK18, LGH^{+}18, LYZ18, LXC19, LLSJ19, LCLY19, LJ19, LMBZ15, LXL17, LC17a, Lia16, LZT17, LL19b, lLLNS16, lLNS17, LSD^{+}17, LHY^{+}19, LZ19, LLW19, LYKW19, LMMS16, LM19b, LTB16b, LC16, LC17b, LD15, LDWZ15, LZT^{+}15, LY16c, LY16b, LWB^{+}16, LK16a, LW17e, LTXB17, LWY18, LTW18, LMZ19, LZS^{+}19, LMB19, LLH19, LM19c, LM19d, LYPP17, LHGF16, LMG19, LQB16, LT17c, LZW^{+}17, LHA15b, LSYF15, LP17a, LWC17, LRGO18, LWTF19, LSCC19, LHQ19, LMGG17, MMNI16, MD17, MD18, Mac16, MC18, MAK15, MZAF17, MS18a, MS18b, MIM^{+}19, MDVM16, MS16b, MMB18, MG17, MK15, MDL16, MA19, MA17, MO18a, MST15, MKS18, MMvR18, MPFL16, MRZG16, MBS19, MHGM^{+}15, MBST17, MTK15, MTK17, MBD15, MCS16, Moh15, Moo17, MF16b, MLL18, MTK^{+}16, MDD^{+}19, MLMM17, MDAB18, ML16, MWYZ16, MN16c, MSA19, MH17, MM16d]. method [MM18, MPMB19, NPB19, NVBDV15, NBZ^{+}19, NWKC16, NPRC15, NJHL18, NS19a, NS19b, NLK^{+}15, NBMB19, NN18, Nis18a, Nis18b, Noe15, NSL16, NLW^{+}16, NRS19, OB19, OT15, Ols15, OM19, ÖPHA15, OD15, PZNG15, PKLC16, PKLC17, PHHR17, PDdG^{+}17, PNZ18, PHHA18, PGCG18, PLC18, PL16a, PL18, PCF15, PSS17, Pan20, PSB^{+}18, PK17, PJC16, PPLC16, Par18b, PWC18a, PKB15, PN18, PR16a, PR16b, PLL15b, PGGW18, PF16, PS15b, PR16c, PLR18, PLPR19, PTT18, PBKK17, PMB18, PG18, PZF16, PCA19, PSP16, QWZ^{+}19a, QWX18, QYJ19, RTG18, RBI18, RVZB15, RRS19b, RRS19a, RBJS15, RW19, RG15, RS16b, RWG18, RW15b, RMLvR18, RZ17, RKRGW17, RXSG15, RXS16, RLGT19, RSBS19, RMBN18, RAMB15, RTO15, RMC15, Rua18, RPC^{+}18, RSB15, SY17, SPX^{+}18, SZM19a, SZM19b, SWC18, SLL19, SYI^{+}19, SXBB15]. method [SWS17, SPD^{+}17, SPB17, SSL^{+}16a, SSVL18, SGMS16, SHA16, SWG^{+}17, SKF15, SKF16, SBG^{+}17, SHKL16, STKL19, SAK18, SF18a, SMT^{+}16, SRBB18, ST17, Sha17b, SPB18, SWMD17a, SWMD17b, SMP16, SWZ15, SW16, SwS16, STV18, SL15, SW15, SZW^{+}16, SSM^{+}17, Shi17, SL16a, ST18b, SP16b, SLY16, SMA^{+}16, SCS16, SYM17, SLVE18, SW18b, SA19, SGD18, SO17, SMD18b, SPP16b, SMLB15, SA15, SBH19, SSA17, SZ15b, SDW16, SZ17, SMSR18, ST18c, STV19, SS16c, SLA^{+}19, SMOM^{+}17, SHW18, SGT16, SGT17, SF16, SHP^{+}16, SCLG15, Stü17, SL16b, SC18b, SPCH16, SWZ17, SWZW19, Sub15, SLZ^{+}17, SGP17b, SCS18, Sun19b, TWN15, TCD17, TCB18, TH18, TW17, TMWF18, TD16a, TD17, TSH17, TAH16, TV19, TST^{+}15, TXKvdV15, TXKvdV16, TLLF15, TMS^{+}19, TRL15, TOR^{+}15, TT16, Tow18, TO15, TMH18, TLB^{+}18]. method [TKF17, URGT18, UG16, VBG16, VCNGP15, VCNOP18, VBF15, Vog17, VK16, Vre17, Vre21, VMC^{+}19, Wac15, WG16a, WY17, WW15, WZ15, WSS^{+}15, WDS15, WE15, WXW15, WSJY16, WRL16a, WRL16b, WRPL17, WLWW17, WW17, WCH^{+}17, WMY18, WC19, WWX19, WZLS19, WDT^{+}19, WLE17, WMYG16, WSN^{+}15, WPB15, WNW^{+}19, WSOW16, WXSJ19, WMS18, WG19, WHE17, WHEK18, WR16, WWGK17, WWGW18, WSB19, WCCB16, WKHT19, WZRZ15, WH16b, WL16, WYA^{+}17a, WYA^{+}17b, WTX17, WA18, WHZ18, Wu19, WAZ19, XJG18, XLY15, Xia15, XZ15, XX16, XDSX17, XX17, XXR18, XOX19, XDLX19, XY18, XY17, XP15, XHC15, XL17b, XWW17, XSL18, XTYL18, XM18, YXW19, YYY^{+}16, YSC^{+}17, YC15, Yan16a, YZZ19, YCPD15, YHQ15, YS15, YXF^{+}16, YY16, YYN^{+}17, YFJ17, YS17, YZW17, YHKPF17, YZT^{+}18, YGJ18, YSYW19, YBSTT19, YG19, YM17b, YJB18, YXX^{+}16, YX15, YM15, YK19, YTW15]. method [YB17, YZZ15, YCS^{+}17, YL17, ZP16, ZCHS15, ZS16, ZTBW19, ZND16, ZFZL15, ZS15, ZLY15, ZL15a, ZB15, ZL15b, ZD15b, ZD15a, ZLJ16, ZLL16a, ZLL16b, ZLL17a, ZZS^{+}17, ZZZ17, ZZ17b, ZD17, ZHA17a, ZJ18, ZC18, ZG18a, ZMCC18, ZF18, ZLFW18, ZWL^{+}19, ZC19a, ZRW19, ZCL19, ZY19, ZZL19, ZZYC19, ZXW^{+}19, ZVO15, ZBZT17, ZY17, ZSX17, ZYD^{+}19a, ZZ19, ZYCK15, ZGD^{+}16, ZZW^{+}16, ZCL17, ZSO^{+}19, ZRT18, ZL15c, ZLX17, ZHW18, ZK18, ZG19, ZYD19b, aKT16, dLDG^{+}18, dLKK19, dTP16, dlHC16, vdLJLV16, ADN19, BDV17, CGJ16, FPT17, GBS15, GAS^{+}18, GLTG15, HWK19, LVTR15, MNR19, Mue18, RHS18, RMF^{+}18, SRS19, TSFS17, TBLJ15, VALT16, dPSS16]. method-of-lines [SWMD17a, SWMD17b]. methodology [Cac15a, Cac15b, DLK17, FKDL17, KYUO15, LSMS17, MNG15a, MJ16, MP19, MN18b, PTT19, PBA^{+}15, RLV16, RDM15]. Methods [BMR19, FFW17, JHPAT17, ABDN19, AAG16, AGKD19, AW18, And16, ADH^{+}16, ALT17, AC16, ALL18, ARTG^{+}19, AÁPB17, ALA16, BH16a, BZ19b, BMR^{+}16, BGN19, BHdD18, BCS19, BGRC19, BDM17, BHP19, BAVC17, BVT18, BGGM15, BCJL17, BZ18, BGHK19, BK16b, BD18, BRW15, CLW18, CC15, Cen19, CDL17, CGL18, CWM^{+}16, Cha18, CE18, CW19, CZ19a, CCZ18, CQ15, CHZ16, CHY16, CS17a, CWW17, CKQT15, CS18a, CCKQ15, CS18b, CRMP16, CLG^{+}19, CVK16, CLX15, CXY19, CEL15, CR18, CK16a, CL19b, Cot18, CWJ18, CCGH19, CHLZ17, DD17b, DDD17, DPW^{+}15, DZ16, Die15, DB16a, DP19, DWZ19, DM17b, DM19, DKK^{+}18, DGL^{+}15, DMM19, DKC15, DZC16, DLWY19, DJD^{+}17, EHXM15, EFT15, EG18b, EARA15, FH17, FWK17, FWK18, FSWW17, FGLW18, FHE15, FPDT17, FRRV16, FL16, FS17b, FAC^{+}19, FCL19, FS19b]. methods [GK19a, GMP15, GO15, Ger17, GFO18, GO16, GFvR18, GLMC16, GA18, GH17b, GY17, GXX17, HKA19, HGR16, HKLW15, HK19b, Heu17, HF18, Hu17, HXB15, HB15a, HB15b, HS18b, HDF18, IC17, JZSX18, JW15b, JL15, JX15, JZ16, JX17, JL17b, JYY18, JL18c, JWH16, JXZ15, JBLO15, JFS17, JSY15, KTK15, KS16a, KDF15, KMS^{+}18, KPKG15, KA15, KADE17, KE15, KK17a, Kla15, KRFV16, KG15, KWHB19, KS15b, LSMS17, LBTCG16, LH17a, Lau17, LR19a, LSL15, LPWK15, LE16, LW17a, LW18, LYC16, LW17b, LPR18, LX18, LZSG19, LYZ15, LY15b, LL16c, LGB16, LR19b, LP16b, LJ16, LSZ18, LMB18, LCCZ19, LKSM17, LDSM19, LHGF19, LYA16, Lot18, LLLN18, LHL15, LP17b, LSI16, MM16a, MRM16, MGT18, MS16a, MK17, MCW16, MNG15b, MVK16, MAP17, MDHC15, MW16b, MW17a, MDM^{+}15, MJ17, MGBG16, MSS16, MBBKTH17]. methods [Moc17, MW17b, MLB18, MSP15, NdlLPCC19, NJ15, NPC15, NDCB17, NN15a, NGY^{+}17, Nor15, OLDN17, O’S19, OWKE16, OKWE17, PP18a, PR17a, PP18b, Pea15, PT18, PAFT19, PDN19, PG18, DDM18, PMS19, PPM^{+}19, PR17b, PVB17, PSP16, QLS^{+}19, QSY16, QDH15, RFPSSA18, RFGSV15, RT16, RS18, RJ19a, RHvR^{+}15, RRMF^{+}19, RWN18, SZN19, SZN20, SGN16, SCN^{+}17, SV19, Say17a, Say17b, SLH18, SWML17, SZ15a, SC18a, SW18a, SS15b, SL19b, SLL16, SLL17, SY18a, Shu16, SDM^{+}17, Sid18, SX15, SE16, SGT16, SPN^{+}19, Sti16, SL16c, SZCL18, Suz18, TSC17, TK15a, TMT17, Teu16, TL15, Tsa15, Tsa16, UL16, VPV^{+}17, Vee16, VN15, VWV17, VLV19, WCN15, WJD16, WTGC16, WSN^{+}18, WVB19, WT19, WGME17, XWB15, XYG19, XGZ19, YSY17, YYL18, Yi18, YK19, YJM19, ZOG19, ZK15, ZM16a]. methods [ZX19, ZJL16, ZT17, ZH15, ZCZ19, ZZT^{+}16, ZXDL17, dFVJ15, dFJN16, dlCGCA17, ADG19, CEH16]. Metric [CGL18, SYV17, CCWY18, SYV14, VLAB18, YLH^{+}19]. Metric-based [CGL18, VLAB18]. metrics [KF15, LTR16]. MHD [BD15a, BBKS16, BK17c, CS16a, DWGW16, DWGW17, DWG^{+}18, FS18, HIN^{+}16, LZ17a, LSZ18, MHHX16, PNZ18, PE15, SY18b, WG16b, WDGW17, WRL18, YFJ17, YFJ18, ZSL^{+}19]. MIB [GZ17]. Micro [GJL19, EN17, HKH^{+}16, MKV^{+}17, SN15, SRBB18, SCQP16, WPB15]. micro-devices [WPB15]. micro-flow [HKH^{+}16]. micro-layer [SN15]. Micro-macro [GJL19]. micro-scale [SRBB18]. micro-swimmers [SCQP16]. microchannels [GZM^{+}17]. microdomain [AZ17]. microflow [HLQ16]. microflows [HH19]. microfluidic [LZ15b]. micromagnetics [KPP^{+}19]. micron [LLA19]. microphysical [SNB^{+}15]. microscale [BRK17]. Microscopic [VS17, FHE15]. microsphere [LJZ15]. microstructure [CPT16, LMM17]. microstructures [HS17a]. microswimmers [Str17]. microwave [BPF^{+}16, HK16b, NOM^{+}17, PKLS17]. midpoint [EMSS17, WH16a]. Mie [GHJ15]. migration [CCFC19, LZ15b, MMNI16, Par15]. MILU [PKJ^{+}18]. Mima [HK15a]. mimetic [GL17, KL17b, KD17b, LPG18, LP18, LMMS16, OvdHVH16, PKF16, PG17, Pei16, PSG19, TC15b]. Minimal [BTD16, LKN17, MP15b, PCX17, WC18, ZD15a]. minimalism [OSKN18]. minimax [HPV16]. Minimisation [Jou15]. minimization [BHZ16, CEL18a, GLZ16, GNZ18, JES15, LL16a, LT17c, PHD16, Tav15, VCEK19]. Minimizing [Iwa15, KPKGH19, Sto16, ZM16b]. Minimum [CGM18, CM15, RSB16, WY17, WA18]. miscibility [KS16c]. miscible [BST^{+}18, CXY19, Kwo19, LW17a, SHLG15, SWML17, TGY18]. Misfit [YLH^{+}19]. missing [CJK^{+}19]. mitigation [CFKK19]. mitral [CVM^{+}19]. Mixed [DS16, DH18a, Fal16, MF16a, RBL16, AVT17, ABN15, AGRB18, BNS17, BSM16, CWF16, Cha16, FC19b, GS15b, GVTQ16, HLA19, JL17b, KKL15, LPG18, LP18, LPR18, LR19b, LMG19, Mel18, MJ17, MR16b, NPP15, NRS19, RB15, SPB17, SY18a, SW18b, SHW18, YK18, YK19]. mixed-curved [YK19]. Mixed-hybrid [MF16a, AVT17]. mixed-primal [AGRB18]. mixers [MKV^{+}17]. Mixing [SMD18a, BLG^{+}16, ES18, VMK^{+}19, WSN^{+}15]. mixture [CZB15, PS14, PS15a]. mixture-energy-consistent [PS14, PS15a]. mixtures [BMY19, HHM17, KL17a, WZRZ15, Wu19, ZYSW16, Zoh17]. MLC [BZ19a]. MLFMM [XB18]. MLFMM-based [XB18]. MLMD [IBML16]. MLRPI [HSC16]. MMALE [CZJ17]. MMSISL [CBM19]. mobile [BVM17b, RZ15]. mobility [BDPM18, DD16a, EJZ17, MS17]. Modal [HB15b, GTL18, SZ15a, WSB19]. Mode [AM19, DDM^{+}19, FCW^{+}18, IG15, KH18, LWY18, LZL^{+}19, ZLL16b, ZLL17a]. Model [BHST17a, JL18b, LMPS15, NP16, SS15a, Sch16a, ARF18, AASPT18, AAG16, ABG^{+}18b, ABG^{+}19b, ASB^{+}15, AEL^{+}15a, AEL^{+}15b, AMB17, AEVW18, AZ16, AP16, Ama15, Ama18, APT17, ABH^{+}19, ADHN15, Ani16, AMM^{+}15, AZ19b, BJO18, BST19, BHdD18, BFI^{+}18, BIR18, BH16b, BM16, BVM17b, BTVB15, BLG^{+}16, BAK19, BKR19, BCG^{+}15, BG16b, BTVC16, CF15, CBA17, CCS18, CPT16, CCP19, CS16a, CL16, CLY^{+}15, CJYZ15, CZL^{+}15, CY19a, CLL19, CYS17, CB19, CGS15, CEHM19, CEH16, CV16a, CHS17, CJC19, CDV17, CGG18, DG16a, DLP19, DMRB19, DPK17, DKTH15, DCD^{+}18, DWW15, DKC15, DW19, EL18, FB17, FST15, FBJS19, FMPT18, FK17, FKY15, GHM15, GMS16, GZ19, GFL17, GCVCHH18, GGT18, GSMR18, Hag25, HFND18, HXLL15, HX16, HK15a, HCVH18, HHRA19, HLU15, HLQ16, HH19, HW15c, HW16b]. model [HW18, HWA19, HTvdH^{+}19, Hwa16, HY17, Ido16, IG15, Jac17a, JN19, JRPPS18, JL15, JD19, JS17, JCNK19, JJ18a, KM16a, KC17a, KL17a, KHP17, KK17a, Kor17, KBG^{+}15, KGP^{+}17, KDPK15, KRK^{+}18, Kwo19, LVB^{+}15, LP18, LZ18, LS15c, LZB^{+}17, LT17a, LWY17, LHMB18, LXC19, LZHM19, LHLL19, LLY18, LMY^{+}19, LZT^{+}15, LXSC16, LGZ^{+}19, LDGH16, LHW^{+}17, MMNI16, MNG15b, MP17, MCS^{+}19, MP15b, MHT^{+}19, MGB^{+}18, MCHL16, MHGL19, MAA18, MBA19, MSF^{+}19, NHM17, NFG15, NMA15, NCP^{+}17, NWKC16, NWZ18, OS16, Pan20, PD17, PM16, PS14, PS15a, PMGW16, PRvdL18, RBY19, RMK15, RKH15, ST16, SK19a, SRBÓ17, SN15, SPD^{+}17, SA16, SAH17, SRBB18, SD17, SYY15, SLB^{+}16, SLB^{+}19, SS16c, SZS15, TWN19, TP16a, TGY18, TTN^{+}16, TS17, TD16b, VST16, VMM19, VCNGP15, WMY16, WW17, WVB19, WSN^{+}15, XWW^{+}16, XYF^{+}17]. model [XZZ15, YFKS15, YCBC15, Yan16b, YH17, YZW17, YSLY19, YBSTT19, Yan17, Yas17, YP17, YY17, YCS^{+}17, ZL15a, ZC15, ZZL19, ZYSW16, ZHWQ18, ZW19, ZC19b, ZXL17, ZWB^{+}18, ZR17, ZWUR16, ZK18, dSPDH15, dPSS16, tEDKT17, ALA16, JL17b]. model-based [FK17]. model-form [XWW^{+}16]. model-order [ZWUR16, dPSS16]. modeled [STG17]. Modeling [BBMN18, BGD19, CSY15, DD17a, DD15, HFM17, PKP^{+}17, PMS15, SSL17, TEP19, TK16, AASRT17, ANL^{+}16, BB17, BLL16, BMR^{+}16, BH16b, BNGI19, BHGK18, BHTT17, CFG16, CW16, CP16, CCB^{+}19, CLvS17, CFPB17, CVM^{+}19, DSS18, Did17, DDH^{+}18, FB17, FSK^{+}16, Fuj19, GH17a, GFC18, GHR17, GQC^{+}19, GW16, HHA15, HGR16, HSK^{+}15, HU18, HMBH15, HKS^{+}16, JL18a, JTR16, JH15, JS16, JJ18b, KZ17, KSV^{+}15, KYKS19, KLC19, KS16c, KZG16, KW16, LYLK17, LHMB16, LWWY18, LMKS15, LGD17, LTXB17, LMM17, LYDB17, LHA16b, MKYZ17, MF17, MGKG17, MH19, MSV^{+}16, MAH16, MF16a, NLFM16, NWB19, NSK^{+}16, PD16a, Ram18, RTO15, SYOS19, SYOS21, SSL^{+}16a, SGC^{+}18a, SBG^{+}17, SMP16, SCQP16, SK15b, Sun19a, TCA16, TUJ19, TS19, TNB21, URL16, Vai15, Vos17, WMY18, WHR19, WB17, WC18, XML17, YG18, YYL16, YLZ^{+}19]. modeling [YL16, YPK16, YYJ^{+}19, Zad11, ZCHS15, ZZDB15, ZW16, ZG17, ZHLZ18, ZLC^{+}18, ZZ18, ZZKP19, Zoh17, dFGS^{+}17]. modelled [Mue18]. Modelling [LZ15b, RPNP18, RZ15, YXF^{+}16, ABG^{+}15, APLK19, BC18a, BPGS16, BHMS18, BB15, DLLV17, DMM19, FITcD19, FBC^{+}16, KMS^{+}18, LL19a, LWZ19, Mel18, MM16c, SS16a, SWS^{+}18, SMG19, SZF15, TAJ^{+}17, YSC^{+}17]. Models [CS18b, ABP^{+}16, AAI16, AS16, ATF16, BTD16, BKS18, BFM19, BLVC17, BH16b, BFNGDNR18, BCC^{+}18, BDMZ19, BK16b, BKRB15, CT15, CDM^{+}16, CCS18, CGK17, CS16c, CKQT15, CNQ^{+}19, CCM15, CMR^{+}16, DD17b, DDM^{+}19, DHH^{+}18, EL19, FOF15, FPT17, FS19a, GHV19, GMS19, GZ19, GLG^{+}19, Gri15, Gri19, GH17b, HAPK15, Hig15, HLQ16, KMD^{+}18, KKP15, KL17a, KS16b, KBF17, LM15a, LK17, LSP19, LPW15, LLL16, LZW19, LTWZ18, LPBR15, LPR19, MHHX16, MCN18, MXL16, MPP15, MRXI17, MTL^{+}17, MLB16, NJL19, Niu16, OTS17, PKW17, PSG19, PT17b, RK18, RKN19, RS15b, RBL16, SZY16, SGC^{+}17, SL19c, SFDE15, SSO^{+}15, SGA^{+}15, SM19b, TYD16, TE19, TSB^{+}18, TB18, VM15, VBG^{+}17a, VD16, WJD16, WTL17, WX17, XTS^{+}16, YNW17, YQNW19, ZA15a, dBIM16, dCMR19, dlCGCA17, DCP15]. moderate [XTYL18]. modern [GFA^{+}16]. modes [VMN^{+}18, GYZ19, KP15c, TBHG18, Tre16, WYLX17]. Modification [BK17a, Lau17, Ols15]. modifications [WS16]. Modified [BDMC15, BTA17, MTJ17, SW17b, WZ18a, ADOP18, HS18b, KVKS19, KDL15, LL19a, PKB15, PR16c, RFPSSA18, SwS16, Svä15, XJ16, YZZ19, ZLL16a, ZLL16b]. module [SDH^{+}16, SKG17]. MoF [CZ16, CZ17, CZ19b, QYJ19]. moist [ZA15a]. molecular [BBW16, BT17b, CGC17, CSCM16, Dav10, Dav15, DZ18, DFS16, FPASS16, Gen15, Han19, JLKF17, KBK15b, MDH19, MD15, QS16, RS17, ST15, SMAG17, SAOW17, TPTT18, WYLX17, WTS^{+}17, YSWW16, YT17, YZW17, YZW^{+}18, ZLH^{+}17, ZD17, ZHWQ18]. molecule [Vai15]. molecules [BKKY19, ELH^{+}16, LAA16, RKL18, SZCL18]. mollified [FHS17]. Moment [ABM16, LGB17, RKO^{+}17b, ZM16b, AS15, CSN18, CEHM19, DPW^{+}15, FITcD19, GHH15, HPC19, HLQ16, HH19, HTvdH^{+}19, LN17, LLSJ19, LH18, Nor15, PCMC19, PPM^{+}19, SBT17, SGP17b, TC15a, TKC15, TLQ16, WT19, WYA^{+}17a, WYA^{+}17b, XX16, XDSX17, XDLX19, JSS15, MKC17]. moment-accelerated [HPC19, PCMC19]. Moment-Based [ABM16, DPW^{+}15, TC15a, TKC15, TLQ16]. Moment-of-fluid [LGB17, JSS15]. moments [AKK^{+}19, DC18a, FLV18, GJL19, HKLZ18, STR15, SL16a, SGP17b, ZLX17, PMF^{+}18]. Momentum [IBML16, ALTR17, BDAA^{+}18, Bra16a, DL15, DS15d, JST17, KDL15, LBZA16, LM16, MR17, MBD15, MBD19, MFG15, OD17, RKH15, TCSM15]. momentum-interpolation [MBD19]. momentum-weighted [BDAA^{+}18]. monatomic [WZRZ15]. Monge [DL17, WBBC16]. Monodomain [CGG18, LZT^{+}15, VLP^{+}16]. monoenergetic [GMP16]. Monolithic [LRGO18, BVMW16, BZ16b, CM16a, HF19, PKLC16, PKLC17, PLC18, PAL^{+}16, ZS16, dLKK19]. monotonic [ZA15b]. monotonicity [DvW15b, MG15a]. monotonicity-preserving [DvW15b]. Monte [BC16b, CSS15, Gho17, Lan19, LPU18, Mac16, MNO^{+}17, AR16a, BP18, BTA17, Cha16, CL17, CSN18, CG15, CW18, CHE^{+}17, Cos16, DPW^{+}15, DG16c, EARA15, EN17, FDKI17, GB15b, GMS16, Gen11, GDS^{+}16, GAJ15, GBU15, Hig17, HC17, HMRG16, ION^{+}17, KM17, KMS^{+}18, KL16, KC17b, KES18, KK17b, KLGO18, LS15a, LBTCG16, LYCC17, LYB18, LB17, LXL17, LWL18, MZTS16, MSS16, NHA18, PJE^{+}16, PUA^{+}15, PDS15, Poë19, PvL19, QPK19, RFPSSA18, RRL19, RKH15, SY17, Swe18, TSR15, WBC^{+}16, WL16, XZZ15, XR17, YC15, YZZ19, Yas17, ZLJ16, Zil15, vdKK16]. Monte-Carlo [Mac16, BP18, Hig17, Poë19, RRL19]. MOOD [BLD15, BDLM18, JYY18]. Mori [PD17, ZV18]. morphologies [SMLB15]. Morral [Tav16]. mortar [BBB^{+}16, FC19b, KK17a]. most [TBHG18]. motion [AAB^{+}16, BT17a, CFSN18, HXLL15, LS16c, MP19, MLB18, PZNG15, PGCG18, RTO15, SKF15, TOR^{+}15, XL17a, ZB15]. motions [MNC19]. motivated [FHA17a]. motors [SDM^{+}17]. Moulton [ZM16a]. movable [IML15, LZHM19]. movement [RPNP18]. Moving [DHH^{+}18, MP19, SYV17, ZHLZ18, ABT16, BOA17, BNK18, BKR19, BD17, CV17, CTJ^{+}17, CHJT17, CGJY19, CE17, Chu17, CBM19, De18, DM16, DIX^{+}18, DCBK15, FPDT17, FPV18, HM19b, HW18, HKS^{+}16, JGS16, LSLA16, LY15b, LD15, LHA16a, LHQ19, LMC19, MTK15, MM18, NJPB17, NH17, ÖPHA15, PNZ18, PN18, PR16a, PF16, PR16c, PLPR19, PC19, QLF16, QSB18, SL17, SYY15, SP19b, STKH15, SYV14, SSA17, SC16, SLdTV18, Sub15, TBO^{+}16, TMH16, Vre17, Vre21, YHQ15, YY17, ZZ17b, ZLL^{+}17b, ZLGS18, dTP16, Gam15]. moving-least-squares [dTP16]. moving-water [CV17]. MPACT [CSK^{+}16]. MPDATA [KS17, WKPS18]. MPFA [AEL^{+}17]. MPI [GBR15, JL18a]. MPP [ADK^{+}17]. MPS [Gam15, TCB18]. MR [HLL^{+}16]. MRAG [RHvR^{+}15]. MRAG-I2D [RHvR^{+}15]. MRI [NJHL18]. MRT [APT17, KGT15, PMGW16, ZZL19]. MRT/TRT [KGT15]. MS [BHTT17, CLNH15]. MS-DFM [BHTT17]. MsRSB [SMT^{+}16]. Müller [Fal16]. Multi [BLK15, BB15, Del15, DP19, GIF18, HSM19, IBML16, JL15, KS16c, LCK16, MKB19, MSS16, MOR18, NFG15, Par15, PVPK17, RFPSSA18, RZ17, RHvR^{+}15, SDM^{+}17, Sun19b, TCB18, TWN19, ZLX17, ABG^{+}15, AEL^{+}15a, AEL^{+}15b, ALO18, APT17, ATC19, AB15, BZ19a, BA15, BZ15, BCJ19, CPSF17, CXH15, CHZ16, CEL^{+}18b, CGP16, CLV19, CHLZ17, Did17, FNNW19, FPASS16, FX18, FB15, FPT17, GMD19, GEZK16, HYK^{+}16, HPC19, HHA15, HHM17, HTZG17, HK19b, HSB16, HHLY17, HTvdH^{+}19, JL18c, JLLZ15, JJ17, KHP17, KJ17a, KKLS17, KLRT15, KSSL18, KC18, KLWQ17, Kwo19, LLD^{+}16, LKB15, LML^{+}16, LKK17a, LKK17b, LLM17, LYB18, LLY15, LW15b, LFDP16, LCLY19, LC17a, LK16a, LH17b, LH18, LHA16b, MZAF17, MCL19, MSD^{+}17, MN18b, MCS^{+}19, MBM^{+}15, NKN^{+}17, NLK^{+}15, Nor15, OT15, PHHA18, PPCK17, PK16, Par17, PN18, PT18, PLWJ16]. multi [PBA^{+}15, QWZ19b, RPK17a, RKN19, RS17, RXSG15, RXS16, SPX^{+}18, SDJU15, Sha17a, SWJG19, SPM^{+}15, SS16c, SD16, SF16, SL16c, SM19b, TC15a, TKC15, TCSM15, TCS17, TZ16, Tav16, VSM16a, VSM16b, Vos17, WDS15, WED15, Wil19, WHZ18, XX16, XDSX17, XDLX19, YZ19, YK18, ZKS^{+}15, ZZ17a, ZYK18, ZF18, ZRW19, ZSO^{+}19, ZS18, ZS19a, ZS19b, ZPE^{+}16, dlCGCA17, IBML16]. multi-block [MSD^{+}17, PLWJ16]. multi-body [PLWJ16]. multi-color [KLWQ17]. Multi-component [LCK16, ATC19, Did17, FB15, HHM17, KS16c, LFDP16, Tav16, Vos17, ZS19b]. multi-continua [CEL^{+}18b]. multi-core [CPSF17]. multi-dielectric [HK19b]. Multi-dimensional [Del15, GIF18, BA15, CGP16, FNNW19, GMD19, HPC19, JL18c, JJ17, LK16a, MBM^{+}15, PK16, QWZ19b, RXSG15, RXS16, SF16, TCSM15, TCS17, WDS15, YK18]. Multi-Domain [IBML16, RZ17, BCJ19, CXH15, LC17a, PBA^{+}15]. Multi-element [JL15]. Multi-fidelity [PVPK17, ZLX17, LKK17b, PPCK17, RPK17a, RKN19, RS17, YZ19, ZYK18]. Multi-Fluid [SDM^{+}17, HTvdH^{+}19, Kwo19, LLD^{+}16, MCS^{+}19, NFG15, PN18, SS16c]. Multi-frequency [Par15, HPC19, NKN^{+}17, Par17, ZZ17a, ZF18]. multi-grid [LML^{+}16, LLM17]. multi-group [SWJG19]. multi-layer [BZ19a, KHP17, LCLY19, SD16]. Multi-level [HSM19, MSS16, KC18, LYB18, LLY15, IBML16]. multi-material [HTZG17, KKLS17, MZAF17, MCL19, PHHA18, VSM16a, VSM16b, Wil19, ZSO^{+}19]. multi-medium [LH17b]. multi-mesh [WHZ18]. multi-moment [LH18, Nor15, XX16, XDSX17, XDLX19]. multi-output [ZKS^{+}15]. Multi-particle [TWN19, LKB15]. multi-phase [ABG^{+}15, AB15, FPT17, HSB16, OT15, SPX^{+}18, Sha17a]. multi-physically [WED15]. multi-physics [MN18b, PT18, SM19b]. multi-point [AEL^{+}15a, AEL^{+}15b]. Multi-rate [MKB19]. multi-regime [CLV19]. multi-relaxation [APT17]. Multi-resolution [BLK15, RHvR^{+}15, TCB18, LKK17a, ZS18, ZS19a]. Multi-scale [BB15, DP19, KS16c, MOR18, ALO18, FX18, HHA15, HHLY17, KJ17a, KLRT15, KSSL18, LHA16b, SDJU15, SPM^{+}15, ZPE^{+}16, dlCGCA17]. Multi-solid [NFG15]. multi-source [NLK^{+}15]. multi-species [HYK^{+}16, TC15a, TKC15, TCSM15]. Multi-stage [RFPSSA18, FPASS16]. multi-step [ZRW19]. Multi-symplectic [Sun19b, CHZ16, CHLZ17, LW15b, SL16c]. multi-term [BZ15, JLLZ15]. multi-zone [GEZK16]. multiblock [ADFG17, MDHC15]. multibody [FS15, MHT^{+}19]. multicanonical [CL17, WL16]. Multicomponent [FSB16, GSS15b, HG17, LWB^{+}16, MZTS16, NF17]. multicontinuum [RBL16]. multicore [AAB^{+}16, RHvR^{+}15]. multicube [LTR16]. Multidimensional [BD15b, BN17, SWMD17a, SWMD17b, SS17c, BD15a, BAGK16, BK16a, BTGM17, BK17b, BGTM18, BK19a, BLD15, BDLM18, DLM18, DCCH19, EG17, KD17a, Kay15, KS15a, MD18, MS15b, Mas18, SJX17, SW16]. multidimensional-like [KD17a]. multidimensions [DLNR18]. multidomain [OVP15]. Multifidelity [KMP^{+}19, QPK19, SGC^{+}18a, YBSTT19]. multifluids [SWPS17]. multifractal [RWG18]. multifrequency [TWM18]. multifrequency-grey [TWM18]. Multigrid [JX15, LMJ18, RM16, ANL^{+}16, BMMP19, BST^{+}18, BCB17, CXX16, CGC17, CR18, CV16b, EJMI18, FGLW18, HF18, HSF19, KLGO18, lLLNS16, lLNS17, LRGO18, MR17, MM16c, MDDM17, PHHR17, Pop15, RWG18, RRMF^{+}19, RWN18, DD17c, Sti16, WMYG16, YM17a, Yi18]. Multigrid-based [RM16, CV16b]. multigrid-framework [MR17]. multigroup [ACJ17, Her16, PCMC19, WKOE17]. Multilayer [BFNGDNR18, Jou15, CDV17, FNGDMNR18, PSG19]. multilayered [Cho19]. Multilevel [KMS^{+}18, Ani19, CvKH16, CFvKH18, FDKI17, Gen15, HCVH18, HXX18, JC17, KLGO18, RS15b, SLR^{+}16, XZ15, ZNS19, BC16b, LPU18]. multimaterial [BMR^{+}16, SDJU15, SO16, dBIM16, JSS15]. Multimodal [FK17, LL15]. multiparameter [PGH15]. Multiparticle [AWS16]. Multiphase [BDPM18, Don18, WSHT15, WX17, YD18, APP^{+}16, APT17, BJ15, CZL^{+}15, CD17, CM16b, CLNH15, CvKH16, DM17a, DIX^{+}18, DB16a, Don17, FBL17, FITcD19, FC19b, GHF19, HFM17, HG17, HW16b, HW18, IM17a, JT18, KP15a, KRK^{+}18, LK17, LRA17, LT15, LZT17, LWX19, LHA15b, MHL17, MF16a, NPB19, NGPB19, OD17, PN17, RW15a, RJ19b, SV19, SPD19, STW16, SCJ^{+}18, SPW18, TP16a, TO19, Vos17, WSY15, WSS^{+}15, WCH^{+}17, WG19, YLZ^{+}19, YXF^{+}16, YCS^{+}17, ZN18, ZC19b, dFVJ15]. Multiphysics [JS16, CGM15, FFJ^{+}19, Sla16, TT17b, WKHT19]. Multiple [BW18a, CDX18b, EARA15, Hig15, PBC^{+}17, AN15, ARTG^{+}19, CC16a, CCZC16, CB19, DNBH15, FBL17, HMRG16, LR19a, LK16a, LY16a, LVL18, OMYvdP^{+}15, PKW17, RC18, SGA^{+}15, WBC^{+}16, WCT18, XB18, Xie15, YNW17, ZZDB15, ZZ19, SP19a]. Multiple-correction [PBC^{+}17]. multiple-direction [LK16a]. multiple-reflection [XB18]. multiple-relaxation-time [FBL17, Xie15]. Multiple-resolution [CDX18b, OMYvdP^{+}15]. multiple-scale [LY16a]. multiple-species [SGA^{+}15]. Multiple-time-stepping [EARA15]. multiplicative [AAE19, DDV^{+}15, HJZC17]. multiplier [FG16]. multiply [HN17a]. multipoint [Shi19]. multipole [AC17, BS19a, HLL^{+}18, JDFS16, LLEK17, RSBS19, TCD17, YS18a, ZGD^{+}16]. multipole-to-local [YS18a]. multireaction [DY19b]. Multiresolution [BT17b, YT17, ZOY^{+}19, BCO^{+}15, BDM17, DBMB15, HW16a, LDSM19]. multiring [GFL17]. Multiscale [ADG19, AASRT17, AEVW18, BLL16, BHTT17, CHT17, CCK^{+}17, CJ17, GV18, GFG^{+}15, GH17b, GMT19, KYKS19, LE16, LYDB17, MT18, MT19, PD16b, SS16a, SBH19, dMRHJ17, BJO18, BZ16a, BM16, LL17, BRK17, CE18, CEL15, CEH16, CELZ18, CEL18a, CPP19, Cot16, CLNH15, DGW18, DD17b, DLR15, EZG16, ELH^{+}16, FAC^{+}19, FCL19, FC19b, GFC18, HF18, JTR16, JL15, KKP15, KZ17, KAR17, KEJ18, KK17a, LJ19, LPBR15, MVKD15, MGKG17, MH19, MTL^{+}17, ML16, NS19a, NS19b, NGS16, PJB^{+}19, PC19, RWG18, SMT^{+}16, SSO^{+}15, SDW16, TPT16, TWH15, TAH16, TRL15, TL15, UHKT19, XCX17, YB17, Zau16, ZS16, ZZDB15, GAS^{+}18, TKB^{+}15]. multiscaling [Lot18]. Multislope [LMG15]. multispecies [DY19b, TCS16a, ZLJ16]. multispeed [LMPS15]. multistep [SSZ19, SL19b, Ter18, VK16]. Multithreaded [RB18]. Multitrace [JHPAT17]. Multitrace/singletrace [JHPAT17]. multitude [QLF16]. multiUQ [TO19]. multivalued [FFW17]. multivariate [BMY19, KM16b, PPM^{+}19]. Multiwavelet [GCVMK15]. Multiwavelet-based [GCVMK15]. murmurs [SBG^{+}17]. MUSCL [BR16, BR15a, BR15b, LMG15]. musculo [KBG^{+}15, KGP^{+}17]. musculo-mechanical [KBG^{+}15, KGP^{+}17]. MUSIC [AJP15, PKLS17]. MUSIC-type [AJP15]. myocardium [VLP^{+}16].

N [Gho17, XWZ^{+}18]. N-SDS [XWZ^{+}18]. N-SDS/MC-IP [XWZ^{+}18]. NACA [FW17]. Naghdi [GPTK19, LM15a, LXC19, Pop15]. Nagumo [LZT^{+}15]. Nano [Eva18, BLL16, CFPB17, HCW15, ZHWQ18]. Nano-particle [Eva18]. nano-pore [ZHWQ18]. nano-transistors [HCW15]. nanocapacitor [SCE^{+}19]. nanocomposites [LYDB17]. nanogap [VCNP18]. nanometer [SSL^{+}16a]. nanoparticles [SAH17]. nanophotonic [LDO^{+}19]. nanopore [DWZ19]. nanopores [MBBKTH17]. nanoscale [SSVL18, YT17]. nanostructured [SU15]. nanostructures [EDW19, HC17, VCNOP18]. nanowires [BDPM18]. narrow [LZ19]. Nash [TZ16]. natural [CB18b, EFO19, EFO20, LDSM19, MTJ18, PKLC16, PKLC17, SL17, WSF17]. Navier [BKR19, HW15a, LR19b, YDLC19, AD15, ALL18, AB17, BTD16, BTVB15, BPD19, BHF15, BC16c, BCJ19, CHOR17, CS16c, CYL^{+}16, CYYL18, CDN17, CCKQ15, CLP16b, DRZ^{+}19, Du18, FWK17, FBW16, FZ19, GN19a, GNK18b, GTG15, HPY18, HGW18, HTMP17, JPSX18, LM18, Ler16, LXC^{+}15, LZB^{+}17, LT17a, LHMB18, LYD19, LM16, LFAR19, MS18b, MLM18, MPFL16, MS18c, MHS16, MR16b, MN18c, NGPB19, NN19, OvdHVH16, PG17, PXLL16, PX16, PCN15a, PCN15b, Pea15, PND16, PDRB17, PBBK15, PMB18, RDM15, SHLG15, SMS16, SLB^{+}16, SLY16, SE16, Stü15, Stü17, Svä15, TD16a, TD17, TXKvdV15, TXKvdV16, UL16, UY19, VV19, WY17, WR15, WZ18a, WWZ19, XWW^{+}16, YXW19, YC17, YTW15, YT19, Zha17c, ZLFW18, ZCL19]. Near [LWZ16, Liu19a, AM18, BFF19, CV15, KW15b, LW15a, Liu16, LZL^{+}17, MS17, Ols15, RPNP18, ST16, SX16]. near-boundary [Ols15]. near-coplanar [KW15b]. Near-field [LWZ16, Liu19a, LW15a]. near-limit [LZL^{+}17]. near-optimal [AM18]. near-wall [MS17]. Nearest [GKMS17, Smi18]. Nearest-neighbor [GKMS17]. nearest-neighbour [Smi18]. nearly [KCHW19, LLSJ19, Tsa16]. nearly-singular [Tsa16]. need [RFGSV15]. negative [OADN19, SiI16, VMK^{+}19, YC15]. neighbor [GKMS17, TST^{+}15]. neighbour [Smi18]. nematic [HHZZ19, KLWQ17, ZYSW16]. Nernst [DWZ19, LW17e]. nerve [MW16a]. Nested [PSMPG17, PRvdL18, LZT17, LM19c, SLY16]. net [CMDL18, PYK19, FBY19, LLD19]. Network [UHKT19, BBB^{+}16, BLVC17, BS19b, FBY19, FKS19, KJ17a, LLD19, RH18, RH19, TB18, VLP^{+}16]. networks [AMJ17, BFM19, BPS16, BK16b, BBBG15, CJK^{+}19, CTM^{+}16, Cot16, EEG^{+}15, FNNB19, GZ19, HZE19, HU18, KEJ18, MWD16, MPT16, MB15, MMW15, Noe15, PVFN15, QWX19, RPK19, SSDN15, SCL19, WHR19, WRL19, YP19, ZLGK19, ZZ18, FBC^{+}16, SHTY19]. Neumann [JTD16, SYOS21, TNB21, ABN15, BK17b, BK19a, CJWS19, Cha16, DGHP17, GBD17, GSL^{+}19, MK15, PKJ^{+}18, PS17, RJ19a, SYOS19, WSY16]. Neural [FBC^{+}16, PYK19, BFM19, CJK^{+}19, FBY19, GZ19, HZE19, HU18, QWX19, RPK19, RH18, RH19, SCL19, TB18, WHR19, WRL19, YP19, ZLGK19]. Neural-net-induced [PYK19]. neuromechanics [PBP18]. neutral [Ama15, Ama18, DDD17, Fon16, GMP16, GBD^{+}15, KKS15, KKS16, Luc15, TSFS17]. neutral-fractional [Luc15]. neutron [ACJ17, BABD16, BK19b, BCG^{+}15, CKKD19, CSK^{+}16, HL16b, JPLL15, Lau17, LBB^{+}17, OWKE16, SWJG19, WKOE17, ZCL17]. neutron/photon [BCG^{+}15]. Newman [CY19b]. Newmark [Fuj19, RGPS17, SSM^{+}17]. Newmark-conformal [Fuj19]. Newton [AB17, ALTR17, CDLR19, CZ19b, LSMS17, LSCC19, PKN17, TMS^{+}19, WNDB19, WKHT19, YSY17, YKMM19, ZHLZ18]. Newtonian [AS17, CSB15, DPRZ17, JN19, RV16, STKL19, TL17, ZLC^{+}18]. NFFT [NPP15, WNW^{+}19]. Nicolson [FBF15, HYL17]. NILSAS [NT19]. NILSS [NW17, NWFT19]. nine [DWG^{+}18]. nine-wave [DWG^{+}18]. Nitsche [GY18, JGS16, ZSX17]. NLT [YXX^{+}16]. NN [SW17b]. no [YS18b, dDPG19]. no-slip [YS18b, dDPG19]. Nodal [QDH15, CM18c, EKEB16, FCL17, GWK16, LWLC17, LSTkM15, TVB^{+}16, WWGK17, XJLQ15, ZS16]. node [JPLL15, PG18, SGP17b, ZY17]. nodes [GZLH19, PR17a]. Noh [VW18, VGZ18]. Noise [YR15, BBK19, CHZ16, CVG18, CHLZ17, DWR18, HJZC17, KH15, MGT18, RRS19b, RRS19a, RKB19, XZJK19, ZLL17a, ZPE^{+}16, ZRE16]. Noise-induced [YR15]. noisy [CWL^{+}16, RPK17a, SWX18, SF16, SS18c, YM19]. Non [ALMJ15, CZBC^{+}18, CEL^{+}18b, HU18, LHGF19, NWFT19, PT17a, RS18, RRD16, SSZC19, WHR19, vdBKD17, ÅN19, AMH^{+}18, AD15, ACCCD^{+}17, ADGN17, ALKZ16, AS17, AB16a, AZ16, ABFR16, AHZ19, AB15, Bat17, BLVC16, BNGI19, BWR15, Blo17, CPX19, CMP19, CC17b, CKK18b, CFF18, CSB15, CwYjS16, DLP19, DRM15, DKTH15, Dom18, DB18, DB16b, EDW19, EDC19, FL18, FN17, FKS19, GM19, GMLD18, GN16, GMA18, GL17, HYK^{+}16, HFND18, HWH^{+}16, HFM17, HKKP16, HWA15, HY16, IM15, JN19, JKM19, KKLS17, KJYC17, KZR15, KBR17, LR19a, LM18, LH15, LB15, LYC16, LW17c, LLJJ18, LJ16, LAA16, MG15a, MK17, MM15, MPP15, OKE17, PK17, PL16b, PBL^{+}19, RZOZ19, RJLW19, STR15, SSL^{+}16a, SL17, Sel15, SS16b, SYM17, ST15, SPP16b, SLA^{+}19, Spe15, SCC19, SK18, TUJ19, TXKvdV15]. non [TSST16, TKF17, VMK^{+}19, WR15, WWRS17, WMS18, WG15, XYF^{+}17, XML17, YS18b, YS15, YY16, YHKPF17, YLLH19, YYJ^{+}19, ZFPB16, ZD15a, ZLC^{+}18, ZPW18, ZZT^{+}16, ZQ17, dlHC16, NW17, NT19]. non-adapted [SL17]. non-adiabatic [BLVC16]. non-aligned [KKLS17]. non-blocking [LH15]. Non-body-fitted [CZBC^{+}18]. non-canonical [ZZT^{+}16]. non-classical [Spe15]. non-conformal [ADGN17, Dom18]. Non-conforming [RRD16, ÅN19, FKS19, GM19]. non-conservation [SPP16b]. non-conservative [BNGI19, CC17b, DB16b]. non-constant [OKE17, WG15]. Non-convex [LHGF19, CFF18, IM15]. non-Debye [MG15a]. Non-deteriorating [PT17a]. non-dissipative [AMH^{+}18, JKM19]. non-equidistant [WWRS17]. non-equilibrium [BWR15, CwYjS16, DRM15, GMLD18, HFM17, HKKP16, MPP15, STR15, WMS18]. non-flat [KJYC17]. non-Fourier [ST15]. non-Gaussian [RZOZ19, ZFPB16]. non-graded [Bat17, SLA^{+}19]. non-gradient [GMA18]. non-Hermitian [ZD15a]. non-homogeneous [HWH^{+}16]. non-hydrostatic [AZ16, EDC19]. non-ideal [PL16b]. non-instantaneous [DLP19]. Non-Intrusive [NWFT19, HU18, WHR19, vdBKD17, Blo17, CPX19, CMP19, HFND18, XYF^{+}17, YYJ^{+}19, NW17, NT19]. non-isothermal [BLVC16, RJLW19, TUJ19, XML17]. non-iterative [LR19a, TKF17, YS15]. Non-linear [ALMJ15, AD15, ALKZ16, GN16, HYK^{+}16, KZR15, PK17, PBL^{+}19, SCC19, YHKPF17]. Non-local [CEL^{+}18b, DKTH15, EDW19, SSL^{+}16a, dlHC16]. non-locality [MK17]. non-matching [FKS19, LLJJ18]. non-negative [VMK^{+}19]. non-Newtonian [AS17, CSB15, JN19, ZLC^{+}18]. Non-normal [RS18]. Non-oscillatory [SSZC19, CKK18b, DB18, HWA15, HY16, LJ16, SK18, ZPW18, ZQ17]. non-overlapping [AB15]. non-periodic [YS18b]. non-polynomial [LW17c, YY16]. non-reflecting [FN17]. non-relativistic [Sel15]. non-slip [LM18]. non-smooth [MM15, YLLH19]. non-stationary [ACCCD^{+}17, TSST16, ZFPB16]. non-symmetric [GL17]. non-tensor [ABFR16, LB15]. non-uniform [AB16a, AHZ19, FL18, LYC16, PL16b, SS16b, SYM17, WR15]. non-uniformly [LAA16]. non-zero [KBR17]. Nonaffine [CS16b]. Nonaffine-parametric [CS16b]. nonconforming [BTT18, MNR19, VPM15, ZZYC19]. nonconservative [Ren19]. Nonequilibrium [PSN^{+}19, SPN^{+}19, DFS16, SKO18]. nonequilibriums [DZ18]. nongradient [YPC19]. nonhydrostatic [SZS15, Yi18]. nonintrusive [APLK19]. nonisothermal [BMT18, KS18b, YG19]. Nonlinear [BGM16, LT15, dPSS16, ACJ17, ATZ16, ABR16, ALT17, AHZ19, AEAM15, ANL^{+}16, BM15, BHKS16, BJTZ15, BZ19b, VMN^{+}18, BM16, BK18, BCJL17, BDMZ19, CBZ18, CBA17, CRW16, CGMH18, CS16b, CM19, CLP16a, CG18b, CHLZ17, CC19c, DSS18, DLL^{+}17, EAAM15, EKEB16, FW18, FSWW17, FBF15, FKR16, GJL19, GS15c, GFL17, Gno17, GWWC17, GXX17, GP16c, HdBH^{+}16, HU18, HHCG15, HJS19, HAH16, Hue15, KM17, KM16b, KC17c, LMS17, LM15a, LL19a, LZ16, LYC16, LGH^{+}18, LZ17b, LZZS15, LW17d, LIW18, LYDB17, LSCC19, MD17, MD18, MBJ16, MBNJ16, MK17, MGKG17, MW17a, MDP18, MMP18, NS19a, NBH18, PE16a, PDN19, PBA^{+}15, DM18, RK18, RPK19, Ren19, SPP^{+}16a, STEK17, STEK22, SAEF17, SWG^{+}17, SY16, SL15, SK15a, SL19b, SGC18b, SY18a, SYM15, SPW18, SS18c, SJH^{+}15, SCS18, TCS17, TT17b]. nonlinear [TMT17, TT19, WH15, WMY18, WL17, WHT18, YSC^{+}17, YYN^{+}17, YNW17, YL16, YJB18, Yeo19, YL17, ZK15, ZTBW19, ZVO15]. Nonlinearly [YSY17]. Nonlocal [MGT18, ATZ16, BJTZ15, CP16, DWW15, DY17, DJLQ18, EMZ16, LS19a, QWZ19b, SMD18b, VCNOP18, WW17, XJ16, ZGJ16, ZK18]. nonoscillatory [BR17, HBR15, MSP19, YC17]. nonparametric [LZ18]. nonplanar [DD16b, WHY18]. nonrelativistic [BZ19b]. nonseparable [ZKS^{+}15]. nonsmooth [XY18]. nonstandard [FBY19]. nonsymmetric [EJMI18]. nonuniform [BJTZ15, JL17a, DV17]. norm [BD16, CM15, CGM18, DBZ17, Mat17, MAvdW18, MO18b]. norm-oriented [BD16]. normal [IM15, RS18]. normalized [HK16b, Rua18]. note [AM17b, HS17b, Sir19, Teu15, YY16, ZW15]. Novel [Mue18, RC18, ZNS19, BTVB15, BND16, DC18b, DWGW16, DWGW17, DvWZ18, FFJT16, FLHA17, HY17, JLKF17, KD17a, KM15, LAEK18, DV17, MMvR18, PM19, PN17, TCL15, VST16, WS16, XTYL18, YTW15, ZL15c, ZRE16]. November [Ano19c, Ano19-29]. nuclear [ABdC^{+}18, DDJ17, GDS^{+}16, HBC^{+}16, HLS19, MTL^{+}17, PBA^{+}15]. nucleate [SN15]. nucleation [FSK^{+}16, KES18, KK17b]. NUFFT [Yan19]. number [ABIR19, BBKS16, BLMY17, BAK19, BFNGDNR18, Bon17, BKG15, DCP15, Eva18, GSN17, HR19, LLFX18, LWL18, MM16a, MA19, MDP^{+}15, MBD15, MBD19, MA16, MDAB18, NL18a, Pan15, RFGSV15, SP15a, SK18, TASA19, WSY15, WDT^{+}19, WC18, WDGW17, WGME17, YSYW19, ZV16]. number/compressible [MDAB18]. numbers [FMPT18, JdR^{+}18, KJ17b, TD17, XTYL18, ZZ19]. numeric [LLD19]. numeric-symbolic [LLD19]. Numerical [APR^{+}15, ALA16, Azi19, BLVC16, BTT18, CRW16, CCZ18, CPSF17, CC17c, CCZ15, CVK16, CV16a, DLLV17, DGHP17, DMRB19, DNOP15, DvW15a, DLWY19, EKSS15, HGR16, HB16, HR19, HX16, KS16b, KYW^{+}16, KYW^{+}18, LW15a, LYD19, LLVF^{+}15, LMM17, LAA16, LM15d, Mac15, MSG18a, MSG18b, MR16a, MC15, MFB18, NKN^{+}17, OTS17, OMYvdP^{+}15, PM16, PSV18, RS15a, RF18, RdM19, STKH15, SAF^{+}19, SDFA17, SPN^{+}19, Str17, SS17c, SZCL18, TBHG18, Tou18, WHL17, WL18, WSH19, WX19, YSWW16, YZW17, YY17, ZB15, ZZ17b, ZZPH18b, ZZPH18a, ZLL^{+}17b, ZS17, dLDG^{+}18, ABI17, AAG16, ABG^{+}18b, ABG^{+}19b, ASB^{+}15, ASWvD19, ADH^{+}16, AM17b, BZ19b, BCB15, BS15a, BDBEE15, BR15a, BCC^{+}18, BK16b, Bou19, Bre18, BC16d, CPdS19, CM15, CW16, CWL^{+}16, CY19a, CSC19, CLL19, CYS17, CSG17, Cho19, CLGA17, CELI15, CL19b]. numerical [CC19c, DM17a, DS15a, DDJ19, DSS18, DLNR18, Dod17, Don15b, DLS15, DL18c, DBMB15, EH14, EH15, EDW19, FNGV18, FW17, FB15, FFJT16, FPV18, GB15a, GJL19, GP16a, GO15, GLS15, GN16, GEZK16, GA18, GGT18, GFW16, HPY18, HW15a, HO15, HZL^{+}15, HM16a, Heu17, HN17b, Hu17, IM17a, Jac17a, JN19, JSVD17, JL16, KTN15, KGT15, KPJ18, KLNH17, KCS^{+}17, KHTZ19, KHC^{+}16, KV16, Lap16, LVB^{+}15, LE16, LRA17, LS15b, LZ15a, LWY17, LZW19, LCLY19, LZ19, LB16, LFT^{+}16, LSYF15, LP17b, LMC19, MOAA15, MS15c, MW16a, MST15, MDMS18, MHZ^{+}15, MN16b, MTK15, MTK17, MA16, MW15, MC17, NMM15, NMM17, NPRC15, NLFM16, NN18, NT16, OC18, PP18a, PC16, PS14, PS15a, PGGW18, PBL^{+}19, PT17a, PWC18b, PZF16, QN19, RW15a, RMK15, RLV16, RMF^{+}18, RJ19a]. numerical [RZ15, SZY16, SLL19, SPD^{+}17, SSL^{+}16a, SWML17, SVG18, SAK18, SRBB18, SMD18a, SYY15, SLL16, SYM15, SM19a, SPP16b, Sov16, SD16, SPRW15, SK15b, Suz18, Svä15, TGS19, TM15a, TCL15, UWH17, Vai15, VPM15, VST16, VLP^{+}16, WMY18, WBBC16, WAF^{+}19, WCL15, WTL17, XYPT16, XLL^{+}17, XS15, XML17, XY18, Yan16b, YFJ17, YFJ18, YDN19, YM17b, YXX^{+}16, ZCHS15, ZSP15, ZN18, ZHS18, ZC19b, ZWYW18, BFFB17]. Numerically [CCPdL19, LDHJ15, Vab15, LZ16]. Numerics [KHP15, LLS15]. Nunziato [CHS17, DG16a, FRRV16, LDGH16, TT16]. NURBS [MH18a, SNSG16]. NURBS-based [SNSG16]. NURBS-enhanced [MH18a]. Nyström [APKP16, CCZC16].

obeying [HK15a]. Object [WW16]. Object-oriented [WW16]. objective [FC16]. objects [GWB^{+}15, LO19, LB16, SUR18, SF18a]. oblique [DMM19]. obliquely [GD19]. observation [Yeo19]. observations [BGHK19, CM19, WHRL19]. observations-based [WHRL19]. observer [CCM15]. obstacle [LW15a]. obstacles [BNM15, BFP18, DM16, HGW18, LR19a, ZZ17a]. Ocean [SS15a, CGSS18, DDM^{+}19, Hig15, Kor17, Liu19a, NWKC16, PP18a, PSG19, SP16a]. oceanographic [FDS^{+}15]. oceans [MDW18]. October [Ano19-28]. octonion [ARTG^{+}19]. octonion-based [ARTG^{+}19]. Octree [MC16, GZLH19, HS18a, JL18a]. octree-based [JL18a]. Octrees [GTG15]. ODE [CFG16, CB15]. ODEs [BK16b, CNW17, OZ17, TSC17]. ODEs/PDEs [OZ17]. Off [HHK15, RKL18, HRJ^{+}16, RS15a, ZWG17]. Off-centered [HHK15]. Off-lattice [RKL18, RS15a, ZWG17]. offline [ABI17, SFDE15]. offline-online [ABI17, SFDE15]. offsetting [uKHGK19]. offshore [CGSS18, GPAO^{+}18]. often [SL19c]. oil [ASB^{+}15, WLC15, YSLY19]. on-the-fly [EZG16]. One [Hue15, PKK18, Ram17, SL16b, TC15c, AR16a, APR^{+}15, AS15, An17, BDB^{+}17, CHJT17, CGJY19, DZ16, Hiv18, LDT19, LSTkM15, LW17e, MN18a, Mag19, MB15, MLB16, SV19, TZSS17, Ter18, VSM16a, WRL16a, ZS19b]. one- [LDT19, ZS19b]. One-dimensional [Hue15, Ram17, AR16a, APR^{+}15, CHJT17, CGJY19, DZ16, Hiv18, LW17e, MN18a, Mag19, MB15, MLB16, SV19, TZSS17, VSM16a, WRL16a]. one-shot [BDB^{+}17]. One-step [PKK18]. One-way [SL16b, TC15c, Ter18]. Online [CPP19, ABI17, CEL15, CEL18a, SFDE15]. only [YD19]. onset [SGN16]. onshore [GPAO^{+}18]. opacities [Gen11, Gho17]. Open [HKH^{+}16, BLS16, BBK19, Don15a, DS15c, GTL18, JSY15, LXC^{+}15, MBM^{+}15, STFK19, WNW^{+}19, YD18]. open-boundary [NYD19]. open-rotor [BBK19]. OpenFOAM [BGV17]. opening [OZ17]. operation [FBC^{+}16, HSF17]. operational [BDBEE15, EE16]. Operator [AMXJ19, BOD19, MM16a, Vos17, BPL19, BNK18, BTVB15, CT15, CGQ18, CARdN19, CGS18, CKQT15, CLX15, DDV^{+}15, FS19a, GP16c, HYK^{+}16, HS17a, HM19c, Kas15, KV16, LSL15, LW17d, LYPP17, LZL^{+}17, MLM18, SZY16, SZ17, SLN15, TCS17, WC19, WDGW17, YS18a, ZOY^{+}19]. Operator-adapted [BOD19]. Operator-based [AMXJ19, Vos17]. operator-split [BPL19]. operator-splitting [KV16]. operators [ÅN19, DBZ17, DWGW17, DY17, LS19a, LN15, LKN17, MN04, MN17, Mat17, MAvdW18, MO18b, OKE17, Pei16, RÖS16, RÖS17, Ran18, SPB18, SKO18, SMD18b, Sub18, TMdO19, Vab15, YDN19]. optical [BKKY19, BCJL17, KLWQ17, Pis18, TT19]. optically [BLL16]. optics [BM15, WT16, XB18]. Optimal [AMPG19, CG19, FYZ^{+}15, FMPT18, FS19b, KDF15, LHMB16, OKE17, RG15, VLV19, VL15, WHRL19, YYL16, AM18, BIR18, BMRA^{+}15, BDB^{+}17, BRW15, BCRS19, CVG19, ETAG15, FPASS16, GLOP19, GS15c, KBD19, Lot18, MM17, SX16, SPM16, SZS15, Tav16, WSJY16, WBBC16, ZILZ15]. optimally [BMC18a, DJD^{+}17]. optimisation [BCO^{+}15, HKJ17, MH18b, MKV^{+}17, RPNP18]. Optimised [RSH^{+}17, LH17a]. Optimization [BZ18, DRP^{+}16, GHH15, RBD17, SGL18, ADE^{+}17, BABD16, BKS18, BMPS18, BDB^{+}17, CGC17, CWWZ17, CF19, DBD^{+}17, DK18a, DK18b, EFHZ17, Fid17, FBC^{+}16, FC16, GSS^{+}19, GJ18, GMA18, GGW17, KKZ15, KPKG15, LDO^{+}19, LLY15, Loz17, LBB^{+}17, MHJ15, MMMS15, NSL16, PPCK17, RN18a, RPC^{+}18, STHW17, SYI^{+}19, TZ16, TMH16, TD16b, Wal16, WHZ18, YYY^{+}16, ZP16, ZC19a, ZHW18]. Optimization-based [DRP^{+}16, RN18a]. Optimized [Bra16c, JLC18, DZR18, JLC15, JCWX19, KGS17, KÁGR18, LTXB17, MAvdW18, MRY19, SLL19, SZ17, YWHP15]. Optimizing [TLR16, CFO18]. orbit [SPCH16]. Orbital [LT17c, Fon16, HPV16, PDdG^{+}17, GS16]. Orbital-free [GS16]. orbital-updating [PDdG^{+}17]. orbitals [DLY17]. Order [BD18, BPD19, DS16, SYV17, TRM16, ÅN19, AHNF15, AD15, APP^{+}16, AGC19, AD17, AMJ17, AMP16, APLK19, ABFR16, ATF16, APKP16, And16, ADK^{+}17, ABH18, ABH^{+}19, ABR16, ABG18c, ARTG^{+}19, ALMJ15, Ata15, BZ19a, BTD16, BD15a, BGS16, BAGK16, BTGM17, BGTM18, BLM18, BNM15, BGN19, BMMP19, Bat17, BKO18, BCS19, BIR18, BM16, BR15a, BH18, BHE^{+}17, BZ18, BND16, BDMZ19, BFT17, BST15, BK16b, BQRX19, BDZ15, BDLM18, BB19, BSM16, Bre18, BPF^{+}16, BTT18, BMR19, BC16c, BCJ19, BCG^{+}15, BFTVC18, Cac15a, Cac15b, CGQ18, CBS18, CJK^{+}19, CC15, Cen19, CGMH18, CCK^{+}17, CKK18b, CNK19, CLY^{+}15, CLC16, CTG16, CJL16, CHY16, CWL^{+}16, CS17a, CZ17, CKT17, CSC19, CHJT17, CZL18, CB19, CLX15, CLTX15, CFST16, CGJ19, CXY19, CC16c, CR18, CG16, CCM17, CLP16b]. order [CWJ18, CCGH17, DS15b, DDJ19, DPO16, DBZ17, DNBH15, DC18b, DSX19, DWGW16, DL17, Die15, DLMDV18, DLC15, DM17b, DM19, Dom18, Don15b, DDH^{+}18, DCD^{+}18, DVP^{+}16, DLL^{+}17, DL18b, DL18c, DYL19, DY19a, DY19b, DPRZ16, DPRZ17, DKK15, EGO19, EDC19, EMSS17, Fal15, FS17a, FFM19, FAZ16, FWK17, FS18, FYZ^{+}15, FBM16, FP16, FRRV16, FHA16, FHA18, GP17, hGwSzS15, GZY16, GH17a, GFC18, GPS17a, GPS17b, GSL18, GS16, GAIN19, GGL^{+}17, GBCF15, GBCF16, GZLH19, GGT15, GEZK16, Gro18, GLK19, GY15, GLW18, GL17, HK19a, HW15a, pHzSrC15, HB16, HSLQ16, HTZG17, HBR15, HW16a, HU18, Hiv18, HN17a, HN17b, HN18, HHRA19, HF18, HLQ16, HH19, HHY16, HW16b, HC18a, HC18b, HJW19, Ism15, IDSG15, JLQX15, JPSX18, JZSX18, JYY18, JD19, JH17, JFS17, JTD16, JC17, KMD^{+}18, KC17a, KW15a, KVKS19, uKHGK19]. order [KRFV16, KYW^{+}16, KYW^{+}18, KFWK17, KC18, LMS17, LBTCG16, LK17, LN17, LSL15, LLP^{+}16, LAL18, LPW15, Ler15, Ler16, LX16, LHMB16, LMC16, LW17c, Li17, LHLL19, LWZ19, LC17a, LWWY18, LGB16, LZ17b, LS16c, LM19a, LM19b, Liu16, LW17d, LTXB17, LSZ18, LTW18, LIW18, LMB19, LLH19, LKSM17, LAK^{+}16, LWJV19, LLLN18, LP17b, LSI16, MLL19, MRM16, MZAF17, MCIGO19, MLM18, MNR17, MNG15b, MR16a, MA17, MRY19, MN15, MN16a, MRN16, MDHC15, MTT19, MKC17, MW16b, MP17, MDM^{+}15, MS18c, MP15b, MH18b, MRXI17, MA16, MMPS17, MB15, MSA19, MM16d, MM18, NMM15, NMC15, NJ15, NN15a, NL17, NRS19, OLDN17, O’S15b, OSKN18, OLHD17, OWKE16, OV17, PXLL16, PX16, PxRS17, PLHA18, PE16a, Pan20, PHRA16, PP18b, PAFT19, PE16b, PYAG19, PMB18, PBC^{+}17, QWX18, QPK19, RXSG15, RXS16]. order [RLGT19, RN18a, RSB16, RJ19b, RA17, RA19, Roy15, Rua18, RRMF^{+}19, RWN18, SZN19, SZN20, SLL19, Say17a, Say17b, SSVL18, Sch16b, SL18, SC18a, Sha17a, SM16, SWLZ15, SG19b, SLL16, SGC18b, Shu16, STG17, SYV14, SY18b, SFDE15, SLN15, Spe15, SPZ18, SSN15, SGT16, SGT17, Sti16, SC18b, SK18, SWZW19, Sub18, SWL19, TLQ15, Tao16, TLQ16, TLH15, TD18, TSH17, TK12, TK15b, Ter18, Tie16, TMH16, TMH18, TD16b, TKP16, Tso18, TR19, TYROG19, URL16, VPV^{+}17, VN15, VWV17, VLV19, VSM16a, VSM16b, Vil19, VAD17, VSC18, VLN^{+}18, VK16, WW15, WLM15, WXW15, WRL16a, WRL16b, WRPL17, WLGD18, WHR19, WKPS18, WSOW16, WKOE17, WR16, WSR15, WL17, WT15, Wu16, WTX17, XYF^{+}17, XDLX19, XY18, XJLQ15, XQ17, XYG19, YC17, YCPD15, Yan16b, YFJ17, YH17, YLD19, YDN19, YLA15]. order [YK18, YK19, YYJ^{+}19, YT19, YL19, ZP16, ZK15, ZL15b, ZC15, ZZZ17, Zha17c, ZLFW18, ZRW19, ZY19, ZCQ19, ZCQ20, ZzSK15, ZY17, ZW19, ZSO^{+}19, ZXL17, ZWYW18, ZQ16b, ZS17, ZQ17, ZS18, ZS19a, ZWUR16, dLDG^{+}18, dPSS16, BB19]. order-adaptive [Cen19]. order/low [CCK^{+}17]. ordering [XL17b]. orders [HCB19, PPCK17, VSM16a, VSM16b]. ordinary [CGS18, HBR15, MTK^{+}16]. ordinate [HHY15, JKE^{+}17, OWKE16, OKWE17]. ordinates [CFKK19, DMAM15, LFRH17, MRM16, Mas18, RW19]. organic [CLZZ19, vdKK16]. organised [LSP19]. orientation [HDF18]. orientation-independent [HDF18]. oriented [AMK17, BAD19, BD16, CPP19, TVB^{+}16, VLAB18, WW16]. origin [DK19]. Orthogonal [APLK19, MN18c, SNK18, AH15, BCSK17, BtTBI18, DA17, FBF15, LWZ19, NLW^{+}16, SLB^{+}19, TSST16]. orthogonal/dynamically [BCSK17]. orthotropic [XOX19]. Ortigueira [Kat16]. oscillating [KZR15, RXS16]. oscillation [APV^{+}18, PSS17]. oscillation-free [APV^{+}18]. oscillations [Bra16c, HZL^{+}15, MSG18a, SPP^{+}16a]. oscillators [SF16]. oscillatory [CDC17, CKK18b, DB18, GCI19, HWA15, HY16, IKS19, JS19, KCW17, LS15a, LJ16, MW16b, SK18, SS18c, SSZC19, Yan19, ZPW18, ZQ17]. OSIRIS [DTA^{+}15]. osmotic [YM17b]. other [CV15, JFS17, TMdO19, WS15b]. outer [LDB19]. outflow [KHHN16, NYD19, ST18a, YD18]. outflow/open [NYD19, YD18]. outflow/open-boundary [NYD19]. outlook [MSV^{+}16]. Output [Fid17, NP16, ZKS^{+}15]. Output-based [Fid17]. outputs [CPX19, VCNGP15]. over-penalized [SZ15b]. Over-Relaxation [AC16]. Overcoming [NMM17, CDOY19]. Overestimated [NF17]. overhang [ADE^{+}17]. overlap [SFP16]. overlapped [Sha17b]. overlapping [ABH18, AB15, DPO16, DY19a, HR17, MPFL16, MP19, XL16, XWZ^{+}18]. overloading [FS19a]. oversampling [SDW16]. Overset [MNC19, AB17, BSM16, CWJ18, HK16a, HM19a, HM19b, KLNH17, MKB19, SPB18, Vre17, Vre21, ZA15b, dLDG^{+}18]. overset-curvilinear [AB17]. overset/Yin [ZA15b]. Oxidation [PSN^{+}19, GMS16].

p [RRMF^{+}19]. p-adaptation [RRMF^{+}19]. PAC [ZR17]. PAC-MAN [ZR17]. packing [DFM17]. packings [KGT15, Zau16]. Padé [KM16b, SKO17]. Padé-type [SKO17]. Pages [Ano19-29, Ano19-28, Kat16]. Painlevé [FFW17]. pair [Mel18, RLH19, Zil15]. Paired [Ver19]. Pairwise [AMB17, LS16a, HLS19, LMZ19, WX17, TP16a]. Pairwise-interaction [AMB17]. pairwise-relaxing [LMZ19]. Palindromic [CSS17, HLTC18]. palm [ASB^{+}15]. panel [FS16]. panel-clustering [FS16]. panels [CPS17]. paper [Pan20]. parabolic [AHZ19, BSK15, BK18, EN18, EDvW17, GT18, GY15, LZ16, LMMS16, MBST17, OZ17, PE16b, SCS18, WN18]. paradigm [Cac15b, PD16a]. Parallel [BVMW16, BK18, BVT18, JBLO15, KJ17b, LZ18, LH15, MMSS15, MGBG16, NVBDV15, RPL^{+}18, SRBÓ17, WDS15, XML17, AM19, AG18, BLS15, BLK19, CWJ18, DFGQ16, DG16b, FHY^{+}19, HM19a, HM19b, Kas15, KJ18, LML^{+}16, LWL18, MGB^{+}18, MGPG19, MFF^{+}19, NN15b, PDdG^{+}17, PJE^{+}16, PP17, PBA^{+}15, PBCR19, PJB^{+}19, RS16a, RKO^{+}17b, SWJG19, SC16, SMOM^{+}17, SS18c, TPTT18, WLC15, WS15a, XZZ15, YM17a, YS17, YGJ18, ZYCK15, vdKK16, vdLJLV16]. parallel-in-time [BLK19]. parallelism [Sla16]. parallelization [JL18a].

Parallelized [KBK15a, GKRB17, OVP15, XWZ^{+}18]. parameter [ABT19, BK16b, CM19, CMH15, CMW16, FMPT18, GMS19, HKBR^{+}19, HXB15, ISP^{+}15, LYLK17, LVL18, MG15b, MNG15b, RBK19, SD17, ST15, YJM19]. parameter-free [CMH15, MG15b, MNG15b]. parameterization [RG15, VD16]. parameters [AABD15, CPT16, Don15b, GB15a, LBB^{+}17, NHM17, PKLS17, SC18a, SCE^{+}19]. Parametric [Gri15, Shi17, ATF16, BJWZ17, BH16b, BHS^{+}18, CS16b, HFND18, TMWF18, TT17a, YQNW19]. parametric/stochastic [HFND18]. Parametrization [GPS17a, GPS17b, BFI^{+}16]. parametrized [CMP19, CLTX15, NMA15, VCNP18]. Parareal [WZ18b, Wu16, WZ17, XHC15, NBH18]. parasitic [MC17]. paraxial [KLWQ17, SwS16]. Parcel [TASA19]. Parcel-number-density [TASA19]. parity [MJ17, WKOE17]. parity-mixed [MJ17]. Part [BGTM18, SZM19b, BN17, SLH18, SGD18, SHP^{+}16, TC15a, TKC15, BD15b, BTGM17, BHST17a, BHST17b, CK16a, DLN15, FNGV18, GPS17a, GPS17b, LB15, MBJ16, MBNJ16, MS18a, MS18b, PSN^{+}19, RRS19b, RRS19a, SZM19a, Say17a, Say17b, SPN^{+}19, VSM16a, VSM16b]. partial [ABDN19, AD17, ADH^{+}16, AEAM15, BZ15, BT15, CGS18, CSD19, CZ16, CM19, CC19b, DLL^{+}17, FBL17, Fal16, GXX17, HO15, JX15, JX17, KNS15, KR17, KS16c, LL16c, Lor19, MS16a, MTBT18, NBH18, Pes15, RK18, RPK19, RMP18, SR16, SS18b, Sub15, Sun19b, TST17, TO15, VCNGP15, VBG^{+}17b, WRL19, WAZ19, XY18, YHKPF17, ZTBW19, ZHWQ18]. partial-low-rank [WAZ19]. partially [MS15a, PD15, QWZ^{+}19a]. Particle [AB15, COdLL18, CLMZ17, FRO17, Gam15, GMWC19, KRK^{+}18, MDL16, PWC18a, TP16a, WZ18a, YLZ^{+}19, YG19, YDCK16, AMB17, AWS16, AF18, AP16, Ama18, BLL19, BLL20, BHdD18, BGRC19, BLK15, BBKS18, BAK19, BGHK19, BKKJ17, BLC^{+}17, Bra16a, BMR19, Cac15b, Cap18, CC19a, CGS15, CCL16, CLM15, Cos16, Cot18, CMR^{+}16, DD15, DTA^{+}15, DPK17, DJL^{+}19, DKC15, EH18, Eva18, GB15a, GMP16, GFA^{+}16, GG15, GAIN19, GBD^{+}15, GAJ15, HPC19, HWH^{+}16, Har18, HSLQ15, HSLQ16, HLS19, HM17, HW19, HM16b, HCLT19, HHL19, ID17, Iwa15, JG19, JLC15, JST17, JL19, KGS17, KES18, KF17, KK16, Kwo19, LKB15, Lap17, LN17, LPWK15, LS15b, LS16a, LLY15, LBTK18, LLW19, LS16c, LSR16, LMY^{+}19, LLB19, LWX19, LY17, LWTF19, MLM18, MRP^{+}15, MCW16, MC16, MPR^{+}18, MHZ^{+}15, MHT^{+}19]. particle [MS17, MMBP19, MBA19, MFG15, NOM^{+}17, NT15, OB19, PLL^{+}15a, PKP^{+}17, PCMC19, PR16c, PMF15, PWP15, PSV18, PC19, PvL19, RBJS15, SWC18, Sel15, SGC^{+}18a, SP16b, SMAG17, SCC19, SE15, Sto17, SPCH16, SGP17b, TWN19, TYD16, TZGW18, TUJ19, dCPDC^{+}17, TOR^{+}15, TASA19, TSFS17, TPB16, TL17, WVB19, WSN^{+}15, WNW^{+}19, WG19, WCCB16, YXW19, YXD^{+}16, ZB15, ZHA17b, ZZPH18b, ZZPH18a, ZWL^{+}19, ZZKF15, ZC19b, ZPE^{+}16, ZRE16, AG18, CT19, DDD17, FHA17a, MNO^{+}17, MSD^{+}17, WSJY16]. Particle-Averaged [TUJ19]. particle-based [GAIN19, LLW19, ZPE^{+}16]. particle-fluid [TZGW18]. Particle-in-Cell [CLMZ17, BLC^{+}17, Bra16a, CC19a, GFA^{+}16, HWH^{+}16, HLS19, MHZ^{+}15, PMF15, SPCH16, dCPDC^{+}17, WCCB16, YXD^{+}16, DDD17, MNO^{+}17, MSD^{+}17, AG18, CT19]. Particle-in-Cloud [WSJY16]. particle-ion [SCC19]. particle-ion/fluid-electron [SCC19]. particle-laden [AMB17, JL19, LWTF19, TUJ19, TASA19]. Particle-Mesh [PWC18a, BGRC19, MLM18, MCW16, RBJS15]. particle-number [BAK19]. particle-particle [LY17]. particle-resolved [CMR^{+}16]. particle/finite [PWP15]. particles [CLM15, DM17a, DSH^{+}16, Fan19, JdR^{+}18, KYKS19, KK17b, LMY^{+}19, LHW^{+}17, NLFM16, OADN19, RFGSV15, SK19a, SKF15, SCL19, SGC^{+}17, Tao16, TP16b, TKF17, WSP17, YC15, Yan16a, aKT16]. particles-fluid [WSP17]. particulate [KLNH17, KSI17, LRZ17, MZ15, WSP17, WTL19, Zoh17]. Partition [FLT18, BHKS16, BMPS18, MIM^{+}19, NJHL18]. Partitioned [CLW18, LPB17, WED15, BHST17a, BHST17b, BHST18, BCM15b, DDV18, LLEK17, LHB^{+}16, LLJJ18, MBHS17, SBHS19, Sla16, ZSM19]. partitioning [FLHA17, LK17, NSB15]. parts [CHD^{+}18, DBZ17, DCCH19, FN17, GWK16, LMN18, MN04, MN17, NN17, NN19, NR17, NG17, NG18, PS15b, RÖS16, RÖS17, Ran18, RWN18, RN18b, SPB18, LKN17]. Pascal [LY16a]. Pasquetti [Sir19]. passage [PTMF18]. Passing [CDX^{+}18a]. passive [CDX18b, HM17, LE16]. past [CPS17, SHW17, SSZC19]. patch [BRK17, GFA^{+}16]. patch-based [GFA^{+}16]. patching [BVS18]. Path [HKKP16, KKZ15, XZJK19, CC17b, Cot16, Gen15, LO16, Opp17, SV17, Zil15]. Path-space [HKKP16]. paths [LB17]. pathways [HHZZ19]. patient [BFI^{+}16, ISP^{+}15, PVFN15]. patient-specific [BFI^{+}16, ISP^{+}15, PVFN15]. pattern [AEAM15, MDH19, SPM16, ZYW16]. patterned [PKB15]. patterning [DMRB19]. patterns [HHZZ19, SFDE15]. Pauli [RMC15]. PAW [STHW17]. PB3D [WSH^{+}17]. PC [HD18, RVK^{+}18, WHRL19]. PC-EnKF [WHRL19]. PC-SAFT [RVK^{+}18]. PCA [VD16]. PCE [PES19]. PCFE [CC17a]. PCM [LFR17]. PDE [AGC19, BSK15, CFG16, CNOS15, CVG19, GBR15, LLD19, MKSP19, PYK19, SGT16, TD16b, VBG^{+}17b, ZYK18]. PDE-based [BSK15, VBG^{+}17b]. PDE-constrained [CVG19, TD16b]. PDE-domain [ZYK18]. PDE-Net [LLD19]. PDE/ODE [CFG16]. PDEs [KHP15, Kat16, GH19, AW16, AL19b, BST19, BFFB17, BN19, BK18, BVT18, BGHK19, LL17, CM15, CNW17, CELZ18, CLP16a, DL17, HL15b, HHLY17, LW15b, LLD19, LTR17, MN18a, MJ17, NVBDV15, NNV19, NRS19, OZ17, PRXC19, PR16c, PLR18, PLPR19, SNK18, Shu16, SGL17, SK19c, TSC17, UHKT19, Wu16, ZILZ15]. PDF [MGT18]. Peaceman [SwS16]. peaked [FYC^{+}18]. pebble [LL18]. PEC [HGR16]. peeling [WSH^{+}17]. peeling-ballooning [WSH^{+}17]. peer [LH17a, SLH18]. penalization [EKSS15, GWC17, HKLW15, MLL18, SYOS19, SYOS21, SHW17, TK15a, TNB21]. penalized [SZ15b]. penalty [CM16a, Fer17, GSN16, LFAR19, OKE17, DM18, QN19, MRRRF18]. penalty-projection [CM16a]. penetrable [APKP16, LO19, PA19]. Peng [KS18b]. percolation [CY19b, LL18]. Perfect [WWRS17, CC19a, FSB16]. perfect-conductor [CC19a]. Perfectly [AL19b, DZR18, SJH^{+}15, BJK17, Che19, CMH15, DKK15, GTL18, Par15, Par17, Par18b, PD16b, SF18a, dCMR19, AL19a, CJH^{+}19, DCCC16]. perform [Bra16c]. Performance [Par17, ZWG17, BJRF18, CL19b, DLN15, ETL17, Ein19, HAPK15, MC15, PKA^{+}16, RGPS17, SY17, Shi19]. Peridynamic [XZT18]. Periodic [LSM19, BM19b, BB15, Cho19, DCA^{+}16, DJV^{+}18, DKTH15, GS16, GD19, HTFL18, HN17a, HN18, LKB15, LB16, MVZ16, NL15, RPC^{+}18, ST17, SHW18, SWZ17, WZLS19, YS18b, ZHS18, Pan15]. periodically [SC18a]. periodicity [NPP15, YS18a]. permeabilities [MTK15]. permeability [BDMC15, JCNK19, LFAR19, MTK17]. persistent [WW16]. perturbation [CDM^{+}16, FCW^{+}18, GFvR18, HN17a, HN17b, HN18, KVKS19, SY17, SKW19, UG16, WHCN17, YM17c]. perturbation-method-based [CDM^{+}16]. perturbative [Fal16]. perturbed [BCM19, CAA18, GRMK15, GFvR18, SP15a, ZG18b]. Peshkov [Jac17a]. petroleum [TH18, YSLY19]. Petrov [BK19b, CBA17, GR18, RDM15, SDW16, WZLS19]. Petviashvili [EO15]. Phase [BG16b, CCP19, HW15c, LJZ15, ZW16, ARF18, ACGR15, AASRT17, ABG^{+}15, ASWvD19, Ani16, AT18, AB15, BCST17, BGN15, BT19, BM19a, BAVC17, BGJ^{+}15, BR17, BHMS18, BDPM18, BKG15, BKKRB16, CFSN18, CDM18, CGK17, CJYZ15, CS16c, CY19a, CKQT15, CYS17, Chi19, Cif19, CS17b, CG16, CM18d, DD16a, DD15, DG18, DLP19, DGMT17, DW19, EHXM15, FGL16, FB17, FMRZ17, Fed17, FPT17, FZ19, FS19a, GGL^{+}17, GHL^{+}16, GGT18, Hag25, HHA15, HHM17, HBR15, HSB16, HTMP17, HHRA19, HCLT19, HW16c, HLA19, HTBG15, JTR16, JS16, JS17, JJ18a, JJ18b, KJ17a, KS16c, KS18b, LVTR15, LSL15, LRA17, LW18, LYM19, LPGT16, LWY17, LSD^{+}17, LY16c, LLB19, LDGH16, LY17, MNG15a, MA19, MT18, MT19, MN16b, MAA18, MDP18, NPRC15, NS19a, NWB19, NLW^{+}16, OTS17, OT15, PL18, Pan20]. phase [PSB^{+}18, PM19, PKB15, PS14, PS15a, PGM17, RWG18, Ren19, RV16, RTO15, RZ15, SPX^{+}18, SHA16, Sha17a, SRBB18, SYY15, SLL16, SA19, Sun19a, Suz18, TH18, TK15a, TGS19, TND18, TASA19, TT16, VS17, VSC18, WJD16, WDT^{+}19, Wic16, WKSS15, WT16, WHZ18, XSL18, Yan16b, YH17, YSY17, YD19, YY17, ZZ17b, ZN18, ZHLZ18, ZWL^{+}19, ZKG19, ZCY^{+}19, ZYSW16, ZYCK15, dJRP^{+}15, tEDKT17, KYKS19, SRS19]. phase-based [NPRC15]. phase-dependent [DD16a]. Phase-field [BG16b, ARF18, BDPM18, CJYZ15, CS16c, CKQT15, DW19, GGT18, Hag25, JTR16, JJ18a, JJ18b, LWY17, LY16c, MAA18, NS19a, OTS17, PKB15, SLL16, WDT^{+}19, Wic16, WHZ18, YY17, ZHLZ18, ZYSW16, ZYCK15]. phase-field-lattice [RTO15]. phase-transition [DLP19]. phased [SFDE15]. phaseless [ZZ17a]. phenomena [Fon16, HJY19, LAK^{+}16, RSH^{+}17, SCN^{+}17]. phenomenon [sCYxL^{+}18, LBZA16, Rod17, Rod18, VBG^{+}15]. Phi [SGL18]. phonon [GW16, RW19]. phononic [DBD^{+}17, WZLS19, ZZW^{+}16]. photoelectrochemical [HGR16]. photon [BCG^{+}15]. photonic [CHL^{+}19, GYZ19, MHJ15, WHZ18]. photosynthetic [Pis18]. Phys [ABG^{+}19b, ASS17, BLL20, CNG17, Dav15, DK18a, DvW19, EFO20, GRT21, Gho17, GBCF16, Hag25, HGN17a, KYW^{+}18, KTK19, MN17, NG18, PS15a, SZN20, SYOS21, STEK22, SWMD17a, SYV17, TK15b, Vre21, ZJS15, ZCQ20]. Physalis [SP16b]. Physical [Don15b, HK19b, LDT19, CZ19a, CHZ16, CWB^{+}19, HX16, LS16b, LLY18, MBA19, SAH17, SCE^{+}19, WT15, WT16, XB18]. physical-based [LLY18]. Physical-constraints-preserving [LDT19, WT15]. Physical-density [HK19b]. physically [HKS^{+}16, PA15, WED15]. Physics [BR16, EH15, HSK^{+}15, Kat16, Pan20, RPK19, XS15, YBSTT19, YB17, ZZKP19, CSCM16, DD17b, FHA17a, GV18, GLG^{+}19, GK19c, GSMR18, HBC^{+}16, HHCG15, MN18b, MBS19, PT18, PRvdL18, RK18, SBG^{+}17, SM19b, WNDB19, XWW^{+}16, YP19, YDN19, ZLGK19, ZR17, dFGS^{+}17]. physics-aware [GK19c]. Physics-based [YB17, GLG^{+}19]. Physics-constrained [ZZKP19]. Physics-informed [RPK19, YBSTT19, XWW^{+}16, YP19, ZLGK19]. physics-motivated [FHA17a]. physiological [BC16a]. PIC [Bar18, CC16a, DTA^{+}15, JL18a, MC16, MAM16, SWHK15]. piece [LTKA15]. piece-wise [LTKA15]. Piecewise [VDPP15, BMPS18, BG19c, MNR17, SUR18, Str18, TD16b, YL16, ZC18, ZC19a, LFR17]. piezoelectric [MDD^{+}19, YG18]. pigments [Pis18]. pin [CLL17]. pin-based [CLL17]. pinch [WRL18]. pins [PBA^{+}15]. pipe [PGGW18]. pipelines [BLVC16, DG18]. PISO [HWK19, NSK^{+}16]. Pitaevskii [ATZ16, ABR16, MBM^{+}15]. pivoted [HKLZ18]. placement [ABdC^{+}18, NHM17]. planar [HK18b]. Planck [HYK^{+}16, KJ17b, KJ18, TC15a, TKC15, CM18b, CCL16, DWZ19, FLT17, GJ15, GAJ15, JG19, LW17e, MS18d, QWZ19b, SV17, SK15b, TCSM15, TCS16a, TCS17]. Plane [ZCZ19, CDL19, CV15, DBD^{+}17, IG15, KPP^{+}19, OLHD17, PDdG^{+}17, RYZ18, TBB^{+}19, ZZW^{+}16]. plane-wave [CDL19, PDdG^{+}17]. planet [LML^{+}16]. planet-plasma [LML^{+}16]. planetary [BLC^{+}17]. planewave [CDM^{+}16, PUA^{+}15]. planewave-based [PUA^{+}15]. plasma [Ama15, Ama18, AG18, BZ16a, BJK17, Bra16a, DS15a, DD17b, DNOP15, DCD^{+}18, ESGS17, FH17, GDS^{+}16, GBC^{+}18, GSMR18, HYK^{+}16, HWH^{+}16, HR17, HLS19, HTvdH^{+}19, IKI15, JL18a, KHTZA16, KB18, KS18a, KHC^{+}16, LML^{+}16, Mar19, MCS^{+}19, MP15b, MP16, PSV18, RKH15, SZ17, SS16c, SR18, SJH^{+}15, TMWF18, TC15d, TSR15, Yan16a, YDN19, ZZDB15, GFA^{+}16, MAM16]. plasma-coupled [TMWF18]. plasma-lunar [HWH^{+}16]. plasma-vacuum [SR18]. plasmas [CBB16, Hig17, JHT^{+}18, KKS15, KKS16, LLD^{+}16, MNO^{+}17, MSF^{+}19, PMS15, PC16, SP16c, TBC^{+}16, VSC18, WCCB16, ZG17, dSPDH15]. plasmon [MML17]. plasmonic [EDW19, Fuj19, VCNP18]. plastic [CHJT17, GSL18, Heu17, Heu19, JN19, KTK15, KDL15, WTL17]. plasticity [NS19a]. platelets [ZZDB15]. plates [BBMN18, PWC18b]. platform [NRZS17, TC15d]. PLIC [LHGF16, SGD18]. PLS [PES19]. PLS-based [PES19]. plume [JL18a]. PML [FYZ^{+}15, PKN17, WSOW16]. PML-truncated [PKN17]. PN [Bou19]. PNP [MBBKTH17]. PNPM [BK19a]. POD [DA17, BFI^{+}16, BFI^{+}18, BIR18, BCG^{+}15, FMPT18, LHLL19, MS16b, RTV17, SSN15, URL16]. POD-based [LHLL19]. POD-Galerkin [BFI^{+}16]. POD/DEIM [SSN15]. POD/kriging [MS16b]. Poincaré [DDV^{+}15, HS17a]. Point [RHS18, AEL^{+}15a, AEL^{+}15b, AMB17, BLL19, BLL20, CSY19, CR18, DDJ17, DZ16, DZ18, EH18, GS15a, HM16b, HSC16, JL18b, KR17, LLB19, MBA19, MWB^{+}15a, MWB^{+}15b, PJC16, POSB16, PR16c, PLR18, PLPR19, WG19, WHEK18, WX18, XM18, ZZS^{+}17, ZD17, Vog17]. point-centered [MWB^{+}15a, MWB^{+}15b]. point-kinetic [DDJ17]. point-mass [WG19]. point-particle [AMB17, BLL19, BLL20, EH18, HM16b, LLB19, MBA19]. point-value [XM18]. points [DZ16, LTXB17, LWTF19, MRRRF18, Teu15, WN17, XL17a]. pointwise [CLL17]. Poisson [ZG17, And16, ABG18c, AC17, BLA^{+}15, Bat17, BPTA16, BG19c, BD16, CGQ18, CCZ18, CG18a, Cot18, CLMZ17, DWZ19, DBMB15, DvWZ18, EG17, EL17, EG18b, FFM19, GWC18, GN19b, HW16a, HF18, JTD16, LY15a, LW17e, Liu19b, MS18a, MDVM16, MNR17, MHH19, NN15b, PKJ^{+}18, RBI18, SH19, SMLB15, SC16, SHW18, Sti16, Tow18, VSC18, WSJY16, WW18, XJ16, YM15, YM17c]. Poisson-like [NN15b]. Poisson-type [BG19c]. polar [AM17b, RBJS15, VBL^{+}16]. polarized [ZND16]. pole [TLH15]. pole-treatment [TLH15]. Pollutant [VST16, PHRA16]. polycrystalline [JTR16]. polydisperse [BVM17b, DAO17, FITcD19, LMY^{+}19, SDM^{+}17, WB16, Zoh17]. polygonal [BMPS18, CSY19, CFF18, HR18a, LM19a, LM19b, QYJ19, Rag15, SDW18, TC15b, ZSW17]. polyhedra [BHP19]. polyhedral [IM15, MDD^{+}19, WHY17]. polyhedron [CZJ17]. polyhedrons [CZ19b]. polymer [Cen19, DD17a, KBK15a, OLB^{+}17, SB18]. polymeric [MS18d, SK15b]. polymers [ADFG17, OLD^{+}16]. Polynomial [ABM16, KS16b, LB15, SG17, ARG^{+}17, ATM^{+}18, AM18, BSN19, GMS19, GNZ18, GLZ19, HD15, HD18, JES15, KSV^{+}15, LW17c, LMBZ15, LGB16, LY16a, LMTC15, NDH19, OB17, PUA^{+}15, PHD16, PSMPG17, SS17b, SL18, SX16, TCA16, TWF19, TE19, TMES19, THS^{+}19, WMM^{+}18, YZ19, YLBL16, YY16, ZNX15, CJC19]. polynomials [BFFB17, BFF19, CGJY19, FFBB16, KKJB16, O’S15b, OTS17, Str18, VBL^{+}16]. polytopal [DDM18]. pool [SN15, WNDB19]. poorly [Bra16c]. population [BAK19, LPWK15, NLFM16, WYA^{+}17b, XZZ15, Yas17, PMF^{+}18]. pore [GMT19, KEJ18, MT18, MT19, ZHWQ18, ZQCT15]. pore-scale [GMT19, MT18]. pores [GHP15]. poro [FFJ^{+}19]. poro-mechanics [FFJ^{+}19]. poroelastic [AEVW18, Buk16, CM18d, DGMT17, HMFJ18, PWC18b, RSBS19, SCLG15, WLE17, XOX19]. poroelasticity [FAC^{+}19, SBH19]. poromechanics [CWF16, DGW18]. porosity [GCVCHH18, JCNK19, OADN19, dFGS^{+}17]. porosity/permeability [JCNK19]. Porous [MSG18a, MSG18b, ABI17, BDMC15, BPS17, BHMS18, BKKRB16, CXY19, CS17b, CLNH15, CvKH16, CFvKH18, DBD^{+}17, DCA^{+}16, DLWY19, FPT17, GAS^{+}18, HSK^{+}15, JT18, JS16, KJ17a, LW18, LH15, LT15, LJ19, LZT17, LFAR19, LNM15, LRGO18, MCN18, MTZ16, ML16, NH17, NSK^{+}16, PF16, PSN^{+}19, QM18, RdM19, RJLW19, SSL17, SPX^{+}18, SWML17, SMT^{+}16, STG17, SPN^{+}19, TWH15, TAH16, UHKT19, VS17, Vos17, WC18, XML17, YSY17, YB17, Zad11, ZZ17b, ZOY^{+}19, dMRHJ17]. port [KML18, TRLK18]. port-Hamiltonian [KML18, TRLK18]. posed [NN19, NN16, NL18b, PND16, dlHC16]. posedness [GMD19, IG15]. position [CCM15, HDF18]. position- [HDF18]. positioning [CZ19b]. positive [CHS17, DWGW16]. Positivity [DWZ19, GY15, MKC17, Par18a, VSM16a, VSM16b, CLTX15, CFST16, JJ17, JJ18a, JJ18b, LAEK18, PAL^{+}16, QWZ19b, RJ19a, SY16, SPZ18, SDW18, WHY18, WMYG16, Zha17c, ZSW17, ZCQ19, ZCQ20]. Positivity-preserving [MKC17, Par18a, VSM16a, VSM16b, CLTX15, CFST16, SPZ18, SDW18, WHY18, WMYG16, Zha17c, ZSW17, ZCQ19, ZCQ20]. possibility [RRL19]. possible [Ant17]. possibly [LL17]. Post [BBW16, CDM^{+}16, MBM^{+}18, KH17]. Post-processing [BBW16, CDM^{+}16, MBM^{+}18]. posterior [CMW16, MTM19]. posteriori [AGRB18, BD17, CNOS15, DS16, DL16, JW15a, KH17, SL18, Vil19]. posteriori-driven [DS16]. potential [BDMC15, CCHL15, CX16, Gen15, HLU15, JGS16, LLEK17, LC19, MVKD15, Moc17, PMF15, RLH19, SMOM^{+}17, SE15, SD16, TRM16, WH16b, YTW15, YT19, ZRT18]. potentials [BTT18, CKK18a, DH18b, DOO17, EMZ16, FS15, IKS19, MTD15, PAFT19, RBGV15, ST18b, SAF^{+}19, SWLW19, SCS18, TST^{+}15, YZW^{+}18]. Potts [PRvdL18]. powder [Zoh17]. power [CT15, CSCM16, HED^{+}16, ZLL16a, ZLL16b]. powers [Vab15]. Poynting [LZL^{+}19]. Poynting-vector [LZL^{+}19]. PPML [KV16]. Prabhakar [MG15a]. Practical [HFND18, BDB^{+}17, VBF15]. Practice [SSZ19, LLL16]. Prager [LEB^{+}17]. pre [PPLC16]. pre-conditioned [PPLC16]. precision [AM19]. precomputable [YD19]. Preconditioned [FSWW17, HP17, LTR17, Pea15, ZJL16, ALT17, HYL17, JW15c, LSCC19, SP19a, SYM15, TWM18, VYP15, YSY17, YGJ18, ZZ19]. preconditioner [AAE17, AAE19, BST^{+}18, CLZ18, DFGQ16, DDV^{+}15, EG16, Kas15, KC17c, LY19, LY16d, RJLW19, SLR^{+}16, TCD17, YFC19]. preconditioners [BMMP19, BVMW16, DM17b, DM19, DMSC16, KCW17, MHHX16, MDDM17, PP18b]. Preconditioning [WNDB19, AMXJ19, CG18a, FFJ^{+}19, HB15b, JTD16, KA18, KPP^{+}19, LM19c, PKJ^{+}18, RM16, XLY15, YM17c]. precursor [KS18a]. predict [DCA^{+}16, YL16]. predicted [CZ19b]. predicted-Newton [CZ19b]. predicting [AEAM15, CSG17, CLG^{+}19, KL18b, LYL19]. Prediction [CP17, DJV^{+}18, BHGK18, BBK19, BHFB19, Eva18, FS17a, IPSG15, NP16, PVPK17, RKN19, TMWF18]. predictions [ALM^{+}17, ID17, KBF17]. Predictive [KZ17, SZK17, CSCM16, KL17a, KZG16, MGKG17, OCSC18, PD16a, RKN19]. predictor [BK16a, Jac17b, PHRA16]. predictor-corrector [PHRA16]. predictors [PSMPG17]. Preface [PC16]. prefactored [RSH^{+}17]. premixed [SWS^{+}18]. prescribed [CRMP16, EJZ17]. presence [BTA17, FP18, GGW17, LT15, NL15, PMS19, RTO15, WS15b]. Preservation [CHZ16, AHNF15, BTGM17, BGTM18, LCF16, MHH19, OV17, PAL^{+}16, RJ19a, VW16]. preserved [LYL19]. Preserving [DD17b, AS15, ADK^{+}17, BK19a, BM19a, BLMY17, BT16, BMC^{+}18b, CZW17, CBZ18, CJWS19, CCBdL15, CCdL15, CCP19, CWS18, CX15, CQL^{+}17, CDN17, CLTX15, CFST16, CEHM19, CXY19, CDV17, DLM18, DvW15b, DL15, DWZ19, DMSC16, DY19a, DY19b, DMTB15, EHXM15, FG17, FLT17, GJL19, GLTB18, GWYS18, GY15, HCB19, HSLQ15, HSLQ16, HDA^{+}18, Hiv18, HW15b, HS18b, JLQX15, JL18c, JXZ15, JL17c, JS17, JJ17, JJ18a, JJ18b, KTK18, KTK19, LW15b, LZW19, LDT19, LY16b, LAEK18, LCK19, Liu19b, Loh17, LHL15, LP17b, MSK18, MD17, MKC17, MGCW18, Nis15, NL17, NSB15, Par18a, QWZ19b, QDRB15, QSY16, RKH15, SS18a, Sch16b, SY16, SL15, SKO18, SPZ18, SL16c, SDW18, SJX15, SJXL15, Suz18, TCS17, TW17, TRLK18, Vel19, VSM16a, VSM16b, WZ15, WY16, WW18, WHY18, WW19, WMYG16, WKOE17, WT15, XJLQ15, XYG19, YJ17]. preserving [YSLY19, YWHP15, ZLJ16, Zha17c, ZSW17, ZY19, ZCQ19, ZCQ20, ZH19, DDD17]. Pressure [DXvW18, DvW19, LR19b, AEL^{+}17, ALL18, Cif19, CLNH15, DWGW16, DS15c, HTFL18, Hig15, HHA16, KTN15, KHHN16, LW18, MS15b, MCN18, MBD19, NF17, RMF^{+}18, STHW17, SS17c, SiI16, TD17, Tou18, XDvW17, XDSX17, ZCHS15, ZZ17b, ZKG19]. Pressure-based [DXvW18, DvW19, TD17, XDvW17]. pressure-corrected [RMF^{+}18]. pressure-correction [ALL18, MBD19]. pressure-density-based [XDSX17]. pressure-dependent [MCN18]. Pressure-induced [LR19b]. pressures [TK12, TK15b]. pressurized [CLB^{+}16]. Preventing [HZL^{+}15]. Primal [RB15, AAE17, AGRB18, Eng18, TC15b]. primal-dual [TC15b]. Primal-mixed [RB15]. primitive [Niu16]. principle [ADK^{+}17, CHY16, CLTX15, DY19a, RMC15, SWPS17, WYZZ18, YL19]. principles [AZK16, FPDT17, FPV18, MN16c]. prior [KKL15, dFGS^{+}17]. priori [BAD19, FAZ16, KF15]. priors [TBLM15]. prism [CLFL17]. probabilistic [GSS^{+}19, GK19c, LZ18, MCS16, SGS^{+}19]. probability [BVM^{+}17a, BC16b, CVK16, DH18a, GHL15, KMP^{+}19, PKW17, RC18, SG16]. probing [PKN17]. problem [AJP15, ABN15, AB19, Alm19, ADP^{+}17, BD15b, BN17, BLM18, BHST17a, BXY17, Bat17, BtTBI18, BDKK17, BD16, BKL17, Cac15b, CCHL15, CC15, CCBF19, CGM15, DGHP17, DvWZ18, FPV18, GPRA18, GP17, Gro18, GP16b, GY17, GY18, HK18b, IKI15, JPLL15, JTD16, uKHGK19, LL19a, LDOK17, LHS^{+}18, LW15a, LMC16, LYPP17, LZW^{+}17, MST15, NDH19, NKN^{+}17, PP19, QN19, RMA17, RZ18, SF18a, TFGK18, TMT17, TV19, TM15b, VW18, WZ15, WL18, WHR19, Xia15, ZF18, CTM^{+}16, RZ15]. Problems [GIF18, LBTCG16, APP^{+}16, ADN19, AB16b, ATF16, ACJ17, Ani19, AHZ19, AR16b, AWJ17, Azi19, BJO18, BCSK17, BSK15, BK17a, BABD16, BJWZ17, BOA17, Bar19, BDB18, BQCG17, BGV17, VMN^{+}18, BS19a, BGL^{+}17, BKR19, BGHK19, BFT17, BCB17, BM19b, BCG^{+}15, BCM15b, BKRB15, CPV16, CDL17, CMP19, CXX16, CH17, CWW17, COV18, CR18, CG18b, CFF18, CGM15, CM16b, CMW16, De18, DPW^{+}15, DCCC16, DL17, DHH^{+}18, Die15, DAO17, DJL^{+}19, DLS15, DZC16, ETL17, EZG16, EN18, EMS^{+}19, EE16, FFJ^{+}19, FPDT17, FK17, FG18, FL16, FCL19, FP18, GM19, GS15c, GWC18, GAIN19, GFvR18, GQC^{+}19, GWE^{+}15, GLTG15, GHH^{+}16, HPC19, HLL^{+}16, HU18, HHCG15, HN17b, HMFJ18, HLTC18, HDF18, IPSG15, IC17, JHPAT17, JL18b, JJ19, JFS17, KVKS19, KS15a, KA18, KW15b, KADE15, KKLS17, KE15]. problems [KCW17, KR17, KFL17, Kou16, LSLA16, LPU18, LW17a, LZ16, LW17c, LJ19, LWZ19, LL19b, LGB16, LMMS16, LM19b, LY16a, LHW^{+}17, Lot18, LMTC15, LSCC19, MZAF17, MS18a, MIM^{+}19, MNR17, MPR^{+}18, MM15, MBS19, MMMS15, MM16c, MTK17, MBD15, MDD^{+}19, MSP16, MFF^{+}19, MW15, MWYZ16, NS19a, NYD19, NN15a, NL18a, NGS16, NL18b, NSL16, NN15b, O’S19, PKLC16, PKLC17, PHHR17, PNZ18, PLC18, PM19, Par15, PCMC19, Pea15, Pes15, POSB16, PTT18, PDN19, PBA^{+}15, PBKK17, PKK18, PGH15, RPK19, RYZ18, RHS18, RSBS19, RRD16, RBGV15, RPC^{+}18, SY17, SYI^{+}19, STEK17, STEK22, STFK19, SSC^{+}16, SPB18, Sir19, Sla16, SL19c, SMLB15, SDW16, SP15b, SPM16, SDW18, SCS18, TBHG18, TV19, TMS^{+}19, Tow18, VMC^{+}19, WHY18, WZLS19, WWRS17, WED15, WSF17, XZ15, XZT18, YK15, YZ19, YL19, ZP16, ZILZ15]. problems [Zha16, ZLL16a, ZGJ16, ZSW17, ZG18b, ZCY^{+}19, ZZYC19, ZLGK19, ZVO15, ZSO^{+}19, ZCZ19, ZZT^{+}16, ZHW18, ZG19, dFJN16]. procedure [BBKS16, EH14, EH15, ED16, GB15a, GN19a, JLC15, JJ18a, KKL15, LFT^{+}16, MRK15, PK16, RÖS16, SW15, XS15, dFVJ15]. procedures [DLP19, HHM17, OS16]. process [AZ19b, CLFL17, DG16c, DW19, ISST18, IM17a, JdR^{+}18, PYK19, PVPK17, PSMPG17, WLL16, WYA^{+}17a, YBSTT19, ZKS^{+}15]. process-based [PSMPG17]. processes [ADH^{+}16, AJVH17, Beg15, CZ19a, CZB15, CMR^{+}16, GLG^{+}19, LP16a, Luc15, MZ15, RPK17b, RKRGW17, SSL17, TBG16, YSWW16, ZQCT15, Zoh17]. Processing [GP18, BBW16, BDMZ19, CDM^{+}16, GK19b, MBM^{+}18]. processors [AAB^{+}16, SGL18]. produce [WDG^{+}17]. product [CLZ18, DM17b, DM19, PP18b]. production [Bra16b]. products [CC17b, CSY15]. profile [ZS15]. profiles [NMM19, WG16a]. Progress [TM17]. projected [VYP15]. Projection [dFJN16, ALL18, BTD16, BVG^{+}16, BPF^{+}16, BFTVC18, CBA17, CM16a, CDM18, CDM19, CBN^{+}16, FWK17, FMPT18, FSB16, GMP15, GTG15, HTBG15, LTB16a, MO18a, PKLC16, PKLC17, PLC18, RSB15, TG17, UL16, WE15, WYA^{+}17a, WYA^{+}17b, CGK17]. projection-based [BTD16]. projective [LMS17]. prolate [HXB15, LC18]. Promising [KJHA19, ASB^{+}15]. proof [EBQ15]. propagation [BS19a, BG19b, CCFC19, CGMH18, CZB15, Chu17, CLQ17, CJH^{+}19, DSS18, FKR16, GFG^{+}15, Hiv18, HL16b, IPSG15, Kim15, KLRT15, LCLY19, LTB16b, LMM17, MP15a, MSS16, Mue18, MNW19, MSH^{+}15, MH17, PS15b, POSB16, RA17, RA19, RSH^{+}17, SCN^{+}17, STEK17, STEK22, SS17b, TLB^{+}18, TEP19, TBG16, VSDW18, WLE17, XOX19, XYG19, ZLL17a, ZZW^{+}16, ZWUR16]. propelled [SCQP16]. propelling [BGRC19]. Proper [APLK19, LWZ19, TSST16, DA17, KPKG15, SLB^{+}19]. Properties [LP17b, AF18, CCRdL17, CHZ16, CFSZ19, GO15, GDFL17, Ism15, JL18a, KGS17, LPG18, LJT16, LGZ^{+}19, MT17, NR17, PWP15, RdM19, SLL19, VV16, VLTPS16, Wil18]. Properties-preserving [LP17b]. property [FHA17b, GWK16, LXC19, NDH19, SZ15b, WY19, WCWY19]. proposal [KGS17]. propulsion [Har18, Moo17]. prospective [CVM^{+}19]. protection [PYAG19, PSN^{+}19, SPN^{+}19]. protein [XR17]. prototype [SSC^{+}16]. provably [Tie18]. pseudo [BCSK17, GWWC17, HLU15, KW15b, KADE17, LWJV19, RN18b, VLV19, WS15a, dlHC16]. pseudo-compressible [WS15a]. pseudo-convergence [KW15b]. pseudo-inverse [BCSK17]. pseudo-potential [HLU15]. pseudo-spectral [GWWC17, KADE17, RN18b, dlHC16]. pseudo-time [LWJV19]. pseudopotential [HW16b]. pseudospectra [RS18]. Pseudospectral [NGY^{+}17, AL19a, HXB15, MH17, PCA19, RRD19]. pseudospectral/discontinuous [MH17]. pulmonary [FK19]. pulsating [LZHM19]. pulse [DHC16, TT19, ZX19]. pulses [TSN16]. pure [YT19]. purely [Cap18, YJ17]. Purkinje [PVFN15, VLP^{+}16]. purpose [AVT17]. pursued [TK16]. pyramid [WHY17, WHY18]. pyramidal [JG15]. pyrolysis [PSN^{+}19].

QTT [BDKK17, BKKY19]. Quad [GTG15]. Quad/Octrees [GTG15]. Quadratic [HR18a, SOS19, CCWY18, HLTC18, LMB19, MHH19, NYD19, TD16b, WT16, XX17]. quadratization [YZW17]. Quadrature [PMF^{+}18, Tsa15, ZGJ16, CFKK19, CS17a, EE16, FLV18, GN16, LLH19, Nis15, NL17, PPM^{+}19, RF18, SS17a, ST18b, Spe15, SGP17b, XX17, ZSP15, aKT16, RKO17a, WK18, WK19]. quadrature-based [PPM^{+}19]. Quadratures [TS18, AS16, Tsa16, ZNX15]. quadrilateral [LTW18, MKSP19, Rag15, ZPW18]. quadrilaterals [EL19]. quadtree [MGB^{+}18, Pop15]. quadtree-adaptive [Pop15]. quadtrees [Bat17]. quality [FAZ16, KF15, WQZ15]. Quantification [KBK15b, AKZ16, AÁPB17, CPX19, CC17a, CE18, CQ15, DP19, DH18b, GS18, HAPK15, KRBW17, KCS^{+}17, LS15c, LLL16, LSD18, MBNJ16, MS16b, MSS16, MCS16, PTT19, RS17, TE19, TB18, TO19, WL16, WSK19, YP19, ZZ18, ZZKP19, vdBKD17]. quantified [WRL19]. quantify [CELI15].

Quantifying [AZK16, BHJ15, GZ19, XWW^{+}16, ZLGK19, GKRB17, LSS16]. quantitative [KBF17, OTS17]. quantities [Loh17, STR15]. quantity [CCHL15]. quantized [CVG18, RKN19]. quantum [CLY^{+}15, CSC19, GP18, LM15d, MFB18, PUA^{+}15, PD16b, Sel15, SCL19, SDFA17, TST^{+}15, WH16b, Wu19, Yan17, ZCZ19]. quantum-accurate [TST^{+}15]. quantum-mechanical [PD16b]. quasi [Ama15, Ama18, BC16c, DDD17, ETAG15, EDK19, FCL17, KKS15, KKS16, KPJ18, KCS^{+}17, LHQ19, MA16, Noe15, NF17, OMLdL16, RLH19, RPC^{+}18, RSB15, Sun19b, TZSS17, TMS^{+}19, YXD^{+}16, ZD15a]. quasi- [YXD^{+}16]. quasi-conservative [NF17]. quasi-DNS [KCS^{+}17]. quasi-Hamiltonian [EDK19]. quasi-interpolation [Sun19b]. quasi-Lagrangian [LHQ19, OMLdL16]. quasi-minimal [ZD15a]. quasi-neutral [Ama15, Ama18, DDD17, KKS15, KKS16]. quasi-optimal [ETAG15]. quasi-periodic [RPC^{+}18]. quasi-spectral [MA16]. quasi-static [FCL17, KPJ18, RSB15]. quasi-stationary [TMS^{+}19]. quasi-unconditionally [BC16c]. quasidiffusion [Ani19]. quasigeostrophic [WGJS19]. quasilinear [MNW19, WTX17]. quasipotential [YPC19]. Quaternion [ARTG^{+}19]. quenching [CSY15]. quintic [ZYW16].

R [Pan20]. Rachford [SwS16]. radar [CW16, CW17, Dod17, SG18]. Radial [KMGR16, SGN16, SNK18, ASS13, ASS17, CS18a, CB19, FHY^{+}19, FBW16, GBvZB16, GA18, HSC16, HXB15, LB15, LHY17, MF17, RRD19, SKS17, Sha17b, SF18b, SW18a, SP15b, WQZ15, WF17, XYPT16, XL17a, YYL18, Mue18]. Radial-Basis-Function [Mue18]. Radiation [DRM15, AMXJ19, BGD19, BHE^{+}17, CG15, CCGH19, CwYjS16, DS15a, DC18a, HR18b, KPKGH19, KL15, LMH16, NT16, PJE^{+}16, PD15, Rag15, RKO^{+}17b, SSM15, SFDE15, SL16b, TWM18, YHQ15, ZM16b]. Radiation-Hydrodynamics [DRM15, BHE^{+}17]. radiations [WT16]. radiative [Ani19, BJRF18, CCGH17, DPW^{+}15, FAY19, FYC^{+}18, HPC19, Her16, ION^{+}17, JL17c, LFRH17, LTKA15, LMG19, MRM16, PCMC19, PCA19, RZZ19, SJX15, SJXL15, SJX17, WED15, ZCQ19, ZCQ20]. radiative-transfer [DPW^{+}15]. radiofrequency [JH15]. radius [CCZ18, XXR18]. Raman [SNB^{+}15]. Raman-lidar-based [SNB^{+}15]. Random [LWL18, OADN19, ADHN15, BSP18, LL17, CLZZ19, CC19b, CLX19, CN16, CELZ18, DGHP17, DH18b, Gri15, Gri19, HKLZ18, HS17a, JXZ15, JL17c, KC17b, KGT15, LS16b, LL16c, Liu19b, LY17, MS15a, MSS16, MN18c, NHA18, PVPK17, PND16, RFGSV15, TST17, TG17, WMY16, WN17, WH16b, YGEM17, ZFPB16, TSFS17]. Randomized [BBB15, ZNX15]. randomly [LZT17]. range [AS16, FMPT18, LXL17, Loh17, LMC19, MDP^{+}15, SG19a]. ranges [SY18b]. rank [AAD16, BDKK17, BKKY19, CL18, FDKI17, FG18, HKLZ18, Jer19, KS16b, Lau17, LO16, RNO19, WAZ19]. Rankine [FKK19]. RANS [CARdN19, DWR18, EKV^{+}16, NMJFM19, PBC^{+}17, WS16, YKMM19]. rapid [DWR18, XYPT16]. rare [CL17, GH17b, MCS16, SPB16]. rarefied [CCL16, DMAM15, DM16, DY16, DLR15, GJ15, KJ17b, LLY18, LXSC16, LWX19, SYI^{+}19, SBT17, STKH15, WT19]. rate [AWJ17, HMRG16, MZAF17, MKB19, Opp17, TWN19, WY19, WCWY19]. rates [BR15a, DD16a, SN19]. ratio [GZLH19, KLC18, LWB^{+}16, NPB19, NGPB19, OM19, Sti16, TKF17, WSY15, WSS^{+}15, WSHT15, WDT^{+}19]. Rational [ABDN19, JB15, VMN^{+}18, WF17]. ratios [Cif19, FB17, KP15a, YM17c]. raw [NKN^{+}17]. Ray [CFKK19, RKO^{+}17b, TSFS17, CLQ17, DC18a, JH15, NLK^{+}15, Sto16, Swe18, WSU^{+}15]. ray-based [CLQ17]. Ray-Moment [RKO^{+}17b]. ray-tracing [DC18a]. Rayleigh [BGM16, CSG17, RLV16, RMLvR18, WC18]. rays [CL19b]. RBC [Ler15]. RBF [MF17, Sha17b, SF18b, BFFB17, BFF19, DA17, FFBB16, GN16, KAR17, KÁGR18, LLH19, PLR18, PLPR19, RBY19, Sha17b, SNK18, TLLF15, XWB15, ZNS19, ZZW^{+}16]. RBF-based [LLH19]. RBF-FD [BFFB17, FFBB16, KÁGR18, PLR18]. RBF-LOI [SNK18]. RBF-spectral [RBY19]. RBF-vortex [XWB15]. RBFs [SNK18, ECC18]. RCS [ALM^{+}17]. re [MMSS15, Wac15]. re-initialization [Wac15]. re-meshing [MMSS15]. reacting [DAO17, MM16a, MM18, SXBB15, TKP16, ZLY15, ZW15]. Reaction [BL18, ADH^{+}16, BTVC16, CLC16, CCM15, Cot16, Cui15, DJL^{+}19, EEG^{+}15, FNNW19, FBF15, FHE15, Hiv18, IZ18, JZ16, JJ17, Kay15, KS15a, Kim15, LPB17, LYB18, LWZ19, LCCZ19, LM19d, MMNI16, MD18, MWD16, MPT16, MW15, MN16c, NCP^{+}17, PD15, RB15, SSDN15, SYM15, SGA^{+}15, TWN15, ZLL^{+}17b, ZWYW18]. reaction-diffusion [CLC16, CCM15, FBF15, FHE15, IZ18, LM19d, MMNI16, RB15, SGA^{+}15, TWN15, ZLL^{+}17b]. reactions [BL18, IZ18, SPB17]. Reactive [SP15a, HGR16, LLD^{+}16, MN18b, MA16, PWP15, RFGSV15, SDM^{+}17, VMK^{+}19, WSN^{+}15, XXR18, TRM16]. reactive-mixing [VMK^{+}19]. reactivity [BHdD18]. Reactor [TCS^{+}16b, CSLL15, DDJ17, HBC^{+}16, HED^{+}16]. Reactors [TK16, TM17, ABdC^{+}18, CLB^{+}16, FBC^{+}16, GDS^{+}16]. Real [ATF16, BHL15, ATC17, ATC19, BLVC16, BW18b, MLI17, MRZG16, MRK15, PD16b]. real-fluid [MLI17]. real-space [BW18b, PD16b]. Real-time [ATF16, BHL15]. realistic [BPS16, CCBF19, CGG18, GBC^{+}18, HR17, TZSS17]. Realizability [CEHM19, AS15, Sch16b]. Realizability-preserving [CEHM19, AS15, Sch16b]. Realizable [LN17]. really [RFGSV15]. Rebay [MRRRF18, QN19]. REBO [TRM16]. receptivity [PBCR19]. recombination [CLGA17, YCBC15]. reconnection [BS15b, DD17c]. reconnection-based [BS15b]. Reconstructed [LLLN18, CZL18, CL19a, HL16a, NLW^{+}16, PL16a]. reconstructing [KYUO15, KBR17, Par17, PR16b]. Reconstruction [ALMJ15, AWJ17, BMPS18, MCGS16, Yam19, AHNF15, AMH^{+}18, APV^{+}18, BAGK16, BB19, BJ16, BMCK15, CLP16b, DIX^{+}18, DF16, DL18b, DJD^{+}17, EDvW17, FRRV16, GV18, GDFL17, HK19a, HY16, KKH18, LGB17, LX18, LAEK18, LAK^{+}16, LHGF16, LMGG17, MDMS18, PK16, PVB17, QN19, RÖS16, RW15b, SR19, SW15, SGD18, SO15, SO16, SIX16, VV16, Vil19, WRPL17, XDLX19, Yeo19, YK19, ZOG19, ZS15, ZSL^{+}19, ZRW19, ZXW^{+}19, LWJV19, ZN16]. reconstruction/differentiation [LAK^{+}16]. reconstructions [SL18, TLQ16]. recovered [BCS19]. Recovering [CJK^{+}19, RZ18, ZZ17a]. recovery [AM18, DH18a, GY17, GY18, GYZ19, JJ19, LYL19, LS19b, SZ15b, XGZ19]. rectangular [CCFC19, CV16a, GKE15, PMGW16, PG18, RMLvR18]. rectangular-grid [CCFC19]. recurrence [FHY^{+}19]. Recursive [HSSZ16, AN15, FBJS19, HS17a, TZSS17]. recycled [GWC17]. Recycling [AdSŚ^{+}15]. red [HS17b, KLC19]. redistancing [LDOK17, RPL^{+}18, dLKK19]. redistributed [HLL^{+}16, PS16]. redistribution [KO17]. Reduced [BIR18, CB19, FS17a, LL16c, WC18, AH15, APLK19, ATF16, AMPG19, BTD16, BM16, BDMZ19, BCG^{+}15, CR17, CCBdL15, CS16a, CMP19, CQ15, CS16b, DC18a, HU18, JL17b, JD19, KMD^{+}18, KM17, KTG16, LK17, LVB^{+}15, LWZ19, LL19b, MLL19, MWD16, MP17, MRXI17, OS16, SFT16, SSN15, TD16b, URL16, WHR19, XYF^{+}17, YYJ^{+}19, vdBKD17, TG17]. reduced-basis [CS16b]. Reduced-order [BIR18, FS17a, ATF16, JD19, KMD^{+}18, MLL19, SSN15, TD16b, URL16, XYF^{+}17]. Reducing [Bra16b, GZM^{+}17, MLB18, BGGM15, XWW^{+}16]. Reduction [BMY19, BMCK15, PG18, AEVW18, An17, AZ19b, BJO18, BBKS18, BGG16, CBA17, CCS18, CLY^{+}15, CVG19, CEH16, CCL16, CMW16, Don18, GBvZB16, GAJ15, HFND18, HKBR^{+}19, JL15, JL18b, Lan19, LYCC17, LS15c, LHLL19, NMA15, NW15, OS16, PQR17, SLB^{+}16, SLB^{+}19, SFDE15, TBG16, VCNGP15, XL17a, ZWUR16, ZPE^{+}16, ZRE16, dPSS16]. reduction-consistent [Don18]. reentry [PLWJ16]. reference [ALA16, BT19, DJD^{+}17, LTR16, ZLGS18]. refined [CC17a, JW15c, SLY16, ZJLC15]. Refinement [BV15, COdLL18, PSB^{+}18, SS15a, SPD19, SL18, AAE19, ACS16, BHS^{+}18, CJC19, DMS17, EH14, EH15, FGL16, GK19b, HS17b, HS18a, HIN^{+}16, HW16c, JW15a, KL15, KLRT15, KJ18, KS16d, LS15c, LS16b, LLJJ18, LNM15, MNG15a, MW17b, MSB^{+}16, NdlLPCC19, SLdTV18, Tie18, WDG^{+}17, WDS15, XS15, dlAC17]. refinement-free [AAE19]. reflecting [FN17, MS15a]. reflection [DCA^{+}16, DJV^{+}18, EG18b, LFRH17, XB18]. reflectionless [Che19]. reflector [BtTBI18]. reformulation [CD17]. refractive [LTKA15]. regime [BZ16a, BZ19b, CGK17, CCZ18, CLV19, DLMDV18, GK19c, HBR15, HJS19, IKS19, LYA16, YSWW16]. regimes [KF17, LLFX18, PLWJ16, TCS16a, ZZX16, ZZX19]. region [JL18c, LDB19]. regional [PLHA18]. regions [Bon17, LH17a, MDT16, RC18, SR16]. registration [ZC15]. Regression [TE19, ARG^{+}17, GJ18, LLL16, NP16, PYK19, PVPK17, SX16, TMES19, YBSTT19]. Regression-based [TE19, LLL16]. regular [Gro18, KGT15, MDM^{+}15]. Regularization [BPM18, WSU^{+}15, BKL17, CLS^{+}18, GZ17, HW16a, KGS17, LT17c, LEB^{+}17, SHW18, UY19]. regularizations [HNS16]. Regularized [Cor18, TB19, VD16, WGME17, WCVF16, CDLR19, CV15, FBJS19, FC19a, GPRA18, NL15, SNB^{+}15, Smi18, WT19]. regularly [YFKS15]. reinforced [LYDB17]. reinitialisation [AGC19]. reinitialization [CD17, JH17, PLHA18, ZY19]. Rejection [CKKD19]. Rejection-based [CKKD19]. related [LS15b]. relation [NG17, NG18, SL15, YWHP15, ZLC^{+}18]. relations [Abg18a, STR15]. relationship [WS16, WSK19]. relative [MCHL16, Opp17]. relativistic [BAGK16, BK16a, CKT17, ION^{+}17, KFF^{+}17, LDT19, PSV18, QSY16, Sel15, SOS19, Teu16, WT15, ZT17]. Relaxation [ACCCDA16, ACCCD^{+}17, KJHA19, AIP17, APT17, CDN17, DLP19, EDC19, FBL17, GSN17, GN19b, GPTK19, HHM17, LMS17, LL19a, LMSK17, MG15a, NGS16, RJ19b, Wu19, Xie15, ZL15c, AC16]. relaxed [EN17, YM17a]. relaxed-Jacobi [YM17a]. relaxing [LMZ19, TEP19]. reliability [KM17]. reliable [RDQ19]. Relinearization [TYROG19]. relocation [SF18a]. remap [BMC^{+}18b, DDJ18, LM19a, MCL19, Bra16b]. remapping [CZJ17, SO15]. remeshed [HKLW15, OB19, RHvR^{+}15]. remeshing [BKKJ17, PA15]. Removal [KW15b, BG19a, CMR^{+}16]. removing [AR16b]. renormalised [BDB18]. renormalization [FYC^{+}18, LMJ18]. replica [BLS15]. Reply [EH15]. represent [MVKD15]. representation [ATM^{+}18, BT19, CDM19, DOO17, JL17b, LC18, LY16c, PES19, SG17, TG17, XX17, ZZKF15]. representations [BBB15, LTR16, QWX18, RBD17, UHKT19]. represented [ML16]. Reproducing [FRO17]. repulsion [LY15c]. repulsive [Rua18]. rescaled [MTD15]. Rescaling [Bon17]. RESCU [MRZG16]. Research [TM17, AK17]. reservoir [AVT17, QWXZ17, vdLJLV16]. reservoirs [ABG^{+}15, DGW18, TH18, YSLY19]. Residual [CEL15, AD15, BC16a, CC16c, CLGA17, CLNH15, GMLD18, GMD19, IC17, KSSL18, LW17a, LSI16, MN15, NS19b, OKWE17, PHRA16, Ric15, ZD15a, ZM16b]. residual-based [KSSL18]. residual-distribution [MN15]. Residual-driven [CEL15, NS19b]. resilient [LKK17a, LKK17b]. resistive [DNOP15]. Resolution [MMW15, AWS16, ANL^{+}16, BZ19a, BTGM17, BGTM18, BST^{+}18, BP18, BLK15, CFSZ19, CDX18b, FS19a, FHA17b, GMA18, iI15, iI17, KQB18, KH17, KK17b, LKK17a, LW17b, LZZS15, MKSP19, NYNYM15, NMM19, OMYvdP^{+}15, PE16a, Poë19, RHvR^{+}15, SSM^{+}17, SWL19, SZS15, TCB18, WLW^{+}18, YDN19, YK18, ZS18, ZS19a]. Resolved [SP16b, CMR^{+}16, FWK18, HF19, JL19, Kim15, KCS^{+}17, KLC19, KBG^{+}15, KSI17, MSP15, MMPS17, PD19, PBP18, SLC^{+}18, SMOM^{+}17, Vre17, Vre21, WSP17, WTL19, WMM^{+}18, ZLY15]. Resolved-particle [SP16b]. resolving [BVM17b, BCG^{+}15, DIX^{+}18, LHO^{+}19, MMPS17, SKG17, PEVG18]. Resonance [CLL17]. resonances [LZ19]. resonant [VMN^{+}18, DS15d, NMJFM19]. respect [TVB^{+}16]. response [BLS16, BCRS19, GTL18, GS18, LWLC17, MN18b, Pis18, ST16]. restrained [TR17]. Restricted [Mac16]. restriction [ML16, SMT^{+}16]. restriction-smoothed [ML16]. restrictions [DvW15a]. result [DvW15a]. results [MBM^{+}18]. retraction [LWTF19]. retrieval [WHRL19]. retrievals [SNB^{+}15]. retrieving [Par18b]. Retrospective [LHMB18]. reversal [RTG18, SSM^{+}17]. reverse [BT17a, CCFC19]. reversed [WSOW16]. reversibility [GKRB17]. reversible [BL18, IZ18, PD15]. review [Che18, GFO18, MGKG17, TMdO19, ZJLC15]. revised [FDS^{+}15]. Revisiting [AAI16, LM15b, LDOK17, Ram18, STFK19, VPV^{+}17]. revolution [EGO19, NBT19]. Reynolds [TUJ19, BBKS16, DCP15, Eva18, FMPT18, GZ19, LZB^{+}17, LHMB18, NL18a, OB19, WSY15, WDT^{+}19, WGME17, XWW^{+}16, XTYL18, ZV16]. Reynolds-averaged [GZ19, LZB^{+}17, LHMB18, XWW^{+}16]. rezoning [SP19b]. rezoning-free [SP19b]. RG [SFT16]. rheology [FNGDMNR18, FC19a, MDP18, SHP^{+}16]. rheology-IBM [SHP^{+}16]. rhombus [CCFC19]. ribbed [MF16b]. Richardson [ACCCD^{+}17, ABFR16, HK18a]. ridges [AKM^{+}19]. Riemann [CTM^{+}16, RZ15, AB19, Ama15, BD15a, BD15b, Bal15, BAGK16, BK16a, BVG^{+}16, BTGM17, BN17, BK17b, BGTM18, BLM18, BK19a, BM19a, BNGI19, CCBF19, DB16b, FFW17, FLW19, GSC19, GP16b, Guo15, GFW16, JZSX18, KW15b, LSTkM15, LDGH16, LZW^{+}17, MS15b, MBD15, NMM19, Niu16, PP19, QSBY19, SW17a, SYY16, SM19a, TM15b, VNA15, ZHA17a]. Riemannian [RNO19]. Riesz [AHKT17, DLC15, MD17, YYN^{+}17]. right [Mac15]. rigid [BHST17a, BHST17b, BHST18, CFSN18, CGRV17, DSH^{+}16, DCA^{+}16, IML15, LTB16a, LMM17, MP19, PN18, PR16a, QYF15, RW15a, Say17a, Say17b, SGMS16, STKH15, TOR^{+}15, WE15]. rigid-body [WE15]. rigidity [LL18]. Ring [ZSL^{+}19, SB18]. rising [NBMB19, WB17]. Ritz [Lot18, RMLvR18]. RKDG [BK17b, ZYD^{+}19a]. roadmap [CSS17]. Robert [MRY19]. Robin [GAS^{+}18, ABG18c, BSP18, BG19c, ED16, LZ17b]. Robinson [KS18b]. robotic [LHY^{+}19]. Robust [BDMZ19, FWK18, GPTK19, KWHB19, LGO17, LSLA16, LDO^{+}19, MHHX16, RWKW16, Sha17a, SWLZ15, Sti16, WN18, YL17, ARG^{+}17, BCSK17, BW18b, CBC^{+}18, CZJ17, CCB^{+}19, CD17, CV16b, DL16, HK19a, HHM17, JSP16, KTN15, LKB15, LZ17a, LSD^{+}17, LM19c, MTK^{+}16, NGPB19, Nis19a, Niu16, NW15, OSKN18, PL18, PGM17, QYJ19, SYY16, Sub15, SGP17b, WR15, YK15]. Robustly [LYB18]. robustness [COdLL18, MDMS18]. rock [dFGS^{+}17]. rock-physics [dFGS^{+}17]. rocket [SDM^{+}17]. rod [AMM^{+}15, CLB^{+}16]. rods [AHHC18]. Roe [BR16, NMM15, ATC17, ATC19, BR15b, Bon17, CC17b, GMD19, LT17a, PM19, SPP16b]. Roe-like [SPP16b]. Roe-MUSCL [BR16, BR15b]. Roe-type [CC17b]. rogue [FS17a, WMY16]. Role [NWB19, BFFB17, BFF19, FFBB16]. Romenski [Jac17a]. Roohi [Pan20]. Rosenbluth [TCSM15, TCS16a, TCS17]. rotated [GH17a]. rotating [ALT17, BC18b, Cai16, CE17, HSM19, LP18, NMM18, RLP16, SGN16, SP16a, SGC18b, TC15b, WLM15, ZL15b]. Rotation [HSB16, ATZ16, BTD16, CFKK19]. Rotational [HW19, CVG18, WYLX17]. rotations [YLBL16]. rotor [BBK19]. rough [EG18b, HHLY17, LWC17, OZ17, XWW17]. roughness [YFKS15]. RPA [LT17b]. RTM [LZL^{+}19]. rule [AW16, ZSP15]. rules [CS17a, Tsa15, ZGJ16, vdBKD17]. run [FKR16]. run-up [FKR16]. Runge [BR16, SLL17, BR15b, BZ18, CCRdL17, CB15, CCGH19, HK18a, HS18b, JH17, LDSM19, MVK16, MW17a, MH18b, NMC15, O’S15b, O’S19, PP17, SV19, VLV19, Ver19, WJD16, WBM15a, ZT17, ZYD19b]. runs [Bre18]. rupture [DD16b]. Rutherford [Hig17].

saddle [JL18b]. Sadegh [HSK^{+}15]. SAFT [RVK^{+}18]. Saint [LAEK18]. Saint-Venant [LAEK18]. Salpeter [BDKK17]. SAMBA [ABM16]. sample [HD18, LB17]. samples [Gri15]. Sampling [PDS15, PT17b, AM18, CKKD19, CW17, CC19b, CCL16, Cot16, FJLC18, FMPT18, Gen15, GHH^{+}16, HD15, KBK15b, LL15, MTM19, Par18b, PRvdL18, RFPSSA18, SX16, SG16, WBC^{+}16, YYL16, RC18]. sandwich [AA19]. SAR [SGP17a]. satisfy [FS15]. Satisfying [SYV17, Abg18a, CHY16, CT19, CHS17, LW17e, SWPS17, SYV14, YL19]. saturated [HSK^{+}15, LL18, SSL17, Zad11]. SAV [SXY18, ZS19b]. save [TP17]. Saving [FJLC18]. SBP [Mat17, MAvdW18, MO18b]. SBT [GRS15]. scalability [EDC16]. Scalable [CWF16, CMW16, DM19, IPSG15, KRBW17, ANL^{+}16, CS16a, EG18a, EJMI18, KC17c, KDPK15, MCN18, MNC19, TCSM15, TLLF15, WLC15, YS17]. Scalar [DM19, IC17, MD16, SKC17, CDX18b, HAX19, HAH16, HS18b, JCWX19, LBZA16, MWZ19, OMYvdP^{+}15, PKN17, RBGV15, SYOS19, SYOS21, SMD18a, SXY18, SWZ17, TWN19, TNB21]. scalar-source [MWZ19]. Scale [HHA15, ABI17, ALO18, AG18, BB17, BKS18, BB15, Cac15a, Cac15b, CGSS18, CPT16, CGC17, CMW16, DLLV17, DKPC15, DNOP15, DPW^{+}15, DP19, EZG16, FX18, GHH15, GMT19, HHLY17, HKS^{+}16, IPSG15, JS19, KJ17a, KLRT15, KS16c, KDPK15, KSSL18, LHS^{+}18, LLM17, LHO^{+}19, LMC16, LXL17, LWL18, LY16a, LHA16b, MT18, MT19, MAH16, MOR18, NS19a, OCSC18, PD17, SKP^{+}15, SSL^{+}16a, SAH17, SDJU15, SRBB18, SDH^{+}16, SSX16, SKW19, SPM^{+}15, SSA17, SLdTV18, Sun19a, VKE^{+}18, WMY16, WSN^{+}18, WB17, XB18, ZKS^{+}15, ZQCT15, ZPE^{+}16, dlCGCA17, PEVG18]. scale-bridging [DPW^{+}15]. Scale-Resolving [PEVG18]. scaled [LZS^{+}19, GBS15]. scales [Hig15, LLA19, MMW15, SDJU15, TBHG18]. Scaling [HSF19, BMC18a, CCDL19, JLQX15, LL16a, LDB19, LY15c, LY17, LT17b, NN18, XXR18]. scalings [JXZ15, JL17c, Liu19b]. scattered [BFP18, CGM18]. scatterer [RTG18]. scatterers [CR17, CCZC16, DKTH15, JHPAT17]. scattering [AJP15, AN15, APKP16, BABD16, BXY17, BNM15, BDMZ19, BM19b, BKL17, CKKD19, CDL17, Cho19, CHE^{+}17, DCCC16, DvB17, DDV^{+}15, EGO19, FM15, FYC^{+}18, GMP16, GD19, GHJ15, GHH^{+}16, HK19b, Hig17, HN17a, HN18, KE15, LKB15, LGO17, LO19, LW15a, LZ19, LB16, LS19b, LBB^{+}17, PA19, Par15, PLL15b, PWC18b, SUR18, SF18a, SLVE18, UWH17, XB18, Yam19, ZFZL15, ZZ17a, ZZ19]. SCDM [DLY17]. SCDM-k [DLY17]. scenarios [BHNS19, WAF^{+}19]. SCF [BW18b]. Scheduled [ACCCDA16, ACCCD^{+}17, KJHA19]. scheme [AIP17, AdRBC16, Ali15, AS15, APKP16, ABH^{+}19, ADH^{+}16, AA15, ATC17, ATC19, AHKT17, BHZ16, BAGK16, BK16a, BEJ15, BK17c, BCS19, BT16, BSP18, BSWG15, BHGK18, Bon17, BCC^{+}18, BQRX19, BDLM18, Bra16b, BCM15b, Buk16, BRK17, BMC^{+}18b, CGK17, CSY19, CX15, CXL16, CTG16, CY19a, CHJT17, CDX18b, CGJ19, CC16c, CLV19, CB18b, CV16b, CHS17, CCM17, CDV17, Cui15, DGW18, DDJ17, DSX19, DSS18, DLN15, DS15c, DL18a, DYL19, DMTB15, EDW19, EMS^{+}19, EH18, EMSS17, FDS^{+}15, FX18, FLT17, FK19, FSB16, GZY19, GOR17, GHR17, GBM16, GGL^{+}17, GT18, GHL15, GHL^{+}16, GX15, GLW18, HK19a, HSM19, HW15a, HCB19, Hiv18, HW15b, HWA15, HAH16, HC18a, HC18b, HLA19, HJY19, HTBG15, Ism15, IDSG15, Jac17a, JN19, JME18, JPSX18, JW15c, JS17, JC17, JJ18b, KGS17, KPP^{+}19, KHC^{+}16, KV16]. scheme [KC18, LL19a, LLP^{+}16, LHO^{+}19, Ler16, LL16b, LHMB16, LT17a, LLWJ18, LLFX18, LC19, LWWY18, LSTkM15, LMKS15, Liu16, LXSC16, LTWZ18, LAEK18, LCK19, LYZ19, LMB19, Liu19b, LWX19, LM15d, LHQ16, LI15, MLI17, MNG15a, MRRRF18, MA19, MN15, MDBCF17, MS18d, MTJ17, MSB^{+}16, MBM^{+}15, NMM15, NS19b, Nis18a, NF17, NT16, OSKN18, PX15, PXLL16, PX16, PL16b, PR19, PGM17, PBBK15, PKK18, Poë19, PS17, PA15, PSV18, PME^{+}15, QWX18, QWZ19b, RRS19b, RRS19a, RDG17, RSSSE18, SNSG16, STK^{+}16, SKF15, SS16b, Sha17a, SFT16, SWPS17, SP19b, SY16, SG19b, SLL17, SOS19, SWJG19, SH19, SWK18, SWHK15, SPZ18, SD16, Sto16, SS17c, SDW18, SK18, SWL19, SJX15, SJXL15, SJX17, TZGW18, Tav16, TK12, TK15b, TM15a, TC15b, TKP16, VST16, Vil19, WR15, WH15, WDS15, WY16, WH16a, WHY17, WYLX17, WLGD18]. scheme [WLW^{+}18, WHY18, WL17, WZL^{+}17, WRL18, XCX17, Xie15, XQ17, XS19, YSWS16, YWS^{+}16, YFJ17, YFJ18, YD19, YLA15, YWHP15, YT19, YZZ15, ZG17, ZSW17, ZFL^{+}19, ZYSW16, ZPW18, ZMZC19, ZW19, ZLGS18, ZWYW18, ZZX16, ZQ16b, ZQ17, ZWG17, ZZX19, dSPDH15, dBIM16, dDPG19, vEKdB16, FRO17]. Schemes [ZQ16a, AHNF15, AMH^{+}18, AD15, Abg18a, AD17, An17, ABH18, ABR16, AHZ19, ALMJ15, ADOP18, ABIR19, BZ19a, BD15a, BGS16, BTGM17, BK17b, BGTM18, BK19a, BC18b, BR15a, BR15b, BR16, BH18, BLMY17, BLD15, BDZ15, BD17, BB19, Bra16c, Bre17, BTVC16, CBZ18, CV17, CCRdL17, CBS18, CCP19, CB15, CC17b, CJYZ15, CS16c, CKT17, sCYxL^{+}18, CLS^{+}18, CCZ15, CYS17, CFSZ19, CCK^{+}18, CLTX15, CTM^{+}16, CK16b, CCDL19, CwYjS16, DS15a, DDJ19, DPO16, DL15, DLMDV18, DL18b, DJLQ18, DB18, DPRZ16, DPRZ17, EMM^{+}18, EL18, EL19, Eng18, EN18, EDC19, EJZ17, Fan16, FNGV18, FLV15, FG17, FLW16, FHA16, FHA17b, FHA18, GJL19, GSS15a, hGwSzS15, GS15a, GMLD18, GMD19, GWK16, GCVMK15, GR15, GSK18, GBCF15, GBCF16, GGT15, GAJ15, GLK19, GGT18, HK18a, HSLQ16, Heu19, HLJ^{+}19]. schemes [JLQX15, JMM19, JZSX18, JSB^{+}19, JLC18, KM17, KW15a, KW15b, KGS17, KGT15, uKHGK19, KFL17, KHTZ19, KYW^{+}16, KYW^{+}18, Kri17, KTK18, KTK19, LN17, Ler15, LX16, LDT19, LM19a, LZZS15, LJ16, LW17d, LCF16, LPR19, LMN18, MCL19, MCW16, MN16a, MSP19, MKC17, MDMS18, MDP^{+}15, MCGS16, NMM15, NMM16, NMM17, NMM18, NMJFM19, NMC15, NLFM16, NL18a, NR17, NG17, NG18, OS15a, OZ17, OV17, PxRS17, Par18a, PS16, PBC^{+}17, RS15a, Rod17, RRD16, RSH^{+}17, Sch16b, SAEF17, SSZ19, SUR18, SMS16, SYY15, SWLZ15, SGC18b, STG17, SY18b, SSM15, Sto17, SGL17, Stü17, SIX16, Svä15, SN19, TLQ15, TLQ16, TLH15, TD18, Ter18, Tso18, Vab18, VW16, VN15, VV16, VLV19, Ver19, VSM16a, VSM16b, WW15, WRL16a, WCL15, Wil18, Wil19, WMM^{+}18, WT15]. schemes [WSF17, XLL^{+}17, XJLQ15, XL16, YC17, YDLC19, YYL16, Yan16b, YH17, YLD19, YL19, ZZK16, ZJLC15, Zha16, Zha17c, ZRW19, ZCQ19, ZZL19, ZCQ20, ZY17, ZSMP19, ZSQ17, ZS17, ZS18, ZH19, ZS19a, dFJN16, vLtTBI17, ZN16]. Schrödinger [ATZ16, ABR16, BM15, BJTZ15, BA15, BPTA16, BCM15a, CBZ18, CQL^{+}17, CV16b, CHLZ17, GMP15, GN16, GWWC17, GT18, GHL15, HZE19, IKS19, LGH^{+}18, LYA16, LHL15, STEK17, STEK22, SL15, WH15, WWRS17]. Schrödinger-like [WWRS17]. Schrödinger-type [GT18]. Schrödinger/Gross [ATZ16, ABR16]. Schur [JTD16]. Schwarz [Kas15, KC17c]. science [AK17]. Scientific [ZWB^{+}18, AKK^{+}19]. scour [RPNP18]. scramjet [CELI15, GSS^{+}19]. scrape [HRJ^{+}16]. scrape-off [HRJ^{+}16]. SDE [AÁPB17]. sDEM [ACGR15]. SDEs [XZJK19, YPC19]. SDPD [DJL^{+}19]. SDS/MC [XWZ^{+}18]. SE [WMS18, ZWL^{+}19]. sea [ALTR17, KDL15, MR17, SRBÓ17, WTL17]. sea-ice [SRBÓ17, WTL17]. Seafloor [EFHZ17]. seamless [iI15]. search [GBCF15, GBCF16]. searching [PKLS17, WTL19]. seas [WMY16]. Second [BHE^{+}17, Cac15a, Cac15b, CKT17, CR18, DS15b, DLMDV18, DVP^{+}16, FYZ^{+}15, Gro18, HHY16, Ism15, LMC16, NDCB17, WXW15, ZzSK15, ABFR16, ABH18, ABG18c, BTGM17, BLM18, Bat17, BR15a, BND16, BST15, BDLM18, CC15, GP17, GAIN19, GBCF15, GBCF16, GZLH19, HW15a, HBR15, LN17, LSL15, LPW15, LC17a, Liu16, LLA19, MZAF17, MN04, MN17, MN15, MH18b, MM18, PHRA16, Roy15, Sha17a, SLL16, WKOE17, Wu16, Yan16b, YH17, YLD19, ZY17, ZW19, ZWYW18]. Second-order [BHE^{+}17, Cac15a, Cac15b, CKT17, DS15b, DLMDV18, DVP^{+}16, FYZ^{+}15, HHY16, LMC16, ZzSK15, ABH18, ABG18c, BTGM17, BR15a, BND16, BST15, BDLM18, GAIN19, GBCF15, GBCF16, GZLH19, LN17, LPW15, LC17a, Liu16, MZAF17, MN15, MM18, Sha17a, WKOE17, Wu16, Yan16b, YH17, YLD19, ZY17]. section [ABT16, Dod17, JDFS16]. sectional [FSK^{+}16]. sections [CV16b, DJD^{+}17, KFL17, LMGG17]. sediment [BVM17b]. sedimentation [AGRB18, BKRB15, KM15, ZZPH18b, ZZPH18a]. sedimentation-consolidation [AGRB18]. Segel [ZM16a]. segmentation [WLWW17]. segments [Cor18]. Seismic [MF17, CZW17, LLY15, LTB16b, LTXB17, MKYZ17, dFGS^{+}17]. selected [DLY17]. selection [FOF15, JES15, KKL15, LHLL19, Xia15]. selective [CGTH18, MTK15, Sub18]. Self [BLL19, BLL20, DK18a, DK18b, LLB19, MNW19, OLD^{+}16, TTN^{+}16, TSR15, ADFG17, BD15b, BVG^{+}16, BN17, Cen19, KLWQ17, LLVF^{+}15, LYPP17, OLB^{+}17, SCQP16, VGZ18, Wal16, WXSJ19]. Self-adaptive [MNW19]. self-adjoint [LYPP17]. self-avoiding [Wal16]. Self-consistent [DK18a, DK18b, OLD^{+}16, TTN^{+}16, TSR15, ADFG17, Cen19, WXSJ19]. self-focusing [KLWQ17]. Self-induced [BLL19, BLL20, LLB19]. self-propelled [SCQP16]. self-similar [BD15b, BVG^{+}16, BN17, VGZ18]. Semi [AGKD19, CT19, Cot18, FSM16, Gam15, GT19, GXX17, KCHW19, LC18, MM16a, STEK17, STEK22, SKW19, WCN15, ZBH^{+}18, AZ19a, BDM17, BFNGDNR18, BQRX19, CQQ16, Cai16, CGQ18, CBM19, CM16b, DLR15, DAO17, DJLQ18, Ein19, FFM19, GS15b, GBD17, HPY18, HAH16, KFWK17, KYPK15, Lap17, LYM19, LXC^{+}15, LZT^{+}15, LYA16, MD15, MTD15, NRZS17, OD17, PBBK15, PKK18, PME^{+}15, RAMB15, SXBB15, SW18a, SFT16, SH19, SLZ^{+}17, TD17, TM15b, WTL19, XXR18, YZW^{+}18, ZBZ^{+}18]. semi-alternating [LZT^{+}15]. Semi-analytic [KCHW19]. Semi-analytical [GT19, LC18, MD15, MTD15, TM15b]. semi-classical [LYA16]. semi-discrete [SLZ^{+}17]. semi-empirical [YZW^{+}18]. semi-explicit [KFWK17]. semi-flexible [NRZS17]. Semi-global [STEK17, STEK22]. Semi-Implicit [CT19, AGKD19, Gam15, GXX17, MM16a, SKW19, WCN15, ZBH^{+}18, AZ19a, BFNGDNR18, BQRX19, Cai16, CM16b, DJLQ18, GS15b, HPY18, Lap17, LYM19, LXC^{+}15, PME^{+}15, RAMB15, SXBB15, TD17, XXR18]. semi-infinite [GBD17, ZBZ^{+}18]. Semi-Lagrangian [Cot18, AZ19a, BDM17, CQQ16, CGQ18, CBM19, DLR15, DAO17, Ein19, FFM19, HAH16, KYPK15, OD17, PBBK15, PKK18, SW18a, SFT16, SH19]. semi-resolved [WTL19]. Semi-spectral [FSM16]. semiclassical [HHY15, IKS19]. semiconductor [FKF17, HW15b, Liu19b]. semiconductors [CLZZ19, vdKK16]. semilinear [HJW19, ZHW18]. Semiparametric [BH16b]. semismooth [YSY17]. sensing [KKZ15, LSD18, THS^{+}19]. sensitivities [Cac15a, Cac15b, KPKG15, Loz17]. Sensitivity [NW17, NWFT19, SD17, AMJ17, ADP^{+}17, Blo17, BW18a, Cac15a, Cac15b, CNW17, GMS19, LSM19, Lia16, NT19, SP19a, SW17b, TCA16, TMWF18, TWF19, WSK19, MBJ16]. Sensor [ABdC^{+}18, NHM17, Fon16, FKK19]. sensors [ST16, ZYK18]. Separable [BPF^{+}16, LT17b, PGH15, TBO^{+}16, ZZT^{+}16]. separated [BBB15, FW17, RBD17]. separation [BG19a, Fon16, GKNA17, HHA15, JL18b, LZL^{+}19, LHA16b]. sequence [MCS^{+}19]. Sequential [CC16b, DKK^{+}18, LPU18, MTJ18, PvL19, SWX18, WHT18, WKHT19, WX18, YNW17, CM18d, MP17, MTJ17, RZOZ19, SGP17a]. Sequential-implicit [WKHT19]. serendipity [HR18a]. series [KKP15, OLV16, YYL16, YL16]. Serre [Pop15]. Serre-Green-Naghdi [Pop15]. Set [LVTR15, SGD18, VALT16, AASRT17, AGC19, AT18, BT19, CSW^{+}19, CWWZ17, CD17, CG16, CM16b, GLTB18, GM16, GHP15, GFO18, GWYS18, HKJ17, JLC15, JGS16, LSMS17, LDO^{+}19, LYKW19, LM19b, LVL18, LSYF15, MGBG16, MW17b, MLMM17, NLK^{+}15, OLB^{+}17, OLD^{+}18, PLHA18, PDN19, PC19, QWX18, STV18, SCJ^{+}18, SSA17, TAR17, Wac15, XSL18, YJM19, YCS^{+}17, ZA15a, ZLH^{+}17, ZC18, ZC19a, ZY19, ZHW18, dLGT^{+}17, dLKK19, AAL15, AB15, BAVC17]. Set/Ghost [LVTR15, GLTB18]. sets [DHH^{+}18, JH17, LN17, STHW17, WX18]. setting [BS19a]. settings [CK16a]. several [GBR15, Shu16]. severe [GZM^{+}17]. severely [QYJ19]. SFO [MAP17]. SGS [LL16b, MNG15b]. Shadowing [NW17, NT19, NWFT19, Blo17, BW18a, LSM19, CNW17, SP19a]. Shafranov [PKF16, RCRF16]. shale [AEVW18]. Shallow [ABT16, NMM18, SP16a, ALKZ16, AB19, BC18b, BNGI19, BFNGDNR18, BHGK18, Cap18, CS18b, CE17, CDV17, DK19, DA17, DMTB15, EL18, EL19, EDK19, GP16a, GIF18, GCVMK15, HSM19, HLJ^{+}19, JJS15, Jou15, KL18a, KDPK15, LMPS15, LPG18, LP18, LCLY19, LDWZ15, LMKS15, LMSK17, LCK19, MDBCF17, NMM15, NMJFM19, PP19, DM18, Ric15, SGC18b, SMSR18, SD16, TC15b, TSB^{+}18, VST16, WHRL19, WWGK17, WWGW18, WG15, WBM^{+}15b, ZA15a, ZED15, NMM17]. Shallow-water [ABT16, KL18a, MDBCF17, ZED15]. Sham [BEJ15, BHL15, CDM^{+}16, HXX18, ZLH^{+}17]. Shape [DBD^{+}17, TS18, Wal16, BFI^{+}16, BCO^{+}15, Bar18, BMPS18, CF19, HKJ17, KPKG15, LDO^{+}19, Loz17, RLH19, SKF16, TBLM15, ZHW18]. shaped [DWZ19, GN16, SUR18, TSN16, TP16b]. shapes [GA18, WHL17]. Shapiro [Fal15, Fal17, GDFL17]. Sharp [GHF19, TGS19, ACS16, FMRZ17, HHA15, HG17, HK16b, LHA15b, MIM^{+}19, SHA16, SP15b, Tsa16, ZD15b, ZMZC19, ZSM19]. sharp-interface [SHA16, ZSM19]. Sharpening [CSN17, HTZG17, MNG15a]. sharpness [LWY17]. shear [BVMW16, BNGI19, CB18a, GIF18, SK19a, TEP19]. sheared [LVB^{+}15]. sheath [dCPDC^{+}17, TTN^{+}16]. sheet [CCB^{+}19, CLvS17, CS18a, CLFL17, IPSG15, MML17]. sheets [AGKD19]. shelf [Jou15]. shell [CLFL17]. shells [GLS15, MKS18, SDMS17, WLM15]. Shift [PJE^{+}16, SZ17]. Shifted [BM19b, MS18a, MS18b, NRS19, RM16, SLR^{+}16, SMSR18]. shifted-Laplacian [RM16]. shifting [KGS17, OMLdL16, PG18]. Shock [CLG^{+}19, Sid18, ZD17, ABIR19, BPM18, sCYxL^{+}18, DM17a, DSX19, DZ16, FKK19, KYW^{+}16, KYW^{+}18, LYZ19, LMC19, LSI16, NMM19, NLL^{+}15, OSKN18, PSS17, RRS19b, RRS19a, RLGT19, RA17, RA19, SGC^{+}17, SGC^{+}18a, SYY16, SM19a, TLB^{+}18, TEP19, WL17, ZS16, ZSMP19, ZXDL17]. shock- [WL17]. shock-capturing [KYW^{+}16, KYW^{+}18, OSKN18, PSS17, RLGT19, ZSMP19]. shock-fitted [RA17, RA19]. shock-fitting [ZXDL17]. shock-particle [SGC^{+}18a]. shock-stable [sCYxL^{+}18]. shock-turbulence [BPM18]. shocked [CJD^{+}17, PME^{+}15, WTS^{+}17]. shocks [SWL19, WS15b, XLL^{+}17, dFVJ15]. shockwave [NMM17]. Shooting [SP19a, BW18a, Die15]. Short [Teu15, HS17b, XYPT16, YY16, ZW15, ZX19]. shot [BDB^{+}17]. shrinkage [KKL15, WYA^{+}17b]. Shu [YY16]. side [FNGDMNR18]. sided [KSM19, SZM19a, SZM19b, SYM17]. sign [DCCC16, DY19b]. sign-changing [DCCC16]. sign-preserving [DY19b]. signal [RKB19]. signal-noise [RKB19]. signaling [CFG16]. signals [CWL^{+}16]. Signed [SCL19, Sel15]. significant [TBHG18]. similar [BD15b, BVG^{+}16, BN17, VGZ18]. similarity [NN18]. Simple [KH17, KC18, AL19a, ATC17, ATC19, DL16, HK15b, HC18b, KBK15b, MHT^{+}19, Niu16, OS15a, OSKN18, RS15b, SM19a, TDC^{+}19, VNA15, ZL15b]. Simplex [EDC16, KHTZA16]. Simplex-in-cell [KHTZA16]. Simplex-stochastic [EDC16]. simplical [MMSS15]. simplicial [MHS16, PR17a, SC16]. Simplification [ZXL17]. Simplified [CVM^{+}19, HE15, LWWY18, MF17, CNW17]. simply [LDL^{+}16]. simulate [CG15, DA17, HM19b, QM18, RFGSV15, ZWUR16]. simulated [YDCK16, ZD17]. Simulating [KS18a, LP16a, AJVH17, Cap18, CL17, CWB^{+}19, CGS15, DvB17, Don15b, ELH^{+}16, GLS15, GWYS18, HXLL15, HJY19, Kwo19, LYB18, MAK15, Moo17, NRZS17, OT15, OD17, PZNG15, PGCG18, QSB18, RL18, SHKL16, STKL19, SMA^{+}16, SDH^{+}16, TK12, TK15b, TND18, TKP16, WMY16, WVB19, dTP16]. Simulation [CS17b, FBM16, GFA^{+}16, GZLH19, GBS15, LSD^{+}17, LM15c, Mac16, MDW18, MM18, OMYvdP^{+}15, PST19, RVK^{+}18, SSVL18, SB18, STW16, SDM^{+}17, SSX16, TM17, dlAC17, AVT17, AAB^{+}16, Ama18, BBKS16, BBMN18, Bar19, BST^{+}18, BT19, BS15a, BR15a, BGJ^{+}15, BKR19, Bra16a, BJ16, CGSS18, CRW16, CB15, CL16, CTJ^{+}17, CPSF17, CH17, CD17, CLB^{+}16, CSCM16, CLGA17, CB18b, CELI15, CMR^{+}16, CLNH15, DLLV17, DD17a, DKPC15, DPW^{+}15, DL17, DHH^{+}18, DH18a, DMS17, DJL^{+}19, DHC16, DEZ16, ESGS17, ESHA16, EKSS15, FDKI17, FG17, Fon16, FK19, FFJT16, FC19a, FRW16, FKY15, FP18, GB15b, GB15a, GWB^{+}15, GV18, GZY16, GPAO^{+}18, GDFL17, GD19, GRS15, GH17b, GBU15, GFW16, HYK^{+}16, Han19, HIN^{+}16, HN17b, HLML17, HY17, IM17a, JKM19, JL19, JD19, KM17, KBK15b]. simulation [KPJ18, KLNH17, KLRT15, KK17b, KS18b, KHC^{+}16, KJ17b, KYPK15, LYCC17, LRA17, LPW15, LZ15a, LWY17, LCLY19, LXL17, LLY18, LLW19, LTB16b, LD15, LLA19, LLB19, LPBR15, LAA16, LDB19, LSYF15, LWC17, LWTF19, LMC19, LEB^{+}17, MC18, MWD16, MNG15a, MPT16, MG15b, MNO^{+}17, MS15c, MW16a, MBM^{+}18, MN16b, MN18b, MRK15, MMBP19, MBD19, MHGL19, MTJ17, MTJ18, MLL18, MZ15, MM16d, MOR18, NOM^{+}17, NYNYM15, NMJFM19, PC16, PGGW18, PYAG19, Pis18, QWXZ17, RKL18, RBY19, RS16b, RWG18, RW15a, RMF^{+}18, RBGV15, SNSG16, SKF16, SD15, SDJU15, SAK18, SSC^{+}16, SMG19, STKH15, SP16b, SCS16, SBH19, SHW17, SF16, Str17, SP16c, SK18, SST^{+}15, TCA16, TGS19, TGY18, TASA19, TT19, TK16, TC15d, VMM19, VV16, VBG^{+}17a, VK18, Vre17, Vre21, WDS15, WSP17, WSOW16, WG19]. simulation [WCCB16, WHT18, WKHT19, XYPT16, XDvW17, XZZ15, XL17b, XR17, XS19, YCBC15, YLZ^{+}19, YSW15, YSWS16, YSWW16, YFJ17, YFJ18, Yas17, YM19, YCS^{+}17, ZFPB16, ZLY15, ZB15, ZDGW16, ZW16, ZZZ17, ZHLZ18, ZOY^{+}19, ZQCT15, dJRP^{+}15, dlCGCA17, vdLJLV16, CWS18, FNP17, LLM17, TABR17]. Simulations [BGRC19, CBS18, Gan15, AWS16, AMXJ19, ALM15, AT18, AG18, BZ19a, BT17a, BLL19, BLL20, BCD^{+}15, BFI^{+}16, BL18, BBB^{+}16, BCB15, BI16, BPS16, BPS17, BBW16, BLK15, BVM17b, BLK19, BLJ17, BPM18, CGQ18, CM18a, CCBdL15, CCdL15, CDM18, CKT17, CS18a, CGJ16, CC16c, CSK^{+}16, CSB15, Cos16, CvKH16, DMAM15, Dav10, Dav15, DM16, DG16c, DB16a, Don15a, DCBK15, DBMB15, DD16b, EMM^{+}18, EFHZ17, ED16, EH18, Fan16, FHS17, Fed17, FPASS16, Fer17, FHE15, FHA16, FKS19, GZM^{+}17, GK19a, GSL18, GDS^{+}16, GBC^{+}18, GJ15, GEZK16, GSS15b, HRJ^{+}16, HBC^{+}16, HWH^{+}16, HTZG17, Heu17, HLS19, HSB16, HHA16, HHL19, HLY15, HLQ16, HLSY16, HMRG16, Ido16, IG15, IBML16, ID17, JSP16, JYY18, JLKF17, KQB18, KHTZA16, KL16, KZR15, KG15, KRK^{+}18, KSI17, KV16, KS16d]. simulations [LM15a, LY15a, LKK17a, LHO^{+}19, LBZA16, LHMB18, LWL18, LGD17, LHY17, LZL^{+}17, LRZ17, MLI17, MCL19, MS16b, Mar19, MC16, MD16, MGPG19, MBBKTH17, MAH16, MSP15, MMW15, NdlLPCC19, NCP^{+}17, OM19, OLD^{+}16, OCSC18, PD19, PP18a, PN17, PP17, PDS15, PBC^{+}17, RLV16, RS17, RKH15, STHW17, SCL19, SS15b, SMD18a, SKG17, SSA17, SWHK15, SMOM^{+}17, SLdTV18, SMAG17, SSL^{+}16b, SHP^{+}16, SKC17, SR18, SAOW17, TBC^{+}16, TLH15, TPT16, dCPDC^{+}17, TRL15, TS17, Tou18, TEP19, TSR15, VWV17, VKE^{+}18, VBF15, WNW^{+}19, WTS^{+}17, WDGW17, WKSS15, WRL18, XWW^{+}16, XXR18, YFC19, YKMM19, YC16, YXD^{+}16, ZV16, ZYW16, ZN18, Zil15, ZPE^{+}16, ZRE16, dLGT^{+}17, vdKK16, PD17, PEVG18]. simulator [VBG^{+}17b, WLC15, YSLY19]. simulators [MRA16]. simultaneous [CB19, GGW17, SLC^{+}18]. sine [CJWS19]. sine-Gordon [CJWS19]. Single [FCW^{+}18, MT18, PS17, ZY17, ZYD^{+}19a, AJP15, ASWvD19, CFSN18, CFST16, DG18, DPK17, EMS^{+}19, FS19a, Hig17, JZ16, LFDP16, NBMB19, RMK15, TB19]. single- [DG18, LFDP16]. Single-cone [PS17]. single-event [Hig17]. Single-material [ZYD^{+}19a]. Single-mode [FCW^{+}18]. Single-node [ZY17]. single-particle [DPK17]. Single-phase [MT18]. single-stage [CFST16]. single-step [CFST16, JZ16]. singletrace [JHPAT17]. singly [ST17]. singular [EG17, GRMK15, KVKS19, MW15, NL15, POSB16, SO17, Tsa15, Tsa16, WHCN17, ZG19]. singularities [SDW16]. singularity [FQZNZ18, GZ17, OvdHVH16, Pru18, ZSL^{+}19]. singularly [CAA18]. sinks [RRL19]. six [HTvdH^{+}19, PS14, PS15a]. six-equation [PS14, PS15a]. six-moment [HTvdH^{+}19]. sixth [CCM17, HC18a]. sixth-order [CCM17, HC18a]. size [CS18a, EMSS17, LN17]. sizing [GPAO^{+}18]. Skeletal [DDM18]. skew [CARdN19, GWE^{+}15, RÖS17]. skew- [GWE^{+}15]. skew-symmetric [CARdN19, RÖS17]. skewed [OS16]. skewness [BMC18a, DvW15b]. skin [DMRB19]. sky [BPF^{+}16]. slab [Sch16a, Sch16b]. slabs [DBD^{+}17]. SLEIPNNIR [PC19]. slender [LC15]. slice [YSC^{+}17]. sliding [ZL15b]. sliding-mesh [ZL15b]. slip [BC18b, HGW18, KLSF15, LM18, SWLW19, YS18b, dDPG19]. slit [CSC19, DWZ19, LZ19]. slit-shaped [DWZ19]. slope [GK18, KH17, Xia15]. slopes [ST16]. sloshing [ABT16]. slow [WT19]. slowing [CLL17]. slowing-down [CLL17]. slug [KL18b]. Small [GK19c, BFP18, CR17, CHE^{+}17, DH18a, Gam15, KS15a, Par18b]. small-angle [CHE^{+}17]. smallest [ZC18]. Smoluchowski [MST15, MZTS16, SWK18, TMS^{+}19]. Smoluchowski-type [TMS^{+}19]. Smooth [iI15, SGT17, YK15, BG19c, Cap18, FQZNZ18, GS15a, GH19, iI17, MM15, RRS19b, RRS19a, RF18, SGT16, TLB^{+}18, YLLH19, YG19]. Smoothed [FRO17, KRK^{+}18, Kwo19, MDL16, MFG15, TP16a, AF18, BLK15, DD15, DJL^{+}19, Iwa15, LLW19, LS16c, LSR16, MPR^{+}18, MDD^{+}19, ML16, NT15, PKP^{+}17, SMT^{+}16, SE15, TOR^{+}15, TPB16, ZHA17b, ZC19b]. smoothers [YM17a]. Smoothing [RBK19, CC16c, OSP17]. smoothness [HC18b, KC18, XDLX19]. snapshot [LHLL19]. snapshots [URL16]. Sobolev [CM15, CGM18]. soft [JKM19, LHY^{+}19, RBK19, TEP19, WSU^{+}15]. soft-adherence [RBK19]. soft-X-ray [WSU^{+}15]. software [YZW^{+}18]. solar [HGR16]. solid [AASRT17, AGBL15, BJWZ17, Bar19, BLG^{+}16, BB15, CH17, CLM15, CLFL17, DBD^{+}17, DRZ^{+}19, HW15c, HW16c, HF19, KTK15, KLC18, LZ15b, MAK15, MPR^{+}18, NFG15, PAL^{+}16, RHS18, SDM^{+}17, Tre16, VM15, WCVF16, XYF^{+}17, YK15, YS15, ZLY15, ZDGW16, ZW16, ZQCT15, aKT16]. solid-air [DBD^{+}17]. solid-fluid [Bar19, HF19, PAL^{+}16]. solid-fluid-interaction [CH17]. solid-liquid [BLG^{+}16, HW15c, HW16c]. solid-solid [KTK15]. solid-state [BJWZ17]. solidification [BGJ^{+}15, CY19a, OTS17, RKRGW17, RTO15, TYD16]. solids [AAI16, BHKS16, DLY17, DPRZ16, DD16b, GSL18, Heu17, Heu19, HF19, JN19, JKM19, QSB18, RJ19b, SBHS19, TEP19]. solitary [AEAM15, SS17c]. soliton [LY16d]. solitons [MW16a]. soluble [BGN15, SCJ^{+}18, XSL18, dLGT^{+}17]. solute [BGJ^{+}15, SMG19, SZCL18, YM17b]. solute-solvent [SZCL18]. Solution [CLP16a, KE15, LHS^{+}18, NLFM16, ASB^{+}15, AB19, ATF16, And16, AA19, AB17, AKM^{+}19, BHL15, BNM15, BBF^{+}17, BK19b, BFFB17, BGV17, BDKK17, BVMW16, BLVC16, BLVC17, BS19b, BDBEE15, BAK19, BCB17, BTT18, CPV16, Cha16, CMH15, CFF18, CLMZ17, DGHP17, DMAM15, EMM^{+}18, EAAM15, EDC19, Eva18, Fal16, FG18, GMP16, GP16a, GN16, HE15, HO15, HP17, IGQ15, JHPAT17, KA15, KF17, KPP^{+}19, LW15a, LWWY18, LB16, LMTC15, MMNI16, MRM16, MNR17, MR16a, MPFL16, MBS19, MFB18, MFF^{+}19, MSH^{+}15, NDH19, NH17, NPRC15, NKN^{+}17, PYK19, PBA^{+}15, PCA19, Rag15, RMP18, RDQ19, RHS18, RZ15, SZ15a, SR16, SWZ15, SW16, SWPS17, SPRW15, SL16b, SV17, TK15a, TD17, TM15b, TO15, UWH17, VSM17, VST16, Wac15, WL18, WBBC16, WSH19, WZR15, YSY17, ZNS19, ZOG19]. solution [ZSP15, ZLL16b, ZCL17, ZLL^{+}17b, ZSM19]. solution-filtering [EMM^{+}18]. Solutions [Gno17, AEAM15, Azi19, BSWG15, CGTH18, CNQ^{+}19, CSN18, GS15a, GIF18, GS15c, GH19, GY15, HPY18, JL16, JS19, LZ17b, MN18a, MKYZ17, MP19, MM17, NDCB17, PX16, RPK17a, RDM15, RMC15, SWZW19, Sub15, Svä15, VBG16, VGZ18, VCNP18, WDG^{+}17, WSJY16, WBM^{+}15b, XZJK19, YK15, YLLH19, ZLL16b, ZLL17a, ZHS18, ZZX16, ZS17, bWAW15]. solvable [HW15a]. solvated [YX15]. solvation [GZ17]. solve [ALTR17, CE17, Chu17, DLK17, DYL19, DBMB15, FQZNZ18, GAIN19, LZ16, LYPP17, MD18, MR17, PMF^{+}18, RYZ18, SLZ^{+}17, TBHG18, YWHP15]. solved [KW15b, LFRH17]. Solvent [EG18a, SZCL18, ZRT18]. Solvent-Excluded [EG18a]. Solver [ABG18c, APV^{+}18, AGBL15, Ama15, AAD16, AB16b, AdSŚ^{+}15, ABT16, AC17, ANL^{+}16, Bal15, BAGK16, BK16a, BVG^{+}16, BN17, BLM18, BDK^{+}17, BP18, BM19a, BNGI19, BWR15, CBB16, CBC^{+}18, CM16a, CG18a, CCB^{+}19, CNQ^{+}19, CGP16, CTM^{+}16, CRZ17, CM18c, CLP16b, CCGH17, DWGW16, DY16, DS15d, DL18b, DB16b, EGO19, EJMI18, FGLW18, Fer17, GRT18, GRT21, GWC18, GN19b, Guo15, GFW16, Har18, HY16, HH19, HSF17, HSF19, JZSX18, KCSW19, KEJ18, KC17c, KFWK17, LTB16a, LKB15, LO19, LZ17a, lLLNS16, LSTkM15, LDGH16, MCIGO19, MS15b, MH18a, MHL17, MHGL19, MM17, NBT19, NPB19, NGPB19, NNW17, Nis19a, Niu16, NN15b, OC18, OVP15, PKF16, PM19, PCMC19, PR16a, PCBG18, Pop15, QSBY19, RCRF16, SKP^{+}15, SZM19a, SHLG15, SP16a, SYY16, STW16, SYM15, SM19a, SPW18, SGD18, SC16, Sti16, SL16b, SST^{+}15]. solver [SK15b, TCS16a, TWH15, Ter18, VLP^{+}16, VKE^{+}18, VNA15, VSC18, WY17, WSY15, WSHT15, WSY16, WWZ19, WNDB19, WS15a, WCCB16, WHT18, XJ16, YSW15, ZHA17a, ZG18b, ZKG19, ZCSZ19, ZBZ^{+}18, ZSM19]. solver-based [BK16a]. solvers [BSK15, BD15a, BTGM17, BK17b, BGTM18, BK19a, BAVC17, BC16c, BCJ19, CDC17, DS15b, DG16a, DWR18, DL18c, EKV^{+}16, Ein19, GSC19, HR19, JCNK19, Jou15, LPGT16, MVZ16, MM16c, NMM19, PA19, PP17, SLL19, SW17a, SPM16, TKB^{+}15, VBG16, XRMM15, ZAK15, ZSL^{+}19, dPSS16]. Solving [BG19c, CCBF19, GMP15, GMS16, GLTG15, HZE19, IKS19, KR17, LHS^{+}19, MBD15, ZG19, AR16a, Alm19, ADH^{+}16, ADP^{+}17, BM15, VMN^{+}18, BZ15, Bou19, CH19, CZ19b, Chi19, CG18b, CC19c, DLNR18, EFO19, EFO20, EE16, FNNW19, GSN16, GP17, GZY19, GBM16, GPTK19, HW16a, HHCG15, HSC16, HF18, HHY15, IKI15, JG19, KKLS17, KDL15, KJHA19, LMS17, LM18, LW15b, LC16, LK16a, LIW18, LLH19, Lot18, LZW^{+}17, LSI16, MW16b, MW17a, Moh15, MLMM17, MBM^{+}15, Noe15, OLV16, PxRS17, PKJ^{+}18, PLR18, PBBK15, RPK19, SNK18, SWLZ15, SP19b, SL19b, SS18b, SWK18, SS16c, SHW18, SGT16, SGT17, TSH17, TP17, TBO^{+}16, Vab15, WR15, WXW15, WA18, WAZ19, XL16, YZZ19, YHKPF17, YJB18, YM17c, ZZK16, ZD15a, Zha16, ZG18a, ZLGK19, ZQ16a, ZQ16b].

SOMAR [SS15a]. Some [BSN19, FLW16, hGwSzS15, FSWW17, GFO18, JMM19, KD17b, MDH19, Pas16]. sonar [EFHZ17]. sound [KPKGH19, LCLY19]. source [ASB^{+}15, ADOP18, BG19a, BCB15, BT16, DH18a, DMTB15, EG17, FQZNZ18, GKNA17, HS18b, MWZ19, NMM16, NLK^{+}15, NL17, RTG18, RZ18, SY17, Tow18]. sources [BS19a, BM19b, EDW19, POSB16, RRL19]. Space [PCF15, SWHV16, VLN^{+}18, AS15, AP16, ATZ16, AL19b, AMPG19, AHKT17, BCST17, BHJ18, BCS19, BZ15, BK18, BHE^{+}17, BW18b, BTWY15, CLZ18, CV16a, CCGH19, CGG18, DM17b, EHXM15, FNNW19, Fid17, FLW19, GAIN19, GLW18, HKKP16, Har18, HLML17, JW15b, JW16, JX15, JX17, JCWX19, KL15, KLRT15, LS16b, LWY17, lLNS17, LCF16, MD17, MN18a, MCL19, MRZG16, MDDM17, PD16b, RRS19b, RRS19a, SWZ15, SWLZ15, SW16, SWPS17, SW18b, SX15, Tav16, TD16a, TD17, TD18, VGF16, VSC18, YYN^{+}17, YFC19, YLLH19, ZJL16, ZBZT17]. space- [LCF16]. space-angle [KL15]. space-charge [AP16]. Space-fractional [PCF15, CLZ18, CV16a, CGG18, FLW19, JW15b, JW16, lLNS17, MD17, MDDM17]. space-staggered [MCL19]. Space-time [VLN^{+}18, CCGH19, DM17b, Fid17, HLML17, LWY17, RRS19b, RRS19a, SWPS17, SW18b, TD17, TD18]. spaced [CB19]. spaces [GMP16, HW19, KCW17, KC17c, YY16]. spacetime [AM17a, NLK^{+}15]. SPARSE [TUJ19, ABM16, CS16b, HLS15, TPA19, WTGC16, ARG^{+}17, ATM^{+}18, AM18, APLK19, BSN19, BGG16, BGHK19, CQ15, CJC19, GNZ18, HLTC18, JW15a, JL17b, KS16b, KJHA19, LZT17, LCCZ19, MJ17, PSP16, SLL19, SS17b, TCA16, TWF19, TE19, WCN15, WKSS15, Yeo19]. Sparse-grid [CS16b]. Sparsifying [LY19, LY16d]. Sparsity [KMD^{+}18, BKL17, SGP17a, YLBL16]. Spatial [CNG99, MDMS18, AA15, BB15, CNG17, DJL^{+}19, FS16, GZY19, KKL15, LPWK15, LZ16, LN15, LK16a, LMG19, NSB15, SP16c, TC15c, VLTPS16, WMY16, WKOE17, XTS^{+}16, Yan16a, YM17b]. spatial-stochastic [KKL15]. spatially [LM15d, MPT16, NHM17, SSVL18]. Spatio [Han16]. Spatio-spectral [Han16]. spatiotemporal [SLZ^{+}17, Tie18, TL15]. spatiotemporal-adaptive [TL15]. SPDEs [KKL15, OB17]. Special [KHP15, KZ17, Kat16, KZG16, TM17, KS15a, LDT19, MRRRF18, WT15, ZT17]. species [CCZ15, HYK^{+}16, MN16c, SGA^{+}15, TC15a, TKC15, TCSM15, WB17]. Specific [PWP15, BFI^{+}16, ISP^{+}15, PVFN15, ST18b, YDCK16]. specification [SD17]. specified [JG19, SAF^{+}19]. spectra [BKKY19, Roy15]. Spectral [ATM^{+}18, CB18a, DMSC16, DZC16, GHM15, GCI19, GA18, LWL17, MK17, MDDM17, NS16, PMS19, PGH15, SZW^{+}16, SX15, TST^{+}15, Tre16, WB16, YJB18, ABP^{+}16, AABD15, AB16b, ALM15, ABR16, ALT17, ALMJ15, BZ19a, BEJ15, BZ16a, BA15, BDBEE15, BGM16, Cai16, CZW17, CL16, CJD^{+}17, CDDL19, CXH15, CLC16, CH17, Cho15, Del15, DY17, EKEB16, FFM19, FBM16, FSM16, GM19, GWWC17, GXX17, HSM19, Han16, HM19c, HB15a, HB15b, HSF17, JAH19, JB15, KC17a, KADE17, KP15c, LPG18, LP18, Li17, LPR18, LMBZ15, LZZS15, LTXB17, LYPP17, MDVM16, MS16a, MG15b, MJ16, MDM^{+}15, MP19, MP16, MJ17, MA16, MSP15, MSP16, MM17, PKF16, PG17, Pas16, PR17a, RBY19, RGW16, RN18b, SZM19a, SZM19b, SHLG15, SC18a, SO17, SMD18b, Sov16, SGT16, Sub15, TH18, TO15, URGT18]. spectral [VSM17, VPV^{+}17, VK16, WLM15, WZ15, WLW^{+}18, WC19, WCL15, WG15, WSB19, WZRZ15, WZR15, Wu19, XXR18, YC17, YDLC19, ZK15, ZL15b, dlHC16, ST17]. Spectral-collocation [LWL17]. spectral-element [CL16, HSF17, Sov16]. spectral-Galerkin [MS16a]. spectral-like [BZ19a, LZZS15]. spectral/ [MSP15]. spectral/hp [MSP16]. Spectrally [Cho19, ZMF15, CC19c, MPFL16]. Spectrally-accurate [Cho19, ZMF15]. SpECTRE [KFF^{+}17]. Spectrometer [SNB^{+}15]. spectrum [FW17, HR19, HLTC18, LW17d]. Specular [LFRH17]. speed [AIP17, BMR^{+}16, BTT18, Ger17, GEZK16, GP16b, KYW^{+}16, KYW^{+}18, MSB^{+}16, PUA^{+}15, QSB18, TR17, WNDB19, ZWG17]. speed-up [TR17]. Speeding [WKSS15]. speeds [TPB16, XDvW17]. SPH [LLW19, LSR16, CZL^{+}15, CMDL18, CLV19, KM15, LHA15a, LHW^{+}17, Mon19, OMLdL16, OSP17, SiI16, SiI17, VM15, VL15, ZBH^{+}18, ZD15b, ZHA17a, ZXW^{+}19]. sphere [BKKJ17, BMC18a, CT15, Cap18, DFM17, GP16a, GPS17b, HSM19, HK16a, iI15, IDSG15, KBK15a, KC17a, KGT15, LP18, LZHM19, MJ16, SP16a, SW18a, SGC18b, SMD18b, SHW17, SSZC19, WBBC16, WCVF16, XWB15, YSW15, YP17, Zau16, ZL15c]. Spheres [HHK15, CV18, IML15, MS15a, PGCG18, SGN16, SFP16, Vre17, Vre21]. Spherical [Ike18, BBKS18, BHGK18, DBSS^{+}19, FCW^{+}18, GT19, Sub18, TC15b, TKF17, WLM15, YGJ18]. spherically [LS19a]. spheroidal [HXB15, LC18]. spin [TPTT18]. spinning [AMM^{+}15, CWJ18, WAF^{+}19]. spinodal [Tav16]. spline [Ein19, FBF15, FGLB16, LW17c, SLVE18, WR15, YZT^{+}18, ZG19]. splines [CZBC^{+}18]. Split [GWK16, AMH^{+}18, BPL19, DGW18, DDM^{+}19, GNK18a, GNK18b, HZL^{+}15, HMRG16, KTK18, KTK19, MS15b, MMB18, SWHK15, WMM^{+}18, BM15]. split-explicit [DDM^{+}19]. split-step [HMRG16, BM15]. Splitting [SLL17, ZZT^{+}16, AMP16, BG19a, BNK18, BNGI19, BCM15a, BND16, CGQ18, CSS17, CGS18, CJL16, CLZ18, CKQT15, CLX15, CEF15, EO15, FYO^{+}15, GO16, GKNA17, GZ18, HTBG15, KKS16, KO17, KV16, LSL15, LHS^{+}19, LZL^{+}17, MM16a, MLM18, MBD19, PL16b, PGM17, PMB18, QHZ^{+}15, RFPSSA18, RS15b, Rua18, SZY16, SLL16, SX15, TT16, TL15, VST16, tEDKT17]. splitting-based [tEDKT17]. splitting-free [KO17]. splittings [BMT16]. spoof [Fuj19]. spray [CPdS19, DAO17, MZ15]. sprays [SDM^{+}17]. spreading [BDG^{+}17, JJS15, LWC17]. spring [iI15, iI17]. spurious [AAL15, DK19, MSG18a, MLB18, ZW15]. squall [MG15b]. square [GLZ19, JL15, MH19]. Squares [CNW17, NW17, NT19, NWFT19, BVG^{+}16, BtTBI18, Blo17, CCFC19, CBA17, CZL18, LYZ18, LJ16, MNR19, MAP17, PLPR19, SX16, SLdTV18, TMWF18, TMH16, VLN^{+}18, ZTBW19, ZNX15, dTP16]. squares/fictitious [HGW18]. squircles [LB15]. sSSA [DJL^{+}19]. Stabilisation [XRMM15]. Stabilised [SL17, EKEB16, Fer17]. Stability [ACJ17, Ani19, CSK^{+}16, DDM^{+}19, DDV18, GSS15a, KL16, MBJ16, MBNJ16, QN19, SZM19b, SwS16, VBG^{+}15, ZOG19, AA15, Ata15, BK17b, BK19a, BR15b, BR16, BHFB19, DLS15, DKK15, FWK17, GZY19, GR15, JS19, LH17a, LS16a, DV17, MF16b, NR17, O’S15b, PJC16, Ram18, RS15a, RS18, SSZ19, SC18a, WSH^{+}17, WSB19]. Stabilization [Sov16, ALMJ15, BTD16, BK19b, BNK18, BC16a, CG19, CLGA17, DRM15, EMM^{+}18, Kla15, LW17a, LMB19, MSK18, Mas18, SF18b, Ter18, VV19]. Stabilized [ASS13, ASS17, DJLQ18, MNG15b, SSC^{+}16, STV18, BHF15, DGMT17, FMPT18, HMFJ18, MVK16, SDMS17, SL18, WDT^{+}19, LWZ19]. Stable [AMH^{+}18, BGN15, BJK17, CNG99, KHTZ19, MDT16, SKP^{+}15, Sto17, WF17, ZN16, dSPDH15, BHST17a, BHST17b, BHST18, Beg15, BCJL17, BKR19, BC16c, CNG17, Cha18, CW19, CJYZ15, CS16c, CS17a, sCYxL^{+}18, CLS^{+}18, CKQT15, CLvS17, CGJ19, CCPdL19, CHD^{+}18, DRZ^{+}19, DCCH19, DC18b, DWGW16, DWGW17, Don15a, DS15c, DB18, DW19, EMS^{+}19, FCL17, FFJT16, FPV18, GN19a, GBCF15, GBCF16, GHL^{+}16, GGT18, GTG15, GX15, HW15a, HHR15, HN17b, HJY19, IC17, LTB16a, LM18, LYM19, LHB^{+}16, LSZ18, MLI17, MBHS17, MDMS18, MXL16, MTD15, NYD19, NN15a, NN17, NL18b, OLDN17, PCN15a, PCN15b, PR19, PND16, PKK18, PMB18, QN19, Ren19, ST18a, SW17a, SLH18, SBHS19, SPB18, SYY15, SLL17, SN19, Tav16, Tie18, UY19, VV16, WJD16, WWGK17, WWGW18, WG15, WG16b, WDGW17, WCCB16, WSF17, Xia15]. stable [YC17, YDLC19, Yan16b, YH17, YLD19, YD19, ZYSW16, ZW19]. Stage [PP17, BJO18, CSS17, CFST16, DSX19, DL18c, FPASS16, LZT17, PXLL16, RFPSSA18]. Stage-parallel [PP17]. Staggered [CCKQ15, DCCH19, AB17, CCM17, DDJ19, GH17a, GZLH19, KSSL18, LHMB16, LWWY18, LMMS16, LCF16, MCL19, MO18b, OLDN17, QWZ^{+}19a, SO15, SZF15, TLQ15, TD16a, TD17, TD18, TRLK18, VW16, Vre17, Vre21, Wil18, YYL16]. Staggered-grid [DCCH19, GH17a, LHMB16, LWWY18, SZF15, YYL16]. staggered-mesh [Wil18]. standard [Fan16, FST15, STG17, VWV17]. stars [HSF19, Lau17, RLP16]. start [Nis19b]. start-up [Nis19b]. started [GWC17, SHW17]. startup [YKMM19]. state [BJTZ15, BJWZ17, BST19, CGTH18, CKK18b, CZL18, CL19a, CDOY19, CCM17, CMW16, DY19b, EFT15, FAY19, HH19, HWA19, KTN15, KS18b, LZ17b, MC15, Noe15, RZOZ19, SE16, SWZW19, TYD16, TST17, TCL15, XZT18, YJM19, ZJLC15, ZZX16, ZS17]. state-based [XZT18]. state-dependent [BST19, CDOY19]. states [ATZ16, ALT17, BKKY19, GLZ16, GZ18, LY17, MCIGO19, OC18, PDS15, Rua18, RKH15, SWZ17]. Static [AF18, DG16b, FCL17, KKLS17, KPJ18, MDD^{+}19, RSB15, ZSO^{+}19]. stationary [ACCCD^{+}17, ALT17, DCBK15, LZ17a, LYZ15, MFF^{+}19, RRL19, TMS^{+}19, TSST16, ZFPB16, ZL15b]. Statistical [LKK17b, RS15b, AKK^{+}19, CSN18, PRvdL18, RL17, Shi19, VCNGP15, ZLX17]. statistically [CM18b]. statistics [DY16, FKF17]. status [MSV^{+}16]. Steady [TYD16, AD15, AA19, CV17, CGTH18, CKK18b, CZL18, CL19a, CCM17, DKPC15, DDV18, DY19b, EFT15, EMS^{+}19, FAY19, HY16, HH19, JL16, KA18, LL16b, LLWJ18, LLSJ19, LZ17b, MC15, MH18a, NdlLPCC19, Noe15, PQR17, RKH15, SE16, SWZW19, TST17, XRMM15, YT19, ZJLC15, ZZX16, ZS17]. Steady-state [TYD16, CGTH18, CZL18, CL19a, CCM17, FAY19, HH19, LZ17b, MC15, SE16, SWZW19, TST17, ZJLC15]. steady-state-preserving [DY19b]. steep [SWMD17a, SWMD17b]. Steepest [MH18b, TP16b, FSWW17]. Stefan [Gro18]. Steklov [DDV^{+}15, HS17a]. stella [BPL19]. Stellar [Lau17]. stellarators [MCIGO19]. stencil [GEZK16, PG18, ZRW19]. stencils [CCFC19, Nis19a]. stenosis [LSCC19]. stenotic [GZM^{+}17]. step [BH18, CC15, CG19, CFST16, DvW15a, DL18a, EMSS17, FW18, HPY18, HC17, HMRG16, JZ16, LM19d, MBSS15, PKK18, SP16c, VL15, WWZ19, WBM15a, ZRW19, BM15]. Stephen [ZJS15]. stepping [BDZ15, CLvS17, CLP16b, DNBH15, EARA15, GSK18, HLJ^{+}19, JMM19, KCHW19, LW17d, LWY18, LDSM19, LWJV19, MM15, Par18a, PR19, QB16, RGPS17, RKB19, Tav16, Tie16, VLV19, XXR18, Yi18, ZZDB15]. stepwise [ARG^{+}17]. steric [DWZ19]. Stiefel [BTD16]. stiff [BP18, CB15, CTM^{+}16, DY19b, HS18b, LTR17, PBKK17, SXBB15, SYM15, TWN15, TM15b, Ver19, ZW15]. stiffly [RT16]. stiffness [AM17b, WHCN17]. Stochastic [AMK17, BHS^{+}18, CL18, CHLZ17, HKJ17, LPWK15, PTMF18, RL18, SE16, ATM^{+}18, ADH^{+}16, AÁPB17, BCSK17, BLK19, BV18, CHZ16, CSN18, Cot16, CMR^{+}16, Dav10, Dav15, DH18a, DJL^{+}19, DEZ16, EDC16, GHV19, GHS19, GH17b, HFND18, HHCG15, HMBH15, HJZC17, HL15b, HJ16, HJS19, ISST18, JL15, JL18b, JL17c, JS17, KM17, KKL15, KK17b, Kou16, LYLK17, LJZ15, LZT17, Liu19b, LM19d, LM15c, MPT16, Moh15, MFB18, NW15, Opp17, PND16, PDN19, PE15, SSDN15, SS18a, SS15b, SGA^{+}15, TWN15, TPT16, TT17b, TAJ^{+}17, VLAB18, VCNGP15, WN18, WHCN17, WH16b, WTX17, XZZ15, YYL18, ZLGK19, ZLX17, dlCGCA17, ACGR15, HSB16]. Stokes [CDN17, HTMP17, TXKvdV15, TXKvdV16, AD15, ALKZ16, AB18, ALL18, AB17, BTD16, BC16a, BTVB15, BLJ17, BKR19, BPD19, BHF15, BC16c, BCJ19, BC16d, CGS18, CHOR17, CS16c, CYL^{+}16, CYYL18, CCKQ15, CGRV17, CV18, CCM17, CLP16b, DRZ^{+}19, Du18, FWK17, FBW16, FZ19, GN19a, GNK18b, GTG15, GMT19, GSMR18, HPY18, HW15a, HTFL18, HM16b, JPSX18, KML18, KLGO18, LM18, Ler16, LXC^{+}15, LZB^{+}17, LT17a, LHMB18, LYD19, LM16, LR19b, LY16a, LFAR19, LRGO18, MS18a, MS18b, MLM18, MPFL16, MS18c, MS17, MBBKTH17, MHS16, MR16b, MN18c, NGPB19, NL15, NN19, OT15, OvdHVH16, PST19, PG17, PXLL16, PX16, PJC16, PCN15a, PCN15b, Pea15, PND16, PDRB17, PBBK15, PMB18, RDM15, SHLG15, SMS16, SLB^{+}16, SLY16, STG17, SE16, Stü15, Stü17, SZCL18, Svä15, TD16a, TD17, TB19, TMH18, UL16]. Stokes [UY19, VV19, WY17, WR15, WZ18a, WWZ19, WCVF16, XWW^{+}16, YXW19, YC17, YDLC19, YS18b, YTW15, YT19, Zha17c, ZLFW18, ZCL19]. Stokes-like [GSMR18]. Stokes-residual [BC16a]. Stokesian [LRZ17, WB16]. stokeslet [Smi18, Cor18]. Stokeslets [CV15]. stopping [RMP18]. storage [CB15]. straight [LBZ16]. strain [LK17, LHY^{+}19, SY18a, TBO^{+}16, WS16]. strand [KW15a]. strategies [BCB17, HD15, JW15a, KRFV16, KJHA19, Kou16, LJT16, PQR17, SSN15, WED15]. strategy [AM18, CPX19, CCdL15, CGC17, DC18b, DBMB15, FC16, HX16, JYY18, JBLO15, OMYvdP^{+}15, OLD^{+}18, PK16, RRMF^{+}19, SX16, SDM^{+}17, WTL19, YKMM19, YK18, ZFZL15, ZLL16b, vLtTBI17]. stratification [CDV17]. stratified [BNM15, Cai16, GDA16, Liu19a, MTD15, MDT16, RS16b, SHLG15, WLM15, vOMB17, SS15a]. stream [ZH19]. streamer [DBMB15]. streamfunction [YT19]. streaming [AM19, BDMC15, YSWW16]. Streamline [WHE17, WHEK18, AA19, BK19b]. streamline-upwind [BK19b]. strength [LK17]. Stress [TUJ19, CDL19, CDDL19, DGW18, EFO19, EFO20, Fal16, LEB^{+}17, WVB19, WS16]. Stress-Equivalent [TUJ19]. stress-free [CDL19, CDDL19]. stress-strain [WS16]. stressed [JTR16]. stresses [BJ16, GLMC16, Wil18]. stretched [LSLA16, MMvR18]. strip [BPGS16]. strong [AHHC18, BJK17, CLMZ17, ESGS17, Guo15, LZHM19, RS16a, SYY16, Stü15, XLL^{+}17]. strong-constraint [RS16a]. Strongly [WE15, ANL^{+}16, CLS^{+}18, GC17, Iwa15, LGH^{+}18, SDM^{+}17, TW17, Tsa15, Tsa16, YS15, Zau16]. strongly-coupled [GC17, YS15]. Structural [ADE^{+}17, BQCG17, MN18b, NDH19]. Structure [CJWS19, CCP19, SKO18, AQ19, BD15b, BVG^{+}16, BN17, BHKS16, BQCG17, BCM15b, Buk16, CGSS18, CZBC^{+}18, CM16a, CDM18, CQL^{+}17, CM16b, CYWL17, DG16a, DG18, DFGQ16, DMSC16, EKSS15, ED16, FW18, FLV15, FLT17, FRW16, GLS15, HDF18, KKP15, KLC18, KC17c, LTB16a, LKB15, LC15, LHS^{+}18, LMC16, LLJJ18, LC19, LHY^{+}19, LGD17, LZS^{+}19, LYZ19, LGZ^{+}19, MD17, MRZG16, MMMS15, MAA18, MTM19, MKV^{+}17, NPB19, NBZ^{+}19, PDdG^{+}17, PHÖ^{+}16, PD16b, PME^{+}15, RO16, RLP16, Say17a, Say17b, SMP16, SLB^{+}19, SMA^{+}16, SPM^{+}15, TRLK18, WCH^{+}17, Wic16, WHZ18, XTYL18, YXF^{+}16, ZC18, dTP16]. Structure-preserving [CJWS19, SKO18, CQL^{+}17, MD17]. Structured [AZ19b, Bal15, FGLW18, FLHA17, GBR15, iI15, LSZ18, Lor19, MSD^{+}17, PF15, RSB16, WHEK18, XWZ^{+}18, YFKS15, ZJ18]. structured/unstructured [XWZ^{+}18]. structures [BC18a, BB15, CGSS18, CWWZ17, DCA^{+}16, DJV^{+}18, FBG15, GC17, GBS15, GLG^{+}19, KML18, KH18, LDL^{+}16, LSS16, Liu16, MHJ15, NJ15, NWB19, PEVG18, RPNP18, SSL^{+}16a, SWZ17, TBLM15, VCNP18, ZHWQ18, ZBZ^{+}18]. studies [DD16a, EH14, EH15, XS15]. study [BTA17, CX15, CCZ15, DNOP15, DLS15, ED16, HM16a, Hu17, KMdB16, KGS17, LHY^{+}19, MOAA15, MVZ16, MC17, NH17, PR16a, RS15a, DD17c, SCQP16, Stü17, TK15a, VM15, WMM^{+}18, ZHS18, ZZPH18b, ZZPH18a, ZFL^{+}19, ZED15, ZS17]. studying [BLC^{+}17, KES18]. Sturm [KADE15].

Sub [Wil18, Ani16, BVG^{+}16, BTGM17, BGTM18, BKS18, BJ16, DZ16, DL16, GSS15a, HF19, LW17c, lLLNS16, MAH16, MMW15, PxRS17, SPCH16, YZZ19]. sub-cell [BTGM17, BGTM18, DL16, PxRS17]. sub-cycling [SPCH16]. sub-diffusion [GSS15a, LW17c, lLLNS16, YZZ19]. sub-element [MMW15]. sub-filter [BJ16]. Sub-grid [Wil18, Ani16, HF19]. sub-grid-scale [MAH16]. sub-points [DZ16]. sub-scale [BKS18]. sub-structure [BVG^{+}16]. subcell [BD17, LMB19, Vil19, YK18]. subcooled [JSVD17]. subcycling [SAOW17]. subdiffusion [ADH^{+}16, ZZK16]. Subdiffusive [NHA18]. Subdivision [LDL^{+}16, CZJ17, PXXZ15, PCX17, PRXC19]. subdomain [vdLJLV16]. subdomain-levelset [vdLJLV16]. subdomains [GDA16]. subgrid [DLLV17, PD17, WB17, TUJ19]. subgrid-scale [DLLV17, WB17]. subject [Cha16]. subjected [CG16]. subluminal [BK16a]. submerged [ZBZ^{+}18]. subordination [NHA18]. subset [CL17]. subsonic [CPS17, FBJS19, PWC18a]. subspace [BTD16, CN16, GRT18, GRT21, Par15]. subspaces [AdSŚ^{+}15, CELI15, GWC17]. substrate [FKY15, HJY19]. substrates [LD15, YFKS15]. subsurface [AVT17, DGW18, JD19, ST16, TE19, TV19, TD16b, TBLM15, YNW17]. subsurfaces [DGMT17]. subzonal [SO16]. subzone [SO15]. successive [YCPD15]. such [ELH^{+}16]. Suitable [DB18, CS17a, GM19, Sti16]. suited [MBM^{+}18]. sum [EMZ16, Ike18]. Summation [MN04, MN17, RÖS16, CHD^{+}18, DBZ17, DCCH19, FN17, GWK16, GKE15, LMN18, NPP15, NN17, NN19, NR17, NG17, NG18, PS15b, RÖS17, Ran18, RWN18, RN18b, SPB18, LKN17]. Summation-by-parts [RÖS16, CHD^{+}18, DBZ17, DCCH19, FN17, GWK16, LMN18, NN17, NN19, NR17, NG17, NG18, PS15b, RÖS17, Ran18, RWN18, RN18b, SPB18, LKN17]. super [Fed17, SLH18, SSM^{+}17]. super-convergent [SLH18]. super-hydrophobic [Fed17]. super-resolution [SSM^{+}17]. superconducting [FBC^{+}16, PK17]. superconductivity [GS15b]. superconductors [LL19a, SKP^{+}15, SSL^{+}16b]. Superconvergence [SZ15b, GSS15a]. Superconvergent [GP17]. supercooled [RW15b]. supercritical [KTN15, PSS17, TK12, TK15b]. supermesh [ADGN17]. supersonic [BCRS19, WG16a]. Supervised [BLK19, PT17b]. SUPG [LWZ19]. SUPG-stabilized [LWZ19]. support [AEL^{+}17]. Suppressing [NT16]. suppression [MAP17]. Surface [BC16d, TP16a, YT17, AAL15, AASPT18, AZ17, ADN19, AS17, AEL^{+}15b, APT17, Ani16, BJWZ17, CCHL15, CSW^{+}19, COV18, CSG17, CDM19, CG16, DDM^{+}19, DDV18, DKTH15, DKC15, EJZ17, FRL15, FMRZ17, FPDT17, FKY15, FPV18, FP18, GOR17, Gen15, GG15, GLMC16, GVTQ16, GP16c, HHA15, HWH^{+}16, HLTC18, HKS^{+}16, JRPPS18, Lau17, LYM19, LC17a, LLY18, LLW19, LS16c, LZ15b, LWC17, LEB^{+}17, MML17, MNG15a, MK15, MDL16, MHS16, NNV19, NWKC16, PRXC19, DM18, QWX18, RZ15, Say17a, Say17b, SR19, SAK18, SL16a, TBO^{+}16, WX17, XYPT16, XX17, XY17, XS19, YFKS15, YCS^{+}17, ZZS^{+}17, ZZKF15]. surface-integral-equation [SL16a]. Surface-tension-driven [BC16d]. Surfaces [EG18a, AASRT17, CLR15, CE17, Chu17, FFW17, Fed17, GA18, GVTQ16, HN17a, HN17b, HN18, KJYC17, LWZ16, PCX17, PKB15, PR16c, PLR18, PLPR19, QS16, RF18, SNK18, SK19c, WXSJ19, XWW17, ZW16]. surfactant [BGN15, DD17a, Gan15, HCLT19, PST19, SA16, SCJ^{+}18, SRS19, ST18c, STV19, XSL18, dJRP^{+}15, dLGT^{+}17]. surfactant-covered [STV19]. surfactant-dependent [Gan15]. surfactant-laden [PST19, SRS19, ST18c]. surfactant-polymer [DD17a]. surfactants [BG16b]. surrogate [CJC19, KSV^{+}15, PKW17, SGC^{+}17, SGC^{+}18a, TCA16, TB18, WL16, ZZ18, ZZKP19]. surrogates [WLL16]. survey [Shu16]. susceptibility [DKTH15]. suspension [CGS15, CV18, FKY15, IML15, TMH18]. suspensions [BRK^{+}18, BLG^{+}16, DKPC15, GLTB18, KQB18, LRZ17, NRZS17, QB16, WB16]. swarm [LLY15]. swarming [GHM15]. SWE [CV17]. sweeping [DG16b, EFT15, EG16, KLWQ17, LQB16]. sweeps [dCMR19]. swept [AW16, MN18a]. swimmer [LHY^{+}19]. swimmers [SCQP16]. swimming [BI16, PBP18]. swirling [ZCSZ19]. switch [DWR18]. switching [HSK^{+}15, Zad11]. Sylvester [HO15]. symbolic [LLD19]. Symmetric [LIW18, BBF^{+}17, CARdN19, FS19b, GPRA18, GL17, LMC16, LS19a, Loh17, RÖS17, RZ17, MRRRF18]. Symmetry [PLL^{+}15a, LWLC17, OV17, VW16, VW18]. Symplectic [EBQ15, MW17a, SCN^{+}17, Web14, ZJS15, CHZ16, CQL^{+}17, CHLZ17, GAN^{+}16, GZY16, LW15b, SL16c, Sun19b, Tao16, TPTT18, ZZT^{+}16]. Synchronized [LK16b]. synchrotron [BGD19]. synthesis [GLG^{+}19, KH18, MCS16]. synthetic [KH15, PBCR19, SG18]. system [BMT16, BZ16a, BPTA16, BQRX19, CCZ15, CV15, DDD17, DLM18, DLMDV18, EL17, EDC19, FDS^{+}15, FFM19, FK19, FS15, GHS19, GX15, HK16a, IM17a, KKS15, KKS16, KLGO18, LCK16, LMKS15, LMSK17, LAEK18, Liu19b, LWX19, LLLN18, LRGO18, MDVM16, MN16a, MRN16, MP16, NBH18, SHLG15, SOS19, SH19, TC15a, TKC15, VLAB18, VLN^{+}18, WCVF16, YM19, YLLH19, ZM16a, ZWYW18]. Systematic [LYLK17, MPP15]. systems [AQ19, AGRB18, ABR16, BN17, BT16, Blo17, BW18a, BV18, BDV17, BRK17, Cac15a, Cac15b, CGS18, CFG16, CB15, CTJ^{+}17, CQL^{+}17, CLS^{+}18, CM19, CCB^{+}19, CLM15, Cos16, Cot18, DL18a, DB16b, EBQ15, EFT15, EJMI18, FDKI17, FOF15, FBF15, FS19b, GAN^{+}16, GKMS17, GS18, GK19c, HKKP16, HM16a, JL18c, KBK15a, KNS15, KM16b, KJHA19, KML18, LMS17, LSM19, LS15b, LPB17, LYZ18, LXL17, Lia16, lLNS17, LSP^{+}18, LJT16, LW17e, LM19d, LPBR15, MSK18, MDH19, MW16b, MW17a, MHH19, MMvR18, NMA15, NL15, NW17, NT19, NWFT19, NW15, NN16, Nor15, PxRS17, Pan15, PTT19, PYAG19, PT17b, PSP16, RRL19, RZ17, Ren19, RB15, SLL19, SS18a, SCL19, SL18, SMS16, SP19a, SPP16b, SMSR18, SS18c, TWN15, TPT16, TT17b, TZSS17, Ver19, VNA15, WN18, WE15]. systems [WCN15, WHCN17, WW18, Web14, WTX17, YS18b, ZNS19, ZJS15, ZD15a, ZCZ19].

T [AMXJ19, DS15a]. TABI [CG18a]. tailored [FHA17b]. Takahashi [Zil15]. taking [LYDB17, SSL^{+}16a]. tallies [Swe18]. tangent [KPP^{+}19]. tangential [AKM^{+}19]. tank [XYPT16]. target [ST18b, ZFZL15]. Targeted [FHA17b, FHA16, FHA18]. targets [ALM^{+}17, Dod17, HZ17]. TAS [GRS15]. task [KFF^{+}17]. task-based [KFF^{+}17]. Tau [ASB^{+}15, BZ15, BDBEE15, VK16, KRFV16]. Tau-estimation [KRFV16]. Taylor [BR16, BR15b, CVG19, GR15, MCIGO19, OC18, RLV16, YYL16, YHKPF17]. Taylor-Green [BR16, BR15b]. Taylor-series [YYL16]. TDIBC [DSSP18]. tearing [DNOP15]. technique [ALM^{+}17, AMP16, CM15, DA17, DWGW17, ESGS17, GDFL17, GRS15, GKE15, GPTK19, Hig17, KHTZA16, KSV^{+}15, KPJ18, LDL^{+}16, LYL19, MNG15a, Ram18, SGL18, SCJ^{+}18, Sir19, TBO^{+}16, ZWUR16, ZXDL17]. techniques [CSN18, GWB^{+}15, GHF19, HR18b, HHR^{+}19, KA18, DV17, MLL19, MDM^{+}15, QWXZ17, SDJU15, SGC^{+}18a, ZNS19, ZSL^{+}19]. technologies [ADE^{+}17]. telegraph [AR16a, Ata15, HB16, KA15]. telegraphic [GCVCHH18]. Temam [BPD19]. Temperature [DJD^{+}17, AJW17, Gen11, Gho17, LL19a, LLB19, Mar19, NF17, SSM15, TCS16a, ZV16]. temperature-dependent [Gen11, Gho17]. temperatures [DJD^{+}17, Lap16, TTN^{+}16]. Tempered [SMC15, Beg15]. temporal [DL18b, GH17a, MSG18a, MM16b, MS18c, PKK18, SKF15, WMY16]. Ten [MKC17]. Ten-Moment [MKC17]. tension [AAL15, AASPT18, APT17, Ani16, BC16d, CSW^{+}19, CDM19, CG16, EJZ17, FMRZ17, FP18, GOR17, HKS^{+}16, JRPPS18, LYM19, LLW19, Say17a, Say17b, SAK18, TP16a, YCS^{+}17, ZZKF15]. tensions [WX17]. Tensor [CRZ17, DM17b, GMS16, MZTS16, VBL^{+}16, ABFR16, BDMC15, BDKK17, BKKY19, BVT18, DM19, EL17, Fal16, GKMS17, GJ18, GL17, HLL^{+}18, LB15, Loh17, LY15c, NNV19, OS16, PP18b, RO16, RBD17, TWN19, VMK^{+}19, WN17]. tensor-based [OS16]. Tensor-product [DM17b, DM19, PP18b]. Tensor-Train [CRZ17, GKMS17, GJ18]. tensor-valued [NNV19]. tensors [Jer19, LBTCG16]. term [ATZ16, ADOP18, BZ15, BT16, FRW16, GZY16, JLLZ15, MDP^{+}15, NL17, Tow18, WYA^{+}17b]. terminal [Die15]. terms [CCPdL19, DMTB15, EG17, FSWW17, HS18b, LVTR15, NMM16, SL16c, VAD17, XOX19]. ternary [ARF18, Hag25, ZDGW16, ZWYW18]. terrain [AZ19a, SWMD17a, SWMD17b]. terrains [MRK15]. tessellation [SC16]. test [HW19, KDF15]. testing [Ani16, OTS17, ZA15a]. tests [AA19, LP16a]. tetrahedral [AG18, BCST17, MWB^{+}15b, MW17b, TD16a, ZQ17]. tetrahedron [WR16]. texturally [GHP15]. th [LHS^{+}18]. their [Beg15, DLC15, FFW17, HKH^{+}16, JL15, LIW18, MKYZ17, WW19, WX17]. theoretic [LSWF16]. Theoretical [MDH19, PBL^{+}19, VPM15]. theories [MGKG17]. Theory [AQ19, CDC17, CDDL19, DG16c, KHP15, KADE15, KADE17, LLL16, ADFG17, AKZ16, BEJ15, Cen19, GCI19, GYZ19, LWLC17, LW17d, MW16b, NGY^{+}17, OLD^{+}16, OLB^{+}17, RXSG15, RXS16, SDFA17, Sto16, WXSJ19, XZT18, ZLH^{+}17, ZCSZ19, GS16]. theory-based [GYZ19, ZCSZ19]. Thermal [ST15, AABD15, Ani19, BGJ^{+}15, CBN^{+}16, DPW^{+}15, EDK19, FST15, FBJS19, GSL^{+}19, HPC19, HGN17a, HGN17b, HCW15, LMH16, Lap16, LNM15, MRM16, PCMC19, PBA^{+}15, PSN^{+}19, SSC^{+}16, SKO18, SSM15, SPN^{+}19, SS17c, TBHG18, TWM18, WSY16, XZT18, YYY^{+}16, ZCSZ19]. thermal-fluid [YYY^{+}16]. thermal-hydraulics [CBN^{+}16, SSC^{+}16]. thermal-nonequilibrium [SKO18]. thermal-solute [BGJ^{+}15]. thermally [BBMN18, HM17, LLA19]. thermo [HWA19, LMC16, MBJ16, MBNJ16]. thermo-acoustic [MBJ16, MBNJ16]. thermo-hydrodynamic [HWA19]. thermo-mechanical [LMC16]. thermoacoustic [AMJ17]. thermocapillary [LZ15b]. thermochemical [MPP15]. thermodynamic [DZ18, LB17]. Thermodynamical [LW17b]. Thermodynamically [KS18b, BKR19, Don18, LZW19]. thermodynamically-consistent [Don18]. thermodynamics [AZK16]. thermoelasticity [GQC^{+}19]. thermomechanical [AQ19, Heu17, KYKS19]. thermostats [Dav10, Dav15, LS16a]. thermoviscous [DSS18]. thick [BPGS16, SP16a, SMA^{+}16]. thickening [JL19]. thickening-based [JL19]. thickness [DGHP17, dTP16]. Thin [Pes15, AASRT17, AJP15, DGHP17, Fuj19, GLS15, GC17, HJY19, JTR16, KHP17, LVB^{+}15, MTK15, QYF15, SF18a, Xia15]. Thin-film [Pes15]. thin-walled [FLV15]. THINC [LH17b, QWX18, XX17, ZMZC19]. Third [CHY16, DY19b, GZY16, HW16b, WKPS18, YL19, CC15, CHJT17, DY19a, MN15, NL17, PX16, VK16, ZLFW18, ZQ17]. Third-order [DY19b, GZY16, HW16b, WKPS18, CHJT17, DY19a, MN15, NL17, PX16, VK16, ZLFW18]. Three [AEL^{+}15b, Bal15, BGJ^{+}15, GS15a, LMSK17, ABI17, AB16b, APT17, AJW17, BHST18, BKP16, BOA17, CWF16, CC15, CP16, CZJ17, CHL^{+}19, CB18b, CGRV17, CM18d, DS15a, DS15b, DvW15b, FB17, FST15, FPDT17, GGL^{+}17, GT19, GQC^{+}19, GR19, HN17a, HHL19, IDSG15, IM15, JGS16, KF15, KA15, KCW17, KSVB18, KS15b, LGO17, LZS^{+}19, MIM^{+}19, Mar19, MHL17, PHHR17, PCN15a, PCN15b, PR16b, RVZB15, RG15, RDG17, RKRGW17, Rod18, SSVL18, SHKL16, STW16, SWJG19, SSA17, TD16a, TSB^{+}18, Tre16, Vee16, VCNOP18, WK19, WSY15, WHY17, YSW15, YSYW19, ZL15a, ZYW16, ZW16, ZCL17, ZL15c]. three-component [STW16]. Three-dimensional [AEL^{+}15b, LMSK17, AB16b, APT17, BOA17, CP16, CZJ17, DvW15b, FPDT17, GGL^{+}17, GT19, GQC^{+}19, HHL19, IDSG15, IM15, JGS16, KF15, KCW17, KS15b, LGO17, LZS^{+}19, MIM^{+}19, Mar19, MHL17, PCN15a, PCN15b, PR16b, RG15, RDG17, RKRGW17, Rod18, SSVL18, SWJG19, SSA17, TD16a, TSB^{+}18, Tre16, VCNOP18, WSY15, WHY17, YSW15, YSYW19, ZL15a, ZYW16, ZCL17, ZL15c]. three-dimensions [GR19]. three-field [CWF16, CC15]. three-material [PR16b]. three-phase [CM18d, FB17, GGL^{+}17, ZW16]. Three-point [GS15a]. three-scale [ABI17]. three-temperature [AJW17]. threshold [EJZ17, WWX19, XWW17]. thresholding [WLWW17]. through-flow [YTW15]. Through-the-wall [CW16, CW17]. TI [LZL^{+}19, bWAW15]. tight [PD16b, YZT^{+}18]. tight-binding [PD16b]. tightly [TPT16, TT17b, TSN16]. tightly-coupled [TPT16, TT17b]. tiling [Tav16]. Time [ALO18, AMP16, BOA17, BCB15, BG16a, DOO17, FM15, GKNA17, KBK15b, LHY17, MM16b, MM15, MH19, MMMS15, MDP18, SPB18, TRL15, Yi18, AM17a, AW16, Ali15, APT17, ATF16, ADHN15, AL19a, AL19b, AR16b, AEAM15, AHKT17, AWJ17, Ata15, BJO18, BHL15, BZ16a, BG19b, BZ15, BDBEE15, BK18, BGD19, BCM15a, BLK19, BSP18, BHE^{+}17, BSWG15, BFNGDNR18, BDZ15, Bre18, BC16c, BTWY15, CCFC19, CR17, CCdL15, CGS18, CG19, CXH15, CLC16, CHY16, CX16, Che18, CLvS17, COdLL18, CGJ19, Chu17, CLQ17, CLP16a, CC16c, CJH^{+}19, CLP16b, CCGH19, CLMZ17, Cui15, DKPC15, DNOP15, DNBH15, DvW15a, DM17b, DGL^{+}15, DL18a, DSSP18, DLL^{+}17, DL18c, DY19b, DBMB15, EMM^{+}18, ETAG15, EARA15, EN17, FW18, FBL17, FFM19, FNNW19, FLV15, FJLC18, Fid17, FN17, GAN^{+}16, GSN16, GSS15a, GS15a, GZY16]. time [GFC18, GMP15, GP16a, GAIN19, GSK18, GHJ15, GQC^{+}19, GN19b, GHH^{+}16, HPC19, HW15a, HB16, HEPG15, Hig15, HLJ^{+}19, HL16b, HLML17, HSC16, HJW19, HTBG15, IKS19, JSP16, JMM19, JG19, JLLZ15, JCWX19, KKP15, KCHW19, KNS15, KPJ18, KLRT15, LH16, Ler15, LZ15a, LZ16, LWY17, LLWJ18, LHLL19, lLNS17, LR19b, LW17d, LWY18, LIW18, LLH19, LY19, LCF16, LM19d, LMM17, LDSM19, LWJV19, LPR19, MBSS15, MN18a, MWD16, MAM16, MRY19, MP15a, MKB19, MMP18, MM17, MH17, MPMB19, NBT19, NHA18, NDCB17, Nor15, OB17, PLC18, PLHA18, Par18a, PCMC19, PHRA16, PR19, PT17a, PSG19, PTMF18, PME^{+}15, QB16, RTG18, RRS19b, RRS19a, RS16a, RDQ19, RGPS17, RM16, RRD16, RKB19, RN18b, RL17, SXBB15, STEK17, STEK22, SS15b, SWZ15, SWLZ15, SW16, SWPS17, SSM^{+}17, SL19b, Shu16, SW18b, SZ17, SWHV16]. time [SP15b, SPRW15, SC18b, SP16c, SWZ17, SS18c, Sub15, SAOW17, SSZC19, TBHG18, Tav16, TD16a, TD17, TD18, TSH17, TP17, Tie16, TBO^{+}16, VLV19, VL15, VLN^{+}18, VK16, WJD16, WSOW16, WBM15a, WR16, WZ18b, Xie15, XXR18, XY18, XHC15, YLA15, ZZK16, ZTBW19, ZZDB15, Zha16, ZJL16, ZBZT17, ZMZC19, ZLL^{+}17b, ZZH16, ZS19b]. Time- [DOO17]. Time-accurate [ALO18, BOA17, MH19, MDP18, BDZ15, DL18c, EMM^{+}18, FFM19]. Time-dependent [BCB15, GKNA17, KBK15b, AL19a, AL19b, AWJ17, BHL15, BOA17, BSWG15, CX16, Chu17, CLP16a, DKPC15, DBMB15, GSN16, HPC19, HL16b, IKS19, LZ15a, lLNS17, LR19b, PLC18, PCMC19, RDQ19, RRD16, STEK17, STEK22, SS15b, Shu16, Sub15, ZTBW19]. time-differencing [WBM15a]. Time-domain [BG16a, BGD19, CLQ17, GFC18, GHJ15, KPJ18, LH16, LHLL19, MMP18, MH17, NBT19, PR19, SWZ17]. Time-filtered [MM16b]. time-fractional [AEAM15, Ata15, CXH15, DLL^{+}17, GSS15a, GS15a, GMP15, HB16, JLLZ15, KNS15, MP15a, XHC15, YLA15, ZZK16]. time-harmonic [BG19b, CJH^{+}19, DGL^{+}15, ETAG15, LY19, RM16, SC18b]. time-integral [ZMZC19]. time-integration [SXBB15]. time-integrator [LLWJ18]. time-integrators [WZ18b]. time-marching [FLV15, PLHA18, SL19b]. time-parallel [RS16a]. time-space [AHKT17, BZ15, BK18, FNNW19, KLRT15, MN18a]. time-space-domain [JCWX19]. time-spectral [MM17]. Time-stable [SPB18]. time-staggered [LCF16]. time-step [DvW15a, FW18]. Time-stepping [MM15, DNBH15, HLJ^{+}19, LW17d, LDSM19, PR19, RKB19, Tie16, VLV19]. times [LM15c]. timescales [Cos16]. timestepping [KMS^{+}18]. timesteps [CS17b]. Toeplitz [KNS15]. Toeplitz-like [KNS15]. TOKAM3X [TBC^{+}16]. tokamak [BDB^{+}17, BHNS19, FH17, FBC^{+}16, GBC^{+}18, HRJ^{+}16, KHC^{+}16, KYPK15, MP15b, MP16, TBC^{+}16, WSU^{+}15]. tokamaks [LBZ16]. tolerant [AD17]. tomography [KBR17, MS15a, NLK^{+}15]. tool [TO19]. tools [LKK17b, LHGF19, VWV17]. topography [MDBCF17, NMM17, PP19, TDC^{+}19, WHRL19, WG15, YR15]. topological [CDLR19, GYZ19, LR19a, LDHJ15, Par17]. topologically [LWC17]. Topology [CWWZ17, MKV^{+}17, NSL16, QDRB15, YYY^{+}16, DK18a, DK18b, GMA18, LDO^{+}19, LSD^{+}17, SYI^{+}19, ZC19a]. tori [ZYW16]. toroidal [CJH^{+}19, OC18, RKH15]. torques [NJPB17]. Torrey [BTWY15, ZBZT17]. Total [HW16c, BKL17, DLMDV18, HW15c, ZC15, ZLGK19]. trace [OKE17, WLK^{+}16]. tracer [BKKJ17]. tracers [HM17]. traces [ABT17, HLL^{+}18, ZND16]. tracing [DC18a, JH15]. track [RRL19]. tracking [AP16, BMRA^{+}15, BHFB19, CTJ^{+}17, CL19b, Fan19, FL18, Gro18, HM16b, IM17a, uKHGK19, PR17b, SCJ^{+}18, SR18, dJRP^{+}15]. tractable [FK19]. traction [FRL15, LXC^{+}15, MS17]. traffic [HY17]. Train [CRZ17, GMS16, GKMS17, GJ18, MZTS16]. trajectories [LDHJ15]. Trajectory [TD16b]. transcendents [FFW17]. transcritical [KTN15, LMPS15, MLI17, RVK^{+}18]. transfer [Ani19, BJRF18, CMP19, CLG^{+}19, CMDL18, DRZ^{+}19, DPW^{+}15, EMS^{+}19, FAY19, FB15, FYC^{+}18, HPC19, HG17, Her16, HDA^{+}18, ION^{+}17, JL17c, LFRH17, LLB19, LMG19, MRM16, MBHS17, MTK15, MSF^{+}19, NBZ^{+}19, PLC18, PCMC19, STK^{+}16, ST15, Sla16, SJX15, SJXL15, SJX17, VBG16, WSP17, WY19, WCWY19, WB17, WED15, YK15, ZCQ19, ZCQ20]. transferable [EN18]. transform [GZY16, Ike18, JB15, YXX^{+}16, GKE15, GN19b]. transformation [HPV16, KM17, LP17a, MOAA15, MSA19, PX15, VDPP15]. transformations [BG16b, ZX19]. transformed [SV17]. transforms [BSK15, Gno17, Yan19, Str18]. transient [BLVC17, BPTA16, CPV16, LSR16, MIM^{+}19, MOAA15, Noe15, QWZ^{+}19a, SK15b, TBHG18, UWH17, YG18]. transistors [HCW15]. Transition [ABG^{+}18b, ABG^{+}19b, HHZZ19, BKG15, DLP19, FMRZ17, GLZ16, GZ18, HHM17, LS16c, NWB19, PBCR19, RZ15, YR15]. transitional [BS15a, FNP17, XWZ^{+}18]. transitions [CCP19, LJZ15, PKB15]. translational [BKP16, WYLX17]. Transmission [SCS16, BHMS18, DCA^{+}16, DJV^{+}18, DGL^{+}15, HK18b, HSSZ16, JHPAT17, NL18b, SC18b, XJG18]. Transmission-line [SCS16]. Transparent [PE16b, Vai15, BNS17]. Transport [BCST17, SZY16, AAI16, ADN19, AEVW18, ASWvD19, ADK^{+}17, BK19b, BHdD18, BIR18, BBKS18, BTA17, BKKJ17, Bou19, BCG^{+}15, BRW15, BKRB15, CPT16, CP17, CLY^{+}15, CSLL15, CEHM19, CSK^{+}16, CK16b, CCGH17, CCGH19, DBSS^{+}19, DAO17, DS15d, EG18b, FS18, FNNW19, FL16, FC19b, GMP16, GLOP19, GBD^{+}15, GW16, HR18a, HR18b, Hiv18, HL16b, HW15b, JSVD17, JXZ15, KM16a, KFL17, KL15, KBG^{+}15, KGP^{+}17, KLGO18, LMH16, LTKA15, LLS15, LT15, LBZA16, LKSM17, Loh17, MG15b, Mas18, MXL16, MP15b, Moc17, MBD19, MFF^{+}19, OMLdL16, OWKE16, PJE^{+}16, PHRA16, PGM17, PBA^{+}15, PDRB17, PCA19, RFGSV15, RW19, RZZ19, RMC15, RL18, SLL19, Sch16a, Sch16b, SWS^{+}18, SWG^{+}17, SW18a, SMG19, SFT16, SWMD17a, SWMD17b, SWJG19, SU15, SPW18, SSM15, Spe15, TWM18, TSB^{+}18, TSFS17, TYROG19, VST16, Vel19, WWR16]. transport [WB17, WKOE17, WBBC16, WSH19, Wu19, XCX17, YB17, ZA15b, ZHA17b, ZHWQ18, ZM16b, ZCL17, vdKK16, PC19]. transport-velocity [ZHA17b]. transported [BMC18a, Ger17]. transpose [ZD15a, CGJ16]. transpose-free [ZD15a]. transverse [CJH^{+}19, DvB17, ZZW^{+}16]. Trapezoidal [AHKT17]. traveling [AA19, Yas17]. traveltime [bWAW15]. Treating [BLVC17]. Treatment [CNG99, SG19a, CB18a, CNG17, CLL17, DDJ18, HL15a, KPKG15, LFRH17, LS15b, LBZA16, MF17, MLB16, OvdHVH16, PSV18, SMLB15, TLH15, TTN^{+}16, Zha16]. treatments [HHY16]. Tree [JdR^{+}18, HS17b, Jer19, KDPK15, MGBG16, SLA^{+}19]. tree-based [MGBG16, SLA^{+}19]. treecode [CG18a]. treecode-accelerated [CG18a]. trees [SPD19]. Trefftz [BBF^{+}17, LK16a]. trends [PSMPG17]. tri [KNS15]. tri-diagonal [KNS15]. Trial [RSSSE18]. Triangular [MSP19, Bar18, BDZ15, CHY16, CXY19, CLFL17, DK19, GK18, HLL^{+}16, KNS15, KD17b, KL18a, LTXB17, LAEK18, MN15, Pas16, QDH15, QN19, XP15, XYG19, ZLFW18, ZCQ19, ZCQ20, ZPW18, ZS19a]. trigger [PBCR19]. triply [HN17b]. troubled [FS17b, RH18, RH19]. troubled-cell [FS17b, RH18]. troubled-cells [RH19]. TRT [KGT15]. true [RLH19]. truly [SH19]. truncated [BDMZ19, LB15, LT17a, PKN17]. truncating [FYZ^{+}15]. Truncation [GR15, Hwa16, LHGF19, RRMF^{+}19, TPA19, ZFZL15]. tsunami [BHGK18]. tsunamis [dlAC17]. TT-M [YLLH19]. Tucker [LMGG17]. tumor [LTWZ18, TT17a]. tumour [dlCGCA17]. tunable [LWY17]. tune [Ant17]. tunneling [DS15d, HCW15]. tunnelling [LYDB17]. turbines [CGSS18, MBST17]. turbomachinery [dLDG^{+}18]. turbulence [BPM18, CGSS18, CM18a, CDX18b, GZ19, HK15a, JYY18, KH15, KYPK15, LSP19, LT17a, LDB19, MAH16, MSP15, MMPS17, OMYvdP^{+}15, SSO^{+}15, TBC^{+}16, TWN19, VBF15, WVB19, WN17, WMM^{+}18, YWS^{+}16]. turbulence/particle [WVB19]. turbulent [BBKS16, BS15a, BKG15, BFTVC18, CCBdL15, CL16, CM19, CV16a, CCPdL19, ESHA16, FWK18, FNP17, FMPT18, FBM16, HM19b, KYUO15, KTN15, KCS^{+}17, KFWK17, KM15, LE16, LHO^{+}19, LZB^{+}17, LHMB18, LDHJ15, MM16a, MP17, MRK15, MWZ19, OVP15, PM16, PGGW18, PEVG18, PBCR19, PWP15, RWG18, SWS^{+}18, SK18, SWL19, TKP16, UG16, WG16a, WMYG16, WSN^{+}15, XWL^{+}16, XS19]. TVD [Sid18, BR15a, DvW15b, Heu17, ZJLC15]. Two [CHCC18, GN19b, JSY15, LEB^{+}17, RMA17, SAH17, TUJ19, Vab18, ACGR15, AASRT17, AAI16, Ama15, ACJ17, Ani16, ADOP18, BJO18, BAGK16, BVG^{+}16, BHJ18, BXY17, BGN15, BT19, BM19a, BH18, BAVC17, BLS16, BHMS18, BTWY15, BKKRB16, CJWS19, CFSN18, CPT16, CDM18, CS16a, CGK17, CSY19, CCZ18, CLZ18, CCZ15, CS18b, Chi19, CLX19, CEHM19, Cif19, CS17b, CG16, CM18c, CLMZ17, CYWL17, DS15a, DS15b, DG18, DCA^{+}16, DLM18, DGMT17, DSX19, DG16c, DL17, DHH^{+}18, DvB17, DL18c, EH18, EDvW17, FR18, FGL16, FS16, FS17a, FZ19, FG19, FS19a, FLW19, GSC19, GZ17, GN16, HHA15, HTFL18, HN17a, HLML17, HM16b, HTMP17, HCLT19, HC17, HLA19, HTBG15, ID17, IGQ15, JPLL15, JS16, JS17, JJ18b, KSM19, KJ17a, KGS17, KHTZ19, KS16c, KS18b, LVTR15, LW18, LYM19, LPGT16, LMC16, LPR18]. two [LXC19, LZT17, lLLNS16, lLNS17, LSD^{+}17, LDT19, LD15, LSTkM15, LCK19, LLH19, LDGH16, MNG15a, Mar19, MA19, MT19, MDDM17, MDP18, Mue18, Niu16, PST19, PXLL16, PxRS17, PL18, PSB^{+}18, PM19, PM16, PR16a, PAL^{+}16, PS14, PS15a, PGM17, PSV18, QYF15, QSBY19, RWG18, RH19, Ren19, RV16, RTG15, RLH19, RZ15, SZM19a, SZM19b, SG18, SHLG15, SHA16, SRBB18, SYM17, SA19, SX15, SWZ17, SJH^{+}15, SLZ^{+}17, SGP17b, Suz18, TH18, TSH17, TGS19, TND18, TASA19, TT16, TBO^{+}16, TEP19, UWH17, VNA15, VSM16b, WRL16b, WDT^{+}19, WLE17, WHE17, WWGK17, WG15, WKSS15, XSL18, XZT18, YSY17, YFC19, YD19, YM17b, ZMF15, ZTBW19, ZLL16a, ZZ17b, ZWL^{+}19, ZKG19, ZBZT17, ZYD19b, dJRP^{+}15, tEDKT17, YK18]. two- [Mar19]. two-channel [DG16c]. two-component [GZ17]. Two-dimensional [JSY15, LEB^{+}17, ADOP18, BVG^{+}16, BLS16, BTWY15, CJWS19, CCZ18, CLZ18, CLMZ17, CYWL17, DCA^{+}16, EDvW17, FS17a, FLW19, GSC19, HTFL18, IGQ15, KHTZ19, LPR18, LXC19, lLLNS16, lLNS17, LDT19, LD15, LCK19, LLH19, MDDM17, PST19, PxRS17, QSBY19, RH19, RLH19, SG18, SX15, SWZ17, SLZ^{+}17, TSH17, TBO^{+}16, VNA15, VSM16b, WRL16b, WLE17, WHE17, ZMF15, ZTBW19, ZLL16a, ZBZT17, YK18]. two-dimensions [ZYD19b]. two-field [CS16a, XZT18]. two-fluid [AAI16, Ama15, BAGK16, FG19, LDGH16, Niu16, RTG15, SJH^{+}15, ZKG19]. two-grid [ACJ17, YFC19]. two-group [JPLL15]. two-layer [CS18b, PM16]. Two-level [Vab18]. two-miscible-layer [SHLG15]. two-moment [CEHM19]. two-node [JPLL15, SGP17b]. two-particle [PSV18]. two-phase [ACGR15, AASRT17, Ani16, BGN15, BT19, BM19a, BAVC17, BHMS18, BKKRB16, CFSN18, CDM18, CGK17, Chi19, Cif19, CS17b, CG16, DG18, DGMT17, FGL16, FZ19, FS19a, HHA15, HTMP17, HCLT19, HLA19, HTBG15, JS16, JS17, JJ18b, KJ17a, KS16c, KS18b, LVTR15, LW18, LYM19, LPGT16, LSD^{+}17, LDGH16, MNG15a, MA19, MT19, MDP18, PL18, PSB^{+}18, PM19, PS14, PS15a, PGM17, RWG18, Ren19, RV16, RZ15, SHA16, SRBB18, SA19, Suz18, TH18, TGS19, TND18, TASA19, TT16, WDT^{+}19, WKSS15, XSL18, YSY17, YD19, ZZ17b, ZWL^{+}19, dJRP^{+}15, tEDKT17]. two-point [CSY19]. Two-scale [SAH17, CPT16, LMC16]. two-sided [KSM19, SZM19a, SZM19b, SYM17]. two-species [CCZ15]. two-stage [BJO18, DSX19, DL18c, LZT17, PXLL16]. two-step [BH18, HC17]. Two-way [TUJ19, EH18, HM16b, ID17, Mue18, PAL^{+}16, QYF15]. type [AAG16, AJP15, ADOP18, BG19c, BDZ15, BTVC16, CC17b, sCYxL^{+}18, DG16a, DL18c, GT18, HHY15, JCWX19, LDGH16, LHQ16, MRY19, MDP^{+}15, RMK15, Rod17, SYY16, Spe15, TMS^{+}19, WBBC16, XLL^{+}17, YZW^{+}18, YLD19, ZHS18, ZX19, ZS18, ZS19a, SW17a, SKO17].

Uehling [PSV18, Wu19, Yan17]. Uhlenbeck [PSV18, Wu19, Yan17]. ULPH [TL17, YLZ^{+}19]. ultra [CKT17, CL19b, DLN15, ION^{+}17]. ultra-high [CL19b]. ultra-relativistic [CKT17, ION^{+}17]. ultrasound [HTBG15]. un-split [MMB18]. unaveraged [ALM15]. unbounded [BNM15, BLS16, CLC16, FH17, GWC18, KADE15, KADE17, LZ16, LC16, MSA19, NGY^{+}17, SHW18]. unbounded-periodic [SHW18]. Uncertain [LSS16, BC18a, CNQ^{+}19, FDKI17, Poë19, SS18a]. uncertainties [AZK16, XWW^{+}16]. Uncertainty [CPX19, CZB15, GS18, PTT19, SS17b, AKZ16, AÁPB17, BHS^{+}18, BHJ15, CC17a, CE18, CQ15, CVG19, CELI15, DP19, DH18b, EH14, EH15, FC16, GMS19, GZ19, GSS^{+}19, HAPK15, HJ16, HJS19, IPSG15, JS17, KRBW17, KSV^{+}15, KBK15b, LS15c, LLL16, LSD18, MS16b, MSS16, PE15, RZOZ19, RMK15, RS17, TE19, TT17a, TBG16, TB18, TO19, WRL19, WL16, WTX17, WSK19, XS15, YP19, YQNW19, ZLGK19, ZZ18, ZZKP19, vdBKD17, MBNJ16]. uncertainty-based [FC16]. Uncollided [HHR^{+}19]. Unconditional [SSZ19]. Unconditionally [GGT18, SLL17, Tav16, WSF17, BC16c, CGJ19, GX15, HW15a, HJY19, LYM19, WCCB16, Yan16b, YH17, YD19]. unconformities [ST16]. under-resolved [FWK18, Kim15, KCS^{+}17, MSP15, MMPS17, WMM^{+}18]. undergoing [AQ19, GLS15]. underlying [ATF16]. underresolved [FBM16]. Uneven [Fal15]. Uneven-order [Fal15]. unfitted [ZSX17]. uniaxial [MDT16]. unidimensional [Heu17]. unification [Sid18]. Unified [BDAA^{+}18, PCBG18, WYLX17, AG16, BT15, CX15, DPRZ16, DPRZ17, FKK19, GSL18, HF19, KG15, KBD19, LS16b, LLFX18, LXSC16, LWX19, PLWJ16, SZM19a, SZM19b, SWJG19, SJX15, SJXL15, SJX17, TZGW18, XCX17, Yan19, ZZX16, ZZX19]. Uniform [An17, AB16a, AHZ19, CLMZ17, FL18, LB15, LYC16, PL16b, SS16b, SYM17, WR15, XDSX17]. Uniformly [CLMZ17, LN15, BZ16a, LAA16, XQ17]. uniquely [HW15a, WDGW17]. uniqueness [AB19]. unit [DJV^{+}18, WC18]. units [GK19b, GP18]. unity [FLT18, MIM^{+}19, NJHL18]. Universal [TKB^{+}15, YS18b, BLM18]. unknown [RZ18, WL18]. Unnormalized [GLOP19]. unresolved [BLG^{+}16]. Unsplit [BP18, CSH15, FGLB16, OD17]. unstable [CGTH18, SWML17]. unstaggered [MM16b]. unsteadiness [LSP19]. unsteady [ALO18, ACS16, ALL18, BK17a, BFT17, CPS17, CLP16b, EFO19, EFO20, Fid17, KPKGH19, Kay15, KA18, Ler15, Ler16, LLWJ18, LHMB18, LWZ19, MN18a, MS16b, MDP^{+}15, MCGS16, MC17, MPMB19, NdlLPCC19, NMJFM19, NDCB17, NSL16, PTT18, RDG17, RPC^{+}18, SPP^{+}16a, SDM^{+}17, Tou18, VPM15, VWV17, WHR19, YL16, Yi18, ZZX19]. Unstructured [ACS16, Har18, SCS16, AAE17, AEL^{+}17, AG18, BD15a, BD15b, BT16, BDZ15, BD17, BDLM18, BB19, BHTT17, Bou19, CFSN18, CBC^{+}18, CGK17, CHY16, CL19a, CSN17, CLTX15, CCM17, CLP16b, DvW15b, DL16, DMTB15, Eng18, EDvW17, FP16, GK18, HM19a, HM19b, Hu17, IZ18, Ism15, IM15, IM17b, JBLO15, KEJ18, KC17c, Kor17, KSI17, KS17, LLD^{+}16, LMG15, LLP^{+}16, LAL18, LL16b, LYKW19, LJ16, MHGM^{+}15, MF16a, ML16, MM17, NOM^{+}17, NYNYM15, Nis15, Nis18a, PX16, PxRS17, PL16a, PM16, PN17, PBC^{+}17, RH19, RSD17, SP19b, SSX16, Stü17, SJX17, SZS15, TD16a, TD17, TD18, TLB^{+}18, Tso18, VST16, VN15, VLV19, WLM15, WRL16a, WRL16b, WRPL17, WWGK17, WRL18, XX16, XDSX17, XX17, XL16, XWZ^{+}18, YDLC19, YSYW19, ZCHS15, ZOG19, ZZZ17, ZRW19, ZKG19, ZPW18, ZYD^{+}19a, ZXDL17]. unstructured [dLDG^{+}18]. unstructured-mesh [KS17, SZS15]. Unsupervised [VMK^{+}19]. updated [TL17, YLZ^{+}19]. updating [PDdG^{+}17]. upper [WSH19]. upscaling [ASWvD19, CEL^{+}18b, VCEK19]. upstream [MBD19]. Upwind [FRRV16, AGBL15, ABH18, BZ19a, BK19b, CKK18b, Fan16, FS18, HC18a, HLA19, LMKS15, LAEK18, Mat17, MO18b, YFJ18, ZH19]. upwind/central [HLA19]. upwinding [Sub18]. UQ [TB18, WRL19]. Use [MTL^{+}17, VBG16, BT17a, DA17, DCCC16, FG17, GMS19, HS17b, LSWF16, MHGL19, Yan19]. Using [CG15, KV16, SNB^{+}15, ATM^{+}18, AGC19, AKK^{+}19, ADGN17, AMJ17, AZK16, APLK19, AN15, ATF16, ABT16, BVM^{+}17a, BCSK17, BJRF18, BCST17, BD15a, BK17b, BK19a, BST^{+}18, BJ15, BDKK17, BNK18, BAVC17, BLS16, BKR19, Bou19, BRW15, CR17, Cap18, CBS18, CJK^{+}19, CP17, CC17b, CE18, CZB15, CZ17, CWWZ17, CCK^{+}18, CRMP16, CDM19, CLL17, CEL18a, CSK^{+}16, CLM15, CV16b, CLP16b, CCGH17, CCGH19, DD17a, DD15, DG16a, DSH^{+}16, DJV^{+}18, DPO16, DC18a, DMS17, Dod17, Dom18, DW19, EST17, EEG^{+}15, ECC18, Eva18, EDvW17, Fan19, Fid17, FGLB16, FBM16, FP16, FSB16, FRRV16, FS19a, FN17, FYC^{+}18, FKY15, Gam15, GSS^{+}19, GH19, GBvZB16, GWC17, GS18, GGL^{+}17, Gno17, GFvR18, GRS15, GBS15, GSN17, HED^{+}16, HZE19, HB16, HR19, HLL^{+}16, HX16, HW16a, HU18]. using [HLQ16, HLL^{+}18, Hue15, JW15a, JES15, JL18a, JL18b, JWH16, KAR17, KW15b, KK17a, uKHGK19, KZR15, KP15b, KDPK15, KSI17, KB19, LMH16, LTB16a, LDO^{+}19, LDOK17, LPG18, LYB18, LWLC17, LT17a, LHS^{+}19, LMBZ15, LHY^{+}19, LYKW19, LSR16, LC17b, LZS^{+}19, LMB19, LDSM19, LT17b, LVL18, LMGG17, LSI16, LBB^{+}17, MLL19, MBSS15, MM16b, MNG15a, MG15b, MH19, MPP15, MTJ18, MCGS16, MFB18, MSP15, MSB^{+}16, MSA19, MM18, MC17, NdlLPCC19, NMM17, NMJFM19, NCP^{+}17, NLK^{+}15, NSL16, Nor15, OB19, OLHD17, OKE17, PKN17, PPCK17, PD16a, PKLS17, PR16a, PGGW18, PS15b, PF15, PD16b, QLS^{+}19, QWX19, QLF16, QN19, RPK17a, RPK17b, RH19, RC18, RS17, RBGV15, RVK^{+}18, RPL^{+}18, RPC^{+}18, SGP17a, SYI^{+}19, SSVL18, SS17b, SAK18, SRBB18, SW18a, SPB18, SFT16, SWMD17a, SWMD17b, SCQP16, STW16, SLL17, SLB^{+}19, SDM^{+}17, SW18b]. using [SL19c, SWHV16, SC16, SGT16, SHP^{+}16, SS17c, SD18, TK12, TK15b, TND18, TVB^{+}16, Tou18, TO15, TBLM15, VLAB18, VSM17, VBL^{+}16, VLN^{+}18, WHR19, WHRL19, WWRS17, WBBC16, WS15a, WF17, WX19, XL17a, XDSX17, XP15, YYY^{+}16, YSC^{+}17, YCPD15, YXF^{+}16, Yan17, YL16, YC16, ZS16, ZB15, ZD17, ZJ18, ZZPH18b, ZZPH18a, ZNX15, aKT16, dPSS16, dlAC17, vdBKD17]. utility [VWV17]. Uzawa [WSF17].

v [CBA17, TCS16a]. vacuum [CSY15, NOM^{+}17, SR18]. valid [RKO17a]. Validation [ION^{+}17, SMA^{+}16, CDDL19, DDJ19, FOF15, GPS17a, GG15, MML17, MPP15, PST19, PT17b, RBY19, SHP^{+}16, SS17c]. validity [JG15]. value [Azi19, BDB18, DGHP17, Die15, DZC16, KADE15, PHHR17, PGH15, WZ15, WL18, XM18, ZG18b]. valued [LM15d, NNV19, Tav15, WF17]. valve [CVM^{+}19]. vanishing [MK17, MSP16]. vapor [BG16b, DD15, FMRZ17]. Vaporization [WY19, WCWY19]. vaporizing [PD19]. VAR [FDS^{+}15]. Variable [CWL^{+}16, SHP^{+}16, WCWY19, ABT16, Ata15, AZ19b, BDAA^{+}18, BFNGDNR18, BTT18, Cui15, EJMI18, EMSS17, GT18, HHRA19, HAX19, HW18, JL18b, LYD19, LLH19, MS16a, Niu16, PPCK17, Pan20, RBI18, Ran18, RLGT19, SP15a, SAK18, SXY18, SK18, TSH17, TPB16, WZ15, WW17, WKPS18, WSF17, YD19, YY17, ZK15, ZzSK15, LMG19]. variable-coefficient [WZ15, WW17]. Variable-density [WCWY19, EJMI18, SP15a]. Variable-order [CWL^{+}16, HHRA19, LLH19, Pan20, TSH17]. variable-property [WCWY19]. variable-separation [JL18b]. variables [GMLD18, Kla15, LK16b, MTJ18, SP15a, WBC^{+}16, ZKG19]. variably [HSK^{+}15, Zad11]. Variance [BBKS18, GAJ15, MWD16, NW15, CVG19, CCL16, KM17, VCNGP15]. Variance-reduced [MWD16]. variant [GBU15, HAX19, ZD15a]. variate [DP19, FDKI17, SWHK15]. Variation [SIX16, BKL17, DSX19, DLMDV18, ZC15]. Variational [Kou16, KTG16, PK17, WRPL17, ZC15, ZSX17, ADP^{+}17, CZBC^{+}18, CZB15, CL19a, CDM19, DLWY19, EBQ15, EE16, FPDT17, FG18, FKDL17, FPV18, GAN^{+}16, GS15c, GM16, GWE^{+}15, HKKP16, HK15b, JJ17, JJ18a, JJ18b, KR17, LWLC17, LWL17, LLJJ18, MCN18, MPR^{+}18, MH19, MWZ19, NBZ^{+}19, RG15, RS16a, RWG18, SWML17, SD17, SSO^{+}15, SWHV16, SSN15, YGEM17, YSY17, ZS16, AMPG19]. variations [GS18, WT16]. varied [WRL19]. various [BMT16]. varying [GDFL17, NHM17, NSK^{+}16, RÖS17, SKF16]. vascular [BFI^{+}16]. Vector [KBR17, BMT16, BGGM15, CLW18, CJL16, CX16, GKE15, HKA19, HN18, LZL^{+}19, Moc17, NNV19, OM19, SAF^{+}19, SE15, Tav15, WF17, YTW15, YT19]. vector- [NNV19]. vector-potential [CX16]. vector-valued [Tav15, WF17]. Velocity [SMS16, BLL19, BLL20, BDG^{+}17, BS15a, CSG17, CVG18, Fal16, HLML17, JLQX15, JLKF17, LM16, MBST17, NF17, OMLdL16, RRM^{+}16, SVG18, SW18b, YSYW19, ZHA17b]. Velocity-correction [SMS16]. velocity-decomposition [MBST17]. velocity-vorticity [BS15a]. Venant [LAEK18]. ventricular [CVM^{+}19]. VERA [TCS^{+}16b]. verifiability [GS15c]. Verification [WS15b, DDJ17, EKV^{+}16, RWKW16, VBG16, VGZ18]. verified [RMBN18]. Versatile [SUR18, AMS17, TBC^{+}16]. version [HZ15]. version/ [HZ15]. versus [MM16a, MZTS16, PR17a]. vertex [AGBL15, GZLH19, MMB18, MF16a, ZSW17]. vertex-centered [ZSW17]. vertex-discontinuous-Galerkin [MF16a]. vertex/edge [GZLH19]. vertex/edge-based [GZLH19]. Vertical [YSC^{+}17, YP17, CK16a, Mue18]. very [BZ19a, GS18, HXB15, Lap16, NMM15, PP18b]. Vesicle [HLSY16, BLJ17, CJYZ15, GGT18, KQB18, PZNG15, QB16, SHKL16, Vee16, Vog17]. vesicles [RVZB15, STKL19, TBLJ15]. vessel [ABT16]. vessels [Gam15]. VI [LWX19]. via [AA19, ALMJ15, BJTZ15, VMN^{+}18, BMPS18, BGRC19, BLK15, BDMZ19, CPX19, CZ19a, CW16, CHZ16, CMW16, DLY17, EFHZ17, FS18, FPDT17, GLG^{+}19, HKLZ18, HMRG16, Ike18, KM16b, KR17, KMP^{+}19, KW16, LM15b, LM15c, MPR^{+}18, NHA18, PK16, PHD16, PR16c, PBCR19, RÖS16, RTV17, RBD17, RRD19, SW15, TG17, VST16, ZZ17b, ZX19]. vibrating [RLH19, ZMF15]. vibration [BPGS16, CLB^{+}16, ZLH^{+}17, ZC18, ZBZ^{+}18]. vibrational [BHJ15, CVG18, WYLX17]. vibroacoustic [BC18a, TP17]. view [AJP15, Par15, Stü15]. Vinokur [GMD19]. violent [SA19]. virtual [BBB^{+}16, CWW17, LJ19, PJC16, TTN^{+}16, ZZYC19, TCS^{+}16b]. viscoelastic [CC15, EFO19, EFO20, GSS15b, HM17, KSI17, LHMB16, MOAA15, MS18d, MLB16, NBMB19, RJ19b, STKL19]. viscoelasticity [YPK16]. viscoplastic [FNGV18, LEB^{+}17]. viscoresistive [HdBH^{+}16]. viscosities [BR15a, YY17]. viscosity [CM18a, CWS18, CJD^{+}17, DRM15, FB17, FRRV16, HIN^{+}16, LWB^{+}16, MK17, MCN18, MG15b, MSP16, RVZB15, RRS19b, RRS19a, Rod17, Rod18, SHP^{+}16, TLB^{+}18, WDT^{+}19]. Viscous [DPRZ16, LAA16, AMM^{+}15, BST^{+}18, BAD19, BLG^{+}16, BKKRB16, CBS18, CJD^{+}17, CX16, FKK19, HEPG15, HGW18, HLS15, HDF18, KDL15, LVTR15, LT15, LC16, LC17b, MS17, Mon19, MCGS16, MM16d, NNW17, OB19, PL18, PPLC16, QSB18, QM18, RBJS15, RXSG15, RAMB15, SGMS16, SKF15, SWLW19, SST^{+}15, SSZC19, TGY18, Tou18, WTL17, Wil18, XDLX19, YSWS16, YXF^{+}16, YZZ15, ZZPH18b, ZZPH18a, ZYSW16, ZW19, ZLGS18, aKT16]. viscous-plastic [WTL17]. visualization [HIN^{+}16, KLA17, MBM^{+}18]. vivo [LGZ^{+}19]. Vlasov [QHZ^{+}15, BDM17, CQQ16, CGQ18, CCZ18, CCZ15, CGJ16, Cot18, CEF15, CLMZ17, DDD17, Del15, DCD^{+}18, EL17, Ein19, FFM19, JG19, LY15a, LHS^{+}19, MDVM16, RTG15, SG19a, SOS19, SH19, SC16, TC15a, TKC15, VK18, VSC18, WSJY16, ZG17]. VOF [CDM18, CDM19, HDA^{+}18, LHGF19, MNG15a, PR16a, MMB18]. VOF-based [PR16a]. voids [BKS18]. Volterra [Moh15]. Volume [ADN19, AGBL15, DG18, FB15, FPT17, HSLQ15, JBLO15, Kat16, MDL16, MHGL19, NT15, RW15b, SYOS19, SYOS21, SAK18, SGD18, TNB21, VW18, WW19, ABG^{+}15, APP^{+}16, AEL^{+}15a, AEL^{+}15b, AA19, AHZ19, ABT16, AMM^{+}15, AM17b, AKM^{+}19, BD15a, Bat17, BGV17, BLVC17, BLMY17, BTVB15, BLD15, BDZ15, BD17, BDLM18, BHTT17, BKRB15, BFTVC18, CSW^{+}19, CCS18, CCZC16, COV18, Cho15, CDX18b, CGP16, CSH15, CHS17, CCM17, DRP^{+}16, DK19, DB16a, DMS17, DDH^{+}18, DVP^{+}16, DMM19, DL16, EKSS15, Eng18, EDvW17, FS18, FLW19, GOR17, GHL15, GLK19, GSL^{+}19, HWK19, HSLQ16, Heu17, HMFJ18, HY16, Hu17, Ism15, IGQ15, IDSG15, IM17b, JME18, JMM19, JW15c, JW16, KKH18, KW15b, Kla15, KB19, KS17, LLD^{+}16, LN17, LAL18, LX16, LL16b, LLSJ19, LZ17b, LYKW19, LDT19, LY16c, LJ16, LHGF16]. volume [LHGF19, MAK15, MDHC15, MH18a, MMvR18, MRK15, MH18b, MSS16, MLB18, NJPB17, NBMB19, Nis15, Nis19a, Nor15, PD19, PxRS17, PL18, PHÖ^{+}16, PS16, PR16b, Pei16, PWP15, QLS^{+}19, RMA17, RKRGW17, RBL16, SPX^{+}18, SR19, SAEF17, SRBB18, SY16, SKO17, SLY16, SYM17, SDH^{+}16, SKG17, SBH19, SFP16, SDW18, Sub18, Tav15, TMT17, TV19, TND18, TVB^{+}16, Tso18, TR19, Vil19, VSC18, WR15, WRL16a, WRL16b, WRPL17, XWL^{+}16, XDvW17, XX16, XL17a, XDSX17, XDLX19, XM18, ZCHS15, ZOG19, ZZZ17, ZG18a, ZSL^{+}19, ZRW19, ZQ17, ZXDL17, vEKdB16, AAL15, BAVC17, CJ17, YK19]. volume-averaged [BTVB15]. volume-conserving [LYKW19]. Volume-of-Fluid [JBLO15, RW15b, SAK18, CSW^{+}19, IM17b, LY16c, PR16b, RKRGW17, SRBB18, TND18]. Volume-of-Fluid-based [FB15]. Volume-preserving [HSLQ15, WW19, HSLQ16]. volume/finite [BFTVC18]. volume/Monte [GDS^{+}16]. volumetric [Swe18, TWN19, WN17]. volumetric-ray-casting [Swe18]. Voronoi [FHA17a, GLTG15, GPG17, MGPG19, PLB18, YGJ18]. Voronoi-based [MGPG19]. Vortex [PWC18a, BPGS16, BGRC19, BR15b, BR16, HKLW15, RHvR^{+}15, XWB15, XY17, ZYW16, GMWC19]. vortex-induced [BPGS16]. vortex-surface [XY17]. vortices [LLM17]. Vorticity [CX16, BS15a, CWS18, CMDL18, KO17, PG17, XWB15, Sid18]. voxelization [PA15]. Vries [LY16b]. vs [CFG16, DLLV17].

Waals [PSS17]. wake [PEVG18]. Wakefield [MAM16, YXD^{+}16]. wakefields [RMLvR18]. Walk [HHK15, ADHN15, BSP18, KC17b, MS15a, RFGSV15]. Walk-on-Spheres [HHK15]. Walks [OADN19, NHA18]. Wall [Don17, CW16, CW17, CV15, DRZ^{+}19, HL15a, HHY15, LDB19, MHT^{+}19, MS17, NL15, PM16, PCN15b, PBCR19, Stü15, SGP17b, VM15, YS18b].

Wall-bounded [Don17, LDB19]. walled [FLV15]. walls [DCBK15, FNGDMNR18]. Walsh [Gno17]. Wang [FJLC18]. Warburton [AMP16]. warm [SP16c]. Wasserstein [CCWY18, YLH^{+}19]. Water [NMM17, NMM18, SP16a, TK16, TM17, AB19, ABT16, BC18b, BNGI19, BFNGDNR18, BHGK18, CV17, Cap18, CS18b, CSLL15, CLB^{+}16, CE17, CSCM16, CK16a, CDV17, DK19, DA17, DMTB15, EL18, EL19, EDK19, EKEB16, EDC19, FS17a, GP16a, GIF18, GCVMK15, HSM19, HLJ^{+}19, JJS15, KL18a, LMPS15, LPG18, LP18, LCLY19, LDWZ15, LMKS15, LY16c, LMSK17, LCK19, MDBCF17, Mue18, NMM15, PP19, DM18, RW15b, Ric15, SGC18b, SMSR18, SD16, TC15b, TSB^{+}18, VST16, WHRL19, WWGK17, WWGW18, WG15, WBM^{+}15b, YM17b, ZA15a, ZED15]. Wave [Luc15, MT17, PS15b, AMN18, ABP^{+}16, AMJ17, An17, ABH18, ADOP18, BJO18, BG19a, BHJ18, BNM15, BS19a, BG19b, BDBEE15, BH18, BGGM15, BTT18, CZW17, CCFC19, CGMH18, CDL19, CDDL19, CNQ^{+}19, CSG17, CLX19, CLQ17, CJH^{+}19, DCA^{+}16, DWG^{+}18, DL18a, DYL19, DKK15, FS16, FKR16, GFG^{+}15, GH17a, GFC18, GK19a, GZY19, GKNA17, GP16b, HK15a, HSC16, HXB15, JCWX19, KS18a, KÁGR18, KLRT15, LC18, LHMB16, LC17a, LWWY18, LGB16, LTB16b, LY16c, LK16a, LZL^{+}19, LMM17, LQB16, MD17, MSS16, MFB18, Mue18, MNW19, MSH^{+}15, MH17, MSF^{+}19, NPB19, PDdG^{+}17, POSB16, RBY19, RM16, RSH^{+}17, SSL17, ST16, SCN^{+}17, SS17a, SF18a, SM16, SZF15, TDC^{+}19, Ter18, TM15a, VSDW18, VMM19, WMY18, WLW^{+}18, WLE17, WSOW16, WAZ19, XYPT16, XOX19, XS19, YYL16, Yas17, YLA15, ZZZ17, ZHS18, ZZ19, ZZW^{+}16, ZCZ19, ZWUR16]. Wave [DJV^{+}18]. wave-based [AMJ17, LGB16, MSF^{+}19]. wave-current [WMY18]. Wave-diffusion [Luc15]. wave-equation [LWWY18]. wave-in-cell [TM15a]. wave-mode [LZL^{+}19]. wave-structure [NPB19]. wavefield [LTXB17]. Waveform [NGS16, BFP18, MKYZ17, PKN17, YLH^{+}19]. waveguide [BCM19, Liu19a]. waveguides [GTL18, RMLvR18, Tre16]. Wavelet [BDV17, DLK17, CWL^{+}16, CYWL17, FBY19, GBD^{+}15, HHRA19, LAK^{+}16, Moh15, NVBDV15, Pan20, RRS19b, RRS19a]. Wavelet-based [DLK17, CYWL17, GBD^{+}15, Moh15]. Wavelets [MVKD15, ABP^{+}16, BOD19, SWHV16]. wavenumber [LN15]. wavepacket [BHFB19]. waves [CGSS18, CDX^{+}18a, DBD^{+}17, DK19, DZR18, DK18a, DK18b, DvW15a, DZ16, DDH^{+}18, DHC16, EKEB16, EDC19, FS17a, GD19, GBS15, GP16c, HN17a, Hu17, HTBG15, IG15, IML15, LC18, LWZ16, LMC19, Mag19, MDW18, MC17, OLHD17, PKN17, PS14, PS15a, SZW^{+}16, SSM^{+}17, SMSR18, SS17c, SWZ17, TLB^{+}18, TEP19, VAD17, VK15, WMY16, WTL17, XJG18, XYPT16, YSC^{+}17, ZD17, ZED15, vOMB17]. way [EH18, HM16b, ID17, Mue18, PAL^{+}16, QYF15, SL16b, TUJ19, Ter18, TP17, TC15c]. Weak [DDJ19, KML18, Svä15, Fal16, FG18, KLC18, LYZ18, LTW18, MWYZ16, SD17, VMC^{+}19, WYZZ18]. weak-constraint [SD17]. weak-coupling [KLC18]. weak-perturbative [Fal16]. weakly [ALA16, CGM15, LM15a, MA19, PMS15, DM18, SHA16, SPP16b, Tsa15, Tsa16, VM15, ZHA17a, ZXW^{+}19]. weakly-ionized [PMS15]. Webb [ZJS15]. wedge [CDL19, CDDL19]. weight [CW19]. weight-adjusted [CW19]. Weighted [LJ16, DM18, ZPW18, ZNX15, BDAA^{+}18, FBL17, GSN17, GLZ19, HLS19, HWA15, KK17b, LPWK15, LZZS15, MSP19, OKWE17, SWL19, WL17, YL16, ZQ17]. Weighted-least-squares [LJ16]. Well [CV17, CCK^{+}18, IG15, JWH16, LX18, LMKS15, LAEK18, MBM^{+}18, NL18b, AASPT18, ABT16, DVP^{+}16, FNGDMNR18, GMD19, GLK19, MDBCF17, NMM18, PN17, PND16, SO17, XCX17]. Well-balanced [CCK^{+}18, LX18, LMKS15, LAEK18, AASPT18, ABT16, FNGDMNR18, GLK19, MDBCF17, NMM18, PN17, XCX17]. Well-conditioned [JWH16, SO17]. well-driven [DVP^{+}16]. Well-posed [NL18b, PND16]. Well-posedness [IG15, GMD19]. Well-suited [MBM^{+}18]. Wendroff [DDJ18, DL18c, FLW16, Heu17, Heu19, JCWX19, LFT^{+}16]. WENO [SZN20, Sid18, AdRBC16, AHZ19, BGS16, BK16a, BQRX19, BB19, Bre17, CLTX15, CGJ16, CGJY19, DLK17, DS15d, DL18b, GGL^{+}17, HAH16, HC18a, HC18b, HLA19, Jac17b, JZ16, uKHGK19, KC18, LX16, NMJFM19, NF17, Nor15, PS16, RLGT19, SZN19, SG19b, Shu16, SH19, TLQ15, TLQ16, WDS15, WLGD18, WT15, ZXW^{+}19, ZQ16a, ZSQ17, ZQ16b, ZS17, ZS18, ZH19, ZS19a, dFJN16, vLtTBI17]. WENO-based [CGJ16]. WENO-solver [DS15d]. WENO-Z [AdRBC16, WLGD18]. Westervelt [SK15a]. wet [LAEK18, PP19, WWGW18]. wet/dry [LAEK18, WWGW18]. wetting [ABT16, HSB16, LGD17, PKB15, Pes15, WWX19, XWW17]. wetting-drying [ABT16]. Wheeler [JdR^{+}18]. white [CHLZ17, XZJK19]. Whitney [KSVB18]. Whole [BMRA^{+}15, ANL^{+}16, MJ16, NCP^{+}17]. wide [SY18b]. wideband [EGO19]. widest [DBD^{+}17]. Wiener [TG17]. Wigner [CSC19, DS15d, FSM16, SD15, SS15b, VSM17]. Willmore [CLS^{+}18]. wind [CGSS18, GPAO^{+}18, MBST17]. Windowed [JDFS16, SL16a]. windows [DH18a]. wing [Moo17]. Winslow [FP16]. wise [LTKA15, MN16c]. within [AAL15, BGV17, JS16, LSS16, LLVF^{+}15, PE16a, RW15a]. without [CGQ18, GMS16, KS18a, KL19, Xia15, ZZKP19]. WLP [WSOW16]. WLP-FDTD [WSOW16]. WLS [CH19, LJ16]. WMLES [DWR18]. Wood [BM19b]. workflow [LBB^{+}17]. wormhole [XYG19]. WOS [HHK15]. wrinkling [HJY19].

X [NLK^{+}15, WSU^{+}15]. X-ray [NLK^{+}15]. XAVM [RWG18]. Xeon [SGL18]. xylose [ASB^{+}15].

Yang [HK16a, AZ16, ZA15b]. Yee [DPO16, LL19a, NT16, dSPDH15]. yield [LK17, LEB^{+}17]. Yin [AZ16, HK16a, ZA15b]. Yuan [YY16].

Z [Pan20, AdRBC16, WLGD18, WRL18]. Z-pinch [WRL18]. Zadeh [HSK^{+}15]. Zakharov [BZ16a]. Zero [PPM^{+}19, HED^{+}16, KBR17]. Zero-flux [PPM^{+}19]. Zhong [HK16a]. Ziff [CY19b]. zirconium [MTL^{+}17]. zonal [BFI^{+}18, PM16, CJ17]. zonation [LVL18]. zone [GEZK16, NWB19]. Zwanzig [PD17, ZV18].

References

Anistratov:2015:ISA

[AA15] Dmitriy Y. Anistratov and Yousry Y. Azmy. Iterative stability analysis of spatial domain decomposition based on block Jacobi algorithm for the diamond-difference scheme. Journal of Computational Physics, 297(??):462–479, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003678.

Andrews:2019:DCS

[AA19] Stephen A. Andrews and Tariq Aslam. On the direct construction of the steady traveling solution to high explosive sandwich, cylinder and aquarium tests via a streamline finite volume approximation. Journal of Computational Physics, 395(??):653–670, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304371.

Aliaga:2016:FBK

[AAB^{+}16] José I. Aliaga, Pedro Alonso, José M. Badía, Pablo Chacón, Davor Davidović, José R. López-Blanco, and Enrique S. Quintana-Ortí. A fast band-Krylov eigensolver for macromolecular functional motion simulation on multicore architectures and graphics processors. Journal of Computational Physics, 309(??):314–323, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000085.

Ahusborde:2015:MSD

[AABD15] E. Ahusborde, M. Azaïez, F. Ben Belgacem, and E. Palomo Del Barrio. Mercer’s spectral decomposition for the characterization of thermal parameters. Journal of Computational Physics, 294(??):1–19, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001928.

Aminfar:2016:FBL

[AAD16] AmirHossein Aminfar, Sivaram Ambikasaran, and Eric Darve. A fast block low-rank dense solver with applications to finite-element matrices. Journal of Computational Physics, 304(??):170–188, January 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006750.

Adrian:2017:HPE

[AAE17] S. B. Adrian, F. P. Andriulli, and T. F. Eibert. A hierarchical preconditioner for the electric field integral equation on unstructured meshes based on primal and dual Haar bases. Journal of Computational Physics, 330(??):365–379, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630599X.

Adrian:2019:RFC

[AAE19] S. B. Adrian, F. P. Andriulli, and T. F. Eibert. On a refinement-free Calderón multiplicative preconditioner for the electric field integral equation. Journal of Computational Physics, 376(??):1232–1252, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306661.

Adimurthi:2016:GTN

[AAG16] A. Adimurthi, Aekta Aggarwal, and G. D. Veerappa Gowda. Godunov-type numerical methods for a model of granular flow. Journal of Computational Physics, 305(??):1083–1118, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006294.

Adeleke:2016:RLF

[AAI16] Najeem Adeleke, Michael Adewumi, and Thaddeus Ityokumbul. Revisiting low-fidelity two-fluid models for gas-solids transport. Journal of Computational Physics, 319(??):28–43, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301577.

Abadie:2015:CES

[AAL15] T. Abadie, J. Aubin, and D. Legendre. On the combined effects of surface tension force calculation and interface advection on spurious currents within Volume of Fluid and Level Set frameworks. Journal of Computational Physics, 297(??):611–636, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003113.

Arnst:2017:ISM

[AÁPB17] M. Arnst, B. Abello Álvarez, J.-P. Ponthot, and R. Boman. Itô-SDE MCMC method for Bayesian characterization of errors associated with data limitations in stochastic expansion methods for uncertainty quantification. Journal of Computational Physics, 349(??):59–79, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305764.

Abu-Al-Saud:2018:CWB

[AASPT18] Moataz O. Abu-Al-Saud, Stéphane Popinet, and Hamdi A. Tchelepi. A conservative and well-balanced surface tension model. Journal of Computational Physics, 371(??):896–913, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301049.

Abu-Al-Saud:2017:MLS

[AASRT17] Moataz O. Abu-Al-Saud, Amir Riaz, and Hamdi A. Tchelepi. Multiscale level-set method for accurate modeling of immiscible two-phase flow with deposited thin films on solid surfaces. Journal of Computational Physics, 333(??):297–320, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116307008.

Archer:2015:NNO

[AB15] Philip J. Archer and Wei Bai. A new non-overlapping concept to improve the Hybrid Particle Level Set method in multi-phase fluid flows. Journal of Computational Physics, 282(??):317–333, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114007761.

Akiki:2016:IBM

[AB16a] G. Akiki and S. Balachandar. Immersed boundary method with non-uniform distribution of Lagrangian markers for a non-uniform Eulerian mesh. Journal of Computational Physics, 307(??):34–59, February 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115007597.

Amlani:2016:FBS

[AB16b] Faisal Amlani and Oscar P. Bruno. An FC-based spectral solver for elastodynamic problems in general three-dimensional domains. Journal of Computational Physics, 307(??):333–354, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008086.

Asgharzadeh:2017:NKM

[AB17] Hafez Asgharzadeh and Iman Borazjani. A Newton–Krylov method with an approximate analytical Jacobian for implicit solution of Navier–Stokes equations on staggered oversetcurvilinear grids with immersed boundaries. Journal of Computational Physics, 331(??):227–256, February 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306258.

Alinovi:2018:BEM

[AB18] Edoardo Alinovi and Alessandro Bottaro. A boundary element method for Stokes flows with interfaces. Journal of Computational Physics, 356(??):261–281, March 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308793.

Aleksyuk:2019:UES

[AB19] Andrey I. Aleksyuk and Vitaly V. Belikov. The uniqueness of the exact solution of the Riemann problem for the shallow water equations with discontinuous bottom. Journal of Computational Physics, 390(??):232–248, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302347.

Argaud:2018:SPN

[ABdC^{+}18] J.-P. Argaud, B. Bouriquet, F. de Caso, H. Gong, Y. Maday, and O. Mula. Sensor placement in nuclear reactors based on the generalized empirical interpolation method. Journal of Computational Physics, 363(??):354–370, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301414.

Aceto:2019:RKM

[ABDN19] L. Aceto, D. Bertaccini, F. Durastante, and P. Novati. Rational Krylov methods for functions of matrices with applications to fractional partial differential equations. Journal of Computational Physics, 396(??):470–482, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304942.

Amore:2016:HOE

[ABFR16] Paolo Amore, John P. Boyd, Francisco M. Fernández, and Boris Rösler. High order eigenvalues for the Helmholtz equation in complicated non-tensor domains through Richardson extrapolation of second order finite differences. Journal of Computational Physics, 312(??):252–271, May 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000796.

Abushaikha:2015:ICV

[ABG^{+}15] Ahmad S. Abushaikha, Martin J. Blunt, Olivier R. Gosselin, Christopher C. Pain, and Matthew D. Jackson. Interface control volume finite element method for modelling multiphase fluid flow in highly heterogeneous and fractured reservoirs. Journal of Computational Physics, 298(??):41–61, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003587.

Abgrall:2016:E

[Abg16]  Rémi Abgrall.  Editorial.  Journal of Computational

Physics, 307(??):A1, February 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115008785.

Abgrall:2018:GFC

[Abg18a] R. Abgrall. A general framework to construct schemes satisfying additional conservation relations. Application to entropy conservative and entropy dissipative schemes. Journal of Computational Physics, 372(??):640–666, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304091.

Afkhami:2018:TNM

[ABG^{+}18b] S. Afkhami, J. Buongiorno, A. Guion, S. Popinet, Y. Saade, R. Scardovelli, and S. Zaleski. Transition in a numerical model of contact line dynamics and forced dewetting. Journal of Computational Physics, 374(??):1061–1093, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305357. See corrigendum

[ABG^{+}19b].

Arias:2018:PEI

[ABG18c] Victoria Arias, Daniil Bochkov, and Frederic Gibou. Poisson equations in irregular domains with Robin boundary conditions — solver with second-order accurate gradients. Journal of Computational Physics, 365(??):1–6, July 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301827.

Abgrall:2019:E

[Abg19a]  Rémi Abgrall.  Editorial.  Journal of Computational

Physics, 390(??):iii–iv, August 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302268.

Afkhami:2019:CTN

[ABG^{+}19b] S. Afkhami, J. Buongiorno, A. Guion, S. Popinet, Y. Saade, R. Scardovelli, and S. Zaleski. Corrigendum to “Transition in a numerical model of contact line dynamics and forced dewetting” [j. comput. phys. 374 (2018) 1061–1093]. Journal of Computational Physics, 382(??):61–64, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306806. See [ABG^{+}18b].

Angel:2018:HOU

[ABH18] Jordan B. Angel, Jeffrey W. Banks, and William D. Henshaw. High-order upwind schemes for the wave equation on overlapping grids: Maxwell’s equations in second-order form. Journal of Computational Physics, 352(??):534–567, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306976.

Angel:2019:HOA

[ABH^{+}19] Jordan B. Angel, Jeffrey W. Banks, William D. Henshaw, Michael J. Jenkinson, Alexander V. Kildishev, Gregor Kovačič, Ludmila J. Prokopeva, and Donald W. Schwendeman. A high-order accurate scheme for Maxwell’s equations with a generalized dispersive material model. Journal of Computational Physics, 378(??):411–444, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307472.

Abdulle:2017:TSO

[ABI17] Assyr Abdulle, Ondrej Budác, and Antoine Imboden. A threescale offline-online numerical method for fluid flow in porous media. Journal of Computational Physics, 337(??):175–202, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300992.

Avgerinos:2019:LIA

[ABIR19] Stavros Avgerinos, Florian Bernard, Angelo Iollo, and Giovanni Russo. Linearly implicit all Mach number shock capturing schemes for the Euler equations. Journal of Computational Physics, 393(??):278–312, September 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302530.

Ahlfeld:2016:SSA

[ABM16] R. Ahlfeld, B. Belkouchi, and F. Montomoli. SAMBA: Sparse Approximation of Moment-Based Arbitrary polynomial chaos. Journal of Computational Physics, 320(??):1–16, September 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301516.

Akhmetgaliyev:2015:BIA

[ABN15] Eldar Akhmetgaliyev, Oscar P. Bruno, and Nilima Nigam. A boundary integral algorithm for the Laplace Dirichlet– Neumann mixed eigenvalue problem. Journal of Computational Physics, 298(??):1–28, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003447.

Adam:2016:AHW

[ABP^{+}16] Alexandros Adam, Andrew G. Buchan, Matthew D. Piggott, Christopher C. Pain, Jon Hill, and Mark A. Goffin. Adaptive Haar wavelets for the angular discretisation of spectral wave models. Journal of Computational Physics, 305(??):521–538, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007238.

Antoine:2016:HOI

[ABR16] Xavier Antoine, Christophe Besse, and Vittorio Rispoli. Highorder IMEX-spectral schemes for computing the dynamics of systems of nonlinear Schrödinger/Gross–Pitaevskii equations. Journal of Computational Physics, 327(??):252–269, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630434X.

Ardakani:2016:SWS

[ABT16] Hamid Alemi Ardakani, Thomas J. Bridges, and Matthew R. Turner. Shallow-water sloshing in a moving vessel with variable cross-section and wetting-drying using an extension of George’s well-balanced finite volume solver. Journal of Computational Physics, 314(??):590–617, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001893.

Antunes:2017:DHE

[ABT17] Pedro R. S. Antunes, Cristian Barbarosie, and Anca-Maria Toader. Detection of holes in an elastic body based on eigenvalues and traces of eigenmodes. Journal of Computational Physics, 333(??):352–368, March 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116307082.

Archibald:2019:DFM

[ABT19] Richard Archibald, Feng Bao, and Xuemin Tu. A direct filter method for parameter estimation. Journal of Computational Physics, 398(??):Article 108871, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305625.

Antuono:2016:DRI

[AC16] M. Antuono and G. Colicchio. Delayed Over-Relaxation for iterative methods. Journal of Computational Physics, 321(??):892–907, September 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302418.

Askham:2017:AFM

[AC17] T. Askham and A. J. Cerfon. An adaptive fast multipole accelerated Poisson solver for complex geometries. Journal of Computational Physics, 344(??):1–22, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730342X.

Ammari:2015:DCE

[ACC^{+}15] Habib Ammari, Junqing Chen, Zhiming Chen, Darko Volkov, and Han Wang. Detection and classification from electromagnetic induction data. Journal of Computational Physics, 301(??):201–217, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005549.

Adsuara:2017:EBS

[ACCCD^{+}17] J. E. Adsuara, I. Cordero-Carrión, P. Cerdá-Durán, V. Mewes, and M. A. Aloy. On the equivalence between the Scheduled Relaxation Jacobi method and Richardson’s non-stationary method. Journal of Computational Physics, 332(??):446–460, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306738.

Adsuara:2016:SRJ

[ACCCDA16] J. E. Adsuara, I. Cordero-Carrión, P. Cerdá-Durán, and M. A. Aloy. Scheduled Relaxation Jacobi method: Improvements and applications. Journal of Computational Physics, 321(??):369–413, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911630198X.

Abgrall:2015:SDE

[ACGR15] R. Abgrall, P. M. Congedo, G. Geraci, and M. G. Rodio. Stochastic Discrete Equation Method (sDEM) for two-phase flows. Journal of Computational Physics, 299(??):281–306, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500399X.

Anistratov:2017:SAN

[ACJ17] Dmitriy Y. Anistratov, Luke R. Cornejo, and Jesse P. Jones. Stability analysis of nonlinear two-grid method for multigroup neutron diffusion problems. Journal of Computational Physics, 346(??):278–294, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730459X.

Angelidis:2016:UCR

[ACS16] Dionysios Angelidis, Saurabh Chawdhary, and Fotis Sotiropoulos. Unstructured Cartesian refinement with sharp interface immersed boundary method for 3D unsteady incompressible flows. Journal of Computational Physics, 325(??):272–300, November 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303850.

Abgrall:2015:LNL

[AD15] R. Abgrall and D. De Santis. Linear and non-linear high order accurate residual distribution schemes for the discretization of the steady compressible Navier–Stokes equations. Journal of Computational Physics, 283(??):329–359, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007931.

Aditya:2017:HOA

[AD17] Konduri Aditya and Diego A. Donzis. High-order asynchronytolerant finite difference schemes for partial differential equations. Journal of Computational Physics, 350(??):550–572, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306149.

Allaire:2017:SOU

[ADE^{+}17] G. Allaire, C. Dapogny, R. Estevez, A. Faure, and G. Michailidis. Structural optimization under overhang constraints imposed by additive manufacturing technologies. Journal of Computational Physics, 351(??):295–328, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307015.

Ackerman:2017:FEA

[ADFG17] David M. Ackerman, Kris Delaney, Glenn H. Fredrickson, and Baskar Ganapathysubramanian. A finite element approach to self-consistent field theory calculations of multiblock polymers. Journal of Computational Physics, 331(??):280–296, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306064.

Abreu:2019:CAG

[ADG19] Eduardo Abreu, Ciro Díaz, and Juan Galvis. A convergence analysis of Generalized Multiscale Finite Element Methods. Journal of Computational Physics, 396(??):303–324, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304802.

Aguerre:2017:CHA

[ADGN17] Horacio J. Aguerre, Santiago Márquez Damián, Juan M. Gimenez, and Norberto M. Nigro. Conservative handling of arbitrary non-conformal interfaces using an efficient supermesh. Journal of Computational Physics, 335(??):21–49, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300281.

Angstmann:2016:SPN

[ADH^{+}16] C. N. Angstmann, I. C. Donnelly, B. I. Henry, B. A. Jacobs, T. A. M. Langlands, and J. A. Nichols. From stochastic processes to numerical methods: a new scheme for solving reaction subdiffusion fractional partial differential equations. Journal of Computational Physics, 307(??):508–534, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007937.

Angstmann:2015:DTR

[ADHN15] C. N. Angstmann, I. C. Donnelly, B. I. Henry, and J. A. Nichols. A discrete time random walk model for anomalous diffusion. Journal of Computational Physics, 293(??):53–69, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400549X.

Anderson:2017:HOL

[ADK^{+}17] R. Anderson, V. Dobrev, Tz. Kolev, D. Kuzmin, M. Quezada de Luna, R. Rieben, and V. Tomov. High-order local maximum principle preserving (MPP) discontinuous Galerkin finite element method for the transport equation. Journal of Computational Physics, 334(??):102–124, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306854.

Aguerre:2019:CIS

[ADN19] Horacio J. Aguerre, Santiago Márquez Damián, and Norberto M. Nigro. Conservative interpolation on surface interfaces for transport problems in the Finite Volume Method. Journal of Computational Physics, 395(??):144–165, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304267.

Audusse:2018:AMG

[ADOP18] Emmanuel Audusse, Minh Hieu Do, Pascal Omnes, and Yohan Penel. Analysis of modified Godunov type schemes for the two-dimensional linear wave equation with Coriolis source term on cartesian meshes. Journal of Computational Physics, 373(??):91–129, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303073.

Arcucci:2017:VDA

[ADP^{+}17] Rossella Arcucci, Luisa D’Amore, Jenny Pistoia, Ralf Toumi, and Almerico Murli. On the variational data assimilation problem solving and sensitivity analysis. Journal of Computational Physics, 335(??):311–326, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300505.

Acker:2016:IWZ

[AdRBC16] F. Acker, R. B. de R. Borges, and B. Costa. An improved WENO-Z scheme. Journal of Computational Physics, 313(??):726–753, May 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116000826.

Amritkar:2015:RKS

[AdSŚ^{+}15] Amit Amritkar, Eric de Sturler, Katarzyna Świrydowicz, Danesh Tafti, and Kapil Ahuja. Recycling Krylov subspaces for CFD applications and a new hybrid recycling solver. Journal of Computational Physics, 303(??):222–237, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006336.

Arqub:2015:CPS

[AEAM15] Omar Abu Arqub, Ahmad El-Ajou, and Shaher Momani. Constructing and predicting solitary pattern solutions for nonlinear time-fractional dispersive partial differential equations. Journal of Computational Physics, 293(??):385–399, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006706.

Ahmed:2015:CVD

[AEL^{+}15a] R. Ahmed, M. G. Edwards, S. Lamine, B. A. H. Huisman, and M. Pal. Control-volume distributed multi-point flux approximation coupled with a lower-dimensional fracture model. Journal of Computational Physics, 284(??):462–489, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008705.

Ahmed:2015:TDC

[AEL^{+}15b] Raheel Ahmed, Michael G. Edwards, Sadok Lamine, Bastiaan A. H. Huisman, and Mayur Pal. Three-dimensional controlvolume distributed multi-point flux approximation coupled with a lower-dimensional surface fracture model. Journal of Computational Physics, 303(??):470–497, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006609.

Ahmed:2017:CMF

[AEL^{+}17] Raheel Ahmed, Michael G. Edwards, Sadok Lamine, Bastiaan A. H. Huisman, and Mayur Pal. CVD–MPFA full pressure support, coupled unstructured discrete fracturematrix Darcy-flux approximations. Journal of Computational Physics, 349(??):265–299, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305491.

Akkutlu:2018:MMR

[AEVW18] I. Yucel Akkutlu, Yalchin Efendiev, Maria Vasilyeva, and Yuhe Wang. Multiscale model reduction for shale gas transport in poroelastic fractured media. Journal of Computational Physics, 353(??):356–376, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307763.

Azis:2019:IBM

[AEvW19] Mohd Hazmil Abdol Azis, Fabien Evrard, and Berend van Wachem. An immersed boundary method for incompressible flows in complex domains. Journal of Computational Physics, 378(??):770–795, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307150.

Alizadehrad:2018:SDP

[AF18] Davod Alizadehrad and Dmitry A. Fedosov. Static and dynamic properties of smoothed dissipative particle dynamics. Journal of Computational Physics, 356(??):303–318, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308859.

Abdi:2016:ECU

[AG16] Daniel S. Abdi and Francis X. Giraldo. Efficient construction of unified continuous and discontinuous Galerkin formulations for the 3D Euler equations. Journal of Computational Physics, 320(??):46–68, September 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301784.

Averkin:2018:PEP

[AG18] Sergey N. Averkin and Nikolaos A. Gatsonis. A parallel electrostatic Particle-in-Cell method on unstructured tetrahedral grids for large-scale bounded collisionless plasma simulations. Journal of Computational Physics, 363(??):178–199, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300871.

Aguirre:2015:UVC

[AGBL15] Miquel Aguirre, Antonio J. Gil, Javier Bonet, and Chun Hean Lee. An upwind vertex centred finite volume solver for Lagrangian solid dynamics. Journal of Computational Physics, 300(??):387–422, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911500474X.

Adams:2019:HOE

[AGC19] Thomas Adams, Stefano Giani, and William M. Coombs. A high-order elliptic PDE based level set reinitialisation method using a discontinuous Galerkin discretisation. Journal of Computational Physics, 379(??):373–391, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307915.

Alben:2019:SIM

[AGKD19] Silas Alben, Alex A. Gorodetsky, Donghak Kim, and Robert D. Deegan. Semi-implicit methods for the dynamics of elastic sheets. Journal of Computational Physics, 399 (??):Article 108952, December 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306576.

Alvarez:2018:PEE

[AGRB18] Mario Alvarez, Gabriel N. Gatica, and Ricardo Ruiz-Baier. A posteriori error estimation for an augmented mixed-primal method applied to sedimentation-consolidation systems. Journal of Computational Physics, 367(??):322–346, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302651.

Abdulle:2015:RBL

[AH15] Assyr Abdulle and Patrick Henning. A reduced basis localized orthogonal decomposition. Journal of Computational Physics, 295(??):379–401, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002624.

Ameline:2018:AEE

[AHHC18] Olivier Ameline, Sinan Haliyo, Xingxi Huang, and Jean A. H. Cognet. Analytical expression of elastic rods at equilibrium under 3D strong anchoring boundary conditions. Journal of Computational Physics, 373(??):736–749, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304820.

Arshad:2017:TST

[AHKT17] Sadia Arshad, Jianfei Huang, Abdul Q. M. Khaliq, and Yifa Tang. Trapezoidal scheme for time-space fractional diffusion equation with Riesz derivative. Journal of Computational Physics, 350(??):1–15, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306150.

Abe:2015:FPH

[AHNF15] Yoshiaki Abe, Takanori Haga, Taku Nonomura, and Kozo Fujii. On the freestream preservation of high-order conservative flux-reconstruction schemes. Journal of Computational Physics, 281(??):28–54, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006937.

Arbogast:2019:FVW

[AHZ19] Todd Arbogast, Chieh-Sen Huang, and Xikai Zhao. Finite volume WENO schemes for nonlinear parabolic problems with degenerate diffusion on non-uniform meshes. Journal of Computational Physics, 399(??):Article 108921, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306266.

Abbate:2017:ASR

[AIP17] Emanuela Abbate, Angelo Iollo, and Gabriella Puppo. An all-speed relaxation scheme for gases and compressible materials. Journal of Computational Physics, 351(??):1–24, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306332.

Ahn:2015:AMT

[AJP15] Chi Young Ahn, Kiwan Jeon, and Won-Kwang Park. Analysis of MUSIC-type imaging functional for single, thin electromagnetic inhomogeneity in limited-view inverse scattering problem. Journal of Computational Physics, 291(??):198–217, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001655.

Aursjo:2017:ILB

[AJVH17] Olav Aursjø, Espen Jettestuen, Jan Ludvig Vinningland, and Aksel Hiorth. An improved lattice Boltzmann method for simulating advective-diffusive processes in fluids. Journal of Computational Physics, 332(??):363–375, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306672.

An:2017:AAA

[AJW17] Hengbin An, Xiaowei Jia, and Homer F. Walker. Anderson acceleration and application to the three-temperature energy equations. Journal of Computational Physics, 347(??):1–19, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730476X.

Abgrall:2017:CSE

[AK17] Rémi Abgrall and Barry Koren. Computational science for energy research. Journal of Computational Physics, 345 (??):A1, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304552.

Aditya:2019:ADS

[AKK^{+}19] Konduri Aditya, Hemanth Kolla, W. Philip Kegelmeyer, Timothy M. Shead, Julia Ling, and Warren L. Davis. Anomaly detection in scientific data using joint statistical moments. Journal of Computational Physics, 387(??):522–538, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911930172X.

Aubry:2019:ESC

[AKM^{+}19] R. Aubry, B. K. Karamete, E. L. Mestreau, C. Jones, and S. Dey. Entropy solution at concave corners and ridges, and volume boundary layer tangential adaptivity. Journal of Computational Physics, 376(??):1–19, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306326.

afKlinteberg:2016:FIE

[aKT16] Ludvig af Klinteberg and Anna-Karin Tornberg. A fast integral equation method for solid particles in viscous flow using quadrature by expansion. Journal of Computational Physics, 326(??):420–445, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304181.

Aldegunde:2016:DEC

[AKZ16] Manuel Aldegunde, James R. Kermode, and Nicholas Zabaras. Development of an exchange-correlation functional with uncertainty quantification capabilities for density functional theory. Journal of Computational Physics, 311(??):173–195, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000425.

Antoine:2019:SPM

[AL19a] Xavier Antoine and Emmanuel Lorin. A simple pseudospectral method for the computation of the time-dependent Dirac equation with Perfectly Matched Layers. Journal of Computational Physics, 395(??):583–601, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304279.

Antoine:2019:TPM

[AL19b] Xavier Antoine and Emmanuel Lorin. Towards perfectly matched layers for time-dependent space fractional PDEs. Journal of Computational Physics, 391(??):59–90, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302712.

Azarnykh:2016:NMW

[ALA16] Dmitrii Azarnykh, Sergey Litvinov, and Nikolaus A. Adams. Numerical methods for the weakly compressible Generalized Langevin Model in Eulerian reference frame. Journal of Computational Physics, 314(??):93–106, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001558.

Alikhanov:2015:NDS

[Ali15] Anatoly A. Alikhanov. A new difference scheme for the time fractional diffusion equation. Journal of Computational Physics, 280(??):424–438, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006676.

Ahlkrona:2016:DCN

[ALKZ16] Josefin Ahlkrona, Per Lötstedt, Nina Kirchner, and Thomas Zwinger. Dynamically coupling the non-linear Stokes equations with the shallow ice approximation in glaciology: Description and first applications of the ISCAL method. Journal of Computational Physics, 308(??):1–19, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500844X.

Aoussou:2018:IPC

[ALL18] Jean Aoussou, Jing Lin, and Pierre F. J. Lermusiaux. Iterated pressure-correction projection methods for the unsteady incompressible Navier–Stokes equations. Journal of Computational Physics, 373(??):940–974, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304406.

Andriyash:2015:SUA

[ALM15] I. A. Andriyash, R. Lehe, and V. Malka. A spectral unaveraged algorithm for free electron laser simulations. Journal of Computational Physics, 282(??):397–409, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007888.

Algar:2017:EHT

[ALM^{+}17] María-Jesús Algar, Lorena Lozano, Javier Morneo, Iván González, and Felipe Cátedra. An efficient hybrid technique in RCS predictions of complex targets at high frequencies. Journal of Computational Physics, 345(??):345–357, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304102.

Almanasreh:2019:FEM

[Alm19] Hasan Almanasreh. Finite element method for solving the Dirac eigenvalue problem with linear basis functions. Journal of Computational Physics, 376(??):1199–1211, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306892.

Asthana:2015:NLS

[ALMJ15] Kartikey Asthana, Manuel R. López-Morales, and Antony Jameson. Non-linear stabilization of high-order flux reconstruction schemes via Fourier-spectral filtering. Journal of Computational Physics, 303(??):269–294, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006403.

Alauzet:2018:TAM

[ALO18] F. Alauzet, A. Loseille, and G. Olivier. Time-accurate multiscale anisotropic mesh adaptation for unsteady flows in CFD. Journal of Computational Physics, 373(??):28–63, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304212.

Antoine:2017:ESC

[ALT17] Xavier Antoine, Antoine Levitt, and Qinglin Tang. Efficient spectral computation of the stationary states of rotating Bose–Einstein condensates by preconditioned nonlinear conjugate gradient methods. Journal of Computational Physics, 343(??):92–109, August 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303194.

Auclair:2017:INM

[ALTR17] J.-P. Auclair, J.-F. Lemieux, L. B. Tremblay, and H. Ritchie. Implementation of Newton’s method with an analytical Jacobian to solve the 1D sea ice momentum equation. Journal of Computational Physics, 340(??):69–84, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301687.

Abedi:2017:ASD

[AM17a] Reza Abedi and Saba Mudaliar. An asynchronous spacetime discontinuous Galerkin finite element method for time domain electromagnetics. Journal of Computational Physics, 351(??):121–144, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306514.

Asaithambi:2017:NCF

[AM17b] Rajapandiyan Asaithambi and Krishnan Mahesh. A note on a conservative finite volume approach to address numerical stiffness in polar meshes. Journal of Computational Physics, 341(??):377–385, July 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302966.

Alemazkoor:2018:NOS

[AM18] Negin Alemazkoor and Hadi Meidani. A near-optimal sampling strategy for sparse recovery of polynomial chaos expansions. Journal of Computational Physics, 371(??):137–151, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303255.

Anantharamu:2019:PSD

[AM19] Sreevatsa Anantharamu and Krishnan Mahesh. A parallel and streaming Dynamic Mode Decomposition algorithm with finite precision error analysis for large data. Journal of Computational Physics, 380(??):355–377, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118308088.

Amano:2015:DFA

[Ama15] Takanobu Amano. Divergence-free approximate Riemann solver for the quasi-neutral two-fluid plasma model. Journal of Computational Physics, 299(??):863–886, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004805.

Amano:2018:GQN

[Ama18] Takanobu Amano. A generalized quasi-neutral fluid-particle hybrid plasma model and its application to energetic-particle-magnetohydrodynamics hybrid simulation. Journal of Computational Physics, 366(??):366–385, August 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302365.

Akiki:2017:PIE

[AMB17] G. Akiki, W. C. Moore, and S. Balachandar. Pairwiseinteraction extended point-particle model for particle-laden flows. Journal of Computational Physics, 351(??):329–357, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306848.

Abe:2018:SND

[AMH^{+}18] Yoshiaki Abe, Issei Morinaka, Takanori Haga, Taku Nonomura, Hisaichi Shibata, and Koji Miyaji. Stable, non-dissipative, and conservative flux-reconstruction schemes in split forms. Journal of Computational Physics, 353(??):193–227, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307453.

Aguilar:2017:ABS

[AMJ17] José G. Aguilar, Luca Magri, and Matthew P. Juniper. Adjoint-based sensitivity analysis of low-order thermoacoustic networks using a wave-based approach. Journal of Computational Physics, 341(??):163–181, July 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730284X.

Ackmann:2017:SGO

[AMK17] Jan Ackmann, Jochem Marotzke, and Peter Korn. Stochastic goal-oriented error estimation with memory. Journal of Computational Physics, 348(??):195–219, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305119.

Arne:2015:FVA

[AMM^{+}15] Walter Arne, Nicole Marheineke, Andreas Meister, Stefan Schiessl, and Raimund Wegener. Finite volume approach for the instationary Cosserat rod model describing the spinning of viscous jets. Journal of Computational Physics, 294(??):20–37, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001977.

Abraham:2018:CFM

[AMN18] David S. Abraham, Alexandre Noll Marques, and Jean-Christophe Nave. A correction function method for the wave equation with interface jump conditions. Journal of Computational Physics, 353(??):281–299, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307684.

Alonso-Mallo:2016:TES

[AMP16] I. Alonso-Mallo and A. M. Portillo. Time exponential splitting technique for the Klein–Gordon equation with hagstromwarburton high-order absorbing boundary conditions. Journal of Computational Physics, 311(??):196–212, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000589.

Arcucci:2019:ORS

[AMPG19] Rossella Arcucci, Laetitia Mottet, Christopher Pain, and Yi-Ke Guo. Optimal reduced space for Variational Data Assimilation. Journal of Computational Physics, 379(??):51–69, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307095.

Al-Marouf:2017:VEB

[AMS17] M. Al-Marouf and R. Samtaney. A versatile embedded boundary adaptive mesh method for compressible flow in complex geometry. Journal of Computational Physics, 337(??):339–378, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730147X.

An:2019:OBP

[AMXJ19] Hengbin An, Zeyao Mo, Xiaowen Xu, and Xiaowei Jia. Operator-based preconditioning for the 2-D 3-T energy equations in radiation hydrodynamics simulations. Journal of Computational Physics, 385(??):51–74, May 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301019.

Amirkulova:2015:AMS

[AN15] Feruza A. Amirkulova and Andrew N. Norris. Acoustic multiple scattering using recursive algorithms. Journal of Computational Physics, 299(??):787–803, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004763.

An:2017:UDR

[An17] Yajun An. Uniform dispersion reduction schemes for the one dimensional wave equation in isotropic media. Journal of Computational Physics, 341(??):13–21, July 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302863.

Aalund:2019:EHO

[ÅN19] Oskar Ålund and Jan Nordström. Encapsulated high order difference operators on curvilinear non-conforming grids. Journal of Computational Physics, 385(??):209–224, May 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301184.

Anderson:2016:HOE

[And16] Christopher R. Anderson. High order expanding domain methods for the solution of Poisson’s equation in infinite domains. Journal of Computational Physics, 314(??):194–205, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001613.

Aniszewski:2016:ITD

[Ani16] Wojciech Aniszewski. Improvements, testing and development of the ADM-τ sub-grid surface tension model for two-phase LES. Journal of Computational Physics, 327(??):389–415, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304545.

Anistratov:2019:SAM

[Ani19] Dmitriy Y. Anistratov. Stability analysis of a multilevel quasidiffusion method for thermal radiative transfer problems. Journal of Computational Physics, 376(??):186–209, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306363.

Augustin:2016:AAH

[ANL^{+}16] Christoph M. Augustin, Aurel Neic, Manfred Liebmann, Anton J. Prassl, Steven A. Niederer, Gundolf Haase, and Gernot Plank. Anatomically accurate high resolution modeling of human whole heart electromechanics: a strongly scalable algebraic multigrid solver method for nonlinear deformation. Journal of Computational Physics, 305(??):622–646, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007226.

Anonymous:2015:Ca

[Ano15a] Anonymous. Contents. Journal of Computational Physics, 280(??):IBC, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114007359.

Anonymous:2015:Cb

[Ano15b] Anonymous. Contents. Journal of Computational Physics, 280(??):OBC, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114007360.

Anonymous:2015:Cc

[Ano15c] Anonymous. Contents. Journal of Computational Physics, 281(??):IBC, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114007724.

Anonymous:2015:Cd

[Ano15d] Anonymous. Contents. Journal of Computational Physics, 281(??):OBC, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114007736.

Anonymous:2015:Ce

[Ano15e] Anonymous. Contents. Journal of Computational Physics, 282(??):IBC, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114008286.

Anonymous:2015:Cf

[Ano15f] Anonymous. Contents. Journal of Computational Physics, 282(??):OBC, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114008298.

Anonymous:2015:Cg

[Ano15g] Anonymous. Contents. Journal of Computational Physics, 283(??):IBC, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114008651.

Anonymous:2015:Ch

[Ano15h] Anonymous. Contents. Journal of Computational Physics, 283(??):OBC, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114008663.

Anonymous:2015:Ci

[Ano15i] Anonymous. Contents. Journal of Computational Physics, 284(??):IBC, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115000455.

Anonymous:2015:Cj

[Ano15j] Anonymous. Contents. Journal of Computational Physics, 284(??):OBC, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115000467.

Anonymous:2015:Ck

[Ano15k] Anonymous. Contents. Journal of Computational Physics, 285(??):IBC, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115000601.

Anonymous:2015:Cl

[Ano15l] Anonymous. Contents. Journal of Computational Physics, 285(??):OBC, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115000613.

Anonymous:2015:Cm

[Ano15m] Anonymous. Contents. Journal of Computational Physics, 286(??):OBC, April 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115000819.

Anonymous:2015:Cn

[Ano15n] Anonymous. Contents. Journal of Computational Physics, 287(??):OBC–, 2015. CODEN JCTPAH. ISSN 0021-9991.

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001291.

Anonymous:2015:Co

[Ano15o] Anonymous. Contents. Journal of Computational Physics, 288(??):OBC, May 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115001370.

Anonymous:2015:Cp

[Ano15p] Anonymous. Contents. Journal of Computational Physics, 289(??):OBC, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115001606.

Anonymous:2015:Cq

[Ano15q] Anonymous. Contents. Journal of Computational Physics, 290(??):IBC, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115001850.

Anonymous:2015:Cr

[Ano15r] Anonymous. Contents. Journal of Computational Physics, 290(??):OBC, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115001862.

Anonymous:2015:Cs

[Ano15s] Anonymous. Contents. Journal of Computational Physics, 291(??):ibc, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115002120.

Anonymous:2015:Ct

[Ano15t] Anonymous. Contents. Journal of Computational Physics, 291(??):obc, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115002132.

Anonymous:2015:Cu

[Ano15u] Anonymous. Contents. Journal of Computational Physics, 292(??):ibc, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115002351.

Anonymous:2015:Cv

[Ano15v] Anonymous. Contents. Journal of Computational Physics, 292(??):obc, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115002363.

Anonymous:2015:Cw

[Ano15w] Anonymous. Contents. Journal of Computational Physics, 293(??):ibc, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911500251X.

Anonymous:2015:Cx

[Ano15x] Anonymous. Contents. Journal of Computational Physics, 293(??):obc, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115002521.

Anonymous:2015:Cy

[Ano15y] Anonymous. Contents. Journal of Computational Physics, 294(??):ibc, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115003174.

Anonymous:2015:Cz

[Ano15z] Anonymous. Contents. Journal of Computational Physics, 294(??):obc, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115003186.

Anonymous:2015:Caa

[Ano15-27] Anonymous. Contents. Journal of Computational Physics, 295(??):ibc, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115003344.

Anonymous:2015:Cab

[Ano15-28] Anonymous. Contents. Journal of Computational Physics, 295(??):obc, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115003356.

Anonymous:2015:Cac

[Ano15-29] Anonymous. Contents. Journal of Computational Physics, 296(??):ibc, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115003502.

Anonymous:2015:Cad

[Ano15-30] Anonymous. Contents. Journal of Computational Physics, 296(??):obc, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115003514.

Anonymous:2015:Cae

[Ano15-31] Anonymous. Contents. Journal of Computational Physics, 297(??):ibc, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004076.

Anonymous:2015:Caf

[Ano15-32] Anonymous. Contents. Journal of Computational Physics, 297(??):obc, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004088.

Anonymous:2015:Cag

[Ano15-33] Anonymous. Contents. Journal of Computational Physics, 298(??):ibc, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004970.

Anonymous:2015:Cah

[Ano15-34] Anonymous. Contents. Journal of Computational Physics, 298(??):obc, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004982.

Anonymous:2015:Cai

[Ano15-35] Anonymous. Contents. Journal of Computational Physics, 299(??):ibc, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115005410.

Anonymous:2015:Caj

[Ano15-36] Anonymous. Contents. Journal of Computational Physics, 299(??):obc, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115005422.

Anonymous:2015:Cak

[Ano15-37] Anonymous. Contents. Journal of Computational Physics, 300(??):ibc, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115006038.

Anonymous:2015:Cal

[Ano15-38] Anonymous. Contents. Journal of Computational Physics, 300(??):obc, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911500604X.

Anonymous:2015:Cam

[Ano15-39] Anonymous. Contents. Journal of Computational Physics, 301(??):ibc, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115006385.

Anonymous:2015:Can

[Ano15-40] Anonymous. Contents. Journal of Computational Physics, 301(??):obc, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115006397.

Anonymous:2015:Cao

[Ano15-41] Anonymous. Contents. Journal of Computational Physics, 302(??):ibc, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911500649X.

Anonymous:2015:Cap

[Ano15-42] Anonymous. Contents. Journal of Computational Physics, 302(??):obc, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115006506.

Anonymous:2015:Caq

[Ano15-43] Anonymous. Contents. Journal of Computational Physics, 303(??):ibc, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115007044.

Anonymous:2015:Car

[Ano15-44] Anonymous. Contents. Journal of Computational Physics, 303(??):obc, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115007056.

Anonymous:2016:Ca

[Ano16a] Anonymous. Contents. Journal of Computational Physics, 304(??):ibc, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115007469.

Anonymous:2016:Cb

[Ano16b] Anonymous. Contents. Journal of Computational Physics, 304(??):obc, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115007470.

Anonymous:2016:Cc

[Ano16c] Anonymous. Contents. Journal of Computational Physics, 305(??):ibc, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115007986.

Anonymous:2016:Cd

[Ano16d] Anonymous. Contents. Journal of Computational Physics, 305(??):obc, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115007998.

Anonymous:2016:Ce

[Ano16e] Anonymous. Contents. Journal of Computational Physics, 306(??):ibc, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911500813X.

Anonymous:2016:Cf

[Ano16f] Anonymous. Contents. Journal of Computational Physics, 306(??):obc, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115008141.

Anonymous:2016:Cg

[Ano16g] Anonymous. Contents. Journal of Computational Physics, 307(??):ibc, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115008761.

Anonymous:2016:Ch

[Ano16h] Anonymous. Contents. Journal of Computational Physics, 307(??):obc, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115008773.

Anonymous:2016:Ci

[Ano16i] Anonymous. Contents. Journal of Computational Physics, 308(??):ibc, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116000243.

Anonymous:2016:Cj

[Ano16j] Anonymous. Contents. Journal of Computational Physics, 308(??):obc, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116000255.

Anonymous:2016:Ck

[Ano16k] Anonymous. Contents. Journal of Computational Physics, 309(??):ibc, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116000474.

Anonymous:2016:Cl

[Ano16l] Anonymous. Contents. Journal of Computational Physics, 309(??):obc, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116000486.

Anonymous:2016:Cm

[Ano16m] Anonymous. Contents. Journal of Computational Physics, 311(??):ibc, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116001091.

Anonymous:2016:Cn

[Ano16n] Anonymous. Contents. Journal of Computational Physics, 311(??):obc, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116001108.

Anonymous:2016:Co

[Ano16o] Anonymous. Contents. Journal of Computational Physics, 312(??):ibc, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116001510.

Anonymous:2016:Cp

[Ano16p] Anonymous. Contents. Journal of Computational Physics, 312(??):obc, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116001522.

Anonymous:2016:Cq

[Ano16q] Anonymous. Contents. Journal of Computational Physics, 313(??):ibc, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116300055.

Anonymous:2016:Cr

[Ano16r] Anonymous. Contents. Journal of Computational Physics, 313(??):obc, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116300067.

Anonymous:2016:Cs

[Ano16s] Anonymous. Contents. Journal of Computational Physics, 314(??):ii, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116300572.

Anonymous:2016:Ct

[Ano16t] Anonymous. Contents. Journal of Computational Physics, 314(??):ibc, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116300559.

Anonymous:2016:Cu

[Ano16u] Anonymous. Contents. Journal of Computational Physics, 314(??):obc, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116300560.

Anonymous:2016:Cv

[Ano16v] Anonymous. Contents. Journal of Computational Physics, 315(??):ibc, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116300961.

Anonymous:2016:Cw

[Ano16w] Anonymous. Contents. Journal of Computational Physics, 315(??):obc, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116300973.

Anonymous:2016:Cx

[Ano16x] Anonymous. Contents. Journal of Computational Physics, 316(??):iv, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116301218.

Anonymous:2016:Cy

[Ano16y] Anonymous. Contents. Journal of Computational Physics, 316(??):ibc, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911630119X.

Anonymous:2016:Cz

[Ano16z] Anonymous. Contents. Journal of Computational Physics, 316(??):obc, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116301206.

Anonymous:2016:Caa

[Ano16-27] Anonymous. Contents. Journal of Computational Physics, 317(??):ibc, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116301462.

Anonymous:2016:Cab

[Ano16-28] Anonymous. Contents. Journal of Computational Physics, 317(??):obc, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116301474.

Anonymous:2016:Cac

[Ano16-29] Anonymous. Contents. Journal of Computational Physics, 318(??):ibc, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116301693.

Anonymous:2016:Cad

[Ano16-30] Anonymous. Contents. Journal of Computational Physics, 318(??):obc, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911630170X.

Anonymous:2016:Cae

[Ano16-31] Anonymous. Contents. Journal of Computational Physics, 319(??):ibc, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911630211X.

Anonymous:2016:Caf

[Ano16-32] Anonymous. Contents. Journal of Computational Physics, 319(??):obc, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116302121.

Anonymous:2016:Cag

[Ano16-33] Anonymous. Contents. Journal of Computational Physics, 320(??):obc, September 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116302200.

Anonymous:2016:Cah

[Ano16-34] Anonymous. Contents. Journal of Computational Physics, 321(??):ibc, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303023.

Anonymous:2016:Cai

[Ano16-35] Anonymous. Contents. Journal of Computational Physics, 321(??):obc, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303035.

Anonymous:2016:Caj

[Ano16-36] Anonymous. Contents. Journal of Computational Physics, 322(??):ibc, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303370.

Anonymous:2016:Cak

[Ano16-37] Anonymous. Contents. Journal of Computational Physics, 322(??):obc, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303382.

Anonymous:2016:Cal

[Ano16-38] Anonymous. Contents. Journal of Computational Physics, 323(??):ibc, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303606.

Anonymous:2016:Cam

[Ano16-39] Anonymous. Contents. Journal of Computational Physics, 323(??):obc, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303618.

Anonymous:2016:Can

[Ano16-40] Anonymous. Contents. Journal of Computational Physics, 324(??):ibc, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303801.

Anonymous:2016:Cao

[Ano16-41] Anonymous. Contents. Journal of Computational Physics, 324(??):obc, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303813.

Anonymous:2016:Cap

[Ano16-42] Anonymous. Contents. Journal of Computational Physics, 325(??):ibc, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304478.

Anonymous:2016:Caq

[Ano16-43] Anonymous. Contents. Journal of Computational Physics, 325(??):obc, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911630448X.

Anonymous:2016:CCe

[Ano16-44] Anonymous. Contents. Journal of Computational Physics, 326(??):ibc, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304909.

Anonymous:2016:CCf

[Ano16-45] Anonymous. Contents. Journal of Computational Physics, 326(??):obc, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304910.

Anonymous:2016:Car

[Ano16-46] Anonymous. Contents. Journal of Computational Physics, 327(??):ibc, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116305460.

Anonymous:2016:Cas

[Ano16-47] Anonymous. Contents. Journal of Computational Physics, 327(??):obc, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116305472.

Anonymous:2016:CCa

[Ano16-48] Anonymous. Contents continued. Journal of Computational Physics, 321(??):iii, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303084.

Anonymous:2016:CCb

[Ano16-49] Anonymous. Contents continued. Journal of Computational Physics, 321(??):iv, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303060.

Anonymous:2016:CCc

[Ano16-50] Anonymous. Contents continued. Journal of Computational Physics, 322(??):ii, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303515.

Anonymous:2016:CCd

[Ano16-51] Anonymous. Contents continued. Journal of Computational Physics, 326(??):iv, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305034.

Anonymous:2016:CCg

[Ano16-52] Anonymous. Contents continued. Journal of Computational Physics, 327(??):iv, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305514.

Anonymous:2017:Ca

[Ano17a] Anonymous. Contents. Journal of Computational Physics, 328(??):ibc, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116305885.

Anonymous:2017:Cb

[Ano17b] Anonymous. Contents. Journal of Computational Physics, 328(??):obc, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116305897.

Anonymous:2017:Cc

[Ano17c] Anonymous. Contents. Journal of Computational Physics, 329(??):obc, January 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306210.

Anonymous:2017:Cd

[Ano17d] Anonymous. Contents. Journal of Computational Physics, 330(??):ibc, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306581.

Anonymous:2017:Ce

[Ano17e] Anonymous. Contents. Journal of Computational Physics, 330(??):obc, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306593.

Anonymous:2017:Cf

[Ano17f] Anonymous. Contents. Journal of Computational Physics, 331(??):ibc, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306921.

Anonymous:2017:Cg

[Ano17g] Anonymous. Contents. Journal of Computational Physics, 331(??):obc, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306933.

Anonymous:2017:Ch

[Ano17h] Anonymous. Contents. Journal of Computational Physics, 332(??):ibc, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300050.

Anonymous:2017:Ci

[Ano17i] Anonymous. Contents. Journal of Computational Physics, 332(??):obc, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300062.

Anonymous:2017:Cj

[Ano17j] Anonymous. Contents. Journal of Computational Physics, 333(??):ibc, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300396.

Anonymous:2017:Ck

[Ano17k] Anonymous. Contents. Journal of Computational Physics, 333(??):obc, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300402.

Anonymous:2017:Cl

[Ano17l] Anonymous. Contents. Journal of Computational Physics, 334(??):ibc, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300906.

Anonymous:2017:Cm

[Ano17m] Anonymous. Contents. Journal of Computational Physics, 334(??):obc, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300918.

Anonymous:2017:Cn

[Ano17n] Anonymous. Contents. Journal of Computational Physics, 335(??):obc, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117301390.

Anonymous:2017:Co

[Ano17o] Anonymous. Contents. Journal of Computational Physics, 336(??):ibc, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911730178X.

Anonymous:2017:Cp

[Ano17p] Anonymous. Contents. Journal of Computational Physics, 336(??):obc, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117301791.

Anonymous:2017:Cq

[Ano17q] Anonymous. Contents. Journal of Computational Physics, 337(??):ibc, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302115.

Anonymous:2017:Cr

[Ano17r] Anonymous. Contents. Journal of Computational Physics, 337(??):obc, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302127.

Anonymous:2017:Cs

[Ano17s] Anonymous. Contents. Journal of Computational Physics, 338(??):ibc, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302450.

Anonymous:2017:Ct

[Ano17t] Anonymous. Contents. Journal of Computational Physics, 338(??):obc, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302462.

Anonymous:2017:Cu

[Ano17u] Anonymous. Contents. Journal of Computational Physics, 339(??):ibc, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302693.

Anonymous:2017:Cv

[Ano17v] Anonymous. Contents. Journal of Computational Physics, 339(??):obc, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911730270X.

Anonymous:2017:Cw

[Ano17w] Anonymous. Contents. Journal of Computational Physics, 340(??):ibc, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303121.

Anonymous:2017:Cx

[Ano17x] Anonymous. Contents. Journal of Computational Physics, 340(??):obc, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303133.

Anonymous:2017:Cy

[Ano17y] Anonymous. Contents. Journal of Computational Physics, 341(??):ibc, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303509.

Anonymous:2017:Cz

[Ano17z] Anonymous. Contents. Journal of Computational Physics, 341(??):obc, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303510.

Anonymous:2017:Caa

[Ano17-27] Anonymous. Contents. Journal of Computational Physics, 342(??):ibc, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303911.

Anonymous:2017:Cab

[Ano17-28] Anonymous. Contents. Journal of Computational Physics, 342(??):obc, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303923.

Anonymous:2017:Cac

[Ano17-29] Anonymous. Contents. Journal of Computational Physics, 343(??):ibc, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304175.

Anonymous:2017:Cad

[Ano17-30] Anonymous. Contents. Journal of Computational Physics, 343(??):obc, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304187.

Anonymous:2017:Cae

[Ano17-31] Anonymous. Contents. Journal of Computational Physics, 344(??):ibc, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304369.

Anonymous:2017:Caf

[Ano17-32] Anonymous. Contents. Journal of Computational Physics, 344(??):obc, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304370.

Anonymous:2017:Cag

[Ano17-33] Anonymous. Contents. Journal of Computational Physics, 345(??):obc, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304849.

Anonymous:2017:Cah

[Ano17-34] Anonymous. Contents. Journal of Computational Physics, 346(??):ibc, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117305259.

Anonymous:2017:Cai

[Ano17-35] Anonymous. Contents. Journal of Computational Physics, 346(??):obc, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117305260.

Anonymous:2017:Caj

[Ano17-36] Anonymous. Contents. Journal of Computational Physics, 347(??):ibc, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117305685.

Anonymous:2017:Cak

[Ano17-37] Anonymous. Contents. Journal of Computational Physics, 347(??):obc, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117305697.

Anonymous:2017:Cam

[Ano17-38] Anonymous. Contents. Journal of Computational Physics, 348(??):ibc, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117306113.

Anonymous:2017:Can

[Ano17-39] Anonymous. Contents. Journal of Computational Physics, 348(??):obc, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117306125.

Anonymous:2017:Cao

[Ano17-40] Anonymous. Contents. Journal of Computational Physics, 349(??):ibc, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117306642.

Anonymous:2017:Cap

[Ano17-41] Anonymous. Contents. Journal of Computational Physics, 349(??):obc, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117306654.

Anonymous:2017:Caq

[Ano17-42] Anonymous. Contents. Journal of Computational Physics, 350(??):ibc, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117307222.

Anonymous:2017:Car

[Ano17-43] Anonymous. Contents. Journal of Computational Physics, 350(??):obc, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117307234.

Anonymous:2017:Cas

[Ano17-44] Anonymous. Contents. Journal of Computational Physics, 351(??):ibc, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911730757X.

Anonymous:2017:Cat

[Ano17-45] Anonymous. Contents. Journal of Computational Physics, 351(??):obc, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117307581.

Anonymous:2017:CCa

[Ano17-46] Anonymous. Contents continued. Journal of Computational Physics, 330(??):iii, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911630660X.

Anonymous:2017:CCb

[Ano17-47] Anonymous. Contents continued. Journal of Computational Physics, 330(??):iv, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306611.

Anonymous:2017:CCc

[Ano17-48] Anonymous. Contents continued. Journal of Computational Physics, 335(??):ii, April 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301419.

Anonymous:2017:CCd

[Ano17-49] Anonymous. Contents continued. Journal of Computational Physics, 335(??):ibc, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301389.

Anonymous:2017:CCe

[Ano17-50] Anonymous. Contents continued. Journal of Computational Physics, 345(??):ii, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304898.

Anonymous:2017:CCf

[Ano17-51] Anonymous. Contents continued. Journal of Computational Physics, 345(??):ibc, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304837.

Anonymous:2017:Cal

[Ano17-52] Anonymous. Contents continued. Journal of Computational Physics, 348(??):iv, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306204.

Anonymous:2017:CCg

[Ano17-53] Anonymous. Contents continued. Journal of Computational Physics, 350(??):ii, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117307246.

Anonymous:2018:Ca

[Ano18a] Anonymous. Contents. Journal of Computational Physics, 352(??):ibc, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117307891.

Anonymous:2018:Cb

[Ano18b] Anonymous. Contents. Journal of Computational Physics, 352(??):obc, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117307908.

Anonymous:2018:Cc

[Ano18c] Anonymous. Contents. Journal of Computational Physics, 353(??):ibc, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117308446.

Anonymous:2018:Cd

[Ano18d] Anonymous. Contents. Journal of Computational Physics, 353(??):obc, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117308458.

Anonymous:2018:EBa

[Ano18e] Anonymous. Editorial Board. Journal of Computational Physics, 354(??):ii, February 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117308902.

Anonymous:2018:EBb

[Ano18f] Anonymous. Editorial Board. Journal of Computational Physics, 355(??):ii, February 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117309117.

Anonymous:2018:EBc

[Ano18g] Anonymous. Editorial Board. Journal of Computational Physics, 356(??):ii, March 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300111.

Anonymous:2018:EBd

[Ano18h] Anonymous. Editorial Board. Journal of Computational Physics, 357(??):ii, March 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300639.

Anonymous:2018:EBe

[Ano18i] Anonymous. Editorial Board. Journal of Computational Physics, 358(??):ii, April 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300706.

Anonymous:2018:EBf

[Ano18j] Anonymous. Editorial Board. Journal of Computational Physics, 359(??):ii, April 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300949.

Anonymous:2018:EBg

[Ano18k] Anonymous. Editorial Board. Journal of Computational Physics, 360(??):ii, May 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301335.

Anonymous:2018:EBh

[Ano18l] Anonymous. Editorial Board. Journal of Computational Physics, 361(??):ii, May 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301517.

Anonymous:2018:EBi

[Ano18m] Anonymous. Editorial Board. Journal of Computational Physics, 362(??):ii, June 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301748.

Anonymous:2018:EBj

[Ano18n] Anonymous. Editorial Board. Journal of Computational Physics, 363(??):ii, June 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302018.

Anonymous:2018:EBk

[Ano18o] Anonymous. Editorial Board. Journal of Computational Physics, 364(??):ii, July 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302481.

Anonymous:2018:EBl

[Ano18p] Anonymous. Editorial Board. Journal of Computational Physics, 367(??):ii, August 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303322.

Anonymous:2018:EBm

[Ano18q] Anonymous. Editorial Board. Journal of Computational Physics, 368(??):ii, September 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303607.

Anonymous:2018:EBo

[Ano18r] Anonymous. Editorial Board. Journal of Computational Physics, 368(??):ii, September 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303607.

Anonymous:2018:EBn

[Ano18s] Anonymous. Editorial Board. Journal of Computational Physics, 369(??):ii, September 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303668.

Anonymous:2018:EBp

[Ano18t] Anonymous. Editorial Board. Journal of Computational Physics, 370(??):ii, October 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303759.

Anonymous:2018:EBq

[Ano18u] Anonymous. Editorial Board. Journal of Computational Physics, 371(??):ii, October 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304698.

Anonymous:2018:EBr

[Ano18v] Anonymous. Editorial Board. Journal of Computational Physics, 372(??):ii, November 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305606.

Anonymous:2018:EBs

[Ano18w] Anonymous. Editorial Board. Journal of Computational Physics, 373(??):ii, November 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830617X.

Anonymous:2018:EBt

[Ano18x] Anonymous. Editorial Board. Journal of Computational Physics, 374(??):ii, December 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306776.

Anonymous:2018:EB

[Ano18y] Anonymous. Editorial Board. Journal of Computational Physics, 375(??):ii, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830740X.

Anonymous:2019:Da

[Ano19a] Anonymous. 1 December 2019. Journal of Computational Physics, 398(??):??, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Anonymous:2019:Db

[Ano19b] Anonymous. 15 December 2019. Journal of Computational Physics, 399(??):??, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Anonymous:2019:N

[Ano19c] Anonymous. 15 November 2019. Journal of Computational Physics, 397(??):??, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Anonymous:2019:EBa

[Ano19d] Anonymous. Editorial Board. Journal of Computational Physics, 376(??):ii, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118307599.

Anonymous:2019:EBb

[Ano19e] Anonymous. Editorial Board. Journal of Computational Physics, 377(??):ii, January 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118307538.

Anonymous:2019:EBc

[Ano19f] Anonymous. Editorial Board. Journal of Computational Physics, 378(??):ii, February 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830799X.

Anonymous:2019:EBd

[Ano19g] Anonymous. Editorial Board. Journal of Computational Physics, 379(??):ii, February 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300324.

Anonymous:2019:EBe

[Ano19h] Anonymous. Editorial Board. Journal of Computational Physics, 380(??):ii, March 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930052X.

Anonymous:2019:EBg

[Ano19i] Anonymous. Editorial Board. Journal of Computational Physics, 381(??):ii, March 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300580.

Anonymous:2019:EBh

[Ano19j] Anonymous. Editorial Board. Journal of Computational Physics, 382(??):ii, April 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300725.

Anonymous:2019:EBi

[Ano19k] Anonymous. Editorial Board. Journal of Computational Physics, 383(??):ii, April 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301263.

Anonymous:2019:EBj

[Ano19l] Anonymous. Editorial Board. Journal of Computational Physics, 384(??):ii, May 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301779.

Anonymous:2019:EBk

[Ano19m] Anonymous. Editorial Board. Journal of Computational Physics, 385(??):ii, May 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301834.

Anonymous:2019:EBl

[Ano19n] Anonymous. Editorial Board. Journal of Computational Physics, 386(??):ii, June 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930261X.

Anonymous:2019:EBm

[Ano19o] Anonymous. Editorial Board. Journal of Computational Physics, 387(??):ii, June 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302670.

Anonymous:2019:EBn

[Ano19p] Anonymous. Editorial Board. Journal of Computational Physics, 388(??):ii, July 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302992.

Anonymous:2019:EBo

[Ano19q] Anonymous. Editorial Board. Journal of Computational Physics, 389(??):ii, July 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303456.

Anonymous:2019:EBp

[Ano19r] Anonymous. Editorial Board. Journal of Computational Physics, 390(??):ii, August 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303663.

Anonymous:2019:EBq

[Ano19s] Anonymous. Editorial Board. Journal of Computational Physics, 391(??):ii, August 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303778.

Anonymous:2019:EBr

[Ano19t] Anonymous. Editorial Board. Journal of Computational Physics, 392(??):ii, September 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303936.

Anonymous:2019:EBs

[Ano19u] Anonymous. Editorial Board. Journal of Computational Physics, 393(??):ii, September 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119304085.

Anonymous:2019:EBt

[Ano19v] Anonymous. Editorial Board. Journal of Computational Physics, 394(??):ii, October 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119305510.

Anonymous:2019:EBu

[Ano19w] Anonymous. Editorial Board. Journal of Computational Physics, 395(??):ii, October 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305704.

Anonymous:2019:EBv

[Ano19x] Anonymous. Editorial Board. Journal of Computational Physics, 396(??):ii, November 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306047.

Anonymous:2019:EBw

[Ano19y] Anonymous. Editorial Board. Journal of Computational Physics, 397(??):Article 108935, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119306400.

Anonymous:2019:EBx

[Ano19z] Anonymous. Editorial Board. Journal of Computational Physics, 398(??):Article 108984, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306898.

Anonymous:2019:EBy

[Ano19-27] Anonymous. Editorial Board. Journal of Computational Physics, 399(??):Article 109037, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119307430.

Anonymous:2019:PO

[Ano19-28] Anonymous. Pages 1-772 (15 October 2019). Journal of Computational Physics, 395(??):1–772, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Anonymous:2019:PN

[Ano19-29] Anonymous. Pages 1-904 (1 November 2019). Journal of Computational Physics, 396(??):1–904, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Antunes:2017:IPT

[Ant17] Pedro R. S. Antunes. Is it possible to tune a drum? Journal of Computational Physics, 338(??):91–106, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301596.

Altsybeyev:2016:AGL

[AP16] V. V. Altsybeyev and V. A. Ponomarev. Application of Gauss’s law space-charge limited emission model in iterative particle tracking method. Journal of Computational Physics, 324(??):62–72, November 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303473.

Anand:2016:EHO

[APKP16] Akash Anand, Ambuj Pandey, B. V. Rathish Kumar, and Jagabandhu Paul. An efficient high-order Nyström scheme for acoustic scattering by inhomogeneous penetrable media with discontinuous material interface. Journal of Computational Physics, 311(??):258–274, April 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911600036X.

Alsayyari:2019:NRO

[APLK19] Fahad Alsayyari, Zoltán Perkó, Danny Lathouwers, and Jan Leen Kloosterman. A nonintrusive reduced order modelling approach using proper orthogonal decomposition and locally adaptive sparse grids. Journal of Computational Physics, 399(??):Article 108912, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306175.

Adam:2016:HOC

[APP^{+}16] A. Adam, D. Pavlidis, J. R. Percival, P. Salinas, Z. Xie, F. Fang, C. C. Pain, A. H. Muggeridge, and M. D. Jackson. Higher-order conservative interpolation between controlvolume meshes: Application to advection and multiphase flow problems with dynamic mesh adaptivity. Journal of Computational Physics, 321(??):512–531, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302030.

Acosta:2015:NMC

[APR^{+}15] Sebastian Acosta, Charles Puelz, Béatrice Rivière, Daniel J. Penny, and Craig G. Rusin. Numerical method of characteristics for one-dimensional blood flow. Journal of Computational Physics, 294(??):96–109, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002004.

Ammar:2017:MTD

[APT17] Sami Ammar, Guillaume Pernaudat, and Jean-Yves Trépanier. A multiphase three-dimensional multi-relaxation time (MRT) lattice Boltzmann model with surface tension adjustment. Journal of Computational Physics, 343(??):73–91, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303248.

Aguerre:2018:OFF

[APV^{+}18] Horacio J. Aguerre, Cesar I. Pairetti, Cesar M. Venier, Santiago Márquez Damián, and Norberto M. Nigro. An oscillationfree flow solver based on flux reconstruction. Journal of Computational Physics, 365(??):135–148, July 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301931.

Abali:2019:TCE

[AQ19] B. E. Abali and A. F. Queiruga. Theory and computation of electromagnetic fields and thermomechanical structure interaction for systems undergoing large deformations. Journal of Computational Physics, 394(??):200–231, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303894.

Acebron:2016:MCM

[AR16a] Juan A. Acebrón and Marco A. Ribeiro. A Monte Carlo method for solving the one-dimensional telegraph equations with boundary conditions. Journal of Computational Physics, 305(??):29–43, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006919.

Arjmand:2016:TDA

[AR16b] Doghonay Arjmand and Olof Runborg. A time dependent approach for removing the cell boundary error in elliptic homogenization problems. Journal of Computational Physics, 314(??):206–227, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001601.

Abadi:2018:CPF

[ARF18] Reza Haghani Hassan Abadi, Mohammad Hassan Rahimian, and Abbas Fakhari. Conservative phase-field lattice-Boltzmann model for ternary fluids. Journal of Computational Physics, 374(??):668–691, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118305060.

See corrigendum [Hag25].

Abraham:2017:RES

[ARG^{+}17] Simon Abraham, Mehrdad Raisee, Ghader Ghorbaniasl, Francesco Contino, and Chris Lacor. A robust and efficient stepwise regression method for building sparse polynomial chaos expansions. Journal of Computational Physics, 332(??):461–474, March 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306684.

Aristizabal:2019:QOB

[ARTG^{+}19] Mauricio Aristizabal, Daniel Ramirez-Tamayo, Manuel Garcia, Andres Aguirre-Mesa, Arturo Montoya, and Harry Millwater. Quaternion and octonion-based finite element analysis methods for computing multiple first order derivatives. Journal of Computational Physics, 397(??):Article 108831, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305157.

Alldredge:2015:RPD

[AS15] Graham Alldredge and Florian Schneider. A realizabilitypreserving discontinuous Galerkin scheme for entropy-based moment closures for linear kinetic equations in one space dimension. Journal of Computational Physics, 295(??):665–684, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002910.

Ambrus:2016:LBM

[AS16] Victor E. Ambrus and Victor Sofonea. Lattice Boltzmann models based on half-range Gauss–Hermite quadratures. Journal of Computational Physics, 316(??):760–788, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300420.

Ahlkrona:2017:MAN

[AS17] Josefin Ahlkrona and Victor Shcherbakov. A meshfree approach to non-Newtonian free surface ice flow: Application to the Haut Glacier d’Arolla. Journal of Computational Physics, 330(??):633–649, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305496.

Ahmadian:2015:TMN

[ASB^{+}15] A. Ahmadian, S. Salahshour, D. Baleanu, H. Amirkhani, and R. Yunus. Tau method for the numerical solution of a fuzzy fractional kinetic model and its application to the oil palm frond as a promising source of xylose. Journal of Computational Physics, 294(??):562–584, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001473.

Almanasreh:2013:SFE

[ASS13] Hasan Almanasreh, Sten Salomonson, and Nils Svanstedt. Stabilized finite element method for the radial Dirac equation. Journal of Computational Physics, 236(??):426–442, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007000. See corrigendum

[ASS17].

Almanasreh:2017:CSF

[ASS17] Hasan Almanasreh, Sten Salomonson, and Nils Svanstedt. Corrigendum to “Stabilized finite element method for the radial Dirac equation” [J. Comput. Phys. 236 (2013) 426– 442]. Journal of Computational Physics, 340(??):712, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301432.

Amanbek:2019:ANH

[ASWvD19] Yerlan Amanbek, Gurpreet Singh, Mary F. Wheeler, and Hans van Duijn. Adaptive numerical homogenization for upscaling single phase flow and transport. Journal of Computational Physics, 387(??):117–133, June 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301317.

Anumolu:2018:GAL

[AT18] Lakshman Anumolu and Mario F. Trujillo. Gradient augmented level set method for phase change simulations. Journal of Computational Physics, 353(??):377–406, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307696.

Atangana:2015:SCT

[Ata15] Abdon Atangana. On the stability and convergence of the time-fractional variable order telegraph equation. Journal of Computational Physics, 293(??):104–114, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008602.

Arabi:2017:SER

[ATC17] Sina Arabi, Jean-Yves Trépanier, and Ricardo Camarero. A simple extension of Roe’s scheme for real gases. Journal of Computational Physics, 329(??):16–28, January 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305757.

Arabi:2019:SER

[ATC19] Sina Arabi, Jean-Yves Trépanier, and Ricardo Camarero. A simple extension of Roe’s scheme for multi-component real gas flows. Journal of Computational Physics, 388(??):178–194, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301901.

Amsallem:2016:RTS

[ATF16] David Amsallem, Radek Tezaur, and Charbel Farhat. Realtime solution of linear computational problems using databases of parametric reduced-order models with arbitrary underlying meshes. Journal of Computational Physics, 326(??):373–397, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303758.

Abraham:2018:SRS

[ATM^{+}18] Simon Abraham, Panagiotis Tsirikoglou, João Miranda, Chris Lacor, Francesco Contino, and Ghader Ghorbaniasl. Spectral representation of stochastic field data using sparse polynomial chaos expansions. Journal of Computational Physics, 367(??):109–120, August 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302419.

Antoine:2016:GSD

[ATZ16] Xavier Antoine, Qinglin Tang, and Yong Zhang. On the ground states and dynamics of space fractional nonlinear Schrödinger/Gross–Pitaevskii equations with rotation term and nonlocal nonlinear interactions. Journal of Computational Physics, 325(??):74–97, November 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303527.

Abushaikha:2017:FIM

[AVT17] Ahmad S. Abushaikha, Denis V. Voskov, and Hamdi A. Tchelepi. Fully implicit mixed-hybrid finite-element discretization for general purpose subsurface reservoir simulation. Journal of Computational Physics, 346(??):514–538, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304874.

Alhubail:2016:SRB

[AW16] Maitham Alhubail and Qiqi Wang. The swept rule for breaking the latency barrier in time advancing PDEs. Journal of Computational Physics, 307(??):110–121, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007664.

Alhawwary:2018:FAE

[AW18] Mohammad Alhawwary and Z. J. Wang. Fourier analysis and evaluation of DG, FD and compact difference methods for conservation laws. Journal of Computational Physics, 373(??):835–862, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304790.

Asthana:2017:CRC

[AWJ17] Kartikey Asthana, Jerry Watkins, and Antony Jameson. On consistency and rate of convergence of flux reconstruction for time-dependent problems. Journal of Computational Physics, 334(??):367–391, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300189.

Alekseeva:2016:HAR

[AWS16] Uliana Alekseeva, Roland G. Winkler, and Godehard Sutmann. Hydrodynamics in adaptive resolution particle simulations: Multiparticle collision dynamics. Journal of Computational Physics, 314(??):14–34, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001376.

Allen:2016:DNH

[AZ16] T. Allen and M. Zerroukat. A deep non-hydrostatic compressible atmospheric model on a Yin–Yang grid. Journal of Computational Physics, 319(??):44–60, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301590.

Adkins:2017:GCD

[AZ17] Melissa R. Adkins and Y. C. Zhou. Geodesic curvature driven surface microdomain formation. Journal of Computational Physics, 345(??):260–274, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304047.

Allen:2019:SLS

[AZ19a] T. Allen and M. Zerroukat. A semi-Lagrangian semi-implicit immersed boundary method for atmospheric flow over complex terrain. Journal of Computational Physics, 397(??):Article 108857, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305418.

Atkinson:2019:SBG

[AZ19b] Steven Atkinson and Nicholas Zabaras. Structured Bayesian Gaussian process latent variable model: Applications to datadriven dimensionality reduction and high-dimensional inversion. Journal of Computational Physics, 383(??):166–195, April 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300397.

Azis:2019:NSH

[Azi19] Moh. Ivan Azis. Numerical solutions for the Helmholtz boundary value problems of anisotropic homogeneous media. Journal of Computational Physics, 381(??):42–51, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300178.

Aldegunde:2016:QUF

[AZK16] Manuel Aldegunde, Nicholas Zabaras, and Jesper Kristensen. Quantifying uncertainties in first-principles alloy thermodynamics using cluster expansions. Journal of Computational Physics, 323(??):17–44, October 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302972.

Bhrawy:2015:FSC

[BA15] A. H. Bhrawy and M. A. Abdelkawy. A fully spectral collocation approximation for multi-dimensional fractional Schrödinger equations. Journal of Computational Physics, 294(??):462–483, August 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115002247.

Bao:2016:HON

[BABD16] Feng Bao, Rick Archibald, Dipanshu Bansal, and Olivier Delaire. Hierarchical optimization for neutron scattering problems. Journal of Computational Physics, 315(??):39–51, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001698.

Belme:2019:PAG

[BAD19] A. Belme, F. Alauzet, and A. Dervieux. An a priori anisotropic goal-oriented error estimate for viscous compressible flow and application to mesh adaptation. Journal of Computational Physics, 376(??):1051–1088, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305801.

Balsara:2016:HOR

[BAGK16] Dinshaw S. Balsara, Takanobu Amano, Sudip Garain, and Jinho Kim. A high-order relativistic two-fluid electrodynamic scheme with consistent reconstruction of electromagnetic fields and a multidimensional Riemann solver for electromagnetism. Journal of Computational Physics, 318(??):169–200, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301334.

Boje:2019:HPN

[BAK19] Astrid Boje, Jethro Akroyd, and Markus Kraft. A hybrid particle-number and particle model for efficient solution of population balance equations. Journal of Computational Physics, 389(??):189–218, July 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302165.

Balsara:2015:TDH

[Bal15] Dinshaw S. Balsara. Three dimensional HLL Riemann solver for conservation laws on structured meshes; application to Euler and magnetohydrodynamic flows. Journal of Computational Physics, 295(??):1–23, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500217X.

Bogner:2015:BCF

[BAR15] Simon Bogner, Regina Ammer, and Ulrich Rüde. Boundary conditions for free interfaces with the lattice Boltzmann method. Journal of Computational Physics, 297(??): 1–12, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115003125.

Barnes:2018:CDP

[Bar18] D. C. Barnes. Continuously differentiable PIC shape functions for triangular meshes. Journal of Computational Physics, 362(??):243–263, June 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300780.

Barton:2019:ICG

[Bar19] Philip T. Barton. An interface-capturing Godunov method for the simulation of compressible solid-fluid problems. Journal of Computational Physics, 390(??):25–50, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930227X.

Batty:2017:CCF

[Bat17] Christopher Batty. A cell-centred finite volume method for the Poisson problem on non-graded quadtrees with second order accurate gradients. Journal of Computational Physics, 331(??):49–72, February 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306271.

Bilger:2017:ETP

[BAVC17] C. Bilger, M. Aboukhedr, K. Vogiatzaki, and R. S. Cant. Evaluation of two-phase flow solvers using Level Set and Volume of Fluid methods. Journal of Computational Physics, 345(??):665–686, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304278.

Buckinx:2015:MSM

[BB15] Geert Buckinx and Martine Baelmans. Multi-scale modelling of flow in periodic solid structures through spatial averaging. Journal of Computational Physics, 291(??):34–51, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001448.

Balogh:2017:CAM

[BB17] Peter Balogh and Prosenjit Bagchi. A computational approach to modeling cellular-scale blood flow in complex geometry. Journal of Computational Physics, 334(??):280–307, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300177.

Boscheri:2019:HOD

[BB19] Walter Boscheri and Dinshaw S. Balsara. High order direct Arbitrary-Lagrangian–Eulerian (ALE) P_{N}P_{M} schemes with WENO Adaptive-Order reconstruction on unstructured meshes. Journal of Computational Physics, 398 (??):Article 108899, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305972.

Biagioni:2015:RID

[BBB15] David J. Biagioni, Daniel Beylkin, and Gregory Beylkin. Randomized interpolative decomposition of separated representations. Journal of Computational Physics, 281(??):116–134, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006913.

Benedetto:2016:HMV

[BBB^{+}16] Matías Fernando Benedetto, Stefano Berrone, Andrea Borio, Sandra Pieraccini, and Stefano Scialò. A hybrid mortar virtual element method for discrete fracture network simulations. Journal of Computational Physics, 306(??):148–166, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007743.

Brock:2015:EIG

[BBBG15] Benjamin Brock, Andrew Belt, Jay Jay Billings, and Mike Guidry. Explicit integration with GPU acceleration for large kinetic networks. Journal of Computational Physics, 302(??):591–602, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006063.

Barucq:2017:STD

[BBF^{+}17] H. Barucq, A. Bendali, M. Fares, V. Mattesi, and S. Tordeux. A symmetric Trefftz–DG formulation based on a local boundary element method for the solution of the Helmholtz equation. Journal of Computational Physics, 330(??):1069–1092, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304831.

Brehm:2019:DIB

[BBK19] C. Brehm, M. F. Barad, and C. C. Kiris. Development of immersed boundary computational aeroacoustic prediction capabilities for open-rotor noise. Journal of Computational Physics, 388(??):690–716, July 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301226.

Bandaru:2016:HFD

[BBKS16] Vinodh Bandaru, Thomas Boeck, Dmitry Krasnov, and Jörg Schumacher. A hybrid finite difference-boundary element procedure for the simulation of turbulent MHD duct flow at finite magnetic Reynolds number. Journal of Computational Physics, 304(??):320–339, January 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006695.

Blanc:2018:VRM

[BBKS18] X. Blanc, C. Bordin, G. Kluth, and G. Samba. Variance reduction method for particle transport equation in spherical geometry. Journal of Computational Physics, 364(??):274–297, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300974.

Bartels:2018:MST

[BBMN18] Sören Bartels, Andrea Bonito, Anastasia H. Muliana, and Ricardo H. Nochetto. Modeling and simulation of thermally actuated bilayer plates. Journal of Computational Physics, 354(??):512–528, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117308112.

Berggren:2018:ABL

[BBN18] Martin Berggren, Anders Bernland, and Daniel Noreland. Acoustic boundary layers as boundary conditions. Journal of Computational Physics, 371(??):633–650, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303838.

Bhardwaj:2016:PPI

[BBW16] U. Bhardwaj, S. Bukkuru, and M. Warrier. Post-processing interstitialcy diffusion from molecular dynamics simulations. Journal of Computational Physics, 305(??):263–275, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500710X.

Bertoglio:2016:SRB

[BC16a] Cristóbal Bertoglio and Alfonso Caiazzo. A Stokes-residual backflow stabilization method applied to physiological flows. Journal of Computational Physics, 313(??):260–278, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001157.

Bierig:2016:APD

[BC16b] Claudio Bierig and Alexey Chernov. Approximation of probability density functions by the Multilevel Monte Carlo Maximum Entropy method. Journal of Computational Physics, 314(??):661–681, June 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001790.

Bruno:2016:HOT

[BC16c] Oscar P. Bruno and Max Cubillos. Higher-order in time “quasi-unconditionally stable” ADI solvers for the compressible Navier–Stokes equations in 2D and 3D curvilinear domains. Journal of Computational Physics, 307(??):476–495, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008281.

Buchak:2016:STD

[BC16d] Peter Buchak and Darren G. Crowdy. Surface-tension-driven Stokes flow: a numerical method based on conformal geometry. Journal of Computational Physics, 317(??):347–361, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300900.

Bajars:2018:BIM

[BC18a] Janis Bajars and David J. Chappell. A boundary integral method for modelling vibroacoustic energy distributions in uncertain built up structures. Journal of Computational Physics, 373(??):130–147, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304455.

Bauer:2018:EEC

[BC18b] W. Bauer and C. J. Cotter. Energy–enstrophy conserving compatible finite element schemes for the rotating shallow water equations with slip boundary conditions. Journal of Computational Physics, 373(??):171–187, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304509.

Bertin:2018:FBF

[BC18c] N. Bertin and L. Capolungo. A FFT-based formulation for discrete dislocation dynamics in heterogeneous media. Journal of Computational Physics, 355(??):366–384, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308586.

Benoit:2015:TDC

[BCB15] J. Benoit, C. Chauvière, and P. Bonnet. Time-dependent current source identification for numerical simulations of Maxwell’s equations. Journal of Computational Physics, 289 (??):116–128, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115001035.

Botti:2017:MAB

[BCB17] L. Botti, A. Colombo, and F. Bassi. h-multigrid agglomeration based solution strategies for discontinuous Galerkin discretizations of incompressible flow problems. Journal of Computational Physics, 347(??):382–415, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305041.

Bonilla:2018:CNS

[BCC^{+}18] Luis L. Bonilla, Ana Carpio, Manuel Carretero, Gema Duro, Mihaela Negreanu, and Filippo Terragni. A convergent numerical scheme for integrodifferential kinetic models of angiogenesis. Journal of Computational Physics, 375(??):1270–1294, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306028.

Balin:2015:BEF

[BCD^{+}15] Nolwenn Balin, Fabien Casenave, François Dubois, Eric Duceau, Stefan Duprey, and Isabelle Terrasse. Boundary element and finite element coupling for aeroacoustics simulations. Journal of Computational Physics, 294(??):274–296, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001990.

Buchan:2015:PRO

[BCG^{+}15] A. G. Buchan, A. A. Calloo, M. G. Goffin, S. Dargaville, F. Fang, C. C. Pain, and I. M. Navon. A POD reduced order model for resolving angular direction in neutron/photon transport problems. Journal of Computational Physics, 296(??):138–157, September 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003009.

Bruno:2019:HOI

[BCJ19] Oscar P. Bruno, Max Cubillos, and Edwin Jimenez. Higherorder implicit-explicit multi-domain compressible Navier– Stokes solvers. Journal of Computational Physics, 391(??): 322–346, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119301500.

Bokil:2017:ESD

[BCJL17] Vrushali A. Bokil, Yingda Cheng, Yan Jiang, and Fengyan Li. Energy stable discontinuous Galerkin methods for Maxwell’s equations in nonlinear optical media. Journal of Computational Physics, 350(??):420–452, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305806.

Blanes:2015:EAB

[BCM15a] Sergio Blanes, Fernando Casas, and Ander Murua. An efficient algorithm based on splitting for the time integration of the Schrödinger equation. Journal of Computational Physics, 303(??):396–412, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006555.

Bukac:2015:PSF

[BCM15b] M. Bukac, S. Canić, and B. Muha. A partitioned scheme for fluid-composite structure interaction problems. Journal of Computational Physics, 281(??):493–517, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400727X.

Borcea:2019:DAI

[BCM19] Liliana Borcea, Fioralba Cakoni, and Shixu Meng. A direct approach to imaging in a waveguide with perturbed geometry.

Journal of Computational Physics, 392(??):556–577, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303262.

Bandara:2015:BEB

[BCO^{+}15] Kosala Bandara, Fehmi Cirak, Günther Of, Olaf Steinbach, and Jan Zapletal. Boundary element based multiresolution shape optimisation in electrostatics. Journal of Computational Physics, 297(??):584–598, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003459.

Bugeat:2019:GOF

[BCRS19] B. Bugeat, J. C. Chassaing, J. C. Robinet, and P. Sagaut. 3D global optimal forcing and response of the supersonic boundary layer. Journal of Computational Physics, 398 (??):Article 108888, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305868.

Bendall:2019:RSA

[BCS19] Thomas M. Bendall, Colin J. Cotter, and Jemma Shipton. The ‘recovered space’ advection scheme for lowestorder compatible finite element methods. Journal of Computational Physics, 390(??):342–358, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302463.

Babaee:2017:RBO

[BCSK17] Hessam Babaee, Minseok Choi, Themistoklis P. Sapsis, and George Em Karniadakis. A robust bi-orthogonal/dynamically-orthogonal method using the covariance pseudo-inverse with application to stochastic flow problems. Journal of Computational Physics, 344(??):303–319, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303364.

Bajars:2017:TPS

[BCST17] Janis Bajars, David J. Chappell, Niels Søndergaard, and Gregor Tanner. Transport of phase space densities through tetrahedral meshes using discrete flow mapping. Journal of Computational Physics, 328(??):95–108, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305174.

Balsara:2015:DFM

[BD15a] Dinshaw S. Balsara and Michael Dumbser. Divergence-free MHD on unstructured meshes using high order finite volume schemes based on multidimensional Riemann solvers. Journal of Computational Physics, 299(??):687–715, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500457X.

Balsara:2015:MRP

[BD15b] Dinshaw S. Balsara and Michael Dumbser. Multidimensional Riemann problem with self-similar internal structure. Part II. Application to hyperbolic conservation laws on unstructured meshes. Journal of Computational Physics, 287 (??):269–292, 2015. CODEN JCTPAH. ISSN 0021-9991.

URL http://www.sciencedirect.com/science/article/
pii/S0021999114007566.

Brethes:2016:ANO

[BD16] Gautier Brèthes and Alain Dervieux. Anisotropic normoriented mesh adaptation for a Poisson problem. Journal of Computational Physics, 322(??):804–826, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630287X.

Boscheri:2017:ALE

[BD17] Walter Boscheri and Michael Dumbser. Arbitrary-Lagrangian– Eulerian discontinuous Galerkin schemes with a posteriori subcell finite volume limiting on moving unstructured meshes. Journal of Computational Physics, 346(??):449–479, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304679.

Botti:2018:AHH

[BD18] Lorenzo Botti and Daniele A. Di Pietro. Assessment of hybrid high-order methods on curved meshes and comparison with discontinuous Galerkin methods. Journal of Computational Physics, 370(??):58–84, October 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303176.

Bartholomew:2018:UFM

[BDAA^{+}18] Paul Bartholomew, Fabian Denner, Mohd Hazmil Abdol-Azis, Andrew Marquis, and Berend G. M. van Wachem. Unified formulation of the momentum-weighted interpolation for collocated variable arrangements. Journal of Computational Physics, 375(??):177–208, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305539.

Blommaert:2017:PGO

[BDB^{+}17] Maarten Blommaert, Wouter Dekeyser, Martine Baelmans, Nicolas R. Gauger, and Detlev Reiter. A practical globalization of one-shot optimization for optimal design of tokamak divertors. Journal of Computational Physics, 328(??):399–412, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305393.

Basic:2018:CRM

[BDB18] Josip Basic, Nastia Degiuli, and Dario Ban. A class of renormalised meshless Laplacians for boundary value problems. Journal of Computational Physics, 354(??):269–287, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308306.

Bhrawy:2015:STA

[BDBEE15] A. H. Bhrawy, E. H. Doha, D. Baleanu, and S. S. Ezz-Eldien. A spectral tau algorithm based on Jacobi operational matrix for numerical solution of time fractional diffusion-wave equations. Journal of Computational Physics, 293(??):142–156, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002265.

Bao:2017:IBM

[BDG^{+}17] Yuanxun Bao, Aleksandar Donev, Boyce E. Griffith, David M. McQueen, and Charles S. Peskin. An immersed boundary method with divergence-free velocity interpolation and force spreading. Journal of Computational Physics, 347 (??):183–206, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304953.

Bruel:2019:LMC

[BDJP19] Pascal Bruel, Simon Delmas, Jonathan Jung, and Vincent Perrier. A low Mach correction able to deal with low Mach acoustics. Journal of Computational Physics, 378(??):723–759, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307460.

Belonosov:2017:ISH

[BDK^{+}17] Mikhail Belonosov, Maxim Dmitriev, Victor Kostin, Dmitry Neklyudov, and Vladimir Tcheverda. An iterative solver for the 3D Helmholtz equation. Journal of Computational Physics, 345(??):330–344, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304011.

Benner:2017:FIS

[BDKK17] Peter Benner, Sergey Dolgov, Venera Khoromskaia, and Boris N. Khoromskij. Fast iterative solution of the Bethe– Salpeter eigenvalue problem using low-rank and QTT tensor approximation. Journal of Computational Physics, 334 (??):221–239, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116307112.

Boscheri:2018:SOC

[BDLM18] Walter Boscheri, Michael Dumbser, Raphaël Loubère, and Pierre-Henri Maire. A second-order cell-centered Lagrangian ADER-MOOD finite volume scheme on multidimensional unstructured meshes for hydrodynamics. Journal of Computational Physics, 358(??):103–129, April 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309348.

Besse:2017:AMS

[BDM17] N. Besse, E. Deriaz, and É. Madaule. Adaptive multiresolution semi-Lagrangian discontinuous Galerkin methods for the Vlasov equations. Journal of Computational Physics, 332(??):376–417, March 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306441.

Bandopadhyay:2015:CSP

[BDMC15] Aditya Bandopadhyay, Debabrata DasGupta, Sushanta K. Mitra, and Suman Chakraborty. Computation of streaming potential in porous media: Modified permeability tensor. Journal of Computational Physics, 300(??):53–69, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004751.

Borcea:2019:RNP

[BDMZ19] Liliana Borcea, Vladimir Druskin, Alexander V. Mamonov, and Mikhail Zaslavsky. Robust nonlinear processing of active array data in inverse scattering via truncated reduced order models. Journal of Computational Physics, 381(??):1–26, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300026.

Bretin:2018:MMC

[BDPM18] Elie Bretin, Alexandre Danescu, José Penuelas, and Simon Masnou. Multiphase mean curvature flows with high mobility contrasts: A phase-field approach, with applications to nanowires. Journal of Computational Physics, 366(??):324– 349, July 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301426.

Brown-Dymkoski:2017:AAW

[BDV17] Eric Brown-Dymkoski and Oleg V. Vasilyev. Adaptive– Anisotropic Wavelet Collocation Method on general curvilinear coordinate systems. Journal of Computational Physics, 333(??):414–426, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116307021.

Boscheri:2015:HOC

[BDZ15] Walter Boscheri, Michael Dumbser, and Olindo Zanotti. High order cell-centered Lagrangian-type finite volume schemes with time-accurate local time stepping on unstructured triangular meshes. Journal of Computational Physics, 291(??):120–150, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500145X.

Beghin:2015:FTS

[Beg15] Luisa Beghin. On fractional tempered stable processes and their governing differential equations. Journal of Computational Physics, 293(??):29–39, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003726.

Banerjee:2015:SSK

[BEJ15] Amartya S. Banerjee, Ryan S. Elliott, and Richard D. James. A spectral scheme for Kohn–Sham density functional theory of clusters. Journal of Computational Physics, 287 (??):226–253, 2015. CODEN JCTPAH. ISSN 0021-9991.

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000704.

Bayona:2019:RPR

[BFF19] Víctor Bayona, Natasha Flyer, and Bengt Fornberg. On the role of polynomials in RBF–FD approximations: III. Behavior near domain boundaries. Journal of Computational Physics, 380(??):378–399, March 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830809X.

Bayona:2017:RPR

[BFFB17] Victor Bayona, Natasha Flyer, Bengt Fornberg, and Gregory A. Barnett. On the role of polynomials in RBF-FD approximations: II. Numerical solution of elliptic PDEs. Journal of Computational Physics, 332(??):257–273, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306490.

Ballarin:2016:FSP

[BFI^{+}16] Francesco Ballarin, Elena Faggiano, Sonia Ippolito, Andrea Manzoni, Alfio Quarteroni, Gianluigi Rozza, and Roberto Scrofani. Fast simulations of patient-specific haemodynamics of coronary artery bypass grafts based on a POD-Galerkin method and a vascular shape parametrization. Journal of Computational Physics, 315(??):609–628, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300304.

Bergmann:2018:ZGF

[BFI^{+}18] Michel Bergmann, Andrea Ferrero, Angelo Iollo, Edoardo Lombardi, Angela Scardigli, and Haysam Telib. A zonal Galerkin-free POD model for incompressible flows. Journal of Computational Physics, 352(??):301–325, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307386.

Beck:2019:DNN

[BFM19] Andrea Beck, David Flad, and Claus-Dieter Munz. Deep neural networks for data-driven LES closure models. Journal of Computational Physics, 398(??):Article 108910, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119306151.

Bonaventura:2018:MSW

[BFNGDNR18] Luca Bonaventura, Enrique D. Fernández-Nieto, José Garres-Díaz, and Gladys Narbona-Reina. Multilayer shallow water models with locally variable number of layers and semi-implicit time discretization. Journal of Computational Physics, 364(??):209–234, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301694.

Barucq:2018:LSO

[BFP18] Hélène Barucq, Florian Faucher, and Ha Pham. Localization of small obstacles from back-scattered data at limited incident angles with full-waveform inversion. Journal of Computational Physics, 370(??):1–24, October 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303036.

Borker:2017:HOD

[BFT17] Raunak Borker, Charbel Farhat, and Radek Tezaur. A high-order discontinuous Galerkin method for unsteady advection-diffusion problems. Journal of Computational Physics, 332(??):520–537, March 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911630674X.

Busto:2018:PHH

[BFTVC18] S. Busto, J. L. Ferrín, E. F. Toro, and M. E. Vázquez-Cendón. A projection hybrid high order finite volume/finite element method for incompressible turbulent flows. Journal of Computational Physics, 353(??):169–192, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307428.

Brambley:2016:TDI

[BG16a] E. J. Brambley and G. Gabard. Time-domain implementation of an impedance boundary condition with boundary layer correction. Journal of Computational Physics, 321(??):755–775, September 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911630225X.

Bueno:2016:LVT

[BG16b] Jesus Bueno and Hector Gomez. Liquid-vapor transformations with surfactants. phase-field model and isogeometric analysis. Journal of Computational Physics, 321(??):797–818, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116302339.

Baffet:2019:WSS

[BG19a] Daniel Baffet and Marcus J. Grote. On wave splitting, source separation and echo removal with absorbing boundary conditions. Journal of Computational Physics, 387(??):589–596, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301731.

Beriot:2019:AAF

[BG19b] Hadrien Bériot and Gwénaël Gabard. Anisotropic adaptivity of the p-FEM for time-harmonic acoustic wave propagation. Journal of Computational Physics, 378(??):234–256, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307319.

Bochkov:2019:SPT

[BG19c] Daniil Bochkov and Frederic Gibou. Solving Poisson-type equations with Robin boundary conditions on piecewise smooth interfaces. Journal of Computational Physics, 376 (??):1156–1198, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118306831.

Bizzozero:2019:MCS

[BGD19] D. A. Bizzozero, E. Gjonaj, and H. De Gersem. Modeling coherent synchrotron radiation with a discontinuous Galerkin time-domain method. Journal of Computational Physics, 394(??):745–758, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303602.

Bohn:2016:SGB

[BGG16] Bastian Bohn, Jochen Garcke, and Michael Griebel. A sparse grid based method for generative dimensionality reduction of high-dimensional data. Journal of Computational Physics, 309(??):1–17, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115008529.

Bokil:2015:DRM

[BGGM15] V. A. Bokil, N. L. Gibson, V. Gyrya, and D. A. McGregor. Dispersion reducing methods for edge discretizations of the electric vector wave equation. Journal of Computational Physics, 287(??):88–109, 2015. CODEN JCTPAH. ISSN

0021-9991. URL http://www.sciencedirect.com/science/
article/pii/S0021999115000728.

Borggaard:2019:GAP

[BGHK19] Jeff Borggaard, Nathan Glatt-Holtz, and Justin Krometis. GPU-accelerated particle methods for evaluation of sparse observations for inverse problems constrained by diffusion PDEs. Journal of Computational Physics, 391(??):142–154, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302803.

Bollada:2015:TDT

[BGJ^{+}15] P. C. Bollada, C. E. Goodyer, P. K. Jimack, A. M. Mullis, and F. W. Yang. Three dimensional thermal-solute phase field simulation of binary alloy solidification. Journal of Computational Physics, 287(??):130–150, 2015. CODEN JCTPAH. ISSN

0021-9991. URL http://www.sciencedirect.com/science/
article/pii/S0021999115000546.

Beskos:2017:GMI

[BGL^{+}17] Alexandros Beskos, Mark Girolami, Shiwei Lan, Patrick E. Farrell, and Andrew M. Stuart. Geometric MCMC for infinite-dimensional inverse problems. Journal of Computational Physics, 335(??):327–351, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116307033.

Brenowitz:2016:NLS

[BGM16] N. D. Brenowitz, D. Giannakis, and A. J. Majda. Nonlinear Laplacian spectral analysis of Rayleigh–Bénard convection. Journal of Computational Physics, 315(??):536–553, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300109.

Barrett:2015:SFE

[BGN15] John W. Barrett, Harald Garcke, and Robert Nürnberg. Stable finite element approximations of two-phase flow with soluble surfactant. Journal of Computational Physics, 297(??):530–564, September 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003630.

Barrett:2019:FEM

[BGN19] John W. Barrett, Harald Garcke, and Robert Nürnberg. Finite element methods for fourth order axisymmetric geometric evolution equations. Journal of Computational Physics, 376(??):733–766, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306636.

Bernier:2019:SPE

[BGRC19] Caroline Bernier, Mattia Gazzola, Renaud Ronsse, and Philippe Chatelain. Simulations of propelling and energy harvesting articulated bodies via vortex particle-mesh methods. Journal of Computational Physics, 392(??):34–55, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302827.

Balsara:2016:ECW

[BGS16] Dinshaw S. Balsara, Sudip Garain, and Chi-Wang Shu. An efficient class of WENO schemes with adaptive order. Journal of Computational Physics, 326(??):780–804, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304211.

Balsara:2018:CEM

[BGTM18] Dinshaw S. Balsara, Sudip Garain, Allen Taflove, and Gino Montecinos. Computational electrodynamics in material media with constraint-preservation, multidimensional Riemann solvers and sub-cell resolution — part II, higher order FVTD schemes. Journal of Computational Physics, 354(??):613–645, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307519.

Beckstein:2017:ESE

[BGV17] Pascal Beckstein, Vladimir Galindo, and Vuko Vukcević. Efficient solution of 3D electromagnetic eddy-current problems within the finite volume framework of OpenFOAM. Journal of Computational Physics, 344(??):623–646, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303741.

Banks:2016:GDM

[BH16a] J. W. Banks and T. Hagstrom. On Galerkin difference methods. Journal of Computational Physics, 313(??):310–327, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001121.

Berry:2016:SMC

[BH16b] Tyrus Berry and John Harlim. Semiparametric modeling: Correcting low-dimensional model error in parametric models. Journal of Computational Physics, 308(??):305–321, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008621.

Bilbao:2018:HOA

[BH18] Stefan Bilbao and Brian Hamilton. Higher-order accurate twostep finite difference schemes for the many-dimensional wave equation. Journal of Computational Physics, 367(??):134–165, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302262.

Beaudoin:2018:APM

[BHdD18] Anthony Beaudoin, Serge Huberson, and Jean-Raynald de Dreuzy. Adapting particle methods to model the dynamics of concentration gradients and chemical reactivity under advective diffusive transport conditions. Journal of Computational Physics, 354(??):196–210, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308045.

Bolding:2017:SOD

[BHE^{+}17] Simon Bolding, Joshua Hansel, Jarrod D. Edwards, Jim E. Morel, and Robert B. Lowrie. Second-order discretization in space and time for radiation-hydrodynamics. Journal of Computational Physics, 338(??):511–526, June 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301663.

Brehm:2015:LSI

[BHF15] C. Brehm, C. Hader, and H. F. Fasel. A locally stabilized immersed boundary method for the compressible Navier–Stokes equations. Journal of Computational Physics, 295(??):475–504, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002697.

Browne:2019:ELW

[BHFB19] Oliver M. F. Browne, Anthony P. Haas, Herman F. Fasel, and Christoph Brehm. An efficient linear wavepacket tracking method for hypersonic boundary-layer stability prediction. Journal of Computational Physics, 380(??):243–268, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307721.

Bonev:2018:DGS

[BHGK18] Boris Bonev, Jan S. Hesthaven, Francis X. Giraldo, and Michal A. Kopera. Discontinuous Galerkin scheme for the spherical shallow water equations with applications to tsunami modeling and prediction. Journal of Computational Physics, 362(??):425–448, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300846.

Butler:2015:QUM

[BHJ15] T. Butler, A. Huhtala, and M. Juntunen. Quantifying uncertainty in material damage from vibrational data. Journal of Computational Physics, 283(??):414–435, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008195.

Banks:2018:GDA

[BHJ18] J. W. Banks, T. Hagstrom, and J. Jacangelo. Galerkin differences for acoustic and elastic wave equations in two space dimensions. Journal of Computational Physics, 372(??):864–892, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304078.

Banks:2016:AMP

[BHKS16] J. W. Banks, W. D. Henshaw, A. K. Kapila, and D. W. Schwendeman. An added-mass partition algorithm for fluid-structure interactions of compressible fluids and nonlinear solids. Journal of Computational Physics, 305(??):1037–1064, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007202.

Bao:2015:RTA

[BHL15] Gang Bao, Guanghui Hu, and Di Liu. Real-time adaptive finite element solution of time-dependent Kohn–Sham equation. Journal of Computational Physics, 281(??):743–758, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007402.

Brenner:2018:HDM

[BHMS18] Konstantin Brenner, Julian Hennicker, Roland Masson, and Pierre Samier. Hybrid-dimensional modelling of two-phase flow through fractured porous media with enhanced matrix fracture transmission conditions. Journal of Computational Physics, 357(??):100–124, March 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308781.

Blum:2019:ADT

[BHNS19] Jacques Blum, Holger Heumann, Eric Nardon, and Xiao Song. Automating the design of tokamak experiment scenarios. Journal of Computational Physics, 394(??):594–614, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303985.

Betcke:2019:ABE

[BHP19] Timo Betcke, Alexander Haberl, and Dirk Praetorius. Adaptive boundary element methods for the computation of the electrostatic capacity on complex polyhedra. Journal of Computational Physics, 397(??):Article 108837, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305212.

Bhaduri:2018:SCA

[BHS^{+}18] Anindya Bhaduri, Yanyan He, Michael D. Shields, Lori Graham-Brady, and Robert M. Kirby. Stochastic collocation approach with adaptive mesh refinement for parametric uncertainty analysis. Journal of Computational Physics, 371(??):732–750, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303814.

Banks:2017:SPFa

[BHST17a] J. W. Banks, W. D. Henshaw, D. W. Schwendeman, and Qi Tang. A stable partitioned FSI algorithm for rigid bodies and incompressible flow. Part I: Model problem analysis. Journal of Computational Physics, 343(??):432–468, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300256.

Banks:2017:SPFb

[BHST17b] J. W. Banks, W. D. Henshaw, D. W. Schwendeman, and Qi Tang. A stable partitioned FSI algorithm for rigid bodies and incompressible flow. Part II: General formulation. Journal of Computational Physics, 343(??):469–500, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303431.

Banks:2018:SPF

[BHST18] J. W. Banks, W. D. Henshaw, D. W. Schwendeman, and Qi Tang. A stable partitioned FSI algorithm for rigid bodies and incompressible flow in three dimensions. Journal of Computational Physics, 373(??):455–492, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304510.

Bosma:2017:MFV

[BHTT17] Sebastian Bosma, Hadi Hajibeygi, Matei Tene, and Hamdi A. Tchelepi. Multiscale finite volume method for discrete fracture modeling on unstructured grids (MS-DFM). Journal of Computational Physics, 351(??):145–164, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306927.

Bajc:2016:MAS

[BHZ16] Iztok Bajc, Frédéric Hecht, and Slobodan Zumer. A mesh adaptivity scheme on the Landau-de Gennes functional minimization case in 3D, and its driving efficiency. Journal of Computational Physics, 321(??):981–996, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001443.

Bergmann:2016:BSS

[BI16] Michel Bergmann and Angelo Iollo. Bioinspired swimming simulations. Journal of Computational Physics, 323(??):310–321, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303175.

Bernard:2018:ROM

[BIR18] Florian Bernard, Angelo Iollo, and Sébastien Riffaud. Reduced-order model for the BGK equation based on POD and optimal transport. Journal of Computational Physics, 373(??):545–570, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304546.

Beig:2015:MIE

[BJ15] Shahaboddin Alahyari Beig and Eric Johnsen. Maintaining interface equilibrium conditions in compressible multiphase flows using interface capturing. Journal of Computational Physics, 302(??):548–566, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006117.

Bull:2016:EFE

[BJ16] Jonathan R. Bull and Antony Jameson. Explicit filtering and exact reconstruction of the sub-filter stresses in large eddy simulation. Journal of Computational Physics, 306(??):117–136, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007779.

Becache:2017:SPM

[BJK17] Eliane Bécache, Patrick Joly, and Maryna Kachanovska. Stable perfectly matched layers for a cold plasma in a strong background magnetic field. Journal of Computational Physics, 341(??):76–101, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302528.

Ba:2018:TSE

[BJO18] Yuming Ba, Lijian Jiang, and Na Ou. A two-stage ensemble Kalman filter based on multiscale model reduction for inverse problems in time fractional diffusion-wave equations. Journal of Computational Physics, 374(??):300–330, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305138.

Badri:2018:HPC

[BJRF18] M. A. Badri, P. Jolivet, B. Rousseau, and Y. Favennec. High performance computation of radiative transfer equation using the finite element method. Journal of Computational Physics, 360(??):74–92, May 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300378.

Bao:2015:CGS

[BJTZ15] Weizhu Bao, Shidong Jiang, Qinglin Tang, and Yong Zhang. Computing the ground state and dynamics of the nonlinear Schrödinger equation with nonlocal interactions via the nonuniform FFT. Journal of Computational Physics, 296(??):72–89, September 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003022.

Bao:2017:PFE

[BJWZ17] Weizhu Bao, Wei Jiang, Yan Wang, and Quan Zhao. A parametric finite element method for solid-state dewetting problems with anisotropic surface energies. Journal of Computational Physics, 330(??):380–400, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306015.

Balsara:2016:SRM

[BK16a] Dinshaw S. Balsara and Jinho Kim. A subluminal relativistic magnetohydrodynamics scheme with ADER–WENO predictor and multidimensional Riemann solver-based corrector. Journal of Computational Physics, 312(??):357–384, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000553.

Borsche:2016:HON

[BK16b] Raul Borsche and Jochen Kall. High order numerical methods for networks of hyperbolic conservation laws coupled with ODEs and lumped parameter models. Journal of Computational Physics, 327(??):678–699, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630496X.

Bakhvalov:2017:MFC

[BK17a] P. A. Bakhvalov and T. K. Kozubskaya. Modification of Flux Correction method for accuracy improvement on unsteady problems. Journal of Computational Physics, 338(??):199–216, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301560.

Balsara:2017:NSA

[BK17b] Dinshaw S. Balsara and Roger Käppeli. Von Neumann stability analysis of globally divergence-free RKDG schemes for the induction equation using multidimensional Riemann solvers. Journal of Computational Physics, 336(??):104–127, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300724.

Basting:2017:FFE

[BK17c] Melanie Basting and Dmitri Kuzmin. An FCT finite element scheme for ideal MHD equations in 1D and 2D. Journal of Computational Physics, 338(??):585–605, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301547.

Biala:2018:PAN

[BK18] T. A. Biala and A. Q. M. Khaliq. Parallel algorithms for nonlinear time-space fractional parabolic PDEs. Journal of Computational Physics, 375(??):135–154, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305631.

Balsara:2019:NSA

[BK19a] Dinshaw S. Balsara and Roger Käppeli. von Neumann stability analysis of globally constraint-preserving DGTD and PNPM schemes for the Maxwell equations using multidimensional Riemann solvers. Journal of Computational Physics, 376(??):1108–1137, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306533.

Bassett:2019:MLP

[BK19b] Brody Bassett and Brian Kiedrowski. Meshless local Petrov– Galerkin solution of the neutron transport equation with streamline-upwind Petrov–Galerkin stabilization. Journal of Computational Physics, 377(??):1–59, January 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306958.

Bukhvostova:2015:LMN

[BKG15] A. Bukhvostova, J. G. M. Kuerten, and B. J. Geurts. Low Mach number algorithm for droplet-laden turbulent channel flow including phase transition. Journal of Computational Physics, 295(??):420–437, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002673.

Bosler:2017:LPM

[BKKJ17] Peter A. Bosler, James Kent, Robert Krasny, and Christiane Jablonowski. A Lagrangian particle method with remeshing for tracer transport on the sphere. Journal of Computational Physics, 340(??):639–654, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730253X.

Burger:2016:DAV

[BKKRB16] Raimund Bürger, Sarvesh Kumar, Kenettinkara Sudarshan Kumar, and Ricardo Ruiz-Baier. Discontinuous approximation of viscous two-phase flow in heterogeneous porous media. Journal of Computational Physics, 321(??):126–150, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301887.

Benner:2019:CDS

[BKKY19] Peter Benner, Venera Khoromskaia, Boris N. Khoromskij, and Chao Yang. Computing the density of states for optical spectra of molecules by low-rank and QTT tensor approximation. Journal of Computational Physics, 382(??):221–239, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300282.

Burgel:2017:SRT

[BKL17] Florian Bürgel, Kamil S. Kazimierski, and Armin Lechleiter. A sparsity regularization and total variation based computational framework for the inverse medium problem in scattering. Journal of Computational Physics, 339(??):1–30, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301985.

Beams:2018:HOF

[BKO18] Natalie N. Beams, Andreas Klöckner, and Luke N. Olson. High-order finite element-integral equation coupling on embedded meshes. Journal of Computational Physics, 375(??):1295–1313, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305552.

Bao:2016:GLI

[BKP16] Yuanxun Bao, Jason Kaye, and Charles S. Peskin. A Gaussianlike immersed-boundary kernel with three continuous derivatives and improved translational invariance. Journal of Computational Physics, 316(??):139–144, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300663.

Bonart:2019:CES

[BKR19] Henning Bonart, Christian Kahle, and Jens-Uwe Repke. Comparison of energy stable simulation of moving contact line problems using a thermodynamically consistent Cahn– Hilliard Navier–Stokes model. Journal of Computational Physics, 399(??):Article 108959, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306643.

Burger:2015:DFV

[BKRB15] Raimund Bürger, Sarvesh Kumar, and Ricardo Ruiz-Baier. Discontinuous finite volume element discretization for coupled flow-transport problems arising in models of sedimentation. Journal of Computational Physics, 299(??):446–471, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004659.

Barlow:2018:COF

[BKS18] Andrew Barlow, Matej Klima, and Mikhail Shashkov. Constrained optimization framework for interface-aware sub-scale dynamics models for voids closure in Lagrangian hydrodynamics. Journal of Computational Physics, 371(??):914–944, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301943.

Bashardanesh:2018:EGF

[BL18] Zahedeh Bashardanesh and Per Lötstedt. Efficient Green’s Function Reaction Dynamics (GFRD) simulations for diffusionlimited, reversible reactions. Journal of Computational Physics, 357(??):78–99, March 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117309191.

Bardazzi:2015:GHM

[BLA^{+}15] A. Bardazzi, C. Lugni, M. Antuono, G. Graziani, and O. M. Faltinsen. Generalized HPC method for the Poisson equation. Journal of Computational Physics, 299(??):630–648, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004714.

Bouffard:2017:PCM

[BLC^{+}17] Mathieu Bouffard, Stéphane Labrosse, Gaël Choblet, Alexandre Fournier, Julien Aubert, and Paul J. Tackley. A particlein-cell method for studying double-diffusive convection in the liquid layers of planetary interiors. Journal of Computational Physics, 346(??):552–571, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304734.

Boscheri:2015:DAL

[BLD15] Walter Boscheri, Raphaël Loubère, and Michael Dumbser. Direct arbitrary-Lagrangian–Eulerian ADER–MOOD finite volume schemes for multidimensional hyperbolic conservation laws. Journal of Computational Physics, 292(??):56–87, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001515.

Blais:2016:DUC

[BLG^{+}16] Bruno Blais, Manon Lassaigne, Christoph Goniva, Louis Fradette, and François Bertrand. Development of an unresolved CFD-DEM model for the flow of viscous suspensions and its application to solid-liquid mixing. Journal of Computational Physics, 318(??):201–221, August 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301358.

Boedec:2017:IFB

[BLJ17] Gwenn Boedec, Marc Leonetti, and Marc Jaeger. Isogeometric FEM–BEM simulations of drop, capsule and vesicle dynamics in Stokes flow. Journal of Computational Physics, 342(??):117–138, August 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302954.

Bian:2015:MRF

[BLK15] Xin Bian, Zhen Li, and George Em Karniadakis. Multiresolution flow simulations by smoothed particle hydrodynamics via domain decomposition. Journal of Computational Physics, 297(??):132–155, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003010.

Blumers:2019:SPT

[BLK19] Ansel L. Blumers, Zhen Li, and George Em Karniadakis. Supervised parallel-in-time algorithm for long-time Lagrangian simulations of stochastic dynamics: Application to hydrodynamics. Journal of Computational Physics, 393(??):214–228, September 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303419.

Bao:2016:MMC

[BLL16] Gang Bao, Di Liu, and Songting Luo. Multiscale modeling and computation of optically manipulated nano devices. Journal of Computational Physics, 316(??):558–572, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300791.

Balachandar:2019:SIV

[BLL19] S. Balachandar, Kai Liu, and Mandar Lakhote. Self-induced velocity correction for improved drag estimation in Euler– Lagrange point-particle simulations. Journal of Computational Physics, 376(??):160–185, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306351. See corrigendum [BLL20].

Balachandar:2020:CSI

[BLL20] S. Balachandar, Kai Liu, and Mandar Lakhote. Corrigendum to “Self-induced velocity correction for improved drag estimation in Euler–Lagrange point-particle simulations” [j. comput. phys. 376 (2019) 160–185]. Journal of Computational Physics, 401(??):Article 108813, January 15, 2020. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119304978. See [BLL19].

Balsara:2018:ESO

[BLM18] Dinshaw S. Balsara, Jiequan Li, and Gino I. Montecinos. An efficient, second order accurate, universal generalized Riemann problem solver based on the HLLI Riemann solver. Journal of Computational Physics, 375(??):1238–1269, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306120.

Bispen:2017:API

[BLMY17] Georgij Bispen, Mária Lukácová-Medvid’ová, and Leonid Yelash. Asymptotic preserving IMEX finite volume schemes for low Mach number Euler equations with gravitation. Journal of Computational Physics, 335(??):222–248, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730030X.

Blonigan:2017:ASA

[Blo17] Patrick J. Blonigan. Adjoint sensitivity analysis of chaotic dynamical systems with non-intrusive least squares shadowing. Journal of Computational Physics, 348(??):803–826, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305739.

Binder:2015:GPR

[BLS15] Andrew Binder, Tony Lelièvre, and Gideon Simpson. A generalized parallel replica dynamics. Journal of Computational Physics, 284(??):595–616, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000030.

Birk:2016:CAR

[BLS16] C. Birk, L. Liu, and Ch. Song. Coupled acoustic response of two-dimensional bounded and unbounded domains using doubly-asymptotic open boundaries. Journal of Computational Physics, 310(??):252–284, April 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008487.

Bermudez:2016:NSN

[BLVC16] Alfredo Bermúdez, Xián López, and M. Elena Vázquez-Cendón. Numerical solution of non-isothermal non-adiabatic flow of real gases in pipelines. Journal of Computational Physics, 323(??):126–148, October 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303151.

Bermudez:2017:TNJ

[BLVC17] Alfredo Bermúdez, Xián López, and M. Elena Vázquez-Cendón. Treating network junctions in finite volume solution of transient gas flow models. Journal of Computational Physics, 344(??):187–209, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303546.

Balac:2015:ESS

[BM15] Stéphane Balac and Fabrice Mahé. An Embedded Split-Step method for solving the nonlinear Schrödinger equation in optics. Journal of Computational Physics, 280(??):295–305, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006548.

Bhattacharjee:2016:NMB

[BM16] Satyaki Bhattacharjee and Karel Matous. A nonlinear manifold-based reduced order model for multiscale analysis of heterogeneous hyperelastic materials. Journal of Computational Physics, 313(??):635–653, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001194.

Bhat:2019:CPR

[BM19a] Sourabh Bhat and J. C. Mandal. Contact preserving Riemann solver for incompressible two-phase flows. Journal of Computational Physics, 379(??):173–191, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830706X.

Bruno:2019:SES

[BM19b] Oscar P. Bruno and Martín Maas. Shifted equivalent sources and FFT acceleration for periodic scattering problems, including wood anomalies. Journal of Computational Physics, 378(??):548–572, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307113.

Budd:2018:SSO

[BMC18a] Chris J. Budd, Andrew T. T. McRae, and Colin J. Cotter. The scaling and skewness of optimally transported meshes on the sphere. Journal of Computational Physics, 375(??):540–564, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305515.

Burton:2018:CEC

[BMC^{+}18b] D. E. Burton, N. R. Morgan, M. R. J. Charest, M. A. Kenamond, and J. Fung. Compatible, energy conserving, bounds preserving remap of hydrodynamic fields for an extended ALE scheme. Journal of Computational Physics, 355(??):492–533, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308550.

Burton:2015:RDL

[BMCK15] D. E. Burton, N. R. Morgan, T. C. Carney, and M. A. Kenamond. Reduction of dissipation in Lagrange cell-centered hydrodynamics (CCH) through corner gradient reconstruction (CGR). Journal of Computational Physics, 299(??):229–280, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004398.

Bastian:2019:MFM

[BMMP19] Peter Bastian, Eike Hermann Müller, Steffen Müthing, and Marian Piatkowski. Matrix-free multigrid blockpreconditioners for higher order discontinuous Galerkin discretisations. Journal of Computational Physics, 394(??): 417–439, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303973.

Beretta:2018:RPC

[BMPS18] Elena Beretta, Stefano Micheletti, Simona Perotto, and Matteo Santacesaria. Reconstruction of a piecewise constant conductivity on a polygonal partition via shape optimization in EIT. Journal of Computational Physics, 353(??):264–280, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307702.

Barlow:2016:ALE

[BMR^{+}16] Andrew J. Barlow, Pierre-Henri Maire, William J. Rider, Robert N. Rieben, and Mikhail J. Shashkov. Arbitrary Lagrangian–Eulerian methods for modeling high-speed compressible multimaterial flows. Journal of Computational Physics, 322(??):603–665, October 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302807.

Brugnano:2019:HOE

[BMR19] Luigi Brugnano, Juan I. Montijano, and Luis Rández. Highorder energy-conserving line integral methods for charged particle dynamics. Journal of Computational Physics, 396 (??):209–227, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0021999119304760.

Blazakis:2015:WCT

[BMRA^{+}15] Konstantinos N. Blazakis, Anotida Madzvamuse, Constantino Carlos Reyes-Aldasoro, Vanessa Styles, and Chandrasekhar Venkataraman. Whole cell tracking through the optimal control of geometric evolution laws. Journal of Computational Physics, 297(??):495–514, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003423.

Balsara:2016:EVF

[BMT16] Dinshaw S. Balsara, Gino I. Montecinos, and Eleuterio F. Toro. Exploring various flux vector splittings for the magnetohydrodynamic system. Journal of Computational Physics, 311(??):1– 21, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000371.

Birgle:2018:DDM

[BMT18] N. Birgle, R. Masson, and L. Trenty. A domain decomposition method to couple nonisothermal compositional gas liquid Darcy and free gas flows. Journal of Computational Physics, 368(??):210–235, September 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302602.

Beylkin:2019:RMM

[BMY19] Gregory Beylkin, Lucas Monzón, and Xinshuo Yang. Reduction of multivariate mixtures and its applications. Journal of Computational Physics, 383(??):94–124, April 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300403.

Balsara:2017:MRP

[BN17] Dinshaw S. Balsara and Boniface Nkonga. Multidimensional Riemann problem with self-similar internal structure — part III — a multidimensional analogue of the HLLI Riemann solver for conservative hyperbolic systems. Journal of Computational Physics, 346(??):25–48, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304217.

Berg:2019:DDD

[BN19] Jens Berg and Kaj Nyström. Data-driven discovery of PDEs in complex datasets. Journal of Computational Physics, 384(??):239–252, May 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300944.

Borah:2016:NSO

[BND16] Kalpajyoti Borah, Ganesh Natarajan, and Anoop K. Dass. A novel second-order flux splitting for ideal magnetohydrodynamics. Journal of Computational Physics, 313(??):159–180, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001248.

Bhole:2019:FSR

[BNGI19] Ashish Bhole, Boniface Nkonga, Sergey Gavrilyuk, and Kseniya Ivanova. Fluctuation splitting Riemann solver for a non-conservative modeling of shear shallow water flow. Journal of Computational Physics, 392(??):205–226, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302797.

Bennett:2018:MBF

[BNK18] W. P. Bennett, N. Nikiforakis, and R. Klein. A moving boundary flux stabilization method for Cartesian cut-cell grids using directional operator splitting. Journal of Computational Physics, 368(??):333–358, September 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830281X.

Barnett:2015:HOB

[BNM15] Alex H. Barnett, Bradley J. Nelson, and J. Matthew Mahoney. High-order boundary integral equation solution of high frequency wave scattering from obstacles in an unbounded linearly stratified medium. Journal of Computational Physics, 297(??):407–426, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500368X.

Besse:2017:DTB

[BNS17] Christophe Besse, Pascal Noble, and David Sanchez. Discrete transparent boundary conditions for the mixed KDV– BBM equation. Journal of Computational Physics, 345(??): 484–509, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304060.

Barral:2017:TAA

[BOA17] N. Barral, G. Olivier, and F. Alauzet. Time-accurate anisotropic mesh adaptation for three-dimensional timedependent problems with body-fitted moving geometries. Journal of Computational Physics, 331(??):157–187, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306155.

Budninskiy:2019:OAW

[BOD19] Max Budninskiy, Houman Owhadi, and Mathieu Desbrun. Operator-adapted wavelets for finite-element differential forms. Journal of Computational Physics, 388(??):144–177, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301354.

Boniface:2017:RRS

[Bon17] Jean-Christophe Boniface. Rescaling of the Roe scheme in low Mach-number flow regions. Journal of Computational Physics, 328(??):177–199, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305095.

Bourhrara:2019:NNM

[Bou19] Lahbib Bourhrara. A new numerical method for solving the Boltzmann transport equation using the PN method and the discontinuous finite elements on unstructured and curved meshes. Journal of Computational Physics, 397(??):Article 108801, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119304851.

Bernede:2018:UMC

[BP18] Adrien Bernede and Gaël Poëtte. An unsplit Monte-Carlo solver for the resolution of the linear Boltzmann equation coupled to (stiff) Bateman equations. Journal of Computational Physics, 354(??):211–241, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307805.

Botti:2019:HHO

[BPD19] Lorenzo Botti, Daniele A. Di Pietro, and Jérôme Droniou. A hybrid high-order method for the incompressible Navier– Stokes equations based on Temam’s device. Journal of Computational Physics, 376(??):786–816, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306715.

Briggs:2016:SPI

[BPF^{+}16] J. P. Briggs, S. J. Pennycook, J. R. Fergusson, J. Jäykkä, and E. P. S. Shellard. Separable projection integrals for higher-order correlators of the cosmic microwave sky: Acceleration by factors exceeding 100. Journal of Computational Physics, 310(??):285–300, April 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000279.

Bao:2016:GTS

[BPGS16] Y. Bao, R. Palacios, M. Graham, and S. Sherwin. Generalized thick strip modelling for vortex-induced vibration of long flexible cylinders. Journal of Computational Physics, 321(??):1079–1097, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302236.

Barnes:2019:SOS

[BPL19] M. Barnes, F. I. Parra, and M. Landreman. stella: an operator-split, implicit-explicit δf-gyrokinetic code for general magnetic field configurations. Journal of Computational Physics, 391(??):365–380, August 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930066X.

Braun:2018:RML

[BPM18] N. O. Braun, D. I. Pullin, and D. I. Meiron. Regularization method for large eddy simulations of shock-turbulence interactions. Journal of Computational Physics, 361(??):231–246, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300743.

Berrone:2016:TEF

[BPS16] Stefano Berrone, Sandra Pieraccini, and Stefano Scialò. Towards effective flow simulations in realistic discrete fracture networks. Journal of Computational Physics, 310(??):181–201, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000103.

Berrone:2017:FSP

[BPS17] Stefano Berrone, Sandra Pieraccini, and Stefano Scialò. Flow simulations in porous media with immersed intersecting fractures. Journal of Computational Physics, 345(??): 768–791, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304424.

Bian:2016:AEB

[BPTA16] Lei Bian, Gang Pang, Shaoqiang Tang, and Anton Arnold. ALmost EXact boundary conditions for transient Schrödinger– Poisson system. Journal of Computational Physics, 313(??): 233–246, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116000899.

Basting:2017:EAM

[BQCG17] Steffen Basting, Annalisa Quaini, Suncica Canić, and Roland Glowinski. Extended ALE method for fluid-structure interaction problems with large structural displacements. Journal of Computational Physics, 331(??):312–336, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306350.

Boscarino:2019:HOS

[BQRX19] Sebastiano Boscarino, Jing-Mei Qiu, Giovanni Russo, and Tao Xiong. A high order semi-implicit IMEX WENO scheme for the all-Mach isentropic Euler system. Journal of Computational Physics, 392(??):594–618, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303109.

Bidadi:2015:INV

[BR15a] Shreyas Bidadi and Sarma L. Rani. Investigation of numerical viscosities and dissipation rates of second-order TVD– MUSCL schemes for implicit large-eddy simulation. Journal of Computational Physics, 281(??):1003–1031, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007451.

Bidadi:2015:SDC

[BR15b] Shreyas Bidadi and Sarma L. Rani. On the stability and diffusive characteristics of Roe-MUSCL and Runge–Kutta schemes for inviscid Taylor-green vortex. Journal of Computational Physics, 299(??):339–351, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004581. See corrigendum [BR16].

Bidadi:2016:CSD

[BR16] Shreyas Bidadi and Sarma L. Rani. Corrigendum to “On the stability and diffusive characteristics of Roe-MUSCL and Runge–Kutta schemes for inviscid Taylor-Green vortex” [Journal of Computational Physics 299 (2015) 339–351]. Journal of Computational Physics, 305(??):1172, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006282. See [BR15b].

Bremer:2017:NPF

[BR17] James Bremer and Vladimir Rokhlin. On the nonoscillatory phase function for Legendre’s differential equation. Journal of Computational Physics, 350(??):326–342, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306186.

Brackbill:2016:EMC

[Bra16a] J. U. Brackbill. On energy and momentum conservation in particle-in-cell plasma simulation. Journal of Computational Physics, 317(??):405–427, July 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630105X.

Braeunig:2016:REP

[Bra16b] Jean-Philippe Braeunig. Reducing the entropy production in a collocated Lagrange–Remap scheme. Journal of Computational Physics, 314(??):127–144, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001595.

Brambley:2016:OFD

[Bra16c] E. J. Brambley. Optimized finite-difference (DRP) schemes perform poorly for decaying or growing oscillations. Journal of Computational Physics, 324(??):258–274, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303436.

Brehm:2017:CBC

[Bre17] C. Brehm. On consistent boundary closures for compact finite-difference WENO schemes. Journal of Computational Physics, 334(??):573–581, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300074.

Bremer:2018:ANE

[Bre18] James Bremer. An algorithm for the numerical evaluation of the associated Legendre functions that runs in time independent of degree and order. Journal of Computational Physics, 360(??):15–38, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830024X.

Bunder:2017:GCM

[BRK17] J. E. Bunder, A. J. Roberts, and I. G. Kevrekidis. Good coupling for the multiscale patch scheme on systems with microscale heterogeneity. Journal of Computational Physics, 337(??):154–174, May 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300979.

Bao:2018:FBI

[BRK^{+}18] Yuanxun Bao, Manas Rachh, Eric E. Keaveny, Leslie Greengard, and Aleksandar Donev. A fluctuating boundary integral method for Brownian suspensions. Journal of Computational Physics, 374(??):1094–1119, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305448.

Budd:2015:GAM

[BRW15] C. J. Budd, R. D. Russell, and E. Walsh. The geometry of radaptive meshes generated using optimal transport methods. Journal of Computational Physics, 282(??):113–137, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007591.

Bhaumik:2015:NVV

[BS15a] Swagata Bhaumik and Tapan K. Sengupta. A new velocityvorticity formulation for direct numerical simulation of 3D transitional and turbulent flows. Journal of Computational Physics, 284(??):230–260, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400847X.

Bo:2015:ARB

[BS15b] Wurigen Bo and Mikhail Shashkov. Adaptive reconnectionbased arbitrary Lagrangian Eulerian method. Journal of Computational Physics, 299(??):902–939, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004775.

Bencomo:2019:DMS

[BS19a] Mario J. Bencomo and William W. Symes. Discretization of multipole sources in a finite difference setting for wave propagation problems. Journal of Computational Physics, 386(??):296– 322, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911930097X.

Bermudez:2019:FES

[BS19b] Alfredo Bermúdez and Mohsen Shabani. Finite element solution of isothermal gas flow in a network. Journal of Computational Physics, 396(??):616–652, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304711.

Bakhos:2015:FAP

[BSK15] Tania Bakhos, Arvind K. Saibaba, and Peter K. Kitanidis. A fast algorithm for parabolic PDE-based inverse problems based on Laplace transforms and flexible Krylov solvers. Journal of Computational Physics, 299(??):940–954, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004520.

Brazell:2016:OMA

[BSM16] Michael J. Brazell, Jayanarayanan Sitaraman, and Dimitri J. Mavriplis. An overset mesh approach for 3D mixed element high-order discretizations. Journal of Computational Physics, 322(??):33–51, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911630256X.

Baptista:2019:SGA

[BSN19] Ricardo Baptista, Valentin Stolbunov, and Prasanth B. Nair. Some greedy algorithms for sparse polynomial chaos expansions. Journal of Computational Physics, 387(??):303–325, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300865.

Boccardo:2018:ISR

[BSP18] Gianluca Boccardo, Igor M. Sokolov, and Amir Paster. An improved scheme for a Robin boundary condition in discrete-time random walk algorithms. Journal of Computational Physics, 374(??):1152–1165, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830531X.

Borovikov:2015:ESO

[BST15] Dmitry Borovikov, Igor V. Sokolov, and Gábor Tóth. An efficient second-order accurate and continuous interpolation for block-adaptive grids. Journal of Computational Physics, 297(??):599–610, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003721.

Becker:2018:HRV

[BST^{+}18] G. Becker, C. M. Siefert, R. S. Tuminaro, H. Sun, D. M. Valiveti, A. Mohan, J. Yin, and H. Huang. High resolution viscous fingering simulation in miscible displacement using a p-adaptive discontinuous Galerkin method with algebraic multigrid preconditioner. Journal of Computational Physics, 374(??):495–514, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304571.

Barajas-Solano:2019:ABM

[BST19] D. A. Barajas-Solano and A. M. Tartakovsky. Approximate Bayesian model inversion for PDEs with heterogeneous and state-dependent coefficients. Journal of Computational Physics, 395(??):247–262, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304176.

Bologna:2015:DHM

[BSWG15] Mauro Bologna, Adam Svenkeson, Bruce J. West, and Paolo Grigolini. Diffusion in heterogeneous media: an iterative scheme for finding approximate solutions to fractional differential equations with time-dependent coefficients. Journal of Computational Physics, 293(??):297–311, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005816.

Bui-Thanh:2015:GUH

[BT15] Tan Bui-Thanh. From Godunov to a unified hybridized discontinuous Galerkin framework for partial differential equations. Journal of Computational Physics, 295(??):114–146, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002557.

Blachere:2016:AAP

[BT16] F. Blachère and R. Turpault. An admissibility and asymptoticpreserving scheme for systems of conservation laws with source term on 2D unstructured meshes. Journal of Computational Physics, 315(??):98–123, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001972.

Bakarji:2017:URB

[BT17a] Joseph Bakarji and Daniel M. Tartakovsky. On the use of reverse Brownian motion to accelerate hybrid simulations. Journal of Computational Physics, 334(??):68–80, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306866.

Biyikli:2017:MMM

[BT17b] Emre Biyikli and Albert C. To. Multiresolution molecular mechanics: Implementation and efficiency. Journal of Computational Physics, 328(??):27–45, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305083.

Bellotti:2019:CLS

[BT19] Thomas Bellotti and Maxime Theillard. A coupled level-set and reference map method for interface representation with applications to two-phase flows simulation. Journal of Computational Physics, 392(??):266–290, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303286.

Borowik:2017:MMC

[BTA17] Piotr Borowik, Jean-Luc Thobel, and Leszek Adamowicz. Modified Monte Carlo method for study of electron transport in degenerate electron gas in the presence of electronelectron interactions, application to graphene. Journal of Computational Physics, 341(??):397–405, July 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302826.

Balajewicz:2016:MSR

[BTD16] Maciej Balajewicz, Irina Tezaur, and Earl Dowell. Minimal subspace rotation on the Stiefel manifold for stabilization and enhancement of projection-based reduced order models for the compressible Navier–Stokes equations. Journal of Computational Physics, 321(??):224–241, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301826.

Balsara:2017:CEM

[BTGM17] Dinshaw S. Balsara, Allen Taflove, Sudip Garain, and Gino Montecinos. Computational electrodynamics in material media with constraint-preservation, multidimensional Riemann solvers and sub-cell resolution — Part I, secondorder FVTD schemes. Journal of Computational Physics, 349(??):604–635, November 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305326.

Britt:2018:NSW

[BTT18] S. Britt, S. Tsynkov, and E. Turkel. Numerical solution of the wave equation with variable wave speed on nonconforming domains by high-order difference potentials. Journal of Computational Physics, 354(??):26–42, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308161.

Beltman:2018:LSM

[BtTBI18] René Beltman, Jan ten Thije Boonkkamp, and Wilbert IJzerman. A least-squares method for the inverse reflector problem in arbitrary orthogonal coordinates. Journal of Computational Physics, 367(??):347–373, August 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302663.

Blais:2015:CLB

[BTVB15] Bruno Blais, Jean-Michel Tucny, David Vidal, and François Bertrand. A conservative lattice Boltzmann model for the volume-averaged Navier–Stokes equations based on a novel collision operator. Journal of Computational Physics, 294(??):258–273, August 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001916.

Busto:2016:DAA

[BTVC16] S. Busto, E. F. Toro, and M. E. Vázquez-Cendón. Design and analysis of ADER-type schemes for model advectiondiffusion-reaction equations. Journal of Computational Physics, 327(??):553–575, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304648.

Bu:2015:FDF

[BTWY15] Weiping Bu, Yifa Tang, Yingchuan Wu, and Jiye Yang. Finite difference/finite element method for two-dimensional space and time fractional Bloch-Torrey equations. Journal of Computational Physics, 293(??):264–279, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004380.

Bukac:2016:LCS

[Buk16] M. Bukac. A loosely-coupled scheme for the interaction between a fluid, elastic structure and poroelastic material. Journal of Computational Physics, 313(??):377–399, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001236.

Barbas:2015:DGM

[BV15] Alfonso Barbas and Pedro Velarde. Development of a Godunov method for Maxwell’s equations with adaptive mesh refinement. Journal of Computational Physics, 300(??):186–201, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500501X.

Brennan:2018:DDC

[BV18] Catherine Brennan and Daniele Venturi. Data-driven closures for stochastic dynamical systems. Journal of Computational Physics, 372(??):281–298, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304169.

Balsara:2016:TDR

[BVG^{+}16] Dinshaw S. Balsara, Jeaniffer Vides, Katharine Gurski, Boniface Nkonga, Michael Dumbser, Sudip Garain, and Edouard Audit. A two-dimensional Riemann solver with self-similar sub-structure — alternative formulation based on least squares projection. Journal of Computational Physics, 304(??):138–161, January 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006762.

Baars:2017:CPD

[BVM^{+}17a] S. Baars, J. P. Viebahn, T. E. Mulder, C. Kuehn, F. W. Wubs, and H. A. Dijkstra. Continuation of probability density functions using a generalized Lyapunov approach. Journal of Computational Physics, 336(??):627–643, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301146.

Biegert:2017:CMG

[BVM17b] Edward Biegert, Bernhard Vowinckel, and Eckart Meiburg. A collision model for grain-resolving simulations of flows over dense, mobile, polydisperse granular sediment beds. Journal of Computational Physics, 340(??):105–127, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302309.

Berger-Vergiat:2016:PPM

[BVMW16] Luc Berger-Vergiat, Colin McAuliffe, and Haim Waisman. Parallel preconditioners for monolithic solution of shear bands. Journal of Computational Physics, 304(??):359–379, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500621X.

Bodroski:2018:GBI

[BVS18] Zarko Bodroski, Nenad Vukmirović, and Srdjan Skrbic. Gaussian basis implementation of the charge patching method. Journal of Computational Physics, 368(??):196–209, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302572.

Boelens:2018:PTM

[BVT18] Arnout M. P. Boelens, Daniele Venturi, and Daniel M. Tartakovsky. Parallel tensor methods for high-dimensional linear PDEs. Journal of Computational Physics, 375(??):519– 539, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118305904.

Blonigan:2018:MSS

[BW18a] Patrick J. Blonigan and Qiqi Wang. Multiple shooting shadowing for sensitivity analysis of chaotic dynamical systems. Journal of Computational Physics, 354(??):447–475, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730791X.

Breuer:2018:MRC

[BW18b] Alex Breuer and Xin-Cindy Wang. More robust Chebyshev filtering for SCF iteration, with applications in real-space DFT. Journal of Computational Physics, 374(??):27–46, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308835.

binWaheed:2015:ETS

[bWAW15] Umair bin Waheed, Tariq Alkhalifah, and Hui Wang. Efficient traveltime solutions of the acoustic TI eikonal equation. Journal of Computational Physics, 282(??):62–76, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400758X.

Birrell:2015:BES

[BWR15] Jeremiah Birrell, Jon Wilkening, and Johann Rafelski. Boltzmann equation solver adapted to emergent chemical nonequilibrium. Journal of Computational Physics, 281(??):896– 916, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911400744X.

Bao:2017:ABE

[BXY17] Gang Bao, Liwei Xu, and Tao Yin. An accurate boundary element method for the exterior elastic scattering problem in two dimensions. Journal of Computational Physics, 348(??):343–363, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305405.

Bhrawy:2015:MBJ

[BZ15] A. H. Bhrawy and M. A. Zaky. A method based on the Jacobi tau approximation for solving multi-term time-space fractional partial differential equations. Journal of Computational Physics, 281(??):876–895, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007487.

Bao:2016:UAM

[BZ16a] Weizhu Bao and Xiaofei Zhao. A uniformly accurate multiscale time integrator spectral method for the Klein–Gordon– Zakharov system in the high-plasma-frequency limit regime. Journal of Computational Physics, 327(??):270–293, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304673.

Brown:2016:MHC

[BZ16b] David A. Brown and David W. Zingg. A monolithic homotopy continuation algorithm with application to computational fluid dynamics. Journal of Computational Physics, 321(??):55–75, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116301760.

Boom:2018:OHO

[BZ18] Pieter D. Boom and David W. Zingg. Optimization of high-order diagonally-implicit Runge–Kutta methods. Journal of Computational Physics, 371(??):168–191, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303206.

Bai:2019:NVH

[BZ19a] Zeyu Bai and Xiaolin Zhong. New very high-order upwind multi-layer compact (MLC) schemes with spectrallike resolution for flow simulations. Journal of Computational Physics, 378(??):63–109, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307162.

Bao:2019:CNM

[BZ19b] Weizhu Bao and Xiaofei Zhao. Comparison of numerical methods for the nonlinear Klein–Gordon equation in the nonrelativistic limit regime. Journal of Computational Physics, 398(??):Article 108886, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305844.

Chatterjee:2018:CSP

[CAA18] Sabyasachi Chatterjee, Amit Acharya, and Zvi Artstein. Computing singularly perturbed differential equations. Journal of Computational Physics, 354(??):417–446, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307787.

Cacuci:2015:SOAa

[Cac15a] Dan G. Cacuci. Second-order adjoint sensitivity analysis methodology (2nd-ASAM) for computing exactly and efficiently first- and second-order sensitivities in large-scale linear systems: I. computational methodology. Journal of Computational Physics, 284(??):687–699, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008596.

Cacuci:2015:SOAb

[Cac15b] Dan G. Cacuci. Second-order adjoint sensitivity analysis methodology (2nd-ASAM) for computing exactly and efficiently first- and second-order sensitivities in large-scale linear systems: II. illustrative application to a paradigm particle diffusion problem. Journal of Computational Physics, 284(??):700–717, March 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400792X.

Cai:2016:SIS

[Cai16] Tao Cai. A semi-implicit spectral method for compressible convection of rotating and density-stratified flows in Cartesian geometry. Journal of Computational Physics, 310(??):342–360, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000309.

Capecelatro:2018:PLM

[Cap18] Jesse Capecelatro. A purely Lagrangian method for simulating the shallow water equations on a sphere using smooth particle hydrodynamics. Journal of Computational Physics, 356(??):174–191, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911730877X.

Capizzano:2019:HRC

[CARdN19] Francesco Capizzano, Luigi Alterio, Serena Russo, and Carlo de Nicola. A hybrid RANS–LES Cartesian method based on a skew-symmetric convective operator. Journal of Computational Physics, 390(??):359–379, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302359.

Cavaglieri:2015:LSI

[CB15] Daniele Cavaglieri and Thomas Bewley. Low-storage implicit/explicit Runge–Kutta schemes for the simulation of stiff high-dimensional ODE systems. Journal of Computational Physics, 286(??):172–193, April 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000352.

Candy:2018:STG

[CB18a] J. Candy and E. A. Belli. Spectral treatment of gyrokinetic shear flow. Journal of Computational Physics, 356 (??):448–457, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117309063.

Collins:2018:GSS

[CB18b] James P. Collins and Peter S. Bernard. A gridfree scheme for simulation of natural convection in three dimensions. Journal of Computational Physics, 369(??):209–224, September 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303048.

Cheng:2019:ROM

[CB19] C. Cheng and A. P. Bunger. Reduced order model for simultaneous growth of multiple closely-spaced radial hydraulic fractures. Journal of Computational Physics, 376(??):228–248, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305989.

Carlberg:2017:GVL

[CBA17] Kevin Carlberg, Matthew Barone, and Harbir Antil. Galerkin v. least-squares Petrov–Galerkin projection in nonlinear model reduction. Journal of Computational Physics, 330(??):693–734, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305319.

Candy:2016:HAE

[CBB16] J. Candy, E. A. Belli, and R. V. Bravenec. A high-accuracy Eulerian gyrokinetic solver for collisional plasmas. Journal of Computational Physics, 324(??):73–93, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303400.

Cavalca:2018:DCA

[CBC^{+}18] D. F. Cavalca, C. Bringhenti, G. B. Campos, J. T. Tomita, and O. F. R. Silva. Development and convergence analysis of an effective and robust implicit Euler solver for 3D unstructured grids. Journal of Computational Physics, 367(??):399–415, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302195.

Cook:2019:MMS

[CBM19] S. P. Cook, C. J. Budd, and T. Melvin. The moving mesh semi-Lagrangian MMSISL method. Journal of Computational Physics, 393(??):484–502, September 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300956.

Christon:2016:HIP

[CBN^{+}16] Mark A. Christon, Jozsef Bakosi, Balasubramanya T. Nadiga, Markus Berndt, Marianne M. François, Alan K. Stagg, Yidong Xia, and Hong Luo. A hybrid incremental projection method for thermal-hydraulics applications. Journal of Computational Physics, 317(??):382–404, July 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301279.

Capuano:2018:SVC

[CBS18] M. Capuano, C. Bogey, and P. D. M. Spelt. Simulations of viscous and compressible gas–gas flows using high-order finite difference schemes. Journal of Computational Physics, 361(??):56–81, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300573.

Cai:2018:DLG

[CBZ18] Jiaxiang Cai, Chuanzhi Bai, and Haihui Zhang. Decoupled local/global energy-preserving schemes for the Ncoupled nonlinear Schrödinger equations. Journal of Computational Physics, 374(??):281–299, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305126.

Castillo:2015:FST

[CC15] E. Castillo and R. Codina. First, second and third order fractional step methods for the three-field viscoelastic flow problem. Journal of Computational Physics, 296(??): 113–137, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115002831.

Chacon:2016:CFI

[CC16a] L. Chacón and G. Chen. A curvilinear, fully implicit, conservative electromagnetic PIC algorithm in multiple dimensions. Journal of Computational Physics, 316(??):578–597, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300717.

Chakraborty:2016:SED

[CC16b] Souvik Chakraborty and Rajib Chowdhury. Sequential experimental design based generalised ANOVA. Journal of Computational Physics, 317(??):15–32, July 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300882.

Cinnella:2016:HOI

[CC16c] P. Cinnella and C. Content. High-order implicit residual smoothing time scheme for direct and large eddy simulations of compressible flows. Journal of Computational Physics, 326(??):1–29, December 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303734.

Chakraborty:2017:EAB

[CC17a] Souvik Chakraborty and Rajib Chowdhury. An efficient algorithm for building locally refined hp-adaptive H-PCFE: Application to uncertainty quantification. Journal of Computational Physics, 351(??):59–79, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306836.

Chalons:2017:NCC

[CC17b] Christophe Chalons and Frédéric Coquel. A new comment on the computation of non-conservative products using Roe-type path conservative schemes. Journal of Computational Physics, 335(??):592–604, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300268.

Chen:2017:NEF

[CC17c] Shu C. Chen and Weng Cho Chew. Numerical electromagnetic frequency domain analysis with discrete exterior calculus. Journal of Computational Physics, 350(??):668–689, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306496.

Chacon:2019:ECP

[CC19a]  L. Chacón and G. Chen.  Energy-conserving perfect-

conductor boundary conditions for an implicit, curvilinear Darwin particle-in-cell algorithm. Journal of Computational Physics, 391(??):216–225, August 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302785.

Chen:2019:NHR

[CC19b] Duan Chen and Wei Cai. An O(N log N) hierarchical random compression method for kernel matrices by sampling partial matrix entries. Journal of Computational Physics, 397(??):Article 108828, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119305121.

Cullen:2019:FSA

[CC19c] Andrew C. Cullen and Simon R. Clarke. A fast, spectrally accurate homotopy based numerical method for solving nonlinear differential equations. Journal of Computational Physics, 385(??):106–118, May 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301160.

Chen:2019:RHS

[CCB^{+}19] Chao Chen, Leopold Cambier, Erik G. Boman, Sivasankaran Rajamanickam, Raymond S. Tuminaro, and Eric Darve. A robust hierarchical solver for ill-conditioned systems with applications to ice sheet modeling. Journal of Computational Physics, 396(??):819–836, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305091.

Capuano:2015:EPT

[CCBdL15] F. Capuano, G. Coppola, G. Balarac, and L. de Luca. Energy preserving turbulent simulations at a reduced computational cost. Journal of Computational Physics, 298(??):480–494, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003976.

Chen:2019:SRP

[CCBF19] Zhuo Chen, Matthew S. B. Coleman, Eric G. Blackman, and Adam Frank. Solving the Riemann problem for realistic astrophysical fluids. Journal of Computational Physics, 388(??):490–517, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301998.

Capuano:2015:ETA

[CCdL15] F. Capuano, G. Coppola, and L. de Luca. An efficient time advancing strategy for energy-preserving simulations. Journal of Computational Physics, 295(??):209–229, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002429.

Crestetto:2019:ACD

[CCDL19] Anaïs Crestetto, Nicolas Crouseilles, Giacomo Dimarco, and Mohammed Lemou. Asymptotically complexity diminishing schemes (ACDS) for kinetic equations in the diffusive scaling. Journal of Computational Physics, 394(??):243–262, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303705.

Cai:2019:LSB

[CCFC19] Xiaohui Cai, Guoxing Chen, Xiaoping Fan, and Yangkang Chen. Least-squares based rectangular-grid cross and rhombus stencils for acoustic wave propagation and reverse time migration. Journal of Computational Physics, 392(??):335– 353, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303146.

Crockatt:2017:AOF

[CCGH17] Michael M. Crockatt, Andrew J. Christlieb, C. Kristopher Garrett, and Cory D. Hauck. An arbitrary-order, fully implicit, hybrid kinetic solver for linear radiative transport using integral deferred correction. Journal of Computational Physics, 346(??):212–241, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730462X.

Crockatt:2019:HMR

[CCGH19] Michael M. Crockatt, Andrew J. Christlieb, C. Kristopher Garrett, and Cory D. Hauck. Hybrid methods for radiation transport using diagonally implicit Runge–Kutta and space-time discontinuous Galerkin time integration. Journal of Computational Physics, 376(??):455–477, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306430.

Cao:2015:IIF

[CCHL15] Yong Cao, Yuchuan Chu, Xiaoming He, and Tao Lin. An iterative immersed finite element method for an electric potential interface problem based on given surface electric quantity. Journal of Computational Physics, 281(??):82–95, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006962.

Chacon:2017:MHO

[CCK^{+}17] L. Chacón, G. Chen, D. A. Knoll, C. Newman, H. Park, W. Taitano, J. A. Willert, and G. Womeldorff. Multiscale high-order/low-order (HOLO) algorithms and applications. Journal of Computational Physics, 330(??):21–45, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305770.

Chertock:2018:WBS

[CCK^{+}18] Alina Chertock, Shumo Cui, Alexander Kurganov, Seyma Nur Özcan, and Eitan Tadmor. Well-balanced schemes for the Euler equations with gravitation: Conservative formulation using global fluxes. Journal of Computational Physics, 358(??):36– 52, April 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117309208.

Cheung:2015:SDG

[CCKQ15] Siu Wun Cheung, Eric Chung, Hyea Hyun Kim, and Yue Qian. Staggered discontinuous Galerkin methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 302(??):251–266, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005513.

Collyer:2016:ISV

[CCL16] B. S. Collyer, C. Connaughton, and D. A. Lockerby. Importance sampling variance reduction for the Fokker–Planck rarefied gas particle method. Journal of Computational Physics, 325(??):116–128, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303503.

Collin:2015:LOR

[CCM15] Annabelle Collin, Dominique Chapelle, and Philippe Moireau. A Luenberger observer for reaction-diffusion models with front position data. Journal of Computational Physics, 300 (??):288–307, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004891.

Costa:2017:SOF

[CCM17] Ricardo Costa, Stéphane Clain, and Gaspar J. Machado. A sixth-order finite volume scheme for the steady-state incompressible Stokes equations on staggered unstructured meshes. Journal of Computational Physics, 349(??):501–527, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305557.

Carrillo:2019:SPS

[CCP19] José A. Carrillo, Young-Pil Choi, and Lorenzo Pareschi. Structure preserving schemes for the continuum Kuramoto model: Phase transitions. Journal of Computational Physics, 376(??):365–389, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830651X.

Coppola:2019:NSF

[CCPdL19] G. Coppola, F. Capuano, S. Pirozzoli, and L. de Luca. Numerically stable formulations of convective terms for turbulent compressible flows. Journal of Computational Physics, 382(??):86– 104, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300245.

Capuano:2017:ERK

[CCRdL17] F. Capuano, G. Coppola, L. Rández, and L. de Luca. Explicit Runge–Kutta schemes for incompressible flow with improved energy-conservation properties. Journal of Computational Physics, 328(??):86–94, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305381.

Carlberg:2018:CMR

[CCS18] Kevin Carlberg, Youngsoo Choi, and Syuzanna Sargsyan. Conservative model reduction for finite-volume models. Journal of Computational Physics, 371(??):280–314, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830319X.

Chen:2018:QWM

[CCWY18] Jing Chen, Yifan Chen, Hao Wu, and Dinghui Yang. The quadratic Wasserstein metric for earthquake location. Journal of Computational Physics, 373(??):188–209, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304443.

Cheng:2015:NST

[CCZ15] Yingda Cheng, Andrew J. Christlieb, and Xinghui Zhong. Numerical study of the two-species Vlasov–Ampère system: Energy-conserving schemes and the current-driven ion-acoustic instability. Journal of Computational Physics, 288(??):66–85, May 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500090X.

Chartier:2018:NMT

[CCZ18] Philippe Chartier, Nicolas Crouseilles, and Xiaofei Zhao. Numerical methods for the two-dimensional Vlasov–Poisson equation in the finite Larmor radius approximation regime. Journal of Computational Physics, 375(??):619–640, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306016.

Chen:2016:AEN

[CCZC16] Duan Chen, Wei Cai, Brian Zinser, and Min Hyung Cho. Accurate and efficient Nyström volume integral equation method for the Maxwell equations for multiple 3-D scatterers. Journal of Computational Physics, 321(??):303–320, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301875.

Chiodi:2017:RCL

[CD17] Robert Chiodi and Olivier Desjardins. A reformulation of the conservative level set reinitialization equation for accurate and robust simulation of complex multiphase flows. Journal of Computational Physics, 343(??):186–200, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303327.

Chaillat:2017:TII

[CDC17] Stéphanie Chaillat, Luca Desiderio, and Patrick Ciarlet. Theory and implementation of H-matrix based iterative and direct solvers for Helmholtz and elastodynamic oscillatory kernels. Journal of Computational Physics, 351(??):165–186, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306721.

Chehade:2019:SFM

[CDDL19] Samar Chehade, Audrey Kamta Djakou, Michel Darmon, and Gilles Lebeau. The spectral functions method for acoustic wave diffraction by a stress-free wedge: Theory and validation. Journal of Computational Physics, 377(??):200–218, January 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307071.

Chaillat:2017:FIB

[CDL17] Stéphanie Chaillat, Marion Darbas, and Frédérique Le Louër. Fast iterative boundary element methods for high-frequency scattering problems in 3D elastodynamics. Journal of Computational Physics, 341(??):429–446, July 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302917.

Chehade:2019:EPW

[CDL19] Samar Chehade, Michel Darmon, and Gilles Lebeau. 2D elastic plane-wave diffraction by a stress-free wedge of arbitrary angle. Journal of Computational Physics, 394(??):532–558, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119304231.

Carpio:2019:WTD

[CDLR19] Ana Carpio, Thomas G. Dimiduk, Frédérique Le Louër, and María Luisa Rapún. When topological derivatives met regularized Gauss–Newton iterations in holographic 3D imaging. Journal of Computational Physics, 388(??):224–251, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302104.

Cances:2016:PMB

[CDM^{+}16] Eric Cancès, Geneviève Dusson, Yvon Maday, Benjamin Stamm, and Martin Vohralík. A perturbation-method-based post-processing for the planewave discretization of Kohn–Sham models. Journal of Computational Physics, 307(??):446–459, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500830X.

Cerroni:2018:PMC

[CDM18] Daniele Cerroni, Roberto Da Vià, and Sandro Manservisi. A projection method for coupling two-phase VOF and fluid structure interaction simulations. Journal of Computational Physics, 354(??):646–671, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308227.

Chirco:2019:VES

[CDM19] L. Chirco, R. Da Vià, and S. Manservisi. VOF evaluation of the surface tension by using variational representation and Galerkin interpolation projection. Journal of Computational Physics, 395(??):537–562, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304449.

Chertock:2017:APM

[CDN17] Alina Chertock, Pierre Degond, and Jochen Neusser. An asymptotic-preserving method for a relaxation of the Navier– Stokes–Korteweg equations. Journal of Computational Physics, 335(??):387–403, April 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300463.

Chow:2019:AOC

[CDOY19] Yat Tin Chow, Jérôme Darbon, Stanley Osher, and Wotao Yin. Algorithm for overcoming the curse of dimensionality for state-dependent Hamilton–Jacobi equations. Journal of Computational Physics, 387(??):376–409, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301093.

Couderc:2017:EAP

[CDV17] F. Couderc, A. Duran, and J.-P. Vila. An explicit asymptotic preserving low Froude scheme for the multilayer shallow water model with density stratification. Journal of Computational Physics, 343(??):235–270, August 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302899.

Chen:2018:PWA

[CDX^{+}18a] Xiang Chen, Adrian Diaz, Liming Xiong, David L. McDowell, and Youping Chen. Passing waves from atomistic to continuum. Journal of Computational Physics, 354(??):393–402, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308057.

Chong:2018:MRS

[CDX18b] Kai Leong Chong, Guangyu Ding, and Ke-Qing Xia. Multipleresolution scheme in finite-volume code for active or passive scalar turbulence. Journal of Computational Physics, 375(??):1045–1058, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306132.

Chun:2017:MMFa

[CE17] S. Chun and C. Eskilsson. Method of moving frames to solve the shallow water equations on arbitrary rotating curved surfaces. Journal of Computational Physics, 333(??):1–23, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306660.

Chan:2018:MLA

[CE18] Shing Chan and Ahmed H. Elsheikh. A machine learning approach for efficient uncertainty quantification using multiscale methods. Journal of Computational Physics, 354(??):493–511, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307933.

Crouseilles:2015:HSV

[CEF15] Nicolas Crouseilles, Lukas Einkemmer, and Erwan Faou. Hamiltonian splitting for the Vlasov–Maxwell equations. Journal of Computational Physics, 283(??):224–240, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007918.

Chung:2016:AMM

[CEH16] Eric Chung, Yalchin Efendiev, and Thomas Y. Hou. Adaptive multiscale model reduction with Generalized Multiscale Finite Element Methods. Journal of Computational Physics, 320(??):69–95, September 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116301097.

Chu:2019:RPD

[CEHM19] Ran Chu, Eirik Endeve, Cory D. Hauck, and Anthony Mezzacappa. Realizability-preserving DG-IMEX method for the two-moment model of fermion transport. Journal of Computational Physics, 389(??):62–93, July 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302207.

Chung:2015:RDO

[CEL15] Eric T. Chung, Yalchin Efendiev, and Wing Tat Leung. Residual-driven online generalized multiscale finite element methods. Journal of Computational Physics, 302(??):176–190, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005744.

Chung:2018:FOG

[CEL18a] Eric T. Chung, Yalchin Efendiev, and Wing Tat Leung. Fast online generalized multiscale finite element method using constraint energy minimization. Journal of Computational Physics, 355(??):450–463, February 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308604.

Chung:2018:NLM

[CEL^{+}18b] Eric T. Chung, Yalchin Efendiev, Wing Tat Leung, Maria Vasilyeva, and Yating Wang. Non-local multi-continua upscaling for flows in heterogeneous fractured media. Journal of Computational Physics, 372(??):22–34, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303474.

Constantine:2015:EAS

[CELI15] P. G. Constantine, M. Emory, J. Larsson, and G. Iaccarino. Exploiting active subspaces to quantify uncertainty in the numerical simulation of the HyShot II scramjet. Journal of Computational Physics, 302(??):1–20, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500580X.

Chung:2018:CBG

[CELZ18] Eric T. Chung, Yalchin Efendiev, Wing Tat Leung, and Zhiwen Zhang. Cluster-based generalized multiscale finite element method for elliptic PDEs with random coefficients. Journal of Computational Physics, 371(??):606–617, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303504.

Ceniceros:2019:EOA

[Cen19] Hector D. Ceniceros. Efficient order-adaptive methods for polymer self-consistent field theory. Journal of Computational Physics, 386(??):9–21, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301445.

Caputo:2015:DFD

[CF15] Michele Caputo and Mauro Fabrizio. Damage and fatigue described by a fractional derivative model. Journal of Computational Physics, 293(??):400–408, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007645.

Chen:2019:DEC

[CF19] Guodong Chen and Krzysztof J. Fidkowski. Discretization error control for constrained aerodynamic shape optimization. Journal of Computational Physics, 387(??):163–185, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911930155X.

Colbrook:2018:FMS

[CFF18] Matthew J. Colbrook, Natasha Flyer, and Bengt Fornberg. On the Fokas method for the solution of elliptic problems in both convex and non-convex polygonal domains. Journal of Computational Physics, 374(??):996–1016, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305278.

Carraro:2016:CVD

[CFG16] Thomas Carraro, Elfriede Friedmann, and Daniel Gerecht. Coupling vs decoupling approaches for PDE/ODE systems modeling intercellular signaling. Journal of Computational Physics, 314(??):522–537, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001728.

Camminady:2019:REM

[CFKK19] Thomas Camminady, Martin Frank, Kerstin Küpper, and Jonas Kusch. Ray effect mitigation for the discrete ordinates method through quadrature rotation. Journal of Computational Physics, 382(??):105–123, April 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300415.

Casas:2018:ABF

[CFO18] G. Casas, A. Ferrer, and E. Oñate. Approximating the Basset force by optimizing the method of van Hinsberg et al. Journal of Computational Physics, 352(??):142–171, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307337.

Cheung:2017:HCA

[CFPB17] James Cheung, Amalie L. Frischknecht, Mauro Perego, and Pavel Bochev. A hybrid, coupled approach for modeling charged fluids from the nano to the mesoscale. Journal of Computational Physics, 348(??):364–384, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305387.

Carmouze:2018:CRB

[CFSN18] Quentin Carmouze, François Fraysse, Richard Saurel, and Boniface Nkonga. Coupling rigid bodies motion with single phase and two-phase compressible flows on unstructured meshes. Journal of Computational Physics, 375(??):1314–1338, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305461.

Christlieb:2016:HOP

[CFST16] Andrew J. Christlieb, Xiao Feng, David C. Seal, and Qi Tang. A high-order positivity-preserving single-stage single-step method for the ideal magnetohydrodynamic equations. Journal of Computational Physics, 316(??):218–242, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300481.

Cheng:2019:AAR

[CFSZ19] Ziqiang Cheng, Jinwei Fang, Chi-Wang Shu, and Mengping Zhang. Assessment of aeroacoustic resolution properties of DG schemes and comparison with DRP schemes. Journal of Computational Physics, 399(??):Article 108960, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119306655.

Cusini:2018:ADM

[CFvKH18] Matteo Cusini, Barnaby Fryer, Cor van Kruijsdijk, and Hadi Hajibeygi. Algebraic dynamic multilevel method for compositional flow in heterogeneous porous media. Journal of Computational Physics, 354(??):593–612, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308197.

Cleveland:2015:UHI

[CG15] Mathew A. Cleveland and Nick Gentile. Using hybrid implicit Monte Carlo diffusion to simulate gray radiation hydrodynamics. Journal of Computational Physics, 291(??):1–19, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001060.

Coquerelle:2016:FOA

[CG16] Mathieu Coquerelle and Stéphane Glockner. A fourth-order accurate curvature computation in a level set framework for two-phase flows subjected to surface tension forces. Journal of Computational Physics, 305(??):838–876, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007548.

Chen:2018:PTA

[CG18a] Jiahui Chen and Weihua Geng. On preconditioning the treecode-accelerated boundary integral (TABI) Poisson– Boltzmann solver. Journal of Computational Physics, 373(??):750–762, November 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304728.

Cohen:2018:EDF

[CG18b] Ido Cohen and Guy Gilboa. Energy dissipating flows for solving nonlinear eigenpair problems. Journal of Computational Physics, 375(??):1138–1158, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306065.

Castillo:2019:OST

[CG19] Paul Castillo and Sergio Gómez. Optimal stabilization and time step constraints for the forward Euler–Local Discontinuous Galerkin method applied to fractional diffusion equations. Journal of Computational Physics, 394(??):503–521, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119304127.

Chen:2017:EMS

[CGC17] Jingrun Chen and Carlos J. García-Cervera. An efficient multigrid strategy for large-scale molecular mechanics optimization. Journal of Computational Physics, 342(??):29–42, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303066.

Cusimano:2018:SFM

[CGG18] N. Cusimano and L. Gerardo-Giorda. A space-fractional monodomain model for cardiac electrophysiology combining anisotropy and heterogeneity on realistic geometries. Journal of Computational Physics, 362(??):409–424, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301165.

Christlieb:2016:WBM

[CGJ16] Andrew Christlieb, Wei Guo, and Yan Jiang. A WENO-based Method of Lines Transpose approach for Vlasov simulations. Journal of Computational Physics, 327(??):337–367, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304697.

Christlieb:2019:KBH

[CGJ19] Andrew Christlieb, Wei Guo, and Yan Jiang. A kernel based high order “explicit” unconditionally stable scheme for time dependent Hamilton–Jacobi equations. Journal of Computational Physics, 379(??):214–236, February 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307812.

Christlieb:2019:MMW

[CGJY19] Andrew Christlieb, Wei Guo, Yan Jiang, and Hyoseon Yang. A moving mesh WENO method based on exponential polynomials for one-dimensional conservation laws. Journal of Computational Physics, 380(??):334–354, March 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118308076.

Chalons:2017:ARL

[CGK17] Christophe Chalons, Mathieu Girardin, and Samuel Kokh. An all-regime Lagrange-Projection like scheme for 2D homogeneous models for two-phase flows on unstructured meshes. Journal of Computational Physics, 335(??):885–904, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730027X.

Chaillat:2018:MBA

[CGL18] Stéphanie Chaillat, Samuel P. Groth, and Adrien Loseille. Metric-based anisotropic mesh adaptation for 3D acoustic boundary element methods. Journal of Computational Physics, 372(??):473–499, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304261.

Corrado:2015:IWC

[CGM15] Cesare Corrado, Jean-Frédéric Gerbeau, and Philippe Moireau. Identification of weakly coupled multiphysics problems. application to the inverse problem of electrocardiography. Journal of Computational Physics, 283(??):271–298, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008031.

Chandrasekaran:2018:MSN

[CGM18] S. Chandrasekaran, C. H. Gorman, and H. N. Mhaskar. Minimum Sobolev norm interpolation of scattered derivative data. Journal of Computational Physics, 365(??):149–172, July 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301669.

Chabot:2018:HOD

[CGMH18] S. Chabot, N. Glinsky, E. D. Mercerat, and L. F. Bonilla Hidalgo. A high-order discontinuous Galerkin method for 1D wave propagation in a nonlinear heterogeneous medium. Journal of Computational Physics, 355(??):191–213, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308513.

Clair:2016:MDF

[CGP16] G. Clair, J.-M. Ghidaglia, and J.-P. Perlat. A multidimensional finite volume cell-centered direct ALE solver for hydrodynamics. Journal of Computational Physics, 326 (??):312–333, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304077.

Cai:2018:HOS

[CGQ18] Xiaofeng Cai, Wei Guo, and Jing-Mei Qiu. A high order semi-Lagrangian discontinuous Galerkin method for Vlasov– Poisson simulations without operator splitting. Journal of Computational Physics, 354(??):529–551, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730815X.

Corona:2017:IEF

[CGRV17] Eduardo Corona, Leslie Greengard, Manas Rachh, and Shravan Veerapaneni. An integral equation formulation for rigid bodies in Stokes flow in three dimensions. Journal of Computational Physics, 332(??):504–519, March 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306714.

Chou:2015:ELM

[CGS15] Yi-Ju Chou, Shih-Hung Gu, and Yun-Chuan Shao. An Euler– Lagrange model for simulating fine particle suspension in liquid flows. Journal of Computational Physics, 299(??):955–973, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004830.

Carichino:2018:EBO

[CGS18] Lucia Carichino, Giovanna Guidoboni, and Marcela Szopos. Energy-based operator splitting approach for the time discretization of coupled systems of partial and ordinary differential equations for fluid flows: The Stokes case. Journal of Computational Physics, 364(??):235–256, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301128.

Calderer:2018:FSI

[CGSS18] Antoni Calderer, Xin Guo, Lian Shen, and Fotis Sotiropoulos. Fluid-structure interaction simulation of floating structures interacting with complex, large-scale ocean waves and atmospheric turbulence with application to floating offshore wind turbines. Journal of Computational Physics, 355(??):144–175, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308331.

Casacuberta:2018:EES

[CGTH18] Jordi Casacuberta, Koen J. Groot, Henry J. Tol, and Stefan Hickel. Effectivity and efficiency of selective frequency damping for the computation of unstable steady-state solutions. Journal of Computational Physics, 375(??):481–497, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305898.

Chen:2017:DSM

[CH17] Li-Chieh Chen and Mei-Jiau Huang. A DFFD simulation method combined with the spectral element method for solid-fluid-interaction problems. Journal of Computational Physics, 330(??):749–769, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305587.

Chen:2019:IWE

[CH19] Li Li Chen and Cong Huang. An improved WLS–ENO method for solving hyperbolic conservation laws. Journal of Computational Physics, 392(??):96–114, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303122.

Chatterjee:2016:NGF

[Cha16] Kausik Chatterjee. A new Green’s function Monte Carlo algorithm for the estimation of the derivative of the solution of Helmholtz equation subject to Neumann and mixed boundary conditions. Journal of Computational Physics, 315(??):264– 272, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300110.

Chan:2018:DEC

[Cha18] Jesse Chan. On discretely entropy conservative and entropy stable discontinuous Galerkin methods. Journal of Computational Physics, 362(??):346–374, June 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301153.

Cho:2018:HFL

[CHCC18] Min Hyung Cho, Jingfang Huang, Dangxing Chen, and Wei Cai. A heterogeneous FMM for layered media Helmholtz equation I: Two layers in R^{2}. Journal of Computational Physics, 369(??):237–251, September 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302997.

Crean:2018:ESS

[CHD^{+}18] Jared Crean, Jason E. Hicken, David C. Del Rey Fernández, David W. Zingg, and Mark H. Carpenter. Entropystable summation-by-parts discretization of the Euler equations on general curved elements. Journal of Computational Physics, 356(??):410–438, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117308999.

Cohen:2017:MCC

[CHE^{+}17] B. I. Cohen, D. P. Higginson, C. D. Eng, W. A. Farmer, A. Friedman, D. P. Grote, and D. J. Larson. Monte Carlo calculation of large and small-angle electron scattering in air. Journal of Computational Physics, 349(??):582–588, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305855.

Chen:2018:RHI

[Che18] Juan Chen. A review of hybrid implicit explicit finite difference time domain method. Journal of Computational Physics, 363(??):256–267, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830144X.

Chern:2019:RDP

[Che19] Albert Chern. A reflectionless discrete perfectly matched layer. Journal of Computational Physics, 381(??):91–109, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300075.

Chiu:2019:CPF

[Chi19] Pao-Hsiung Chiu. A coupled phase field framework for solving incompressible two-phase flows. Journal of Computational Physics, 392(??):115–140, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303225.

Cheng:2017:TOM

[CHJT17] Jun-Bo Cheng, Weizhang Huang, Song Jiang, and Baolin Tian. A third-order moving mesh cell-centered scheme for one-dimensional elastic-plastic flows. Journal of Computational Physics, 349(??):137–153, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305892.

Chou:2019:FEB

[CHL^{+}19] So-Hsiang Chou, Tsung-Ming Huang, Tiexiang Li, Jia-Wei Lin, and Wen-Wei Lin. A finite element based fast eigensolver for three dimensional anisotropic photonic crystals. Journal of Computational Physics, 386(??):611–631, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301469.

Cui:2017:SSM

[CHLZ17] Jianbo Cui, Jialin Hong, Zhihui Liu, and Weien Zhou. Stochastic symplectic and multi-symplectic methods for nonlinear Schrödinger equation with white noise dispersion. Journal of Computational Physics, 342(??):267–285, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303005.

Choi:2015:HSD

[Cho15] Jung J. Choi. Hybrid spectral difference/embedded finite volume method for conservation laws. Journal of Computational Physics, 295(??):285–306, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002594.

Cho:2019:SAN

[Cho19] Min Hyung Cho. Spectrally-accurate numerical method for acoustic scattering from doubly-periodic 3D multilayered media. Journal of Computational Physics, 393(??):46–58, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303353.

Charnyi:2017:CLN

[CHOR17] Sergey Charnyi, Timo Heister, Maxim A. Olshanskii, and Leo G. Rebholz. On conservation laws of Navier–Stokes Galerkin discretizations. Journal of Computational Physics, 337(??):289–308, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911730133X.

Coquel:2017:PES

[CHS17] Frédéric Coquel, Jean-Marc Hérard, and Khaled Saleh. A positive and entropy-satisfying finite volume scheme for the Baer–Nunziato model. Journal of Computational Physics, 330(??):401–435, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306039.

Castelletto:2017:MFE

[CHT17] Nicola Castelletto, Hadi Hajibeygi, and Hamdi A. Tchelepi. Multiscale finite-element method for linear elastic geomechanics. Journal of Computational Physics, 331(??):337–356, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306362.

Chun:2017:MMFb

[Chu17] Sehun Chun. Method of moving frames to solve timedependent Maxwell’s equations on anisotropic curved surfaces: Applications to invisible cloak and ELF propagation. Journal of Computational Physics, 340(??):85–104, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302267.

Chen:2016:TOM

[CHY16] Zheng Chen, Hongying Huang, and Jue Yan. Third order maximum-principle-satisfying direct discontinuous Galerkin methods for time dependent convection diffusion equations on unstructured triangular meshes. Journal of Computational Physics, 308(??):198–217, March 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500858X.

Chen:2016:PPP

[CHZ16] Chuchu Chen, Jialin Hong, and Liying Zhang. Preservation of physical properties of stochastic Maxwell equations with additive noise via stochastic multi-symplectic methods. Journal of Computational Physics, 306(??):500–519, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007925.

Cifani:2019:ACC

[Cif19] Paolo Cifani. Analysis of a constant-coefficient pressure equation method for fast computations of two-phase flows at high density ratios. Journal of Computational Physics, 398(??):Article 108904, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306096.

Cortinovis:2017:ZMF

[CJ17] Davide Cortinovis and Patrick Jenny. Zonal Multiscale Finite-Volume framework. Journal of Computational Physics, 337 (??):84–97, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300682.

Cortesi:2019:KSP

[CJC19] Andrea F. Cortesi, Ghina Jannoun, and Pietro M. Congedo. Kriging-sparse Polynomial Dimensional Decomposition surrogate model with adaptive refinement. Journal of Computational Physics, 380(??):212–242, March 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307186.

Chaudhuri:2017:EDS

[CJD^{+}17] A. Chaudhuri, G. B. Jacobs, W. S. Don, H. Abbassi, and F. Mashayek. Explicit discontinuous spectral element method with entropy generation based artificial viscosity for shocked viscous flows. Journal of Computational Physics, 332(??):99–117, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306349.

Colas:2019:PML

[CJH^{+}19] L. Colas, J. Jacquot, J. Hillairet, W. Helou, W. Tierens, S. Heuraux, E. Faudot, L. Lu, and G. Urbanczyk. Perfectly Matched Layers for time-harmonic transverse electric wave propagation in cylindrical and toroidal gyrotropic media. Journal of Computational Physics, 389(??):94–110, July 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301342.

Carlberg:2019:RMC

[CJK^{+}19] Kevin T. Carlberg, Antony Jameson, Mykel J. Kochenderfer, Jeremy Morton, Liqian Peng, and Freddie D. Witherden. Recovering missing CFD data for high-order discretizations using deep neural networks and dynamics learning. Journal of Computational Physics, 395(??):105–124, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119303857.

Chen:2016:HAH

[CJL16] Yibing Chen, Song Jiang, and Na Liu. HFVS: an arbitrary high order approach based on flux vector splitting. Journal of Computational Physics, 322(??):708–722, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302832.

Cai:2019:SPA

[CJWS19] Wenjun Cai, Chaolong Jiang, Yushun Wang, and Yongzhong Song. Structure-preserving algorithms for the two-dimensional sine-Gordon equation with Neumann boundary conditions. Journal of Computational Physics, 395(??):166–185, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304000.

Chen:2015:DES

[CJYZ15] Rui Chen, Guanghua Ji, Xiaofeng Yang, and Hui Zhang. Decoupled energy stable schemes for phase-field vesicle membrane model. Journal of Computational Physics, 302(??):509–523, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500618X.

Conroy:2016:HDG

[CK16a] Colton J. Conroy and Ethan J. Kubatko. hp discontinuous Galerkin methods for the vertical extent of the water column in coastal settings. Part I: Barotropic forcing. Journal of Computational Physics, 305(??):1147–1171, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007159.

Cotter:2016:EDG

[CK16b] C. J. Cotter and D. Kuzmin. Embedded discontinuous Galerkin transport schemes with localised limiters. Journal of Computational Physics, 311(??):363–373, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000759.

Carvalho:2018:AAC

[CKK18a] Camille Carvalho, Shilpa Khatri, and Arnold D. Kim. Asymptotic analysis for close evaluation of layer potentials. Journal of Computational Physics, 355(??):327–341, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308537.

Chamarthi:2018:HOU

[CKK18b] Amareshwara Sainadh Chamarthi, Kimiya Komurasaki, and Rei Kawashima. High-order upwind and non-oscillatory approach for steady state diffusion, advection-diffusion and application to magnetized electrons. Journal of Computational Physics, 374(??):1120–1151, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305412.

Cai:2019:RBS

[CKKD19] X.-X. Cai, T. Kittelmann, E. Klinkby, and J. I. Márquez Damián. Rejection-based sampling of inelastic neutron scattering. Journal of Computational Physics, 380(??):400–407, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307885.

Cheng:2015:FSE

[CKQT15] Yuanzhen Cheng, Alexander Kurganov, Zhuolin Qu, and Tao Tang. Fast and stable explicit operator splitting methods for phase-field models. Journal of Computational Physics, 303(??):45–65, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005859.

Chen:2017:SOA

[CKT17] Yaping Chen, Yangyu Kuang, and Huazhong Tang. Secondorder accurate genuine BGK schemes for the ultra-relativistic flow simulations. Journal of Computational Physics, 349(??): 300–327, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117305934.

Chapelier:2016:SED

[CL16] J.-B. Chapelier and G. Lodato. A spectral-element dynamic model for the Large-Eddy simulation of turbulent flows. Journal of Computational Physics, 321(??):279–302, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301966.

Chen:2017:SMM

[CL17] Xinjuan Chen and Jinglai Li. A subset multicanonical Monte Carlo method for simulating rare failure events. Journal of Computational Physics, 344(??):23–35, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303303.

Cao:2018:SDL

[CL18] Yu Cao and Jianfeng Lu. Stochastic dynamical lowrank approximation method. Journal of Computational Physics, 372(??):564–586, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304364.

Cheng:2019:VRD

[CL19a] Jian Cheng and Tiegang Liu. A variational reconstructed discontinuous Galerkin method for the steady-state compressible flows on unstructured grids. Journal of Computational Physics, 380(??):65–87, March 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830785X.

Costa:2019:APN

[CL19b] R. P. Costa, Jr. and M. A. Leigui de Oliveira. Analysis of the performance of numerical integration methods for the tracking of ultra-high energy cosmic rays. Journal of Computational Physics, 392(??):432–443, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303110.

Christon:2016:LES

[CLB^{+}16] Mark A. Christon, Roger Lu, Jozsef Bakosi, Balasubramanya T. Nadiga, Zeses Karoutas, and Markus Berndt. Large-eddy simulation, fuel rod vibration and grid-to-rod fretting in pressurized water reactors. Journal of Computational Physics, 322(??):142–161, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302674.

Chen:2016:FDS

[CLC16] Hu Chen, Shujuan Lü, and Wenping Chen. Finite difference/spectral approximations for the distributed order time fractional reaction-diffusion equation on an unbounded domain. Journal of Computational Physics, 315(??):84–97, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001960.

Cui:2017:TPS

[CLFL17] Xiangyang Cui, She Li, Hui Feng, and Guangyao Li. A triangular prism solid and shell interactive mapping element for electromagnetic sheet metal forming process. Journal of Computational Physics, 336(??):192–211, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301079.

Ching:2019:SCD

[CLG^{+}19] Eric J. Ching, Yu Lv, Peter Gnoffo, Michael Barnhardt, and Matthias Ihme. Shock capturing for discontinuous Galerkin methods with application to predicting heat transfer in hypersonic flows. Journal of Computational Physics, 376(??):54–75, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306107.

Citro:2017:ESA

[CLGA17] V. Citro, P. Luchini, F. Giannetti, and F. Auteri. Efficient stabilization and acceleration of numerical simulation of fluid flows by residual recombination. Journal of Computational Physics, 344(??):234–246, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303698.

Choi:2017:RTU

[CLL17] Sooyoung Choi, Changho Lee, and Deokjung Lee. Resonance treatment using pin-based pointwise energy slowing-down method. Journal of Computational Physics, 330(??):134–155, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305927.

Chen:2019:NMC

[CLL19] Hongxu Chen, Qin Li, and Jianfeng Lu. A numerical method for coupling the BGK model and Euler equations through the linearized Knudsen layer. Journal of Computational Physics, 398(??):Article 108893, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305911.

Connington:2015:IFI

[CLM15] Kevin W. Connington, Taehun Lee, and Jeffrey F. Morris. Interaction of fluid interfaces with immersed solid particles using the lattice Boltzmann method for liquid-gas-particle systems. Journal of Computational Physics, 283(??):453–477, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008079.

Cui:2016:DIL

[CLM16] Tiangang Cui, Kody J. H. Law, and Youssef M. Marzouk. Dimension-independent likelihood-informed MCMC. Journal of Computational Physics, 304(??):109–137, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006701.

Crouseilles:2017:UAP

[CLMZ17] Nicolas Crouseilles, Mohammed Lemou, Florian Méhats, and Xiaofei Zhao. Uniformly accurate particle-in-cell method for the long time solution of the two-dimensional Vlasov– Poisson equation with uniform strong magnetic field. Journal of Computational Physics, 346(??):172–190, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304564.

Cusini:2015:CPR

[CLNH15] Matteo Cusini, Alexander A. Lukyanov, Jostein Natvig, and Hadi Hajibeygi. Constrained pressure residual multiscale (CPR-MS) method for fully implicit simulation of multiphase flow in porous media. Journal of Computational Physics, 299(??):472–486, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004647.

Cibotarica:2016:SNT

[CLP16a] Alexandru Cibotarica, James V. Lambers, and Elisabeth M. Palchak. Solution of nonlinear time-dependent PDEs through componentwise approximation of matrix functions. Journal of Computational Physics, 321(??):1120–1143, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302492.

Cox:2016:HOS

[CLP16b] Christopher Cox, Chunlei Liang, and Michael W. Plesniak. A high-order solver for unsteady incompressible Navier–Stokes equations using the flux reconstruction method on unstructured grids with implicit dual time stepping. Journal of Computational Physics, 314(??):414–435, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001686.

Chung:2017:RBI

[CLQ17] Eric T. Chung, Chi Yeung Lam, and Jianliang Qian. A raybased IPDG method for high-frequency time-domain acoustic wave propagation in inhomogeneous media. Journal of Computational Physics, 348(??):660–682, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305569.

Cheung:2015:LMM

[CLR15] Ka Chun Cheung, Leevan Ling, and Steven J. Ruuth. A localized meshless method for diffusion on folded surfaces. Journal of Computational Physics, 297(??):194–206, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003551.

Chen:2018:EES

[CLS^{+}18] Ying Chen, John Lowengrub, Jie Shen, Cheng Wang, and Steven Wise. Efficient energy stable schemes for isotropic and strongly anisotropic Cahn–Hilliard systems with the Willmore regularization. Journal of Computational Physics, 365(??):56– 73, July 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301840.

Christlieb:2015:HOP

[CLTX15] Andrew J. Christlieb, Yuan Liu, Qi Tang, and Zhengfu Xu. High order parametrized maximum-principle-preserving and positivity-preserving WENO schemes on unstructured meshes. Journal of Computational Physics, 281(??):334–351, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007116.

Colle:2019:AMR

[CLV19] A. Collé, J. Limido, and J.-P. Vila. An accurate multiregime SPH scheme for barotropic flows. Journal of Computational Physics, 388(??):561–600, July 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302116.

Cheng:2017:AST

[CLvS17] Gong Cheng, Per Lötstedt, and Lina von Sydow. Accurate and stable time stepping in ice sheet modeling. Journal of Computational Physics, 329(??):29–47, January 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630568X.

Cai:2018:PAV

[CLW18] Wenjun Cai, Haochen Li, and Yushun Wang. Partitioned averaged vector field methods. Journal of Computational Physics, 370(??):25–42, October 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303012.

Christlieb:2015:HOO

[CLX15] Andrew J. Christlieb, Yuan Liu, and Zhengfu Xu. High order operator splitting methods based on an integral deferred correction framework. Journal of Computational Physics, 294(??):224–242, August 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001795.

Chou:2019:ECG

[CLX19] Ching-Shan Chou, Yukun Li, and Dongbin Xiu. Energy conserving Galerkin approximation of two dimensional wave equations with random coefficients. Journal of Computational Physics, 381(??):52–66, March 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118308143.

Chen:2015:AFQ

[CLY^{+}15] Quan Chen, Jun Li, Chiyung Yam, Yu Zhang, Ngai Wong, and Guanhua Chen. An approximate framework for quantum transport calculation with model order reduction. Journal of Computational Physics, 286(??):49–61, April 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000364.

Chen:2018:KPS

[CLZ18] Hao Chen, Wen Lv, and Tongtong Zhang. A Kronecker product splitting preconditioner for two-dimensional space-fractional diffusion equations. Journal of Computational Physics, 360(??):1–14, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300445.

Chen:2019:EED

[CLZZ19] Jingrun Chen, Ling Lin, Zhiwen Zhang, and Xiang Zhou. Estimation of exciton diffusion lengths of organic semiconductors in random domains. Journal of Computational Physics, 376(??):894–912, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830665X.

Chandrasekaran:2015:MSN

[CM15] S. Chandrasekaran and H. N. Mhaskar. A minimum Sobolev norm technique for the numerical discretization of PDEs. Journal of Computational Physics, 299(??):649–666, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004702.

Cerroni:2016:PPA

[CM16a] D. Cerroni and S. Manservisi. A penalty-projection algorithm for a monolithic fluid-structure interaction solver. Journal of Computational Physics, 313(??):13–30, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911600111X.

Cottet:2016:SIL

[CM16b] Georges-Henri Cottet and Emmanuel Maitre. A semiimplicit level set method for multiphase flows and fluidstructure interaction problems. Journal of Computational Physics, 314(??):80–92, June 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001546.

Campos:2018:EAB

[CM18a] A. Campos and B. Morgan. The effect of artificial bulk viscosity in simulations of forced compressible turbulence. Journal of Computational Physics, 371(??):111–121, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830336X.

Chen:2018:ESA

[CM18b] Nan Chen and Andrew J. Majda. Efficient statistically accurate algorithms for the Fokker–Planck equation in large dimensions. Journal of Computational Physics, 354(??):242–268, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307751.

Corot:2018:NNS

[CM18c] T. Corot and B. Mercier. A new nodal solver for the two dimensional Lagrangian hydrodynamics. Journal of Computational Physics, 353(??):1–25, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730726X.

Correa:2018:NSM

[CM18d] Maicon R. Correa and Marcio A. Murad. A new sequential method for three-phase immiscible flow in poroelastic media. Journal of Computational Physics, 373(??):493–532, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304479.

Chen:2019:NEP

[CM19] Nan Chen and Andrew J. Majda. A new efficient parameter estimation algorithm for high-dimensional complex nonlinear turbulent dynamical systems with partial observations. Journal of Computational Physics, 397(??):Article 108836, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305200.

Chiron:2018:CSF

[CMDL18] L. Chiron, S. Marrone, A. Di Mascio, and D. Le Touzé. Coupled SPH–FV method with net vorticity and mass transfer. Journal of Computational Physics, 364(??):111–136, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301438.

Cimpeanu:2015:PFP

[CMH15] Radu Cimpeanu, Anton Martinsson, and Matthias Heil. A parameter-free perfectly matched layer formulation for the finite-element-based solution of the Helmholtz equation. Journal of Computational Physics, 296(??):329–347, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003241.

Chakir:2019:NIR

[CMP19] R. Chakir, Y. Maday, and P. Parnaudeau. A non-intrusive reduced basis approach for parametrized heat transfer problems. Journal of Computational Physics, 376(??):617–633, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306570.

Curtis:2016:ASS

[CMR^{+}16] J. H. Curtis, M. D. Michelotti, N. Riemer, M. T. Heath, and M. West. Accelerated simulation of stochastic particle removal processes in particle-resolved aerosol models. Journal of Computational Physics, 322(??):21–32, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302546.

Cui:2016:SPA

[CMW16] Tiangang Cui, Youssef Marzouk, and Karen Willcox. Scalable posterior approximations for large-scale Bayesian inverse problems via likelihood-informed parameter and state reduction. Journal of Computational Physics, 315(??):363–387, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300183.

Chowdhary:2016:BEK

[CN16] Kenny Chowdhary and Habib N. Najm. Bayesian estimation of Karhunen–Loève expansions; a random subspace approach. Journal of Computational Physics, 319(??):280–293, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001285.

Carpenter:1999:SCI

[CNG99] Mark H. Carpenter, Jan Nordström, and David Gottlieb. A stable and conservative interface treatment of arbitrary spatial accuracy. Journal of Computational Physics, 148(2):341–365, January 20, 1999. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999198961149. See corri-

gendum [CNG17].

Carpenter:2017:CSC

[CNG17] Mark H. Carpenter, Jan Nordström, and David Gottlieb. Corrigendum to “A stable and conservative interface treatment of arbitrary spatial accuracy” [J. Comput. Phys. 148 (1999) 341– 365]. Journal of Computational Physics, 351(??):534, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307179. See [CNG99].

Chamarthi:2019:FOH

[CNK19] Amareshwara Sainadh Chamarthi, Hiroaki Nishikawa, and Kimiya Komurasaki. First order hyperbolic approach for anisotropic diffusion equation. Journal of Computational Physics, 396(??):243–263, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119304723.

Chen:2015:PAF

[CNOS15] Long Chen, Ricardo H. Nochetto, Enrique Otárola, and Abner J. Salgado. A PDE approach to fractional diffusion: a posteriori error analysis. Journal of Computational Physics, 293(??):339–358, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000029.

Cheng:2019:ESC

[CNQ^{+}19] Ming Cheng, Akil Narayan, Yi Qin, Peng Wang, Xinghui Zhong, and Xueyu Zhu. An efficient solver for cumulative density function-based solutions of uncertain kinematic wave models. Journal of Computational Physics, 382(??):138–151, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300257.

Chater:2017:SLS

[CNW17] Mario Chater, Angxiu Ni, and Qiqi Wang. Simplified Least Squares Shadowing sensitivity analysis for chaotic ODEs and PDEs. Journal of Computational Physics, 329(??):126–140, January 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305332.

Chiron:2018:AIA

[COdLL18] L. Chiron, G. Oger, M. de Leffe, and D. Le Touzé. Analysis and improvements of Adaptive Particle Refinement (APR) through CPU time, accuracy and robustness considerations. Journal of Computational Physics, 354(??):552–575, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308082.

Cortez:2018:RSS

[Cor18] Ricardo Cortez. Regularized Stokeslet segments. Journal of Computational Physics, 375(??):783–796, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305886.

Costa:2016:MTM

[Cos16] Liborio I. Costa. Meaningful timescales from Monte Carlo simulations of particle systems with hard-core interactions. Journal of Computational Physics, 326(??):773–779, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304375.

Cotter:2016:CAE

[Cot16] Simon L. Cotter. Constrained approximation of effective generators for multiscale stochastic reaction networks and application to conditioned path sampling. Journal of Computational Physics, 323(??):265–282, October 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303308.

Cottet:2018:SLP

[Cot18] Georges-Henri Cottet. Semi-Lagrangian particle methods for high-dimensional Vlasov–Poisson systems. Journal of Computational Physics, 366(??):362–375, July 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302080.

Chernyshenko:2018:HFV

[COV18] Alexey Y. Chernyshenko, Maxim A. Olshanskii, and Yuri V. Vassilevski. A hybrid finite volume — finite element method for bulk-surface coupled problems. Journal of Computational Physics, 352(??):516–533, January 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307374.

Chen:2016:NDF

[CP16] Wen Chen and Guofei Pang. A new definition of fractional Laplacian with application to modeling three-dimensional nonlocal heat conduction. Journal of Computational Physics, 309(??):350–367, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116000048.

Celik:2017:PDE

[CP17] Ismail B. Celik and Don Roscoe Parsons III. Prediction of discretization error using the error transport equation. Journal of Computational Physics, 339(??):96–125, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301614.

Carpio:2019:AAL

[CPdS19] Jaime Carpio, Juan Luis Prieto, and Pedro Galán del Sastre. An anisotropic adaptive, Lagrange–Galerkin numerical method for spray combustion. Journal of Computational Physics, 381(??):246–274, March 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300038.

Chung:2019:OBC

[CPP19] Eric Chung, Sara Pollock, and Sai-Mang Pun. Online basis construction for goal-oriented adaptivity in the generalized multiscale finite element method. Journal of Computational Physics, 393(??):59–73, September 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303341.

Colera:2017:EFD

[CPS17] Manuel Colera and Miguel Pérez-Saborid. An efficient finite differences method for the computation of compressible, subsonic, unsteady flows past airfoils and panels. Journal of Computational Physics, 345(??):596–617, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304291.

Chekhovskoy:2017:NAS

[CPSF17] I. S. Chekhovskoy, V. I. Paasonen, O. V. Shtyrina, and M. P. Fedoruk. Numerical approaches to simulation of multi-core fibers. Journal of Computational Physics, 334(??):31–44, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116307240.

Carr:2016:EDM

[CPT16] E. J. Carr, P. Perré, and I. W. Turner. The extended distributed microstructure model for gradient-driven transport: a two-scale model for bypassing effective parameters. Journal of Computational Physics, 327(??):810–829, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304971.

Carpio:2016:LAA

[CPV16] Jaime Carpio, Juan Luis Prieto, and Marcos Vera. A local anisotropic adaptive algorithm for the solution of low-Mach transient combustion problems. Journal of Computational Physics, 306(??):19–42, February 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007512.

Canuto:2019:UQD

[CPX19] Claudio Canuto, Sandra Pieraccini, and Dongbin Xiu. Uncertainty quantification of discontinuous outputs via a nonintrusive bifidelity strategy. Journal of Computational Physics, 398(??):Article 108885, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305832.

Chen:2015:NAH

[CQ15] Peng Chen and Alfio Quarteroni. A new algorithm for highdimensional uncertainty quantification based on dimensionadaptive sparse grid approximation and reduced basis methods. Journal of Computational Physics, 298(??):176–193, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003861.

Chen:2017:CSS

[CQL^{+}17] Qiang Chen, Hong Qin, Jian Liu, Jianyuan Xiao, Ruili Zhang, Yang He, and Yulei Wang. Canonical symplectic structure and structure-preserving geometric algorithms for Schrödinger–Maxwell systems. Journal of Computational Physics, 349(??):441–452, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306046.

Cai:2016:CSL

[CQQ16] Xiaofeng Cai, Jianxian Qiu, and Jing-Mei Qiu. A conservative semi-Lagrangian HWENO method for the Vlasov equation. Journal of Computational Physics, 323(??):95–114, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303163.

Cakoni:2017:DIS

[CR17] Fioralba Cakoni and Jacob D. Rezac. Direct imaging of small scatterers using reduced time dependent data. Journal of Computational Physics, 338(??):371–387, June 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730164X.

Coco:2018:SOF

[CR18] Armando Coco and Giovanni Russo. Second order finitedifference ghost-point multigrid methods for elliptic problems with discontinuous coefficients on an arbitrary interface. Journal of Computational Physics, 361(??):299–330, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300263.

Chicco-Ruiz:2016:APM

[CRMP16] Aníbal Chicco-Ruiz, Pedro Morin, and M. Sebastian Pauletti. An algorithm for prescribed mean curvature using isogeometric methods. Journal of Computational Physics, 317(??):185–203, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300444.

Carrillo:2016:NSN

[CRW16] José A. Carrillo, Helene Ranetbauer, and Marie-Therese Wolfram. Numerical simulation of nonlinear continuity equations by evolving diffeomorphisms. Journal of Computational Physics, 327(??):186–202, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304612.

Corona:2017:TTA

[CRZ17] Eduardo Corona, Abtin Rahimian, and Denis Zorin. A tensortrain accelerated solver for integral equations in complex geometries. Journal of Computational Physics, 334(??):145–169, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116307185.

Chacon:2016:SFI

[CS16a] L. Chacón and A. Stanier. A scalable, fully implicit algorithm for the reduced two-field low-β extended MHD model. Journal of Computational Physics, 326(??):763–772, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304193.

Chen:2016:SGR

[CS16b] Peng Chen and Christoph Schwab. Sparse-grid, reducedbasis Bayesian inversion: Nonaffine-parametric nonlinear equations. Journal of Computational Physics, 316(??):470–503, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001273.

Chen:2016:EAE

[CS16c] Ying Chen and Jie Shen. Efficient, adaptive energy stable schemes for the incompressible Cahn–Hilliard Navier– Stokes phase-field models. Journal of Computational Physics, 308(??):40–56, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115008219.

Chen:2017:ESH

[CS17a] Tianheng Chen and Chi-Wang Shu. Entropy stable high order discontinuous Galerkin methods with suitable quadrature rules for hyperbolic conservation laws. Journal of Computational Physics, 345(??):427–461, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730400X.

Cogswell:2017:SIT

[CS17b] Daniel A. Cogswell and Michael L. Szulczewski. Simulation of incompressible two-phase flow in porous media with large timesteps. Journal of Computational Physics, 345(??):856–865, September 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304527.

Cheng:2018:ARB

[CS18a] Gong Cheng and Victor Shcherbakov. Anisotropic radial basis function methods for continental size ice sheet simulations. Journal of Computational Physics, 372(??):161–177, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911830398X.

Chiapolino:2018:MMT

[CS18b] Alexandre Chiapolino and Richard Saurel. Models and methods for two-layer shallow water flows. Journal of Computational Physics, 371(??):1043–1066, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303425.

Conrad:2015:ANN

[CSB15] Daniel Conrad, Andreas Schneider, and Martin Böhle. Accuracy of non-Newtonian lattice Boltzmann simulations. Journal of Computational Physics, 301(??):218–229, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005288.

Chen:2019:HOE

[CSC19] Zhenzhu Chen, Sihong Shao, and Wei Cai. A high order efficient numerical method for 4-D Wigner equation of quantum double-slit interferences. Journal of Computational Physics, 396(??):54–71, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304553.

Cipcigan:2016:ECG

[CSCM16] Flaviu S. Cipcigan, Vlad P. Sokhan, Jason Crain, and Glenn J. Martyna. Electronic coarse graining enhances the predictive power of molecular simulation allowing challenges in water physics to be addressed. Journal of Computational Physics, 326(??):222–233, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303874.

Castiglioni:2019:EAD

[CSD19] Giacomo Castiglioni, Guangrui Sun, and J. Andrzej Domaradzki. On the estimation of artificial dissipation and dispersion errors in a generic partial differential equation. Journal of Computational Physics, 397(??):Article 108843, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305273.

Cherry:2017:NMP

[CSG17] Matthew R. Cherry, Shamachary Sathish, and Ramana Grandhi. A numerical method for predicting Rayleigh surface wave velocity in anisotropic crystals. Journal of Computational Physics, 351(??):108–120, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306526.

Comminal:2015:CCU

[CSH15] Raphaël Comminal, Jon Spangenberg, and Jesper Henri Hattel. Cellwise conservative unsplit advection for the volume of fluid method. Journal of Computational Physics, 283 (??):582–608, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114008109.

Collins:2016:SAN

[CSK^{+}16] Benjamin Collins, Shane Stimpson, Blake W. Kelley, Mitchell T. H. Young, Brendan Kochunas, Aaron Graham, Edward W. Larsen, Thomas Downar, and Andrew Godfrey. Stability and accuracy of 3D neutron transport simulations using the 2D/1D method in MPACT. Journal of Computational Physics, 326(??):612–628, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303710.

Choi:2015:IIT

[CSLL15] Sooyoung Choi, Kord Smith, Hyun Chul Lee, and Deokjung Lee. Impact of inflow transport approximation on light water reactor analysis. Journal of Computational Physics, 299(??):352–373, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004507.

Chiapolino:2017:SDI

[CSN17] Alexandre Chiapolino, Richard Saurel, and Boniface Nkonga. Sharpening diffuse interfaces with compressible fluids on unstructured meshes. Journal of Computational Physics, 340 (??):389–417, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302371.

Chowdhary:2018:ESM

[CSN18] Kenny Chowdhary, Cosmin Safta, and Habib N. Najm. Enhancing statistical moment calculations for stochastic Galerkin solutions with Monte Carlo techniques. Journal of Computational Physics, 374(??):1017–1030, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304595.

Campos:2015:ECG

[CSS15] Cédric M. Campos and J. M. Sanz-Serna. Extra Chance Generalized Hybrid Monte Carlo. Journal of Computational Physics, 281(??):365–374, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006731.

Campos:2017:PSS

[CSS17] Cédric M. Campos and J. M. Sanz-Serna. Palindromic 3stage splitting integrators, a roadmap. Journal of Computational Physics, 346(??):340–355, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304515.

Cao:2019:CVF

[CSW^{+}19] Zhizhu Cao, Dongliang Sun, Jinjia Wei, Bo Yu, and Jingfa Li. A coupled volume-of-fluid and level set method based on general curvilinear grids with accurate surface tension calculation. Journal of Computational Physics, 396(??):799–818, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119305017.

Choi:2015:MQB

[CSY15] Sungjin Choi, D. Scott Stewart, and Sunhee Yoo. Modeling of the quenching of blast products from energetic materials by expansion into vacuum. Journal of Computational Physics, 296(??):158–183, September 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002399.

Chang:2019:ITP

[CSY19] Lina Chang, Zhiqiang Sheng, and Guangwei Yuan. An improvement of the two-point flux approximation scheme on polygonal meshes. Journal of Computational Physics, 392 (??):187–204, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119302918.

Cai:2015:ALB

[CT15] Zhenning Cai and Manuel Torrilhon. Approximation of the linearized Boltzmann collision operator for hard-sphere and inverse-power-law models. Journal of Computational Physics, 295(??):617–643, August 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115002879.

Chen:2019:GLS

[CT19] Yuxi Chen and Gábor Tóth. Gauss’s Law satisfying Energy-Conserving Semi-Implicit Particle-in-Cell method. Journal of Computational Physics, 386(??):632–652, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301494.

Chen:2016:FOF

[CTG16] Yuxi Chen, Gábor Tóth, and Tamas I. Gombosi. A fifth-order finite difference scheme for hyperbolic equations on block-adaptive curvilinear grids. Journal of Computational Physics, 305(??):604–621, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007378.

Charin:2017:MMI

[CTJ^{+}17] A. H. L. M. Charin, Z. Tuković, H. Jasak, L. F. L. R. Silva, and P. L. C. Lage. A moving mesh interface tracking method for simulation of liquid-liquid systems. Journal of Computational Physics, 334(??):419–441, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300219.

Contarino:2016:JGR

[CTM^{+}16] Christian Contarino, Eleuterio F. Toro, Gino I. Montecinos, Raul Borsche, and Jochen Kall. Junction-Generalized Riemann Problem for stiff hyperbolic balance laws in networks: an implicit solver and ADER schemes. Journal of Computational Physics, 315(??):409–433, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116002011.

Cui:2015:CES

[Cui15] Mingrong Cui. Compact exponential scheme for the time fractional convection-diffusion reaction equation with variable coefficients. Journal of Computational Physics, 280(??):143–163, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006482.

Cortez:2015:GSI

[CV15] Ricardo Cortez and Douglas Varela. A general system of images for regularized Stokeslets and other elements near a plane wall. Journal of Computational Physics, 285(??):41–54, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000236.

Churbanov:2016:NIS

[CV16a] Alexander G. Churbanov and Petr N. Vabishchevich. Numerical investigation of a space-fractional model of turbulent fluid flow in rectangular ducts. Journal of Computational Physics, 321(??):846–859, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302340.

Cools:2016:FRC

[CV16b] S. Cools and W. Vanroose. A fast and robust computational method for the ionization cross sections of the driven Schrödinger equation using an O(N) multigrid-based scheme. Journal of Computational Physics, 308(??):20–39, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008384.

Caleffi:2017:WBS

[CV17] Valerio Caleffi and Alessandro Valiani. Well balancing of the SWE schemes for moving-water steady flows. Journal of Computational Physics, 342(??):85–116, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303029.

Corona:2018:BIE

[CV18] Eduardo Corona and Shravan Veerapaneni. Boundary integral equation analysis for suspension of spheres in Stokes flow. Journal of Computational Physics, 362(??):327–345, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300998.

Clarke:2018:LND

[CVG18] Peter Clarke, Philip Varghese, and David Goldstein. A low noise discrete velocity method for the Boltzmann equation with quantized rotational and vibrational energy. Journal of Computational Physics, 352(??):326–340, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306460.

Chen:2019:TAV

[CVG19] Peng Chen, Umberto Villa, and Omar Ghattas. Taylor approximation and variance reduction for PDE-constrained optimal control under uncertainty. Journal of Computational Physics, 385(??):163–186, May 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301056.

Cho:2016:NMH

[CVK16] H. Cho, D. Venturi, and G. E. Karniadakis. Numerical methods for high-dimensional probability density function equations. Journal of Computational Physics, 305(??):817–837, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007068.

Cusini:2016:ADM

[CvKH16] Matteo Cusini, Cor van Kruijsdijk, and Hadi Hajibeygi. Algebraic dynamic multilevel (ADM) method for fully implicit simulations of multiphase flow in porous media. Journal of Computational Physics, 314(??):60–79, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001583.

Collia:2019:SMV

[CVM^{+}19] Dario Collia, Marija Vukicevic, Valentina Meschini, Luigino Zovatto, and Gianni Pedrizzetti. Simplified mitral valve modeling for prospective clinical application of left ventricular fluid dynamics. Journal of Computational Physics, 398(??):Article 108895, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305935.

Charnley:2016:TWR

[CW16] Matthew Charnley and Aihua Wood. Through-the-wall radar detection analysis via numerical modeling of Maxwell’s equations. Journal of Computational Physics, 313(??):532–548, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001170.

Charnley:2017:LSM

[CW17] Matthew Charnley and Aihua Wood. A linear sampling method for through-the-wall radar detection. Journal of Computational Physics, 347(??):147–159, October 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304886.

Cleveland:2018:CIM

[CW18] M. A. Cleveland and A. B. Wollaber. Corrected implicit Monte Carlo. Journal of Computational Physics, 359(??):20– 44, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117309324.

Chan:2019:DES

[CW19] Jesse Chan and Lucas C. Wilcox. On discretely entropy stable weight-adjusted discontinuous Galerkin methods: curvilinear meshes. Journal of Computational Physics, 378(??):366–393, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307289.

Chen:2019:ACM

[CWB^{+}19] Feng Chen, Jiawei Wang, Can Ba, Xubin Bai, Tianyu Dong, and Xikui Ma. Adaptive cell method with constitutive matrix correction for simulating physical field on a coarse grid. Journal of Computational Physics, 397(??):Article 108820, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305042.

Castelletto:2016:SAT

[CWF16] Nicola Castelletto, Joshua A. White, and Massimiliano Ferronato. Scalable algorithms for three-field mixed finite element coupled poromechanics. Journal of Computational Physics, 327(??):894–918, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304843.

Crabill:2018:PDC

[CWJ18] J. Crabill, F. D. Witherden, and A. Jameson. A parallel direct cut algorithm for high-order overset methods with application to a spinning golf ball. Journal of Computational Physics, 374(??):692–723, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303450.

Chen:2016:VOF

[CWL^{+}16] Yi-Ming Chen, Yan-Qiao Wei, Da-Yan Liu, Driss Boutat, and Xiu-Kai Chen. Variable-order fractional numerical differentiation for noisy signals by wavelet denoising. Journal of Computational Physics, 311(??):338–347, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911600067X.

Chan:2016:GAD

[CWM^{+}16] Jesse Chan, Zheng Wang, Axel Modave, Jean-François Remacle, and T. Warburton. GPU-accelerated discontinuous Galerkin methods on hybrid meshes. Journal of Computational Physics, 318(??):142–168, August 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300353.

Chapelier:2018:CVP

[CWS18] J.-B. Chapelier, B. Wasistho, and C. Scalo. A coherent vorticity preserving eddy-viscosity correction for Large-Eddy Simulation. Journal of Computational Physics, 359(??):164–182, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300226.

Chen:2017:IFM

[CWW17] Long Chen, Huayi Wei, and Min Wen. An interface-fitted mesh generator and virtual element methods for elliptic interface problems. Journal of Computational Physics, 334(??):327–348, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300141.

Chen:2017:TOH

[CWWZ17] Feifei Chen, Yiqiang Wang, Michael Yu Wang, and Y. F. Zhang. Topology optimization of hyperelastic structures using a level set method. Journal of Computational Physics, 351(??):437–454, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117307003.

Cui:2016:AAD

[CwYjS16] Xia Cui, Guang wei Yuan, and Zhi jun Shen. Asymptotic analysis of discrete schemes for non-equilibrium radiation diffusion. Journal of Computational Physics, 313(??):415–429, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001339.

Chen:2015:CSA

[CX15] Songze Chen and Kun Xu. A comparative study of an asymptotic preserving scheme and unified gas-kinetic scheme in continuum flow limit. Journal of Computational Physics, 288(??):52–65, May 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000844.

Chen:2016:VVP

[CX16] Yu Chen and Xilin Xie. Vorticity vector-potential method for 3D viscous incompressible flows in time-dependent curvilinear coordinates. Journal of Computational Physics, 312(??):50– 81, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000747.

Chen:2015:MDS

[CXH15] Feng Chen, Qinwu Xu, and Jan S. Hesthaven. A multidomain spectral method for time-fractional differential equations. Journal of Computational Physics, 293(??):157–172, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006986.

Chen:2016:CGM

[CXL16] Songze Chen, Kun Xu, and Zhihui Li. Cartesian grid method for gas kinetic scheme on irregular geometries. Journal of Computational Physics, 326(??):862–877, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304326.

Chen:2016:FMM

[CXX16] Hongtao Chen, Hehu Xie, and Fei Xu. A full multigrid method for eigenvalue problems. Journal of Computational Physics, 322(??):747–759, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302881.

Chuenjarern:2019:HOB

[CXY19] Nattaporn Chuenjarern, Ziyao Xu, and Yang Yang. High-order bound-preserving discontinuous Galerkin methods for compressible miscible displacements in porous media on triangular meshes. Journal of Computational Physics, 378(??):110–128, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307216.

Chen:2019:ENS

[CY19a] Chuanjun Chen and Xiaofeng Yang. Efficient numerical scheme for a dendritic solidification phase field model with melt convection. Journal of Computational Physics, 388(??):41–62, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302001.

Choi:2019:NZA

[CY19b] Jeong-Ok Choi and Unjong Yu. Newman–Ziff algorithm for the bootstrap percolation: Application to the Archimedean lattices. Journal of Computational Physics, 386(??):1–8, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301147.

Cheng:2016:DDG

[CYL^{+}16] Jian Cheng, Xiaoquan Yang, Xiaodong Liu, Tiegang Liu, and Hong Luo. A direct discontinuous Galerkin method for the compressible Navier–Stokes equations on arbitrary grids. Journal of Computational Physics, 327(??):484–502, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304703.

Cheng:2017:EAN

[CYS17] Qing Cheng, Xiaofeng Yang, and Jie Shen. Efficient and accurate numerical schemes for a hydro-dynamically coupled phase field diblock copolymer model. Journal of Computational Physics, 341(??):44–60, July 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302814.

Cui:2017:HWB

[CYWL17] Xiongwei Cui, Xiongliang Yao, Zhikai Wang, and Minghao Liu. A hybrid wavelet-based adaptive immersed boundary finite-difference lattice Boltzmann method for twodimensional fluid–structure interaction. Journal of Computational Physics, 333(??):24–48, March 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306726.

Cheng:2018:ADA

[CYYL18] Jian Cheng, Huiqiang Yue, Shengjiao Yu, and Tiegang Liu. Analysis and development of adjoint-based h-adaptive direct discontinuous Galerkin method for the compressible Navier–Stokes equations. Journal of Computational Physics, 362(??):305–326, June 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830113X.

Chen:2016:IMM

[CZ16] Xiang Chen and Xiong Zhang. An improved 3D MoF method based on analytical partial derivatives. Journal of Computational Physics, 326(??):156–170, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304089.

Chen:2017:IMM

[CZ17] Xiang Chen and Xiong Zhang. An improved 2D MoF method by using high order derivatives. Journal of Computational Physics, 349(??):176–190, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306022.

Chang:2019:IPP

[CZ19a] Haibin Chang and Dongxiao Zhang. Identification of physical processes via combined data-driven and dataassimilation methods. Journal of Computational Physics, 393(??):337–350, September 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930333X.

Chen:2019:PNM

[CZ19b] Xiang Chen and Xiong Zhang. A predicted-Newton’s method for solving the interface positioning equation in the MoF method on general polyhedrons. Journal of Computational Physics, 384(??):60–76, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300877.

Chen:2015:UPU

[CZB15] Peng Chen, Nicholas Zabaras, and Ilias Bilionis. Uncertainty propagation using infinite mixture of Gaussian processes and variational Bayesian inference. Journal of Computational Physics, 284(??):291–333, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008456.

Casquero:2018:NBF

[CZBC^{+}18] Hugo Casquero, Yongjie Jessica Zhang, Carles Bona-Casas, Lisandro Dalcin, and Hector Gomez. Non-body-fitted fluidstructure interaction: Divergence-conforming B-splines, fullyimplicit dynamics, and variational formulation. Journal of Computational Physics, 374(??):625–653, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304819.

Chen:2017:REP

[CZJ17] Xiang Chen, Xiong Zhang, and Zupeng Jia. A robust and efficient polyhedron subdivision and intersection algorithm for three-dimensional MMALE remapping. Journal of Computational Physics, 338(??):1–17, June 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301225.

Chen:2015:SMM

[CZL^{+}15] Z. Chen, Z. Zong, M. B. Liu, L. Zou, H. T. Li, and C. Shu. An SPH model for multiphase flows with complex interfaces and large density differences. Journal of Computational Physics, 283(??):169–188, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007992.

Cheng:2018:HOC

[CZL18] Jian Cheng, Fan Zhang, and Tiegang Liu. A high order compact least-squares reconstructed discontinuous Galerkin method for the steady-state compressible flows on hybrid grids. Journal of Computational Physics, 362(??):95–113, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300883.

Cai:2017:DPS

[CZW17] Wenjun Cai, Huai Zhang, and Yushun Wang. Dissipationpreserving spectral element method for damped seismic wave equations. Journal of Computational Physics, 350(??):260–279, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306290.

Dehghan:2017:UPO

[DA17] Mehdi Dehghan and Mostafa Abbaszadeh. The use of proper orthogonal decomposition (POD) meshless RBF–FD technique to simulate the shallow water equations. Journal of Computational Physics, 351(??):478–510, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306575.

Doisneau:2017:SLT

[DAO17] François Doisneau, Marco Arienti, and Joseph C. Oefelein. A semi-Lagrangian transport method for kinetic problems with application to dense-to-dilute polydisperse reacting spray flows. Journal of Computational Physics, 329(??):48–72, January 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630540X.

Davidchack:2010:DEM

[Dav10] Ruslan L. Davidchack. Discretization errors in molecular dynamics simulations with deterministic and stochastic thermostats. Journal of Computational Physics, 229(24):9323– 9346, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110004985.

See corrigendum [Dav15].

Davidchack:2015:CDE

[Dav15] Ruslan L. Davidchack. Corrigendum to “Discretization errors in molecular dynamics simulations with deterministic and stochastic thermostats” [J. Comput. Phys. 229 (2010) 9323– 9346]. Journal of Computational Physics, 298(??):816, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004489. See [Dav10].

Diggs:2016:EMC

[DB16a] Angela Diggs and S. Balachandar. Evaluation of methods for calculating volume fraction in Eulerian–Lagrangian multiphase flow simulations. Journal of Computational Physics, 313(??):775–798, May 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001388.

Dumbser:2016:NEF

[DB16b] Michael Dumbser and Dinshaw S. Balsara. A new efficient formulation of the HLLEM Riemann solver for general conservative and non-conservative hyperbolic systems. Journal of Computational Physics, 304(??):275–319, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006786.

Dubey:2018:SDC

[DB18] Ritesh Kumar Dubey and Biswarup Biswas. Suitable diffusion for constructing non-oscillatory entropy stable schemes. Journal of Computational Physics, 372(??):912–930, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302626.

DAlessandro:2017:SOS

[DBD^{+}17] Luca D’Alessandro, Bichoy Bahr, Luca Daniel, Dana Weinstein, and Raffaele Ardito. Shape optimization of solidair porous phononic crystal slabs with widest full 3D bandgap for in-plane acoustic waves. Journal of Computational Physics, 344(??):465–484, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303935.

deBrauer:2016:CSC

[dBIM16] Alexia de Brauer, Angelo Iollo, and Thomas Milcent. A Cartesian scheme for compressible multimaterial models in 3D. Journal of Computational Physics, 313(??):121–143, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000966.

Duarte:2015:NSD

[DBMB15] Max Duarte, Zdenek Bonaventura, Marc Massot, and Anne Bourdon. A numerical strategy to discretize and solve the Poisson equation on dynamically adapted multiresolution grids for time-dependent streamer discharge simulations. Journal of Computational Physics, 289(??):129–148, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001084.

Dargaville:2019:AAS

[DBSS^{+}19] S. Dargaville, A. G. Buchan, R. P. Smedley-Stevenson, P. N. Smith, and C. C. Pain. Angular adaptivity with spherical harmonics for Boltzmann transport. Journal of Computational Physics, 397(??):Article 108846, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305303.

DelReyFernandez:2017:CCD

[DBZ17] David C. Del Rey Fernández, Pieter D. Boom, and David W. Zingg. Corner-corrected diagonal-norm summation-by-parts operators for the first derivative with increased order of accuracy. Journal of Computational Physics, 330(??):902–923, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305599.

Delgado:2018:CRT

[DC18a] Carlos Delgado and Manuel Felipe Cátedra. Combination of ray-tracing and the method of moments for electromagnetic radiation analysis using reduced meshes. Journal of Computational Physics, 361(??):412–423, May 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300500.

Deng:2018:NSD

[DC18b] Xiaogang Deng and Yaming Chen. A novel strategy for deriving high-order stable boundary closures based on global conservation, I: Basic formulas. Journal of Computational Physics, 372(??):80–106, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303905.

Deckers:2016:WBM

[DCA^{+}16] Elke Deckers, Claus Claeys, Onur Atak, Jean-Philippe Groby, Olivier Dazel, and Wim Desmet. A wave based method to predict the absorption, reflection and transmission coefficient of two-dimensional rigid frame porous structures with periodic inclusions. Journal of Computational Physics, 312(??):115–138, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000619.

Dorschner:2015:GAM

[DCBK15] B. Dorschner, S. S. Chikatamarla, F. Bösch, and I. V. Karlin. Grad’s approximation for moving and stationary walls in entropic lattice Boltzmann simulations. Journal of Computational Physics, 295(??):340–354, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002636.

Dhia:2016:UPM

[DCCC16] A.-S. Bonnet-Ben Dhia, C. Carvalho, L. Chesnel, and P. Ciarlet, Jr. On the use of Perfectly Matched Layers at corners for scattering problems with sign-changing coefficients. Journal of Computational Physics, 322(??):224–247, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302625.

DelReyFernandez:2019:SGE

[DCCH19] David C. Del Rey Fernández, Jared Crean, Mark H. Carpenter, and Jason E. Hicken. Staggered-grid entropystable multidimensional summation-by-parts discretizations on curvilinear coordinates. Journal of Computational Physics, 392(??):161–186, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930275X.

Dorr:2018:HOD

[DCD^{+}18] Milo R. Dorr, Phillip Colella, Mikhail A. Dorf, Debojyoti Ghosh, Jeffrey A. F. Hittinger, and Peter O. Schwartz. High-order discretization of a gyrokinetic Vlasov model in edge plasma geometry. Journal of Computational Physics, 373(??):605–630, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304637.

deConchard:2019:PML

[dCMR19] Antoine Vermeil de Conchard, Huina Mao, and Romain Rumpler. A perfectly matched layer formulation adapted for fast frequency sweeps of exterior acoustics finite element models. Journal of Computational Physics, 398(??):Article 108878, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119305753.

Delmotte:2015:GFB

[DCP15] Blaise Delmotte, Eric Climent, and Franck Plouraboué. A general formulation of Bead Models applied to flexible fibers and active filaments at low Reynolds number. Journal of Computational Physics, 286(??):14–37, April 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000303.

Tejero-del-Caz:2017:IIE

[dCPDC^{+}17] A. Tejero del Caz, J. I. Fernández Palop, J. M. Díaz-Cabrera, G. F. Regodón, R. Carmona-Cabezas, and J. Ballesteros. Ion injection in electrostatic particle-in-cell simulations of the ion sheath. Journal of Computational Physics, 350(??):747–758, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306770.

Das:2015:MLV

[DD15] A. K. Das and P. K. Das. Modeling of liquid-vapor phase change using smoothed particle hydrodynamics. Journal of Computational Physics, 303(??):125–145, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006191.

Dai:2016:CSC

[DD16a] Shibin Dai and Qiang Du. Computational studies of coarsening rates for the Cahn–Hilliard equation with phase-dependent diffusion mobility. Journal of Computational Physics, 310(??):85– 108, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911600019X.

Duru:2016:DER

[DD16b] Kenneth Duru and Eric M. Dunham. Dynamic earthquake rupture simulations on nonplanar faults embedded in 3D geometrically complex, heterogeneous elastic solids. Journal of Computational Physics, 305(??):185–207, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006853.

Daripa:2017:MSS

[DD17a] Prabir Daripa and Sourav Dutta. Modeling and simulation of surfactant-polymer flooding using a new hybrid method. Journal of Computational Physics, 335(??):249–282, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300542.

Degond:2017:APM

[DD17b] Pierre Degond and Fabrice Deluzet. Asymptotic-preserving methods and multiscale models for plasma physics. Journal of Computational Physics, 336(??):429–457, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730102X.

Sarto:2017:MAA

[DD17c] D. Del Sarto and E. Deriaz. A multigrid AMR algorithm for the study of magnetic reconnection. Journal of Computational Physics, 351(??):511–533, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306277.

Degond:2017:APP

[DDD17] P. Degond, F. Deluzet, and D. Doyen. Asymptotic-Preserving Particle-In-Cell methods for the Vlasov–Maxwell system in the quasi-neutral limit. Journal of Computational Physics, 330(??):467–492, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306040.

Dorf:2018:HOF

[DDH^{+}18] M. Dorf, M. Dorr, J. Hittinger, W. Lee, and D. Ghosh. High-order finite-volume modeling of drift waves. Journal of Computational Physics, 373(??):446–454, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304649.

Demaziere:2017:DPK

[DDJ17] C. Demazière, V. Dykin, and K. Jareteg. Development of a point-kinetic verification scheme for nuclear reactor applications. Journal of Computational Physics, 339(??):396–411, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302152.

Dakin:2018:ILW

[DDJ18] Gautier Dakin, Bruno Després, and Stéphane Jaouen. Inverse Lax–Wendroff boundary treatment for compressible Lagrangeremap hydrodynamics on Cartesian grids. Journal of Computational Physics, 353(??):228–257, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307520.

Dakin:2019:HOS

[DDJ19] Gautier Dakin, Bruno Després, and Stéphane Jaouen. Highorder staggered schemes for compressible hydrodynamics. weak consistency and numerical validation. Journal of Computational Physics, 376(??):339–364, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830648X.

Pietro:2018:DSG

[DDM18] Daniele A. Di Pietro, Jérôme Droniou, and Gianmarco Manzini. Discontinuous Skeletal Gradient Discretisation methods on polytopal meshes. Journal of Computational Physics, 355(??):397–425, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117308562.

Demange:2019:SAS

[DDM^{+}19] Jérémie Demange, Laurent Debreu, Patrick Marchesiello, Florian Lemarié, Eric Blayo, and Christopher Eldred. Stability analysis of split-explicit free surface ocean models: Implication of the depth-independent barotropic mode approximation. Journal of Computational Physics, 398(??):Article 108875, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119305662.

deDiego:2019:INS

[dDPG19] G. G. de Diego, A. Palha, and M. Gerritsma. Inclusion of no-slip boundary conditions in the MEEVC scheme. Journal of Computational Physics, 378(??):615–633, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307691.

Dobbelaere:2015:CMP

[DDV^{+}15] D. Dobbelaere, D. De Zutter, J. Van Hese, J. Sercu, T. Boonen, and H. Rogier. A Calderón multiplicative preconditioner for the electromagnetic Poincaré–Steklov operator of a heterogeneous domain with scattering applications. Journal of Computational Physics, 303(??):355–371, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006610.

Demeester:2018:SAP

[DDV18] Toon Demeester, Joris Degroote, and Jan Vierendeels. Stability analysis of a partitioned iterative method for steady free surface flow. Journal of Computational Physics, 354 (??):387–392, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117308203.

De:2018:DII

[De18] Arnab Kr. De. A diffuse interface immersed boundary method for complex moving boundary problems. Journal of Computational Physics, 365(??):226–251, August 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302225.

Delzanno:2015:MDF

[Del15] G. L. Delzanno. Multi-dimensional, fully-implicit, spectral method for the Vlasov–Maxwell equations with exact conservation laws in discrete form. Journal of Computational Physics, 301(??):338–356, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004738.

Duncan:2016:HFS

[DEZ16] Andrew Duncan, Radek Erban, and Konstantinos Zygalakis. Hybrid framework for the simulation of stochastic chemical kinetics. Journal of Computational Physics, 326(??):398–419, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303916.

Diot:2016:IRM

[DF16] Steven Diot and Marianne M. François. An interface reconstruction method based on an analytical formula for 3D arbitrary convex cells. Journal of Computational Physics, 305(??):63–74, January 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006749.

Deparis:2016:FBP

[DFGQ16] Simone Deparis, Davide Forti, Gwenol Grandperrin, and Alfio Quarteroni. FaCSI: a block parallel preconditioner for fluidstructure interaction in hemodynamics. Journal of Computational Physics, 327(??):700–718, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304983.

deFigueiredo:2017:BSI

[dFGS^{+}17] Leandro Passos de Figueiredo, Dario Grana, Marcio Santos, Wagner Figueiredo, Mauro Roisenberg, and Guenther Schwedersky Neto. Bayesian seismic inversion based on rockphysics prior modeling for the joint estimation of acoustic impedance, porosity and lithofacies. Journal of Computational Physics, 336(??):128–142, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301067.

deFrutos:2016:PMI

[dFJN16] Javier de Frutos, Volker John, and Julia Novo. Projection methods for incompressible flow problems with WENO finite difference schemes. Journal of Computational Physics, 309(??):368–386, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008608.

Degond:2017:DAH

[DFM17] Pierre Degond, Marina A. Ferreira, and Sebastien Motsch. Damped Arrow-Hurwicz algorithm for sphere packing. Journal of Computational Physics, 332(??):47–65, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306398.

Dobson:2016:CLA

[DFS16] Matthew Dobson, Ian Fox, and Alexandra Saracino. Cell list algorithms for nonequilibrium molecular dynamics. Journal of Computational Physics, 315(??):211–220, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300213.

deFrahan:2015:NLP

[dFVJ15] Marc T. Henry de Frahan, Sreenivas Varadan, and Eric Johnsen. A new limiting procedure for discontinuous Galerkin methods applied to compressible multiphase flows with shocks and interfaces. Journal of Computational Physics, 280(??):489–509, January 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114006664.

Daude:2016:CBN

[DG16a] F. Daude and P. Galon. On the computation of the Baer-Nunziato model using ALE formulation with HLL- and HLLCtype solvers towards fluid-structure interactions. Journal of Computational Physics, 304(??):189–230, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006774.

Detrixhe:2016:HMP

[DG16b] Miles Detrixhe and Frédéric Gibou. Hybrid massively parallel fast sweeping method for static Hamilton–Jacobi equations. Journal of Computational Physics, 322(??):199–223, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116302480.

Dhiman:2016:CDI

[DG16c] Isha Dhiman and Arvind Kumar Gupta. Collective dynamics of an inhomogeneous two-channel exclusion process: Theory and Monte Carlo simulations. Journal of Computational Physics, 309(??):227–240, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000115.

Daude:2018:FVA

[DG18] F. Daude and P. Galon. A finite-volume approach for compressible single- and two-phase flows in flexible pipelines with fluid-structure interaction. Journal of Computational Physics, 362(??):375–408, June 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301207.

Dambrine:2017:NSH

[DGHP17] M. Dambrine, I. Greff, H. Harbrecht, and B. Puig. Numerical solution of the homogeneous Neumann boundary value problem on domains with a thin layer of random thickness. Journal of Computational Physics, 330(??):943–959, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305484.

Dolean:2015:ETC

[DGL^{+}15] Victorita Dolean, Martin J. Gander, Stephane Lanteri, Jin-Fa Lee, and Zhen Peng. Effective transmission conditions for domain decomposition methods applied to the timeharmonic curl-curl Maxwell’s equations. Journal of Computational Physics, 280(??):232–247, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006603.

Deng:2017:LCS

[DGMT17] Q. Deng, V. Ginting, B. McCaskill, and P. Torsu. A locally conservative stabilized continuous Galerkin finite element method for two-phase flow in poroelastic subsurfaces. Journal of Computational Physics, 347(??):78–98, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304692.

Dana:2018:MFS

[DGW18] Saumik Dana, Benjamin Ganis, and Mary F. Wheeler. A multiscale fixed stress split iterative scheme for coupled flow and poromechanics in deep subsurface reservoirs. Journal of Computational Physics, 352(??):1–22, January 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730709X.

Dobramysl:2018:MAS

[DH18a] U. Dobramysl and D. Holcman. Mixed analytical-stochastic simulation method for the recovery of a Brownian gradient source from probability fluxes to small windows. Journal of Computational Physics, 355(??):22–36, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308252.

Dolz:2018:HMA

[DH18b] Jürgen Dölz and Helmut Harbrecht. Hierarchical matrix approximation for the uncertainty quantification of potentials on random domains. Journal of Computational Physics, 371(??):506–527, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303498.

Du:2016:FAS

[DHC16] Tao Du, Dan Hu, and David Cai. A fast algorithm for the simulation of arterial pulse waves. Journal of Computational Physics, 314(??):450–464, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001881.

DiPietro:2018:MMS

[DHH^{+}18] Kelsey L. DiPietro, Ronald D. Haynes, Weizhang Huang, Alan E. Lindsay, and Yufei Yu. Moving mesh simulation of contact sets in two dimensional models of elasticelectrostatic deflection problems. Journal of Computational Physics, 375(??):763–782, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305862.

Diddens:2017:DFE

[Did17] Christian Diddens. Detailed finite element method modeling of evaporating multi-component droplets. Journal of Computational Physics, 340(??):670–687, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302504.

Diethelm:2015:IES

[Die15] Kai Diethelm. Increasing the efficiency of shooting methods for terminal value problems of fractional order. Journal of Computational Physics, 293(??):135–141, July 15, 2015.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007426.

Deng:2018:HFD

[DIX^{+}18] Xi Deng, Satoshi Inaba, Bin Xie, Keh-Ming Shyue, and Feng Xiao. High fidelity discontinuity-resolving reconstruction for compressible multiphase flows with moving interfaces. Journal of Computational Physics, 371(??):945–966, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301967.

Ducru:2017:KRM

[DJD^{+}17] Pablo Ducru, Colin Josey, Karia Dibert, Vladimir Sobes, Benoit Forget, and Kord Smith. Kernel reconstruction methods for Doppler broadening — temperature interpolation by linear combination of reference cross sections at optimally chosen temperatures. Journal of Computational Physics, 335(??):535–557, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300554.

Drawert:2019:HSD

[DJL^{+}19] Brian Drawert, Bruno Jacob, Zhen Li, Tau-Mu Yi, and Linda Petzold. A hybrid smoothed dissipative particle dynamics (SDPD) spatial stochastic simulation algorithm (sSSA) for advection-diffusion-reaction problems. Journal of Computational Physics, 378(??):1–17, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307101.

Du:2018:SLS

[DJLQ18] Qiang Du, Lili Ju, Xiao Li, and Zhonghua Qiao. Stabilized linear semi-implicit schemes for the nonlocal Cahn–Hilliard equation. Journal of Computational Physics, 363(??):39–54, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301050.

deJesus:2015:FTA

[dJRP^{+}15] Wellington C. de Jesus, Alexandre M. Roma, Márcio R. Pivello, Millena M. Villar, and Aristeu da Silveira-Neto. A 3D front-tracking approach for simulation of a twophase fluid with insoluble surfactant. Journal of Computational Physics, 281(??):403–420, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007037.

Deckers:2018:PTR

[DJV^{+}18] Elke Deckers, Stijn Jonckheere, Lucas Van Belle, Claus Claeys, and Wim Desmet. Prediction of transmission, reflection and absorption coefficients of periodic structures using a hybrid Wave Based–Finite Element unit cell method. Journal of Computational Physics, 356(??):282–302, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308768.

Deng:2018:CT

[DK18a] Yongbo Deng and Jan G. Korvink. Corrigendum to “Selfconsistent adjoint analysis for topology optimization of electromagnetic waves” [J. Comput. Phys. 361 (2018) 353–376]. Journal of Computational Physics, 370(??):85, October 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301864. See [DK18b].

Deng:2018:SCA

[DK18b] Yongbo Deng and Jan G. Korvink. Self-consistent adjoint analysis for topology optimization of electromagnetic waves. Journal of Computational Physics, 361(??):353–376, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830055X. See corrigendum

[DK18a].

Danilov:2019:GOS

[DK19] S. Danilov and A. Kutsenko. On the geometric origin of spurious waves in finite-volume discretizations of shallow water equations on triangular meshes. Journal of Computational Physics, 398(??):Article 108891, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305893.

Duan:2015:CCS

[DKC15] Guangtao Duan, Seiichi Koshizuka, and Bin Chen. A contoured continuum surface force model for particle methods. Journal of Computational Physics, 298(??):280–304, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003836.

Duru:2015:BCS

[DKK15] Kenneth Duru, Jeremy E. Kozdon, and Gunilla Kreiss. Boundary conditions and stability of a perfectly matched layer for the elastic wave equation in first order form. Journal of Computational Physics, 303(??):372–395, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006567.

Dobrev:2018:SLC

[DKK^{+}18] V. Dobrev, Tz. Kolev, D. Kuzmin, R. Rieben, and V. Tomov. Sequential limiting in continuous and discontinuous Galerkin methods for the Euler equations. Journal of Computational Physics, 356(??):372–390, March 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308963.

Delmotte:2015:LSS

[DKPC15] Blaise Delmotte, Eric E. Keaveny, Franck Plouraboué, and Eric Climent. Large-scale simulation of steady and timedependent active suspensions with the force-coupling method. Journal of Computational Physics, 302(??):524–547, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006130.

Dimitriadis:2015:GNL

[DKTH15] Alexandros I. Dimitriadis, Nikolaos V. Kantartzis, Theodoros D. Tsiboukis, and Christian Hafner. Generalized non-local surface susceptibility model and Fresnel coefficients for the characterization of periodic metafilms with bianisotropic scatterers. Journal of Computational Physics, 281(??):251–268, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007104.

Despres:2015:AMP

[DL15] B. Després and E. Labourasse. Angular momentum preserving cell-centered Lagrangian and Eulerian schemes on arbitrary grids. Journal of Computational Physics, 290(??):28–54, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001023.

Dumbser:2016:SRA

[DL16] Michael Dumbser and Raphaël Loubère. A simple robust and accurate a posteriori sub-cell finite volume limiter for the discontinuous Galerkin method on unstructured meshes. Journal of Computational Physics, 319(??):163–199, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301292.

DiPietro:2017:MAS

[DL17] Kelsey L. DiPietro and Alan E. Lindsay. Monge–Ampére simulation of fourth order PDEs in two dimensions with application to elastic-electrostatic contact problems. Journal of Computational Physics, 349(??):328–350, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306034.

Dong:2018:LTS

[DL18a] Haitao Dong and Fujun Liu. Large time step wave adding scheme for systems of hyperbolic conservation laws. Journal of Computational Physics, 374(??):331–360, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304777.

Du:2018:HWR

[DL18b] Zhifang Du and Jiequan Li. A Hermite WENO reconstruction for fourth order temporal accurate schemes based on the GRP solver for hyperbolic conservation laws. Journal of Computational Physics, 355(??):385–396, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308616.

Du:2018:TSF

[DL18c] Zhifang Du and Jiequan Li. A two-stage fourth order timeaccurate discretization for Lax–Wendroff type flow solvers II. High order numerical boundary conditions. Journal of Computational Physics, 369(??):125–147, September 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302948.

delaAsuncion:2017:STG

[dlAC17] M. de la Asunción and M. J. Castro. Simulation of tsunamis generated by landslides using adaptive mesh refinement on GPU. Journal of Computational Physics, 345(??):91–110, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303856.

Ding:2015:HOA

[DLC15] Hengfei Ding, Changpin Li, and YangQuan Chen. Highorder algorithms for Riesz derivative and their applications (II). Journal of Computational Physics, 293(??):218–237, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004148.

delaCruz:2017:CGH

[dlCGCA17] Roberto de la Cruz, Pilar Guerrero, Juan Calvo, and Tomás Alarcón. Coarse-graining and hybrid methods for efficient simulation of stochastic multi-scale models of tumour growth. Journal of Computational Physics, 350(??):974–991, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306782.

deLaborderie:2018:NAH

[dLDG^{+}18] J. de Laborderie, F. Duchaine, L. Gicquel, O. Vermorel, G. Wang, and S. Moreau. Numerical analysis of a high-order unstructured overset grid method for compressible LES of turbomachinery. Journal of Computational Physics, 363(??):371– 398, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301281.

deLangavant:2017:LSS

[dLGT^{+}17] Charles Cleret de Langavant, Arthur Guittet, Maxime Theillard, Fernando Temprano-Coleto, and Frédéric Gibou. Levelset simulations of soluble surfactant driven flows. Journal of Computational Physics, 348(??):271–297, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305053.

delaHoz:2016:PSM

[dlHC16] Francisco de la Hoz and Carlota M. Cuesta. A pseudo-spectral method for a non-local KdV–Burgers equation posed on R. Journal of Computational Physics, 311(??):45–61, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000395.

Do:2017:WBA

[DLK17] Seongju Do, Haojun Li, and Myungjoo Kang. Waveletbased adaptation methodology combined with finite difference WENO to solve ideal magnetohydrodynamics. Journal of Computational Physics, 339(??):482–499, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302231.

deLuna:2019:MCL

[dLKK19] Manuel Quezada de Luna, Dmitri Kuzmin, and Christopher E. Kees. A monolithic conservative level set method with built-in redistancing. Journal of Computational Physics, 379(??):262–278, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118307897.

Du:2017:LDG

[DLL^{+}17] Yanwei Du, Yang Liu, Hong Li, Zhichao Fang, and Siriguleng He. Local discontinuous Galerkin method for a nonlinear time-fractional fourth-order partial differential equation. Journal of Computational Physics, 344(??):108–126, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303662.

Dairay:2017:NDV

[DLLV17] Thibault Dairay, Eric Lamballais, Sylvain Laizet, and John Christos Vassilicos. Numerical dissipation vs. subgridscale modelling for large eddy simulation. Journal of Computational Physics, 337(??):252–274, May 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301298.

DelPino:2018:APM

[DLM18] S. Del Pino, E. Labourasse, and G. Morel. An asymptotic preserving multidimensional ALE method for a system of two compressible flows coupled with friction. Journal of Computational Physics, 363(??):268–301, June 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300986.

Dimarco:2018:SOI

[DLMDV18] Giacomo Dimarco, Raphaël Loubère, Victor Michel-Dansac, and Marie-Hélène Vignal. Second-order implicit-explicit total variation diminishing schemes for the Euler system in the low Mach regime. Journal of Computational Physics, 372(??):178–201, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304005.

Dimarco:2015:TUE

[DLN15] Giacomo Dimarco, Raphaël Loubère, and Jacek Narski. Towards an ultra efficient kinetic scheme. Part III: Highperformance-computing. Journal of Computational Physics, 284(??):22–39, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114008407.

Dimarco:2018:ENM

[DLNR18] Giacomo Dimarco, Raphaël Loubère, Jacek Narski, and Thomas Rey. An efficient numerical method for solving the Boltzmann equation in multidimensions. Journal of Computational Physics, 353(??):46–81, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307489.

DeLorenzo:2019:HPT

[DLP19] M. De Lorenzo, Ph. Lafon, and M. Pelanti. A hyperbolic phase-transition model with non-instantaneous EoSindependent relaxation procedures. Journal of Computational Physics, 379(??):279–308, February 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307903.

Dimarco:2015:MFS

[DLR15] Giacomo Dimarco, Raphaël Loubère, and Vittorio Rispoli. A multiscale fast semi-Lagrangian method for rarefied gas dynamics. Journal of Computational Physics, 291(??):99–119, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001011.

Du:2015:NSS

[DLS15] Kui Du, Buyang Li, and Weiwei Sun. A numerical study on the stability of a class of Helmholtz problems. Journal of Computational Physics, 287(??):46–59, 2015. CODEN JCTPAH. ISSN

0021-9991. URL http://www.sciencedirect.com/science/
article/pii/S0021999115000698.

Duan:2019:NMP

[DLWY19] Chenghua Duan, Chun Liu, Cheng Wang, and Xingye Yue. Numerical methods for porous medium equation by an energetic variational approach. Journal of Computational Physics, 385(??):13–32, May 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301135.

Damle:2017:SKL

[DLY17] Anil Damle, Lin Lin, and Lexing Ying. SCDM-k: Localized orbitals for solids via selected columns of the density matrix. Journal of Computational Physics, 334(??):1–15, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116307215.

Dechriste:2016:CCC

[DM16] G. Dechristé and L. Mieussens. A Cartesian cut cell method for rarefied flow simulations around moving obstacles. Journal of Computational Physics, 314(??):465–488, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001765.

Dahal:2017:NMS

[DM17a] Jeevan Dahal and Jacob A. McFarland. A numerical method for shock driven multiphase flow with evaporating particles. Journal of Computational Physics, 344(??):210–233, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303625.

Diosady:2017:TPP

[DM17b] Laslo T. Diosady and Scott M. Murman. Tensor-product preconditioners for higher-order space-time discontinuous Galerkin methods. Journal of Computational Physics, 330 (??):296–318, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306088.

Pietro:2018:WIP

[DM18] Daniele A. Di Pietro and Fabien Marche. Weighted interior penalty discretization of fully nonlinear and weakly dispersive free surface shallow water flows. Journal of Computational Physics, 355(??):285–309, February 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730846X.

Diosady:2019:STP

[DM19] Laslo T. Diosady and Scott M. Murman. Scalable tensorproduct preconditioners for high-order finite-element methods: Scalar equations. Journal of Computational Physics, 394(??):759–776, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119302931.

Das:2015:COM

[DMAM15] Shankhadeep Das, Sanjay R. Mathur, Alina Alexeenko, and Jayathi Y. Murthy. A coupled ordinates method for solution acceleration of rarefied gas dynamics simulations. Journal of Computational Physics, 289(??):96–115, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001059.

Droniou:2019:DAF

[DMM19] Jérôme Droniou, Matej Medla, and Karol Mikula. Design and analysis of finite volume methods for elliptic equations with oblique derivatives; application to Earth gravity field modelling. Journal of Computational Physics, 398 (??):Article 108876, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S002199911930573X.

DeOliveiraVilaca:2019:NAM

[DMRB19] Luis Miguel De Oliveira Vilaca, Michel C. Milinkovitch, and Ricardo Ruiz-Baier. Numerical approximation of a 3D mechanochemical interface model for skin patterning. Journal of Computational Physics, 384(??):383–404, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300488.

deMoraes:2017:MGC

[dMRHJ17] Rafael J. de Moraes, José R. P. Rodrigues, Hadi Hajibeygi, and Jan Dirk Jansen. Multiscale gradient computation for flow in heterogeneous porous media. Journal of Computational Physics, 336(??):644–663, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301171.

Dobravec:2017:CAF

[DMS17] Tadej Dobravec, Bostjan Mavric, and Bozidar Sarler. A cellular automaton–finite volume method for the simulation of dendritic and eutectic growth in binary alloys using an adaptive mesh refinement. Journal of Computational Physics, 349(??):351–375, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730582X.

Donatelli:2016:SAS

[DMSC16] Marco Donatelli, Mariarosa Mazza, and Stefano Serra-Capizzano. Spectral analysis and structure preserving preconditioners for fractional diffusion equations. Journal of Computational Physics, 307(??):262–279, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008177.

Duran:2015:APS

[DMTB15] A. Duran, F. Marche, R. Turpault, and C. Berthon. Asymptotic preserving scheme for the shallow water equations with source terms on unstructured meshes. Journal of Computational Physics, 287(??):184–206, 2015. CODEN JCT-

PAH. ISSN 0021-9991. URL http://www.sciencedirect.
com/science/article/pii/S0021999115000686.

Demirel:2015:EMT

[DNBH15] Abdullah Demirel, Jens Niegemann, Kurt Busch, and Marlis Hochbruck. Efficient multiple time-stepping algorithms of higher order. Journal of Computational Physics, 285(??): 133–148, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115000224.

Deluzet:2015:NSP

[DNOP15] Fabrice Deluzet, Claudia Negulescu, Maurizio Ottaviani, and Stefan Possanner. Numerical study of the plasma tearing instability on the resistive time scale. Journal of Computational Physics, 280(??):602–625, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006779.

Dodig:2017:BIM

[Dod17] H. Dodig. A boundary integral method for numerical computation of radar cross section of 3D targets using hybrid BEM/FEM with edge elements. Journal of Computational Physics, 348(??):790–802, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730551X.

Domino:2018:DON

[Dom18] Stefan P. Domino. Design-order, non-conformal low-Mach fluid algorithms using a hybrid CVFEM/DG approach. Journal of Computational Physics, 359(??):331–351, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300172.

Dong:2015:CLE

[Don15a] S. Dong. A convective-like energy-stable open boundary condition for simulations of incompressible flows. Journal of Computational Physics, 302(??):300–328, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006105.

Dong:2015:PFN

[Don15b] S. Dong. Physical formulation and numerical algorithm for simulating N immiscible incompressible fluids involving general order parameters. Journal of Computational Physics, 283(??):98–128, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114008018.

Dong:2017:WBM

[Don17] S. Dong. Wall-bounded multiphase flows of N immiscible incompressible fluids: Consistency and contact-angle boundary condition. Journal of Computational Physics, 338(??):21–67, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301511.

Dong:2018:MFI

[Don18] S. Dong. Multiphase flows of N immiscible incompressible fluids: A reduction-consistent and thermodynamicallyconsistent formulation and associated algorithm. Journal of Computational Physics, 361(??):1–49, May 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300512.

Drozdov:2017:TME

[DOO17] Grigory Drozdov, Igor Ostanin, and Ivan Oseledets. Timeand memory-efficient representation of complex mesoscale potentials. Journal of Computational Physics, 343(??):110–114, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303352.

Dimarco:2019:MSC

[DP19] Giacomo Dimarco and Lorenzo Pareschi. Multi-scale control variate methods for uncertainty quantification in kinetic equations. Journal of Computational Physics, 388(??):63–89, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301718.

Deng:2017:ASP

[DPK17] Mingge Deng, Wenxiao Pan, and George Em Karniadakis. Anisotropic single-particle dissipative particle dynamics model. Journal of Computational Physics, 336(??):481–491, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300499.

Deimert:2016:CEF

[DPO16] C. Deimert, M. E. Potter, and M. Okoniewski. Collocated electrodynamic FDTD schemes using overlapping Yee grids and higher-order Hodge duals. Journal of Computational Physics, 326(??):629–649, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304053.

Dumbser:2016:HOA

[DPRZ16] Michael Dumbser, Ilya Peshkov, Evgeniy Romenski, and Olindo Zanotti. High order ADER schemes for a unified first order hyperbolic formulation of continuum mechanics: Viscous heat-conducting fluids and elastic solids. Journal of Computational Physics, 314(??):824–862, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000693.

Dumbser:2017:HOA

[DPRZ17] Michael Dumbser, Ilya Peshkov, Evgeniy Romenski, and Olindo Zanotti. High order ADER schemes for a unified first order hyperbolic formulation of Newtonian continuum mechanics coupled with electro-dynamics. Journal of Computational Physics, 348(??):298–342, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305284.

dePando:2016:NMO

[dPSS16] Miguel Fosas de Pando, Peter J. Schmid, and Denis Sipp. Nonlinear model-order reduction for compressible flow solvers using the Discrete Empirical Interpolation Method. Journal of Computational Physics, 324(??):194–209, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303448.

Densmore:2015:MCS

[DPW^{+}15] J. D. Densmore, H. Park, A. B. Wollaber, R. M. Rauenzahn, and D. A. Knoll. Monte Carlo simulation methods in momentbased scale-bridging algorithms for thermal radiative-transfer problems. Journal of Computational Physics, 284(??):40–58, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008377.

Delchini:2015:EBA

[DRM15] Marc O. Delchini, Jean C. Ragusa, and Jim Morel. Entropybased artificial viscosity stabilization for non-equilibrium grey radiation-hydrodynamics. Journal of Computational Physics, 296(??):293–313, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911500296X.

DElia:2016:OBM

[DRP^{+}16] Marta D’Elia, Denis Ridzal, Kara J. Peterson, Pavel Bochev, and Mikhail Shashkov. Optimization-based mesh correction with volume and convexity constraints. Journal of Computational Physics, 313(??):455–477, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001224.

Dalcin:2019:CES

[DRZ^{+}19] Lisandro Dalcin, Diego Rojas, Stefano Zampini, David C. Del Rey Fernández, Mark H. Carpenter, and Matteo Parsani. Conservative and entropy stable solid wall boundary conditions for the compressible Navier–Stokes equations: Adiabatic wall and heat entropy transfer. Journal of Computational Physics, 397(??):Article 108775, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119304590.

Dai:2015:IDA

[DS15a] William W. Dai and Anthony J. Scannapieco. Interfaceand discontinuity-aware numerical schemes for plasma 3-T radiation diffusion in two and three dimensions. Journal of Computational Physics, 300(??):643–664, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004866.

Dai:2015:SOA

[DS15b] William W. Dai and Anthony J. Scannapieco. Secondorder accurate interface- and discontinuity-aware diffusion solvers in two and three dimensions. Journal of Computational Physics, 281(??):982–1002, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007220.

Dong:2015:PCS

[DS15c] S. Dong and J. Shen. A pressure correction scheme for generalized form of energy-stable open boundary conditions for incompressible flows. Journal of Computational Physics, 291(??):254–278, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001485.

Dorda:2015:WSC

[DS15d] Antonius Dorda and Ferdinand Schürrer. A WENOsolver combined with adaptive momentum discretization for the Wigner transport equation and its application to resonant tunneling diodes. Journal of Computational Physics, 284(??):95–116, March 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114008432.

DiPietro:2016:PDA

[DS16] Daniele A. Di Pietro and Ruben Specogna. An a posterioridriven adaptive mixed high-order method with application to electrostatics. Journal of Computational Physics, 326 (??):35–55, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303989.

DeCorato:2016:FEF

[DSH^{+}16] M. De Corato, J. J. M. Slot, M. Hütter, G. D’Avino, P. L. Maffettone, and M. A. Hulsen. Finite element formulation of fluctuating hydrodynamics for fluids filled with rigid particles using boundary fitted meshes. Journal of Computational Physics, 316(??):632–651, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300869.

daSilva:2015:SEC

[dSPDH15] Filipe da Silva, Martin Campos Pinto, Bruno Després, and Stéphane Heuraux. Stable explicit coupling of the Yee scheme with a linear current model in fluctuating magnetized plasmas. Journal of Computational Physics, 295(??):24–45, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002417.

Diaz:2018:CNS

[DSS18] Manuel A. Diaz, Maxim A. Solovchuk, and Tony W. H. Sheu. A conservative numerical scheme for modeling nonlinear acoustic propagation in thermoviscous homogeneous media. Journal of Computational Physics, 363(??):200–230, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300810.

Douasbin:2018:DTD

[DSSP18] Q. Douasbin, C. Scalo, L. Selle, and T. Poinsot. Delayed-time domain impedance boundary conditions (D-TDIBC). Journal of Computational Physics, 371(??):50–66, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830295X.

Deng:2019:FOS

[DSX19] Xi Deng, Yuya Shimizu, and Feng Xiao. A fifth-order shock capturing scheme with two-stage boundary variation diminishing algorithm. Journal of Computational Physics, 386(??):323– 349, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911930141X.

Davidson:2015:IHP

[DTA^{+}15] A. Davidson, A. Tableman, W. An, F. S. Tsung, W. Lu, J. Vieira, R. A. Fonseca, L. O. Silva, and W. B. Mori. Implementation of a hybrid particle code with a PIC description in r–z and a gridless description in φ into OSIRIS. Journal of Computational Physics, 281(??):1063–1077, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007529.

deTullio:2016:MLS

[dTP16] M. D. de Tullio and G. Pascazio. A moving-least-squares immersed boundary method for simulating the fluid-structure interaction of elastic bodies with arbitrary thickness. Journal of Computational Physics, 325(??):201–225, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303692.

Du:2018:IBE

[Du18] Yongle Du. Implicit boundary equations for conservative Navier–Stokes equations. Journal of Computational Physics, 375(??):641–658, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306041.

Londersele:2017:DSA

[DV17] Arne Van Londersele and Daniël De Zutter and Dries Vande Ginste. An in-depth stability analysis of nonuniform FDTD combined with novel local implicitization techniques. Journal of Computational Physics, 342(??):177–193, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303078.

Dilz:2017:DIE

[DvB17] R. J. Dilz and M. C. van Beurden. A domain integral equation approach for simulating two dimensional transverse electric scattering in a layered medium with a Gabor frame discretization. Journal of Computational Physics, 345(??):528–542, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304096.

Dotlic:2016:SOA

[DVP^{+}16] M. Dotlić, D. Vidović, B. Pokorni, M. Pusić, and M. Dimkić. Second-order accurate finite volume method for well-driven flows. Journal of Computational Physics, 307(??):460–475, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008402.

Denner:2015:NTS

[DvW15a] Fabian Denner and Berend G. M. van Wachem. Numerical time-step restrictions as a result of capillary waves. Journal of Computational Physics, 285(??):24–40, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500025X.

Denner:2015:TDT

[DvW15b] Fabian Denner and Berend G. M. van Wachem. TVD differencing on three-dimensional unstructured meshes with monotonicity-preserving correction of mesh skewness. Journal of Computational Physics, 298(??):466–479, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003939.

Denner:2019:CPB

[DvW19] Fabian Denner and Berend G. M. van Wachem. Corrigendum to “Pressure-based algorithm for compressible interfacial flows with acoustically-conservative interface discretisation” [J. Comput. Phys. 367 (2018) 192–234]. Journal of Computational Physics, 381(??):290–291, March 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307356. See [DXvW18].

Duo:2018:NAF

[DvWZ18] Siwei Duo, Hans Werner van Wyk, and Yanzhi Zhang. A novel and accurate finite difference method for the fractional Laplacian and the fractional Poisson problem. Journal of Computational Physics, 355(??):233–252, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308495.

Duo:2019:FPF

[DW19] Siwei Duo and Hong Wang. A fractional phase-field model using an infinitesimal generator of α stable Lévy process. Journal of Computational Physics, 384(??):253–269, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301007.

Derigs:2018:IGM

[DWG^{+}18] Dominik Derigs, Andrew R. Winters, Gregor J. Gassner, Stefanie Walch, and Marvin Bohm. Ideal GLM–MHD: About the entropy consistent nine-wave magnetic field divergence diminishing ideal magnetohydrodynamics equations. Journal of Computational Physics, 364(??):420–467, July 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301463.

Derigs:2016:NHO

[DWGW16] Dominik Derigs, Andrew R. Winters, Gregor J. Gassner, and Stefanie Walch. A novel high-order, entropy stable, 3D AMR MHD solver with guaranteed positive pressure. Journal of Computational Physics, 317(??):223–256, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301036.

Derigs:2017:NAT

[DWGW17] Dominik Derigs, Andrew R. Winters, Gregor J. Gassner, and Stefanie Walch. A novel averaging technique for discrete entropy-stable dissipation operators for ideal MHD. Journal of Computational Physics, 330(??):624–632, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305630.

Deck:2018:RLN

[DWR18] Sébastien Deck, Pierre-Elie Weiss, and Nicolas Renard. A rapid and low noise switch from RANS to WMLES on curvilinear grids with compressible flow solvers. Journal of Computational Physics, 363(??):231–255, June 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301104.

Du:2015:FMG

[DWW15] Ning Du, Hong Wang, and Che Wang. A fast method for a generalized nonlocal elastic model. Journal of Computational Physics, 297(??):72–83, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003277.

Ding:2019:PPF

[DWZ19] Jie Ding, Zhongming Wang, and Shenggao Zhou. Positivity preserving finite difference methods for Poisson–Nernst– Planck equations with steric interactions: Application to slit-shaped nanopore conductance. Journal of Computational Physics, 397(??):Article 108864, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305558.

Denner:2018:PBA

[DXvW18] Fabian Denner, Cheng-Nian Xiao, and Berend G. M. van Wachem. Pressure-based algorithm for compressible interfacial flows with acoustically-conservative interface discretisation. Journal of Computational Physics, 367(??):192–234, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302535. See corri-

gendum [DvW19].

Diaz:2016:EDS

[DY16] Manuel A. Diaz and Jaw-Yen Yang. An efficient direct solver for rarefied gas flows with arbitrary statistics. Journal of Computational Physics, 305(??):127–149, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005835.

Du:2017:FAI

[DY17] Qiang Du and Jiang Yang. Fast and accurate implementation of Fourier spectral approximations of nonlocal diffusion operators and its applications. Journal of Computational Physics, 332(??):118–134, March 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306143.

Du:2019:MPP

[DY19a] Jie Du and Yang Yang. Maximum-principle-preserving thirdorder local discontinuous Galerkin method for convectiondiffusion equations on overlapping meshes. Journal of Computational Physics, 377(??):117–141, January 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307010.

Du:2019:TOC

[DY19b] Jie Du and Yang Yang. Third-order conservative signpreserving and steady-state-preserving time integrations and applications in stiff multispecies and multireaction detonations. Journal of Computational Physics, 395(??):489–510, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304486.

Du:2019:HOS

[DYL19] Ruilian Du, Yubin Yan, and Zongqi Liang. A high-order scheme to approximate the Caputo fractional derivative and its application to solve the fractional diffusion wave equation. Journal of Computational Physics, 376(??):1312–1330, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306685.

Dhakal:2016:MPM

[DZ16] Tilak R. Dhakal and Duan Z. Zhang. Material point methods applied to one-dimensional shock waves and dual domain material point method with sub-points. Journal of Computational Physics, 325(??):301–313, November 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303904.

Dhakal:2018:CDD

[DZ18] Tilak R. Dhakal and Duan Z. Zhang. Combining dual domain material point method with molecular dynamics for thermodynamic nonequilibriums. Journal of Computational Physics, 374(??):984–995, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305291.

Duan:2016:SAM

[DZC16] Beiping Duan, Zhoushun Zheng, and Wen Cao. Spectral approximation methods and error estimates for Caputo fractional derivative with applications to initial-value problems. Journal of Computational Physics, 319(??):108–128, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301541.

Darvish:2018:OHC

[DZR18] Amirashkan Darvish, Bijan Zakeri, and Nafiseh Radkani. An optimized hybrid Convolutional Perfectly Matched Layer for efficient absorption of electromagnetic waves. Journal of Computational Physics, 356(??):31–45, March 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308689.

El-Ajou:2015:AAS

[EAAM15] Ahmad El-Ajou, Omar Abu Arqub, and Shaher Momani. Approximate analytical solution of the nonlinear fractional KdV– Burgers equation: a new iterative algorithm. Journal of Computational Physics, 293(??):81–95, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005506.

Escribano:2015:MTS

[EARA15] Bruno Escribano, Elena Akhmatskaya, Sebastian Reich, and Jon M. Azpiroz. Multiple-time-stepping generalized hybrid Monte Carlo methods. Journal of Computational Physics, 280(??):1–20, January 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114006433.

Ellison:2015:CSI

[EBQ15] C. L. Ellison, J. W. Burby, and H. Qin. Comment on “Symplectic integration of magnetic systems”: a proof that the Boris algorithm is not variational. Journal of Computational Physics, 301(??):489–493, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005884. See [Web14].

Esmaeilbeigi:2018:FHI

[ECC18] M. Esmaeilbeigi, O. Chatrabgoun, and M. Cheraghi. Fractional Hermite interpolation using RBFs in high dimensions over irregular domains with application. Journal of Computational Physics, 375(??):1091–1120, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306077.

Errera:2016:CSC

[ED16] M.-P. Errera and F. Duchaine. Comparative study of coupling coefficients in Dirichlet–Robin procedure for fluidstructure aerothermal simulations. Journal of Computational Physics, 312(??):218–234, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116000760.

Edeling:2016:SSC

[EDC16] W. N. Edeling, R. P. Dwight, and P. Cinnella. Simplexstochastic collocation method with improved scalability. Journal of Computational Physics, 310(??):301–328, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008530.

Escalante:2019:EHR

[EDC19] C. Escalante, M. Dumbser, and M. J. Castro. An efficient hyperbolic relaxation system for dispersive non-hydrostatic water waves and its solution with high order discontinuous Galerkin schemes. Journal of Computational Physics, 394(??):385–416, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303730.

Eldred:2019:QHD

[EDK19] Christopher Eldred, Thomas Dubos, and Evaggelos Kritsikis. A quasi-Hamiltonian discretization of the thermal shallow water equations. Journal of Computational Physics, 379(??):1–31, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307058.

Evrard:2017:ECV

[EDvW17] Fabien Evrard, Fabian Denner, and Berend van Wachem. Estimation of curvature from volume fractions using parabolic reconstruction on two-dimensional unstructured meshes. Journal of Computational Physics, 351(??):271–294, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306940.

Eremin:2019:NSD

[EDW19] Yuri Eremin, Adrian Doicu, and Thomas Wriedt. The numerical scheme of the discrete sources method to analyze 3D plasmonic nanostructures accounting for the non-local effect. Journal of Computational Physics, 388(??):357–370, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302013.

Ezz-Eldien:2016:NQA

[EE16] S. S. Ezz-Eldien. New quadrature approach based on operational matrix for solving a class of fractional variational problems. Journal of Computational Physics, 317(??):362–381, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300912.

Errami:2015:DHB

[EEG^{+}15] Hassan Errami, Markus Eiswirth, Dima Grigoriev, Werner M. Seiler, Thomas Sturm, and Andreas Weber. Detection of Hopf bifurcations in chemical reaction networks using convex coordinates. Journal of Computational Physics, 291(??):279–302, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001400.

Engquist:2017:SIS

[EFHZ17] Björn Engquist, Christina Frederick, Quyen Huynh, and Haomin Zhou. Seafloor identification in sonar imagery via simulations of Helmholtz equations and discrete optimization. Journal of Computational Physics, 338(??):477–492, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301912.

Evans:2019:ANS

[EFO19] Jonathan D. Evans, Hugo L. França, and Cassio M. Oishi. Application of the natural stress formulation for solving unsteady viscoelastic contraction flows. Journal of Computational Physics, 388(??):462–489, July 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301627. See corrigendum [EFO20].

Evans:2020:CAN

[EFO20] Jonathan D. Evans, Hugo L. França, and Cassio M. Oishi. Corrigendum to “Application of the natural stress formulation for solving unsteady viscoelastic contraction flows” [J. Comput. Phys. 388 (462–489)]. Journal of Computational Physics, 404(??):Article 108693, March 1, 2020. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303511. See [EFO19].

Engquist:2015:FSM

[EFT15] Björn Engquist, Brittany D. Froese, and Yen-Hsi Richard Tsai. Fast sweeping methods for hyperbolic systems of conservation laws at steady state II. Journal of Computational Physics, 286(??):70–86, April 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000327.

Eslaminia:2016:DSP

[EG16] Mehran Eslaminia and Murthy N. Guddati. A double-sweeping preconditioner for the Helmholtz equation. Journal of Computational Physics, 314(??):800–823, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001741.

Egan:2017:GDM

[EG17] Raphael Egan and Frédéric Gibou. Geometric discretization of the multidimensional Dirac delta distribution — application to the Poisson equation with singular source terms. Journal of Computational Physics, 346(??):71–90, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304485.

Egan:2018:FSA

[EG18a] Raphael Egan and Frédéric Gibou. Fast and scalable algorithms for constructing solvent-excluded surfaces of large biomolecules. Journal of Computational Physics, 374(??): 91–120, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118304960.

Escalante:2018:GMB

[EG18b] José A. Morales Escalante and Irene M. Gamba. Galerkin methods for Boltzmann–Poisson transport with reflection conditions on rough boundaries. Journal of Computational Physics, 363(??):302–328, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301244.

Epstein:2019:HOW

[EGO19] Charles L. Epstein, Leslie Greengard, and Michael O’Neil. A high-order wideband direct solver for electromagnetic scattering from bodies of revolution. Journal of Computational Physics, 387(??):205–229, June 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301585.

Eca:2014:PEN

[EH14] L. Eça and M. Hoekstra. A procedure for the estimation of the numerical uncertainty of CFD calculations based on grid refinement studies. Journal of Computational Physics, 262(??):104– 130, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000278. See com-

ment [XS15] and reply [EH15].

Eca:2015:RCP

[EH15] L. Eça and M. Hoekstra. Reply to comment on “A procedure for the estimation of the numerical uncertainty of CFD calculations based on grid refinement studies” (L. Eça and M. Hoekstra, Journal of Computational Physics 262 (2014) 104–130). Journal of Computational Physics, 301(??):487–488, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500577X. See [EH14,

XS15].

Esmaily:2018:CST

[EH18] M. Esmaily and J. A. K. Horwitz. A correction scheme for twoway coupled point-particle simulations on anisotropic grids. Journal of Computational Physics, 375(??):960–982, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911830603X.

Endeve:2015:BPD

[EHXM15] Eirik Endeve, Cory D. Hauck, Yulong Xing, and Anthony Mezzacappa. Bound-preserving discontinuous Galerkin methods for conservative phase space advection in curvilinear coordinates. Journal of Computational Physics, 287 (??):151–183, 2015. CODEN JCTPAH. ISSN 0021-9991.

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000662.

Einkemmer:2019:PCS

[Ein19] Lukas Einkemmer. A performance comparison of semi-Lagrangian discontinuous Galerkin and spline based Vlasov solvers in four dimensions. Journal of Computational Physics, 376(??):937–951, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306697.

Esmaily:2018:SGM

[EJMI18] M. Esmaily, L. Jofre, A. Mani, and G. Iaccarino. A scalable geometric multigrid solver for nonsymmetric elliptic systems with application to variable-density flows. Journal of Computational Physics, 357(??):142–158, March 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730918X.

Esedoglu:2017:KPS

[EJZ17] Selim Esedoğlu, Matt Jacobs, and Pengbo Zhang. Kernels with prescribed surface tension & mobility for threshold dynamics schemes. Journal of Computational Physics, 337(??):62–83, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730116X.

Engsig-Karup:2016:SNS

[EKEB16] A. P. Engsig-Karup, C. Eskilsson, and D. Bigoni. A stabilised nodal spectral element method for fully nonlinear water waves. Journal of Computational Physics, 318(??):1–21, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301255.

Engels:2015:NSF

[EKSS15] Thomas Engels, Dmitry Kolomenskiy, Kai Schneider, and Jörn Sesterhenn. Numerical simulation of fluid-structure interaction with the volume penalization method. Journal of Computational Physics, 281(??):96–115, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006792.

Eca:2016:CVR

[EKV^{+}16] L. Eça, C. M. Klaij, G. Vaz, M. Hoekstra, and F. S. Pereira. On code verification of RANS solvers. Journal of Computational Physics, 310(??):418–439, April 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000036.

Ehrlacher:2017:DAT

[EL17] Virginie Ehrlacher and Damiano Lombardi. A dynamical adaptive tensor method for the Vlasov–Poisson system. Journal of Computational Physics, 339(??):285–306, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302024.

Eldred:2018:DAC

[EL18] Christopher Eldred and Daniel Y. Le Roux. Dispersion analysis of compatible Galerkin schemes for the 1D shallow water model. Journal of Computational Physics, 371(??):779–800, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303851.

Eldred:2019:DAC

[EL19] Christopher Eldred and Daniel Y. Le Roux. Dispersion analysis of compatible Galerkin schemes on quadrilaterals for shallow water models. Journal of Computational Physics, 387(??):539– 568, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301202.

Ervik:2016:MMS

[ELH^{+}16] Åsmund Ervik, Morten Olsen Lysgaard, Carmelo Herdes, Guadalupe Jiménez-Serratos, Erich A. Müller, Svend Tollak Munkejord, and Bernhard Müller. A multiscale method for simulating fluid interfaces covered with large molecules such as asphaltenes. Journal of Computational Physics, 327(??):576–611, December 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304600.

Edoh:2018:CAD

[EMM^{+}18] Ayaboe K. Edoh, Nathan L. Mundis, Charles L. Merkle, Ann R. Karagozian, and Venkateswaran Sankaran. Comparison of artificial-dissipation and solution-filtering stabilization schemes for time-accurate simulations. Journal of Computational Physics, 375(??):1424–1450, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305424.

Errera:2019:SSS

[EMS^{+}19] Marc-Paul Errera, Rocco Moretti, Rami Salem, Yohann Bachelier, Thomas Arrivé, and Minh Nguyen. A single stable scheme for steady conjugate heat transfer problems. Journal of Computational Physics, 394(??):491–502, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303808.

Exl:2017:EEM

[EMSS17] Lukas Exl, Norbert J. Mauser, Thomas Schrefl, and Dieter Suess. The extrapolated explicit midpoint scheme for variable order and step size controlled integration of the Landau–Lifschitz–Gilbert equation. Journal of Computational Physics, 346(??):14–24, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304503.

Exl:2016:AEC

[EMZ16] Lukas Exl, Norbert J. Mauser, and Yong Zhang. Accurate and efficient computation of nonlocal potentials based on Gaussian-sum approximation. Journal of Computational Physics, 327(??):629–642, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304661.

Eskandari:2017:TRM

[EN17] M. Eskandari and S. S. Nourazar. On the time relaxed Monte Carlo computations for the lid-driven micro cavity flow. Journal of Computational Physics, 343(??):355–367, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302048.

Eriksson:2018:FDS

[EN18] Sofia Eriksson and Jan Nordström. Finite difference schemes with transferable interfaces for parabolic problems. Journal of Computational Physics, 375(??):935–949, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305849.

Engwirda:2018:GPD

[Eng18] Darren Engwirda. Generalised primal–dual grids for unstructured co-volume schemes. Journal of Computational Physics, 375(??):155–176, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304868.

Einkemmer:2015:SAK

[EO15] Lukas Einkemmer and Alexander Ostermann. A splitting approach for the Kadomtsev–Petviashvili equation. Journal of Computational Physics, 299(??):716–730, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004696.

Eggl:2018:GBF

[ES18] M. F. Eggl and P. J. Schmid. A gradient-based framework for maximizing mixing in binary fluids. Journal of Computational Physics, 368(??):131–153, September 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302559.

Ebersohn:2017:KST

[ESGS17] Frans H. Ebersohn, J. P. Sheehan, Alec D. Gallimore, and John V. Shebalin. Kinetic simulation technique for plasma flow in strong external magnetic field. Journal of Computational Physics, 351(??):358–375, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306800.

Egerer:2016:EIM

[ESHA16] Christian P. Egerer, Steffen J. Schmidt, Stefan Hickel, and Nikolaus A. Adams. Efficient implicit LES method for the simulation of turbulent cavitating flows. Journal of Computational Physics, 316(??):453–469, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300626.

Egger:2017:EFM

[EST17] H. Egger, T. Seitz, and C. Tropea. Enhancement of flow measurements using fluid-dynamic constraints. Journal of Computational Physics, 344(??):558–574, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303686.

ElBouajaji:2015:QOD

[ETAG15] M. El Bouajaji, B. Thierry, X. Antoine, and C. Geuzaine. A quasi-optimal domain decomposition algorithm for the time-harmonic Maxwell’s equations. Journal of Computational Physics, 294(??):38–57, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001965.

Einkemmer:2017:PEI

[ETL17] Lukas Einkemmer, Mayya Tokman, and John Loffeld. On the performance of exponential integrators for problems in magnetohydrodynamics. Journal of Computational Physics, 330(??):550–565, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306131.

Evans:2018:NPD

[Eva18] B. Evans. Nano-particle drag prediction at low Reynolds number using a direct Boltzmann-BGK solution approach. Journal of Computational Physics, 352(??):123–141, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306988.

Ellam:2016:BAM

[EZG16] Louis Ellam, Nicholas Zabaras, and Mark Girolami. A Bayesian approach to multiscale inverse problems with onthe-fly scale determination. Journal of Computational Physics, 326(??):115–140, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303886.

Fu:2019:CMM

[FAC^{+}19] Shubin Fu, Robert Altmann, Eric T. Chung, Roland Maier, Daniel Petersheim, and Sai-Mang Pun. Computational multiscale methods for linear poroelasticity with high contrast. Journal of Computational Physics, 395(??):286–297, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304358.

Falissard:2015:UOD

[Fal15] F. Falissard. Uneven-order decentered Shapiro filters for boundary filtering. Journal of Computational Physics, 292 (??):168–175, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115001229.

Faliagas:2016:MWP

[Fal16] A. C. Faliagas. Mixed weak-perturbative solution method for Maxwell’s equations of diffusion with Müller’s partial stress tensor in the low velocity limit. Journal of Computational Physics, 308(??):322–346, March 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008657.

Falissard:2017:CLG

[Fal17] F. Falissard. Chebyshev-like generalized Shapiro filters for high-accuracy flow computations. Journal of Computational Physics, 336(??):595–607, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300839.

Fan:2016:SUC

[Fan16] Ping Fan. The standard upwind compact difference schemes for incompressible flow simulations. Journal of Computational Physics, 322(??):74–112, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302558.

Fan:2019:NFT

[Fan19] Ping Fan. A new front tracking method using particles with extended location information. Journal of Computational Physics, 378(??):497–547, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307265.

Fan:2019:FAI

[FAY19] Yuwei Fan, Jing An, and Lexing Ying. Fast algorithms for integral formulations of steady-state radiative transfer equation. Journal of Computational Physics, 380(??):191–211, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118308106.

Fattah:2016:PGQ

[FAZ16] Ryu Fattah, David Angland, and Xin Zhang. A priori grid quality estimation for high-order finite differencing. Journal of Computational Physics, 315(??):629–643, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300286.

Fleckenstein:2015:VFB

[FB15] Stefan Fleckenstein and Dieter Bothe. A volume-of-fluid-based numerical method for multi-component mass transfer with local volume changes. Journal of Computational Physics, 301(??):35–58, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005306.

Fakhari:2017:DIM

[FB17] Abbas Fakhari and Diogo Bolster. Diffuse interface modeling of three-phase contact line dynamics on curved boundaries: a lattice Boltzmann model for large density and viscosity ratios. Journal of Computational Physics, 334(??):620–638, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300414.

Froio:2016:DOA

[FBC^{+}16] A. Froio, R. Bonifetto, S. Carli, A. Quartararo, L. Savoldi, and R. Zanino. Design and optimization of Artificial Neural Networks for the modelling of superconducting magnets operation in tokamak fusion reactors. Journal of Computational Physics, 321(??):476–491, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301735.

Fernandes:2015:AEC

[FBF15] Ryan I. Fernandes, Bernard Bialecki, and Graeme Fairweather. An ADI extrapolated Crank–Nicolson orthogonal spline collocation method for nonlinear reaction-diffusion systems on evolving domains. Journal of Computational Physics, 299(??):561–580, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115004611.

Fortin:2015:MEA

[FBG15] A. Fortin, T. Briffard, and A. Garon. A more efficient anisotropic mesh adaptation for the computation of Lagrangian coherent structures. Journal of Computational Physics, 285(??):100–110, March 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000145.

Feng:2019:HRR

[FBJS19] Yongliang Feng, Pierre Boivin, Jérôme Jacob, and Pierre Sagaut. Hybrid recursive regularized thermal lattice Boltzmann model for high subsonic compressible flows. Journal of Computational Physics, 394(??):82–99, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303626.

Fakhari:2017:WMR

[FBL17] Abbas Fakhari, Diogo Bolster, and Li-Shi Luo. A weighted multiple-relaxation-time lattice Boltzmann method for multiphase flows and its application to partial coalescence cascades. Journal of Computational Physics, 341(??):22–43, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302632.

Flad:2016:SUT

[FBM16] David Flad, Andrea Beck, and Claus-Dieter Munz. Simulation of underresolved turbulent flows by adaptive filtering using the high order discontinuous Galerkin spectral element method. Journal of Computational Physics, 313(??):1–12, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500827X.

Flyer:2016:EFD

[FBW16] Natasha Flyer, Gregory A. Barnett, and Louis J. Wicker. Enhancing finite differences with radial basis functions: Experiments on the Navier–Stokes equations. Journal of Computational Physics, 316(??):39–62, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300195.

Fan:2019:BNN

[FBY19] Yuwei Fan, Cindy Orozco Bohorquez, and Lexing Ying. BCR-Net: a neural network based on the nonstandard wavelet form. Journal of Computational Physics, 384(??):1–15, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300762.

Fusi:2016:ASE

[FC16]  F. Fusi and P. M. Congedo.  An adaptive strategy

on the error of the objective functions for uncertaintybased derivative-free optimization. Journal of Computational Physics, 309(??):241–266, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911600005X.

Franci:2019:RRG

[FC19a] Alessandro Franci and Massimiliano Cremonesi. 3D regularized μ(I)-rheology for granular flows simulation. Journal of Computational Physics, 378(??):257–277, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307290.

Fu:2019:LGM

[FC19b] Shubin Fu and Eric T. Chung. A local-global multiscale mortar mixed finite element method for multiphase transport in heterogeneous media. Journal of Computational Physics, 399(??):Article 108906, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306114.

Feng:2017:SNI

[FCL17] Hui Feng, Xiangyang Cui, and Guangyao Li. A stable nodal integration method for static and quasi-static electromagnetic field computation. Journal of Computational Physics, 336(??):580–594, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301158.

Fu:2019:EMM

[FCL19] Shubin Fu, Eric Chung, and Guanglian Li. Edge multiscale methods for elliptic problems with heterogeneous coefficients. Journal of Computational Physics, 396(??):228–242, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304139.

Flaig:2018:SMP

[FCW^{+}18] M. Flaig, D. Clark, C. Weber, D. L. Youngs, and B. Thornber. Single-mode perturbation growth in an idealized spherical implosion. Journal of Computational Physics, 371(??):801–819, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303929.

Fairbanks:2017:LRC

[FDKI17] Hillary R. Fairbanks, Alireza Doostan, Christian Ketelsen, and Gianluca Iaccarino. A low-rank control variate for multilevel Monte Carlo simulation of high-dimensional uncertain systems. Journal of Computational Physics, 341(??):121–139, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302619.

Farina:2015:RSC

[FDS^{+}15] R. Farina, S. Dobricic, A. Storto, S. Masina, and S. Cuomo. A revised scheme to compute horizontal covariances in an oceanographic 3D-VAR assimilation system. Journal of Computational Physics, 284(??):631–647, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000042.

Fedeli:2017:CSP

[Fed17] Livio Fedeli. Computer simulations of phase field drops on super-hydrophobic surfaces. Journal of Computational Physics, 344(??):247–259, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911730356X.

Ferrer:2017:IPS

[Fer17] Esteban Ferrer. An interior penalty stabilised incompressible discontinuous Galerkin-Fourier solver for implicit large eddy simulations. Journal of Computational Physics, 348 (??):754–775, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117305570.

Flyer:2016:RPR

[FFBB16] Natasha Flyer, Bengt Fornberg, Victor Bayona, and Gregory A. Barnett. On the role of polynomials in RBF-FD approximations: I. Interpolation and accuracy. Journal of Computational Physics, 321(??):21–38, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301632.

Ferronato:2019:GPF

[FFJ^{+}19] Massimiliano Ferronato, Andrea Franceschini, Carlo Janna, Nicola Castelletto, and Hamdi A. Tchelepi. A general preconditioning framework for coupled multiphysics problems with application to contact- and poro-mechanics. Journal of Computational Physics, 398(??):Article 108887, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305856.

Franceschini:2016:NLA

[FFJT16] Andrea Franceschini, Massimiliano Ferronato, Carlo Janna, and Pietro Teatini. A novel Lagrangian approach for the stable numerical simulation of fault and fracture mechanics. Journal of Computational Physics, 314(??):503–521, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001844.

Fatone:2019:AOT

[FFM19] L. Fatone, D. Funaro, and G. Manzini. Arbitrary-order time-accurate semi-Lagrangian spectral approximations of the Vlasov–Poisson system. Journal of Computational Physics, 384(??):349–375, May 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300452.

Fasondini:2017:MCM

[FFW17] Marco Fasondini, Bengt Fornberg, and J. A. C. Weideman. Methods for the computation of the multivalued Painlevé transcendents on their Riemann surfaces. Journal of Computational Physics, 344(??):36–50, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303595.

Feldman:2016:EIB

[FG16] Yuri Feldman and Yosef Gulberg. An extension of the immersed boundary method based on the distributed Lagrange multiplier approach. Journal of Computational Physics, 322(??):248–266, October 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302649.

Flad:2017:UKE

[FG17] David Flad and Gregor Gassner. On the use of kinetic energy preserving DG-schemes for large eddy simulation. Journal of Computational Physics, 350(??):782–795, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730654X.

Freitag:2018:LRA

[FG18] Melina A. Freitag and Daniel L. H. Green. A lowrank approach to the solution of weak constraint variational data assimilation problems. Journal of Computational Physics, 357(??):263–281, March 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309336.

Frantzis:2019:EMT

[FG19] C. Frantzis and D. G. E. Grigoriadis. An efficient method for two-fluid incompressible flows appropriate for the immersed boundary method. Journal of Computational Physics, 376(??):28–53, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306375.

Fakhari:2016:MCL

[FGL16] Abbas Fakhari, Martin Geier, and Taehun Lee. A massconserving lattice Boltzmann method with dynamic grid refinement for immiscible two-phase flows. Journal of Computational Physics, 315(??):434–457, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300237.

Fillion-Gourdeau:2016:GMU

[FGLB16] F. Fillion-Gourdeau, E. Lorin, and A. D. Bandrauk. Galerkin method for unsplit 3-D Dirac equation using atomically/ kinetically balanced B-spline basis. Journal of Computational Physics, 307(??):122–145, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007640.

Feng:2018:MCA

[FGLW18] Wenqiang Feng, Zhenlin Guo, John S. Lowengrub, and Steven M. Wise. A mass-conservative adaptive FAS multigrid solver for cell-centered finite difference methods on block-structured, locally-cartesian grids. Journal of Computational Physics, 352(??):463–497, January 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307404.

Faugeras:2017:FBC

[FH17] Blaise Faugeras and Holger Heumann. FEM–BEM coupling methods for tokamak plasma axisymmetric free-boundary equilibrium computations in unbounded domains. Journal of Computational Physics, 343(??):201–216, August 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303261.

Fu:2016:FHO

[FHA16] Lin Fu, Xiangyu Y. Hu, and Nikolaus A. Adams. A family of high-order targeted ENO schemes for compressible-fluid simulations. Journal of Computational Physics, 305(??):333–359, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007147.

Fu:2017:PMC

[FHA17a] Lin Fu, Xiangyu Y. Hu, and Nikolaus A. Adams. A physicsmotivated Centroidal Voronoi Particle domain decomposition method. Journal of Computational Physics, 335(??):718–735, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300670.

Fu:2017:TES

[FHA17b] Lin Fu, Xiangyu Y. Hu, and Nikolaus A. Adams. Targeted ENO schemes with tailored resolution property for hyperbolic conservation laws. Journal of Computational Physics, 349(??):97–121, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305624.

Fu:2018:NCA

[FHA18] Lin Fu, Xiangyu Y. Hu, and Nikolaus A. Adams. A new class of adaptive high-order targeted ENO schemes for hyperbolic conservation laws. Journal of Computational Physics, 374(??):724–751, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305047.

Flegg:2015:CMC

[FHE15] Mark B. Flegg, Stefan Hellander, and Radek Erban. Convergence of methods for coupling of microscopic and mesoscopic reaction-diffusion simulations. Journal of Computational Physics, 289(??):1–17, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000340.

Fath:2017:FMI

[FHS17] L. Fath, M. Hochbruck, and C. V. Singh. A fast mollified impulse method for biomolecular atomistic simulations. Journal of Computational Physics, 333(??):180–198, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306787.

Fang:2019:ERB

[FHY^{+}19] Hong Fang, Yikun Hu, Caihui Yu, Ming Tie, Jie Liu, and Chunye Gong. An efficient radial basis functions mesh deformation with greedy algorithm based on recurrence Choleskey decomposition and parallel computing. Journal of Computational Physics, 377(??):183–199, January 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830696X.

Fidkowski:2017:OBS

[Fid17] Krzysztof J. Fidkowski. Output-based space-time mesh optimization for unsteady flows using continuous-in-time adjoints. Journal of Computational Physics, 341(??):258–277, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302760.

Forgues:2019:GMM

[FITcD19] François Forgues, Lucian Ivan, Alexandre Trottier, and James G. M. c Donald. A Gaussian moment method for polydisperse multiphase flow modelling. Journal of Computational Physics, 398(??):Article 108839, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305236.

Ferreira:2018:WLS

[FJLC18] L. S. Ferreira, L. N. Jorge, S. A. Leão, and A. A. Caparica. Wang–Landau sampling: Saving CPU time. Journal of Computational Physics, 358(??):130–134, April 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830007X.

Franck:2017:MHD

[FK17] I. M. Franck and P. S. Koutsourelakis. Multimodal, high-dimensional, model-based, Bayesian inverse problems with applications in biomechanics. Journal of Computational Physics, 329(??):91–125, January 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630537X.

Foy:2019:CTS

[FK19] Brody H. Foy and David Kay. A computationally tractable scheme for simulation of the human pulmonary system. Journal of Computational Physics, 388(??):371–393, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301974.

Fuentes:2017:CVF

[FKDL17] Federico Fuentes, Brendan Keith, Leszek Demkowicz, and Patrick Le Tallec. Coupled variational formulations of linear elasticity and the DPG methodology. Journal of Computational Physics, 348(??):715–731, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305594.

Farrell:2017:CAC

[FKF17] Patricio Farrell, Thomas Koprucki, and Jürgen Fuhrmann. Computational and analytical comparison of flux discretizations for the semiconductor device equations beyond Boltzmann statistics. Journal of Computational Physics, 346 (??):497–513, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117304680.

Fujimoto:2019:CED

[FKK19] Takeshi R. Fujimoto, Taro Kawasaki, and Keiichi Kitamura. Canny-edge-detection/Rankine–Hugoniot-conditions unified shock sensor for inviscid and viscous flows. Journal of Computational Physics, 396(??):264–279, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304796.

Filippini:2016:FGN

[FKR16] A. G. Filippini, M. Kazolea, and M. Ricchiuto. A flexible genuinely nonlinear approach for nonlinear wave propagation, breaking and run-up. Journal of Computational Physics, 310(??):381–417, April 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116000358.

Fumagalli:2019:CNC

[FKS19] Alessio Fumagalli, Eirik Keilegavlen, and Stefano Scialò. Conforming, non-conforming and non-matching discretization couplings in discrete fracture network simulations. Journal of Computational Physics, 376(??):694–712, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306508.

Fujita:2015:DSD

[FKY15] Masahiro Fujita, Osamu Koike, and Yukio Yamaguchi. Direct simulation of drying colloidal suspension on substrate using immersed free surface model. Journal of Computational Physics, 281(??):421–448, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007244.

Fu:2016:CCF

[FL16] Kai Fu and Dong Liang. The conservative characteristic FD methods for atmospheric aerosol transport problems. Journal of Computational Physics, 305(??):494–520, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007263.

Febres:2018:EFT

[FL18] Mijail Febres and Dominique Legendre. Enhancement of a 2D front-tracking algorithm with a non-uniform distribution of Lagrangian markers. Journal of Computational Physics, 358(??):173–200, April 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117309075.

Fu:2017:NPM

[FLHA17] Lin Fu, Sergej Litvinov, Xiangyu Y. Hu, and Nikolaus A. Adams. A novel partitioning method for blockstructured adaptive meshes. Journal of Computational Physics, 341(??):447–473, July 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306027.

Foster:2017:SPS

[FLT17] Erich L. Foster, Jérôme Lohéac, and Minh-Binh Tran. A structure preserving scheme for the Kolmogorov–Fokker–Planck equation. Journal of Computational Physics, 330(??):319–339, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305940.

Fryklund:2018:PUE

[FLT18] Fredrik Fryklund, Erik Lehto, and Anna-Karin Tornberg. Partition of unity extension of functions on complex domains. Journal of Computational Physics, 375(??):57–79, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305345.

Fernandez:2015:FDT

[FLV15] Miguel A. Fernández, Mikel Landajuela, and Marina Vidrascu. Fully decoupled time-marching schemes for incompressible fluid/thin-walled structure interaction. Journal of Computational Physics, 297(??):156–181, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003289.

Fox:2018:CHQ

[FLV18] Rodney O. Fox, Frédérique Laurent, and Aymeric Vié. Conditional hyperbolic quadrature method of moments for kinetic equations. Journal of Computational Physics, 365(??):269–293, July 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301852.

Fridrich:2016:SCC

[FLW16] David Fridrich, Richard Liska, and Burton Wendroff. Some cell-centered Lagrangian Lax–Wendroff HLL hybrid schemes. Journal of Computational Physics, 326(??):878–892, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304363.

Fu:2019:FVM

[FLW19] Hongfei Fu, Huan Liu, and Hong Wang. A finite volume method for two-dimensional Riemann–Liouville spacefractional diffusion equation and its efficient implementation. Journal of Computational Physics, 388(??):316–334, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911930213X.

Fan:2015:TDS

[FM15] Li Fan and Peter Monk. Time dependent scattering from a grating. Journal of Computational Physics, 302(??):97–113, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005677.

Fick:2018:SPM

[FMPT18] Lambert Fick, Yvon Maday, Anthony T. Patera, and Tommaso Taddei. A stabilized POD model for turbulent flows over a range of Reynolds numbers: Optimal parameter sampling and constrained projection. Journal of Computational Physics, 371(??):214–243, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303279.

Fechter:2017:SIM

[FMRZ17] Stefan Fechter, Claus-Dieter Munz, Christian Rohde, and Christoph Zeiler. A sharp interface method for compressible liquid-vapor flow with phase transition and surface tension. Journal of Computational Physics, 336(??):347–374, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300943.

Frenander:2017:CNR

[FN17] Hannes Frenander and Jan Nordström. Constructing nonreflecting boundary conditions using summation-by-parts in time. Journal of Computational Physics, 331(??):38–48, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306301.

Fernandez-Nieto:2018:GFR

[FNGDMNR18] E. D. Fernández-Nieto, J. Garres-Díaz, A. Mangeney, and G. Narbona-Reina. 2D granular flows with the μ(I) rheology and side walls friction: a well-balanced multilayer discretization. Journal of Computational Physics, 356(??):192–219, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308823.

Fernandez-Nieto:2018:ENS

[FNGV18] Enrique D. Fernández-Nieto, José M. Gallardo, and Paul Vigneaux. Efficient numerical schemes for viscoplastic avalanches. Part 2: the 2D case. Journal of Computational Physics, 353(??):460–490, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307271.

Fourno:2019:FNC

[FNNB19] André Fourno, Tri-Dat Ngo, Benoit Noetinger, and Christian La Borderie. FraC: a new conforming mesh method for discrete fracture networks. Journal of Computational Physics, 376(??):713–732, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306624.

Feng:2019:TSF

[FNNW19] Dianlei Feng, Insa Neuweiler, Udo Nackenhorst, and Thomas Wick. A time-space flux-corrected transport finite element formulation for solving multi-dimensional advection– diffusion-reaction equations. Journal of Computational Physics, 396(??):31–53, November 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119304619.

Fernandez:2017:HDG

[FNP17] P. Fernandez, N. C. Nguyen, and J. Peraire. The hybridized discontinuous Galerkin method for Implicit Large-Eddy Simulation of transitional turbulent flows. Journal of Computational Physics, 336(??):308–329, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301080.

Farrell:2015:BFA

[FOF15] Kathryn Farrell, J. Tinsley Oden, and Danial Faghihi. A Bayesian framework for adaptive selection, calibration, and validation of coarse-grained models of atomistic systems. Journal of Computational Physics, 295(??):189–208, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002430.

Font:2016:CAO

[Fon16] Gabriel I. Font. Computational acceleration of orbital neutral sensor ionizer simulation through phenomena separation. Journal of Computational Physics, 316(??):1–9, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001327.

Fornberg:2016:FCL

[For16] Bengt Fornberg. Fast calculation of Laurent expansions for matrix inverses. Journal of Computational Physics, 326 (??):722–732, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304429.

Fortunato:2016:HOU

[FP16] Meire Fortunato and Per-Olof Persson. High-order unstructured curved mesh generation using the Winslow equations. Journal of Computational Physics, 307(??):1–14, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007603.

Fuster:2018:AMM

[FP18] Daniel Fuster and Stéphane Popinet. An all-Mach method for the simulation of bubble dynamics problems in the presence of surface tension. Journal of Computational Physics, 374(??):752–768, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305187.

Fernandez-Pendas:2016:AMS

[FPASS16] Mario Fernández-Pendás, Elena Akhmatskaya, and J. M. Sanz-Serna. Adaptive multi-stage integrators for optimal energy conservation in molecular simulations. Journal of Computational Physics, 327(??):434–449, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304569.

Fraggedakis:2017:DTD

[FPDT17] D. Fraggedakis, J. Papaioannou, Y. Dimakopoulos, and J. Tsamopoulos. Discretization of three-dimensional free surface flows and moving boundary problems via elliptic grid methods based on variational principles. Journal of Computational Physics, 344(??):127–150, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730339X.

Foy:2017:MFV

[FPT17] Brody H. Foy, Patrick Perré, and Ian Turner. The Meshfree Finite Volume Method with application to multiphase porous media models. Journal of Computational Physics, 333(??):369–386, March 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116307094.

Fumagalli:2018:FSP

[FPV18] Ivan Fumagalli, Nicola Parolini, and Marco Verani. On a free-surface problem with moving contact line: From variational principles to stable numerical approximations. Journal of Computational Physics, 355(??):253–284, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308318.

Fang:2018:HAS

[FQZNZ18] Jun Fang, Jianliang Qian, Leonardo Zepeda-Núñez, and Hongkai Zhao. A hybrid approach to solve the highfrequency Helmholtz equation with source singularity in smooth heterogeneous media. Journal of Computational Physics, 371(??):261–279, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301633.

Fai:2018:LIB

[FR18] Thomas G. Fai and Chris H. Rycroft. Lubricated immersed boundary method in two dimensions. Journal of Computational Physics, 356(??):319–339, March 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308677.

Fan:2015:ACM

[FRL15] Houfu Fan, Bo Ren, and Shaofan Li. An adhesive contact mechanics formulation based on atomistically induced surface traction. Journal of Computational Physics, 302(??):420–438, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005628.

Frontiere:2017:CCR

[FRO17] Nicholas Frontiere, Cody D. Raskin, and J. Michael Owen. CRKSPH — a Conservative Reproducing Kernel Smoothed Particle Hydrodynamics Scheme. Journal of Computational Physics, 332(??):160–209, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306453.

Fraysse:2016:UMB

[FRRV16] F. Fraysse, C. Redondo, G. Rubio, and E. Valero. Upwind methods for the Baer–Nunziato equations and higher-order reconstruction using artificial viscosity. Journal of Computational Physics, 326(??):805–827, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304314.

Frei:2016:LTS

[FRW16] S. Frei, T. Richter, and T. Wick. Long-term simulation of large deformation, mechano-chemical fluid-structure interactions in ALE and fully Eulerian coordinates. Journal of Computational Physics, 321(??):874–891, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302406.

Fu:2015:HFE

[FS15] Yao Fu and Jeong-Hoon Song. Heat flux expressions that satisfy the conservation laws in atomistic system involving multibody potentials. Journal of Computational Physics, 294(??):191–207, August 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115002053.

Falletta:2016:PCM

[FS16] Silvia Falletta and Stefan A. Sauter. The panel-clustering method for the wave equation in two spatial dimensions. Journal of Computational Physics, 305(??):217–243, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007093.

Farazmand:2017:ROP

[FS17a] Mohammad Farazmand and Themistoklis P. Sapsis. Reducedorder prediction of rogue waves in two-dimensional deep-water waves. Journal of Computational Physics, 340(??):418–434, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302553.

Fu:2017:NTC

[FS17b] Guosheng Fu and Chi-Wang Shu. A new troubled-cell indicator for discontinuous Galerkin methods for hyperbolic conservation laws. Journal of Computational Physics, 347(??):305–327, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305004.

Felker:2018:FOA

[FS18] Kyle Gerard Felker and James M. Stone. A fourthorder accurate finite volume method for ideal MHD via upwind constrained transport. Journal of Computational Physics, 375(??):1365–1400, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305485.

Fraysse:2019:ADU

[FS19a] François Fraysse and Richard Saurel. Automatic differentiation using operator overloading (ADOO) for implicit resolution of hyperbolic single phase and two-phase flow models. Journal of Computational Physics, 399(??):Article 108942, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119306473.

Fu:2019:OEC

[FS19b] Guosheng Fu and Chi-Wang Shu. Optimal energy-conserving discontinuous Galerkin methods for linear symmetric hyperbolic systems. Journal of Computational Physics, 394 (??):329–363, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119304024.

Franchina:2016:MGF

[FSB16] N. Franchina, M. Savini, and F. Bassi. Multicomponent gas flow computations by a discontinuous Galerkin scheme using L^{2}-projection of perfect gas EOS. Journal of Computational Physics, 315(??):302–322, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300249.

Frederix:2016:CBS

[FSK^{+}16] E. M. A. Frederix, M. Stanic, A. K. Kuczaj, M. Nordlund, and B. J. Geurts. Characteristics-based sectional modeling of aerosol nucleation and condensation. Journal of Computational Physics, 326(??):499–515, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630417X.

Furtmaier:2016:SSM

[FSM16] O. Furtmaier, S. Succi, and M. Mendoza. Semi-spectral method for the Wigner equation. Journal of Computational Physics, 305(??):1015–1036, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007639.

Feng:2015:TDL

[FST15] Yongliang Feng, Pierre Sagaut, and Wenquan Tao. A three dimensional lattice model for thermal compressible flow on standard lattices. Journal of Computational Physics, 303(??):514–529, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005926.

Feng:2017:PSD

[FSWW17] Wenqiang Feng, Abner J. Salgado, Cheng Wang, and Steven M. Wise. Preconditioned steepest descent methods for some nonlinear elliptic equations involving p-Laplacian terms. Journal of Computational Physics, 334(??):45–67, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116307100.

Fujita:2019:HNC

[Fuj19] Kazuhiro Fujita. Hybrid Newmark-conformal FDTD modeling of thin spoof plasmonic metamaterials. Journal of Computational Physics, 376(??):390–410, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306521.

Fernandez:2017:LSS

[FW17] P. Fernandez and Q. Wang. Lyapunov spectrum of the separated flow around the NACA 0012 airfoil and its dependence on numerical discretization. Journal of Computational Physics, 350(??):453–469, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730637X.

Failer:2018:ATS

[FW18] Lukas Failer and Thomas Wick. Adaptive time-step control for nonlinear fluid-structure interaction. Journal of Computational Physics, 366(??):448–477, August 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302377.

Fehn:2017:SPM

[FWK17] Niklas Fehn, Wolfgang A. Wall, and Martin Kronbichler. On the stability of projection methods for the incompressible Navier–Stokes equations based on high-order discontinuous Galerkin discretizations. Journal of Computational Physics, 351(??):392–421, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306915.

Fehn:2018:RED

[FWK18] Niklas Fehn, Wolfgang A. Wall, and Martin Kronbichler. Robust and efficient discontinuous Galerkin methods for underresolved turbulent incompressible flows. Journal of Computational Physics, 372(??):667–693, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304157.

Filbet:2018:HDG

[FX18] Francis Filbet and Tao Xiong. A hybrid discontinuous Galerkin scheme for multi-scale kinetic equations. Journal of Computational Physics, 372(??):841–863, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830442X.

Fujii:2018:AEC

[FYC^{+}18] Hiroyuki Fujii, Yukio Yamada, Go Chiba, Yoko Hoshi, Kazumichi Kobayashi, and Masao Watanabe. Accurate and efficient computation of the 3D radiative transfer equation in highly forward-peaked scattering media using a renormalization approach. Journal of Computational Physics, 374(??):591–604, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305084.

Fukumitsu:2015:NDS

[FYO^{+}15] Kohei Fukumitsu, Takashi Yabe, Youichi Ogata, Tetsu Oami, and Tomomasa Ohkubo. A new directional-splitting CIP interpolation with high accuracy and low memory consumption. Journal of Computational Physics, 286(??):62–69, April 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008687.

Feng:2015:SOP

[FYZ^{+}15] Naixing Feng, Yongqing Yue, Chunhui Zhu, Liangtian Wan, and Qing Huo Liu. Second-order PML: Optimal choice of n^{th}-order PML for truncating FDTD domains. Journal of Computational Physics, 285(??):71–83, March 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000194.

Frachon:2019:CFE

[FZ19] Thomas Frachon and Sara Zahedi. A cut finite element method for incompressible two-phase Navier–Stokes flows. Journal of Computational Physics, 384(??):77–98, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300798.

Gross:2018:HFC

[GA18] B. J. Gross and P. J. Atzberger. Hydrodynamic flows on curved surfaces: Spectral numerical methods for radial manifold shapes. Journal of Computational Physics, 371(??):663–689, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303917.

Gimenez:2019:SOT

[GAIN19] Juan M. Gimenez, Horacio J. Aguerre, Sergio R. Idelsohn, and Norberto M. Nigro. A second-order in time and space particle-based method to solve flow problems on arbitrary meshes. Journal of Computational Physics, 380(??):295–310, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307782.

Gorji:2015:VRF

[GAJ15] M. Hossein Gorji, Nemanja Andric, and Patrick Jenny. Variance reduction for Fokker–Planck based particle Monte Carlo schemes. Journal of Computational Physics, 295(??):644–664, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002545.

Gambaruto:2015:CHS

[Gam15] Alberto M. Gambaruto. Computational haemodynamics of small vessels using the Moving Particle Semi-implicit (MPS) method. Journal of Computational Physics, 302(??):68–96, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005665.

Ganesan:2015:SID

[Gan15] Sashikumaar Ganesan. Simulations of impinging droplets with surfactant-dependent dynamic contact angle. Journal of Computational Physics, 301(??):178–200, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005537.

Gagarina:2016:VST

[GAN^{+}16] E. Gagarina, V. R. Ambati, S. Nurijanyan, J. J. W. van der Vegt, and O. Bokhove. On variational and symplectic time integrators for Hamiltonian systems. Journal of Computational Physics, 306(??):370–389, February 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007895.

Guiraldello:2018:MRC

[GAS^{+}18] Rafael T. Guiraldello, Roberto F. Ausas, Fabricio S. Sousa, Felipe Pereira, and Gustavo C. Buscaglia. The Multiscale Robin Coupled Method for flows in porous media. Journal of Computational Physics, 355(??):1–21, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730829X.

Galitzine:2015:APN

[GB15a] Cyril Galitzine and Iain D. Boyd. An adaptive procedure for the numerical parameters of a particle simulation. Journal of Computational Physics, 281(??):449–472, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007268.

Galitzine:2015:ACD

[GB15b] Cyril Galitzine and Iain D. Boyd. An analysis of the convergence of the direct simulation Monte Carlo method. Journal of Computational Physics, 289(??):196–223, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001096.

Giorgiani:2018:HDG

[GBC^{+}18] G. Giorgiani, H. Bufferand, G. Ciraolo, P. Ghendrih, F. Schwander, E. Serre, and P. Tamain. A hybrid discontinuous Galerkin method for tokamak edge plasma simulations in global realistic geometry. Journal of Computational Physics, 374(??):515–532, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304893.

Golbert:2015:SMS

[GBCF15] D. R. Golbert, P. J. Blanco, A. Clausse, and R. A. Feijóo. On the search of more stable second-order lattice-Boltzmann schemes in confined flows. Journal of Computational Physics, 294(??):605–618, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999115002260. See corrigendum [GBCF16].

Golbert:2016:CSM

[GBCF16] D. R. Golbert, P. J. Blanco, A. Clausse, and R. A. Feijóo. Corrigendum to “On the search of more stable second-order lattice-Boltzmann schemes in confined flows” [J. Comp. Phys. 294 (2015) 605–618]. Journal of Computational Physics, 311(??):374, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116000139. See [GBCF15].

Goffin:2015:GBA

[GBD^{+}15] Mark A. Goffin, Andrew G. Buchan, Steven Dargaville, Christopher C. Pain, Paul N. Smith, and Richard P. Smedley-Stevenson. Goal-based angular adaptivity applied to a waveletbased discretisation of the neutral particle transport equation. Journal of Computational Physics, 281(??):1032–1062, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007517.

Godoy:2017:DNF

[GBD17] Eduardo Godoy, Valeria Boccardo, and Mario Durán. A Dirichlet-to-Neumann finite element method for axisymmetric elastostatics in a semi-infinite domain. Journal of Computational Physics, 328(??):1–26, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304934.

Georges:2016:GCC

[GBM16] Gabriel Georges, Jérôme Breil, and Pierre-Henri Maire. A 3D GCL compatible cell-centered Lagrangian scheme for solving gas dynamics equations. Journal of Computational Physics, 305(??):921–941, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007172.

Garain:2015:CCF

[GBR15] Sudip Garain, Dinshaw S. Balsara, and John Reid. Comparing Coarray Fortran (CAF) with MPI for several structured mesh PDE applications. Journal of Computational Physics, 297(??):237–253, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911500354X.

Gravenkamp:2015:SEG

[GBS15] Hauke Gravenkamp, Carolin Birk, and Chongmin Song. Simulation of elastic guided waves interacting with defects in arbitrarily long structures using the Scaled Boundary Finite Element Method. Journal of Computational Physics, 295(??):438–455, August 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115002880.

Guo:2015:LMC

[GBU15] Gang Guo, Arne Bittig, and Adelinde Uhrmacher. Lattice Monte Carlo simulation of Galilei variant anomalous diffusion. Journal of Computational Physics, 288(??):167–180, May 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500087X.

Gillebaart:2016:ARB

[GBvZB16] T. Gillebaart, D. S. Blom, A. H. van Zuijlen, and H. Bijl. Adaptive radial basis function mesh deformation using data reduction. Journal of Computational Physics, 321(??):997– 1025, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116301814.

Goza:2017:SCI

[GC17] Andres Goza and Tim Colonius. A strongly-coupled immersedboundary formulation for thin elastic structures. Journal of Computational Physics, 336(??):401–411, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301201.

Gao:2019:SCH

[GCI19] Jing Gao, Marissa Condon, and Arieh Iserles. Spectral computation of highly oscillatory integral equations in laser theory. Journal of Computational Physics, 395(??):351–381, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S002199911930453X.

Gonzalez-Calderon:2018:AMT

[GCVCHH18] Alfredo González-Calderón, Luis X. Vivas-Cruz, and Erik César Herrera-Hernández. Application of the θ-method to a telegraphic model of fluid flow in a dual-porosity medium. Journal of Computational Physics, 352(??):426–444, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306733.

Gerhard:2015:MBG

[GCVMK15] Nils Gerhard, Daniel Caviedes-Voullième, Siegfried Müller, and Georges Kesserwani. Multiwavelet-based grid adaptation with discontinuous Galerkin schemes for shallow water equations. Journal of Computational Physics, 301(??):265–288, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005574.

Golub:2019:BIE

[GD19] Mikhail V. Golub and Olga V. Doroshenko. Boundary integral equation method for simulation scattering of elastic waves obliquely incident to a doubly periodic array of interface delaminations. Journal of Computational Physics, 376(??):675–693, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306600.

Guddati:2016:ECT

[GDA16] Murthy N. Guddati, Vladimir Druskin, and Ali Vaziri Astaneh. Exponential convergence through linear finite element discretization of stratified subdomains. Journal of Computational Physics, 322(??):429–447, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302704.

Ghigo:2017:LSH

[GDFL17] A. R. Ghigo, O. Delestre, J.-M. Fullana, and P.-Y. Lagrée. Low-Shapiro hydrostatic reconstruction technique for blood flow simulation in large arteries with varying geometrical and mechanical properties. Journal of Computational Physics, 331(??):108–136, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306246.

Ghoos:2016:ACC

[GDS^{+}16] K. Ghoos, W. Dekeyser, G. Samaey, P. Börner, and M. Baelmans. Accuracy and convergence of coupled finite-volume/Monte Carlo codes for plasma edge simulations of nuclear fusion reactors. Journal of Computational Physics, 322(??):162–182, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302741.

Gentile:2011:IET

[Gen11] N. A. Gentile. Including the effects of temperature-dependent opacities in the implicit Monte Carlo algorithm. Journal of Computational Physics, 230(12):5100–5114, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100180X. See comments

[Gho17].

Geng:2015:AIP

[Gen15] Hua Y. Geng. Accelerating ab initio path integral molecular dynamics with multilevel sampling of potential surface. Journal of Computational Physics, 283(??):299–311, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008146.

Gerlinger:2017:LTM

[Ger17] Peter Gerlinger. Lagrangian transported MDF methods for compressible high speed flows. Journal of Computational Physics, 339(??):68–95, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301523.

Greene:2016:HOM

[GEZK16] Patrick T. Greene, Jeff D. Eldredge, Xiaolin Zhong, and John Kim. A high-order multi-zone cut-stencil method for numerical simulations of high-speed flows over complex geometries. Journal of Computational Physics, 316(??):652–681, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630078X.

Germaschewski:2016:PSC

[GFA^{+}16] Kai Germaschewski, William Fox, Stephen Abbott, Narges Ahmadi, Kristofor Maynard, Liang Wang, Hartmut Ruhl, and Amitava Bhattacharjee. The Plasma Simulation Code: a modern particle-in-cell code with patch-based load-balancing. Journal of Computational Physics, 318(??):305–326, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301413.

Gao:2018:HOM

[GFC18] Kai Gao, Shubin Fu, and Eric T. Chung. A high-order multiscale finite-element method for time-domain acoustic-wave modeling. Journal of Computational Physics, 360(??):120–136, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300421.

Gao:2015:GMF

[GFG^{+}15] Kai Gao, Shubin Fu, Richard L. Gibson, Jr., Eric T. Chung, and Yalchin Efendiev. Generalized Multiscale Finite-Element Method (GMsFEM) for elastic wave propagation in heterogeneous, anisotropic media. Journal of Computational Physics, 295(??):161–188, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002405.

Ghigo:2017:NMM

[GFL17] Arthur R. Ghigo, Jose-Maria Fullana, and Pierre-Yves Lagrée. A 2D nonlinear multiring model for blood flow in large elastic arteries. Journal of Computational Physics, 350(??):136–165, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306162.

Gibou:2018:RLS

[GFO18] Frederic Gibou, Ronald Fedkiw, and Stanley Osher. A review of level-set methods and some recent applications. Journal of Computational Physics, 353(??):82–109, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307441.

Gorgizadeh:2018:ECC

[GFvR18] Shahnam Gorgizadeh, Thomas Flisgen, and Ursula van Rienen. Eigenmode computation of cavities with perturbed geometry using matrix perturbation methods applied on generalized eigenvalue problems. Journal of Computational Physics, 364(??):347–364, July 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301645.

Guo:2016:HRS

[GFW16] Xiaocheng Guo, Vladimir Florinski, and Chi Wang. The HLLD Riemann solver based on magnetic field decomposition method for the numerical simulation of magneto-hydrodynamics. Journal of Computational Physics, 327(??):543–552, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304788.

Gimenez:2015:EVL

[GG15] Juan M. Gimenez and Leo M. González. An extended validation of the last generation of particle finite element method for free surface flows. Journal of Computational Physics, 284(??):186–205, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008420.

Gjennestad:2017:CTD

[GGL^{+}17] Magnus Aa. Gjennestad, Andrea Gruber, Karl Yngve Lervåg, Øyvind Johansen, Åsmund Ervik, Morten Hammer, and Svend Tollak Munkejord. Computation of three-dimensional three-phase flow of carbon dioxide using a high-order WENO scheme. Journal of Computational Physics, 348(??):1–22, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305181.

Gordon:2015:CHO

[GGT15] Dan Gordon, Rachel Gordon, and Eli Turkel. Compact high order schemes with gradient-direction derivatives for absorbing boundary conditions. Journal of Computational Physics, 297(??):295–315, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003617.

Guillen-Gonzalez:2018:UES

[GGT18] F. Guillén-González and G. Tierra. Unconditionally energy stable numerical schemes for phase-field vesicle membrane model. Journal of Computational Physics, 354(??):67–85, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308276.

Gunther:2017:FSA

[GGW17] Stefanie Günther, Nicolas R. Gauger, and Qiqi Wang. A framework for simultaneous aerodynamic design optimization in the presence of chaos. Journal of Computational Physics, 328(??):387–398, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305411.

Gao:2017:IRS

[GH17a] Kai Gao and Lianjie Huang. An improved rotated staggeredgrid finite-difference method with fourth-order temporal accuracy for elastic-wave modeling in anisotropic media. Journal of Computational Physics, 350(??):361–386, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306344.

Guerrier:2017:MMS

[GH17b] C. Guerrier and D. Holcman. Multiscale models and stochastic simulation methods for computing rare but key binding events in cell biology. Journal of Computational Physics, 340(??):617– 638, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302590.

Gilani:2019:ASL

[GH19] Faheem Gilani and John Harlim. Approximating solutions of linear elliptic PDE’s on a smooth manifold using local kernel. Journal of Computational Physics, 395(??):563–582, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304425.

Gibou:2019:SIA

[GHF19] Frederic Gibou, David Hyde, and Ron Fedkiw. Sharp interface approaches and deep learning techniques for multiphase flows. Journal of Computational Physics, 380(??):442–463, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303371.

Garrett:2015:OLS

[GHH15] C. Kristopher Garrett, Cory Hauck, and Judith Hill. Optimization and large scale computation of an entropy-based moment closure. Journal of Computational Physics, 302(??):573–590, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005896.

Guo:2016:TDS

[GHH^{+}16] Yukun Guo, Dietmar Hömberg, Guanghui Hu, Jingzhi Li, and Hongyu Liu. A time domain sampling method for inverse acoustic scattering problems. Journal of Computational Physics, 314(??):647–660, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001984.

Greengard:2015:ELM

[GHJ15] Leslie Greengard, Thomas Hagstrom, and Shidong Jiang. Extension of the Lorenz–Mie–Debye method for electromagnetic scattering to the time-domain. Journal of Computational Physics, 299(??):98–105, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004544.

Gordon:2015:AFP

[GHL15] D. F. Gordon, B. Hafizi, and A. S. Landsman. Amplitude flux, probability flux, and gauge invariance in the finite volume scheme for the Schrödinger equation. Journal of Computational Physics, 280(??):457–464, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006901.

Guan:2016:ESH

[GHL^{+}16] Zhen Guan, Vili Heinonen, John Lowengrub, Cheng Wang, and Steven M. Wise. An energy stable, hexagonal finite difference scheme for the 2D phase field crystal amplitude equations. Journal of Computational Physics, 321(??):1026–1054, September 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302327.

Gamba:2015:SMK

[GHM15] Irene M. Gamba, Jeffrey R. Haack, and Sebastien Motsch. Spectral method for a kinetic swarming model. Journal of Computational Physics, 297(??):32–46, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002892.

Ghosh:2017:CIE

[Gho17] Karabi Ghosh. Comments on “Including the effects of temperature-dependent opacities in the implicit Monte Carlo algorithm” by N. A. Gentile [J. Comput. Phys. 230 (2011) 5100–5114]. Journal of Computational Physics, 330(??):111– 113, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116305915.

See [Gen11].

Ghanbarzadeh:2015:LSM

[GHP15] Soheil Ghanbarzadeh, Marc A. Hesse, and Masa Prodanović. A level set method for materials with texturally equilibrated pores. Journal of Computational Physics, 297(??):480–494, September 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003575.

Garrick:2017:ICS

[GHR17] Daniel P. Garrick, Wyatt A. Hagen, and Jonathan D. Regele. An interface capturing scheme for modeling atomization in compressible flows. Journal of Computational Physics, 344(??):260–280, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303674.

Gerster:2019:HSG

[GHS19] Stephan Gerster, Michael Herty, and Aleksey Sikstel. Hyperbolic stochastic Galerkin formulation for the p-system. Journal of Computational Physics, 395(??):186–204, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304012.

Ganesh:2019:EAC

[GHV19] M. Ganesh, S. C. Hawkins, and D. Volkov. An efficient algorithm for a class of stochastic forward and inverse Maxwell models in R^{3}. Journal of Computational Physics, 398(??):Article 108881, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305789.

Gibson:2018:AIL

[Gib18] Peter C. Gibson. Acoustic imaging of layered media. Journal of Computational Physics, 372(??):524–545, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304315.

Gavrilyuk:2018:MDS

[GIF18] S. Gavrilyuk, K. Ivanova, and N. Favrie. Multi-dimensional shear shallow water flows: Problems and solutions. Journal of Computational Physics, 365(??):252–280, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302250.

Gorji:2015:FPD

[GJ15] M. Hossein Gorji and Patrick Jenny. Fokker–Planck DSMC algorithm for simulations of rarefied gas flows. Journal of Computational Physics, 287(??):110–129, 2015. CODEN JCT-

PAH. ISSN 0021-9991. URL http://www.sciencedirect.
com/science/article/pii/S0021999115000558.

Gorodetsky:2018:GBO

[GJ18] Alex A. Gorodetsky and John D. Jakeman. Gradient-based optimization for regression in the functional tensor-train format. Journal of Computational Physics, 374(??):1219–1238, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305321.

Gamba:2019:MMD

[GJL19] Irene M. Gamba, Shi Jin, and Liu Liu. Micro-macro decomposition based asymptotic-preserving numerical schemes and numerical moments conservation for collisional nonlinear kinetic equations. Journal of Computational Physics, 382(??):264–290, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300439.

Giuliani:2018:ASL

[GK18] Andrew Giuliani and Lilia Krivodonova. Analysis of slope limiters on unstructured triangular meshes. Journal of Computational Physics, 374(??):1–26, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304923.

Gao:2019:CFE

[GK19a] Longfei Gao and David Keyes. Combining finite element and finite difference methods for isotropic elastic wave simulations in an energy-conserving manner. Journal of Computational Physics, 378(??):665–685, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307757.

Giuliani:2019:AMR

[GK19b] Andrew Giuliani and Lilia Krivodonova. Adaptive mesh refinement on graphics processing units for applications in gas dynamics. Journal of Computational Physics, 381(??):67–90, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118308155.

Grigo:2019:PAP

[GK19c] Constantin Grigo and Phaedon-Stelios Koutsourelakis. A physics-aware, probabilistic machine learning framework for coarse-graining high-dimensional systems in the small data regime. Journal of Computational Physics, 397(??):Article 108842, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305261.

Gruber:2015:FFT

[GKE15] M. E. Gruber, C. Koenen, and T. F. Eibert. A Fast Fourier Transform accelerated Ewald summation technique for the vector electromagnetic rectangular cavity Green’s function. Journal of Computational Physics, 280(??):570–578, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006949.

Gelss:2017:NNI

[GKMS17] Patrick Gelß, Stefan Klus, Sebastian Matera, and Christof Schütte. Nearest-neighbor interaction systems in the tensortrain format. Journal of Computational Physics, 341(??):140– 162, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302784.

Grote:2017:TDW

[GKNA17] Marcus J. Grote, Marie Kray, Frédéric Nataf, and Franck Assous. Time-dependent wave splitting and source separation. Journal of Computational Physics, 330(??):981–996, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305198.

Gourgoulias:2017:ICQ

[GKRB17] Konstantinos Gourgoulias, Markos A. Katsoulakis, and Luc Rey-Bellet. Information criteria for quantifying loss of reversibility in parallelized KMC. Journal of Computational Physics, 328(??):438–454, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305290.

Gyrya:2017:AOM

[GL17] V. Gyrya and K. Lipnikov. The arbitrary order mimetic finite difference method for a diffusion equation with a non-symmetric diffusion tensor. Journal of Computational Physics, 348(??):549–566, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305272.

Gregory:2019:SDI

[GLG^{+}19] Alastair Gregory, F. Din-Houn Lau, Mark Girolami, Liam J. Butler, and Mohammed Z. E. B. Elshafie. The synthesis of data from instrumented structures and physicsbased models via Gaussian processes. Journal of Computational Physics, 392(??):248–265, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303183.

Grosheintz-Laval:2019:HOW

[GLK19] L. Grosheintz-Laval and R. Käppeli. High-order well-balanced finite volume schemes for the Euler equations with gravitation. Journal of Computational Physics, 378(??):324–343, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307368.

Goza:2016:ACS

[GLMC16] Andres Goza, Sebastian Liska, Benjamin Morley, and Tim Colonius. Accurate computation of surface stresses and forces with immersed boundary methods. Journal of Computational Physics, 321(??):860–873, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630239X.

Gangbo:2019:UOT

[GLOP19] Wilfrid Gangbo, Wuchen Li, Stanley Osher, and Michael Puthawala. Unnormalized optimal transport. Journal of Computational Physics, 399(??):Article 108940, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S002199911930645X.

Gilmanov:2015:NAS

[GLS15] Anvar Gilmanov, Trung Bao Le, and Fotis Sotiropoulos. A numerical approach for simulating fluid structure interaction of flexible thin shells undergoing arbitrarily large deformations in complex domains. Journal of Computational Physics, 300(??):814–843, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005252.

Ge:2018:EMP

[GLTB18] Zhouyang Ge, Jean-Christophe Loiseau, Outi Tammisola, and Luca Brandt. An efficient mass-preserving interfacecorrection level set/ghost fluid method for droplet suspensions under depletion forces. Journal of Computational Physics, 353(??):435–459, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308136.

Guittet:2015:SEP

[GLTG15] Arthur Guittet, Mathieu Lepilliez, Sebastien Tanguy, and Frédéric Gibou. Solving elliptic problems with discontinuities on irregular domains — the Voronoi Interface Method. Journal of Computational Physics, 298(??):747–765, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004234.

Guo:2018:FOS

[GLW18] Xu Guo, Yutian Li, and Hong Wang. A fourth-order scheme for space fractional diffusion equations. Journal of Computational Physics, 373(??):410–424, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830192X.

Gao:2016:IMA

[GLZ16] Weiguo Gao, Jing Leng, and Xiang Zhou. Iterative minimization algorithm for efficient calculations of transition states. Journal of Computational Physics, 309(??):69–87, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008815.

Guo:2019:DDP

[GLZ19] Ling Guo, Yongle Liu, and Tao Zhou. Data-driven polynomial chaos expansions: a weighted least-square approximation. Journal of Computational Physics, 381(??):129–145, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300014.

Gejadze:2016:DCS

[GM16] I. Yu. Gejadze and P.-O. Malaterre. Design of the control set in the framework of variational data assimilation. Journal of Computational Physics, 325(??):358–379, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303862.

Galagusz:2019:IDD

[GM19] Ryan Galagusz and Steve McFee. An iterative domain decomposition, spectral finite element method on non-conforming meshes suitable for high frequency Helmholtz problems. Journal of Computational Physics, 379(??):132–172, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307344.

Guirguis:2018:HRN

[GMA18] David Guirguis, William W. Melek, and Mohamed F. Aly. High-resolution non-gradient topology optimization. Journal of Computational Physics, 372(??):107–125, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304030.

Garicano-Mena:2019:WPM

[GMD19] Jesús Garicano-Mena and Gérard Degrez. On the wellposedness of the multi-dimensional Roe–Liu–Vinokur linearization for residual distribution schemes. Journal of Computational Physics, 378(??):760–769, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307253.

Garicano-Mena:2018:EVB

[GMLD18] Jesús Garicano-Mena, Andrea Lani, and Gérard Degrez. An entropy-variables-based formulation of residual distribution schemes for non-equilibrium flows. Journal of Computational Physics, 362(??):163–189, June 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301025.

Garrappa:2015:STF

[GMP15] Roberto Garrappa, Igor Moret, and Marina Popolizio. Solving the time-fractional Schrödinger equation by Krylov projection methods. Journal of Computational Physics, 293(??):115–134, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006597.

Ganapol:2016:SMN

[GMP16] B. D. Ganapol, D. Mostacci, and A. Previti. A solution of the monoenergetic neutral particle transport equation for adjacent half-spaces with anisotropic scattering. Journal of Computational Physics, 316(??):814–845, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001212.

Gelss:2016:SME

[GMS16] Patrick Gelß, Sebastian Matera, and Christof Schütte. Solving the master equation without kinetic Monte Carlo: Tensor train approximations for a CO oxidation model. Journal of Computational Physics, 314(??):489–502, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001777.

Gejadze:2019:UDP

[GMS19] I. Gejadze, P.-O. Malaterre, and V. Shutyaev. On the use of derivatives in the polynomial chaos based global sensitivity and uncertainty analysis applied to the distributed parameter models. Journal of Computational Physics, 381(??):218–245, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911930004X.

Guo:2019:MFP

[GMT19] Bo Guo, Yashar Mehmani, and Hamdi A. Tchelepi. Multiscale formulation of pore-scale compressible Darcy–Stokes flow. Journal of Computational Physics, 397(??):Article 108849, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305339.

Gillis:2019:IIV

[GMWC19] T. Gillis, Y. Marichal, G. Winckelmans, and P. Chatelain. A 2D immersed interface Vortex Particle–Mesh method. Journal of Computational Physics, 394(??):700–718, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303717.

Golbabai:2016:CNS

[GN16] Ahmad Golbabai and Ahmad Nikpour. Computing a numerical solution of two dimensional non-linear Schrödinger equation on complexly shaped domains by RBF based differential quadrature method. Journal of Computational Physics, 322(??):586–602, October 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302820.

Ghasemi:2019:ESC

[GN19a] Fatemeh Ghasemi and Jan Nordström. An energy stable coupling procedure for the compressible and incompressible Navier–Stokes equations. Journal of Computational Physics, 396(??):280–302, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305078.

Guan:2019:TRT

[GN19b] Yifei Guan and Igor Novosselov. Two relaxation time lattice Boltzmann method coupled to Fast Fourier Transform Poisson solver: Application to electroconvective flow. Journal of Computational Physics, 397(??):Article 108830, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305145.

Gokhale:2018:DSCa

[GNK18a] Nandan Gokhale, Nikos Nikiforakis, and Rupert Klein. A dimensionally split Cartesian cut cell method for hyperbolic conservation laws. Journal of Computational Physics, 364(??):186–208, July 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301578.

Gokhale:2018:DSCb

[GNK18b] Nandan Gokhale, Nikos Nikiforakis, and Rupert Klein. A dimensionally split Cartesian cut cell method for the compressible Navier–Stokes equations. Journal of Computational Physics, 375(??):1205–1219, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306235.

Gnoffo:2017:SND

[Gno17] Peter A. Gnoffo. Solutions of nonlinear differential equations with feature detection using fast Walsh transforms. Journal of Computational Physics, 338(??):620–649, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302036.

Guo:2018:GEM

[GNZ18] Ling Guo, Akil Narayan, and Tao Zhou. A gradient enhanced l_{1}-minimization for sparse approximation of polynomial chaos expansions. Journal of Computational Physics, 367(??):49–64, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302420.

Gelebart:2015:FMP

[GO15] Lionel Gélébart and Franck Ouaki. Filtering material properties to improve FFT-based methods for numerical homogenization. Journal of Computational Physics, 294(??):90–95, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500203X.

Glasner:2016:IAC

[GO16] Karl Glasner and Saulo Orizaga. Improving the accuracy of convexity splitting methods for gradient flow equations. Journal of Computational Physics, 315(??):52–64, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001947.

Garrick:2017:FVH

[GOR17] Daniel P. Garrick, Mark Owkes, and Jonathan D. Regele. A finite-volume HLLC-based scheme for compressible interfacial flows with surface tension. Journal of Computational Physics, 339(??):46–67, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301948.

Gaudreault:2016:EET

[GP16a] Stéphane Gaudreault and Janusz A. Pudykiewicz. An efficient exponential time integration method for the numerical solution of the shallow water equations on the sphere. Journal of Computational Physics, 322(??):827–848, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302911.

Guermond:2016:FEA

[GP16b] Jean-Luc Guermond and Bojan Popov. Fast estimation from above of the maximum wave speed in the Riemann problem for the Euler equations. Journal of Computational Physics, 321(??):908–926, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301991.

Guyenne:2016:OEM

[GP16c] Philippe Guyenne and Emilian I. Parau. An operator expansion method for computing nonlinear surface waves on a ferrofluid jet. Journal of Computational Physics, 321(??):414–434, September 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302005.

Gallinato:2017:SSO

[GP17] Olivier Gallinato and Clair Poignard. Superconvergent second order Cartesian method for solving free boundary problem for invadopodia formation. Journal of Computational Physics, 339(??):412–431, June 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301973.

Gardas:2018:CDQ

[GP18] Bartlomiej Gardas and Andrzej Ptok. Counting defects in quantum computers with Graphics Processing Units. Journal of Computational Physics, 366(??):320–326, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302304.

Gargallo-Peiro:2018:MGS

[GPAO^{+}18] Abel Gargallo-Peiró, Matias Avila, Herbert Owen, Luis Prieto-Godino, and Arnau Folch. Mesh generation, sizing and convergence for onshore and offshore wind farm Atmospheric Boundary Layer flow simulation with actuator discs. Journal of Computational Physics, 375(??):209–227, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305540.

Guittet:2017:VIA

[GPG17] Arthur Guittet, Clair Poignard, and Frederic Gibou. A Voronoi interface approach to cell aggregate electropermeabilization. Journal of Computational Physics, 332(??):143–159, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306404.

G:2018:CRS

[GPRA18] John E. Ortiz G., Axelle Pillain, Lyes Rahmouni, and Francesco P. Andriulli. A Calderon regularized symmetric formulation for the electroencephalography forward problem. Journal of Computational Physics, 375(??):291–306, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305102.

Geier:2017:PCLa

[GPS17a] Martin Geier, Andrea Pasquali, and Martin Schönherr. Parametrization of the cumulant lattice Boltzmann method for fourth order accurate diffusion. Part I: Derivation and validation. Journal of Computational Physics, 348(??):862–888, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304230.

Geier:2017:PCLb

[GPS17b] Martin Geier, Andrea Pasquali, and Martin Schönherr. Parametrization of the cumulant lattice Boltzmann method for fourth order accurate diffusion. Part II: Application to flow around a sphere at drag crisis. Journal of Computational Physics, 348(??):889–898, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305065.

Guermond:2019:RER

[GPTK19] Jean-Luc Guermond, Bojan Popov, Eric Tovar, and Chris Kees. Robust explicit relaxation technique for solving the Green–Naghdi equations. Journal of Computational Physics, 399(??):Article 108917, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306229.

Gu:2019:GFD

[GQC^{+}19] Yan Gu, Wenzhen Qu, Wen Chen, Lina Song, and Chuanzeng Zhang. The generalized finite difference method for long-time dynamic modeling of three-dimensional coupled thermoelasticity problems. Journal of Computational Physics, 384(??):42– 59, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300750.

Ginzburg:2015:TET

[GR15] Irina Ginzburg and Laetitia Roux. Truncation effect on Taylor–Aris dispersion in lattice Boltzmann schemes: Accuracy towards stability. Journal of Computational Physics, 299(??):974–1003, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115004623.

Gamba:2018:GPA

[GR18] Irene M. Gamba and Sergej Rjasanow. Galerkin–Petrov approach for the Boltzmann equation. Journal of Computational Physics, 366(??):341–365, August 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830233X.

Guo:2019:IFT

[GR19] Zhenyu Guo and Albert C. Reynolds. INSIM–FT in three-dimensions with gravity. Journal of Computational Physics, 380(??):143–169, March 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830812X.

Grigoriu:2015:PMS

[Gri15] M. Grigoriu. Parametric models for samples of random functions. Journal of Computational Physics, 297(??):47– 71, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115003101.

Grigoriu:2019:FDM

[Gri19] M. Grigoriu. Finite dimensional models for random functions. Journal of Computational Physics, 376(??):1253–1272, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306314.

Gonzalez-Rodriguez:2015:LEI

[GRMK15] Pedro Gonzalez-Rodriguez, Miguel Moscoso, and Manuel Kindelan. Laurent expansion of the inverse of perturbed, singular matrices. Journal of Computational Physics, 299 (??):307–319, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004519.

Groot:2018:SOF

[Gro18] Robert D. Groot. Second order front tracking algorithm for Stefan problem on a regular grid. Journal of Computational Physics, 372(??):956–971, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302845.

Goshayeshi:2015:DSH

[GRS15] Bijan Goshayeshi, Ehsan Roohi, and Stefan Stefanov. DSMC simulation of hypersonic flows using an improved SBT–TAS technique. Journal of Computational Physics, 303(??):28– 44, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115006208.

Gaudreault:2018:KFA

[GRT18] Stéphane Gaudreault, Greg Rainwater, and Mayya Tokman. KIOPS: a fast adaptive Krylov subspace solver for exponential integrators. Journal of Computational Physics, 372(??):236–255, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999118304042. See corrigendum [GRT21].

Gaudreault:2021:CKF

[GRT21] Stéphane Gaudreault, Greg Rainwater, and Mayya Tokman. Corrigendum to “KIOPS: a fast adaptive Krylov subspace solver for exponential integrators” [J. Comput. Phys. 372 (2018) 236–255]. Journal of Computational Physics, 441(??):Article 110443, September 15, 2021. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999121003387. See [GRT18].

Gao:2015:TPC

[GS15a] Guang-Hua Gao and Hai-Wei Sun. Three-point combined compact difference schemes for time-fractional advectiondiffusion equations with smooth solutions. Journal of Computational Physics, 298(??):520–538, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003940.

Gao:2015:EFL

[GS15b] Huadong Gao and Weiwei Sun. An efficient fully linearized semi-implicit Galerkin-mixed FEM for the dynamical Ginzburg–Landau equations of superconductivity. Journal of Computational Physics, 294(??):329–345, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002181.

Gejadze:2015:GVO

[GS15c] I. Yu. Gejadze and V. Shutyaev. On gauss-verifiability of optimal solutions in variational data assimilation problems with nonlinear dynamics. Journal of Computational Physics, 280(??):439–456, January 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114006688.

Ghosh:2016:HOF

[GS16] Swarnava Ghosh and Phanish Suryanarayana. Higherorder finite-difference formulation of periodic Orbital-free Density Functional Theory. Journal of Computational Physics, 307(??):634–652, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115008463.

Giovanis:2018:UQC

[GS18] D. G. Giovanis and M. D. Shields. Uncertainty quantification for complex systems with very high dimensional response using Grassmann manifold variations. Journal of Computational Physics, 364(??):393–415, July 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830161X.

Gallardo:2019:CTD

[GSC19] José M. Gallardo, Kleiton A. Schneider, and Manuel J. Castro. On a class of two-dimensional incomplete Riemann solvers. Journal of Computational Physics, 386(??):541–567, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301512.

Gnedin:2018:ECF

[GSK18] Nickolay Y. Gnedin, Vadim A. Semenov, and Andrey V. Kravtsov. Enforcing the Courant–Friedrichs–Lewy condition in explicitly conservative local time stepping schemes. Journal of Computational Physics, 359(??):93–105, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300184.

Ghaisas:2018:UHO

[GSL18] Niranjan S. Ghaisas, Akshay Subramaniam, and Sanjiva K. Lele. A unified high-order Eulerian method for continuum simulations of fluid flow and of elastic-plastic deformations in solids. Journal of Computational Physics, 371(??):452–482, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303449.

Guo:2019:IHF

[GSL^{+}19] Tongqing Guo, Ennan Shen, Zhiliang Lu, Yan Wang, and Lu Dong. Implicit heat flux correction-based immersed boundary-finite volume method for thermal flows with Neumann boundary conditions. Journal of Computational Physics, 386(??):64–83, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301330.

Gutierrez-Santacreu:2018:FED

[GSMR18] Juan Vicente Gutiérrez-Santacreu, Omar Maj, and Marco Restelli. Finite element discretization of a Stokes-like model arising in plasma physics. Journal of Computational Physics, 373(??):811–834, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304248.

Galagusz:2016:FPM

[GSN16] Ryan Galagusz, David Shirokoff, and Jean-Christophe Nave. A Fourier penalty method for solving the time-dependent Maxwell’s equations in domains with curved boundaries. Journal of Computational Physics, 306(??):167–198, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007718.

Greene:2017:DMA

[GSN17] Patrick T. Greene, Samuel P. Schofield, and Robert Nourgaliev. Dynamic mesh adaptation for front evolution using discontinuous Galerkin based weighted condition number relaxation. Journal of Computational Physics, 335(??):664–687, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300657.

Gao:2015:SCF

[GSS15a] Guang-Hua Gao, Hai-Wei Sun, and Zhi-Zhong Sun. Stability and convergence of finite difference schemes for a class of timefractional sub-diffusion equations based on certain superconvergence. Journal of Computational Physics, 280(??):510–528, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400669X.

Gupta:2015:HLB

[GSS15b] A. Gupta, M. Sbragaglia, and A. Scagliarini. Hybrid Lattice Boltzmann/Finite Difference simulations of viscoelastic multicomponent flows in confined geometries. Journal of Computational Physics, 291(??):177–197, June 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001394.

Ghanem:2019:DOS

[GSS^{+}19] R. G. Ghanem, C. Soize, C. Safta, X. Huan, G. Lacaze, J. C. Oefelein, and H. N. Najm. Design optimization of a scramjet under uncertainty using probabilistic learning on manifolds. Journal of Computational Physics, 399(??):Article 108930, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119306357.

Gordin:2018:CDS

[GT18] Vladimir A. Gordin and Evgenii A. Tsymbalov. Compact difference scheme for parabolic and Schrödinger-type equations with variable coefficients. Journal of Computational Physics, 375(??):1451–1468, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305965.

Grebenkov:2019:SAC

[GT19] Denis S. Grebenkov and Sergey D. Traytak. Semi-analytical computation of Laplacian Green functions in three-dimensional domains with disconnected spherical boundaries. Journal of Computational Physics, 379(??):91–117, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307009.

Guittet:2015:SPM

[GTG15] Arthur Guittet, Maxime Theillard, and Frédéric Gibou. A stable projection method for the incompressible Navier–Stokes equations on arbitrary geometries and adaptive quad/octrees. Journal of Computational Physics, 292(??):215–238, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001710.

Gallezot:2018:MAB

[GTL18] M. Gallezot, F. Treyssède, and L. Laguerre. A modal approach based on perfectly matched layers for the forced response of elastic open waveguides. Journal of Computational Physics, 356(??):391–409, March 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309026.

Guo:2015:EHR

[Guo15] Xiaocheng Guo. An extended HLLC Riemann solver for the magneto-hydrodynamics including strong internal magnetic field. Journal of Computational Physics, 290(??):352– 363, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001187.

Ganapathy:2018:MRP

[GV18] Chandrashekar Ganapathy and Denis Voskov. Multiscale reconstruction in physics for compositional simulation. Journal of Computational Physics, 375(??):747–762, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305771.

Guo:2016:ESB

[GVTQ16] Jianwei Guo, Stéphanie Veran-Tissoires, and Michel Quintard. Effective surface and boundary conditions for heterogeneous surfaces with mixed boundary conditions. Journal of Computational Physics, 305(??):942–963, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007275.

Guo:2016:LBM

[GW16] Yangyu Guo and Moran Wang. Lattice Boltzmann modeling of phonon transport. Journal of Computational Physics, 315(??):1–15, June 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001935.

Gan:2015:CET

[GWB^{+}15] Zecheng Gan, Huanxin Wu, Kipton Barros, Zhenli Xu, and Erik Luijten. Comparison of efficient techniques for the simulation of dielectric objects in electrolytes. Journal of Computational Physics, 291(??):317–333, June 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001667.

Gillis:2017:EIP

[GWC17] T. Gillis, G. Winckelmans, and P. Chatelain. An efficient iterative penalization method using recycled Krylov subspaces and its application to impulsively started flows. Journal of Computational Physics, 347(??):490–505, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730517X.

Gillis:2018:FII

[GWC18] T. Gillis, G. Winckelmans, and P. Chatelain. Fast immersed interface Poisson solver for 3D unbounded problems around arbitrary geometries. Journal of Computational Physics, 354(??):403–416, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308094.

Guha:2015:VBA

[GWE^{+}15] Nilabja Guha, Xiaoqing Wu, Yalchin Efendiev, Bangti Jin, and Bani K. Mallick. A variational Bayesian approach for inverse problems with skew-t error distributions. Journal of Computational Physics, 301(??):377–393, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500515X.

Gassner:2016:SFN

[GWK16] Gregor J. Gassner, Andrew R. Winters, and David A. Kopriva. Split form nodal discontinuous Galerkin schemes with summation-by-parts property for the compressible Euler equations. Journal of Computational Physics, 327(??):39–66, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304259.

Gong:2017:CFP

[GWWC17] Yuezheng Gong, Qi Wang, Yushun Wang, and Jiaxiang Cai. A conservative Fourier pseudo-spectral method for the nonlinear Schrödinger equation. Journal of Computational Physics, 328(??):354–370, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305204.

Gu:2018:IPL

[GWYS18] Z. H. Gu, H. L. Wen, C. H. Yu, and Tony W. H. Sheu. Interface-preserving level set method for simulating dam-break flows. Journal of Computational Physics, 374(??):249–280, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305205.

Guo:2015:EUE

[GX15] Ruihan Guo and Yan Xu. An efficient, unconditionally energy stable local discontinuous Galerkin scheme for the Cahn–Hilliard-Brinkman system. Journal of Computational Physics, 298(??):387–405, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003927.

Guo:2017:SIS

[GXX17] Ruihan Guo, Yinhua Xia, and Yan Xu. Semi-implicit spectral deferred correction methods for highly nonlinear partial differential equations. Journal of Computational Physics, 338(??):269–284, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301626.

Guo:2015:PPH

[GY15] Li Guo and Yang Yang. Positivity preserving high-order local discontinuous Galerkin method for parabolic equations with blow-up solutions. Journal of Computational Physics, 289(??):181–195, May 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001114.

Guo:2017:GRE

[GY17] Hailong Guo and Xu Yang. Gradient recovery for elliptic interface problem: II. Immersed finite element methods. Journal of Computational Physics, 338(??):606–619, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301900.

Guo:2018:GRE

[GY18] Hailong Guo and Xu Yang. Gradient recovery for elliptic interface problem: III. Nitsche’s method. Journal of Computational Physics, 356(??):46–63, March 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308690.

Guo:2019:BTB

[GYZ19] Hailong Guo, Xu Yang, and Yi Zhu. Bloch theorybased gradient recovery method for computing topological edge modes in photonic graphene. Journal of Computational Physics, 379(??):403–420, February 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307824.

Geng:2017:TCM

[GZ17] Weihua Geng and Shan Zhao. A two-component Matched Interface and Boundary (MIB) regularization for charge singularity in implicit solvation. Journal of Computational Physics, 351(??):25–39, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306861.

Gu:2018:CSM

[GZ18] Shuting Gu and Xiang Zhou. Convex splitting method for the calculation of transition states of energy functional. Journal of Computational Physics, 353(??):417–434, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307817.

Geneva:2019:QMF

[GZ19] Nicholas Geneva and Nicholas Zabaras. Quantifying model form uncertainty in Reynolds-averaged turbulence models with Bayesian deep neural networks. Journal of Computational Physics, 383(??):125–147, April 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300464.

Gomez:2019:SHD

[GZLH19] Pablo Gómez, Claudio Zanzi, Joaquín López, and Julio Hernández. Simulation of high density ratio interfacial flows on cell vertex/edge-based staggered octree grids with secondorder discretization at irregular nodes. Journal of Computational Physics, 376(??):478–507, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306454.

Gao:2017:REC

[GZM^{+}17] Chao Gao, Peng Zhang, Gil Marom, Yuefan Deng, and Danny Bluestein. Reducing the effects of compressibility in DPD-based blood flow simulations through severe stenotic microchannels. Journal of Computational Physics, 335(??): 812–827, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300785.

Gao:2016:TOS

[GZY16] Yingjie Gao, Jinhai Zhang, and Zhenxing Yao. Thirdorder symplectic integration method with inverse time dispersion transform for long-term simulation. Journal of Computational Physics, 314(??):436–449, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001832.

Gao:2019:ESL

[GZY19] Yingjie Gao, Jinhai Zhang, and Zhenxing Yao. Extending the stability limit of explicit scheme with spatial filtering for solving wave equations. Journal of Computational Physics, 397(??):Article 108853, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305376.

Haghani:2025:CCP

[Hag25] Reza Haghani. Corrigendum to “Conservative phase-field lattice-Boltzmann model for ternary fluids” [J. Comput. Phys. 374 (2018) 668–691]. Journal of Computational Physics, 523(??):??, February 15, 2025. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999124009549.

Huang:2016:SLF

[HAH16] Chieh-Sen Huang, Todd Arbogast, and Chen-Hui Hung. A semi-Lagrangian finite difference WENO scheme for scalar nonlinear conservation laws. Journal of Computational Physics, 322(??):559–585, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302522.

Hanasoge:2016:SSC

[Han16] Shravan M. Hanasoge. Spatio-spectral concentration of convolutions. Journal of Computational Physics, 313(??):674–686, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001406.

Han:2019:EMC

[Han19] Minsub Han. Exchange of macromolecules and colloids in a dense medium: a molecular simulation method. Journal of Computational Physics, 395(??):263–274, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304334.

Hadjidoukas:2015:HPC

[HAPK15] P. E. Hadjidoukas, P. Angelikopoulos, C. Papadimitriou, and P. Koumoutsakos. Π4U: a high performance computing framework for Bayesian uncertainty quantification of complex models. Journal of Computational Physics, 284(??):1–21, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008134.

Harris:2018:UCK

[Har18] Robert E. Harris. Unstructured continuum-kinetic solver with fluid-particle interaction for in-space propulsion contaminant dispersal. Journal of Computational Physics, 373 (??):64–90, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118304327.

Hou:2019:VSA

[HAX19] Dianming Hou, Mejdi Azaiez, and Chuanju Xu. A variant of scalar auxiliary variable approaches for gradient flows. Journal of Computational Physics, 395(??):307–332, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S002199911930381X.

Huang:2015:CFS

[HB15a] Zhu Huang and John P. Boyd. Chebyshev–Fourier spectral methods in bipolar coordinates. Journal of Computational Physics, 295(??):46–64, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002235.

Huang:2015:MPG

[HB15b] Zhu Huang and John P. Boyd. Modal preconditioning of Galerkin spectral methods: Dual bookkeeping for the Delves–Freeman iteration. Journal of Computational Physics, 300(??):1–4, November 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005173.

Hashemi:2016:NAH

[HB16] M. S. Hashemi and D. Baleanu. Numerical approximation of higher-order time-fractional telegraph equation by using a combination of a geometric approach and method of line. Journal of Computational Physics, 316(??):10–20, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300419.

Hamilton:2016:ACA

[HBC^{+}16] Steven Hamilton, Mark Berrill, Kevin Clarno, Roger Pawlowski, Alex Toth, C. T. Kelley, Thomas Evans, and Bobby Philip. An assessment of coupling algorithms for nuclear reactor core physics simulations. Journal of Computational Physics, 311(??):241–257, April 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000668.

Heitman:2015:ENP

[HBR15] Zhu Heitman, James Bremer, and Vladimir Rokhlin. On the existence of nonoscillatory phase functions for second order ordinary differential equations in the high-frequency regime. Journal of Computational Physics, 290(??):1–27, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000984.

Hua:2017:ETS

[HC17] Yu-Chao Hua and Bing-Yang Cao. An efficient two-step Monte Carlo method for heat conduction in nanostructures. Journal of Computational Physics, 342(??):253–266, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303212.

Huang:2018:NAC

[HC18a] Cong Huang and Li Li Chen. A new adaptively centralupwind sixth-order WENO scheme. Journal of Computational Physics, 357(??):1–15, March 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117309269.

Huang:2018:SSI

[HC18b] Cong Huang and Li Li Chen. A simple smoothness indicator for the WENO scheme with adaptive order. Journal of Computational Physics, 352(??):498–515, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730743X.

Hazra:2019:GCP

[HCB19] Arijit Hazra, Praveen Chandrashekar, and Dinshaw S. Balsara. Globally constraint-preserving FR/DG scheme for Maxwell’s equations at all orders. Journal of Computational Physics, 394(??):298–328, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119304048.

Hsu:2019:CGB

[HCLT19] Shih-Hsuan Hsu, Jay Chu, Ming-Chih Lai, and Richard Tsai. A coupled grid based particle and implicit boundary integral method for two-phase flows with insoluble surfactant. Journal of Computational Physics, 395(??):747–764, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304541.

HosseiniMehr:2018:ADM

[HCVH18] Mousa HosseiniMehr, Matteo Cusini, Cornelis Vuik, and Hadi Hajibeygi. Algebraic dynamic multilevel method for embedded discrete fracture model (F-ADM). Journal of Computational Physics, 373(??):324–345, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830456X.

Hu:2015:IGT

[HCW15] Langhua Hu, Duan Chen, and Guo-Wei Wei. Impact of geometric, thermal and tunneling effects on nano-transistors. Journal of Computational Physics, 290(??):169–187, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500114X.

Hampton:2015:CSP

[HD15] Jerrad Hampton and Alireza Doostan. Compressive sampling of polynomial chaos expansions: Convergence analysis and sampling strategies. Journal of Computational Physics, 280(??):363–386, January 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400655X.

Hampton:2018:BAS

[HD18] Jerrad Hampton and Alireza Doostan. Basis adaptive sample efficient polynomial chaos (BASE-PC). Journal of Computational Physics, 371(??):20–49, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301955.

Hill:2018:BPI

[HDA^{+}18] S. Hill, D. Deising, T. Acher, H. Klein, D. Bothe, and H. Marschall. Boundedness-preserving implicit correction of mesh-induced errors for VOF based heat and mass transfer. Journal of Computational Physics, 352(??):285–300, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306873.

Haverkort:2016:IFV

[HdBH^{+}16] J. W. Haverkort, H. J. de Blank, G. T. A. Huysmans, J. Pratt, and B. Koren. Implementation of the full viscoresistive magnetohydrodynamic equations in a nonlinear finite element code. Journal of Computational Physics, 316(??):281–302, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300390.

Huang:2018:FPO

[HDF18] Daniel Z. Huang, Dante De Santis, and Charbel Farhat. A family of position- and orientation-independent embedded boundary methods for viscous flow and fluid-structure interaction problems. Journal of Computational Physics, 365(??):74–104, July 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301888.

Hamilton:2015:ESS

[HE15] Steven P. Hamilton and Thomas M. Evans. Efficient solution of the simplified P N equations. Journal of Computational Physics, 284(??):155–170, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008225.

Hamilton:2016:HZP

[HED^{+}16] Steven P. Hamilton, Thomas M. Evans, Gregory G. Davidson, Seth R. Johnson, Tara M. Pandya, and Andrew T. Godfrey. Hot zero power reactor calculations using the Insilico code. Journal of Computational Physics, 314(??):700–711, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001856.

Hay:2015:HAT

[HEPG15] A. Hay, S. Etienne, D. Pelletier, and A. Garon. hp-adaptive time integration based on the BDF for viscous flows. Journal of Computational Physics, 291(??):151–176, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001692.

Hermeline:2016:DMR

[Her16] F. Hermeline. A discretization of the multigroup P_{N} radiative transfer equation on general meshes. Journal of Computational Physics, 313(??):549–582, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001303.

Heuze:2017:LWT

[Heu17] Thomas Heuzé. Lax–Wendroff and TVD finite volume methods for unidimensional thermomechanical numerical simulations of impacts on elastic-plastic solids. Journal of Computational Physics, 346(??):369–388, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304722.

Heuze:2019:LWS

[Heu19] Thomas Heuzé. Lax–Wendroff schemes for elastic-plastic solids. Journal of Computational Physics, 396(??):89–105, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304589.

Hosseinverdi:2018:EHO

[HF18] Shirzad Hosseinverdi and Hermann F. Fasel. An efficient, high-order method for solving Poisson equation for immersed boundaries: Combination of compact difference and multiscale multigrid methods. Journal of Computational Physics, 374(??):912–940, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830528X.

Hyde:2019:UAM

[HF19] David A. B. Hyde and Ronald Fedkiw. A unified approach to monolithic solid-fluid coupling of sub-grid and more resolved solids. Journal of Computational Physics, 390(??):490–526, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302323.

Hank:2017:MHN

[HFM17] Sarah Hank, Nicolas Favrie, and Jacques Massoni. Modeling hyperelasticity in non-equilibrium multiphase flows. Journal of Computational Physics, 330(??):65–91, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305800.

Hampton:2018:PEB

[HFND18] Jerrad Hampton, Hillary R. Fairbanks, Akil Narayan, and Alireza Doostan. Practical error bounds for a non-intrusive bi-fidelity approach to parametric/stochastic model reduction. Journal of Computational Physics, 368(??):315–332, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302298.

He:2017:SIM

[HG17] Ping He and Ahmed F. Ghoniem. A sharp interface method for coupling multiphase flow, heat transfer and multicomponent mass transfer with interphase diffusion. Journal of Computational Physics, 332(??):316–332, March 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630643X.

Homsi:2017:CCE

[HGN17a] L. Homsi, C. Geuzaine, and L. Noels. Corrigendum to “A coupled electro-thermal Discontinuous Galerkin method” [J. Comput. Phys. 348 (2017) 231–258]. Journal of Computational Physics, 351(??):536–538, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306812. See [HGN17b].

Homsi:2017:CET

[HGN17b] L. Homsi, C. Geuzaine, and L. Noels. A coupled electrothermal Discontinuous Galerkin method. Journal of Computational Physics, 348(??):231–258, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999117305363. See corrigendum [HGN17a].

Harmon:2016:NAB

[HGR16] Michael Harmon, Irene M. Gamba, and Kui Ren. Numerical algorithms based on Galerkin methods for the modeling of reactive interfaces in photoelectrochemical (PEC) solar cells. Journal of Computational Physics, 327(??):140–167, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303825.

He:2018:LSF

[HGW18] Qiaolin He, Roland Glowinski, and Xiao-Ping Wang. A leastsquares/fictitious domain method for incompressible viscous flow around obstacles with Navier slip boundary condition. Journal of Computational Physics, 365(??):281–297, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302274.

Gao:2015:SHO

[hGwSzS15] Guang hua Gao, Hai wei Sun, and Zhi zhong Sun. Some highorder difference schemes for the distributed-order differential equations. Journal of Computational Physics, 298(??):337–359, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003873.

Hu:2019:ESS

[HH19] Zhicheng Hu and Guanghui Hu. An efficient steady-state solver for microflows with high-order moment model. Journal of Computational Physics, 392(??):462–482, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303031.

Han:2015:SSM

[HHA15] L. H. Han, X. Y. Hu, and N. A. Adams. Scale separation for multi-scale modeling of free-surface and two-phase flows with the conservative sharp interface method. Journal of Computational Physics, 280(??):387–403, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006755.

Hirabayashi:2016:NFM

[HHA16] Kota Hirabayashi, Masahiro Hoshino, and Takanobu Amano. A new framework for magnetohydrodynamic simulations with anisotropic pressure. Journal of Computational Physics, 327(??):851–872, December 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304855.

Heydari:2015:ECM

[HHCG15] M. H. Heydari, M. R. Hooshmandasl, C. Cattani, and F. M. Maalek Ghaini. An efficient computational method for solving nonlinear stochastic Itô integral equations: Application for stochastic problems in physics. Journal of Computational Physics, 283(??):148–168, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008043.

Hwang:2015:CWS

[HHK15] Chi-Ok Hwang, Sungpyo Hong, and Jinwoo Kim. Off-centered “Walk-on-Spheres” (WOS) algorithm. Journal of Computational Physics, 303(??):331–335, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006646.

Hsu:2019:CII

[HHL19] Shih-Hsuan Hsu, Wei-Fan Hu, and Ming-Chih Lai. A coupled immersed interface and grid based particle method for three-dimensional electrohydrodynamic simulations. Journal of Computational Physics, 398(??):Article 108903, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119306084.

Hou:2017:IAM

[HHLY17] Thomas Y. Hou, Feng-Nan Hwang, Pengfei Liu, and Chien-Chou Yao. An iteratively adaptive multi-scale finite element method for elliptic PDEs with rough coefficients. Journal of Computational Physics, 336(??):375–400, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300955.

Han:2017:ERR

[HHM17] Ee Han, Maren Hantke, and Siegfried Müller. Efficient and robust relaxation procedures for multi-component mixtures including phase transition. Journal of Computational Physics, 338(??):217–239, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301699.

Harlim:2015:AMC

[HHR15] John Harlim, Hoon Hong, and Jacob L. Robbins. An algebraic method for constructing stable and consistent autoregressive filters. Journal of Computational Physics, 283(??):241–257, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008110.

Hanus:2019:UFT

[HHR^{+}19] Milan Hanus, Logan H. Harbour, Jean C. Ragusa, Michael P. Adams, and Marvin L. Adams. Uncollided flux techniques for arbitrary finite element meshes. Journal of Computational Physics, 398(??):Article 108848, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305327.

Hosseininia:2019:CWM

[HHRA19] M. Hosseininia, M. H. Heydari, R. Roohi, and Z. Avazzadeh. A computational wavelet method for variable-order fractional model of dual phase lag bioheat equation. Journal of Computational Physics, 395(??):1–18, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304322. See comment [Pan20].

Huang:2015:CDO

[HHY15] Juan-Chen Huang, Tse-Yang Hsieh, and Jaw-Yen Yang. A conservative discrete ordinate method for solving semiclassical Boltzmann-BGK equation with Maxwell type wall boundary condition. Journal of Computational Physics, 290(??):112–131, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001072.

Huang:2016:SOC

[HHY16] Juntao Huang, Zexi Hu, and Wen-An Yong. Secondorder curved boundary treatments of the lattice Boltzmann method for convection-diffusion equations. Journal of Computational Physics, 310(??):26–44, April 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000097.

Han:2019:TPB

[HHZZ19] Yucen Han, Yucheng Hu, Pingwen Zhang, and Lei Zhang. Transition pathways between defect patterns in confined nematic liquid crystals. Journal of Computational Physics, 396 (??):1–11, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S002199911930436X.

Higdon:2015:MTS

[Hig15] Robert L. Higdon. Multiple time scales and pressure forcing in discontinuous Galerkin approximations to layered ocean models. Journal of Computational Physics, 295(??):230–260, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002569.

Higginson:2017:FAM

[Hig17] Drew P. Higginson. A full-angle Monte-Carlo scattering technique including cumulative and single-event Rutherford scattering in plasmas. Journal of Computational Physics, 349(??):589–603, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305879.

Hatori:2016:LLA

[HIN^{+}16] Tomoharu Hatori, Atsushi M. Ito, Masanori Nunami, Hideyuki Usui, and Hideaki Miura. Level-by-level artificial viscosity and visualization for MHD simulation with adaptive mesh refinement. Journal of Computational Physics, 319(??):231–241, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301504.

Hivert:2018:FOA

[Hiv18] Hélène Hivert. A first-order asymptotic preserving scheme for front propagation in a one-dimensional kinetic reaction– transport equation. Journal of Computational Physics, 367 (??):253–278, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118302614.

Hu:2016:SGM

[HJ16] Jingwei Hu and Shi Jin. A stochastic Galerkin method for the Boltzmann equation with uncertainty. Journal of Computational Physics, 315(??):150–168, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001996.

Hu:2019:SGA

[HJS19] Jingwei Hu, Shi Jin, and Ruiwen Shu. On stochastic Galerkin approximation of the nonlinear Boltzmann equation with uncertainty in the fluid regime. Journal of Computational Physics, 397(??):Article 108838, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305224.

Huang:2019:FCT

[HJW19] Jianguo Huang, Lili Ju, and Bo Wu. A fast compact time integrator method for a family of general order semilinear evolution equations. Journal of Computational Physics, 393(??):313–336, September 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303389.

Huang:2019:UES

[HJY19] Qiong-Ao Huang, Wei Jiang, and Jerry Zhijian Yang. An unconditionally energy stable scheme for simulating wrinkling phenomena of elastic thin films on a compliant substrate. Journal of Computational Physics, 388(??):123–143, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301925.

Hong:2017:ECM

[HJZC17] Jialin Hong, Lihai Ji, Liying Zhang, and Jiaxiang Cai. An energy-conserving method for stochastic Maxwell equations with multiplicative noise. Journal of Computational Physics, 351(??):216–229, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306903.

Held:2015:LBM

[HK15a] M. Held and A. Kendl. Lattice Boltzmann model for collisionless electrostatic drift wave turbulence obeying Charney– Hasegawa–Mima dynamics. Journal of Computational Physics, 298(??):622–635, October 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115004155.

Huang:2015:GDS

[HK15b] Weizhang Huang and Lennard Kamenski. A geometric discretization and a simple implementation for variational mesh generation and adaptation. Journal of Computational Physics, 301(??):322–337, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005598.

Hayashi:2016:YYZ

[HK16a] Hiroshi Hayashi and Akira Kageyama. Yin–Yang–Zhong grid: an overset grid system for a sphere. Journal of Computational Physics, 305(??):895–905, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007561.

Helsing:2016:DNE

[HK16b] Johan Helsing and Anders Karlsson. Determination of normalized electric eigenfields in microwave cavities with sharp edges. Journal of Computational Physics, 304(??):465–486, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006634.

Havasi:2018:REL

[HK18a] Ágnes Havasi and Ehsan Kazemi. On Richardson extrapolation for low-dissipation low-dispersion diagonally implicit Runge– Kutta schemes. Journal of Computational Physics, 358(??):21– 35, April 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300032.

Helsing:2018:HTP

[HK18b] Johan Helsing and Anders Karlsson. On a Helmholtz transmission problem in planar domains with corners. Journal of Computational Physics, 371(??):315–332, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830353X.

Haga:2019:RAL

[HK19a] Takanori Haga and Soshi Kawai. On a robust and accurate localized artificial diffusivity scheme for the highorder flux-reconstruction method. Journal of Computational Physics, 376(??):534–563, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306545.

Helsing:2019:PDI

[HK19b] Johan Helsing and Anders Karlsson. Physical-density integral equation methods for scattering from multi-dielectric cylinders. Journal of Computational Physics, 387(??):14–29, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301676.

Hajduk:2019:NDV

[HKA19] Hennes Hajduk, Dmitri Kuzmin, and Vadym Aizinger. New directional vector limiters for discontinuous Galerkin methods. Journal of Computational Physics, 384(??):308–325, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911930083X.

Holiday:2019:MLP

[HKBR^{+}19] Alexander Holiday, Mahdi Kooshkbaghi, Juan M. Bello-Rivas, C. William Gear, Antonios Zagaris, and Ioannis G. Kevrekidis. Manifold learning for parameter reduction. Journal of Computational Physics, 392(??):419–431, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302487.

Hirschler:2016:OBC

[HKH^{+}16] Manuel Hirschler, Philip Kunz, Manuel Huber, Friedemann Hahn, and Ulrich Nieken. Open boundary conditions for ISPH and their application to micro-flow. Journal of Computational Physics, 307(??):614–633, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008438.

Hedges:2017:SLS

[HKJ17] Lester O. Hedges, H. Alicia Kim, and Robert L. Jack. Stochastic level-set method for shape optimisation. Journal of Computational Physics, 348(??):82–107, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305120.

Harmandaris:2016:PSV

[HKKP16] Vagelis Harmandaris, Evangelia Kalligiannaki, Markos Katsoulakis, and Petr Plechác. Path-space variational inference for non-equilibrium coarse-grained systems. Journal of Computational Physics, 314(??):355–383, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911600173X.

Hejlesen:2015:IBP

[HKLW15] Mads Mølholm Hejlesen, Petros Koumoutsakos, Anthony Leonard, and Jens Honoré Walther. Iterative Brinkman penalization for remeshed vortex methods. Journal of Computational Physics, 280(??):547–562, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006652.

Hao:2018:CMM

[HKLZ18] Yongle Hao, Fengdai Kang, Jingzhi Li, and Kai Zhang. Computation of moments for Maxwell’s equations with random interfaces via pivoted low-rank approximation. Journal of Computational Physics, 371(??):1–19, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302961.

Huber:2016:PBM

[HKS^{+}16] M. Huber, F. Keller, W. Säckel, M. Hirschler, P. Kunz, S. M. Hassanizadeh, and U. Nieken. On the physically based modeling of surface tension and moving contact lines with dynamic contact angles on the continuum scale. Journal of Computational Physics, 310(??):459–477, April 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000383.

Hartmann:2015:GAC

[HL15a] Ralf Hartmann and Tobias Leicht. Generalized adjoint consistent treatment of wall boundary conditions for compressible flows. Journal of Computational Physics, 300(??):754–778, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004878.

Hou:2015:HSF

[HL15b] Thomas Y. Hou and Pengfei Liu. A heterogeneous stochastic FEM framework for elliptic PDEs. Journal of Computational Physics, 281(??):942–969, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007025.

Halashi:2016:RDG

[HL16a] Behrouz Karami Halashi and Hong Luo. A reconstructed discontinuous Galerkin method for magnetohydrodynamics on arbitrary grids. Journal of Computational Physics, 326 (??):258–277, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304132.

Hoffman:2016:TDN

[HL16b] Adam J. Hoffman and John C. Lee. A time-dependent neutron transport method of characteristics formulation with time derivative propagation. Journal of Computational Physics, 307(??):696–714, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007160.

Huang:2019:MUC

[HLA19] Ziyang Huang, Guang Lin, and Arezoo M. Ardekani. A mixed upwind/central WENO scheme for incompressible two-phase flows. Journal of Computational Physics, 387(??):455–480, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301603.

Huang:2019:EFB

[HLCL19] Tsung-Ming Huang, Tiexiang Li, Ruey-Lin Chern, and Wen-Wei Lin. Electromagnetic field behavior of 3D Maxwell’s equations for chiral media. Journal of Computational Physics, 379(??):118–131, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118307708.

Hoang:2019:CEL

[HLJ^{+}19] Thi-Thao-Phuong Hoang, Wei Leng, Lili Ju, Zhu Wang, and Konstantin Pieper. Conservative explicit local time-stepping schemes for the shallow water equations. Journal of Computational Physics, 382(??):152–176, April 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300233.

He:2016:MRF

[HLL^{+}16] Z. C. He, Eric Li, G. R. Liu, G. Y. Li, and A. G. Cheng. A mass-redistributed finite element method (MR–FEM) for acoustic problems using triangular mesh. Journal of Computational Physics, 323(??):149–170, October 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303205.

Huang:2018:IEC

[HLL^{+}18] He Huang, Li-Shi Luo, Rui Li, Jie Chen, and He Zhang. Improve the efficiency of the Cartesian tensor based fast multipole method for Coulomb interaction using the traces. Journal of Computational Physics, 371(??):122–136, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303280.

Horstmann:2017:HSC

[HLML17] Jan Tobias Horstmann, Thomas Le Garrec, Daniel-Ciprian Mincu, and Emmanuel Lévêque. Hybrid simulation combining two space-time discretization of the discrete-velocity Boltzmann equation. Journal of Computational Physics, 349(??):399–414, November 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306009.

Hu:2016:AMS

[HLQ16] Zhicheng Hu, Ruo Li, and Zhonghua Qiao. Acceleration for microflow simulations of high-order moment models by using lower-order model correction. Journal of Computational Physics, 327(??):225–244, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304636.

Hou:2015:SLE

[HLS15] Thomas Y. Hou, Qin Li, and Hayden Schaeffer. Sparse + low-energy decomposition for viscous conservation laws. Journal of Computational Physics, 288(??):150–166, May 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000893.

Higginson:2019:PNF

[HLS19] Drew Pitney Higginson, Anthony Link, and Andrea Schmidt. A pairwise nuclear fusion algorithm for weighted particlein-cell plasma simulations. Journal of Computational Physics, 388(??):439–453, July 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302037.

Hu:2016:VES

[HLSY16] Wei-Fan Hu, Ming-Chih Lai, Yunchang Seol, and Yuan-Nan Young. Vesicle electrohydrodynamic simulations by coupling immersed boundary and immersed interface method. Journal of Computational Physics, 317(??):66–81, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630081X.

Huang:2018:CFS

[HLTC18] Tsung-Ming Huang, Wen-Wei Lin, Heng Tian, and Guan-Hua Chen. Computing the full spectrum of large sparse palindromic quadratic eigenvalue problems arising from surface Green’s function calculations. Journal of Computational Physics, 356(??):340–355, March 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308951.

Hu:2015:FMP

[HLU15] Anjie Hu, Longjian Li, and Rizwan Uddin. Force method in a pseudo-potential lattice Boltzmann model. Journal of Computational Physics, 294(??):78–89, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001436.

Hu:2015:HIB

[HLY15] Wei-Fan Hu, Ming-Chih Lai, and Yuan-Nan Young. A hybrid immersed boundary and immersed interface method for electrohydrodynamic simulations. Journal of Computational Physics, 282(??):47–61, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007578.

Herrada:2016:NMS

[HM16a] M. A. Herrada and J. M. Montanero. A numerical method to study the dynamics of capillary fluid systems. Journal of Computational Physics, 306(??):137–147, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007883.

Horwitz:2016:ACS

[HM16b] J. A. K. Horwitz and A. Mani. Accurate calculation of Stokes drag for point-particle tracking in two-way coupled flows. Journal of Computational Physics, 318(??):85–109, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300808.

Hohenegger:2017:FPD

[HM17] Christel Hohenegger and Scott A. McKinley. Fluid-particle dynamics for passive tracers advected by a thermally fluctuating viscoelastic medium. Journal of Computational Physics, 340(??):688–711, July 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302541.

Horne:2019:MPUa

[HM19a] Wyatt James Horne and Krishnan Mahesh. A massivelyparallel, unstructured overset method for mesh connectivity. Journal of Computational Physics, 376(??):585–596, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306557.

Horne:2019:MPUb

[HM19b] Wyatt James Horne and Krishnan Mahesh. A massivelyparallel, unstructured overset method to simulate moving bodies in turbulent flows. Journal of Computational Physics, 397(??):Article 108790, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119304747.

Hu:2019:FSM

[HM19c] Jingwei Hu and Zheng Ma. A fast spectral method for the inelastic Boltzmann collision operator and application to heated granular gases. Journal of Computational Physics, 385(??):119–134, May 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930107X.

Hoang:2015:CES

[HMBH15] Tuan L. Hoang, Jaime Marian, Vasily V. Bulatov, and Peter Hosemann. Computationally-efficient stochastic cluster dynamics method for modeling damage accumulation in irradiated materials. Journal of Computational Physics, 300 (??):254–268, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115005148.

Honorio:2018:SEB

[HMFJ18] Hermínio T. Honório, Clovis R. Maliska, Massimiliano Ferronato, and Carlo Janna. A stabilized element-based finite volume method for poroelastic problems. Journal of Computational Physics, 364(??):49–72, July 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301621.

Huthmacher:2016:SSM

[HMRG16] Klaus Huthmacher, Andreas K. Molberg, Bärbel Rethfeld, and Jeremy R. Gulley. A split-step method to include electron-electron collisions via Monte Carlo in multiple rate equation simulations. Journal of Computational Physics, 322(??):535–546, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302686.

Hong:2017:HOP

[HN17a] Youngjoon Hong and David P. Nicholls. A high-order perturbation of surfaces method for scattering of linear waves by periodic multiply layered gratings in two and three dimensions. Journal of Computational Physics, 345(??):162–188, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303868.

Hong:2017:SHO

[HN17b] Youngjoon Hong and David P. Nicholls. A stable highorder perturbation of surfaces method for numerical simulation of diffraction problems in triply layered media. Journal of Computational Physics, 330(??):1043–1068, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305654.

Hong:2018:HOP

[HN18] Youngjoon Hong and David P. Nicholls. A high-order perturbation of surfaces method for vector electromagnetic scattering by doubly layered periodic crossed gratings. Journal of Computational Physics, 372(??):748–772, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304194.

Hosseini:2016:RDD

[HNS16] Bamdad Hosseini, Nilima Nigam, and John M. Stockie. On regularizations of the Dirac delta distribution. Journal of Computational Physics, 305(??):423–447, January 15, 2016.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007317.

Harker:2015:SEN

[HO15] Matthew Harker and Paul O’Leary. Sylvester equations and the numerical solution of partial fractional differential equations. Journal of Computational Physics, 293(??):370–384, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400850X.

Hejranfar:2017:PCB

[HP17] Kazem Hejranfar and Kaveh Parseh. Preconditioned characteristic boundary conditions based on artificial compressibility method for solution of incompressible flows. Journal of Computational Physics, 345(??):543–564, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303832.

Hammer:2019:MDM

[HPC19] Hans Hammer, HyeongKae Park, and Luis Chacón. A multi-dimensional, moment-accelerated deterministic particle method for time-dependent, multi-frequency thermal radiative transfer problems. Journal of Computational Physics, 386(??):653–674, June 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301524.

Helmich-Paris:2016:IMA

[HPV16] Benjamin Helmich-Paris and Lucas Visscher. Improvements on the minimax algorithm for the Laplace transformation of orbital energy denominators. Journal of Computational Physics, 321(??):927–931, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302364.

Ha:2018:GAS

[HPY18] Sanghyun Ha, Junshin Park, and Donghyun You. A GPUaccelerated semi-implicit fractional-step method for numerical solutions of incompressible Navier–Stokes equations. Journal of Computational Physics, 352(??):246–264, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307283.

Heumann:2017:FEM

[HR17] Holger Heumann and Francesca Rapetti. A finite element method with overlapping meshes for free-boundary axisymmetric plasma equilibria in realistic geometries. Journal of Computational Physics, 334(??):522–540, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300165.

Hackemack:2018:QSD

[HR18a] Michael W. Hackemack and Jean C. Ragusa. Quadratic serendipity discontinuous finite element discretization for S_{N} transport on arbitrary polygonal grids. Journal of Computational Physics, 374(??):188–212, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303401.

Hansel:2018:FCT

[HR18b] Joshua E. Hansel and Jean C. Ragusa. Flux-corrected transport techniques applied to the radiation transport equation discretized with continuous finite elements. Journal of Computational Physics, 354(??):179–195, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307829.

Hassanaly:2019:NCL

[HR19] Malik Hassanaly and Venkat Raman. Numerical convergence of the Lyapunov spectrum computed using low Mach number solvers. Journal of Computational Physics, 386(??):467–485, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300981.

Halpern:2016:GCT

[HRJ^{+}16] F. D. Halpern, P. Ricci, S. Jolliet, J. Loizu, J. Morales, A. Mosetto, F. Musil, F. Riva, T. M. Tran, and C. Wersal. The GBS code for tokamak scrape-off layer simulations. Journal of Computational Physics, 315(??):388–408, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001923.

Hardin:2017:FFE

[HS17a] Thomas J. Hardin and Christopher A. Schuh. Fast finite element calculation of effective conductivity of random continuum microstructures: the recursive Poincaré–Steklov operator method. Journal of Computational Physics, 342(??):1–12, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302929.

Hasbestan:2017:SNU

[HS17b] Jaber J. Hasbestan and Inanc Senocak. A short note on the use of the red–black tree in Cartesian adaptive mesh refinement algorithms. Journal of Computational Physics, 351(??):473–477, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117307295.

Hasbestan:2018:BOG

[HS18a] Jaber J. Hasbestan and Inanc Senocak. Binarized-octree generation for Cartesian adaptive mesh refinement around immersed geometries. Journal of Computational Physics, 368 (??):179–195, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911830264X.

Huang:2018:BPM

[HS18b] Juntao Huang and Chi-Wang Shu. Bound-preserving modified exponential Runge–Kutta discontinuous Galerkin methods for scalar hyperbolic equations with stiff source terms. Journal of Computational Physics, 361(??):111–135, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300731.

Hiller:2016:SRD

[HSB16] Thomas Hiller, Marta Sanchez de La Lama, and Martin Brinkmann. Stochastic Rotation Dynamics simulations of wetting multi-phase flows. Journal of Computational Physics, 315(??):554–576, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300328.

Hosseini:2016:LRP

[HSC16] Vahid Reza Hosseini, Elyas Shivanian, and Wen Chen. Local radial point interpolation (MLRPI) method for solving time fractional diffusion-wave equation with damping. Journal of Computational Physics, 312(??):307–332, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000942.

Huismann:2017:FFS

[HSF17] Immo Huismann, Jörg Stiller, and Jochen Fröhlich. Factorizing the factorization — a spectral-element solver for elliptic equations with linear operation count. Journal of Computational Physics, 346(??):437–448, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304576.

Huismann:2019:SSL

[HSF19] Immo Huismann, Jörg Stiller, and Jochen Fröhlich. Scaling to the stars — a linearly scaling elliptic solver for p-multigrid. Journal of Computational Physics, 398(??):Article 108868, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119305595.

He:2015:CMC

[HSK^{+}15] Wenkui He, Haibing Shao, Olaf Kolditz, Wenqing Wang, and Thomas Kalbacher. Comments on “A mass-conservative switching algorithm for modeling fluid flow in variably saturated porous media, K. Sadegh Zadeh, Journal of Computational Physics, 230 (2011)”. Journal of Computational Physics, 295(??):815–820, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500265X. See [Zad11].

He:2015:VPA

[HSLQ15] Yang He, Yajuan Sun, Jian Liu, and Hong Qin. Volumepreserving algorithms for charged particle dynamics. Journal of Computational Physics, 281(??):135–147, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007141.

He:2016:HOV

[HSLQ16] Yang He, Yajuan Sun, Jian Liu, and Hong Qin. Higher order volume-preserving schemes for charged particle dynamics. Journal of Computational Physics, 305(??):172–184, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007081.

Hamon:2019:MLS

[HSM19] François P. Hamon, Martin Schreiber, and Michael L. Minion. Multi-level spectral deferred corrections scheme for the shallow water equations on the rotating sphere. Journal of Computational Physics, 376(??):435–454, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306442.

Huang:2016:RIM

[HSSZ16] Ruihao Huang, Allan A. Struthers, Jiguang Sun, and Ruming Zhang. Recursive integral method for transmission eigenvalues. Journal of Computational Physics, 327(??):830–840, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304946.

Huber:2015:TSP

[HTBG15] Grégory Huber, Sébastien Tanguy, Jean-Christophe Béra, and Bruno Gilles. A time splitting projection scheme for compressible two-phase flows. Application to the interaction of bubbles with ultrasound waves. Journal of Computational Physics, 302(??):439–468, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006129.

Hassard:2018:EBE

[HTFL18] Patrick Hassard, Ian Turner, Troy Farrell, and Daniel Lester. An efficient boundary element formulation for doubly-periodic two-dimensional Stokes flow with pressure boundary conditions. Journal of Computational Physics, 365(??):18–36, July 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308860.

Hosseini:2017:IAN

[HTMP17] Babak S. Hosseini, Stefan Turek, Matthias Möller, and Christian Palmes. Isogeometric analysis of the Navier– Stokes–Cahn–Hilliard equations with application to incompressible two-phase flows. Journal of Computational Physics, 348(??):171–194, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305375.

Huang:2019:SMM

[HTvdH^{+}19] Zhenguang Huang, Gábor Tóth, Bart van der Holst, Yuxi Chen, and Tamas Gombosi. A six-moment multi-fluid plasma model. Journal of Computational Physics, 387(??):134–153, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301408.

He:2017:CBI

[HTZG17] Zhiwei He, Baolin Tian, Yousheng Zhang, and Fujie Gao. Characteristic-based and interface-sharpening algorithm for high-order simulations of immiscible compressible multimaterial flows. Journal of Computational Physics, 333(??): 247–268, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306970.

Hu:2017:NSD

[Hu17] Guanghui Hu. A numerical study of 2D detonation waves with adaptive finite volume methods on unstructured grids.

Journal of Computational Physics, 331(??):297–311, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306337.

Hesthaven:2018:NIR

[HU18] J. S. Hesthaven and S. Ubbiali. Non-intrusive reduced order modeling of nonlinear problems using neural networks. Journal of Computational Physics, 363(??):55–78, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301190.

Huet:2015:ODC

[Hue15] Maxime Huet. One-dimensional characteristic boundary conditions using nonlinear invariants. Journal of Computational Physics, 283(??):312–328, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008183.

Han:2015:SOT

[HW15a] Daozhi Han and Xiaoming Wang. A second order in time, uniquely solvable, unconditionally stable numerical scheme for Cahn–Hilliard–Navier–Stokes equation. Journal of Computational Physics, 290(??):139–156, June 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001163.

Hu:2015:APS

[HW15b]  Jingwei Hu and Li Wang.  An asymptotic-preserving

scheme for the semiconductor Boltzmann equation toward the energy-transport limit. Journal of Computational Physics, 281(??):806–824, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007384.

Huang:2015:PIE

[HW15c] Rongzong Huang and Huiying Wu. Phase interface effects in the total enthalpy-based lattice Boltzmann model for solid-liquid phase change. Journal of Computational Physics, 294(??):346–362, August 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115002259.

Hejlesen:2016:MMS

[HW16a] Mads Mølholm Hejlesen and Jens Honoré Walther. A multiresolution method for solving the Poisson equation using high order regularization. Journal of Computational Physics, 326(??):188–196, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304107.

Huang:2016:TOA

[HW16b] Rongzong Huang and Huiying Wu. Third-order analysis of pseudopotential lattice Boltzmann model for multiphase flow. Journal of Computational Physics, 327(??):121–139, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630451X.

Huang:2016:TEB

[HW16c] Rongzong Huang and Huiying Wu. Total enthalpy-based lattice Boltzmann method with adaptive mesh refinement for solid-liquid phase change. Journal of Computational Physics, 315(??):65–83, June 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001959.

Huang:2018:LBM

[HW18] Jizu Huang and Xiao-Ping Wang. A lattice Boltzmann model for multiphase flows with moving contact line and variable density. Journal of Computational Physics, 353(??):26–45, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307398.

Hollbacher:2019:RTS

[HW19] Susanne Höllbacher and Gabriel Wittum. Rotational test spaces for a fully-implicit FVM and FEM for the DNS of fluid-particle interaction. Journal of Computational Physics, 393(??):186–213, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303298.

Hu:2015:ELD

[HWA15] X. Y. Hu, B. Wang, and N. A. Adams. An efficient lowdissipation hybrid weighted essentially non-oscillatory scheme. Journal of Computational Physics, 301(??):415–424, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005719.

Hwang:2016:MMT

[Hwa16] Yao-Hsin Hwang. Macroscopic model and truncation error of discrete Boltzmann method. Journal of Computational Physics, 322(??):52–73, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302583.

Huang:2019:LBM

[HWA19] Rongzong Huang, Huiying Wu, and Nikolaus A. Adams. Lattice Boltzmann model with adjustable equation of state for coupled thermo-hydrodynamic flows. Journal of Computational Physics, 392(??):227–247, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302906.

Han:2016:IFE

[HWH^{+}16] Daoru Han, Pu Wang, Xiaoming He, Tao Lin, and Joseph Wang. A 3D immersed finite element method with nonhomogeneous interface flux jump for applications in particlein-cell simulations of plasma-lunar surface interactions. Journal of Computational Physics, 321(??):965–980, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302029.

Haghshenas:2019:FVG

[HWK19] Majid Haghshenas, James A. Wilson, and Ranganathan Kumar. Finite volume Ghost Fluid Method implementation of interfacial forces in PISO loop. Journal of Computational Physics, 376(??):20–27, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306259.

He:2016:NSM

[HX16] Yanyan He and Dongbin Xiu. Numerical strategy for model correction using physical constraints. Journal of Computational Physics, 313(??):617–634, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001261.

Huang:2015:ARB

[HXB15] Zhu Huang, Jianping Xiao, and John P. Boyd. Adaptive radial basis function and Hermite function pseudospectral methods for computing eigenvalues of the prolate spheroidal wave equation for very large bandwidth parameter. Journal of Computational Physics, 281(??):269–284, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007062.

Hao:2015:FDM

[HXLL15] Wenrui Hao, Zhiliang Xu, Chun Liu, and Guang Lin. A fictitious domain method with a hybrid cell model for simulating motion of cells in fluid flow. Journal of Computational Physics, 280(??):345–362, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006561.

Hu:2018:MCA

[HXX18] Guanghui Hu, Hehu Xie, and Fei Xu. A multilevel correction adaptive finite element method for Kohn–Sham equation. Journal of Computational Physics, 355(??):436–449, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308628.

Huang:2015:BCL

[HY15] Juntao Huang and Wen-An Yong. Boundary conditions of the lattice Boltzmann method for convection-diffusion equations. Journal of Computational Physics, 300(??):70–91, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004908.

Hu:2016:AFV

[HY16] Guanghui Hu and Nianyu Yi. An adaptive finite volume solver for steady Euler equations with non-oscillatory k-exact reconstruction. Journal of Computational Physics, 312(??):235–251, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000735.

Hwang:2017:CSN

[HY17] Yao-Hsin Hwang and Jui-Ling Yu. Construction and simulation of a novel continuous traffic flow model. Journal of Computational Physics, 350(??):927–950, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306551.

Hager:2016:FNL

[HYK^{+}16] Robert Hager, E. S. Yoon, S. Ku, E. F. D’Azevedo, P. H. Worley, and C. S. Chang. A fully non-linear multi-species Fokker–Planck–Landau collision operator for simulation of fusion plasma. Journal of Computational Physics, 315(??):644– 660, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300298.

Hu:2017:APC

[HYL17] Zixi Hu, Zhewei Yao, and Jinglai Li. On an adaptive preconditioned Crank–Nicolson MCMC algorithm for infinite dimensional Bayesian inference. Journal of Computational Physics, 332(??):492–503, March 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306106.

Huang:2015:CVH

[HZ15] Can Huang and Zhimin Zhang. Convergence of a p-version/hp-version method for fractional differential equations. Journal of Computational Physics, 286(??):118–127, April 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000297.

Hou:2017:IIM

[HZ17] Songming Hou and Sui Zhang. An improved imaging method for extended targets. Journal of Computational Physics, 333(??):321–330, March 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116307239.

Han:2019:SME

[HZE19] Jiequn Han, Linfeng Zhang, and Weinan E. Solving manyelectron Schrödinger equation using deep neural networks. Journal of Computational Physics, 399(??):Article 108929, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119306345.

He:2015:PNO

[HZL^{+}15] Zhiwei He, Yousheng Zhang, Xinliang Li, Li Li, and Baolin Tian. Preventing numerical oscillations in the flux-split based finite difference method for compressible flows with discontinuities. Journal of Computational Physics, 300(??):269–287, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005021.

Innocenti:2016:MCM

[IBML16] M. E. Innocenti, A. Beck, S. Markidis, and G. Lapenta. Momentum conservation in Multi-Level Multi-Domain (MLMD) simulations. Journal of Computational Physics, 312(??):14– 18, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000905.

Ismail:2017:DES

[IC17] Farzad Ismail and Hossain Chizari. Developments of entropy-stable residual distribution methods for conservation laws I: Scalar problems. Journal of Computational Physics, 330(??):1093–1115, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305733.

Ireland:2017:IPD

[ID17] Peter J. Ireland and Olivier Desjardins. Improving particle drag predictions in Euler–Lagrange simulations with two-way coupling. Journal of Computational Physics, 338(??):405–430, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301730.

Idomura:2016:NHK

[Ido16] Y. Idomura. A new hybrid kinetic electron model for full-f gyrokinetic simulations. Journal of Computational Physics, 313(??):511–531, May 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001297.

Ivan:2015:HOC

[IDSG15] L. Ivan, H. De Sterck, A. Susanto, and C. P. T. Groth. Highorder central ENO finite-volume scheme for hyperbolic conservation laws on three-dimensional cubed-sphere grids. Journal of Computational Physics, 282(??):157–182, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007542.

Imbert-Gerard:2015:WPG

[IG15] Lise-Marie Imbert-Gérard. Well-posedness and generalized plane waves simulations of a 2D mode conversion model. Journal of Computational Physics, 303(??):105–124, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006269.

Isola:2015:FVS

[IGQ15] D. Isola, A. Guardone, and G. Quaranta. Finite-volume solution of two-dimensional compressible flows over dynamic adaptive grids. Journal of Computational Physics, 285(??):1–23, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000091.

Iga:2015:SSS

[iI15] Shin ichi Iga. Smooth, seamless, and structured grid generation with flexibility in resolution distribution on a sphere based on conformal mapping and the spring dynamics method. Journal of Computational Physics, 297(??):381–406, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003605.

Iga:2017:EEG

[iI17] Shin ichi Iga. An equatorially enhanced grid with smooth resolution distribution generated by a spring dynamics method. Journal of Computational Physics, 330(??):794–809, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305150.

Ikeno:2018:SBT

[Ike18] Hidekazu Ikeno. Spherical Bessel transform via exponential sum approximation of spherical Bessel function. Journal of Computational Physics, 355(??):426–435, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308549.

Imazawa:2015:MMS

[IKI15] Ryota Imazawa, Yasunori Kawano, and Kiyoshi Itami. Meshless method for solving fixed boundary problem of plasma equilibrium. Journal of Computational Physics, 292(??):208–214, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001904.

Iserles:2019:SSE

[IKS19] Arieh Iserles, Karolina Kropielnicka, and Pranav Singh. Solving Schrödinger equation in semiclassical regime with highly oscillatory time-dependent potentials. Journal of Computational Physics, 376(??):564–584, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306491.

Ivey:2015:AIN

[IM15] Christopher B. Ivey and Parviz Moin. Accurate interface normal and curvature estimates on three-dimensional unstructured non-convex polyhedral meshes. Journal of Computational Physics, 300(??):365–386, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005082.

Irfan:2017:FTM

[IM17a] Muhammad Irfan and Metin Muradoglu. A front tracking method for direct numerical simulation of evaporation process in a multiphase system. Journal of Computational Physics, 337(??):132–153, May 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301304.

Ivey:2017:CBV

[IM17b] Christopher B. Ivey and Parviz Moin. Conservative and bounded volume-of-fluid advection on unstructured grids. Journal of Computational Physics, 350(??):387–419, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306356.

Imbert:2015:FDM

[IML15] D. Imbert, S. McNamara, and Y. Le Gonidec. Fictitious domain method for acoustic waves through a granular suspension of movable rigid spheres. Journal of Computational Physics, 280(??):676–691, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006809.

Ishii:2017:VRT

[ION^{+}17] Ayako Ishii, Naofumi Ohnishi, Hiroki Nagakura, Hirotaka Ito, and Shoichi Yamada. Validation of radiative transfer computation with Monte Carlo method for ultra-relativistic background flow. Journal of Computational Physics, 348(??):612–633, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305466.

Isaac:2015:SEA

[IPSG15] Tobin Isaac, Noemi Petra, Georg Stadler, and Omar Ghattas. Scalable and efficient algorithms for the propagation of uncertainty from data through inference to prediction for large-scale problems, with application to flow of the Antarctic ice sheet. Journal of Computational Physics, 296(??):348–368, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003046.

Ismagilov:2015:SOF

[Ism15] Timur Z. Ismagilov. Second order finite volume scheme for Maxwell’s equations with discontinuous electromagnetic properties on unstructured meshes. Journal of Computational Physics, 282(??):33–42, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007530.

Itu:2015:PEF

[ISP^{+}15] Lucian Itu, Puneet Sharma, Tiziano Passerini, Ali Kamen, Constantin Suciu, and Dorin Comaniciu. A parameter estimation framework for patient-specific hemodynamic computations. Journal of Computational Physics, 281(??):316–333, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007165.

Ii:2018:CSA

[ISST18] Satoshi Ii, Kazuya Shimizu, Kazuyasu Sugiyama, and Shu Takagi. Continuum and stochastic approach for cell adhesion process based on Eulerian fluid-capsule coupling with Lagrangian markers. Journal of Computational Physics, 374(??):769–786, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305229.

Iwasaki:2015:MDE

[Iwa15] Kazunari Iwasaki. Minimizing dispersive errors in smoothed particle magnetohydrodynamics for strongly magnetized medium. Journal of Computational Physics, 302(??):359–373, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006154.

Isaacson:2018:UMC

[IZ18] Samuel A. Isaacson and Ying Zhang. An unstructured mesh convergent reaction-diffusion master equation for reversible reactions. Journal of Computational Physics, 374(??):954–983, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304972.

Jackson:2017:FNS

[Jac17a] Haran Jackson. A fast numerical scheme for the Godunov– Peshkov–Romenski model of continuum mechanics. Journal of Computational Physics, 348(??):514–533, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305636.

Jackson:2017:EAW

[Jac17b] Haran Jackson. On the eigenvalues of the ADER–WENO Galerkin predictor. Journal of Computational Physics, 333 (??):409–413, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117300086.

Jaiswal:2019:DGF

[JAH19] Shashank Jaiswal, Alina A. Alexeenko, and Jingwei Hu. A discontinuous Galerkin fast spectral method for the full Boltzmann equation with general collision kernels. Journal of Computational Physics, 378(??):178–208, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307198.

Javaloyes:2015:RCS

[JB15] J. Javaloyes and S. Balle. Rational Chebyshev spectral transform for the dynamics of broad-area laser diodes. Journal of Computational Physics, 298(??):801–815, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004246.

Jofre:2015:PLB

[JBLO15] Lluís Jofre, Ricard Borrell, Oriol Lehmkuhl, and Assensi Oliva. Parallel load balancing strategy for volume-of-fluid methods on 3-D unstructured meshes. Journal of Computational Physics, 282(??):269–288, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400761X.

Jedouaa:2019:EIC

[JBM19] M. Jedouaa, C.-H. Bruneau, and E. Maitre. An efficient interface capturing method for a large collection of interacting bodies immersed in a fluid. Journal of Computational Physics, 378(??):143–177, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307241.

Joshi:2017:HOM

[JC17] S. M. Joshi and A. Chatterjee. Higher-order multilevel framework for ADER scheme in computational aeroacoustics. Journal of Computational Physics, 338(??):388–404, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301651.

Joshaghani:2019:CBS

[JCNK19] M. S. Joshaghani, J. Chang, K. B. Nakshatrala, and M. G. Knepley. Composable block solvers for the fourfield double porosity/permeability model. Journal of Computational Physics, 386(??):428–466, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301378.

Jing:2019:HET

[JCWX19] Hao Jing, Yushu Chen, Jian Wang, and Wei Xue. A highly efficient time-space-domain optimized method with Lax–Wendroff type time discretization for the scalar wave equation. Journal of Computational Physics, 393(??):1–28, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303195.

Jiang:2019:IDA

[JD19] Rui Jiang and Louis J. Durlofsky. Implementation and detailed assessment of a GNAT reduced-order model for subsurface flow simulation. Journal of Computational Physics, 379(??):192–213, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118307836.

Josey:2016:WMC

[JDFS16] C. Josey, P. Ducru, B. Forget, and K. Smith. Windowed multipole for cross section Doppler broadening. Journal of Computational Physics, 307(??):715–727, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500532X.

Jansen:2018:TCC

[JdR^{+}18] O. Jansen, E. d’Humières, X. Ribeyre, S. Jequier, and V. T. Tikhonchuk. Tree code for collision detection of large numbers of particles applied to the Breit–Wheeler process. Journal of Computational Physics, 355(??):582–596, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308598.

Jermyn:2019:ETD

[Jer19] Adam S. Jermyn. Efficient tree decomposition of high-rank tensors. Journal of Computational Physics, 377(??):142–154, January 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306934.

Jakeman:2015:EIM

[JES15] J. D. Jakeman, M. S. Eldred, and K. Sargsyan. Enhancing ℓ_{1}- minimization estimates of polynomial chaos expansions using basis selection. Journal of Computational Physics, 289(??):18– 34, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000959.

Josey:2017:HOM

[JFS17] C. Josey, B. Forget, and K. Smith. High order methods for the integration of the Bateman equations and other problems of the form of y′ = F (y, t)y. Journal of Computational Physics, 350(??):296–313, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730596X.

Johnen:2015:GVC

[JG15] A. Johnen and C. Geuzaine. Geometrical validity of curvilinear pyramidal finite elements. Journal of Computational Physics, 299(??):124–129, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004313.

Jenny:2019:APT

[JG19] Patrick Jenny and Hossein Gorji. Accurate particle time integration for solving Vlasov–Fokker–Planck equations with specified electromagnetic fields. Journal of Computational Physics, 387(??):430–445, June 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301573.

Johansson:2016:TDC

[JGS16] A. Johansson, M. Garzon, and J. A. Sethian. A threedimensional coupled Nitsche and level set method for electrohydrodynamic potential flows in moving domains. Journal of Computational Physics, 309(??):88–111, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008451.

Jenkins:2015:CEM

[JH15] Thomas G. Jenkins and Eric D. Held. Coupling extended magnetohydrodynamic fluid codes with radiofrequency ray tracing codes for fusion modeling. Journal of Computational Physics, 297(??):427–441, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003691.

Jibben:2017:AOR

[JH17] Z. Jibben and M. Herrmann. An arbitrary-order Runge– Kutta discontinuous Galerkin approach to reinitialization for banded conservative level sets. Journal of Computational Physics, 349(??):453–473, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730606X.

Jerez-Hanckes:2017:MSF

[JHPAT17] Carlos Jerez-Hanckes, Carlos Pérez-Arancibia, and Catalin Turc. Multitrace/singletrace formulations and domain decomposition methods for the solution of Helmholtz transmission problems for bounded composite scatterers. Journal of Computational Physics, 350(??):343–360, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306319.

Juno:2018:DGA

[JHT^{+}18] J. Juno, A. Hakim, J. TenBarge, E. Shi, and W. Dorland. Discontinuous Galerkin algorithms for fully kinetic plasmas. Journal of Computational Physics, 353(??):110–147, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307477.

Joshi:2017:PPV

[JJ17] Vaibhav Joshi and Rajeev K. Jaiman. A positivity preserving variational method for multi-dimensional convectiondiffusion-reaction equation. Journal of Computational Physics, 339(??):247–284, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117301924.

Joshi:2018:AVP

[JJ18a] Vaibhav Joshi and Rajeev K. Jaiman. An adaptive variational procedure for the conservative and positivity preserving Allen–Cahn phase-field model. Journal of Computational Physics, 366(??):478–504, August 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302389.

Joshi:2018:PPC

[JJ18b] Vaibhav Joshi and Rajeev K. Jaiman. A positivity preserving and conservative variational scheme for phase-field modeling of two-phase flows. Journal of Computational Physics, 360(??):137–166, May 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830038X.

Johnson:2019:CGR

[JJ19] Philip E. Johnson and Eric Johnsen. The compact gradient recovery discontinuous Galerkin method for diffusion problems. Journal of Computational Physics, 398(??):Article 108872, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119305637.

Jarecka:2015:SDS

[JJS15] Dorota Jarecka, Anna Jaruga, and Piotr K. Smolarkiewicz. A spreading drop of shallow water. Journal of Computational Physics, 289(??):53–61, May 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000649.

Jeffers:2017:GBH

[JKE^{+}17] R. S. Jeffers, J. Kópházi, M. D. Eaton, F. Févotte, F. Hülsemann, and J. Ragusa. Goal-based h-adaptivity of the 1-D diamond difference discrete ordinate method. Journal of Computational Physics, 335(??):179–200, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300530.

Jain:2019:CND

[JKM19] Suhas S. Jain, Ken Kamrin, and Ali Mani. A conservative and non-dissipative Eulerian formulation for the simulation of soft solids in fluids. Journal of Computational Physics, 399(??):Article 108922, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306278.

Jiang:2015:MEL

[JL15] Lijian Jiang and Xinping Li. Multi-element least square HDMR methods and their applications for stochastic multiscale model reduction. Journal of Computational Physics, 294(??):439–461, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002302.

Jiang:2016:AAN

[JL16] Shidong Jiang and Li-Shi Luo. Analysis and accurate numerical solutions of the integral equation derived from the linearized BGKW equation for the steady Couette flow. Journal of Computational Physics, 316(??):416–434, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300432.

Jang:2017:IBM

[JL17a] Juwon Jang and Changhoon Lee. An immersed boundary method for nonuniform grids. Journal of Computational Physics, 341(??):1–12, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302851.

Jiang:2017:MSR

[JL17b] Lijian Jiang and Qiuqi Li. Model’s sparse representation based on reduced mixed GMsFE basis methods. Journal of Computational Physics, 338(??):285–312, June 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301584.

Jin:2017:APS

[JL17c] Shi Jin and Hanqing Lu. An asymptotic-preserving stochastic Galerkin method for the radiative heat transfer equations with random inputs and diffusive scalings. Journal of Computational Physics, 334(??):182–206, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306878.

Jambunathan:2018:COB

[JL18a] Revathi Jambunathan and Deborah A. Levin. CHAOS: an octree-based PIC–DSMC code for modeling of electron kinetic properties in a plasma plume using MPI– CUDA parallelization. Journal of Computational Physics, 373(??):571–604, November 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304601.

Jiang:2018:MRM

[JL18b] Lijian Jiang and Qiuqi Li. Model reduction method using variable-separation for stochastic saddle point problems. Journal of Computational Physics, 354(??):43–66, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308239.

Jiang:2018:IRP

[JL18c] Yi Jiang and Hailiang Liu. Invariant-region-preserving DG methods for multi-dimensional hyperbolic conservation law systems, with an application to compressible Euler equations. Journal of Computational Physics, 373(??):385–409, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301566.

Jiang:2019:BTB

[JL19] Maoqiang Jiang and Zhaohui Liu. A boundary thickeningbased direct forcing immersed boundary method for fully resolved simulation of particle-laden flows. Journal of Computational Physics, 390(??):203–231, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930230X.

Jiang:2015:FPL

[JLC15] Liang Jiang, Fengbin Liu, and Darong Chen. A fast particle level set method with optimized particle correction procedure for interface capturing. Journal of Computational Physics, 299(??):804–819, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004374.

Jin:2018:OLD

[JLC18] Yao Jin, Fei Liao, and Jinsheng Cai. Optimized low-dissipation and low-dispersion schemes for compressible flows. Journal of Computational Physics, 371(??):820–849, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303802.

Jin:2017:NEC

[JLKF17] Hanhui Jin, Ningning Liu, Xiaoke Ku, and Jianren Fan. A novel energy conversion based method for velocity correction in molecular dynamics simulations. Journal of Computational Physics, 336(??):569–579, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301213.

Jin:2015:GFE

[JLLZ15] Bangti Jin, Raytcho Lazarov, Yikan Liu, and Zhi Zhou. The Galerkin finite element method for a multi-term timefractional diffusion equation. Journal of Computational Physics, 281(??):825–843, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114007396.

Jang:2015:HOA

[JLQX15] Juhi Jang, Fengyan Li, Jing-Mei Qiu, and Tao Xiong. High order asymptotic preserving DG–IMEX schemes for discretevelocity kinetic equations in a diffusive scaling. Journal of Computational Physics, 281(??):199–224, January 15, 2015.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007074.

Jacobs:2018:ADV

[JME18] Matt Jacobs, Ekaterina Merkurjev, and Selim Esedoğlu. Auction dynamics: a volume constrained MBO scheme. Journal of Computational Physics, 354(??):288–310, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308033.

Jeanmasson:2019:SEL

[JMM19] G. Jeanmasson, I. Mary, and L. Mieussens. On some explicit local time stepping finite volume schemes for CFD. Journal of Computational Physics, 397(??):Article 108818, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305029.

Jackson:2019:NSN

[JN19] Haran Jackson and Nikos Nikiforakis. A numerical scheme for non-Newtonian fluids and plastic solids under the GPR model. Journal of Computational Physics, 387(??):410–429, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301421.

Jouvet:2015:MSS

[Jou15] Guillaume Jouvet. Multilayer shallow shelf approximation: Minimisation formulation, finite element solvers and applications. Journal of Computational Physics, 287(??):60–76, 2015. CODEN JCTPAH. ISSN 0021-9991. URL http://www.

sciencedirect.com/science/article/pii/S0021999115000674.

Jeong:2015:CAT

[JPLL15] Yongjin Jeong, Jinsu Park, Hyun Chul Lee, and Deokjung Lee. Convergence analysis of two-node CMFD method for two-group neutron diffusion eigenvalue problem. Journal of Computational Physics, 302(??):239–250, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005847.

Ji:2018:CFO

[JPSX18] Xing Ji, Liang Pan, Wei Shyy, and Kun Xu. A compact fourthorder gas-kinetic scheme for the Euler and Navier–Stokes equations. Journal of Computational Physics, 372(??):446–472, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304121.

Jarauta:2018:IST

[JRPPS18] Alex Jarauta, Pavel Ryzhakov, Jordi Pons-Prats, and Marc Secanell. An implicit surface tension model for the analysis of droplet dynamics. Journal of Computational Physics, 374(??):1196–1218, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305217.

Jin:2016:MMT

[JS16] Miaomiao Jin and Michael Short. Multiphysics modeling of two-phase film boiling within porous corrosion deposits. Journal of Computational Physics, 316(??):504–518, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001650.

Jin:2017:SAP

[JS17] Shi Jin and Ruiwen Shu. A stochastic asymptotic-preserving scheme for a kinetic-fluid model for disperse two-phase flows with uncertainty. Journal of Computational Physics, 335 (??):905–924, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911730075X.

Jones:2019:ASD

[JS19] Tiffany N. Jones and Qin Sheng. Asymptotic stability of a dual-scale compact method for approximating highly oscillatory Helmholtz solutions. Journal of Computational Physics, 392(??):403–418, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930292X.

Jiang:2019:DAF

[JSB^{+}19] Yan Jiang, Puttha Sakkaplangkul, Vrushali A. Bokil, Yingda Cheng, and Fengyan Li. Dispersion analysis of finite difference and discontinuous Galerkin schemes for Maxwell’s equations in linear Lorentz media. Journal of Computational Physics, 394(??):100–135, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303535.

Jaensch:2016:RFC

[JSP16] S. Jaensch, C. Sovardi, and W. Polifke. On the robust, flexible and consistent implementation of time domain impedance boundary conditions for compressible flow simulations. Journal of Computational Physics, 314(??):145–159, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001625.

Jemison:2015:FCM

[JSS15] Matthew Jemison, Mark Sussman, and Mikhail Shashkov. Filament capturing with the Multimaterial Moment-of-Fluid method. Journal of Computational Physics, 285(??):149–172, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000182.

Jiang:2017:AMC

[JST17] Chenfanfu Jiang, Craig Schroeder, and Joseph Teran. An angular momentum conserving affine-particle-in-cell method. Journal of Computational Physics, 338(??):137–164, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301535.

Jareteg:2017:NFB

[JSVD17] Klas Jareteg, Srdjan Sasic, Paolo Vinai, and Christophe Demazière. A numerical framework for bubble transport in a subcooled fluid flow. Journal of Computational Physics, 345(??):373–403, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304084.

Jung:2015:TDC

[JSY15] Narina Jung, Hae Won Seo, and Chun Sang Yoo. Twodimensional characteristic boundary conditions for open boundaries in the lattice Boltzmann methods. Journal of Computational Physics, 302(??):191–199, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005720.

Jiang:2018:DBC

[JT18] Jiamin Jiang and Hamdi A. Tchelepi. Dissipation-based continuation method for multiphase flow in heterogeneous porous media. Journal of Computational Physics, 375(??): 307–336, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118305758.

Joshi:2016:DAP

[JTD16] Sumedh M. Joshi, Greg N. Thomsen, and Peter J. Diamessis. Deflation-accelerated preconditioning of the Poisson– Neumann Schur problem on long domains with a high-order discontinuous element-based collocation method. Journal of Computational Physics, 313(??):209–232, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000978.

Jamshidian:2016:MCF

[JTR16] M. Jamshidian, P. Thamburaja, and T. Rabczuk. A multiscale coupled finite-element and phase-field framework to modeling stressed grain growth in polycrystalline thin films. Journal of Computational Physics, 327(??):779–798, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630482X.

Jakeman:2015:EAS

[JW15a] J. D. Jakeman and T. Wildey. Enhancing adaptive sparse grid approximations and improving refinement strategies using adjoint-based a posteriori error estimates. Journal of Computational Physics, 280(??):54–71, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006500.

Jia:2015:FFD

[JW15b] Jinhong Jia and Hong Wang. Fast finite difference methods for space-fractional diffusion equations with fractional derivative boundary conditions. Journal of Computational Physics, 293(??):359–369, July 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005750.

Jia:2015:PFF

[JW15c] Jinhong Jia and Hong Wang. A preconditioned fast finite volume scheme for a fractional differential equation discretized on a locally refined composite mesh. Journal of Computational Physics, 299(??):842–862, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004258.

Jia:2016:FFV

[JW16] Jinhong Jia and Hong Wang. A fast finite volume method for conservative space-fractional diffusion equations in convex domains. Journal of Computational Physics, 310(??):63–84, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000164.

Jiao:2016:WCF

[JWH16] Yujian Jiao, Li-Lian Wang, and Can Huang. Wellconditioned fractional collocation methods using fractional Birkhoff interpolation basis. Journal of Computational Physics, 305(??):1–28, January 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006932.

Jiang:2015:MMS

[JX15] Yingjun Jiang and Xuejun Xu. Multigrid methods for space fractional partial differential equations. Journal of Computational Physics, 302(??):374–392, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005872.

Jiang:2017:DDM

[JX17] Yingjun Jiang and Xuejun Xu. Domain decomposition methods for space fractional partial differential equations. Journal of Computational Physics, 350(??):573–589, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306472.

Jin:2015:APM

[JXZ15] Shi Jin, Dongbin Xiu, and Xueyu Zhu. Asymptoticpreserving methods for hyperbolic and transport equations with random inputs and diffusive scalings. Journal of Computational Physics, 289(??):35–52, May 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000935.

Jiang:2018:EMH

[JYY18] Zhen-Hua Jiang, Chao Yan, and Jian Yu. Efficient methods with higher order interpolation and MOOD strategy for compressible turbulence simulations. Journal of Computational Physics, 371(??):528–550, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303966.

Jiang:2016:KSS

[JZ16] Tian Jiang and Yong-Tao Zhang. Krylov single-step implicit integration factor WENO methods for advectiondiffusion-reaction equations. Journal of Computational Physics, 311(??):22–44, April 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116000292.

Ji:2018:FHO

[JZSX18] Xing Ji, Fengxiang Zhao, Wei Shyy, and Kun Xu. A family of high-order gas-kinetic schemes and its comparison with Riemann solver based high-order methods. Journal of Computational Physics, 356(??):150–173, March 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308744.

Kew:2015:NEG

[KA15] Lee Ming Kew and Norhashidah Hj. Mohd Ali. New explicit group iterative methods in the solution of three dimensional hyperbolic telegraph equations. Journal of Computational Physics, 294(??):382–404, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002077.

Ke:2018:NPT

[KA18] G. Ke and E. Aulisa. New preconditioning techniques for the steady and unsteady buoyancy driven flow problems. Journal of Computational Physics, 371(??):244–260, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303462.

Khosravian-Arab:2015:FSL

[KADE15] Hassan Khosravian-Arab, Mehdi Dehghan, and M. R. Eslahchi. Fractional Sturm–Liouville boundary value problems in unbounded domains: Theory and applications. Journal of Computational Physics, 299(??):526–560, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004283.

Khosravian-Arab:2017:FSP

[KADE17] Hassan Khosravian-Arab, Mehdi Dehghan, and M. R. Eslahchi. Fractional spectral and pseudo-spectral methods in unbounded domains: Theory and applications. Journal of Computational Physics, 338(??):527–566, June 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301638.

Kindelan:2018:FOR

[KÁGR18] Manuel Kindelan, Diego Álvarez, and Pedro Gonzalez-Rodriguez. Frequency optimized RBF-FD for wave equations. Journal of Computational Physics, 371(??):564–580, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830384X.

Kedward:2017:EEM

[KAR17] L. Kedward, C. B. Allen, and T. C. S. Rendall. Efficient and exact mesh deformation using multiscale RBF interpolation. Journal of Computational Physics, 345(??):732–751, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304254.

Kasperski:2015:PSP

[Kas15] G. Kasperski. A parallel Schwarz preconditioner for the Chebyshev Gauss–Lobatto collocation (d^{2}/dx^{2} −h^{2}) operator. Journal of Computational Physics, 296(??):101–112, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500306X.

Katugampola:2016:CWF

[Kat16] Udita N. Katugampola. Correction to “What is a fractional derivative?” by Ortigueira and Machado [Journal of Computational Physics, Volume 293, 15 July 2015, Pages 4–13. Special issue on Fractional PDEs]. Journal of Computational Physics, 321(??):1255–1257, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301978.

Kaya:2015:FDAa

[Kay15] Adem Kaya. Finite difference approximations of multidimensional unsteady convection–diffusion–reaction equations. Journal of Computational Physics, 285(??):331–349, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000285.

Kawazura:2018:HGI

[KB18] Y. Kawazura and M. Barnes. A hybrid gyrokinetic ion and isothermal electron fluid code for astrophysical plasma. Journal of Computational Physics, 360(??):57–73, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300366.

Kromer:2019:HAC

[KB19] Johannes Kromer and Dieter Bothe. Highly accurate computation of volume fractions using differential geometry. Journal of Computational Physics, 396(??):761–784, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304899.

Kumari:2019:UAD

[KBD19] Komal Kumari, Raktim Bhattacharya, and Diego A. Donzis. A unified approach for deriving optimal finite differences. Journal of Computational Physics, 399(??):Article 108957, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S002199911930662X.

Kursawe:2017:IIC

[KBF17] Jochen Kursawe, Ruth E. Baker, and Alexander G. Fletcher. Impact of implementation choices on quantitative predictions of cell-based computational models. Journal of Computational Physics, 345(??):752–767, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304412.

Kou:2015:FRA

[KBG^{+}15] Wenjun Kou, Amneet Pal Singh Bhalla, Boyce E. Griffith, John E. Pandolfino, Peter J. Kahrilas, and Neelesh A. Patankar. A fully resolved active musculo-mechanical model for esophageal transport. Journal of Computational Physics, 298(??):446–465, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003897.

Kampmann:2015:PEC

[KBK15a] Tobias A. Kampmann, Horst-Holger Boltz, and Jan Kierfeld. Parallelized event chain algorithm for dense hard sphere and polymer systems. Journal of Computational Physics, 281(??):864–875, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114007475.

Kim:2015:QSU

[KBK15b] Changho Kim, Oleg Borodin, and George Em Karniadakis. Quantification of sampling uncertainty for molecular dynamics simulation: Time-dependent diffusion coefficient in simple fluids. Journal of Computational Physics, 302(??):485–508, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006142.

Koulouri:2017:VTR

[KBR17] Alexandra Koulouri, Mike Brookes, and Ville Rimpiläinen. Vector tomography for reconstructing electric fields with nonzero divergence in bounded domains. Journal of Computational Physics, 329(??):73–90, January 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305356.

Kang:2017:EIH

[KC17a] Hyun-Gyu Kang and Hyeong-Bin Cheong. An efficient implementation of a high-order filter for a cubed-sphere spectral element model. Journal of Computational Physics, 332(??):66– 82, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306428.

Keady:2017:IRW

[KC17b] Kendra P. Keady and Mathew A. Cleveland. An improved random walk algorithm for the implicit Monte Carlo method. Journal of Computational Physics, 328(??):160–176, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304776.

Kong:2017:SNF

[KC17c] Fande Kong and Xiao-Chuan Cai. A scalable nonlinear fluidstructure interaction solver based on a Schwarz preconditioner with isogeometric unstructured coarse spaces in 3D.

Journal of Computational Physics, 340(??):498–518, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302383.

Kumar:2018:SSI

[KC18] Rakesh Kumar and Praveen Chandrashekar. Simple smoothness indicator and multi-level adaptive order WENO scheme for hyperbolic conservation laws. Journal of Computational Physics, 375(??):1059–1090, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306272.

Karakus:2019:DGD

[KCHW19] A. Karakus, N. Chalmers, J. S. Hesthaven, and T. Warburton. Discontinuous Galerkin discretizations of the Boltzmann– BGK equations for nearly incompressible flows: Semi-analytic time stepping and absorbing boundary layers. Journal of Computational Physics, 390(??):175–202, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302335.

Komen:2017:QMN

[KCS^{+}17] E. M. J. Komen, L. H. Camilo, A. Shams, B. J. Geurts, and B. Koren. A quantification method for numerical dissipation in quasi-DNS and under-resolved DNS, and effects of numerical dissipation in quasi-DNS and under-resolved DNS of turbulent channel flows. Journal of Computational Physics, 345(??):565–595, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304059.

Karakus:2019:GAD

[KCSW19] A. Karakus, N. Chalmers, K. ’Swirydowicz, and T. Warburton. A GPU accelerated discontinuous Galerkin incompressible flow solver. Journal of Computational Physics, 390(??):380–404, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302438.

Kim:2017:BFD

[KCW17] Hyea Hyun Kim, Eric Chung, and Junxian Wang. BDDC and FETI-DP preconditioners with adaptive coarse spaces for three-dimensional elliptic problems with oscillatory and high contrast coefficients. Journal of Computational Physics, 349(??):191–214, November 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305740.

Kalita:2017:NHA

[KD17a] Paragmoni Kalita and Anoop K. Dass. A novel hybrid approach with multidimensional-like effects for compressible flow computations. Journal of Computational Physics, 340(??):55– 68, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302280.

Korn:2017:EDA

[KD17b] P. Korn and S. Danilov. Elementary dispersion analysis of some mimetic discretizations on triangular C-grids. Journal of Computational Physics, 330(??):156–172, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305678.

Kast:2015:OTF

[KDF15] Steven M. Kast, Johann P. S. Dahm, and Krzysztof J. Fidkowski. Optimal test functions for boundary accuracy in discontinuous finite element methods. Journal of Computational Physics, 298(??):360–386, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003885.

Kimmritz:2015:CME

[KDL15] Madlen Kimmritz, Sergey Danilov, and Martin Losch. On the convergence of the modified elastic-viscous-plastic method for solving the sea ice momentum equation. Journal of Computational Physics, 296(??):90–100, September 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003083.

Krause:2015:TLS

[KDPK15] Dorian Krause, Thomas Dickopf, Mark Potse, and Rolf Krause. Towards a large-scale scalable adaptive heart model using shallow tree meshes. Journal of Computational Physics, 298(??):79–94, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500323X.

Kilic:2015:SIS

[KE15] Emre Kiliç and Thomas F. Eibert. Solution of 3D inverse scattering problems by combined inverse equivalent current and finite element methods. Journal of Computational Physics, 288(??):131–149, May 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000650.

Khayrat:2018:AMF

[KEJ18] Karim Khayrat, Robert Epp, and Patrick Jenny. Approximate multiscale flow solver for unstructured pore networks. Journal of Computational Physics, 372(??):62–79, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303528.

Kohn:2018:PMC

[KES18] Christoph Köhn, Martin Bødker Enghoff, and Henrik Svensmark. A 3D particle Monte Carlo approach to studying nucleation. Journal of Computational Physics, 363(??):30–38, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301141.

Kallinderis:2015:PMQ

[KF15] Y. Kallinderis and S. Fotia. A priori mesh quality metrics for three-dimensional hybrid grids. Journal of Computational Physics, 280(??):465–488, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400672X.

Kong:2017:SAF

[KF17] Bo Kong and Rodney O. Fox. A solution algorithm for fluidparticle flows across all flow regimes. Journal of Computational Physics, 344(??):575–594, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303820.

Kidder:2017:STB

[KFF^{+}17] Lawrence E. Kidder, Scott E. Field, François Foucart, Erik Schnetter, Saul A. Teukolsky, Andy Bohn, Nils Deppe, Peter Diener, François Hébert, Jonas Lippuner, Jonah Miller, Christian D. Ott, Mark A. Scheel, and Trevor Vincent. SpECTRE: a task-based discontinuous Galerkin code for relativistic astrophysics. Journal of Computational Physics, 335(??):84–114, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300098.

Kochunas:2017:FAI

[KFL17] Brendan Kochunas, Andrew Fitzgerald, and Edward Larsen. Fourier analysis of iteration schemes for k-eigenvalue transport problems with flux-dependent cross sections. Journal of Computational Physics, 345(??):294–307, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304035.

Krank:2017:HOS

[KFWK17] Benjamin Krank, Niklas Fehn, Wolfgang A. Wall, and Martin Kronbichler. A high-order semi-explicit discontinuous Galerkin solver for 3D incompressible flow with application to DNS and LES of turbulent channel flow. Journal of Computational Physics, 348(??):634–659, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305478.

Kopera:2015:MCU

[KG15] Michal A. Kopera and Francis X. Giraldo. Mass conservation of the unified continuous and discontinuous elementbased Galerkin methods on dynamically adaptive grids with application to atmospheric simulations. Journal of Computational Physics, 297(??):90–103, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003290.

Kou:2017:CMB

[KGP^{+}17] Wenjun Kou, Boyce E. Griffith, John E. Pandolfino, Peter J. Kahrilas, and Neelesh A. Patankar. A continuum mechanicsbased musculo-mechanical model for esophageal transport. Journal of Computational Physics, 348(??):433–459, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305338.

Khayyer:2017:CSA

[KGS17] Abbas Khayyer, Hitoshi Gotoh, and Yuma Shimizu. Comparative study on accuracy and conservation properties of two particle regularization schemes and proposal of an optimized particle shifting scheme in ISPH context. Journal of Computational Physics, 332(??):236–256, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306465.

Khirevich:2015:CFG

[KGT15] Siarhei Khirevich, Irina Ginzburg, and Ulrich Tallarek. Coarse- and fine-grid numerical behavior of MRT/TRT lattice-Boltzmann schemes in regular and random sphere packings. Journal of Computational Physics, 281(??):708–742, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007207.

Kim:2015:ASE

[KH15] Jae Wook Kim and Sina Haeri. An advanced synthetic eddy method for the computation of aerofoil turbulence interaction noise. Journal of Computational Physics, 287 (??):1–17, 2015. CODEN JCTPAH. ISSN 0021-9991.

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000534.

Kitamura:2017:SPS

[KH17] Keiichi Kitamura and Atsushi Hashimoto. Simple a posteriori slope limiter (Post Limiter) for high resolution and efficient flow computations. Journal of Computational Physics, 341(??):313–340, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302735.

Krattiger:2018:GBM

[KH18] Dimitri Krattiger and Mahmoud I. Hussein. Generalized Bloch mode synthesis for accelerated calculation of elastic band structures. Journal of Computational Physics, 357(??):183–205, March 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117309002.

Ku:2016:NHL

[KHC^{+}16] S. Ku, R. Hager, C. S. Chang, J. M. Kwon, and S. E. Parker. A new hybrid-Lagrangian numerical scheme for gyrokinetic simulation of tokamak edge plasma. Journal of Computational Physics, 315(??):467–475, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300274.

Kunz:2016:IOD

[KHHN16] P. Kunz, M. Hirschler, M. Huber, and U. Nieken. Inflow/outflow with Dirichlet boundary conditions for pressure in ISPH. Journal of Computational Physics, 326(??):171–187, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630403X.

Karniadakis:2015:SIF

[KHP15] George Em Karniadakis, Jan S. Hesthaven, and Igor Podlubny. Special issue on “Fractional PDEs: Theory, Numerics, and Applications”. Journal of Computational Physics, 293(??):1–3, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002533.

Kay:2017:MLI

[KHP17] E. D. Kay, S. Hibberd, and H. Power. A multi-layer integral model for locally-heated thin film flow. Journal of Computational Physics, 336(??):51–68, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300827.

Kong:2019:SEN

[KHTZ19] Linghua Kong, Yuqi Hong, Nana Tian, and Wenying Zhou. Stable and efficient numerical schemes for two-dimensional Maxwell equations in lossy medium. Journal of Computational Physics, 397(??):Article 108703, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303614.

Kates-Harbeck:2016:SCT

[KHTZA16] Julian Kates-Harbeck, Samuel Totorica, Jonathan Zrake, and Tom Abel. Simplex-in-cell technique for collisionless plasma simulations. Journal of Computational Physics, 304 (??):231–251, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115006816.

Kim:2015:FPF

[Kim15] Seung Hyun Kim. A front propagation formulation for under-resolved reaction fronts. Journal of Computational Physics, 285(??):193–207, March 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000108.

Khayrat:2017:MSN

[KJ17a] Karim Khayrat and Patrick Jenny. A multi-scale network method for two-phase flow in porous media. Journal of Computational Physics, 342(??):194–210, August 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302942.

Kuchlin:2017:PFP

[KJ17b] Stephan Küchlin and Patrick Jenny. Parallel Fokker–Planck– DSMC algorithm for rarefied gas flow simulation in complex domains at all Knudsen numbers. Journal of Computational Physics, 328(??):258–277, January 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305162.

Kuchlin:2018:AMR

[KJ18] Stephan Küchlin and Patrick Jenny. Automatic mesh refinement and parallel load balancing for Fokker–Planck–DSMC algorithm. Journal of Computational Physics, 363(??):140–157, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301384.

Kong:2019:ASR

[KJHA19] Qian Kong, Yan-Fei Jing, Ting-Zhu Huang, and Heng-Bin An. Acceleration of the Scheduled Relaxation Jacobi method: Promising strategies for solving large, sparse linear systems. Journal of Computational Physics, 397(??):Article 108862, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305467.

Kim:2017:FDM

[KJYC17] Junseok Kim, Darae Jeong, Seong-Deog Yang, and Yongho Choi. A finite difference method for a conservative Allen– Cahn equation on non-flat surfaces. Journal of Computational Physics, 334(??):170–181, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300104.

Kozynchenko:2016:IAE

[KK16] Alexander I. Kozynchenko and Sergey A. Kozynchenko. On improving the algorithm efficiency in the particle–particle force calculations. Journal of Computational Physics, 320 (??):40–45, September 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116301747.

Kim:2017:AMC

[KK17a] Hyea Hyun Kim and Ji Eun Kim. Approximation of macroscopic conductivity for a multiscale model by using mortar methods. Journal of Computational Physics, 336(??):275–287, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301043.

Kotalczyk:2017:MCM

[KK17b] G. Kotalczyk and F. E. Kruis. A Monte Carlo method for the simulation of coagulation and nucleation based on weighted particles and the concepts of stochastic resolution and merging. Journal of Computational Physics, 340(??):276–296, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730236X.

Kannan:2018:CRF

[KKH18] Karthik Kannan, Dominic Kedelty, and Marcus Herrmann. An in-cell reconstruction finite volume method for flows of compressible immiscible fluids. Journal of Computational Physics, 373(??):784–810, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304613.

Kim:2016:EEM

[KKJB16] Philsu Kim, Junghan Kim, WonKyu Jung, and Sunyoung Bu. An error embedded method based on generalized Chebyshev polynomials. Journal of Computational Physics, 306(??):55–72, February 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115007615.

Karagiannis:2015:BMS

[KKL15] Georgios Karagiannis, Bledar A. Konomi, and Guang Lin. A Bayesian mixed shrinkage prior procedure for spatialstochastic basis selection and evaluation of gPC expansions: Applications to elliptic SPDEs. Journal of Computational Physics, 284(??):528–546, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008511.

Kikinzon:2017:ASC

[KKLS17] Evgeny Kikinzon, Yuri Kuznetsov, Konstatin Lipnikov, and Mikhail Shashkov. Approximate static condensation algorithm for solving multi-material diffusion problems on meshes non-aligned with material interfaces. Journal of Computational Physics, 347(??):416–436, October 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305028.

Kalliadasis:2015:NFE

[KKP15] S. Kalliadasis, S. Krumscheid, and G. A. Pavliotis. A new framework for extracting coarse-grained models from time series with multiscale structure. Journal of Computational Physics, 296(??):314–328, September 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003204.

Kawashima:2015:HES

[KKS15] Rei Kawashima, Kimiya Komurasaki, and Tony Schönherr. A hyperbolic-equation system approach for magnetized electron fluids in quasi-neutral plasmas. Journal of Computational Physics, 284(??):59–69, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008419.

Kawashima:2016:FSM

[KKS16] Rei Kawashima, Kimiya Komurasaki, and Tony Schönherr. A flux-splitting method for hyperbolic-equation system of magnetized electron fluids in quasi-neutral plasmas. Journal of Computational Physics, 310(??):202–212, April 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000073.

Kang:2015:POL

[KKZ15] Sung Ha Kang, Seong Jun Kim, and Haomin Zhou. Path optimization with limited sensing ability. Journal of Computational Physics, 299(??):887–901, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004829.

Kophazi:2015:SAD

[KL15] József Kópházi and Danny Lathouwers. A space-angle DGFEM approach for the Boltzmann radiation transport equation with local angular refinement. Journal of Computational Physics, 297(??):637–668, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003654.

Keady:2016:SMC

[KL16] Kendra P. Keady and Edward W. Larsen. Stability of Monte Carlo k-eigenvalue simulations with CMFD feedback. Journal of Computational Physics, 321(??):947–964, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302273.

Karagiannis:2017:BCC

[KL17a] Georgios Karagiannis and Guang Lin. On the Bayesian calibration of computer model mixtures through experimental data, and the design of predictive models. Journal of Computational Physics, 342(??):139–160, August 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302747.

Kim:2017:MFD

[KL17b] Eugenia Kim and Konstantin Lipnikov. The mimetic finite difference method for the Landau–Lifshitz equation. Journal of Computational Physics, 328(??):109–130, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305149.

Korn:2018:CDS

[KL18a] Peter Korn and Leonidas Linardakis. A conservative discretization of the shallow-water equations on triangular grids.

Journal of Computational Physics, 375(??):871–900, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305953.

Krasnopolsky:2018:CFI

[KL18b] Boris I. Krasnopolsky and Alexander A. Lukyanov. A conservative fully implicit algorithm for predicting slug flows. Journal of Computational Physics, 355(??):597–619, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308707.

Kulkarni:2019:ACE

[KL19] Chinmay S. Kulkarni and Pierre F. J. Lermusiaux. Advection without compounding errors through flow map composition. Journal of Computational Physics, 398(??):Article 108859, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119305431.

Klaij:2015:SFV

[Kla15] C. M. Klaij. On the stabilization of finite volume methods with co-located variables for incompressible flow. Journal of Computational Physics, 297(??):84–89, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500340X.

Kallinderis:2017:FFD

[KLA17] Yannis Kallinderis, Eleni M. Lymperopoulou, and Panagiotis Antonellis. Flow feature detection for grid adaptation and flow visualization. Journal of Computational Physics, 341(??):182– 207, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302644.

Kim:2018:WCI

[KLC18] Woojin Kim, Injae Lee, and Haecheon Choi. A weakcoupling immersed boundary method for fluid-structure interaction with low density ratio of solid to fluid. Journal of Computational Physics, 359(??):296–311, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300056.

Kotsalos:2019:BCG

[KLC19] Christos Kotsalos, Jonas Latt, and Bastien Chopard. Bridging the computational gap between mesoscopic and continuum modeling of red blood cells for fully resolved blood flow. Journal of Computational Physics, 398(??):Article 108905, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119306102.

Kumar:2018:MMM

[KLGO18] Prashant Kumar, Peiyao Luo, Francisco J. Gaspar, and Cornelis W. Oosterlee. A multigrid multilevel Monte Carlo method for transport in the Darcy–Stokes system. Journal of Computational Physics, 371(??):382–408, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303553.

Koblitz:2017:DNS

[KLNH17] A. R. Koblitz, S. Lovett, N. Nikiforakis, and W. D. Henshaw. Direct numerical simulation of particulate flows with an overset grid method. Journal of Computational Physics, 343(??):414–431, August 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117303376.

Kostin:2015:LTS

[KLRT15] Victor Kostin, Vadim Lisitsa, Galina Reshetova, and Vladimir Tcheverda. Local time-space mesh refinement for simulation of elastic wave propagation in multi-scale media. Journal of Computational Physics, 281(??):669–689, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007293.

Kempe:2015:IFS

[KLSF15] Tobias Kempe, Matthias Lennartz, Stephan Schwarz, and Jochen Fröhlich. Imposing the free-slip condition with a continuous forcing immersed boundary method. Journal of Computational Physics, 282(??):183–209, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007670.

Kwan:2017:FHS

[KLWQ17] Wingfai Kwan, Shingyu Leung, Xiao-Ping Wang, and Jianliang Qian. A fast Huygens sweeping method for capturing paraxial multi-color optical self-focusing in nematic liquid crystals. Journal of Computational Physics, 348(??):108–138, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730520X.

Kwon:2015:NSM

[KM15] Jihoe Kwon and J. J. Monaghan. A novel SPH method for sedimentation in a turbulent fluid. Journal of Computational Physics, 300(??):520–532, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004386.

Kammerer:2016:KMT

[KM16a] Clotilde Fermanian Kammerer and Florian Méhats. A kinetic model for the transport of electrons in a graphene layer. Journal of Computational Physics, 327(??):450–483, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304223.

Keshavarzzadeh:2016:IDN

[KM16b] V. Keshavarzzadeh and S. F. Masri. Identification of discontinuous nonlinear systems via a multivariate Padé approach. Journal of Computational Physics, 306(??):520–545, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007913.

Kanjilal:2017:GTB

[KM17] Oindrila Kanjilal and C. S. Manohar. Girsanov’s transformation based variance reduced Monte Carlo simulation schemes for reliability estimation in nonlinear stochastic dynamics.

Journal of Computational Physics, 341(??):278–294, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302486.

Kaiser:2018:SEC

[KMD^{+}18] Eurika Kaiser, Marek Morzyński, Guillaume Daviller, J. Nathan Kutz, Bingni W. Brunton, and Steven L. Brunton. Sparsity enabled cluster reduced-order models for control. Journal of Computational Physics, 352(??):388–409, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307301.

Kastner:2016:IAC

[KMdB16] Markus Kästner, Philipp Metsch, and René de Borst. Isogeometric analysis of the Cahn–Hilliard equation — a convergence study. Journal of Computational Physics, 305(??):360–371, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500724X.

Kindelan:2016:RBF

[KMGR16] Manuel Kindelan, Miguel Moscoso, and Pedro González-Rodríguez. Radial basis function interpolation in the limit of increasingly flat basis functions. Journal of Computational Physics, 307(??):225–242, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008347.

Kotyczka:2018:WFS

[KML18] Paul Kotyczka, Bernhard Maschke, and Laurent Lefèvre. Weak form of Stokes–Dirac structures and geometric discretization of port-Hamiltonian systems. Journal of Computational Physics, 361(??):442–476, May 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300822.

Kramer:2019:MPE

[KMP^{+}19] Boris Kramer, Alexandre Noll Marques, Benjamin Peherstorfer, Umberto Villa, and Karen Willcox. Multifidelity probability estimation via fusion of estimators. Journal of Computational Physics, 392(??):385–402, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303249.

Katsiolides:2018:MMC

[KMS^{+}18] Grigoris Katsiolides, Eike H. Müller, Robert Scheichl, Tony Shardlow, Michael B. Giles, and David J. Thomson. Multilevel Monte Carlo and improved timestepping methods in atmospheric dispersion modelling. Journal of Computational Physics, 354(??):320–343, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307945.

Ke:2015:FDM

[KNS15] Rihuan Ke, Michael K. Ng, and Hai-Wei Sun. A fast direct method for block triangular Toeplitz-like with tri-diagonal block systems from time-fractional partial differential equations. Journal of Computational Physics, 303(??):203–211, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006415.

Kirchhart:2017:SFV

[KO17] M. Kirchhart and S. Obi. A splitting-free vorticity redistribution method. Journal of Computational Physics, 330 (??):282–295, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306003.

Korn:2017:FUG

[Kor17] Peter Korn. Formulation of an unstructured grid model for global ocean dynamics. Journal of Computational Physics, 339(??):525–552, June 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301961.

Koutsourelakis:2016:VBS

[Kou16] P. S. Koutsourelakis. Variational Bayesian strategies for highdimensional, stochastic design problems. Journal of Computational Physics, 308(??):124–152, March 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008505.

Kim:2015:LBM

[KP15a] Seung Hyun Kim and Heinz Pitsch. On the lattice Boltzmann method for multiphase flows with large density ratios. Journal of Computational Physics, 303(??):19–27, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006221.

Knaus:2015:CAF

[KP15b] R. Knaus and C. Pantano. A computational approach to flame hole dynamics using an embedded manifold approach. Journal of Computational Physics, 296(??):209–240, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002946.

Kornet:2015:MSD

[KP15c] Kacper Kornet and Alban Pothérat. A method for spectral DNS of low Rm channel flows based on the least dissipative modes. Journal of Computational Physics, 298(??):266–279, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003526.

Kim:2018:ANT

[KPJ18] Minhyuk Kim, SangWook Park, and Hyun-Kyo Jung. An advanced numerical technique for a quasi-static electromagnetic field simulation based on the finite-difference time-domain method. Journal of Computational Physics, 373(??):917–923, November 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304935.

Kavvadias:2015:PTG

[KPKG15] I. S. Kavvadias, E. M. Papoutsis-Kiachagias, and K. C. Giannakoglou. On the proper treatment of grid sensitivities in continuous adjoint methods for shape optimization. Journal of Computational Physics, 301(??):1–18, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005318.

Kapellos:2019:UCA

[KPKGH19] Christos S. Kapellos, Evangelos M. Papoutsis-Kiachagias, Kyriakos C. Giannakoglou, and Michael Hartmann. The unsteady continuous adjoint method for minimizing flow-induced sound radiation. Journal of Computational Physics, 392 (??):368–384, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303092.

Kraus:2019:ISP

[KPP^{+}19] Johannes Kraus, Carl-Martin Pfeiler, Dirk Praetorius, Michele Ruggeri, and Bernhard Stiftner. Iterative solution and preconditioning for the tangent plane scheme in computational micromagnetics. Journal of Computational Physics, 398(??):Article 108866, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305571.

Kabacaoglu:2018:LRS

[KQB18] Gökberk Kabacaoğlu, Bryan Quaife, and George Biros. Lowresolution simulations of vesicle suspensions in 2D. Journal of Computational Physics, 357(??):43–77, March 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309178.

King:2017:SVP

[KR17] Nathan D. King and Steven J. Ruuth. Solving variational problems and partial differential equations that map between manifolds via the closest point method. Journal of Computational Physics, 336(??):330–346, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301122.

Katsoulakis:2017:SII

[KRBW17] Markos A. Katsoulakis, Luc Rey-Bellet, and Jie Wang. Scalable information inequalities for uncertainty quantification. Journal of Computational Physics, 336(??):513–545, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301134.

Kompenhans:2016:ASH

[KRFV16] Moritz Kompenhans, Gonzalo Rubio, Esteban Ferrer, and Eusebio Valero. Adaptation strategies for high order discontinuous Galerkin methods based on Tau-estimation. Journal of Computational Physics, 306(??):216–236, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500772X.

Kriel:2017:EAF

[Kri17] A. J. Kriel. Error analysis of flux limiter schemes at extrema. Journal of Computational Physics, 328(??):371–386, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305228.

Krimi:2018:SPH

[KRK^{+}18] Abdelkader Krimi, Mehdi Rezoug, Sofiane Khelladi, Xesús Nogueira, Michael Deligant, and Luis Ramírez. Smoothed particle hydrodynamics: A consistent model for interfacial multiphase fluid flow simulations. Journal of Computational Physics, 358(??):53–87, April 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308811.

Kaya:2015:FDAb

[KS15a] Adem Kaya and Ali Sendur. Finite difference approximations of multidimensional convection-diffusion-reaction problems with small diffusion on a special grid. Journal of Computational Physics, 300(??):574–591, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005240.

Kuwata:2015:ALB

[KS15b] Y. Kuwata and K. Suga. Anomaly of the lattice Boltzmann methods in three-dimensional cylindrical flows. Journal of Computational Physics, 280(??):563–569, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006767.

Kaessmair:2016:CCA

[KS16a] S. Kaessmair and P. Steinmann. Comparative computational analysis of the Cahn–Hilliard equation with emphasis on C^{1}-continuous methods. Journal of Computational Physics, 322(??):783–803, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302844.

Konakli:2016:PMM

[KS16b] Katerina Konakli and Bruno Sudret. Polynomial meta-models with canonical low-rank approximations: Numerical insights and comparison to sparse polynomial chaos expansions. Journal of Computational Physics, 321(??):1144–1169, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302303.

Kou:2016:MSD

[KS16c] Jisheng Kou and Shuyu Sun. Multi-scale diffuse interface modeling of multi-component two-phase flow with partial miscibility. Journal of Computational Physics, 318(??):349–372, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301103.

Kuwata:2016:ICG

[KS16d] Y. Kuwata and K. Suga. Imbalance-correction grid-refinement method for lattice Boltzmann flow simulations. Journal of Computational Physics, 311(??):348–362, April 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000620.

Kuhnlein:2017:UMF

[KS17]  Christian Kühnlein and Piotr K. Smolarkiewicz.  An

unstructured-mesh finite-volume MPDATA for compressible atmospheric dynamics. Journal of Computational Physics, 334(??):16–30, April 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116307227.

Kilian:2018:SIM

[KS18a] Patrick Kilian and Felix Spanier. Simulating the injection of magnetized plasma without electromagnetic precursor wave. Journal of Computational Physics, 353(??):258–263, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307507.

Kou:2018:TCS

[KS18b] Jisheng Kou and Shuyu Sun. Thermodynamically consistent simulation of nonisothermal diffuse-interface two-phase flow with Peng–Robinson equation of state. Journal of Computational Physics, 371(??):581–605, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303565.

Krishnan:2017:FRV

[KSI17] Sreenath Krishnan, Eric S. G. Shaqfeh, and Gianluca Iaccarino. Fully resolved viscoelastic particulate simulations using unstructured grids. Journal of Computational Physics, 338(??):313–338, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301717.

Kelly:2019:BCT

[KSM19] James F. Kelly, Harish Sankaranarayanan, and Mark M. Meerschaert. Boundary conditions for two-sided fractional diffusion. Journal of Computational Physics, 376(??):1089–1107, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306673.

Kucharik:2018:MSR

[KSSL18] M. Kucharik, G. Scovazzi, M. Shashkov, and R. Loubère. A multi-scale residual-based anti-hourglass control for compatible staggered Lagrangian hydrodynamics. Journal of Computational Physics, 354(??):1–25, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308173.

Kersaudy:2015:NSM

[KSV^{+}15] Pierric Kersaudy, Bruno Sudret, Nadège Varsier, Odile Picon, and Joe Wiart. A new surrogate modeling technique combining kriging and polynomial chaos expansions — application to uncertainty analysis in computational dosimetry. Journal of Computational Physics, 286(??):103–117, April 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000388.

Kramer:2018:FCD

[KSVB18] Richard M. J. Kramer, Christopher M. Siefert, Thomas E. Voth, and Pavel B. Bochev. Formulation and computation of dynamic, interface-compatible Whitney complexes in three dimensions. Journal of Computational Physics, 359(??):45–76, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117309312.

Kraus:2016:VIR

[KTG16] Michael Kraus, Emanuele Tassi, and Daniela Grasso. Variational integrators for reduced magnetohydrodynamics. Journal of Computational Physics, 321(??):435–458, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301929.

Kaboudian:2015:GSM

[KTK15] A. Kaboudian, P. Tavallali, and B. C. Khoo. The ghost solid methods for the elastic-plastic solid-solid interface and the θcriterion. Journal of Computational Physics, 302(??):618–652, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006166.

Kuya:2018:KEE

[KTK18] Yuichi Kuya, Kosuke Totani, and Soshi Kawai. Kinetic energy and entropy preserving schemes for compressible flows by split convective forms. Journal of Computational Physics, 375(??):823–853, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999118305916. See corrigendum [KTK19].

Kuya:2019:CKE

[KTK19] Yuichi Kuya, Kosuke Totani, and Soshi Kawai. Corrigendum to “Kinetic energy and entropy preserving schemes for compressible flows by split convective forms” [j. comput. phys. 375 (2018) 823–853]. Journal of Computational Physics, 391(??):397, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302517. See [KTK18].

Kawai:2015:RAN

[KTN15] Soshi Kawai, Hiroshi Terashima, and Hideyo Negishi. A robust and accurate numerical method for transcritical turbulent flows at supercritical pressure with an arbitrary equation of state. Journal of Computational Physics, 300(??):116–135, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004921.

Kulikov:2016:UPA

[KV16] Igor Kulikov and Eduard Vorobyov. Using the PPML approach for constructing a low-dissipation, operator-splitting scheme for numerical simulations of hydrodynamic flows. Journal of Computational Physics, 317(??):318–346, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301127.

Kaushik:2019:MGM

[KVKS19] Aditya Kaushik, Anil K. Vashishth, Vijayant Kumar, and Manju Sharma. A modified graded mesh and higher order finite element approximation for singular perturbation problems. Journal of Computational Physics, 395(??):275–285, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304292.

Katz:2015:HOF

[KW15a] Aaron Katz and Dalon Work. High-order flux correction/ finite difference schemes for strand grids. Journal of Computational Physics, 282(??):360–380, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007773.

Kercher:2015:RPC

[KW15b] A. D. Kercher and R. S. Weigel. Removal of pseudoconvergence in coplanar and near-coplanar Riemann problems of ideal magnetohydrodynamics solved using finite volume schemes. Journal of Computational Physics, 283(??):23–36, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400789X.

Krank:2016:NAW

[KW16] Benjamin Krank and Wolfgang A. Wall. A new approach to wall modeling in LES of incompressible flow via function enrichment. Journal of Computational Physics, 316(??):94–116, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630033X.

Kozdon:2019:RAH

[KWHB19] Jeremy E. Kozdon, Lucas C. Wilcox, Thomas Hagstrom, and Jeffrey W. Banks. Robust approaches to handling complex geometries with Galerkin difference methods. Journal of Computational Physics, 392(??):483–510, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302773.

Kwon:2019:SPH

[Kwo19] Jihoe Kwon. Smoothed particle hydrodynamics model for simulating miscible multi-fluid flow. Journal of Computational Physics, 384(??):114–133, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118307952.

Kodjo:2019:MMT

[KYKS19] Kossi-Mensah Kodjo, Julien Yvonnet, Mustapha Karkri, and Karam Sab. Multiscale modeling of the thermomechanical behavior in heterogeneous media embedding Phase Change Materials particles. Journal of Computational Physics, 378(??):303–323, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307320.

Kwon:2015:DSL

[KYPK15] Jae-Min Kwon, Dokkyun Yi, Xiangfan Piao, and Philsu Kim. Development of semi-Lagrangian gyrokinetic code for full-f turbulence simulation in general tokamak geometry. Journal of Computational Physics, 283(??):518–540, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008341.

Kato:2015:DAM

[KYUO15] Hiroshi Kato, Akira Yoshizawa, Genta Ueno, and Shigeru Obayashi. A data assimilation methodology for reconstructing turbulent flows around aircraft. Journal of Computational Physics, 283(??):559–581, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008213.

Kotov:2016:NDC

[KYW^{+}16] D. V. Kotov, H. C. Yee, A. A. Wray, B. Sjögreen, and A. G. Kritsuk. Numerical dissipation control in high order shock-capturing schemes for LES of low speed flows. Journal of Computational Physics, 307(??):189–202, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500769X. See corrigendum

[KYW^{+}18].

Kotov:2018:CND

[KYW^{+}18] D. V. Kotov, H. C. Yee, A. A. Wray, Björn Sjögreen, and A. G. Kritsuk. Corrigendum to “Numerical dissipation control in high order shock-capturing schemes for LES of low speed flows” [J. Comput. Phys. 307 (2016) 189–202]. Journal of Computational Physics, 352(??):637, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307416. See [KYW^{+}16].

Katsoulakis:2017:SIP

[KZ17] M. A. Katsoulakis and N. Zabaras. Special issue: Predictive multiscale materials modeling. Journal of Computational Physics, 338(??):18–20, June 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301481.

Koutsourelakis:2016:SIB

[KZG16] P. S. Koutsourelakis, N. Zabaras, and M. Girolami. Special issue: Big data and predictive computational modeling. Journal of Computational Physics, 321(??):1252–1254, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001807.

King:2015:BCS

[KZR15] J. R. C. King, A. M. Ziolkowski, and M. Ruffert. Boundary conditions for simulations of oscillating bubbles using the non-linear acoustic approximation. Journal of Computational Physics, 284(??):273–290, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008547.

Lowengrub:2016:NSE

[LAA16] John Lowengrub, Jun Allard, and Sebastian Aland. Numerical simulation of endocytosis: Viscous flow driven by membranes with non-uniformly distributed curvature-inducing molecules. Journal of Computational Physics, 309(??):112–128, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008803.

Liu:2018:WBP

[LAEK18] Xin Liu, Jason Albright, Yekaterina Epshteyn, and Alexander Kurganov. Well-balanced positivity preserving centralupwind scheme with a novel wet/dry reconstruction on triangular grids for the Saint-Venant system. Journal of Computational Physics, 374(??):213–236, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304996.

Lombardini:2016:HOW

[LAK^{+}16] Richard Lombardini, Ramiro Acevedo, Alexander Kuczala, Kerry P. Keys, Carl P. Goodrich, and Bruce R. Johnson. Higher-order wavelet reconstruction/differentiation filters and Gibbs phenomena. Journal of Computational Physics, 305(??):244–262, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115007111.

Lee:2018:CCH

[LAL18] Euntaek Lee, Hyung Taek Ahn, and Hong Luo. Cellcentered high-order hyperbolic finite volume method for diffusion equation on unstructured grids. Journal of Computational Physics, 355(??):464–491, February 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308185.

Lan:2019:ADR

[Lan19] Shiwei Lan. Adaptive dimension reduction to accelerate infinite-dimensional geometric Markov Chain Monte Carlo. Journal of Computational Physics, 392(??):71–95, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911930289X.

Lappa:2016:MNF

[Lap16] Marcello Lappa. A mathematical and numerical framework for the analysis of compressible thermal convection in gases at very high temperatures. Journal of Computational Physics, 313(??):687–712, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001340.

Lapenta:2017:EEC

[Lap17] Giovanni Lapenta. Exactly energy conserving semi-implicit particle in cell formulation. Journal of Computational Physics, 334(??):349–366, April 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300128.

Lau:2017:SSL

[Lau17] Stephen R. Lau. Stellar surface as low-rank modification in iterative methods for binary neutron stars. Journal of Computational Physics, 348(??):460–481, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730534X.

Li:2015:ANT

[LB15] Shan Li and John P. Boyd. Approximation on non-tensor domains including squircles, Part III: Polynomial hyperinterpolation and radial basis function interpolation on Chebyshevlike grids and truncated uniform grids. Journal of Computational Physics, 281(??):653–668, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007177.

Liu:2016:ENS

[LB16] Yuxiang Liu and Alex H. Barnett. Efficient numerical solution of acoustic scattering from doubly-periodic arrays of axisymmetric objects. Journal of Computational Physics, 324(??):226–245, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303540.

Levien:2017:CSP

[LB17] Ethan Levien and Paul C. Bressloff. Coupling sample paths to the thermodynamic limit in Monte Carlo estimators with applications to gene expression. Journal of Computational Physics, 346(??):1–13, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304436.

Lynch:2017:AAW

[LBB^{+}17] Vickie E. Lynch, Jose M. Borreguero, Debsindhu Bhowmik, Panchapakesan Ganesh, Bobby G. Sumpter, Thomas E. Proffen, and Monojoy Goswami. An automated analysis workflow for optimization of force-field parameters using neutron scattering data. Journal of Computational Physics, 340(??):128– 137, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302401.

Lan:2016:EHO

[LBTCG16] Shiwei Lan, Tan Bui-Thanh, Mike Christie, and Mark Girolami. Emulation of higher-order tensors in manifold Monte Carlo methods for Bayesian inverse problems. Journal of Computational Physics, 308(??):81–101, March 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008517.

Li:2018:DPD

[LBTK18] Zhen Li, Xin Bian, Yu-Hang Tang, and George Em Karniadakis. A dissipative particle dynamics method for arbitrarily complex geometries. Journal of Computational Physics, 355(??):534–547, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117308525.

Li:2016:CSF

[LBZ16] Meng Li, Boris N. Breizman, and Linjin Zheng. Canonical straight field line magnetic flux coordinates for tokamaks. Journal of Computational Physics, 326(??):334–341, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304168.

Li:2016:ITG

[LBZA16] Qi Li, Elie Bou-Zeid, and William Anderson. The impact and treatment of the Gibbs phenomenon in immersed boundary method simulations of momentum and scalar transport. Journal of Computational Physics, 310(??):237–251, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000140.

Lee:2015:DFI

[LC15] Injae Lee and Haecheon Choi. A discrete-forcing immersed boundary method for the fluid-structure interaction of an elastic slender body. Journal of Computational Physics, 280(??):529–546, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006640.

Liska:2016:FLG

[LC16] Sebastian Liska and Tim Colonius. A fast lattice Green’s function method for solving viscous incompressible flows on unbounded domains. Journal of Computational Physics, 316(??):360–384, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911630064X.

Liang:2017:NMD

[LC17a] Hui Liang and Xiaobo Chen. A new multi-domain method based on an analytical control surface for linear and secondorder mean drift wave loads on floating bodies. Journal of Computational Physics, 347(??):506–532, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305168.

Liska:2017:FIB

[LC17b] Sebastian Liska and Tim Colonius. A fast immersed boundary method for external incompressible viscous flows using lattice Green’s functions. Journal of Computational Physics, 331(??):257–279, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911630626X.

Lee:2018:SAK

[LC18] Gibbeum Lee and Yeunwoo Cho. Semi-analytical Karhunen– Loeve representation of irregular waves based on the prolate spheroidal wave functions. Journal of Computational Physics, 352(??):172–189, January 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306824.

Li:2019:DGS

[LC19] Xiaoxu Li and Huajie Chen. A discontinuous Galerkin scheme for full-potential electronic structure calculations. Journal of Computational Physics, 385(??):33–50, May 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301172.

Liu:2019:KII

[LCCZ19] Yuan Liu, Yingda Cheng, Shanqin Chen, and Yong-Tao Zhang. Krylov implicit integration factor discontinuous Galerkin methods on sparse grids for high dimensional reaction–diffusion equations. Journal of Computational Physics, 388(??):90–102, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302049.

Llor:2016:EPE

[LCF16] Antoine Llor, Alexandra Claisse, and Christophe Fochesato. Energy preservation and entropy in Lagrangian space- and time-staggered hydrodynamic schemes. Journal of Computational Physics, 309(??):324–349, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008633.

Li:2016:MCC

[LCK16] Yibao Li, Jung-Il Choi, and Junseok Kim. Multi-component Cahn–Hilliard system with different boundary conditions in complex domains. Journal of Computational Physics, 323 (??):1–16, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303126.

Liu:2019:APS

[LCK19] Xin Liu, Alina Chertock, and Alexander Kurganov. An asymptotic preserving scheme for the two-dimensional shallow water equations with Coriolis forces. Journal of Computational Physics, 391(??):259–279, August 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302815.

Li:2019:FML

[LCLY19] Chengxi Li, Bryce K. Campbell, Yuming Liu, and Dick K. P. Yue. A fast multi-layer boundary element method for direct numerical simulation of sound propagation in shallow water environments. Journal of Computational Physics, 392(??):694–712, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303213.

Liu:2015:DII

[LD15] Hao-Ran Liu and Hang Ding. A diffuse-interface immersedboundary method for two-dimensional simulation of flows with moving contact lines on curved substrates. Journal of Computational Physics, 294(??):484–502, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500220X.

Lozano-Duran:2019:ESL

[LDB19] Adrián Lozano-Durán and Hyunji Jane Bae. Error scaling of large-eddy simulation in the outer region of wallbounded turbulence. Journal of Computational Physics, 392 (??):532–555, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911930316X.

Lochon:2016:HTR

[LDGH16] H. Lochon, F. Daude, P. Galon, and J.-M. Hérard. HLLCtype Riemann solver with approximated two-phase contact for the computation of the Baer–Nunziato two-fluid model. Journal of Computational Physics, 326(??):733–762, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304272.

Lozano-Duran:2015:NAC

[LDHJ15] Adrián Lozano-Durán, Markus Holzner, and Javier Jiménez. Numerically accurate computation of the conditional trajectories of the topological invariants in turbulent flows. Journal of Computational Physics, 295(??):805–814, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002934.

Li:2016:SBI

[LDL^{+}16] Jie Li, Daniel Dault, Beibei Liu, Yiying Tong, and Balasubramaniam Shanker. Subdivision based isogeometric analysis technique for electric field integral equations for simply connected structures. Journal of Computational Physics, 319(??):145–162, August 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116300407.

Lebbe:2019:RST

[LDO^{+}19] N. Lebbe, C. Dapogny, E. Oudet, K. Hassan, and A. Gliere. Robust shape and topology optimization of nanophotonic devices using the level set method. Journal of Computational Physics, 395(??):710–746, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304656.

Lee:2017:RRP

[LDOK17] Byungjoon Lee, Jérôme Darbon, Stanley Osher, and Myungjoo Kang. Revisiting the redistancing problem using the Hopf–Lax formula. Journal of Computational Physics, 330(??):268–281, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305903.

Lopes:2019:LTS

[LDSM19] Müller Moreira Lopes, Margarete Oliveira Domingues, Kai Schneider, and Odim Mendes. Local time-stepping for adaptive multiresolution using natural extension of Runge–Kutta methods. Journal of Computational Physics, 382(??):291–318, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300191.

Ling:2019:PCP

[LDT19] Dan Ling, Junming Duan, and Huazhong Tang. Physicalconstraints-preserving Lagrangian finite volume schemes for one- and two-dimensional special relativistic hydrodynamics. Journal of Computational Physics, 396(??):507–543, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304632.

Liu:2015:LBM

[LDWZ15] Haifei Liu, Yu Ding, Hongda Wang, and Jie Zhang. Lattice Boltzmann method for the age concentration equation in shallow water. Journal of Computational Physics, 299(??):613–629, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004672.

Lee:2016:MNM

[LE16] Yoonsang Lee and Bjorn Engquist. Multiscale numerical methods for passive advection-diffusion in incompressible turbulent flow fields. Journal of Computational Physics, 317(??):33–46, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301000.

Lusso:2017:TDS

[LEB^{+}17] Christelle Lusso, Alexandre Ern, François Bouchut, Anne Mangeney, Maxime Farin, and Olivier Roche. Two-dimensional simulation by regularization of free surface viscoplastic flows with Drucker–Prager yield stress and application to granular collapse. Journal of Computational Physics, 333(??):387–408, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306982.

Lerat:2015:HOT

[Ler15] A. Lerat. A high-order time formulation of the RBC schemes for unsteady compressible Euler equations. Journal of Computational Physics, 303(??):251–268, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006440.

Lerat:2016:EHO

[Ler16] A. Lerat. An efficient high-order compact scheme for the unsteady compressible Euler and Navier–Stokes equations. Journal of Computational Physics, 322(??):365–386, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302753.

Liu:2019:IPD

[LFAR19] Chen Liu, Florian Frank, Faruk O. Alpak, and Béatrice Rivière. An interior penalty discontinuous Galerkin approach for 3D incompressible Navier–Stokes equation for permeability estimation of porous media. Journal of Computational Physics, 396(??):669–686, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304607.

Li:2016:IBLa

[LFDP16] Zhe Li, Julien Favier, Umberto D’Ortona, and Sébastien Poncet. An immersed boundary-lattice Boltzmann method for single- and multi-component fluid flows. Journal of Computational Physics, 304(??):424–440, January 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006907.

Lee:2017:PCM

[LFR17] Dongwook Lee, Hugues Faller, and Adam Reyes. The Piecewise Cubic Method (PCM) for computational fluid dynamics. Journal of Computational Physics, 341(??):230–257, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302759.

LeHardy:2017:SRT

[LFRH17] D. Le Hardy, Y. Favennec, B. Rousseau, and F. Hecht. Specular reflection treatment for the 3D radiative transfer equation solved with the discrete ordinates method. Journal of Computational Physics, 334(??):541–572, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300293.

Lu:2016:ILW

[LFT^{+}16] Jianfang Lu, Jinwei Fang, Sirui Tan, Chi-Wang Shu, and Mengping Zhang. Inverse Lax–Wendroff procedure for numerical boundary conditions of convection-diffusion equations. Journal of Computational Physics, 317(??):276–300, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301140.

Lieu:2016:CHO

[LGB16] Alice Lieu, Gwénaël Gabard, and Hadrien Bériot. A comparison of high-order polynomial and wave-based methods for Helmholtz problems. Journal of Computational Physics, 321(??):105–125, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116301905.

Lemoine:2017:MFA

[LGB17] Antoine Lemoine, Stéphane Glockner, and Jérôme Breil. Moment-of-fluid analytic reconstruction on 2D Cartesian grids. Journal of Computational Physics, 328(??):131–139, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305113.

Liu:2017:FSI

[LGD17] Hao-Ran Liu, Peng Gao, and Hang Ding. Fluid-structure interaction involving dynamic wetting: 2D modeling and simulations. Journal of Computational Physics, 348(??):45–65, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305193.

Li:2018:FLC

[LGH^{+}18] Meng Li, Xian-Ming Gu, Chengming Huang, Mingfa Fei, and Guoyu Zhang. A fast linearized conservative finite element method for the strongly coupled nonlinear fractional Schrödinger equations. Journal of Computational Physics, 358(??):256–282, April 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300044.

Lai:2017:RIF

[LGO17] Jun Lai, Leslie Greengard, and Michael O’Neil. Robust integral formulations for electromagnetic scattering from three-dimensional cavities. Journal of Computational Physics, 345(??):1–16, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303777.

Liu:2019:IGC

[LGZ^{+}19] Geng Liu, Biao Geng, Xudong Zheng, Qian Xue, Haibo Dong, and George V. Lauder. An image-guided computational approach to inversely determine in vivo material properties and model flow-structure interactions of fish fins. Journal of Computational Physics, 392(??):578–593, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303158.

Lemmer:2015:PDD

[LH15] Andreas Lemmer and Rudolf Hilfer. Parallel domain decomposition method with non-blocking communication for flow through porous media. Journal of Computational Physics, 281(??):970–981, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114006007.

LaGrone:2016:DAB

[LH16] John LaGrone and Thomas Hagstrom. Double absorbing boundaries for finite-difference time-domain electromagnetics. Journal of Computational Physics, 326(??):650–665, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304260.

Lang:2017:EBI

[LH17a] Jens Lang and Willem Hundsdorfer. Extrapolation-based implicit–explicit peer methods with optimised stability regions. Journal of Computational Physics, 337(??):203–215, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301274.

Liu:2017:ATG

[LH17b] Cheng Liu and Changhong Hu. Adaptive THINC–GFM for compressible multi-medium flows. Journal of Computational Physics, 342(??):43–65, August 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303030.

Liu:2018:AMM

[LH18] Cheng Liu and Changhong Hu. An adaptive multi-moment FVM approach for incompressible flows. Journal of Computational Physics, 359(??):239–262, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300160.

Litvinov:2015:TCC

[LHA15a] S. Litvinov, X. Y. Hu, and N. A. Adams. Towards consistence and convergence of conservative SPH approximations. Journal of Computational Physics, 301(??):394–401, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005690.

Luo:2015:CSI

[LHA15b] J. Luo, X. Y. Hu, and N. A. Adams. A conservative sharp interface method for incompressible multiphase flows. Journal of Computational Physics, 284(??):547–565, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008675.

Luo:2016:CBC

[LHA16a] J. Luo, X. Y. Hu, and N. A. Adams. Curvature boundary condition for a moving contact line. Journal of Computational Physics, 310(??):329–341, April 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000322.

Luo:2016:EFS

[LHA16b] J. Luo, X. Y. Hu, and N. A. Adams. Efficient formulation of scale separation for multi-scale modeling of interfacial flows. Journal of Computational Physics, 308(??):411–420, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007846.

Li:2016:SPF

[LHB^{+}16] L. Li, W. D. Henshaw, J. W. Banks, D. W. Schwendeman, and A. Main. A stable partitioned FSI algorithm for incompressible flow and deforming beams. Journal of Computational Physics, 312(??):272–306, May 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000565.

Lopez:2016:NVC

[LHGF16] J. López, J. Hernández, P. Gómez, and F. Faura. A new volume conservation enforcement method for PLIC reconstruction in general convex grids. Journal of Computational Physics, 316(??):338–359, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630050X.

Lopez:2019:NCA

[LHGF19] Joaquín López, Julio Hernández, Pablo Gómez, and Félix Faura. Non-convex analytical and geometrical tools for volume truncation, initialization and conservation enforcement in VOF methods. Journal of Computational Physics, 392 (??):666–693, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303080.

Lu:2015:MPD

[LHL15] Wenying Lu, Yunqing Huang, and Hailiang Liu. Mass preserving discontinuous Galerkin methods for Schrödinger equations. Journal of Computational Physics, 282(??):210–226, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007669.

Li:2019:PBM

[LHLL19] Kun Li, Ting-Zhu Huang, Liang Li, and Stéphane Lanteri. POD-based model order reduction with an adaptive snapshot selection for a discontinuous Galerkin approximation of the time-domain Maxwell’s equations. Journal of Computational Physics, 396(??):106–128, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304814.

Li:2016:OFO

[LHMB16] Y. Li, B. Han, L. Métivier, and R. Brossier. Optimal fourthorder staggered-grid finite-difference scheme for 3D frequencydomain viscoelastic wave modeling. Journal of Computational Physics, 321(??):1055–1078, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302431.

Li:2018:RCA

[LHMB18] Zhiyong Li, Jesse B. Hoagg, Alexandre Martin, and Sean C. C. Bailey. Retrospective cost adaptive Reynolds-averaged Navier– Stokes k-ω model for data-driven unsteady turbulent simulations. Journal of Computational Physics, 357(??):353–374, March 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308756.

Lehmkuhl:2019:LDF

[LHO^{+}19] O. Lehmkuhl, G. Houzeaux, H. Owen, G. Chrysokentis, and I. Rodriguez. A low-dissipation finite element scheme for scale resolving simulations of turbulent flows. Journal of Computational Physics, 390(??):51–65, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302372.

Luo:2016:HLH

[LHQ16] Dongmi Luo, Weizhang Huang, and Jianxian Qiu. A hybrid LDG–HWENO scheme for KdV-type equations. Journal of Computational Physics, 313(??):754–774, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001364.

Luo:2019:QLM

[LHQ19] Dongmi Luo, Weizhang Huang, and Jianxian Qiu. A quasi-Lagrangian moving mesh discontinuous Galerkin method for hyperbolic conservation laws. Journal of Computational Physics, 396(??):544–578, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304693.

Lee:2018:SEP

[LHS^{+}18] Dongjin Lee, Takeo Hoshi, Tomohiro Sogabe, Yuto Miyatake, and Shao-Liang Zhang. Solution of the k-th eigenvalue problem in large-scale electronic structure calculations. Journal of Computational Physics, 371(??):618–632, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303796.

Li:2019:SVM

[LHS^{+}19] Yingzhe Li, Yang He, Yajuan Sun, Jitse Niesen, Hong Qin, and Jian Liu. Solving the Vlasov–Maxwell equations using Hamiltonian splitting. Journal of Computational Physics, 396(??):381–399, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304784.

Long:2017:ABG

[LHW^{+}17] Ting Long, Dean Hu, Detao Wan, Chen Zhuang, and Gang Yang. An arbitrary boundary with ghost particles incorporated in coupled FEM–SPH model for FSI problems. Journal of Computational Physics, 350(??):166–183, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306253.

Liu:2017:TEA

[LHY17] Wen Liu, ChengDe Huang, and Guowei Yang. Time efficient aeroelastic simulations based on radial basis functions. Journal of Computational Physics, 330(??):810–827, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630571X.

Lin:2019:FSI

[LHY^{+}19] Zhaowu Lin, Andrew Hess, Zhaosheng Yu, Shengqiang Cai, and Tong Gao. A fluid-structure interaction study of soft robotic swimmer using a fictitious domain/active-strain method. Journal of Computational Physics, 376(??):1138– 1155, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118306727.

Lv:2015:EBD

[LI15] Yu Lv and Matthias Ihme. Entropy-bounded discontinuous Galerkin scheme for Euler equations. Journal of Computational Physics, 295(??):715–739, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500282X.

Li:2017:HOS

[Li17] Weidong Li. High order spectral difference lattice Boltzmann method for incompressible hydrodynamics. Journal of Computational Physics, 345(??):618–636, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304229.

Liao:2016:ESA

[Lia16] Haitao Liao. Efficient sensitivity analysis method for chaotic dynamical systems. Journal of Computational Physics, 313(??):57–75, May 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911600070X.

Liu:2016:SOC

[Liu16] Jie Liu. A second-order changing-connectivity ALE scheme and its application to FSI with large convection of fluids and near contact of structures. Journal of Computational Physics, 304(??):380–423, January 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006798.

Liu:2019:NFI

[Liu19a] Keji Liu. Near-field imaging of inhomogeneities in a stratified ocean waveguide. Journal of Computational Physics, 398(??):Article 108901, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305996.

Liu:2019:SAP

[Liu19b] Liu Liu. A stochastic asymptotic-preserving scheme for the bipolar semiconductor Boltzmann–Poisson system with random inputs and diffusive scalings. Journal of Computational Physics, 376(??):634–659, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306582.

Liu:2018:SAH

[LIW18] Changying Liu, Arieh Iserles, and Xinyuan Wu. Symmetric and arbitrarily high-order Birkhoff–Hermite time integrators and their long-time behaviour for solving nonlinear Klein–Gordon equations. Journal of Computational Physics, 356(??):1–30, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308240.

Liu:2016:WEW

[LJ16] Hongxu Liu and Xiangmin Jiao. WLS–ENO: Weightedleast-squares based essentially non-oscillatory schemes for finite volume methods on unstructured meshes. Journal of Computational Physics, 314(??):749–773, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001911.

Li:2019:MVE

[LJ19] Qiuqi Li and Lijian Jiang. A multiscale virtual element method for elliptic problems in heterogeneous porous media. Journal of Computational Physics, 388(??):394–415, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302141.

Ling:2016:MLS

[LJT16] Julia Ling, Reese Jones, and Jeremy Templeton. Machine learning strategies for systems with invariance properties. Journal of Computational Physics, 318(??):22–35, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301309.

Li:2015:PTM

[LJZ15] Xiao Li, Guanghua Ji, and Hui Zhang. Phase transitions of macromolecular microsphere composite hydrogels based on the stochastic Cahn–Hilliard equation. Journal of Computational Physics, 283(??):81–97, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007943.

Liu:2016:MDT

[LK16a] Chein-Shan Liu and Chung-Lun Kuo. A multiple-direction Trefftz method for solving the multi-dimensional wave equation in an arbitrary spatial domain. Journal of Computational Physics, 321(??):39–54, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301759.

Lohmann:2016:SFL

[LK16b] Christoph Lohmann and Dmitri Kuzmin. Synchronized flux limiting for gas dynamics variables. Journal of Computational Physics, 326(??):973–990, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304399.

Latypov:2017:DDR

[LK17] Marat I. Latypov and Surya R. Kalidindi. Data-driven reduced order models for effective yield strength and partitioning of strain in multiphase materials. Journal of Computational Physics, 346(??):242–261, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304588.

Lai:2015:FRS

[LKB15] Jun Lai, Motoki Kobayashi, and Alex Barnett. A fast and robust solver for the scattering from a layered periodic structure containing multi-particle inclusions. Journal of Computational Physics, 298(??):194–208, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003848.

Lee:2017:GCF

[LKK17a] Seungjoon Lee, Ioannis G. Kevrekidis, and George Em Karniadakis. A general CFD framework for fault-resilient simulations based on multi-resolution information fusion. Journal of Computational Physics, 347(??):290–304, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304989.

Lee:2017:REC

[LKK17b] Seungjoon Lee, Ioannis G. Kevrekidis, and George Em Karniadakis. A resilient and efficient CFD framework: Statistical learning tools for multi-fidelity and heterogeneous information fusion. Journal of Computational Physics, 344 (??):516–533, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303960.

Linders:2017:SPO

[LKN17] Viktor Linders, Marco Kupiainen, and Jan Nordström. Summation-by-Parts operators with minimal dispersion error for coarse grid flow calculations. Journal of Computational Physics, 340(??):160–176, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302346.

Lohmann:2017:FCTa

[LKSM17] Christoph Lohmann, Dmitri Kuzmin, John N. Shadid, and Sibusiso Mabuza. Flux-corrected transport algorithms for continuous Galerkin methods based on high order Bernstein finite elements. Journal of Computational Physics, 344(??):151–186, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117303388.

Li:2015:AIS

[LL15] Weixuan Li and Guang Lin. An adaptive importance sampling algorithm for Bayesian inversion with multimodal distributions. Journal of Computational Physics, 294(??):173–190, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002028.

Lai:2016:LDM

[LL16a] Rongjie Lai and Jianfeng Lu. Localized density matrix minimization and linear-scaling algorithms. Journal of Computational Physics, 315(??):194–210, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300146.

Li:2016:IBLb

[LL16b] Weidong Li and Li-Shi Luo. An implicit block LU-SGS finite-volume lattice-Boltzmann scheme for steady flows on arbitrary unstructured meshes. Journal of Computational Physics, 327(??):503–518, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304594.

Liao:2016:RBA

[LL16c] Qifeng Liao and Guang Lin. Reduced basis ANOVA methods for partial differential equations with high-dimensional random inputs. Journal of Computational Physics, 317(??):148–164, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300754.

Bris:2017:ECA

[LL17] Claude Le Bris and Frédéric Legoll. Examples of computational approaches for elliptic, possibly multiscale PDEs with random inputs. Journal of Computational Physics, 328(??):455–473, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305253.

Lester:2018:FPG

[LL18] Daniel Lester and Ruru Li. The frictional pebble game: An algorithm for rigidity percolation in saturated frictional assemblies. Journal of Computational Physics, 369(??):225– 236, September 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118303085.

Law:2019:NRF

[LL19a] Y.-M. Law and M. Laforest. A nonlinear relaxation formulation of the p-curl problem modelling high-temperature superconductors: a modified Yee’s scheme. Journal of Computational Physics, 378(??):591–614, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830771X.

Liao:2019:ARB

[LL19b] Qifeng Liao and Jinglai Li. An adaptive reduced basis ANOVA method for high-dimensional Bayesian inverse problems. Journal of Computational Physics, 396(??):364–380, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S002199911930467X.

Liu:2019:EST

[LLA19] Kai Liu, John Lowengrub, and Jun Allard. Efficient simulation of thermally fluctuating biopolymers immersed in fluids on 1-micron, 1-second scales. Journal of Computational Physics, 386(??):248–263, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300889.

Liu:2019:SIT

[LLB19] Kai Liu, Mandar Lakhote, and S. Balachandar. Self-induced temperature correction for inter-phase heat transfer in Euler– Lagrange point-particle simulation. Journal of Computational Physics, 396(??):596–615, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304772.

Laguna:2016:FIF

[LLD^{+}16] A. Alvarez Laguna, A. Lani, H. Deconinck, N. N. Mansour, and S. Poedts. A fully-implicit finite-volume method for multi-fluid reactive and collisional magnetized plasmas on unstructured meshes. Journal of Computational Physics, 318(??):252–276, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301139.

Long:2019:PNL

[LLD19] Zichao Long, Yiping Lu, and Bin Dong. PDE-Net 2.0: Learning PDEs from data with a numeric-symbolic hybrid deep network. Journal of Computational Physics, 399 (??):Article 108925, December 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306308.

Lee:2017:APM

[LLEK17] M. Lee, K. Leiter, C. Eisner, and J. Knap. Atom-partitioned multipole expansions for electrostatic potential boundary conditions. Journal of Computational Physics, 328(??):344–353, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305101.

Li:2018:UGK

[LLFX18] Shiyi Li, Qibing Li, Song Fu, and Kun Xu. A unified gaskinetic scheme for axisymmetric flow in all Knudsen number regimes. Journal of Computational Physics, 365(??):144–169, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302183.

Liu:2019:RBD

[LLH19] Jianming Liu, Xinkai Li, and Xiuling Hu. A RBF-based differential quadrature method for solving two-dimensional variable-order time fractional advection–diffusion equation. Journal of Computational Physics, 384(??):222–238, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300920.

Li:2018:CRM

[LLJJ18] Yulong Li, Yun Zhi Law, Vaibhav Joshi, and Rajeev K. Jaiman. A 3D common-refinement method for non-matching meshes in partitioned variational fluid-structure analysis. Journal of Computational Physics, 374(??):163–187, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303231.

Li:2016:IRB

[LLL16] Weixuan Li, Guang Lin, and Bing Li. Inverse regression-based uncertainty quantification algorithms for high-dimensional models: Theory and practice. Journal of Computational Physics, 321(??):259–278, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116301851.

Lou:2018:RDG

[LLLN18] Jialin Lou, Lingquan Li, Hong Luo, and Hiroaki Nishikawa. Reconstructed discontinuous Galerkin methods for linear advection-diffusion equations based on first-order hyperbolic system. Journal of Computational Physics, 369(??):103– 124, September 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118302912.

Lin:2016:FAA

[lLLNS16] Xue lei Lin, Xin Lu, Micheal K. Ng, and Hai-Wei Sun. A fast accurate approximation method with multigrid solver for two-dimensional fractional sub-diffusion equation. Journal of Computational Physics, 323(??):204–218, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303266.

Legrand:2017:MGF

[LLM17] Nicolas Legrand, Ghislain Lartigue, and Vincent Moureau. A multi-grid framework for the extraction of large-scale vortices in Large-Eddy Simulation. Journal of Computational Physics, 349(??):528–560, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306010.

Lin:2017:MML

[lLNS17] Xue lei Lin, Michael K. Ng, and Hai-Wei Sun. A multigrid method for linear systems arising from time-dependent two-dimensional space-fractional diffusion equations. Journal of Computational Physics, 336(??):69–86, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301018.

Lee:2016:HOC

[LLP^{+}16] D. Lee, R. Lowrie, M. Petersen, T. Ringler, and M. Hecht. A high order characteristic discontinuous Galerkin scheme for advection on unstructured meshes. Journal of Computational Physics, 324(??):289–302, November 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303539.

Li:2015:DAD

[LLS15] Qin Li, Jianfeng Lu, and Weiran Sun. Diffusion approximations and domain decomposition method of linear transport equations: Asymptotics and numerics. Journal of Computational Physics, 292(??):141–167, July 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001503.

Li:2019:CMB

[LLSJ19] Weidong Li, Wei Li, Pai Song, and Hao Ji. A conservationmoment-based implicit finite volume lattice Boltzmann method for steady nearly incompressible flows. Journal of Computational Physics, 398(??):Article 108882, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119305807.

Liu:2015:NII

[LLVF^{+}15] Fang Liu, Lin Lin, Derek Vigil-Fowler, Johannes Lischner, Alexander F. Kemper, Sahar Sharifzadeh, Felipe H. da Jornada, Jack Deslippe, Chao Yang, Jeffrey B. Neaton, and Steven G. Louie. Numerical integration for ab initio manyelectron self energy calculations within the GW approximation. Journal of Computational Physics, 286(??):1–13, April 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000273.

Lin:2019:PBF

[LLW19] Yixin Lin, G. R. Liu, and Guangyu Wang. A particle-based free surface detection method and its application to the surface tension effects simulation in smoothed particle hydrodynamics (SPH). Journal of Computational Physics, 383(??):196–206, April 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300385.

Li:2018:ETI

[LLWJ18] Shu-Jie Li, Li-Shi Luo, Z. J. Wang, and Lili Ju. An exponential time-integrator scheme for steady and unsteady inviscid flows. Journal of Computational Physics, 365(??):206–225, July 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301803.

Li:2015:FGA

[LLY15] Jinglai Li, Guang Lin, and Xu Yang. A frozen Gaussian approximation-based multi-level particle swarm optimization for seismic inversion. Journal of Computational Physics, 296(??):58–71, September 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003071.

Liang:2018:PBG

[LLY18] Tengfei Liang, Qi Li, and Wenjing Ye. A physical-based gas-surface interaction model for rarefied gas flow simulation. Journal of Computational Physics, 352(??):105–122, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306423.

Lannes:2015:NCF

[LM15a] D. Lannes and F. Marche. A new class of fully nonlinear and weakly dispersive Green–Naghdi models for efficient 2-D simulations. Journal of Computational Physics, 282(??):238–268, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007748.

LeFloch:2015:RMC

[LM15b] Philippe G. LeFloch and Jean-Marc Mercier. Revisiting the method of characteristics via a convex hull algorithm. Journal of Computational Physics, 298(??):95–112, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003770.

Lotstedt:2015:SSD

[LM15c] Per Lötstedt and Lina Meinecke. Simulation of stochastic diffusion via first exit times. Journal of Computational Physics, 300(??):862–886, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005276.

Lu:2015:NSS

[LM15d] Jianfeng Lu and Christian B. Mendl. Numerical scheme for a spatially inhomogeneous matrix-valued quantum Boltzmann equation. Journal of Computational Physics, 291(??):303–316, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001679.

Linke:2016:VED

[LM16] A. Linke and C. Merdon. On velocity errors due to irrotational forces in the Navier–Stokes momentum balance. Journal of Computational Physics, 313(??):654–661, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911600142X.

Lee:2018:ESM

[LM18] Byungjoon Lee and Chohong Min. An energy-stable method for solving the incompressible Navier–Stokes equations with non-slip boundary condition. Journal of Computational Physics, 360(??):104–119, May 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300408.

Lipnikov:2019:HOC

[LM19a] Konstantin Lipnikov and Nathaniel Morgan. A high-order conservative remap for discontinuous Galerkin schemes on curvilinear polygonal meshes. Journal of Computational Physics, 399(??):Article 108931, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306369.

Lipnikov:2019:HOD

[LM19b] Konstantin Lipnikov and Nathaniel Morgan. A high-order discontinuous Galerkin method for level set problems on polygonal meshes. Journal of Computational Physics, 397(??):Article 108834, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305182.

Liu:2019:REI

[LM19c] Ju Liu and Alison L. Marsden. A robust and efficient iterative method for hyper-elastodynamics with nested block preconditioning. Journal of Computational Physics, 383(??):72–93, April 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300440.

Lo:2019:HSM

[LM19d] Wing-Cheong Lo and Shaokun Mao. A hybrid stochastic method with adaptive time step control for reaction-diffusion systems. Journal of Computational Physics, 379(??):392–402, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307873.

Liu:2018:LDG

[LMB18] Xiaodong Liu, Nathaniel R. Morgan, and Donald E. Burton. Lagrangian discontinuous Galerkin hydrodynamic methods in axisymmetric coordinates. Journal of Computational Physics, 373(??):253–283, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304522.

Liu:2019:HOL

[LMB19] Xiaodong Liu, Nathaniel R. Morgan, and Donald E. Burton. A high-order Lagrangian discontinuous Galerkin hydrodynamic method for quadratic cells using a subcell mesh stabilization scheme. Journal of Computational Physics, 386(??):110–157, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301196.

Liang:2015:NSD

[LMBZ15] Xie Liang, Xu Min, Zhang Bin, and Qiu Zihua. A new spectral difference method using hierarchical polynomial bases for hyperbolic conservation laws. Journal of Computational Physics, 284(??):434–461, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007608.

Li:2016:SOT

[LMC16] Zhi-Hui Li, Qiang Ma, and Junzhi Cui. Second-order twoscale finite element algorithm for dynamic thermo-mechanical coupling problem in symmetric structure. Journal of Computational Physics, 314(??):712–748, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001868.

Luquet:2019:LRN

[LMC19] David Luquet, Régis Marchiano, and François Coulouvrat. Long range numerical simulation of acoustical shock waves in a 3D moving heterogeneous and absorbing medium. Journal of Computational Physics, 379(??):237–261, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307861.

LeTouze:2015:MMM

[LMG15] C. Le Touze, A. Murrone, and H. Guillard. Multislope MUSCL method for general unstructured meshes. Journal of Computational Physics, 284(??):389–418, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008493.

Lou:2019:VEF

[LMG19] Jijie Lou, Jim E. Morel, and N. A. Gentile. A Variable Eddington Factor method for the 1-D grey radiative transfer equations with discontinuous Galerkin and mixed finite-element spatial differencing. Journal of Computational Physics, 393(??):258–277, September 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303377.

Luu:2017:NMR

[LMGG17] Thi Hieu Luu, Yvon Maday, Matthieu Guillo, and Pierre Guérin. A new method for reconstruction of cross-sections using Tucker decomposition. Journal of Computational Physics, 345(??):189–206, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117303947.

Laboure:2016:IFH

[LMH16] Vincent M. Laboure, Ryan G. McClarren, and Cory D. Hauck. Implicit filtered P_{N} for high-energy density thermal radiation transport using discontinuous Galerkin finite elements. Journal of Computational Physics, 321(??):624–643, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301917.

Lubasch:2018:MR

[LMJ18] Michael Lubasch, Pierre Moinier, and Dieter Jaksch. Multigrid renormalization. Journal of Computational Physics, 372 (??):587–602, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118304431.

Liu:2015:WBC

[LMKS15] Xin Liu, Abdolmajid Mohammadian, Alexander Kurganov, and Julio Angel Infante Sedano. Well-balanced centralupwind scheme for a fully coupled shallow water system modeling flows over erodible bed. Journal of Computational Physics, 300(??):202–218, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500488X.

Leclercq:2016:MPH

[LML^{+}16] L. Leclercq, R. Modolo, F. Leblanc, S. Hess, and M. Mancini. 3D magnetospheric parallel hybrid multi-grid method applied to planet-plasma interactions. Journal of Computational Physics, 309(??):295–313, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000061.

Lombard:2017:NMA

[LMM17] Bruno Lombard, Agnès Maurel, and Jean-Jacques Marigo. Numerical modeling of the acoustic wave propagation across a homogenized rigid microstructure in the time domain. Journal of Computational Physics, 335(??):558–577, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300529.

Lipnikov:2016:MFD

[LMMS16] Konstantin Lipnikov, Gianmarco Manzini, J. David Moulton, and Mikhail Shashkov. The mimetic finite difference method for elliptic and parabolic problems with a staggered discretization of diffusion coefficient. Journal of Computational Physics, 305(??):111–126, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500707X.

Lundquist:2018:HFC

[LMN18] Tomas Lundquist, Arnaud Malan, and Jan Nordström. A hybrid framework for coupling arbitrary summation-by-parts schemes on general meshes. Journal of Computational Physics, 362(??):49–68, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301001.

LaRocca:2015:MDB

[LMPS15] Michele La Rocca, Andrea Montessori, Pietro Prestininzi, and Sauro Succi. A multispeed discrete Boltzmann model for transcritical 2D shallow water flows. Journal of Computational Physics, 284(??):117–132, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008468.

Lafitte:2017:HOR

[LMS17] Pauline Lafitte, Ward Melis, and Giovanni Samaey. A highorder relaxation method with projective integration for solving nonlinear systems of hyperbolic conservation laws. Journal of Computational Physics, 340(??):1–25, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730222X.

Liu:2017:TDS

[LMSK17] Xin Liu, Abdolmajid Mohammadian, Julio Ángel Infante Sedano, and Alexander Kurganov. Three-dimensional shallow water system: a relaxation approach. Journal of Computational Physics, 333(??):160–179, March 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306842.

Lu:2015:LPC

[LMTC15] Fei Lu, Matthias Morzfeld, Xuemin Tu, and Alexandre J. Chorin. Limitations of polynomial chaos expansions in the Bayesian solution of inverse problems. Journal of Computational Physics, 282(??):138–147, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007621.

Lindberg:2019:DPM

[LMY^{+}19] Casper S. Lindberg, Manoel Y. Manuputty, Edward K. Y. Yapp, Jethro Akroyd, Rong Xu, and Markus Kraft. A detailed particle model for polydisperse aggregate particles. Journal of Computational Physics, 397(??):Article 108799, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119304838.

Liu:2019:APR

[LMZ19] Xiaoxing Liu, Koji Morita, and Shuai Zhang. An ALE pairwise-relaxing meshless method for compressible flows. Journal of Computational Physics, 387(??):1–13, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301548.

Linders:2015:UBW

[LN15] Viktor Linders and Jan Nordström. Uniformly best wavenumber approximations by spatial central difference operators. Journal of Computational Physics, 300(??):695–709, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005227.

Laurent:2017:RSO

[LN17] F. Laurent and T. T. Nguyen. Realizable second-order finite-volume schemes for the advection of moment sets of the particle size distribution. Journal of Computational Physics, 337(??):309–338, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301493.

Lovett:2015:AMR

[LNM15] Sean Lovett, Nikolaos Nikiforakis, and Franck Monmont. Adaptive mesh refinement for compressible thermal flow in porous media. Journal of Computational Physics, 280(??): 21–36, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006536.

Litsarev:2016:LRA

[LO16] Mikhail S. Litsarev and Ivan V. Oseledets. A low-rank approach to the computation of path integrals. Journal of Computational Physics, 305(??):557–574, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007494.

Lai:2019:FAD

[LO19] Jun Lai and Michael O’Neil. An FFT-accelerated direct solver for electromagnetic scattering from penetrable axisymmetric objects. Journal of Computational Physics, 390(??):152–174, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302384.

Lohmann:2017:FCTb

[Loh17] Christoph Lohmann. Flux-corrected transport algorithms preserving the eigenvalue range of symmetric tensor quantities. Journal of Computational Physics, 350(??):907–926, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730668X.

Lorin:2019:SDE

[Lor19] E. Lorin. From structured data to evolution linear partial differential equations. Journal of Computational Physics, 393(??):162–185, September 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302955.

Lotfi:2018:CER

[Lot18] Ali Lotfi. Combination of epsilon and Ritz methods with multiscaling basis for solving a class of fractional optimal control problems. Journal of Computational Physics, 365(??):107–119, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302110.

Lozano:2017:MSB

[Loz17] Carlos Lozano. On mesh sensitivities and boundary formulas for discrete adjoint-based gradients in inviscid aerodynamic shape optimization. Journal of Computational Physics, 346(??):403–436, October 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304709.

Lejay:2016:SDP

[LP16a] Antoine Lejay and Géraldine Pichot. Simulating diffusion processes in discontinuous media: Benchmark tests. Journal of Computational Physics, 314(??):384–413, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001534.

Liu:2016:AED

[LP16b] Hailiang Liu and Michael Pollack. Alternating evolution discontinuous Galerkin methods for convection-diffusion equations. Journal of Computational Physics, 307(??):574–592, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008360.

Luo:2017:AMB

[LP17a] Songting Luo and Nicholas Payne. An asymptotic method based on a Hopf–Cole transformation for a kinetic BGK equation in the hyperbolic limit. Journal of Computational Physics, 341(??):295–312, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302991.

Luo:2017:PPH

[LP17b] Songting Luo and Nicholas Payne. Properties-preserving high order numerical methods for a kinetic eikonal equation. Journal of Computational Physics, 331(??):73–89, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306325.

Lee:2018:MMS

[LP18] D. Lee and A. Palha. A mixed mimetic spectral element model of the rotating shallow water equations on the cubed sphere. Journal of Computational Physics, 375(??):240–262, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305734.

Lenarda:2017:PCA

[LPB17] Pietro Lenarda, Marco Paggi, and Ricardo Ruiz Baier. Partitioned coupling of advection–diffusion–reaction systems and Brinkman flows. Journal of Computational Physics, 344(??):281–302, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117303807.

Lockerby:2015:ACH

[LPBR15] Duncan A. Lockerby, Alexander Patronis, Matthew K. Borg, and Jason M. Reese. Asynchronous coupling of hybrid models for efficient simulation of multiscale systems. Journal of Computational Physics, 284(??):261–272, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008523.

Lee:2018:DCP

[LPG18] D. Lee, A. Palha, and M. Gerritsma. Discrete conservation properties for shallow water flows using mixed mimetic spectral elements. Journal of Computational Physics, 357(??):282–304, March 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117309166.

Lepilliez:2016:TPF

[LPGT16] Mathieu Lepilliez, Elena Roxana Popescu, Frederic Gibou, and Sébastien Tanguy. On two-phase flow solvers in irregular domains with contact line. Journal of Computational Physics, 321(??):1217–1251, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302388.

Li:2018:DGS

[LPR18] Wanai Li, Jianhua Pan, and Yu-Xin Ren. The discontinuous Galerkin spectral element methods for compressible flows on two-dimensional mixed grids. Journal of Computational Physics, 364(??):314–346, July 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301451.

Luan:2019:FDE

[LPR19] Vu Thai Luan, Janusz A. Pudykiewicz, and Daniel R. Reynolds. Further development of efficient and accurate time integration schemes for meteorological models. Journal of Computational Physics, 376(??):817–837, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306818.

Latz:2018:MSM

[LPU18] Jonas Latz, Iason Papaioannou, and Elisabeth Ullmann. Multilevel Sequential 2 Monte Carlo for Bayesian inverse problems. Journal of Computational Physics, 368(??):154–178, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302286.

Leguebe:2015:SOC

[LPW15] M. Leguèbe, C. Poignard, and L. Weynans. A second-order Cartesian method for the simulation of electropermeabilization cell models. Journal of Computational Physics, 292(??):114– 140, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001758.

Lee:2015:SWP

[LPWK15] Kok Foong Lee, Robert I. A. Patterson, Wolfgang Wagner, and Markus Kraft. Stochastic weighted particle methods for population balance equations with coagulation, fragmentation and spatial inhomogeneity. Journal of Computational Physics, 303(??):1–18, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006245.

Lu:2016:BEF

[LQB16] Wangtao Lu, Jianliang Qian, and Robert Burridge. Babich’s expansion and the fast Huygens sweeping method for the Helmholtz wave equation at high frequencies. Journal of Computational Physics, 313(??):478–510, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001200.

LeLouer:2019:DMI

[LR19a] F. Le Louër and M.-L. Rapún. Detection of multiple impedance obstacles by non-iterative topological gradient based methods. Journal of Computational Physics, 388(??):534–560, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302062.

Linke:2019:PIL

[LR19b] Alexander Linke and Leo G. Rebholz. Pressure-induced locking in mixed methods for time-dependent (Navier–)Stokes equations. Journal of Computational Physics, 388(??):350–356, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301937.

Lee:2017:DNS

[LRA17] Moon Soo Lee, Amir Riaz, and Vikrant Aute. Direct numerical simulation of incompressible multiphase flow with phase change. Journal of Computational Physics, 344(??): 381–418, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303613.

Luo:2018:MMM

[LRGO18] P. Luo, C. Rodrigo, F. J. Gaspar, and C. W. Oosterlee. Monolithic multigrid method for the coupled Stokes flow and deformable porous medium system. Journal of Computational Physics, 353(??):148–168, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307350.

Lu:2017:CAS

[LRZ17] Libin Lu, Abtin Rahimian, and Denis Zorin. Contactaware simulations of particulate Stokesian suspensions. Journal of Computational Physics, 347(??):160–182, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730493X.

Ladiges:2015:FDM

[LS15a] Daniel R. Ladiges and John E. Sader. Frequency-domain Monte Carlo method for linear oscillatory gas flows. Journal of Computational Physics, 284(??):351–366, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008535.

Leimkuhler:2015:NTD

[LS15b] Benedict Leimkuhler and Xiaocheng Shang. On the numerical treatment of dissipative particle dynamics and related systems. Journal of Computational Physics, 280(??):72–95, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006445.

Li:2015:MRU

[LS15c] Jing Li and Panos Stinis. Mesh refinement for uncertainty quantification through model reduction. Journal of Computational Physics, 280(??):164–183, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006573.

Leimkuhler:2016:PAT

[LS16a] Benedict Leimkuhler and Xiaocheng Shang. Pairwise adaptive thermostats for improved accuracy and stability in dissipative particle dynamics. Journal of Computational Physics, 324(??):174–193, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303291.

Li:2016:UFM

[LS16b] Jing Li and Panos Stinis. A unified framework for mesh refinement in random and physical space. Journal of Computational Physics, 323(??):243–264, October 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303229.

Lind:2016:HOE

[LS16c] S. J. Lind and P. K. Stansby. High-order Eulerian incompressible smoothed particle hydrodynamics with transition to Lagrangian free-surface motion. Journal of Computational Physics, 326(??):290–311, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304041.

Li:2019:FAA

[LS19a] Yu Li and Richard Mikaël Slevinsky. Fast and accurate algorithms for the computation of spherically symmetric nonlocal diffusion operators on lattices. Journal of Computational Physics, 397(??):Article 108870, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305613.

Liu:2019:DRI

[LS19b] Xiaodong Liu and Jiguang Sun. Data recovery in inverse scattering: From limited-aperture to full-aperture. Journal of Computational Physics, 386(??):350–364, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307034.

Luo:2019:NEP

[LSCC19] Li Luo, Wen-Shin Shiu, Rongliang Chen, and Xiao-Chuan Cai. A nonlinear elimination preconditioned inexact Newton method for blood flow problems in human artery with stenosis. Journal of Computational Physics, 399(??):Article 108926, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S002199911930631X.

Lin:2017:SCT

[LSD^{+}17] Jian-Yu Lin, Yi Shen, Hang Ding, Nan-Sheng Liu, and Xi-Yun Lu. Simulation of compressible two-phase flows with topology change of fluid-fluid interface by a robust cut-cell method. Journal of Computational Physics, 328(??):140–159, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305216.

Liang:2018:DDC

[LSD18] Hong Liang, Qi Sun, and Qiang Du. Data-driven compressive sensing and applications in uncertainty quantification. Journal of Computational Physics, 374(??):787–802, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305199.

Lv:2016:ERS

[LSI16] Yu Lv, Yee Chee See, and Matthias Ihme. An entropyresidual shock detector for solving conservation laws using high-order discontinuous Galerkin methods. Journal of Computational Physics, 322(??):448–472, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302777.

Lee:2015:FSO

[LSL15] Hyun Geun Lee, Jaemin Shin, and June-Yub Lee. First and second order operator splitting methods for the phase field crystal equation. Journal of Computational Physics, 299(??):82–91, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004362.

Landry:2016:RMM

[LSLA16] Jonathan Landry, Azzeddine Soulaïmani, Edward Luke, and Amine Ben Haj Ali. Robust moving mesh algorithms for hybrid stretched meshes: Application to moving boundaries problems. Journal of Computational Physics, 326(??):691–721, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630420X.

Lasagna:2019:PSS

[LSM19] Davide Lasagna, Ati Sharma, and Johan Meyers. Periodic shadowing sensitivity analysis of chaotic systems. Journal of Computational Physics, 391(??):119–141, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302542.

Laadhari:2017:FIM

[LSMS17] Aymen Laadhari, Pierre Saramito, Chaouqi Misbah, and Gábor Székely. Fully implicit methodology for the dynamics of biomembranes and capillary interfaces by combining the level set and Newton methods. Journal of Computational Physics, 343(??):271–299, August 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302905.

Lindgren:2018:IEA

[LSP^{+}18] Eric B. Lindgren, Anthony J. Stace, Etienne Polack, Yvon Maday, Benjamin Stamm, and Elena Besley. An integral equation approach to calculate electrostatic interactions in many-body dielectric systems. Journal of Computational Physics, 371(??):712–731, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303930.

Lav:2019:FDD

[LSP19] Chitrarth Lav, Richard D. Sandberg, and Jimmy Philip. A framework to develop data-driven turbulence models for flows with organised unsteadiness. Journal of Computational Physics, 383(??):148–165, April 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300476.

Lind:2016:ICF

[LSR16] S. J. Lind, P. K. Stansby, and B. D. Rogers. Incompressiblecompressible flows with a transient discontinuous interface using smoothed particle hydrodynamics (SPH). Journal of Computational Physics, 309(??):129–147, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008207.

Lipton:2016:ULQ

[LSS16] Robert Lipton, Paul Sinz, and Michael Stuebner. Uncertain loading and quantifying maximum energy concentration within composite structures. Journal of Computational Physics, 325(??):38–52, November 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302893.

Liu:2015:TDN

[LSTkM15] Yan Liu, Weidong Shen, Baolin Tian, and De kang Mao. A two dimensional nodal Riemann solver based on one dimensional Riemann solver for a cell-centered Lagrangian scheme. Journal of Computational Physics, 284(??):566–594, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008481.

Lewis:2016:ITA

[LSWF16] Allison Lewis, Ralph Smith, Brian Williams, and Victor Figueroa. An information theoretic approach to use highfidelity codes to calibrate low-fidelity codes. Journal of Computational Physics, 324(??):24–43, November 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303412.

Luo:2015:MCL

[LSYF15] Kun Luo, Changxiao Shao, Yue Yang, and Jianren Fan. A mass conserving level set method for detailed numerical simulation of liquid atomization. Journal of Computational Physics, 298(??):495–519, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003952.

Liu:2018:ESH

[LSZ18] Yong Liu, Chi-Wang Shu, and Mengping Zhang. Entropy stable high order discontinuous Galerkin methods for ideal compressible MHD on structured meshes. Journal of Computational Physics, 354(??):163–178, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308100.

Li:2015:NAM

[LT15] Boxiao Li and Hamdi A. Tchelepi. Nonlinear analysis of multiphase transport in porous media in the presence of viscous, buoyancy, and capillary forces. Journal of Computational Physics, 297(??):104–131, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003381.

Li:2017:ITM

[LT17a] Chung-Gang Li and Makoto Tsubokura. An implicit turbulence model for low-Mach Roe scheme using truncated Navier–Stokes equations. Journal of Computational Physics, 345(??):462–474, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304072.

Lu:2017:CSAb

[LT17b] Jianfeng Lu and Kyle Thicke. Cubic scaling algorithms for RPA correlation using interpolative separable density fitting. Journal of Computational Physics, 351(??):187–202, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730671X.

Lu:2017:OMM

[LT17c]  Jianfeng Lu and Kyle Thicke.  Orbital minimization

method with l^{1} regularization. Journal of Computational Physics, 336(??):87–103, May 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300980.

Lacis:2016:SFS

[LTB16a] Uğis Lacis, Kunihiko Taira, and Shervin Bagheri. A stable fluid-structure-interaction solver for low-density rigid bodies using the immersed boundary projection method. Journal of Computational Physics, 305(??):300–318, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007184.

Lisitsa:2016:CDG

[LTB16b] Vadim Lisitsa, Vladimir Tcheverda, and Charlotte Botter. Combination of the discontinuous Galerkin method with finite differences for simulation of seismic wave propagation. Journal of Computational Physics, 311(??):142–157, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000590.

Lehtikangas:2015:FEA

[LTKA15] O. Lehtikangas, T. Tarvainen, A. D. Kim, and S. R. Arridge. Finite element approximation of the radiative transport equation in a medium with piece-wise constant refractive index. Journal of Computational Physics, 282(??):345–359, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007876.

Lindblom:2016:CRM

[LTR16] Lee Lindblom, Nicholas W. Taylor, and Oliver Rinne. Constructing reference metrics on multicube representations of arbitrary manifolds. Journal of Computational Physics, 313(??):31–56, May 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116000930.

Luan:2017:PIE

[LTR17] Vu Thai Luan, Mayya Tokman, and Greg Rainwater. Preconditioned implicit-exponential integrators (IMEXP) for stiff PDEs. Journal of Computational Physics, 335(??):846–864, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300700.

Liu:2018:LOW

[LTW18] Jiangguo Liu, Simon Tavener, and Zhuoran Wang. The lowest-order weak Galerkin finite element method for the Darcy equation on quadrilateral and hybrid meshes. Journal of Computational Physics, 359(??):312–330, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300020.

Liu:2018:AFC

[LTWZ18] Jian-Guo Liu, Min Tang, Li Wang, and Zhennan Zhou. An accurate front capturing scheme for tumor growth models with a free boundary limit. Journal of Computational Physics, 364(??):73–94, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301657.

Liu:2017:HOT

[LTXB17] Youshan Liu, Jiwen Teng, Tao Xu, and José Badal. Higherorder triangular spectral element method with optimized cubature points for seismic wavefield modeling. Journal of Computational Physics, 336(??):458–480, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300852.

Luchko:2015:WDD

[Luc15] Yuri Luchko. Wave-diffusion dualism of the neutral-fractional processes. Journal of Computational Physics, 293(??):40–52, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004124.

Lavalle:2015:NRM

[LVB^{+}15] G. Lavalle, J.-P. Vila, G. Blanchard, C. Laurent, and F. Charru. A numerical reduced model for thin liquid films sheared by a gas flow. Journal of Computational Physics, 301(??):119–140, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005458.

Lu:2018:IAP

[LVL18] Zhiming Lu, Velimir V. Vesselinov, and Hongzhuan Lei. Identifying arbitrary parameter zonation using multiple level set functions. Journal of Computational Physics, 364(??):257–273, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301682.

Lalanne:2015:CVT

[LVTR15] Benjamin Lalanne, Lucia Rueda Villegas, Sébastien Tanguy, and Frédéric Risso. On the computation of viscous terms for incompressible two-phase flows with Level Set/ Ghost Fluid Method. Journal of Computational Physics, 301(??):289–307, November 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911500563X.

Li:2015:NSI

[LW15a] Peijun Li and Yuliang Wang. Numerical solution of an inverse obstacle scattering problem with near-field data. Journal of Computational Physics, 290(??):157–168, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001230.

Li:2015:GLE

[LW15b] Yu-Wen Li and Xinyuan Wu. General local energy-preserving integrators for solving multi-symplectic Hamiltonian PDEs. Journal of Computational Physics, 301(??):141–166, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005501.

Lee:2017:AEG

[LW17a] Sanghyun Lee and Mary F. Wheeler. Adaptive enriched Galerkin methods for miscible displacement problems with entropy residual stabilization. Journal of Computational Physics, 331(??):19–37, February 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305952.

Li:2017:TEH

[LW17b] Jiequan Li and Yue Wang. Thermodynamical effects and high resolution methods for compressible fluid flows. Journal of Computational Physics, 343(??):340–354, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303273.

Li:2017:HON

[LW17c] Xuhao Li and Patricia J. Y. Wong. A higher order nonpolynomial spline method for fractional sub-diffusion problems. Journal of Computational Physics, 328(??):46–65, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304995.

Liu:2017:AHO

[LW17d] Changying Liu and Xinyuan Wu. Arbitrarily high-order time-stepping schemes based on the operator spectrum theory for high-dimensional nonlinear Klein–Gordon equations. Journal of Computational Physics, 340(??):243–275, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302334.

Liu:2017:FES

[LW17e] Hailiang Liu and Zhongming Wang. A free energy satisfying discontinuous Galerkin method for one-dimensional Poisson–Nernst–Planck systems. Journal of Computational Physics, 328(??):413–437, January 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305022.

Lee:2018:EGM

[LW18] Sanghyun Lee and Mary F. Wheeler. Enriched Galerkin methods for two-phase flow in porous media with capillary pressure. Journal of Computational Physics, 367(??):65–86, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301918.

Liu:2016:LBM

[LWB^{+}16] Haihu Liu, Lei Wu, Yan Ba, Guang Xi, and Yonghao Zhang. A lattice Boltzmann method for axisymmetric multicomponent flows with high viscosity ratio. Journal of Computational Physics, 327(??):873–893, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305009.

Luo:2017:EFE

[LWC17] Li Luo, Xiao-Ping Wang, and Xiao-Chuan Cai. An efficient finite element method for simulation of droplet spreading on a topologically rough surface. Journal of Computational Physics, 349(??):233–252, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305818.

Loppi:2019:LAP

[LWJV19] N. A. Loppi, F. D. Witherden, A. Jameson, and P. E. Vincent. Locally adaptive pseudo-time stepping for high-order Flux Reconstruction. Journal of Computational Physics, 399(??):Article 108913, December 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306187.

Li:2017:SCV

[LWL17]  Yiqun Li, Boying Wu, and Melvin Leok.  Spectral-

collocation variational integrators. Journal of Computational Physics, 332(??):83–98, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306489.

Lin:2018:RNG

[LWL18] Y. Lin, F. Wang, and B. Liu. Random number generators for large-scale parallel Monte Carlo simulations on FPGA. Journal of Computational Physics, 360(??):93–103, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300391.

Li:2017:BDV

[LWLC17] Zhipeng Li, Hongchun Wu, Yunzhao Li, and Liangzhi Cao. Block-diagonalization of the variational nodal response matrix using the symmetry group theory. Journal of Computational Physics, 351(??):230–253, December 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306897.

Luo:2019:IDF

[LWTF19] Kun Luo, Zhuo Wang, Junhua Tan, and Jianren Fan. An improved direct-forcing immersed boundary method with inward retraction of Lagrangian points for simulation of particle-laden flows. Journal of Computational Physics, 376(??):210–227, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306399.

Liang:2018:SSG

[LWWY18] Wenquan Liang, Xiu Wu, Yanfei Wang, and Changchun Yang. A simplified staggered-grid finite-difference scheme and its linear solution for the first-order acoustic wave-equation modeling. Journal of Computational Physics, 374(??):863–872, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305333.

Liu:2019:UGK

[LWX19] Chang Liu, Zhao Wang, and Kun Xu. A unified gaskinetic scheme for continuum and rarefied flows VI: Dilute disperse gas-particle multiphase system. Journal of Computational Physics, 386(??):264–295, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300890.

Li:2017:STF

[LWY17] Zheng Li, Hong Wang, and Danping Yang. A space-time fractional phase-field model with tunable sharpness and decay behavior and its efficient numerical simulation. Journal of Computational Physics, 347(??):20–38, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304904.

Liu:2018:DME

[LWY18] Yi Liu, Gang Wang, and Zhengyin Ye. Dynamic mode extrapolation to improve the efficiency of dual time stepping method. Journal of Computational Physics, 352(??):190–212, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307039.

Li:2016:NFI

[LWZ16] Peijun Li, Yuliang Wang, and Yue Zhao. Near-field imaging of biperiodic surfaces for elastic waves. Journal of Computational Physics, 324(??):1–23, November 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303254.

Li:2019:POD

[LWZ19] Richen Li, Qingbiao Wu, and Shengfeng Zhu. Proper orthogonal decomposition with SUPG-stabilized isogeometric analysis for reduced order modelling of unsteady convectiondominated convection-diffusion-reaction problems. Journal of Computational Physics, 387(??):280–302, June 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301688.

Li:2016:HOF

[LX16] Gang Li and Yulong Xing. High order finite volume WENO schemes for the Euler equations under gravitational fields. Journal of Computational Physics, 316(??):145–163, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630047X.

Li:2018:WBD

[LX18] Gang Li and Yulong Xing. Well-balanced discontinuous Galerkin methods with hydrostatic reconstruction for the Euler equations with gravitation. Journal of Computational Physics, 352(??):445–462, January 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307362.

Li:2015:SIA

[LXC^{+}15] Zhilin Li, Li Xiao, Qin Cai, Hongkai Zhao, and Ray Luo. A semi-implicit augmented IIM for Navier–Stokes equations with open, traction, or free boundary conditions. Journal of Computational Physics, 297(??):182–193, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003216.

Li:2019:CFM

[LXC19] Maojun Li, Liwei Xu, and Yongping Cheng. A CDG-FE method for the two-dimensional Green-Naghdi model with the enhanced dispersive property. Journal of Computational Physics, 399(??):Article 108953, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306588.

Liang:2017:GBL

[LXL17] Yihao Liang, Xiangjun Xing, and Yaohang Li. A GPUbased large-scale Monte Carlo simulation method for systems with long-range interactions. Journal of Computational Physics, 338(??):252–268, June 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301729.

Liu:2016:UGK

[LXSC16] Chang Liu, Kun Xu, Quanhua Sun, and Qingdong Cai. A unified gas-kinetic scheme for continuum and rarefied flows IV: Full Boltzmann and model equations. Journal of Computational Physics, 314(??):305–340, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001662.

Larson:2015:FMB

[LY15a] David J. Larson and Christopher V. Young. A finite mass based method for Vlasov–Poisson simulations. Journal of Computational Physics, 284(??):171–185, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008390.

Liao:2015:EGCb

[LY15b] Fei Liao and Zhengyin Ye. Extending geometric conservation law to cell-centered finite difference methods on moving and deforming grids. Journal of Computational Physics, 303(??):212–221, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006257.

Lu:2015:CER

[LY15c] Jianfeng Lu and Lexing Ying. Compression of the electron repulsion integral tensor in tensor hypercontraction format with cubic scaling cost. Journal of Computational Physics, 302(??):329–335, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006075.

Liu:2016:MSP

[LY16a] Chein-Shan Liu and D. L. Young. A multiple-scale Pascal polynomial for 2D Stokes and inverse Cauchy–Stokes problems. Journal of Computational Physics, 312(??):1–13, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000711.

Liu:2016:HPD

[LY16b] Hailiang Liu and Nianyu Yi. A Hamiltonian preserving discontinuous Galerkin method for the generalized Korteweg–de Vries equation. Journal of Computational Physics, 321(??):776–796, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302352.

Liu:2016:CPF

[LY16c]  Yu Liu and Xiping Yu.  A coupled phase-field and

volume-of-fluid method for accurate representation of limiting water wave deformation. Journal of Computational Physics, 321(??):459–475, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302042.

Lu:2016:SPS

[LY16d]  Jianfeng Lu and Lexing Ying.  Sparsifying precondi-

tioner for soliton calculations. Journal of Computational Physics, 315(??):458–466, June 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116300262.

Lu:2017:CSAa

[LY17] Jianfeng Lu and Haizhao Yang. A cubic scaling algorithm for excited states calculations in particle-particle random phase approximation. Journal of Computational Physics, 340(??):297–308, July 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302565.

Liu:2019:SPT

[LY19] Fei Liu and Lexing Ying. Sparsifying preconditioner for the time-harmonic Maxwell’s equations. Journal of Computational Physics, 376(??):913–923, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306612.

Lorin:2016:FGA

[LYA16] E. Lorin, X. Yang, and X. Antoine. Frozen Gaussian approximation based domain decomposition methods for the linear Schrödinger equation beyond the semi-classical regime. Journal of Computational Physics, 315(??):221–237, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000991.

Lester:2018:RSB

[LYB18] Christopher Lester, Christian A. Yates, and Ruth E. Baker. Robustly simulating biochemical reaction kinetics using multilevel Monte Carlo approaches. Journal of Computational Physics, 375(??):1401–1423, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304236.

Li:2016:FDM

[LYC16] Changpin Li, Qian Yi, and An Chen. Finite difference methods with non-uniform meshes for nonlinear fractional differential equations. Journal of Computational Physics, 316(??):614– 631, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300857.

Le:2017:MCS

[LYCC17] Hai P. Le, Bokai Yan, Russel E. Caflisch, and Jean-Luc Cambier. Monte Carlo simulation of excitation and ionization collisions with complexity reduction. Journal of Computational Physics, 346(??):480–496, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304746.

Lin:2019:NAI

[LYD19] Lianlei Lin, Zhiguo Yang, and Suchuan Dong. Numerical approximation of incompressible Navier–Stokes equations based on an auxiliary energy variable. Journal of Computational Physics, 388(??):1–22, July 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301950.

Lu:2017:MMN

[LYDB17] Xiaoxin Lu, Julien Yvonnet, Fabrice Detrez, and Jinbo Bai. Multiscale modeling of nonlinear electric conductivity in graphene-reinforced nanocomposites taking into account tunnelling effect. Journal of Computational Physics, 337(??):116– 131, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300797.

Lin:2019:VCB

[LYKW19] Stephen Lin, Jinhui Yan, Dmitriy Kats, and Gregory J. Wagner. A volume-conserving balanced-force level set method on unstructured meshes using a control volume finite element formulation. Journal of Computational Physics, 380(??):119–142, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307769.

Li:2019:FJP

[LYL19] Jiao Li, Jinyong Ying, and Benzhuo Lu. A flux-jump preserved gradient recovery technique for accurately predicting the electrostatic field of an immersed biomolecule. Journal of Computational Physics, 396(??):193–208, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304577.

Lei:2017:SPI

[LYLK17] Huan Lei, Xiu Yang, Zhen Li, and George Em Karniadakis. Systematic parameter inference in stochastic mesoscopic modeling. Journal of Computational Physics, 330(??):571–593, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305277.

Lee:2019:SIU

[LYM19] Byungjoon Lee, Gangjoon Yoon, and Chohong Min. A semi-implicit and unconditionally stable approximation of the surface tension in two-phase fluids. Journal of Computational Physics, 397(??):Article 108829, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305133.

Lopez-Yela:2017:FEM

[LYPP17] A. López-Yela and J. M. Pérez-Pardo. Finite element method to solve the spectral problem for arbitrary self-adjoint extensions of the Laplace–Beltrami operator on manifolds with a boundary. Journal of Computational Physics, 347(??):235–260, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304977.

Liao:2015:EGCa

[LYZ15] Fei Liao, Zhengyin Ye, and Lingxia Zhang. Extending geometric conservation law to cell-centered finite difference methods on stationary grids. Journal of Computational Physics, 284(??):419–433, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008572.

Li:2018:WGL

[LYZ18] Jichun Li, Xiu Ye, and Shangyou Zhang. A weak Galerkin least-squares finite element method for div–curl systems. Journal of Computational Physics, 363(??):79–86, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301189.

Liu:2019:EGK

[LYZ19] Sha Liu, Yuehan Yang, and Chengwen Zhong. An extended gas-kinetic scheme for shock structure calculations. Journal of Computational Physics, 390(??):1–24, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302499.

Li:2015:NAN

[LZ15a] Buyang Li and Zhimin Zhang. A new approach for numerical simulation of the time-dependent Ginzburg–Landau equations. Journal of Computational Physics, 303(??):238–250, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006579.

Liu:2015:MTM

[LZ15b] Haihu Liu and Yonghao Zhang. Modelling thermocapillary migration of a microfluidic droplet on a solid surface. Journal of Computational Physics, 280(??):37–53, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006512.

Li:2016:EIN

[LZ16] Dongfang Li and Jiwei Zhang. Efficient implementation to numerically solve the nonlinear time fractional parabolic problems on unbounded spatial domain. Journal of Computational Physics, 322(??):415–428, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302716.

Li:2017:RSF

[LZ17a] Lingxiao Li and Weiying Zheng. A robust solver for the finite element approximation of stationary incompressible MHD equations in 3D. Journal of Computational Physics, 351(??):254–270, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911730685X.

Lin:2017:HOF

[LZ17b] Zhi Lin and Qinghai Zhang. High-order finite-volume solutions of the steady-state advection-diffusion equation with nonlinear Robin boundary conditions. Journal of Computational Physics, 345(??):358–372, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303984.

Lee:2018:PPG

[LZ18] Wonjung Lee and Nicholas Zabaras. Parallel probabilistic graphical model approach for nonparametric Bayesian inference. Journal of Computational Physics, 372(??):546–563, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304352.

Lin:2019:IEM

[LZ19] Junshan Lin and Hai Zhang. An integral equation method for numerical computation of scattering resonances in a narrow metallic slit. Journal of Computational Physics, 385(??):75– 105, May 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301044.

Li:2017:DDA

[LZB^{+}17] Zhiyong Li, Huaibao Zhang, Sean C. C. Bailey, Jesse B. Hoagg, and Alexandre Martin. A data-driven adaptive Reynoldsaveraged Navier–Stokes k–ω model for turbulent flow. Journal of Computational Physics, 345(??):111–131, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303789.

Li:2019:FCM

[LZHM19] Shuai Li, A-Man Zhang, Rui Han, and Qingwei Ma. 3D full coupling model for strong interaction between a pulsating bubble and a movable sphere. Journal of Computational Physics, 392(??):713–731, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303250.

Lu:2017:AOS

[LZL^{+}17] Zhen Lu, Hua Zhou, Shan Li, Zhuyin Ren, Tianfeng Lu, and Chung K. Law. Analysis of operator splitting errors for near-limit flame simulations. Journal of Computational Physics, 335(??):578–591, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300608.

Liu:2019:FPV

[LZL^{+}19] Qiancheng Liu, Jianfeng Zhang, Yongming Lu, Hongwei Gao, Shaolin Liu, and Hao Zhang. Fast Poynting-vector based wave-mode separation and RTM in 2D elastic TI media. Journal of Computational Physics, 381(??):27–41, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300051.

Liu:2019:ATD

[LZS^{+}19] Lei Liu, Junqi Zhang, Chongmin Song, Carolin Birk, Albert A. Saputra, and Wei Gao. Automatic three-dimensional acoustic-structure interaction analysis using the scaled boundary finite element method. Journal of Computational Physics, 395(??):432–460, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119304413.

Li:2019:AFC

[LZSG19] Yu Li, Tao Zhang, Shuyu Sun, and Xin Gao. Accelerating flash calculation through deep learning methods. Journal of Computational Physics, 394(??):153–165, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303596.

Liu:2015:SAD

[LZT^{+}15] F. Liu, P. Zhuang, I. Turner, V. Anh, and K. Burrage. A semialternating direction method for a 2-D fractional FitzHugh– Nagumo monodomain model on an approximate irregular domain. Journal of Computational Physics, 293(??):252–263, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004045.

Liao:2017:TSA

[LZT17] Qinzhuo Liao, Dongxiao Zhang, and Hamdi Tchelepi. A twostage adaptive stochastic collocation method on nested sparse grids for multiphase flow in randomly heterogeneous porous media. Journal of Computational Physics, 330(??):828–845, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305691.

Lu:2017:RPB

[LZW^{+}17] Haitian Lu, Jun Zhu, Chunwu Wang, Donghong Wang, and Ning Zhao. A Riemann problem based method for solving compressible and incompressible flows. Journal of Computational Physics, 330(??):1–20, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305551.

Li:2019:EEP

[LZW19] Jun Li, Jia Zhao, and Qi Wang. Energy and entropy preserving numerical approximations of thermodynamically consistent crystal growth models. Journal of Computational Physics, 382(??):202–220, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300142.

Liu:2015:NCC

[LZZS15] Xuliang Liu, Shuhai Zhang, Hanxin Zhang, and Chi-Wang Shu. A new class of central compact schemes with spectrallike resolution II: Hybrid weighted nonlinear schemes. Journal of Computational Physics, 284(??):133–154, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008444.

Motheau:2016:HON

[MA16] E. Motheau and J. Abraham. A high-order numerical algorithm for DNS of low-Mach-number reactive flows with detailed chemistry and quasi-spectral accuracy. Journal of Computational Physics, 313(??):430–454, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001315.

Mattsson:2017:HOA

[MA17] Ken Mattsson and Martin Almquist. A high-order accurate embedded boundary method for first order hyperbolic equations. Journal of Computational Physics, 334(??):255–279, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306969.

Matsushita:2019:WCS

[MA19] Shintaro Matsushita and Takayuki Aoki. A weakly compressible scheme with a diffuse-interface method for low Mach number two-phase flows. Journal of Computational Physics, 376(??):838–862, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830682X.

Mokbel:2018:PFM

[MAA18] Dominic Mokbel, Helmut Abels, and Sebastian Aland. A phase-field model for fluid-structure interaction. Journal of Computational Physics, 372(??):823–840, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304418.

Machado:2015:NCL

[Mac15] J. Tenreiro Machado. Numerical calculation of the left and right fractional derivatives. Journal of Computational Physics, 293(??):96–103, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003751.

Macrossan:2016:RCL

[Mac16] Michael N. Macrossan. Restricted collision list method for faster Direct Simulation Monte-Carlo (DSMC) collisions. Journal of Computational Physics, 319(??):1–8, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301528.

Magri:2019:ACD

[Mag19] Luca Magri. Adjoint characteristic decomposition of onedimensional waves. Journal of Computational Physics, 388 (??):454–461, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119302153.

Miura:2016:HES

[MAH16] Hideaki Miura, Keisuke Araki, and Fujihiro Hamba. Hall effects and sub-grid-scale modeling in magnetohydrodynamic turbulence simulations. Journal of Computational Physics, 316(??):385–395, July 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116300651.

Mahady:2015:VFM

[MAK15] Kyle Mahady, Shahriar Afkhami, and Lou Kondic. A volume of fluid method for simulating fluid/fluid interfaces in contact with solid boundaries. Journal of Computational Physics, 294(??):243–257, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002065.

Massimo:2016:CTE

[MAM16] F. Massimo, S. Atzeni, and A. Marocchino. Comparisons of time explicit hybrid kinetic-fluid code architect for Plasma Wakefield Acceleration with a full PIC code. Journal of Computational Physics, 327(??):841–850, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305010.

Martins:2017:CCL

[MAP17] Diogo M. C. Martins, Duarte M. S. Albuquerque, and José C. F. Pereira. Continuity constrained least-squares interpolation for SFO suppression in immersed boundary methods. Journal of Computational Physics, 336(??):608–626, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301195.

Marskar:2019:ACE

[Mar19] Robert Marskar. An adaptive Cartesian embedded boundary approach for fluid simulations of two- and three-dimensional low temperature plasma filaments in complex geometries. Journal of Computational Physics, 388(??):624–654, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302190.

Masiello:2018:FSC

[Mas18] Emiliano Masiello. “On the fly” stabilization of the Coarse-Mesh Finite Difference acceleration for multidimensional discrete-ordinates transport calculations. Journal of Computational Physics, 373(??):1–27, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304340.

Mattsson:2017:DNU

[Mat17] Ken Mattsson. Diagonal-norm upwind SBP operators. Journal of Computational Physics, 335(??):283–310, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730058X.

Mattsson:2018:BOD

[MAvdW18] Ken Mattsson, Martin Almquist, and Edwin van der Weide. Boundary optimized diagonal-norm SBP operators. Journal of Computational Physics, 374(??):1261–1266, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303887.

Muller:2015:HOA

[MB15] Lucas O. Müller and Pablo J. Blanco. A high order approximation of hyperbolic conservation laws in networks: Application to one-dimensional blood flow. Journal of Computational Physics, 300(??):423–437, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005094.

Moore:2019:HPP

[MBA19] W. C. Moore, S. Balachandar, and G. Akiki. A hybrid pointparticle force model that combines physical and data-driven approaches. Journal of Computational Physics, 385(??):187– 208, May 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301111.

Mitscha-Baude:2017:AIM

[MBBKTH17] Gregor Mitscha-Baude, Andreas Buttinger-Kreuzhuber, Gerhard Tulzer, and Clemens Heitzinger. Adaptive and iterative methods for simulations of nanopores with the PNP–Stokes equations. Journal of Computational Physics, 338(??):452–476, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301870.

Moguen:2015:SLM

[MBD15] Yann Moguen, Pascal Bruel, and Erik Dick. Solving low Mach number Riemann problems by a momentum interpolation method. Journal of Computational Physics, 298(??):741–746, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004350.

Moguen:2019:CMI

[MBD19] Yann Moguen, Pascal Bruel, and Erik Dick. A combined momentum-interpolation and advection upstream splitting pressure-correction algorithm for simulation of convective and acoustic transport at all levels of Mach number. Journal of Computational Physics, 384(??):16–41, May 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300804.

Meng:2017:SAP

[MBHS17] F. Meng, J. W. Banks, W. D. Henshaw, and D. W. Schwendeman. A stable and accurate partitioned algorithm for conjugate heat transfer. Journal of Computational Physics, 344(??):51–85, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303315.

Magri:2016:SATa

[MBJ16] Luca Magri, Michael Bauerheim, and Matthew P. Juniper. Stability analysis of thermo-acoustic nonlinear eigenproblems in annular combustors. Part I. Sensitivity. Journal of Computational Physics, 325(??):395–410, November 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303278.

Moxley:2015:GFS

[MBM^{+}15] Frederick Ira Moxley III, Tim Byrnes, Baoling Ma, Yun Yan, and Weizhong Dai. A G-FDTD scheme for solving multidimensional open dissipative Gross–Pitaevskii equations. Journal of Computational Physics, 282(??):303–316, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007839.

Maunoury:2018:WSA

[MBM^{+}18] Matthieu Maunoury, Christophe Besse, Vincent Mouysset, Sébastien Pernet, and Pol-André Haas. Well-suited and adaptive post-processing for the visualization of hp simulation results. Journal of Computational Physics, 375(??):1179–1204, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306119.

Magri:2016:SATb

[MBNJ16] Luca Magri, Michael Bauerheim, Franck Nicoud, and Matthew P. Juniper. Stability analysis of thermo-acoustic nonlinear eigenproblems in annular combustors. Part II. Uncertainty quantification. Journal of Computational Physics, 325(??):411–421, November 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304004.

Mirfatah:2019:SPP

[MBS19] S. M. Mirfatah, B. Boroomand, and E. Soleimanifar. On the solution of 3D problems in physics: From the geometry definition in CAD to the solution by a meshless method. Journal of Computational Physics, 393(??):351–374, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303328.

MacDonald:2015:ECG

[MBSS15] Christopher L. MacDonald, Nirupama Bhattacharya, Brian P. Sprouse, and Gabriel A. Silva. Efficient computation of the Grünwald–Letnikov fractional diffusion derivative using adaptive time step memory. Journal of Computational Physics, 297(??):221–236, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003058.

Mittal:2017:PVD

[MBST17] Anshul Mittal, W. Roger Briley, Kidambi Sreenivas, and Lafayette K. Taylor. A parabolic velocity-decomposition method for wind turbines. Journal of Computational Physics, 330(??):650–667, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305368.

Medale:2015:HPC

[MC15] Marc Medale and Bruno Cochelin. High performance computations of steady-state bifurcations in 3D incompressible fluid flows by asymptotic numerical method. Journal of Computational Physics, 299(??):581–596, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004660.

Martin:2016:OPM

[MC16] Robert Scott Martin and Jean-Luc Cambier. Octree particle management for DSMC and PIC simulations. Journal of Computational Physics, 327(??):943–966, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000280.

Murashige:2017:NSP

[MC17] Sunao Murashige and Wooyoung Choi. A numerical study on parasitic capillary waves using unsteady conformal mapping. Journal of Computational Physics, 328(??):234–257, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305137.

Maeda:2018:ELM

[MC18] Kazuki Maeda and Tim Colonius. Eulerian–Lagrangian method for simulation of cloud cavitation. Journal of Computational Physics, 371(??):994–1017, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303358.

Mons:2016:RUV

[MCGS16] V. Mons, J.-C. Chassaing, T. Gomez, and P. Sagaut. Reconstruction of unsteady viscous flows using data assimilation schemes. Journal of Computational Physics, 316(??):255–280, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300638.

Miyawaki:2016:CIB

[MCHL16] Shinjiro Miyawaki, Sanghun Choi, Eric A. Hoffman, and Ching-Long Lin. A 4DCT imaging-based breathing lung model with relative hysteresis. Journal of Computational Physics, 326(??):76–90, December 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303965.

Malhotra:2019:TSS

[MCIGO19] Dhairya Malhotra, Antoine Cerfon, Lise-Marie Imbert-Gérard, and Michael O’Neil. Taylor states in stellarators: a fast high-order boundary integral solver. Journal of Computational Physics, 397(??):Article 108791, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119304759.

Marboeuf:2019:CEC

[MCL19] Alexis Marboeuf, Alexandra Claisse, and Patrick Le Tallec. Conservative and entropy controlled remap for multi-material ALE simulations with space-staggered schemes. Journal of Computational Physics, 390(??):66–92, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302505.

Mapakshi:2018:SVI

[MCN18] N. K. Mapakshi, J. Chang, and K. B. Nakshatrala. A scalable variational inequality approach for flow through porous media models with pressure-dependent viscosity. Journal of Computational Physics, 359(??):137–163, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300329.

Mohamad:2016:PDS

[MCS16] Mustafa A. Mohamad, Will Cousins, and Themistoklis P. Sapsis. A probabilistic decomposition-synthesis method for the quantification of rare events due to internal instabilities. Journal of Computational Physics, 322(??):288–308, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302728.

Miller:2019:IES

[MCS^{+}19] S. T. Miller, E. C. Cyr, J. N. Shadid, R. M. J. Kramer, E. G. Phillips, S. Conde, and R. P. Pawlowski. IMEX and exact sequence discretization of the multi-fluid plasma model. Journal of Computational Physics, 397(??):Article 108806, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119304905.

Marichal:2016:III

[MCW16] Yves Marichal, Philippe Chatelain, and Grégoire Winckelmans. Immersed interface interpolation schemes for particlemesh methods. Journal of Computational Physics, 326(??): 947–972, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304417.

Michels:2015:SAA

[MD15] Dominik L. Michels and Mathieu Desbrun. A semi-analytical approach to molecular dynamics. Journal of Computational Physics, 303(??):336–354, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006713.

Matheou:2016:SEL

[MD16] Georgios Matheou and Paul E. Dimotakis. Scalar excursions in large-eddy simulations. Journal of Computational Physics, 327(??):97–120, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303928.

Macias-Diaz:2017:SPM

[MD17] J. E. Macías-Díaz. A structure-preserving method for a class of nonlinear dissipative wave equations with Riesz space-fractional derivatives. Journal of Computational Physics, 351(??):40–58, December 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306885.

Macias-Diaz:2018:DCM

[MD18] J. E. Macías-Díaz. A dynamically consistent method to solve nonlinear multidimensional advection–reaction equations with fractional diffusion. Journal of Computational Physics, 365(??):71–88, August 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302146.

Motheau:2018:HAL

[MDAB18] Emmanuel Motheau, Max Duarte, Ann Almgren, and John B. Bell. A hybrid adaptive low-Mach number/ compressible method: Euler equations. Journal of Computational Physics, 372(??):1027–1047, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300469.

Michel-Dansac:2017:WBS

[MDBCF17] Victor Michel-Dansac, Christophe Berthon, Stéphane Clain, and Françoise Foucher. A well-balanced scheme for the shallow-water equations with topography or manning friction. Journal of Computational Physics, 335(??):115–154, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300190.

Moro:2019:FSC

[MDD^{+}19] Federico Moro, Daniele Desideri, Alberto Doria, Alvise Maschio, Cristian Medé, and Lorenzo Codecasa. A facesmoothed cell method for static and dynamic piezoelectric coupled problems on polyhedral meshes. Journal of Computational Physics, 386(??):84–109, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301299.

Moghaderi:2017:SAM

[MDDM17] Hamid Moghaderi, Mehdi Dehghan, Marco Donatelli, and Mariarosa Mazza. Spectral analysis and multigrid preconditioners for two-dimensional space-fractional diffusion equations. Journal of Computational Physics, 350(??):992–1011, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306459.

Macias-Diaz:2019:ASA

[MDH19] J. E. Macías-Díaz and A. S. Hendy. Algorithm for some anomalously diffusive hyperbolic systems in molecular dynamics: Theoretical analysis and pattern formation. Journal of Computational Physics, 397(??):Article 108863, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305479.

McCorquodale:2015:HOF

[MDHC15] P. McCorquodale, M. R. Dorr, J. A. F. Hittinger, and P. Colella. High-order finite-volume methods for hyperbolic conservation laws on mapped multiblock grids. Journal of Computational Physics, 288(??):181–195, May 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500008X.

Marrone:2016:CSP

[MDL16] S. Marrone, A. Di Mascio, and D. Le Touzé. Coupling of smoothed particle hydrodynamics with finite volume method for free-surface flows. Journal of Computational Physics, 310(??):161–180, April 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008074.

Mengaldo:2015:DTH

[MDM^{+}15] G. Mengaldo, D. De Grazia, D. Moxey, P. E. Vincent, and S. J. Sherwin. Dealiasing techniques for high-order spectral element methods on regular and irregular grids. Journal of Computational Physics, 299(??):56–81, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004301.

Mengaldo:2018:SEA

[MDMS18] Gianmarco Mengaldo, Daniele De Grazia, Rodrigo C. Moura, and Spencer J. Sherwin. Spatial eigensolution analysis of energy-stable flux reconstruction schemes and influence of the numerical flux on accuracy and robustness. Journal of Computational Physics, 358(??):1–20, April 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309051.

Moguen:2015:GTS

[MDP^{+}15] Yann Moguen, Simon Delmas, Vincent Perrier, Pascal Bruel, and Erik Dick. Godunov-type schemes with an inertia term for unsteady full Mach number range flow calculations. Journal of Computational Physics, 281(??):556–590, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007232.

Monsorno:2018:TAC

[MDP18] Davide Monsorno, Athanassios A. Dimas, and Miltiadis V. Papalexandris. Time-accurate calculation of two-phase granular flows exhibiting compaction, dilatancy and nonlinear rheology. Journal of Computational Physics, 372(??):799–822, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911830439X.

Moon:2016:SEG

[MDT16] H. Moon, B. Donderici, and F. L. Teixeira. Stable evaluation of Green’s functions in cylindrically stratified regions with uniaxial anisotropic layers. Journal of Computational Physics, 325(??):174–200, November 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303680.

Manzini:2016:LFS

[MDVM16] G. Manzini, G. L. Delzanno, J. Vencels, and S. Markidis. A Legendre–Fourier spectral method with exact conservation laws for the Vlasov–Poisson system. Journal of Computational Physics, 317(??):82–107, July 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300705.

Mattsson:2018:SAF

[MDW18] Ken Mattsson, Eric M. Dunham, and Jonatan Werpers. Simulation of acoustic and flexural-gravity waves in ice-covered oceans. Journal of Computational Physics, 373(??):230– 252, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118304388.

Melvin:2018:DAM

[Mel18] Thomas Melvin. Dispersion analysis of the P_{n} − P_{n−1}^{DG} mixed finite element pair for atmospheric modelling. Journal of Computational Physics, 355(??):342–365, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308574.

Moortgat:2016:MHV

[MF16a] Joachim Moortgat and Abbas Firoozabadi. Mixed-hybrid and vertex-discontinuous-Galerkin finite element modeling of multiphase compositional flow on 3D unstructured grids. Journal of Computational Physics, 315(??):476–500, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300171.

Moradi:2016:MAS

[MF16b] H. V. Moradi and J. M. Floryan. A method for analysis of stability of flows in ribbed annuli. Journal of Computational Physics, 314(??):35–59, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001418.

Martin:2017:SMR

[MF17] Bradley Martin and Bengt Fornberg. Seismic modeling with radial basis function-generated finite differences (RBF–FD)

— a simplified treatment of interfaces. Journal of Computational Physics, 335(??):828–845, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300815.

Mora:2018:NSS

[MFB18] C. M. Mora, J. Fernández, and R. Biscay. Numerical solution of stochastic quantum master equations using stochastic interacting wave functions. Journal of Computational Physics, 367(??):28–48, August 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302122.

Moustafa:2019:EPS

[MFF^{+}19] Salli Moustafa, François Févotte, Mathieu Faverge, Laurent Plagne, and Pierre Ramet. Efficient parallel solution of the 3D stationary Boltzmann transport equation for diffusive problems. Journal of Computational Physics, 388(??):335–349, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302025.

Muller:2015:SDP

[MFG15] Kathrin Müller, Dmitry A. Fedosov, and Gerhard Gompper. Smoothed dissipative particle dynamics with angular momentum conservation. Journal of Computational Physics, 281(??):301–315, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114006998.

Mainardi:2015:CMP

[MG15a] Francesco Mainardi and Roberto Garrappa. On complete monotonicity of the Prabhakar function and non-Debye relaxation in dielectrics. Journal of Computational Physics, 293(??):70–80, July 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400552X.

Marras:2015:PFD

[MG15b] Simone Marras and Francis X. Giraldo. A parameter-free dynamic alternative to hyper-viscosity for coupled transport equations: Application to the simulation of 3D squall lines using spectral elements. Journal of Computational Physics, 283(??):360–373, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114008158.

Marie:2017:AFL

[MG17] Simon Marié and Xavier Gloerfelt. Adaptive filtering for the lattice Boltzmann method. Journal of Computational Physics, 333(??):212–226, March 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306702.

Mistani:2018:IDM

[MGB^{+}18] Pouria Mistani, Arthur Guittet, Daniil Bochkov, Joshua Schneider, Dionisios Margetis, Christian Ratsch, and Frederic Gibou. The island dynamics model on parallel quadtree grids. Journal of Computational Physics, 361(??):150–166, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300767.

Mirzadeh:2016:PLS

[MGBG16] Mohammad Mirzadeh, Arthur Guittet, Carsten Burstedde, and Frederic Gibou. Parallel level-set methods on adaptive tree-based grids. Journal of Computational Physics, 322 (??):345–364, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911630242X.

Mu:2018:ELE

[MGCW18] Zhenguo Mu, Yuezheng Gong, Wenjun Cai, and Yushun Wang. Efficient local energy dissipation preserving algorithms for the Cahn–Hilliard equation. Journal of Computational Physics, 374(??):654–667, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305266.

Matous:2017:RPN

[MGKG17] Karel Matous, Marc G. D. Geers, Varvara G. Kouznetsova, and Andrew Gillman. A review of predictive nonlinear theories for multiscale modeling of heterogeneous materials. Journal of Computational Physics, 330(??):192–220, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305782.

Mistani:2019:PVB

[MGPG19] Pouria Mistani, Arthur Guittet, Clair Poignard, and Frederic Gibou. A parallel Voronoi-based approach for mesoscale simulations of cell aggregate electropermeabilization. Journal of Computational Physics, 380(??):48–64, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118308052.

Maltba:2018:NPM

[MGT18] Tyler Maltba, Pierre A. Gremaud, and Daniel M. Tartakovsky. Nonlocal PDF methods for Langevin equations with colored noise. Journal of Computational Physics, 367(??):87–101, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302390.

Munoz:2017:HFP

[MH17] Raúl Pagán Muñoz and Maarten Hornikx. Hybrid Fourier pseudospectral/discontinuous Galerkin time-domain method for wave propagation. Journal of Computational Physics, 348 (??):416–432, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117305545.

Meng:2018:NEF

[MH18a] Xucheng Meng and Guanghui Hu. A NURBS-enhanced finite volume solver for steady Euler equations. Journal of Computational Physics, 359(??):77–92, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730935X.

Misev:2018:SDO

[MH18b] Cyril Misev and Nicholas J. Hills. Steepest descent optimisation of Runge–Kutta coefficients for second order implicit finite volume CFD codes. Journal of Computational Physics, 354(??):576–592, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306599.

Meliga:2019:TAC

[MH19] P. Meliga and E. Hachem. Time-accurate calculation and bifurcation analysis of the incompressible flow over a square cavity using variational multiscale modeling. Journal of Computational Physics, 376(??):952–972, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306387.

Mohand:2019:UFM

[MHGL19] H. Si Hadj Mohand, H. Hoang, G. Galliero, and D. Legendre. On the use of a friction model in a volume of fluid solver for the simulation of dynamic contact lines. Journal of Computational Physics, 393(??):29–45, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303304.

Misztal:2015:DAL

[MHGM^{+}15] Marek Krzysztof Misztal, Anier Hernandez-Garcia, Rastin Matin, Henning Osholm Sørensen, and Joachim Mathiesen. Detailed analysis of the lattice Boltzmann method on unstructured grids. Journal of Computational Physics, 297(??):316–339, September 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003538.

Mei:2019:EIQ

[MHH19] Lijie Mei, Li Huang, and Shixiang Huang. Exponential integrators with quadratic energy preservation for linear Poisson systems. Journal of Computational Physics, 387(??):446–454, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301883.

Ma:2016:RPI

[MHHX16] Yicong Ma, Kaibo Hu, Xiaozhe Hu, and Jinchao Xu. Robust preconditioners for incompressible MHD models. Journal of Computational Physics, 316(??):721–746, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300602.

Meng:2015:BDE

[MHJ15] Fei Meng, Xiaodong Huang, and Baohua Jia. Bi-directional evolutionary optimization for photonic band gap structures. Journal of Computational Physics, 302(??):393–404, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005914.

Miao:2017:CTD

[MHL17] Sha Miao, Kelli Hendrickson, and Yuming Liu. Computation of three-dimensional multiphase flow dynamics by Fully-Coupled Immersed Flow (FCIF) solver. Journal of Computational Physics, 350(??):97–116, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306198.

Mohamed:2016:DEC

[MHS16] Mamdouh S. Mohamed, Anil N. Hirani, and Ravi Samtaney. Discrete exterior calculus discretization of incompressible Navier–Stokes equations over surface simplicial meshes. Journal of Computational Physics, 312(??):175–191, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000929.

Mishra:2019:SAM

[MHT^{+}19] A. Mishra, A. Hemeda, M. Torabi, J. Palko, S. Goyal, and Y. Ma. A simple analytical model of complex wall in multibody dissipative particle dynamics. Journal of Computational Physics, 396(??):416–426, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S002199911930484X.

Meyers:2015:NDE

[MHZ^{+}15] M. D. Meyers, C.-K. Huang, Y. Zeng, S. A. Yi, and B. J. Albright. On the numerical dispersion of electromagnetic particle-in-cell code: Finite grid instability. Journal of Computational Physics, 297(??):565–583, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500371X.

Malek:2019:PUF

[MIM^{+}19] Mustapha Malek, Nouh Izem, M. Shadi Mohamed, Mohammed Seaid, and Omar Laghrouche. A partition of unity finite element method for three-dimensional transient diffusion problems with sharp gradients. Journal of Computational Physics, 396(??):702–717, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S002199911930470X.

Marti:2016:FSM

[MJ16] P. Marti and A. Jackson. A fully spectral methodology for magnetohydrodynamic calculations in a whole sphere. Journal of Computational Physics, 305(??):403–422, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007330.

Miquel:2017:HCF

[MJ17] Benjamin Miquel and Keith Julien. Hybrid Chebyshev function bases for sparse spectral methods in parity-mixed PDEs on an infinite domain. Journal of Computational Physics, 349(??):474–500, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306058.

Markl:2015:FSN

[MK15] Matthias Markl and Carolin Körner. Free surface Neumann boundary condition for the advection-diffusion lattice Boltzmann method. Journal of Computational Physics, 301(??):230–246, November 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005604.

Mao:2017:FBE

[MK17] Zhiping Mao and George Em Karniadakis. Fractional Burgers equation with nonlinear non-locality: Spectral vanishing viscosity and local discontinuous Galerkin methods. Journal of Computational Physics, 336(??):143–163, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300645.

Mikida:2019:MRT

[MKB19] Cory Mikida, Andreas Klöckner, and Daniel Bodony. Multirate time integration on overset meshes. Journal of Computational Physics, 396(??):325–346, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304280.

Meena:2017:PPH

[MKC17] Asha Kumari Meena, Harish Kumar, and Praveen Chandrashekar. Positivity-preserving high-order discontinuous Galerkin schemes for Ten-Moment Gaussian closure equations. Journal of Computational Physics, 339(??):370–395, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730219X.

Maxian:2018:CEB

[MKS18] Ondrej Maxian, Andrew T. Kassen, and Wanda Strychalski. A continuous energy-based immersed boundary method for elastic shells. Journal of Computational Physics, 371(??):333–362, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303541.

Marcon:2019:HRP

[MKSP19] Julian Marcon, David A. Kopriva, Spencer J. Sherwin, and Joaquim Peiró. A high resolution PDE approach to quadrilateral mesh generation. Journal of Computational Physics, 399(??):Article 108918, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306230.

Munk:2017:TOM

[MKV^{+}17] David J. Munk, Timoleon Kipouros, Gareth A. Vio, Grant P. Steven, and Geoffrey T. Parks. Topology optimisation of micro fluidic mixers considering fluid-structure interactions with a coupled lattice Boltzmann algorithm. Journal of Computational Physics, 349(??):11–32, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730579X.

Malovichko:2017:ASA

[MKYZ17] M. Malovichko, N. Khokhlov, N. Yavich, and M. Zhdanov. Approximate solutions of acoustic 3D integral equation and their application to seismic modeling and full-waveform inversion. Journal of Computational Physics, 346(??):318–339, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304667.

Moyner:2016:MRS

[ML16] Olav Møyner and Knut-Andreas Lie. A multiscale restrictionsmoothed basis method for high contrast porous media represented on unstructured grids. Journal of Computational Physics, 304(??):46–71, January 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006725.

Muller:2016:CTV

[MLB16] Lucas O. Müller, Günter Leugering, and Pablo J. Blanco. Consistent treatment of viscoelastic effects at junctions in one-dimensional blood flow models. Journal of Computational Physics, 314(??):167–193, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001649.

Morgan:2018:RSM

[MLB18] Nathaniel R. Morgan, Xiaodong Liu, and Donald E. Burton. Reducing spurious mesh motion in Lagrangian finite volume and discontinuous Galerkin hydrodynamic methods. Journal of Computational Physics, 372(??):35–61, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303863.

Ma:2017:ESH

[MLI17] Peter C. Ma, Yu Lv, and Matthias Ihme. An entropy-stable hybrid scheme for simulations of transcritical real-fluid flows. Journal of Computational Physics, 340(??):330–357, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302176.

Morente:2018:PMS

[MLL18] Antoine Morente, Jérôme Laviéville, and Dominique Legendre. A penalization method for the simulation of bubbly flows. Journal of Computational Physics, 374(??):563–590, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305035.

Ma:2019:CGL

[MLL19] Lina Ma, Xiantao Li, and Chun Liu. Coarse-graining Langevin dynamics using reduced-order techniques. Journal of Computational Physics, 380(??):170–190, March 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307794.

Maljaars:2018:HDG

[MLM18] Jakob M. Maljaars, Robert Jan Labeur, and Matthias Möller. A hybridized discontinuous Galerkin framework for high-order particle-mesh operator splitting of the incompressible Navier–Stokes equations. Journal of Computational Physics, 358(??):150–172, April 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309300.

Moroney:2017:EFL

[MLMM17] Timothy J. Moroney, Dylan R. Lusmore, Scott W. Mc-Cue, and D. L. Sean McElwain. Extending fields in a level set method by solving a biharmonic equation. Journal of Computational Physics, 343(??):170–185, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303285.

McLean:2015:TSE

[MM15] William McLean and Kassem Mustapha. Time-stepping error bounds for fractional diffusion problems with non-smooth initial data. Journal of Computational Physics, 293(??):201–217, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006263.

MacArt:2016:SII

[MM16a] Jonathan F. MacArt and Michael E. Mueller. Semi-implicit iterative methods for low Mach number turbulent reacting flows: Operator splitting versus approximate factorization. Journal of Computational Physics, 326(??):569–595, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304284.

Mahalov:2016:TFL

[MM16b] A. Mahalov and M. Moustaoui. Time-filtered leapfrog integration of Maxwell equations using unstaggered temporal grids. Journal of Computational Physics, 325(??):98–115, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303655.

Mitchell:2016:HLI

[MM16c] Lawrence Mitchell and Eike Hermann Müller. High level implementation of geometric multigrid solvers for finite element problems: Applications in atmospheric modelling. Journal of Computational Physics, 327(??):1–18, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304582.

Muralidharan:2016:HOA

[MM16d] Balaji Muralidharan and Suresh Menon. A high-order adaptive Cartesian cut-cell method for simulation of compressible viscous flow over immersed bodies. Journal of Computational Physics, 321(??):342–368, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301954.

Mundis:2017:TOS

[MM17] Nathan L. Mundis and Dimitri J. Mavriplis. Toward an optimal solver for time-spectral fluid-dynamic and aeroelastic solutions on unstructured meshes. Journal of Computational Physics, 345(??):132–161, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303558.

Muralidharan:2018:SMB

[MM18] Balaji Muralidharan and Suresh Menon. Simulation of moving boundaries interacting with compressible reacting flows using a second-order adaptive Cartesian cut-cell method. Journal of Computational Physics, 357(??):230–262, March 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117309245.

Maric:2018:ESF

[MMB18] Tomislav Marić, Holger Marschall, and Dieter Bothe. An enhanced un-split face-vertex flux-based VoF method. Journal of Computational Physics, 371(??):967–993, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302432.

Mitchell:2019:EFB

[MMBP19] Matthew S. Mitchell, Matthew T. Miecnikowski, Gregory Beylkin, and Scott E. Parker. Efficient Fourier basis particle simulation. Journal of Computational Physics, 396(??):

837–847, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S002199911930508X.

Maier:2017:DES

[MML17] Matthias Maier, Dionisios Margetis, and Mitchell Luskin. Dipole excitation of surface plasmon on a conducting sheet: Finite element approximation and validation. Journal of Computational Physics, 339(??):126–145, June 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302012.

Mishra:2015:TDA

[MMMS15] Asitav Mishra, Karthik Mani, Dimitri Mavriplis, and Jay Sitaraman. Time dependent adjoint-based optimization for coupled fluid-structure problems. Journal of Computational Physics, 292(??):253–271, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001461.

MacDonald:2016:CMC

[MMNI16] G. MacDonald, J. A. Mackenzie, M. Nolan, and R. H. Insall. A computational method for the coupled solution of reaction-diffusion equations on evolving domains and manifolds: Application to a model of cell migration and chemotaxis. Journal of Computational Physics, 309(??):207–226, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008578.

Monteghetti:2018:EAD

[MMP18] Florian Monteghetti, Denis Matignon, and Estelle Piot. Energy analysis and discretization of nonlinear impedance boundary conditions for the time-domain linearized Euler equations. Journal of Computational Physics, 375(??):393–426, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305680.

Moura:2017:ERC

[MMPS17] R. C. Moura, G. Mengaldo, J. Peiró, and S. J. Sherwin. On the eddy-resolving capability of high-order discontinuous Galerkin approaches to implicit LES/under-resolved DNS of Euler turbulence. Journal of Computational Physics, 330(??):615–623, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305642.

Menon:2015:PAS

[MMSS15] Sandeep Menon, Kyle G. Mooney, K. G. Stapf, and David P. Schmidt. Parallel adaptive simplical re-meshing for deforming domain CFD computations. Journal of Computational Physics, 298(??):62–78, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003782.

Merrick:2018:NFV

[MMvR18] D. G. Merrick, A. G. Malan, and J. A. van Rooyen. A novel finite volume discretization method for advectiondiffusion systems on stretched meshes. Journal of Computational Physics, 362(??):220–242, June 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301074.

Muller:2015:RSE

[MMW15] Kei W. Müller, Christoph Meier, and Wolfgang A. Wall. Resolution of sub-element length scales in Brownian dynamics simulations of biopolymer networks with geometrically exact beam finite elements. Journal of Computational Physics, 303(??):185–202, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006312.

Mattsson:2004:SPO

[MN04] Ken Mattsson and Jan Nordström. Summation by parts operators for finite difference approximations of second derivatives. Journal of Computational Physics, 199(2):503–540, September 20, 2004. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999104000932. See corri-

gendum [MN17].

Mazaheri:2015:ISO

[MN15] Alireza Mazaheri and Hiroaki Nishikawa. Improved secondorder hyperbolic residual-distribution scheme and its extension to third-order on arbitrary triangular grids. Journal of Computational Physics, 300(??):455–491, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005070.

Mazaheri:2016:EHO

[MN16a] Alireza Mazaheri and Hiroaki Nishikawa. Efficient high-order discontinuous Galerkin schemes with first-order hyperbolic advection-diffusion system approach. Journal of Computational Physics, 321(??):729–754, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302315.

Michael:2016:HFN

[MN16b] L. Michael and N. Nikiforakis. A hybrid formulation for the numerical simulation of condensed phase explosives. Journal of Computational Physics, 316(??):193–217, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300493.

Mudunuru:2016:EMP

[MN16c] M. K. Mudunuru and K. B. Nakshatrala. On enforcing maximum principles and achieving element-wise species balance for advection–diffusion-reaction equations under the finite element method. Journal of Computational Physics, 305(??):448–493, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007135.

Mattsson:2017:CSP

[MN17] Ken Mattsson and Jan Nordström. Corrigendum to “Summation by parts operators for finite difference approximations of second derivatives” [J. Comput. Phys. 199 (2004) 503–540].

Journal of Computational Physics, 351(??):535, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307118. See [MN04].

Magee:2018:ASO

[MN18a] Daniel J. Magee and Kyle E. Niemeyer. Accelerating solutions of one-dimensional unsteady PDEs with GPUbased swept time-space decomposition. Journal of Computational Physics, 357(??):338–352, March 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309221.

Michael:2018:MPM

[MN18b] L. Michael and N. Nikiforakis. A multi-physics methodology for the simulation of reactive flow and elastoplastic structural response. Journal of Computational Physics, 367(??):1–27, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301979.

Musharbash:2018:DDO

[MN18c] Eleonora Musharbash and Fabio Nobile. Dual dynamically orthogonal approximation of incompressible Navier Stokes equations with random boundary conditions. Journal of Computational Physics, 354(??):135–162, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307349.

Martin:2019:OGA

[MNC19] J. Ezequiel Martin, Ralph W. Noack, and Pablo M. Carrica. Overset grid assembly approach for scalable computational fluid dynamics with body motions. Journal of Computational Physics, 390(??):297–305, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302426.

Malgarinos:2015:CLA

[MNG15a] Ilias Malgarinos, Nikolaos Nikolopoulos, and Manolis Gavaises. Coupling a local adaptive grid refinement technique with an interface sharpening scheme for the simulation of two-phase flow and free-surface flows using VOF methodology. Journal of Computational Physics, 300(??):732–753, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005215.

Marras:2015:SHO

[MNG15b] Simone Marras, Murtazo Nazarov, and Francis X. Giraldo. Stabilized high-order Galerkin methods based on a parameterfree dynamic SGS model for LES. Journal of Computational Physics, 301(??):77–101, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004799.

Mattei:2017:FIP

[MNO^{+}17] S. Mattei, K. Nishida, M. Onai, J. Lettry, M. Q. Tran, and A. Hatayama. A fully-implicit Particle-In-Cell Monte Carlo collision code for the simulation of inductively coupled plasmas. Journal of Computational Physics, 350(??):891–906, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306745.

Marques:2017:HOS

[MNR17] Alexandre Noll Marques, Jean-Christophe Nave, and Rodolfo Ruben Rosales. High order solution of Poisson problems with piecewise constant coefficients and interface jumps. Journal of Computational Physics, 335(??):497–515, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300451.

Marques:2019:IJC

[MNR19] Alexandre Noll Marques, Jean-Christophe Nave, and Rodolfo Ruben Rosales. Imposing jump conditions on nonconforming interfaces for the Correction Function Method: a least squares approach. Journal of Computational Physics, 397(??):Article 108869, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305601.

Muhr:2019:SAA

[MNW19] Markus Muhr, Vanja Nikolić, and Barbara Wohlmuth. Self-adaptive absorbing boundary conditions for quasilinear acoustic wave propagation. Journal of Computational Physics, 388(??):279–299, July 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302086.

Mattsson:2018:IPM

[MO18a] Ken Mattsson and Pelle Olsson. An improved projection method. Journal of Computational Physics, 372(??):349–372, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911830408X.

Mattsson:2018:CDN

[MO18b] Ken Mattsson and Ossian O’Reilly. Compatible diagonalnorm staggered and upwind SBP operators. Journal of Computational Physics, 352(??):52–75, January 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307040.

Martins:2015:NSK

[MOAA15] F. P. Martins, C. M. Oishi, A. M. Afonso, and M. A. Alves. A numerical study of the kernel-conformation transformation for transient viscoelastic fluid flows. Journal of Computational Physics, 302(??):653–673, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005653.

Mocz:2017:CBC

[Moc17] Philip Mocz. Correspondence between constrained transport and vector potential methods for magnetohydrodynamics. Journal of Computational Physics, 328(??):221–233, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304806.

Mohammadi:2015:WBC

[Moh15] Fakhrodin Mohammadi. A wavelet-based computational method for solving stochastic Itô–Volterra integral equations. Journal of Computational Physics, 298(??):254–265, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003915.

Monaghan:2019:ISE

[Mon19] J. J. Monaghan. On the integration of the SPH equations for a highly viscous fluid. Journal of Computational Physics, 394(??):166–176, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930350X.

Moore:2017:FCM

[Moo17] M. Nicholas J. Moore. A fast Chebyshev method for simulating flexible-wing propulsion. Journal of Computational Physics, 345(??):792–817, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730445X.

Musehane:2018:MSS

[MOR18] Ndivhuwo M. Musehane, Oliver F. Oxtoby, and B. Daya Reddy. Multi-scale simulation of droplet-droplet interaction and coalescence. Journal of Computational Physics, 373(??):924–939, November 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304881.

Mentrelli:2015:FPA

[MP15a] Andrea Mentrelli and Gianni Pagnini. Front propagation in anomalous diffusive media governed by time-fractional diffusion. Journal of Computational Physics, 293(??):427–441, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008237.

Minjeaud:2015:HOA

[MP15b] Sébastian Minjeaud and Richard Pasquetti. High order approximation of a tokamak edge plasma transport minimal model with Bohm boundary conditions. Journal of Computational Physics, 285(??):84–87, March 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008729.

Minjeaud:2016:FSE

[MP16] Sebastian Minjeaud and Richard Pasquetti. Fourier-spectral element approximation of the ion-electron Braginskii system with application to tokamak edge plasma in divertor configuration. Journal of Computational Physics, 321(??):492–511, September 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302017.

Meldi:2017:ROM

[MP17] M. Meldi and A. Poux. A reduced order model based on Kalman filtering for sequential data assimilation of turbulent flows. Journal of Computational Physics, 347(??):207–234, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304965.

Merrill:2019:MOG

[MP19] Brandon E. Merrill and Yulia T. Peet. Moving overlapping grid methodology of spectral accuracy for incompressible flow solutions around rigid bodies in motion. Journal of Computational Physics, 390(??):121–151, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301068.

Merrill:2016:SAM

[MPFL16] Brandon E. Merrill, Yulia T. Peet, Paul F. Fischer, and James W. Lottes. A spectrally accurate method for overlapping grid solution of incompressible Navier–Stokes equations. Journal of Computational Physics, 307(??):60–93, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008050.

Muscat:2019:CIE

[MPMB19] Laurent Muscat, Guillaume Puigt, Marc Montagnac, and Pierre Brenner. A coupled implicit-explicit time integration method for compressible unsteady flows. Journal of Computational Physics, 398(??):Article 108883, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305819.

Miki:2015:SVN

[MPP15] Kenji Miki, Marco Panesi, and Serge Prudhomme. Systematic validation of non-equilibrium thermochemical models using Bayesian inference. Journal of Computational Physics, 298(??):125–144, October 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003393.

Mazhar:2018:DVA

[MPR^{+}18] H. Mazhar, A. Pazouki, M. Rakhsha, P. Jayakumar, and D. Negrut. A differential variational approach for handling fluidsolid interaction problems via smoothed particle hydrodynamics. Journal of Computational Physics, 371(??):92–110, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911830305X.

Marchetti:2016:HEH

[MPT16] Luca Marchetti, Corrado Priami, and Vo Hong Thanh. HRSSA — efficient hybrid stochastic simulation for spatially homogeneous biochemical reaction networks. Journal of Computational Physics, 317(??):301–317, July 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301115.

Mashayekhi:2016:NSD

[MR16a] S. Mashayekhi and M. Razzaghi. Numerical solution of distributed order fractional differential equations by hybrid functions. Journal of Computational Physics, 315(??):169–181, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630016X.

Murali:2016:NMB

[MR16b] Avinaash Murali and R. G. Rajagopalan. A new mixed basis Navier–Stokes formulation for incompressible flows over complex geometries. Journal of Computational Physics, 307 (??):378–400, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115008311.

Mehlmann:2017:FEM

[MR17] C. Mehlmann and T. Richter. A finite element multigridframework to solve the sea ice momentum equation. Journal of Computational Physics, 348(??):847–861, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305752.

Machac:2016:EDS

[MRA16] David Machac, Peter Reichert, and Carlo Albert. Emulation of dynamic simulators with application to hydrology. Journal of Computational Physics, 313(??):352–366, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001169.

Mirkov:2015:IFV

[MRK15] Nikola Mirkov, Boško Rašuo, and Saša Kenjere. On the improved finite volume procedure for simulation of turbulent flows over real complex terrains. Journal of Computational Physics, 287(??):18–45, 2015. CODEN JCTPAH. ISSN

0021-9991. URL http://www.sciencedirect.com/science/
article/pii/S0021999115000625.

Maginot:2016:HOS

[MRM16] Peter G. Maginot, Jean C. Ragusa, and Jim E. Morel. High-order solution methods for grey discrete ordinates thermal radiative transfer. Journal of Computational Physics, 327(??):719–746, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304764.

Mazaheri:2016:FOH

[MRN16] Alireza Mazaheri, Mario Ricchiuto, and Hiroaki Nishikawa. A first-order hyperbolic system approach for dispersion. Journal of Computational Physics, 321(??):593–605, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302261.

Manoli:2015:IPF

[MRP^{+}15] Gabriele Manoli, Matteo Rossi, Damiano Pasetto, Rita Deiana, Stefano Ferraris, Giorgio Cassiani, and Mario Putti. An iterative particle filter approach for coupled hydro-geophysical inversion of a controlled infiltration experiment. Journal of Computational Physics, 283(??):37–51, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007979.

Manzanero:2018:BRS

[MRRRF18] Juan Manzanero, Andrés M. Rueda-Ramírez, Gonzalo Rubio, and Esteban Ferrer. The Bassi Rebay 1 scheme is a special case of the Symmetric Interior Penalty formulation for discontinuous Galerkin discretisations with Gauss–Lobatto points. Journal of Computational Physics, 363(??):1–10, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301177.

Mohebujjaman:2017:EBM

[MRXI17] Muhammad Mohebujjaman, Leo G. Rebholz, Xuping Xie, and Traian Iliescu. Energy balance and mass conservation in reduced order models of fluid flows. Journal of Computational Physics, 346(??):262–277, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304643.

Maurya:2019:HOO

[MRY19] Praveen K. Maurya, Manoj K. Rajpoot, and Vivek S. Yadav. Higher-order optimized hybrid Robert–Asselin type time filters. Journal of Computational Physics, 399 (??):Article 108941, December 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306461.

Michaud-Rioux:2016:RRS

[MRZG16] Vincent Michaud-Rioux, Lei Zhang, and Hong Guo. RESCU: a real space electronic structure method. Journal of Computational Physics, 307(??):593–613, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008335.

Maire:2015:PRR

[MS15a] Sylvain Maire and Martin Simon. A partially reflecting random walk on spheres algorithm for electrical impedance tomography. Journal of Computational Physics, 303(??):413–430, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006671.

Mandal:2015:GMC

[MS15b] J. C. Mandal and V. Sharma. A genuinely multidimensional convective pressure flux split Riemann solver for Euler equations. Journal of Computational Physics, 297(??):669–688, September 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003733.

Mattsson:2015:HFN

[MS15c] Ken Mattsson and Vidar Stiernström. High-fidelity numerical simulation of the dynamic beam equation. Journal of Computational Physics, 286(??):194–213, April 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000522.

Mao:2016:ESG

[MS16a] Zhiping Mao and Jie Shen. Efficient spectral-Galerkin methods for fractional partial differential equations with variable coefficients. Journal of Computational Physics, 307(??):243–261, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007871.

Margheri:2016:HAA

[MS16b] Luca Margheri and Pierre Sagaut. A hybrid anchored-ANOVA — POD/kriging method for uncertainty quantification in unsteady high-fidelity CFD simulations. Journal of Computational Physics, 324(??):137–173, November 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630331X.

Mitchell:2017:GTI

[MS17] William H. Mitchell and Saverio E. Spagnolie. A generalized traction integral equation for Stokes flow, with applications to near-wall particle mobility and viscous erosion. Journal of Computational Physics, 333(??):462–482, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116307070.

Main:2018:SBMa

[MS18a] A. Main and G. Scovazzi. The shifted boundary method for embedded domain computations. Part I: Poisson and Stokes problems. Journal of Computational Physics, 372(??):972–995, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307799.

Main:2018:SBMb

[MS18b] A. Main and G. Scovazzi. The shifted boundary method for embedded domain computations. Part II: Linear advectiondiffusion and incompressible Navier–Stokes equations. Journal of Computational Physics, 372(??):996–1026, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300330.

Minion:2018:HOT

[MS18c] M. L. Minion and R. I. Saye. Higher-order temporal integration for the incompressible Navier–Stokes equations in bounded domains. Journal of Computational Physics, 375(??):797–822, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305874.

Mizerova:2018:CSF

[MS18d] Hana Mizerová and Bangwei She. A conservative scheme for the Fokker–Planck equation with applications to viscoelastic polymeric fluids. Journal of Computational Physics, 374 (??):941–953, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118305382.

Mumtaz:2019:EHO

[MSA19] Faisal Mumtaz, Hamed Saidaoui, and Fahhad H. Alharbi. Efficient high order method for differential equations in unbounded domains using generalized coordinate transformation. Journal of Computational Physics, 381(??):275–289, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300117.

Moura:2016:LEA

[MSB^{+}16] R. C. Moura, A. F. C. Silva, E. D. V. Bigarella, A. L. Fazenda, and M. A. Ortega. Lyapunov exponents and adaptive mesh refinement for high-speed flows using a discontinuous Galerkin scheme. Journal of Computational Physics, 319(??):9–27, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301565.

Meierbachtol:2017:EPC

[MSD^{+}17] Collin S. Meierbachtol, Daniil Svyatskiy, Gian Luca Delzanno, Louis J. Vernon, and J. David Moulton. An electrostatic Particle-In-Cell code on multi-block structured meshes. Journal of Computational Physics, 350(??):796–823, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306757.

Myatt:2019:LDW

[MSF^{+}19] Jason F. Myatt, John G. Shaw, Russell K. Follett, Dana H. Edgell, Dustin H. Froula, John P. Palastro, and Valeri N.

Goncharov. LPSE: a 3-d wave-based model of cross-beam energy transfer in laser-irradiated plasmas. Journal of Computational Physics, 399(??):Article 108916, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306217.

Maddix:2018:NAD

[MSG18a] Danielle C. Maddix, Luiz Sampaio, and Margot Gerritsen. Numerical artifacts in the discontinuous Generalized Porous Medium Equation: How to avoid spurious temporal oscillations. Journal of Computational Physics, 368(??):277–298, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830278X.

Maddix:2018:NAG

[MSG18b] Danielle C. Maddix, Luiz Sampaio, and Margot Gerritsen. Numerical artifacts in the Generalized Porous Medium Equation: Why harmonic averaging itself is not to blame. Journal of Computational Physics, 361(??):280–298, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830086X.

Mulloth:2015:HAS

[MSH^{+}15] Akhil Mulloth, Nilesh Sawant, Ijlal Haider, Nidhi Sharma, and Tapan K. Sengupta. High accuracy solution of bidirectional wave propagation in continuum mechanics. Journal of Computational Physics, 298(??):209–236, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003745.

Mabuza:2018:LBP

[MSK18] Sibusiso Mabuza, John N. Shadid, and Dmitri Kuzmin. Local bounds preserving stabilization for continuous Galerkin discretization of hyperbolic systems. Journal of Computational Physics, 361(??):82–110, May 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300585.

Moura:2015:LDD

[MSP15] R. C. Moura, S. J. Sherwin, and J. Peiró. Linear dispersiondiffusion analysis and its application to under-resolved turbulence simulations using discontinuous Galerkin spectral/hp methods. Journal of Computational Physics, 298(??):695–710, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004179.

Moura:2016:EAS

[MSP16] R. C. Moura, S. J. Sherwin, and J. Peiró. Eigensolution analysis of spectral/hp continuous Galerkin approximations to advection-diffusion problems: Insights into spectral vanishing viscosity. Journal of Computational Physics, 307(??):401–422, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008256.

Mazaheri:2019:BCW

[MSP19] Alireza Mazaheri, Chi-Wang Shu, and Vincent Perrier. Bounded and compact weighted essentially nonoscillatory limiters for discontinuous Galerkin schemes: Triangular elements. Journal of Computational Physics, 395(??):461–488, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304310.

Mishra:2016:MLM

[MSS16] S. Mishra, Ch. Schwab, and J. Sukys. Multi-level Monte Carlo finite volume methods for uncertainty quantification of acoustic wave propagation in random heterogeneous layered medium. Journal of Computational Physics, 312(??):192–217, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000681.

Matveev:2015:FNM

[MST15] S. A. Matveev, A. P. Smirnov, and E. E. Tyrtyshnikov. A fast numerical method for the Cauchy problem for the Smoluchowski equation. Journal of Computational Physics, 282(??):23–32, February 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114007554.

Mittal:2016:CMC

[MSV^{+}16] Rajat Mittal, Jung Hee Seo, Vijay Vedula, Young J. Choi, Hang Liu, H. Howie Huang, Saurabh Jain, Laurent Younes, Theodore Abraham, and Richard T. George. Computational modeling of cardiac hemodynamics: Current status and future outlook. Journal of Computational Physics, 305(??):1065–1082, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115007627.

Melvin:2017:WDP

[MT17] Thomas Melvin and John Thuburn. Wave dispersion properties of compound finite elements. Journal of Computational Physics, 338(??):68–90, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301183.

Mehmani:2018:MCP

[MT18] Yashar Mehmani and Hamdi A. Tchelepi. Multiscale computation of pore-scale fluid dynamics: Single-phase flow. Journal of Computational Physics, 375(??):1469–1487, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830576X.

Mehmani:2019:MFT

[MT19] Yashar Mehmani and Hamdi A. Tchelepi. Multiscale formulation of two-phase flow at the pore scale. Journal of Computational Physics, 389(??):164–188, July 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302189.

Muralikrishnan:2018:IIH

[MTBT18] Sriramkrishnan Muralikrishnan, Minh-Binh Tran, and Tan Bui-Thanh. An improved iterative HDG approach for partial differential equations. Journal of Computational Physics, 367(??):295–321, August 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302584.

Moon:2015:CPC

[MTD15] Haksu Moon, Fernando L. Teixeira, and Burkay Donderici. Computation of potentials from current electrodes in cylindrically stratified media: a stable, rescaled semi-analytical formulation. Journal of Computational Physics, 280(??):692–709, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006974.

Moncorge:2017:MSF

[MTJ17] A. Moncorgé, H. A. Tchelepi, and P. Jenny. Modified sequential fully implicit scheme for compositional flow simulation. Journal of Computational Physics, 337(??):98–115, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301250.

Moncorge:2018:SFI

[MTJ18] A. Moncorgé, H. A. Tchelepi, and P. Jenny. Sequential fully implicit formulation for compositional simulation using natural variables. Journal of Computational Physics, 371(??):690–711, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303711.

Miyauchi:2015:NMM

[MTK15] Suguru Miyauchi, Shintaro Takeuchi, and Takeo Kajishima. A numerical method for mass transfer by a thin moving membrane with selective permeabilities. Journal of Computational Physics, 284(??):490–504, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008717.

Morii:2016:EFR

[MTK^{+}16] Youhi Morii, Hiroshi Terashima, Mitsuo Koshi, Taro Shimizu, and Eiji Shima. ERENA: a fast and robust Jacobian-free integration method for ordinary differential equations of chemical kinetics. Journal of Computational Physics, 322(??):547–558, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116302479.

Miyauchi:2017:NMI

[MTK17] Suguru Miyauchi, Shintaro Takeuchi, and Takeo Kajishima. A numerical method for interaction problems between fluid and membranes with arbitrary permeability for fluid. Journal of Computational Physics, 345(??):33–57, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303753.

Montgomery:2017:UMZ

[MTL^{+}17] Robert Montgomery, Carlos Tomé, Wenfeng Liu, Alankar Alankar, Gopinath Subramanian, and Christopher Stanek. Use of multiscale zirconium alloy deformation models in nuclear fuel behavior analysis. Journal of Computational Physics, 328(??):278–300, January 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304727.

Morzfeld:2019:LMS

[MTM19] M. Morzfeld, X. T. Tong, and Y. M. Marzouk. Localization for MCMC: sampling high-dimensional posterior distributions with local structure. Journal of Computational Physics, 380(??):1–28, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118308040.

Medvinsky:2019:DIH

[MTT19] M. Medvinsky, S. Tsynkov, and E. Turkel. Direct implementation of high order BGT artificial boundary conditions. Journal of Computational Physics, 376(??):98–128, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306429.

Masson:2016:CCL

[MTZ16] R. Masson, L. Trenty, and Y. Zhang. Coupling compositional liquid gas Darcy and free gas flows at porous and free-flow domains interface. Journal of Computational Physics, 321(??):708–728, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302285.

Mueller:2018:NTW

[Mue18] A. Mueller. Novel two-way artificial boundary condition for 2D vertical water wave propagation modelled with Radial-Basis-Function Collocation Method. Journal of Computational Physics, 359(??):283–295, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300275.

Martin-Vaquero:2016:ESE

[MVK16] J. Martín-Vaquero and B. Kleefeld. Extrapolated stabilized explicit Runge–Kutta methods. Journal of Computational Physics, 326(??):141–155, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303990.

Maiolo:2015:WBF

[MVKD15] M. Maiolo, A. Vancheri, R. Krause, and A. Danani. Wavelets as basis functions to represent the coarse-graining potential in multiscale coarse graining approach. Journal of Computational Physics, 300(??):592–604, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004842.

Mishra:2016:CSL

[MVZ16] Nachiketa Mishra, Jaroslav Vondrejc, and Jan Zeman. A comparative study on low-memory iterative solvers for FFTbased homogenization of periodic media. Journal of Computational Physics, 321(??):151–168, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301863.

Moyles:2015:NFS

[MW15] Iain Moyles and Brian Wetton. A numerical framework for singular limits of a class of reaction diffusion problems. Journal of Computational Physics, 300(??):308–326, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005069.

Mattsson:2016:HFN

[MW16a] Ken Mattsson and Jonatan Werpers. High-fidelity numerical simulation of solitons in the nerve axon. Journal of Computational Physics, 305(??):793–816, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500741X.

Mei:2016:CAO

[MW16b] Lijie Mei and Xinyuan Wu. The construction of arbitrary order ERKN methods based on group theory for solving oscillatory Hamiltonian systems with applications. Journal of Computational Physics, 323(??):171–190, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630328X.

Mei:2017:SER

[MW17a] Lijie Mei and Xinyuan Wu. Symplectic exponential Runge– Kutta methods for solving nonlinear Hamiltonian systems. Journal of Computational Physics, 338(??):567–584, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730205X.

Morgan:2017:LSM

[MW17b] Nathaniel R. Morgan and Jacob I. Waltz. 3D level set methods for evolving fronts on tetrahedral meshes with adaptive mesh refinement. Journal of Computational Physics, 336(??):492– 512, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301237.

Morgan:2015:GLP

[MWB^{+}15a] Nathaniel R. Morgan, Jacob I. Waltz, Donald E. Burton, Marc R. Charest, Thomas R. Canfield, and John G. Wohlbier. A Godunov-like point-centered essentially Lagrangian hydrodynamic approach. Journal of Computational Physics, 281(??):614–652, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400730X.

Morgan:2015:PCA

[MWB^{+}15b] Nathaniel R. Morgan, Jacob I. Waltz, Donald E. Burton, Marc R. Charest, Thomas R. Canfield, and John G. Wohlbier. A point-centered arbitrary Lagrangian Eulerian hydrodynamic approach for tetrahedral meshes. Journal of Computational Physics, 290(??):239–273, June 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000947.

Maginnis:2016:VRS

[MWD16] P. A. Maginnis, M. West, and G. E. Dullerud. Variancereduced simulation of lattice discrete-time Markov chains with applications in reaction networks. Journal of Computational Physics, 322(??):400–414, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302443.

Mu:2016:NWG

[MWYZ16] Lin Mu, Junping Wang, Xiu Ye, and Shan Zhao. A new weak Galerkin finite element method for elliptic interface problems. Journal of Computational Physics, 325(??):157–173, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303746.

Mons:2019:KEE

[MWZ19] Vincent Mons, Qi Wang, and Tamer A. Zaki. Kriging-enhanced ensemble variational data assimilation for scalar-source identification in turbulent environments. Journal of Computational Physics, 398(??):Article 108856, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305406.

Metti:2016:ESD

[MXL16] Maximilian S. Metti, Jinchao Xu, and Chun Liu. Energetically stable discretizations for charge transport and electrokinetic models. Journal of Computational Physics, 306(??):1–18, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007305.

Mukherjee:2015:DEB

[MZ15] Debanjan Mukherjee and Tarek I. Zohdi. A discrete element based simulation framework to investigate particulate spray deposition processes. Journal of Computational Physics, 290(??):298–317, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001047.

Main:2017:EFM

[MZAF17] Alex Main, Xianyi Zeng, Philip Avery, and Charbel Farhat. An enhanced FIVER method for multi-material flow problems with second-order convergence rate. Journal of Computational Physics, 329(??):141–172, January 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305265.

Matveev:2016:TTV

[MZTS16] Sergey A. Matveev, Dmitry A. Zheltkov, Eugene E. Tyrtyshnikov, and Alexander P. Smirnov. Tensor train versus Monte Carlo for the multicomponent Smoluchowski coagulation equation. Journal of Computational Physics, 316(??):164–179, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300675.

Nielsen:2018:CAA

[NBH18] Allan S. Nielsen, Gilles Brunner, and Jan S. Hesthaven. Communication-aware adaptive Parareal with application to a nonlinear hyperbolic system of partial differential equations.

Journal of Computational Physics, 371(??):483–505, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302894.

Niethammer:2019:EVF

[NBMB19] Matthias Niethammer, Günter Brenn, Holger Marschall, and Dieter Bothe. An extended volume of fluid method and its application to single bubbles rising in a viscoelastic liquid. Journal of Computational Physics, 387(??):326–355, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911930138X.

Na:2019:FET

[NBT19] Dong-Yeop Na, Ben-Hur V. Borges, and Fernando L. Teixeira. Finite element time-domain body-of-revolution Maxwell solver based on discrete exterior calculus. Journal of Computational Physics, 376(??):249–275, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306247.

Nestola:2019:IBM

[NBZ^{+}19] Maria Giuseppina Chiara Nestola, Barna Becsek, Hadi Zolfaghari, Patrick Zulian, Dario De Marinis, Rolf Krause, and Dominik Obrist. An immersed boundary method for fluidstructure interaction based on variational transfer. Journal of Computational Physics, 398(??):Article 108884, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119305820.

Neic:2017:ECE

[NCP^{+}17] Aurel Neic, Fernando O. Campos, Anton J. Prassl, Steven A. Niederer, Martin J. Bishop, Edward J. Vigmond, and Gernot Plank. Efficient computation of electrograms and ECGs in human whole heart simulations using a reaction–eikonal model. Journal of Computational Physics, 346(??):191–211, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304655.

Nigro:2017:SDT

[NDCB17] A. Nigro, C. De Bartolo, A. Crivellini, and F. Bassi. Second derivative time integration methods for discontinuous Galerkin solutions of unsteady compressible flows. Journal of Computational Physics, 350(??):493–517, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306307.

Nath:2019:IPC

[NDH19] Kamaljyoti Nath, Anjan Dutta, and Budhaditya Hazra. An iterative polynomial chaos approach for solution of structural mechanics problem with Gaussian material property. Journal of Computational Physics, 390(??):425–451, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302475.

Naddei:2019:CRI

[NdlLPCC19] Fabio Naddei, Marta de la Llave Plata, Vincent Couaillier, and Frédéric Coquel. A comparison of refinement indicators for p-adaptive simulations of steady and unsteady flows using discontinuous Galerkin methods. Journal of Computational Physics, 376(??):508–533, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306478.

Nonomura:2017:CFD

[NF17] Taku Nonomura and Kozo Fujii. Characteristic finite-difference WENO scheme for multicomponent compressible fluid analysis: Overestimated quasi-conservative formulation maintaining equilibriums of velocity, pressure, and temperature. Journal of Computational Physics, 340(??):358–388, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301572.

Ndanou:2015:MSM

[NFG15] S. Ndanou, N. Favrie, and S. Gavrilyuk. Multi-solid and multi-fluid diffuse interface model: Applications to dynamic fracture and fragmentation. Journal of Computational Physics, 295(??):523–555, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002806.

Nordstrom:2017:RBC

[NG17] Jan Nordström and Fatemeh Ghasemi. On the relation between conservation and dual consistency for summation-by-parts schemes. Journal of Computational Physics, 344(??): 437–439, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303601.

See corrigendum [NG18].

Nordstrom:2018:CT

[NG18] Jan Nordström and Fatemeh Ghasemi. Corrigendum to “On the relation between conservation and dual consistency for summation-by-parts schemes” [J. Comput. Phys. 344 (2017) 437–439]. Journal of Computational Physics, 360(??):247, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301293. See [NG17].

Nangia:2019:RIN

[NGPB19] Nishant Nangia, Boyce E. Griffith, Neelesh A. Patankar, and Amneet Pal Singh Bhalla. A robust incompressible Navier– Stokes solver for high density ratio multiphase flows. Journal of Computational Physics, 390(??):548–594, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302256.

Niyonzima:2016:WRC

[NGS16] I. Niyonzima, C. Geuzaine, and S. Schöps. Waveform relaxation for the computational homogenization of multiscale magnetoquasistatic problems. Journal of Computational Physics, 327(??):416–433, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304235.

Nold:2017:PMD

[NGY^{+}17] Andreas Nold, Benjamin D. Goddard, Peter Yatsyshin, Nikos Savva, and Serafim Kalliadasis. Pseudospectral methods for density functional theory in bounded and unbounded domains. Journal of Computational Physics, 334(??):639–664, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306775.

Ngo:2017:SMM

[NH17] Cuong Ngo and Weizhang Huang. A study on moving mesh finite element solution of the porous medium equation. Journal of Computational Physics, 331(??):357–380, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306374.

Nichols:2018:SDT

[NHA18] J. A. Nichols, B. I. Henry, and C. N. Angstmann. Subdiffusive discrete time random walks via Monte Carlo and subordination. Journal of Computational Physics, 372(??):373–384, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304224.

Nath:2017:SPC

[NHM17] Paromita Nath, Zhen Hu, and Sankaran Mahadevan. Sensor placement for calibration of spatially varying model parameters. Journal of Computational Physics, 343(??):150–169, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303042.

Nishikawa:2015:APB

[Nis15] Hiroaki Nishikawa. Accuracy-preserving boundary flux quadrature for finite-volume discretization on unstructured grids. Journal of Computational Physics, 281(??):518–555, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007153.

Nishikawa:2018:HDS

[Nis18a] Hiroaki Nishikawa. From hyperbolic diffusion scheme to gradient method: Implicit Green–Gauss gradients for unstructured grids. Journal of Computational Physics, 372(??):126–160, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303978.

Nishikawa:2018:HMD

[Nis18b] Hiroaki Nishikawa. On hyperbolic method for diffusion with discontinuous coefficients. Journal of Computational Physics, 367(??):102–108, August 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302444.

Nishikawa:2019:EGS

[Nis19a] Hiroaki Nishikawa. Efficient gradient stencils for robust implicit finite-volume solver convergence on distorted grids. Journal of Computational Physics, 386(??):486–501, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301433.

Nishikawa:2019:LSE

[Nis19b] Hiroaki Nishikawa. On large start-up error of BDF2. Journal of Computational Physics, 392(??):456–461, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303237.

Niu:2016:CTF

[Niu16] Yang-Yao Niu. Computations of two-fluid models based on a simple and robust hybrid primitive variable Riemann solver with AUSMD. Journal of Computational Physics, 308(??):389–410, March 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115008645.

Nelson:2015:DFL

[NJ15] Daniel A. Nelson and Gustaaf B. Jacobs. DG–FTLE: Lagrangian coherent structures with high-order discontinuous-Galerkin methods. Journal of Computational Physics, 295 (??):65–86, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115001953.

Nguyen:2018:PUF

[NJHL18] Van-Dang Nguyen, Johan Jansson, Johan Hoffman, and Jing-Rebecca Li. A partition of unity finite element method for computational diffusion MRI. Journal of Computational Physics, 375(??):271–290, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305709.

Nadiga:2019:LBA

[NJL19] Balasubramanya Nadiga, Chiyu Jiang, and Daniel Livescu. Leveraging Bayesian analysis to improve accuracy of approximate models. Journal of Computational Physics, 394 (??):280–297, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303407.

Nangia:2017:MCV

[NJPB17] Nishant Nangia, Hans Johansen, Neelesh A. Patankar, and Amneet Pal Singh Bhalla. A moving control volume approach to computing hydrodynamic forces and torques on immersed bodies. Journal of Computational Physics, 347(??):437–462, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305016.

Nguyen:2017:NSC

[NKN^{+}17] Dinh-Liem Nguyen, Michael V. Klibanov, Loc H. Nguyen, Aleksandr E. Kolesov, Michael A. Fiddy, and Hui Liu. Numerical solution of a coefficient inverse problem with multifrequency experimental raw data by a globally convergent algorithm. Journal of Computational Physics, 345(??):17– 32, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303844.

Nguyen:2015:CSR

[NL15] Hoang-Ngan Nguyen and Karin Leiderman. Computation of the singular and regularized image systems for doublyperiodic Stokes flow in the presence of a wall. Journal of Computational Physics, 297(??):442–461, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003642.

Nishikawa:2017:APS

[NL17] Hiroaki Nishikawa and Yi Liu. Accuracy-preserving source term quadrature for third-order edge-based discretization. Journal of Computational Physics, 344(??):595–622, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303637.

Nishikawa:2018:HAD

[NL18a] Hiroaki Nishikawa and Yi Liu. Hyperbolic advection-diffusion schemes for high-Reynolds-number boundary-layer problems. Journal of Computational Physics, 352(??):23–51, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730699X.

Nordstrom:2018:WPS

[NL18b] Jan Nordström and Viktor Linders. Well-posed and stable transmission problems. Journal of Computational Physics, 364(??):95–110, July 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301475.

Nguyen:2016:SPB

[NLFM16] T. T. Nguyen, F. Laurent, R. O. Fox, and M. Massot. Solution of population balance equations in applications with fine particles: Mathematical modeling and numerical schemes. Journal of Computational Physics, 325(??):129–156, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303667.

Niemi:2015:DMS

[NLK^{+}15] Esa Niemi, Matti Lassas, Aki Kallonen, Lauri Harhanen, Keijo Hämäläinen, and Samuli Siltanen. Dynamic multisource X-ray tomography using a spacetime level set method. Journal of Computational Physics, 291(??):218–237, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001527.

Noumir:2015:FML

[NLL^{+}15] Y. Noumir, A. Le Guilcher, N. Lardjane, R. Monneau, and A. Sarrazin. A fast-marching like algorithm for geometrical shock dynamics. Journal of Computational Physics, 284(??):206–229, March 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114008365.

Nourgaliev:2016:FIO

[NLW^{+}16] R. Nourgaliev, H. Luo, B. Weston, A. Anderson, S. Schofield, T. Dunn, and J.-P. Delplanque. Fully-implicit orthogonal reconstructed discontinuous Galerkin method for fluid dynamics with phase change. Journal of Computational Physics, 305(??):964–996, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911500738X.

Negri:2015:EMR

[NMA15] Federico Negri, Andrea Manzoni, and David Amsallem. Efficient model reduction of parametrized systems by matrix discrete empirical interpolation. Journal of Computational Physics, 303(??):431–454, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006543.

Nazari:2015:HOL

[NMC15] Farshid Nazari, Abdolmajid Mohammadian, and Martin Charron. High-order low-dissipation low-dispersion diagonally implicit Runge–Kutta schemes. Journal of Computational Physics, 286(??):38–48, April 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000248.

Navas-Montilla:2019:DAU

[NMJFM19] A. Navas-Montilla, C. Juez, M. J. Franca, and J. Murillo. Depth-averaged unsteady RANS simulation of resonant shallow flows in lateral cavities using augmented WENO–ADER schemes. Journal of Computational Physics, 395(??):511–536, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304450.

Navas-Montilla:2015:EBN

[NMM15] A. Navas-Montilla and J. Murillo. Energy balanced numerical schemes with very high order. the Augmented Roe Flux ADER scheme. application to the shallow water equations. Journal of Computational Physics, 290(??):188–218, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001217.

Navas-Montilla:2016:AEE

[NMM16] A. Navas-Montilla and J. Murillo. Asymptotically and exactly energy balanced augmented flux-ADER schemes with application to hyperbolic conservation laws with geometric source terms. Journal of Computational Physics, 317(??):108–147, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301024.

Navas-Montilla:2017:ONS

[NMM17] A. Navas-Montilla and J. Murillo. Overcoming numerical shockwave anomalies using energy balanced numerical schemes. Application to the Shallow Water Equations with discontinuous topography. Journal of Computational Physics, 340(??):575–616, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302589.

Navas-Montilla:2018:WBA

[NMM18] A. Navas-Montilla and J. Murillo. 2D well-balanced augmented ADER schemes for the shallow water equations with bed elevation and extension to the rotating frame. Journal of Computational Physics, 372(??):316–348, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304170.

Navas-Montilla:2019:IRS

[NMM19] A. Navas-Montilla and J. Murillo. Improved Riemann solvers for an accurate resolution of 1D and 2D shock profiles with application to hydraulic jumps. Journal of Computational Physics, 378(??):445–476, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307496.

Nikkar:2015:FDE

[NN15a] Samira Nikkar and Jan Nordström. Fully discrete energy stable high order finite difference methods for hyperbolic problems in deforming domains. Journal of Computational Physics, 291(??):82–98, June 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000972.

Notay:2015:MPS

[NN15b] Yvan Notay and Artem Napov. A massively parallel solver for discrete Poisson-like problems. Journal of Computational Physics, 281(??):237–250, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007256.

Nordstrom:2016:HSE

[NN16] Jan Nordström and Samira Nikkar. Hyperbolic systems of equations posed on erroneous curved domains. Journal of Computational Physics, 308(??):438–442, March 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008670.

Nikkar:2017:FDS

[NN17] Samira Nikkar and Jan Nordström. A fully discrete, stable and conservative summation-by-parts formulation for deforming interfaces. Journal of Computational Physics, 339(??):500–524, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730150X.

Nishikawa:2018:DSN

[NN18] Hiroaki Nishikawa and Yoshitaka Nakashima. Dimensional scaling and numerical similarity in hyperbolic method for diffusion. Journal of Computational Physics, 355(??):121–143, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308367.

Nikkar:2019:DCS

[NN19] Samira Nikkar and Jan Nordström. A dual consistent summation-by-parts formulation for the linearized incompressible Navier–Stokes equations posed on deforming domains. Journal of Computational Physics, 376(??):322–338, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306004.

Nestler:2019:FEA

[NNV19] Michael Nestler, Ingo Nitschke, and Axel Voigt. A finite element approach for vector- and tensor-valued surface PDEs. Journal of Computational Physics, 389(??):48–61, July 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301895.

Nishikawa:2017:EHF

[NNW17] Hiroaki Nishikawa, Yoshitaka Nakashima, and Norihiko Watanabe. Effects of high-frequency damping on iterative convergence of implicit viscous solver. Journal of Computational Physics, 348(??):66–81, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305296.

Noetinger:2015:QSS

[Noe15] B. Noetinger. A quasi steady state method for solving transient Darcy flow in complex 3D fractured networks accounting for matrix to fracture flow. Journal of Computational Physics, 283(??):205–223, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008006.

Na:2017:ACC

[NOM^{+}17] Dong-Yeop Na, Yuri A. Omelchenko, Haksu Moon, Ben-Hur V. Borges, and Fernando L. Teixeira. Axisymmetric charge-conservative electromagnetic particle simulation algorithm on unstructured grids: Application to microwave vacuum electronic devices. Journal of Computational Physics, 346(??):295–317, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304618.

Norman:2015:HWL

[Nor15] Matthew R. Norman. Hermite WENO limiting for multimoment finite-volume methods using the ADER–DT time discretization for 1-D systems of conservation laws. Journal of Computational Physics, 282(??):381–396, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400775X.

Nguyen:2016:GFR

[NP16] N. C. Nguyen and J. Peraire. Gaussian functional regression for output prediction: Model assimilation and experimental design. Journal of Computational Physics, 309(??):52–68, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008542.

Nangia:2019:DIB

[NPB19] Nishant Nangia, Neelesh A. Patankar, and Amneet Pal Singh Bhalla. A DLM immersed boundary method based wavestructure interaction solver for high density ratio multiphase flows. Journal of Computational Physics, 398 (??):Article 108804, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119304887.

Nguyen:2015:CED

[NPC15] N. C. Nguyen, J. Peraire, and B. Cockburn. A class of embedded discontinuous Galerkin methods for computational fluid dynamics. Journal of Computational Physics, 302(??):674–692, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006178.

Nestler:2015:FES

[NPP15] Franziska Nestler, Michael Pippig, and Daniel Potts. Fast Ewald summation based on NFFT with mixed periodicity. Journal of Computational Physics, 285(??):280–315, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500011X.

Nguyen:2015:PBH

[NPRC15] N. C. Nguyen, J. Peraire, F. Reitich, and B. Cockburn. A phase-based hybridizable discontinuous Galerkin method for the numerical solution of the Helmholtz equation. Journal of Computational Physics, 290(??):318–335, June 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000637.

Nordstrom:2017:CSP

[NR17] Jan Nordström and Andrea A. Ruggiu. On conservation and stability properties for summation-by-parts schemes. Journal of Computational Physics, 344(??):451–464, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303716.

Nouveau:2019:HOG

[NRS19] L. Nouveau, M. Ricchiuto, and G. Scovazzi. High-order gradients with the shifted boundary method: an embedded enriched mixed formulation for elliptic PDEs. Journal of Computational Physics, 398(??):Article 108898, December 1, 2019.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305960.

Nazockdast:2017:FPS

[NRZS17] Ehssan Nazockdast, Abtin Rahimian, Denis Zorin, and Michael Shelley. A fast platform for simulating semi-flexible fiber suspensions applied to cell mechanics. Journal of Computational Physics, 329(??):173–209, January 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305241.

Nagel:2016:SLE

[NS16] Joseph B. Nagel and Bruno Sudret. Spectral likelihood expansions for Bayesian inference. Journal of Computational Physics, 309(??):267–294, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008669.

Nguyen:2019:MFE

[NS19a] Lam H. Nguyen and Dominik Schillinger. The multiscale finite element method for nonlinear continuum localization problems at full fine-scale fidelity, illustrated through phase-field fracture and plasticity. Journal of Computational Physics, 396(??):129–160, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304668.

Nguyen:2019:RDL

[NS19b] Lam H. Nguyen and Dominik Schillinger. A residual-driven local iterative corrector scheme for the multiscale finite element method. Journal of Computational Physics, 377(??):60–88, January 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306971.

Nivarti:2015:MPA

[NSB15] Girish V. Nivarti, M. Mahdi Salehi, and W. Kendal Bushe. A mesh partitioning algorithm for preserving spatial locality in arbitrary geometries. Journal of Computational Physics, 281(??):352–364, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114007049.

Nordlund:2016:IPA

[NSK^{+}16] M. Nordlund, M. Stanic, A. K. Kuczaj, E. M. A. Frederix, and B. J. Geurts. Improved PISO algorithms for modeling density varying flow in conjugate fluid-porous domains. Journal of Computational Physics, 306(??):199–215, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007755.

Norgaard:2016:TOU

[NSL16] Sebastian Nørgaard, Ole Sigmund, and Boyan Lazarov. Topology optimization of unsteady flow problems using the lattice Boltzmann method. Journal of Computational Physics, 307(??):291–307, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115008426.

Nair:2015:VCI

[NT15] Prapanch Nair and Gaurav Tomar. Volume conservation issues in incompressible smoothed particle hydrodynamics. Journal of Computational Physics, 297(??):689–699, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003769.

Nuter:2016:SNC

[NT16] Rachel Nuter and Vladimir Tikhonchuk. Suppressing the numerical Cherenkov radiation in the Yee numerical scheme. Journal of Computational Physics, 305(??):664–676, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007366.

Ni:2019:ASA

[NT19] Angxiu Ni and Chaitanya Talnikar. Adjoint sensitivity analysis on chaotic dynamical systems by Non-Intrusive Least Squares Adjoint Shadowing (NILSAS). Journal of Computational Physics, 395(??):690–709, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304437.

Nejadmalayeri:2015:PAW

[NVBDV15] Alireza Nejadmalayeri, Alexei Vezolainen, Eric Brown-Dymkoski, and Oleg V. Vasilyev. Parallel adaptive wavelet collocation method for PDEs. Journal of Computational Physics, 298(??):237–253, October 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003629.

Nordstrom:2015:VRT

[NW15] Jan Nordström and Markus Wahlsten. Variance reduction through robust design of boundary conditions for stochastic hyperbolic systems of equations. Journal of Computational Physics, 282(??):1–22, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007499.

Ni:2017:SAC

[NW17] Angxiu Ni and Qiqi Wang. Sensitivity analysis on chaotic dynamical systems by Non-Intrusive Least Squares Shadowing (NILSS). Journal of Computational Physics, 347(??):56–77, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304783.

Nguyen:2019:RIT

[NWB19] Thanh-Tung Nguyen, Danièle Waldmann, and Tinh Quoc Bui. Role of interfacial transition zone in phase field modeling of fracture in layered heterogeneous structures. Journal of Computational Physics, 386(??):585–610, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301391.

Ni:2019:SAC

[NWFT19] Angxiu Ni, Qiqi Wang, Pablo Fernández, and Chaitanya Talnikar. Sensitivity analysis on chaotic dynamical systems by Finite Difference Non-Intrusive Least Squares Shadowing (FD-NILSS). Journal of Computational Physics, 394(??):615–631, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119304115.

Newman:2016:CAI

[NWKC16] Christopher Newman, Geoffrey Womeldorff, Dana A. Knoll, and Luis Chacón. A communication-avoiding implicit–explicit method for a free-surface ocean model. Journal of Computational Physics, 305(??):877–894, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007482.

Nochetto:2018:EML

[NWZ18] Ricardo H. Nochetto, Shawn W. Walker, and Wujun Zhang. The Ericksen model of liquid crystals with colloidal and electric effects. Journal of Computational Physics, 352(??):568–601, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306952.

Ni:2019:ESB

[NYD19] Naxian Ni, Zhiguo Yang, and Suchuan Dong. Energy-stable boundary conditions based on a quadratic form: Applications to outflow/open-boundary problems in incompressible flows. Journal of Computational Physics, 391(??):179–215, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302761.

Najafi-Yazdi:2015:HRD

[NYNYM15] A. Najafi-Yazdi, M. Najafi-Yazdi, and L. Mongeau. A high resolution differential filter for large eddy simulation: Toward explicit filtering on unstructured grids. Journal of Computational Physics, 292(??):272–286, July 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001898.

Oukili:2019:RWN

[OADN19] H. Oukili, R. Ababou, G. Debenest, and B. Noetinger. Random walks with negative particles for discontinuous diffusion and porosity. Journal of Computational Physics, 396(??):687–701, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304917.

Ozen:2017:DPC

[OB17] H. Cagan Ozen and Guillaume Bal. A dynamical polynomial chaos approach for long-time evolution of SPDEs. Journal of Computational Physics, 343(??):300–323, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303339.

Obeidat:2019:IBA

[OB19] Anas Obeidat and Stéphane P. A. Bordas. An implicit boundary approach for viscous compressible high Reynolds flows using a hybrid remeshed particle hydrodynamics method. Journal of Computational Physics, 391(??):347–364, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300993.

ONeil:2018:IEB

[OC18] Michael O’Neil and Antoine J. Cerfon. An integral equationbased numerical solver for Taylor states in toroidal geometries. Journal of Computational Physics, 359(??):263–282, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300147.

Owkes:2018:ICE

[OCSC18] Mark Owkes, Eric Cauble, Jacob Senecal, and Robert A. Currie. Importance of curvature evaluation scale for predictive simulations of dynamic gas-liquid interfaces. Journal of Computational Physics, 365(??):37–55, July 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301700.

Owkes:2015:MDH

[OD15] Mark Owkes and Olivier Desjardins. A mesh-decoupled height function method for computing interface curvature. Journal of Computational Physics, 281(??):285–300, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007189.

Owkes:2017:MMC

[OD17] Mark Owkes and Olivier Desjardins. A mass and momentum conserving unsplit semi-Lagrangian framework for simulating multiphase flows. Journal of Computational Physics, 332(??):21–46, March 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306386.

Owens:2017:OTI

[OKE17] A. R. Owens, J. Kópházi, and M. D. Eaton. Optimal trace inequality constants for interior penalty discontinuous Galerkin discretisations of elliptic operators using arbitrary elements with non-constant Jacobians. Journal of Computational Physics, 350(??):847–870, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306794.

Owens:2017:EDM

[OKWE17] A. R. Owens, J. Kópházi, J. A. Welch, and M. D. Eaton. Energy dependent mesh adaptivity of discontinuous isogeometric discrete ordinate methods with dual weighted residual error estimators. Journal of Computational Physics, 335(??):352–386, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300517.

Ouaknin:2017:FLS

[OLB^{+}17] Gaddiel Ouaknin, Nabil Laachi, Daniil Bochkov, Kris Delaney, Glenn H. Fredrickson, and Frederic Gibou. Functional level-set derivative for a polymer self consistent field theory Hamiltonian. Journal of Computational Physics, 345(??):207–223, September 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304126.

Ouaknin:2016:SCF

[OLD^{+}16] Gaddiel Ouaknin, Nabil Laachi, Kris Delaney, Glenn H. Fredrickson, and Frederic Gibou. Self-consistent field theory simulations of polymers on arbitrary domains. Journal of Computational Physics, 327(??):168–185, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304351.

Ouaknin:2018:LSS

[OLD^{+}18] Gaddiel Y. Ouaknin, Nabil Laachi, Kris Delaney, Glenn H. Fredrickson, and Frederic Gibou. Level-set strategy for inverse DSA-lithography. Journal of Computational Physics, 375(??):1159–1178, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306211.

OReilly:2017:ESH

[OLDN17] Ossian O’Reilly, Tomas Lundquist, Eric M. Dunham, and Jan Nordström. Energy stable and high-order-accurate finite difference methods on staggered grids. Journal of Computational Physics, 346(??):572–589, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304758.

Orgis:2017:BWP

[OLHD17] Thomas Orgis, Matthias Läuter, Dörthe Handorf, and Klaus Dethloff. Baroclinic waves on the β plane using low-order Discontinuous Galerkin discretisation. Journal of Computational Physics, 339(??):461–481, June 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302243.

Olson:2015:NBM

[Ols15] Nathan M. Olson. A near-boundary modification for the link bounce-back boundary condition in the lattice Boltzmann method. Journal of Computational Physics, 301(??):102–110, November 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005483.

Osnabrugge:2016:CBS

[OLV16] Gerwin Osnabrugge, Saroch Leedumrongwatthanakun, and Ivo M. Vellekoop. A convergent Born series for solving the inhomogeneous Helmholtz equation in arbitrarily large media. Journal of Computational Physics, 322(??):113–124, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116302595.

Ortigueira:2015:WFD

[OM15] Manuel D. Ortigueira and J. A. Tenreiro Machado. What is a fractional derivative? Journal of Computational Physics, 293(??):4–13, July 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005129.

Ong:2019:IBM

[OM19] Chia Rui Ong and Hiroaki Miura. Immersed boundary method with irrotational discrete delta vector for droplet simulations of large density ratio. Journal of Computational Physics, 391(??):280–302, August 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302724.

Oger:2016:SAI

[OMLdL16] G. Oger, S. Marrone, D. Le Touzé, and M. de Leffe. SPH accuracy improvement through the combination of a quasi-Lagrangian shifting transport velocity and consistent ALE formalisms. Journal of Computational Physics, 313(??):76–98, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001030.

Ostilla-Monico:2015:MRS

[OMYvdP^{+}15] R. Ostilla-Monico, Yantao Yang, E. P. van der Poel, D. Lohse, and R. Verzicco. A multiple-resolution strategy for direct numerical simulation of scalar turbulence. Journal of Computational Physics, 301(??):308–321, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005586.

Orley:2015:CEB

[ÖPHA15] Felix Örley, Vito Pasquariello, Stefan Hickel, and Nikolaus A. Adams. Cut-element based immersed boundary method for moving geometries in compressible liquid flows with cavitation. Journal of Computational Physics, 283(??):1–22, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007906.

Opper:2017:ERE

[Opp17] Manfred Opper. An estimator for the relative entropy rate of path measures for stochastic differential equations. Journal of Computational Physics, 330(??):127–133, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306076.

Oberman:2015:FSH

[OS15a] Adam M. Oberman and Tiago Salvador. Filtered schemes for Hamilton–Jacobi equations: a simple construction of convergent accurate difference schemes. Journal of Computational Physics, 284(??):367–388, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008560.

OSullivan:2015:CHO

[O’S15b] Stephen O’Sullivan. A class of high-order Runge–Kutta– Chebyshev stability polynomials. Journal of Computational Physics, 300(??):665–678, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005033.

Ohlberger:2016:ASI

[OS16] Mario Ohlberger and Kathrin Smetana. Approximation of skewed interfaces with tensor-based model reduction procedures: Application to the reduced basis hierarchical model reduction approach. Journal of Computational Physics, 321(??):1185–1205, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302467.

OSullivan:2019:RKG

[O’S19] Stephen O’Sullivan. Runge–Kutta–Gegenbauer explicit methods for advection–diffusion problems. Journal of Computational Physics, 388(??):209–223, July 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301706.

Ohwada:2018:SRE

[OSKN18] Taku Ohwada, Yuki Shibata, Takuma Kato, and Taichi Nakamura. A simple, robust and efficient high-order accurate shock-capturing scheme for compressible flows: Towards minimalism. Journal of Computational Physics, 362(??):131–162, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301013.

Olejnik:2017:SDS

[OSP17] Michal Olejnik, Kamil Szewc, and Jacek Pozorski. SPH with dynamical smoothing length adjustment based on the local flow kinematics. Journal of Computational Physics, 348(??):23–44, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305314.

Ojala:2015:AIE

[OT15] Rikard Ojala and Anna-Karin Tornberg. An accurate integral equation method for simulating multi-phase Stokes flow. Journal of Computational Physics, 298(??):145–160, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003812.

Ohno:2017:NTQ

[OTS17] Munekazu Ohno, Tomohiro Takaki, and Yasushi Shibuta. Numerical testing of quantitative phase-field models with different polynomials for isothermal solidification in binary alloys. Journal of Computational Physics, 335(??):621–636, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300694.

Ozbenli:2017:HOA

[OV17] Ersin Ozbenli and Prakash Vedula. High order accurate finite difference schemes based on symmetry preservation. Journal of Computational Physics, 349(??):376–398, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305946.

Oud:2016:FCM

[OvdHVH16] G. T. Oud, D. R. van der Heul, C. Vuik, and R. A. W. M. Henkes. A fully conservative mimetic discretization of the Navier–Stokes equations in cylindrical coordinates with associated singularity treatment. Journal of Computational Physics, 325(??):314–337, November 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303953.

Oguic:2015:PMC

[OVP15] Romain Oguic, Stéphane Viazzo, and Sébastien Poncet. A parallelized multidomain compact solver for incompressible turbulent flows in cylindrical geometries. Journal of Computational Physics, 300(??):710–731, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005203.

Owens:2016:DIA

[OWKE16] A. R. Owens, J. A. Welch, J. Kópházi, and M. D. Eaton. Discontinuous isogeometric analysis methods for the firstorder form of the neutron transport equation with discrete ordinate (S_{N}) angular discretisation. Journal of Computational Physics, 315(??):501–535, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300250.

Owhadi:2019:KFL

[OY19] Houman Owhadi and Gene Ryan Yoo. Kernel Flows: From learning kernels from data into the abyss. Journal of Computational Physics, 389(??):22–47, July 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302232.

Owhadi:2017:GOC

[OZ17] Houman Owhadi and Lei Zhang. Gamblets for opening the complexity-bottleneck of implicit schemes for hyperbolic and parabolic ODEs/PDEs with rough coefficients. Journal of Computational Physics, 347(??):99–128, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304916.

Powell:2015:EGR

[PA15] Devon Powell and Tom Abel. An exact general remeshing scheme applied to physically conservative voxelization. Journal of Computational Physics, 297(??):340–356, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003563.

Pandey:2019:ICF

[PA19] Ambuj Pandey and Akash Anand. Improved convergence of fast integral equation solvers for acoustic scattering by inhomogeneous penetrable media with discontinuous material interface. Journal of Computational Physics, 376(??):767–785, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306648.

Perez-Arancibia:2019:HDI

[PAFT19] Carlos Pérez-Arancibia, Luiz M. Faria, and Catalin Turc. Harmonic density interpolation methods for high-order evaluation of Laplace layer potentials in 2D and 3D. Journal of Computational Physics, 376(??):411–434, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306594.

Patkar:2016:TPP

[PAL^{+}16] Saket Patkar, Mridul Aanjaneya, Wenlong Lu, Michael Lentine, and Ronald Fedkiw. Towards positivity preservation for monolithic two-way solid-fluid coupling. Journal of Computational Physics, 312(??):82–114, May 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000644.

Panagiotou:2015:LNS

[Pan15] E. Panagiotou. The linking number in systems with Periodic Boundary Conditions. Journal of Computational Physics, 300(??):533–573, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005112.

Pantokratoras:2020:CPC

[Pan20] Asterios Pantokratoras. Comment on the paper “A computational wavelet method for variable-order fractional model of dual phase lag bioheat equation, M. Hosseininia, M.H. Heydari, R. Roohi, Z. Avazzadeh, Journal of Computational Physics 395 (2019) 1–18”. Journal of Computational Physics, 413(??):Article 109431, July 15, 2020. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999120302059. See [HHRA19].

Park:2015:MFS

[Par15] Won-Kwang Park. Multi-frequency subspace migration for imaging of perfectly conducting, arc-like cracks in full- and limited-view inverse scattering problems. Journal of Computational Physics, 283(??):52–80, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007980.

Park:2017:PAM

[Par17] Won-Kwang Park. Performance analysis of multi-frequency topological derivative for reconstructing perfectly conducting cracks. Journal of Computational Physics, 335(??):865–884, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301006.

Parent:2018:PPD

[Par18a] Bernard Parent. Positivity-preserving dual time stepping schemes for gas dynamics. Journal of Computational Physics, 361(??):391–411, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300561.

Park:2018:DSM

[Par18b] Won-Kwang Park. Direct sampling method for retrieving small perfectly conducting cracks. Journal of Computational Physics, 373(??):648–661, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304753.

Pasquetti:2016:CSI

[Pas16] Richard Pasquetti. Comparison of some isoparametric mappings for curved triangular spectral elements. Journal of Computational Physics, 316(??):573–577, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300845.

Philip:2015:PMD

[PBA^{+}15] Bobby Philip, Mark A. Berrill, Srikanth Allu, Steven P. Hamilton, Rahul S. Sampath, Kevin T. Clarno, and Gary A. Dilts. A parallel multi-domain solution methodology applied to nonlinear thermal transport problems in nuclear fuel pins. Journal of Computational Physics, 286(??):143–171, April 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000339.

Piao:2015:IFB

[PBBK15] Xiangfan Piao, Sunyoung Bu, Soyoon Bak, and Philsu Kim. An iteration free backward semi-Lagrangian scheme for solving incompressible Navier–Stokes equations. Journal of Computational Physics, 283(??):189–204, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400802X.

Pont:2017:MCH

[PBC^{+}17] Grégoire Pont, Pierre Brenner, Paola Cinnella, Bruno Maugars, and Jean-Christophe Robinet. Multiple-correction hybrid k-exact schemes for high-order compressible RANS–LES simulations on fully unstructured grids. Journal of Computational Physics, 350(??):45–83, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306137.

Picella:2019:SFT

[PBCR19] Francesco Picella, Michele Alessandro Bucci, Stefania Cherubini, and Jean-Christophe Robinet. A synthetic forcing to trigger laminar-turbulent transition in parallel wall bounded flows via receptivity. Journal of Computational Physics, 393(??):92–116, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930244X.

Piao:2017:EFC

[PBKK17] Xiangfan Piao, Sunyoung Bu, Dojin Kim, and Philsu Kim. An embedded formula of the Chebyshev collocation method for stiff problems. Journal of Computational Physics, 351(??):376–391, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117307064.

Peshkov:2019:TNC

[PBL^{+}19] Ilya Peshkov, Walter Boscheri, Raphaël Loubère, Evgeniy Romenski, and Michael Dumbser. Theoretical and numerical comparison of hyperelastic and hypoelastic formulations for Eulerian non-linear elastoplasticity. Journal of Computational Physics, 387(??):481–521, June 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301561.

Patel:2018:NCB

[PBP18] Namrata K. Patel, Amneet Pal Singh Bhalla, and Neelesh A. Patankar. A new constraint-based formulation for hydrodynamically resolved computational neuromechanics of swimming animals. Journal of Computational Physics, 375(??):684–716, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305643.

Parker:2016:PAN

[PC16] Scott E. Parker and Luis Chacon. Preface to advances in numerical simulation of plasmas. Journal of Computational Physics, 322(??):849, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302947.

Prieto:2019:SMA

[PC19] Juan Luis Prieto and Jaime Carpio. A-SLEIPNNIR: a multiscale, anisotropic adaptive, particle level set framework for moving interfaces. Transport equation applications. Journal of Computational Physics, 377(??):89–116, January 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306983.

Powell:2019:PMS

[PCA19] Samuel Powell, Ben T. Cox, and Simon R. Arridge. A pseudospectral method for solution of the radiative transport equation. Journal of Computational Physics, 384(??):376–382, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300658.

Pont:2018:USF

[PCBG18] Arnau Pont, Ramon Codina, Joan Baiges, and Oriol Guasch. Unified solver for fluid dynamics and aeroacoustics in isentropic gas flows. Journal of Computational Physics, 363(??):11–29, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301116.

Pang:2015:SFA

[PCF15] Guofei Pang, Wen Chen, and Zhuojia Fu. Space-fractional advection-dispersion equations by the Kansa method. Journal of Computational Physics, 293(??):280–296, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005130.

Park:2019:MMA

[PCMC19] H. Park, L. Chacón, A. Matsekh, and G. Chen. A multigroup moment-accelerated deterministic particle solver for 1-D time-dependent thermal radiative transfer problems. Journal of Computational Physics, 388(??):416–438, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301986.

Parsani:2015:ESD

[PCN15a] Matteo Parsani, Mark H. Carpenter, and Eric J. Nielsen. Entropy stable discontinuous interfaces coupling for the threedimensional compressible Navier–Stokes equations. Journal of Computational Physics, 290(??):132–138, June 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001126.

Parsani:2015:ESW

[PCN15b] Matteo Parsani, Mark H. Carpenter, and Eric J. Nielsen. Entropy stable wall boundary conditions for the threedimensional compressible Navier–Stokes equations. Journal of Computational Physics, 292(??):88–113, July 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001734.

Pan:2017:IFE

[PCX17] Qing Pan, Chong Chen, and Guoliang Xu. Isogeometric finite element approximation of minimal surfaces based on extended loop subdivision. Journal of Computational Physics, 343(??):324–339, August 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117303017.

Plante:2015:GFP

[PD15] Ianik Plante and Luc Devroye. On the Green’s function of the partially diffusion-controlled reversible ABCD reaction for radiation chemistry codes. Journal of Computational Physics, 297(??):515–529, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003253.

Parish:2016:PDD

[PD16a] Eric J. Parish and Karthik Duraisamy. A paradigm for datadriven predictive modeling using field inversion and machine learning. Journal of Computational Physics, 305(??):758–774, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007524.

Pourmatin:2016:MRS

[PD16b] Hossein Pourmatin and Kaushik Dayal. Multiscale real-space quantum-mechanical tight-binding calculations of electronic structure in crystals with defects using perfectly matched layers. Journal of Computational Physics, 323(??):115–125, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303199.

Parish:2017:DSS

[PD17] Eric J. Parish and Karthik Duraisamy. A dynamic subgrid scale model for Large Eddy Simulations based on the Mori–Zwanzig formalism. Journal of Computational Physics, 349(??):154–175, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305612.

Palmore:2019:VFF

[PD19] John Palmore and Olivier Desjardins. A volume of fluid framework for interface-resolved simulations of vaporizing liquid-gas flows. Journal of Computational Physics, 399 (??):Article 108954, December 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S002199911930659X.

Pan:2017:POU

[PDdG^{+}17] Yan Pan, Xiaoying Dai, Stefano de Gironcoli, Xin-Gao Gong, Gian-Marco Rignanese, and Aihui Zhou. A parallel orbital-updating based plane-wave basis method for electronic structure calculations. Journal of Computational Physics, 348(??):482–492, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305417.

Pettersson:2019:LSM

[PDN19] Per Pettersson, Alireza Doostan, and Jan Nordström. Level set methods for stochastic discontinuity detection in nonlinear problems. Journal of Computational Physics, 392(??): 511–531, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303067.

Phillips:2017:ETE

[PDRB17] Tyrone S. Phillips, Joseph M. Derlaga, Christopher J. Roy, and Jeff Borggaard. Error transport equation boundary conditions for the Euler and Navier–Stokes equations. Journal of Computational Physics, 330(??):46–64, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305812.

Pazzona:2015:IAM

[PDS15] Federico G. Pazzona, Pierfranco Demontis, and Giuseppe B. Suffritti. Improving the acceptance in Monte Carlo simulations: Sampling through intermediate states. Journal of Computational Physics, 295(??):556–568, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002600.

Phillips:2015:SAU

[PE15] Edward G. Phillips and Howard C. Elman. A stochastic approach to uncertainty in the equations of MHD kinematics. Journal of Computational Physics, 284(??):334–350, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008092.

Panourgias:2016:NFH

[PE16a] Konstantinos T. Panourgias and John A. Ekaterinaris. A nonlinear filter for high order discontinuous Galerkin discretizations with discontinuity resolution within the cell. Journal of Computational Physics, 326(??):234–257, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304065.

Petrov:2016:TBC

[PE16b] P. S. Petrov and M. Ehrhardt. Transparent boundary conditions for iterative high-order parabolic equations. Journal of Computational Physics, 313(??):144–158, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001017.

Pearson:2015:PIM

[Pea15] John W. Pearson. Preconditioned iterative methods for Navier–Stokes control problems. Journal of Computational Physics, 292(??):194–207, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911500176X.

Peixoto:2016:AAM

[Pei16] Pedro S. Peixoto. Accuracy analysis of mimetic finite volume operators on geodesic grids and a consistent alternative. Journal of Computational Physics, 310(??):127–160, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000267.

Peschka:2015:TFF

[Pes15] Dirk Peschka. Thin-film free boundary problems for partial wetting. Journal of Computational Physics, 295(??):770–778, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002983.

Papaioannou:2019:PBA

[PES19] Iason Papaioannou, Max Ehre, and Daniel Straub. PLS-based adaptation for efficient PCE representation in high dimensions. Journal of Computational Physics, 387(??):186–204, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301639.

Pereira:2018:CSR

[PEVG18] Filipe S. Pereira, Luís Eça, Guilherme Vaz, and Sharath S. Girimaji. Challenges in Scale-Resolving Simulations of turbulent wake flows with coherent structures. Journal of Computational Physics, 363(??):98–115, June 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301219.

Pletzer:2015:CIE

[PF15] Alexander Pletzer and David Fillmore. Conservative interpolation of edge and face data on n dimensional structured grids using differential forms. Journal of Computational Physics, 302(??):21–40, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005562.

Pepona:2016:CIB

[PF16] Marianna Pepona and Julien Favier. A coupled Immersed Boundary-Lattice Boltzmann method for incompressible flows through moving porous media. Journal of Computational Physics, 321(??):1170–1184, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302510.

Palha:2017:MEE

[PG17] A. Palha and M. Gerritsma. A mass, energy, enstrophy and vorticity conserving (MEEVC) mimetic spectral element discretization for the 2D incompressible Navier–Stokes equations. Journal of Computational Physics, 328(??):200–220, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305071.

Picot:2018:RDS

[PG18] Joris Picot and Stéphane Glockner. Reduction of the discretization stencil of direct forcing immersed boundary methods on rectangular cells: The ghost node shifting method. Journal of Computational Physics, 364(??):18–48, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301360.

Pan:2018:DFM

[PGCG18] Tsorng-Whay Pan, Aixia Guo, Shang-Huan Chiu, and Roland Glowinski. A 3D DLM/FD method for simulating the motion of spheres and ellipsoids under creeping flow conditions. Journal of Computational Physics, 352(??):410–425, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307027.

Peng:2018:DNS

[PGGW18] Cheng Peng, Nicholas Geneva, Zhaoli Guo, and Lian-Ping Wang. Direct numerical simulation of turbulent pipe flow using the lattice Boltzmann method. Journal of Computational Physics, 357(??):16–42, March 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730904X.

Plestenjak:2015:SCM

[PGH15] Bor Plestenjak, Calin I. Gheorghiu, and Michiel E. Hochstenbach. Spectral collocation for multiparameter eigenvalue problems arising from separable boundary value problems. Journal of Computational Physics, 298(??):585–601, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004015.

Peluchon:2017:RIE

[PGM17] S. Peluchon, G. Gallice, and L. Mieussens. A robust implicit-explicit acoustic-transport splitting scheme for twophase flows. Journal of Computational Physics, 339(??):328– 355, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302061.

Peng:2016:PCE

[PHD16] Ji Peng, Jerrad Hampton, and Alireza Doostan. On polynomial chaos expansion via gradient-enhanced l_{1}-minimization. Journal of Computational Physics, 310(??):440–458, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008682.

Pan:2018:CII

[PHHA18] Shucheng Pan, Luhui Han, Xiangyu Hu, and Nikolaus A. Adams. A conservative interface-interaction method for compressible multi-material flows. Journal of Computational Physics, 371(??):870–895, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300834.

Pan:2017:NEC

[PHHR17] Kejia Pan, Dongdong He, Hongling Hu, and Zhengyong Ren. A new extrapolation cascadic multigrid method for three dimensional elliptic boundary value problems. Journal of Computational Physics, 344(??):499–515, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303571.

Pasquariello:2016:CCF

[PHÖ^{+}16] Vito Pasquariello, Georg Hammerl, Felix Örley, Stefan Hickel, Caroline Danowski, Alexander Popp, Wolfgang A. Wall, and Nikolaus A. Adams. A cut-cell finite volume–finite element coupling approach for fluid-structure interaction in compressible flow. Journal of Computational Physics, 307(??):670–695, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008323.

Pavan:2016:SOR

[PHRA16] S. Pavan, J.-M. Hervouet, M. Ricchiuto, and R. Ata. A second order residual based predictor-corrector approach for time dependent pollutant transport. Journal of Computational Physics, 318(??):122–141, August 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301085.

Hao:2015:FOA

[pHzSrC15] Zhao peng Hao, Zhi zhong Sun, and Wan rong Cao. A fourthorder approximation of fractional derivatives with its applications. Journal of Computational Physics, 281(??):787–805, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007414.

Pinaud:2015:ALD

[Pin15] Olivier Pinaud. Absorbing layers for the Dirac equation. Journal of Computational Physics, 289(??):169–180, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001199.

Pishchalnikov:2018:ADE

[Pis18] Roman Pishchalnikov. Application of the differential evolution for simulation of the linear optical response of photosynthetic pigments. Journal of Computational Physics, 372(??):603–615, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304182.

Pozzetti:2019:PDG

[PJB^{+}19] Gabriele Pozzetti, Hrvoje Jasak, Xavier Besseron, Alban Rousset, and Bernhard Peters. A parallel dual-grid multiscale approach to CFD–DEM couplings. Journal of Computational Physics, 378(??):708–722, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307745.

Park:2016:ESV

[PJC16] Seong-Kwan Park, Gahyung Jo, and Hi Jun Choe. Existence and stability in the virtual interpolation point method for the Stokes equations. Journal of Computational Physics, 307(??):535–549, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008165.

Pandya:2016:ICB

[PJE^{+}16] Tara M. Pandya, Seth R. Johnson, Thomas M. Evans, Gregory G. Davidson, Steven P. Hamilton, and Andrew T. Godfrey. Implementation, capabilities, and benchmarking of shift, a massively parallel Monte Carlo radiation transport code. Journal of Computational Physics, 308(??):239–272, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008566.

Park:2016:HMD

[PK16] Jin Seok Park and Chongam Kim. Hierarchical multidimensional limiting strategy for correction procedure via reconstruction. Journal of Computational Physics, 308(??):57– 80, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008396.

Pardo:2017:CNL

[PK17] Enric Pardo and Milan Kapolka. 3D computation of nonlinear eddy currents: Variational method and superconducting cubic bulk. Journal of Computational Physics, 344(??): 339–363, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303704.

Pieper:2016:HPI

[PKA^{+}16] Andreas Pieper, Moritz Kreutzer, Andreas Alvermann, Martin Galgon, Holger Fehske, Georg Hager, Bruno Lang, and Gerhard Wellein. High-performance implementation of Chebyshev filter diagonalization for interior eigenvalue computations. Journal of Computational Physics, 325(??):226–243, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303837.

Pashos:2015:MPF

[PKB15] G. Pashos, G. Kokkoris, and A. G. Boudouvis. A modified phase-field method for the investigation of wetting transitions of droplets on patterned surfaces. Journal of Computational Physics, 283(??):258–270, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008080.

Palha:2016:MSE

[PKF16] A. Palha, B. Koren, and F. Felici. A mimetic spectral element solver for the Grad–Shafranov equation. Journal of Computational Physics, 316(??):63–93, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300341.

Park:2018:EMP

[PKJ^{+}18] Yesom Park, Jeongho Kim, Jinwook Jung, Euntaek Lee, and Chohong Min. An efficient MILU preconditioning for solving the 2D Poisson equation with Neumann boundary condition. Journal of Computational Physics, 356(??):115–126, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308665.

Piao:2018:OSS

[PKK18] Xiangfan Piao, Philsu Kim, and Dojin Kim. One-step L(α)- stable temporal integration for the backward semi-Lagrangian scheme and its application in guiding center problems. Journal of Computational Physics, 366(??):327–340, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302353.

Pan:2016:DMP

[PKLC16] Xiaomin Pan, Kyoungyoun Kim, Changhoon Lee, and Jung-Il Choi. A decoupled monolithic projection method for natural convection problems. Journal of Computational Physics, 314(??):160–166, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001716.

Pan:2017:FDM

[PKLC17] Xiaomin Pan, Kyoungyoun Kim, Changhoon Lee, and Jung-Il Choi. Fully decoupled monolithic projection method for natural convection problems. Journal of Computational Physics, 334(??):582–606, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300323.

Park:2017:MAL

[PKLS17] Won-Kwang Park, Hwa Pyung Kim, Kwang-Jae Lee, and Seong-Ho Son. MUSIC algorithm for location searching of dielectric anomalies from S-parameters using microwave imaging. Journal of Computational Physics, 348(??):259–270, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305430.

Pakravan:2017:GNF

[PKN17] Alireza Pakravan, Jun Won Kang, and Craig M. Newtson. A Gauss–Newton full-waveform inversion in PML-truncated domains using scalar probing waves. Journal of Computational Physics, 350(??):824–846, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306769.

Pan:2017:MEF

[PKP^{+}17] Wenxiao Pan, Kyungjoo Kim, Mauro Perego, Alexandre M. Tartakovsky, and Michael L. Parks. Modeling electrokinetic flows by consistent implicit incompressible smoothed particle hydrodynamics. Journal of Computational Physics, 334(??):125–144, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116307069.

Peherstorfer:2017:CMS

[PKW17] Benjamin Peherstorfer, Boris Kramer, and Karen Willcox. Combining multiple surrogate models to accelerate failure probability estimation with expensive high-fidelity models. Journal of Computational Physics, 341(??):61–75, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302838.

Pandare:2016:HRD

[PL16a] Aditya K. Pandare and Hong Luo. A hybrid reconstructed discontinuous Galerkin and continuous Galerkin finite element method for incompressible flows on unstructured grids. Journal of Computational Physics, 322(??):491–510, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302819.

Patel:2016:NSS

[PL16b] Saumil Patel and Taehun Lee. A new splitting scheme to the discrete Boltzmann equation for non-ideal gases on non-uniform meshes. Journal of Computational Physics, 327(??):799–809, December 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304818.

Pandare:2018:REF

[PL18] Aditya K. Pandare and Hong Luo. A robust and efficient finite volume method for compressible inviscid and viscous two-phase flows. Journal of Computational Physics, 371(??):67–91, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303188.

Petkova:2018:FAV

[PLB18] Maya A. Petkova, Guillaume Laibe, and Ian A. Bonnell. Fast and accurate Voronoi density gridding from Lagrangian hydrodynamics data. Journal of Computational Physics, 353(??):300–315, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307775.

Pan:2018:EMP

[PLC18] Xiaomin Pan, Changhoon Lee, and Jung-Il Choi. Efficient monolithic projection method for time-dependent conjugate heat transfer problems. Journal of Computational Physics, 369(??):191–208, September 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303024.

Pan:2018:HOT

[PLHA18] Shucheng Pan, Xiuxiu Lyu, Xiangyu Y. Hu, and Nikolaus A. Adams. High-order time-marching reinitialization for regional level-set functions. Journal of Computational Physics, 354(??):311–319, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308215.

Pal:2015:SBC

[PLL^{+}15a] Souvik Pal, Chuanjin Lan, Zhen Li, E. Daniel Hirleman, and Yanbao Ma. Symmetry boundary condition in dissipative particle dynamics. Journal of Computational Physics, 292(??):287–299, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001722.

Peng:2015:BIE

[PLL15b] Zhen Peng, Kheng-Hwee Lim, and Jin-Fa Lee. A boundary integral equation domain decomposition method for electromagnetic scattering from large and deep cavities. Journal of Computational Physics, 280(??):626–642, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006925.

Petras:2019:LSI

[PLPR19] A. Petras, L. Ling, C. Piret, and S. J. Ruuth. A leastsquares implicit RBF–FD closest point method and applications to PDEs on moving surfaces. Journal of Computational Physics, 381(??):146–161, March 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300129.

Petras:2018:RFC

[PLR18] A. Petras, L. Ling, and S. J. Ruuth. An RBF-FD closest point method for solving PDEs on surfaces. Journal of Computational Physics, 370(??):43–57, October 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830322X.

Peng:2016:IGK

[PLWJ16] Ao-Ping Peng, Zhi-Hui Li, Jun-Lin Wu, and Xin-Yu Jiang. Implicit gas-kinetic unified algorithm based on multi-block docking grid for multi-body reentry flows covering all flow regimes. Journal of Computational Physics, 327(??):919–942, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304715.

Park:2016:NAI

[PM16] George Ilhwan Park and Parviz Moin. Numerical aspects and implementation of a two-layer zonal wall model for LES of compressible turbulent flows on unstructured meshes. Journal of Computational Physics, 305(??):589–603, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007500.

Parameswaran:2019:NRS

[PM19] S. Parameswaran and J. C. Mandal. A novel Roe solver for incompressible two-phase flow problems. Journal of Computational Physics, 390(??):405–424, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302451.

Piatkowski:2018:SHO

[PMB18] Marian Piatkowski, Steffen Müthing, and Peter Bastian. A stable and high-order accurate discontinuous Galerkin based splitting method for the incompressible Navier–Stokes equations. Journal of Computational Physics, 356(??):220–239, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308732.

Puscas:2015:TSI

[PME^{+}15] Maria Adela Puscas, Laurent Monasse, Alexandre Ern, Christian Tenaud, Christian Mariotti, and Virginie Daru. A time semi-implicit scheme for the energy-balanced coupling of a shocked fluid flow with a deformable structure. Journal of Computational Physics, 296(??):241–262, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002582.

Pfeiffer:2015:HDC

[PMF15] M. Pfeiffer, C.-D. Munz, and S. Fasoulas. Hyperbolic divergence cleaning, the electrostatic limit, and potential boundary conditions for particle-in-cell codes. Journal of Computational Physics, 294(??):547–561, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002272.

Pigou:2018:NDE

[PMF^{+}18] Maxime Pigou, Jérôme Morchain, Pascal Fede, Marie-Isabelle Penet, and Geoffrey Laronze. New developments of the Extended Quadrature Method of Moments to solve Population Balance Equations. Journal of Computational Physics, 365(??):243–268, July 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301876.

Peng:2016:HCM

[PMGW16] Cheng Peng, Haoda Min, Zhaoli Guo, and Lian-Ping Wang. A hydrodynamically-consistent MRT lattice Boltzmann model on a 2D rectangular grid. Journal of Computational Physics, 326(??):893–912, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304521.

Parent:2015:MWI

[PMS15] Bernard Parent, Sergey O. Macheret, and Mikhail N. Shneider. Modeling weakly-ionized plasmas in magnetic field: a new computationally-efficient approach. Journal of Computational Physics, 300(??):779–799, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500529X.

Piotrowska:2019:SMP

[PMS19] Joanna Piotrowska, Jonah M. Miller, and Erik Schnetter. Spectral methods in the presence of discontinuities. Journal of Computational Physics, 390(??):527–547, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302311.

Patel:2017:NCW

[PN17] Jitendra Kumar Patel and Ganesh Natarajan. A novel consistent and well-balanced algorithm for simulations of multiphase flows on unstructured grids. Journal of Computational Physics, 350(??):207–236, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306289.

Patel:2018:DII

[PN18] Jitendra Kumar Patel and Ganesh Natarajan. Diffuse interface immersed boundary method for multi-fluid flows with arbitrarily moving rigid bodies. Journal of Computational Physics, 360(??):202–228, May 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300342.

Pettersson:2016:WPS

[PND16] Per Pettersson, Jan Nordström, and Alireza Doostan. A wellposed and stable stochastic Galerkin formulation of the incompressible Navier–Stokes equations with random data. Journal of Computational Physics, 306(??):92–116, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007676.

Pan:2018:CCI

[PNZ18] Jun-Hua Pan, Ming-Jiu Ni, and Nian-Mei Zhang. A consistent and conservative immersed boundary method for MHD flows and moving boundary problems. Journal of Computational Physics, 373(??):425–445, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309282.

Poette:2019:GIM

[Poë19] Gaël Poëtte. A gPC-intrusive Monte-Carlo scheme for the resolution of the uncertain linear Boltzmann equation. Journal of Computational Physics, 385(??):135–162, May 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930110X.

Popinet:2015:QAM

[Pop15] Stéphane Popinet. A quadtree-adaptive multigrid solver for the Serre-green-Naghdi equations. Journal of Computational Physics, 302(??):336–358, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005902.

Petersson:2016:DSP

[POSB16] N. Anders Petersson, Ossian O’Reilly, Björn Sjögreen, and Samuel Bydlon. Discretizing singular point sources in hyperbolic wave propagation problems. Journal of Computational Physics, 321(??):532–555, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302054.

Pazner:2017:SPF

[PP17] Will Pazner and Per-Olof Persson. Stage-parallel fully implicit Runge–Kutta solvers for discontinuous Galerkin fluid simulations. Journal of Computational Physics, 335(??):700–717, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300669.

Paparella:2018:LNM

[PP18a] Francesco Paparella and Marina Popolizio. Lagrangian numerical methods for ocean biogeochemical simulations. Journal of Computational Physics, 360(??):229–246, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830041X.

Pazner:2018:ATP

[PP18b] Will Pazner and Per-Olof Persson. Approximate tensorproduct preconditioners for very high order discontinuous Galerkin methods. Journal of Computational Physics, 354 (??):344–369, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117307830.

Pares:2019:RPS

[PP19] Carlos Parés and Ernesto Pimentel. The Riemann problem for the shallow water equations with discontinuous topography: the wet–dry case. Journal of Computational Physics, 378(??):344–365, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307459.

Pang:2017:DVF

[PPCK17] Guofei Pang, Paris Perdikaris, Wei Cai, and George Em Karniadakis. Discovering variable fractional orders of advection-dispersion equations from field data using multifidelity Bayesian optimization. Journal of Computational Physics, 348(??):694–714, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305600.

Park:2016:PCI

[PPLC16] Hyunwook Park, Xiaomin Pan, Changhoon Lee, and Jung-Il Choi. A pre-conditioned implicit direct forcing based immersed boundary method for incompressible viscous flows. Journal of Computational Physics, 314(??):774–799, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911600187X.

Pollack:2019:ZFA

[PPM^{+}19] Martin Pollack, Michele Pütz, Daniele L. Marchisio, Michael Oevermann, and Christian Hasse. Zero-flux approximations for multivariate quadrature-based moment methods. Journal of Computational Physics, 398(??):Article 108879, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119305765.

Pitton:2017:CRS

[PQR17] Giuseppe Pitton, Annalisa Quaini, and Gianluigi Rozza. Computational reduction strategies for the detection of steady bifurcations in incompressible fluid-dynamics: Applications to Coanda effect in cardiology. Journal of Computational Physics, 344(??):534–557, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303790.

Pathak:2016:FEV

[PR16a] Ashish Pathak and Mehdi Raessi. A 3D, fully Eulerian, VOF-based solver to study the interaction between two fluids and moving rigid bodies using the fictitious domain method. Journal of Computational Physics, 311(??):87–113, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000334.

Pathak:2016:TDV

[PR16b] Ashish Pathak and Mehdi Raessi. A three-dimensional volume-of-fluid method for reconstructing and advecting threematerial interfaces forming contact lines. Journal of Computational Physics, 307(??):550–573, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008232.

Petras:2016:PMS

[PR16c] A. Petras and S. J. Ruuth. PDEs on moving surfaces via the closest point method and a modified grid based particle method. Journal of Computational Physics, 312(??):139–156, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000814.

Pasquetti:2017:CVF

[PR17a] Richard Pasquetti and Francesca Rapetti. Cubature versus Fekete–Gauss nodes for spectral element methods on simplicial meshes. Journal of Computational Physics, 347(??):463–466, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305302.

Popovic:2017:AFI

[PR17b] Jelena Popovic and Olof Runborg. Adaptive fast interface tracking methods. Journal of Computational Physics, 337(??):42–61, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117301109.

Pederson:2019:SFD

[PR19] Dylan M. Pederson and Laxminarayan L. Raja. A stable finite-difference time-domain scheme for local timestepping on an adaptive mesh. Journal of Computational Physics, 394(??):456–476, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303870.

Prusa:2018:CCS

[Pru18] Joseph M. Prusa. Computation at a coordinate singularity. Journal of Computational Physics, 361(??):331–352, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300548.

Pfeifenberger:2018:NSS

[PRvdL18] Manuel J. Pfeifenberger, Michael Rumetshofer, and Wolfgang von der Linden. Nested sampling, statistical physics and the Potts model. Journal of Computational Physics, 375(??):368–392, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305813.

Pan:2019:IAS

[PRXC19] Qing Pan, Timon Rabczuk, Gang Xu, and Chong Chen. Isogeometric analysis for surface PDEs with extended loop subdivision. Journal of Computational Physics, 398(??):Article 108892, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S002199911930590X.

Pelanti:2014:MEC

[PS14] Marica Pelanti and Keh-Ming Shyue. A mixture-energy-consistent six-equation two-phase numerical model for fluids with interfaces, cavitation and evaporation waves. Journal of Computational Physics, 259(??):331–357, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008024. See corrigendum

[PS15a].

Pelanti:2015:CME

[PS15a] Marica Pelanti and Keh-Ming Shyue. Corrigendum to “A mixture-energy-consistent six-equation two-phase numerical model for fluids with interfaces, cavitation and evaporation waves” [J. Comput. Phys. 259 (2014) 331–357]. Journal of Computational Physics, 288(??):196–197, May 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000406. See [PS14].

Petersson:2015:WPA

[PS15b] N. Anders Petersson and Björn Sjögreen. Wave propagation in anisotropic elastic materials and curvilinear coordinates using a summation-by-parts finite difference method. Journal of Computational Physics, 299(??):820–841, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004684.

Pathak:2016:AFV

[PS16] Harshavardhana S. Pathak and Ratnesh K. Shukla. Adaptive finite-volume WENO schemes on dynamically redistributed grids for compressible Euler equations. Journal of Computational Physics, 319(??):200–230, August 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301346.

Potz:2017:SCF

[PS17] Walter Pötz and Magdalena Schreilechner. Single-cone finite difference scheme for the (2+1)d Dirac von Neumann equation. Journal of Computational Physics, 348(??):591–611, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305454.

Papoutsakis:2018:EAM

[PSB^{+}18] Andreas Papoutsakis, Sergei S. Sazhin, Steven Begg, Ionut Danaila, and Francky Luddens. An efficient Adaptive Mesh Refinement (AMR) algorithm for the Discontinuous Galerkin method: Applications for the computation of compressible twophase flows. Journal of Computational Physics, 363(??):399– 427, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301372.

Pieper:2019:ETD

[PSG19] Konstantin Pieper, K. Chad Sockwell, and Max Gunzburger. Exponential time differencing for mimetic multilayer ocean models. Journal of Computational Physics, 398(??):Article 108900, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305984.

Perrin:2017:NPT

[PSMPG17] G. Perrin, C. Soize, S. Marque-Pucheu, and J. Garnier. Nested polynomial trends for the improvement of Gaussian process-based predictors. Journal of Computational Physics, 346(??):389–402, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304448.

Poovathingal:2019:NFT

[PSN^{+}19] Savio Poovathingal, Eric C. Stern, Ioannis Nompelis, Thomas E. Schwartzentruber, and Graham V. Candler. Nonequilibrium flow through porous thermal protection materials, Part II: Oxidation and pyrolysis. Journal of Computational Physics, 380(??):427–441, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301268.

Pratapa:2016:AAJ

[PSP16] Phanisri P. Pratapa, Phanish Suryanarayana, and John E. Pask. Anderson acceleration of the Jacobi iterative method: an efficient alternative to Krylov methods for large, sparse linear systems. Journal of Computational Physics, 306(??):43–54, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007585.

Pantano:2017:OFS

[PSS17] C. Pantano, R. Saurel, and T. Schmitt. An oscillation free shock-capturing method for compressible van der Waals supercritical fluid flows. Journal of Computational Physics, 335(??):780–811, April 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300736.

Paalsson:2019:SVS

[PST19] Sara Pålsson, Michael Siegel, and Anna-Karin Tornberg. Simulation and validation of surfactant-laden drops in two-dimensional Stokes flow. Journal of Computational Physics, 386(??):218–247, June 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300932.

Prakapenia:2018:NST

[PSV18] M. A. Prakapenia, I. A. Siutsou, and G. V. Vereshchagin. Numerical scheme for treatment of Uehling–Uhlenbeck equation for two-particle interactions in relativistic plasma. Journal of Computational Physics, 373(??):533–544, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304650.

Petropavlovsky:2017:NDT

[PT17a] S. Petropavlovsky and S. Tsynkov. Non-deteriorating time domain numerical algorithms for Maxwell’s electrodynamics. Journal of Computational Physics, 336(??):1–35, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300840.

Portman:2017:SAV

[PT17b] Nataliya Portman and Isaac Tamblyn. Sampling algorithms for validation of supervised learning models for Ising-like systems. Journal of Computational Physics, 350(??):871–890, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304990.

Peles:2018:AMM

[PT18] Oren Peles and Eli Turkel. Acceleration methods for multi-physics compressible flow. Journal of Computational Physics, 358(??):201–234, April 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117307490.

Pierro:2018:SFP

[PTMF18] Vincenzo Pierro, Luigi Troiano, Elena Mejuto, and Giovanni Filatrella. Stochastic first passage time accelerated with CUDA. Journal of Computational Physics, 361(??):136–149, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300494.

Petropavlovsky:2018:MBE

[PTT18] S. Petropavlovsky, S. Tsynkov, and E. Turkel. A method of boundary equations for unsteady hyperbolic problems in 3D. Journal of Computational Physics, 365(??):294–323, July 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302055.

Petrella:2019:UQM

[PTT19] M. Petrella, S. Tokareva, and E. F. Toro. Uncertainty quantification methodology for hyperbolic systems with application to blood flow in arteries. Journal of Computational Physics, 386(??):405–427, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301305.

Parker:2015:CPA

[PUA^{+}15] William D. Parker, C. J. Umrigar, Dario Alfè, F. R. Petruzielo, Richard G. Hennig, and John W. Wilkins. Comparison of polynomial approximations to speed up planewave-based quantum Monte Carlo calculations. Journal of Computational Physics, 287(??):77–87, 2015. CODEN JCTPAH. ISSN

0021-9991. URL http://www.sciencedirect.com/science/
article/pii/S0021999115000510.

Posa:2017:ARL

[PVB17] Antonio Posa, Marcos Vanella, and Elias Balaras. An adaptive reconstruction for Lagrangian, direct-forcing, immersedboundary methods. Journal of Computational Physics, 351(??):422–436, December 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117307076.

Palamara:2015:EAG

[PVFN15] Simone Palamara, Christian Vergara, Elena Faggiano, and Fabio Nobile. An effective algorithm for the generation of patient-specific Purkinje networks in computational electrocardiology. Journal of Computational Physics, 283(??):495–517, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008055.

Pulido:2019:SMC

[PvL19] Manuel Pulido and Peter Jan van Leeuwen. Sequential Monte Carlo with kernel embedded mappings: the mapping particle filter. Journal of Computational Physics, 396(??):400–415, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304681.

Parussini:2017:MFG

[PVPK17] L. Parussini, D. Venturi, P. Perdikaris, and G. E. Karniadakis. Multi-fidelity Gaussian process regression for prediction of random fields. Journal of Computational Physics, 336(??):36–50, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300633.

Parmentier:2018:VPM

[PWC18a] Philippe Parmentier, Grégoire Winckelmans, and Philippe Chatelain. A vortex particle-mesh method for subsonic compressible flows. Journal of Computational Physics, 354(??): 692–716, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117308070.

Pimenta:2018:FNF

[PWC18b] Cristiano Pimenta, William R. Wolf, and André V. G. Cavalieri. A fast numerical framework to compute acoustic scattering by poroelastic plates of arbitrary geometry. Journal of Computational Physics, 373(??):763–783, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304807.

Popov:2015:SVC

[PWP15] Pavel P. Popov, Haifeng Wang, and Stephen B. Pope. Specific volume coupling and convergence properties in hybrid particle/finite volume algorithms for turbulent reactive flows. Journal of Computational Physics, 294(??):110–126, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001205.

Pan:2015:GCT

[PX15] Liang Pan and Kun Xu. Generalized coordinate transformation and gas-kinetic scheme. Journal of Computational Physics, 287(??):207–225, 2015. CODEN JCTPAH. ISSN

0021-9991. URL http://www.sciencedirect.com/science/
article/pii/S0021999115000716.

Pan:2016:TOC

[PX16] Liang Pan and Kun Xu. A third-order compact gas-kinetic scheme on unstructured meshes for compressible Navier–Stokes solutions. Journal of Computational Physics, 318(??):327–348, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301401.

Pan:2016:EAT

[PXLL16] Liang Pan, Kun Xu, Qibing Li, and Jiequan Li. An efficient and accurate two-stage fourth-order gas-kinetic scheme for the Euler and Navier–Stokes equations. Journal of Computational Physics, 326(??):197–221, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304119.

Pan:2017:HOS

[PxRS17] Jianhua Pan, Yu xin Ren, and Yutao Sun. High order subcell finite volume schemes for solving hyperbolic conservation laws II: Extension to two-dimensional systems on unstructured grids. Journal of Computational Physics, 338(??):165–198, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301559.

Pan:2015:IAB

[PXXZ15] Qing Pan, Guoliang Xu, Gang Xu, and Yongjie Zhang. Isogeometric analysis based on extended Loop’s subdivision. Journal of Computational Physics, 299(??):731–746, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500443X.

Pfeiffer:2019:HOS

[PYAG19] Robert A. Pfeiffer, John C. Young, Robert J. Adams, and Stephen D. Gedney. Higher-order simulation of impressed current cathodic protection systems. Journal of Computational Physics, 394(??):522–531, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119304152.

Pang:2019:NNI

[PYK19] Guofei Pang, Liu Yang, and George Em Karniadakis. Neural-net-induced Gaussian process regression for function approximation and PDE solution. Journal of Computational Physics, 384(??):270–288, May 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301032.

Podvigina:2016:CLM

[PZF16] O. Podvigina, V. Zheligovsky, and U. Frisch. The Cauchy– Lagrangian method for numerical analysis of Euler flow. Journal of Computational Physics, 306(??):320–342, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007858.

Pan:2015:DFI

[PZNG15] Tsorng-Whay Pan, Shihai Zhao, Xiting Niu, and Roland Glowinski. A DLM/FD/IB method for simulating compound vesicle motion under creeping flow condition. Journal of Computational Physics, 300(??):241–253, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005100.

Quaife:2016:ATS

[QB16] Bryan Quaife and George Biros. Adaptive time stepping for vesicle suspensions. Journal of Computational Physics, 306(??):478–499, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115007901.

Qiu:2015:NDG

[QDH15] Liangliang Qiu, Weihua Deng, and Jan S. Hesthaven. Nodal discontinuous Galerkin methods for fractional diffusion equations on 2D domain with triangular meshes. Journal of Computational Physics, 298(??):678–694, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004192.

Qin:2015:TPA

[QDRB15] Zhipeng Qin, Keegan Delaney, Amir Riaz, and Elias Balaras. Topology preserving advection of implicit interfaces on Cartesian grids. Journal of Computational Physics, 290(??):219–238, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000996.

Qin:2015:CHS

[QHZ^{+}15] Hong Qin, Yang He, Ruili Zhang, Jian Liu, Jianyuan Xiao, and Yulei Wang. Comment on “Hamiltonian splitting for the Vlasov–Maxwell equations”. Journal of Computational Physics, 297(??):721–723, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003265.

Qiu:2016:ADC

[QLF16] Linhai Qiu, Wenlong Lu, and Ronald Fedkiw. An adaptive discretization of compressible flow using a multitude of moving Cartesian grids. Journal of Computational Physics, 305(??):75–110, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006890.

Qi:2019:CCV

[QLS^{+}19] Yinghe Qi, Jiacai Lu, Ruben Scardovelli, Stéphane Zaleski, and Grétar Tryggvason. Computing curvature for volume of fluid methods using machine learning. Journal of Computational Physics, 377(??):155–161, January 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307046.

Quaife:2018:BIF

[QM18] Bryan D. Quaife and M. Nicholas J. Moore. A boundaryintegral framework to simulate viscous erosion of a porous medium. Journal of Computational Physics, 375(??):1–21, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304984.

Quaegebeur:2019:SES

[QN19] Samuel Quaegebeur and Siva Nadarajah. Stability of energy stable flux reconstruction for the diffusion problem using the interior penalty and Bassi and Rebay II numerical fluxes for linear triangular elements. Journal of Computational Physics, 380(??):88–118, March 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118308131.

Quaglino:2019:HDH

[QPK19] A. Quaglino, S. Pezzuto, and R. Krause. High-dimensional and higher-order multifidelity Monte Carlo estimators. Journal of Computational Physics, 388(??):300–315, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302098.

Quan:2016:MAC

[QS16] Chaoyu Quan and Benjamin Stamm. Mathematical analysis and calculation of molecular surfaces. Journal of Computational Physics, 322(??):760–782, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302868.

Qu:2018:IBF

[QSB18] Yegao Qu, Ruchao Shi, and Romesh C. Batra. An immersed boundary formulation for simulating high-speed compressible viscous flows with moving solids. Journal of Computational Physics, 354(??):672–691, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308124.

Qu:2019:GTD

[QSBY19] Feng Qu, Di Sun, Junqiang Bai, and Chao Yan. A genuinely two-dimensional Riemann solver for compressible flows in curvilinear coordinates. Journal of Computational Physics, 386(??):47–63, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301470.

Qin:2016:BPD

[QSY16] Tong Qin, Chi-Wang Shu, and Yang Yang. Bound-preserving discontinuous Galerkin methods for relativistic hydrodynamics. Journal of Computational Physics, 315(??):323–347, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300201.

Qian:2018:CTL

[QWX18] Longgen Qian, Yanhong Wei, and Feng Xiao. Coupled THINC and level set method: a conservative interface capturing scheme with high-order surface representations. Journal of Computational Physics, 373(??):284–303, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304534.

Qin:2019:DDG

[QWX19] Tong Qin, Kailiang Wu, and Dongbin Xiu. Data driven governing equations approximation using deep neural networks. Journal of Computational Physics, 395(??):620–635, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304504.

Qiao:2017:ADT

[QWXZ17] Changhe Qiao, Shuhong Wu, Jinchao Xu, and Chen-Song Zhang. Analytical decoupling techniques for fully implicit reservoir simulation. Journal of Computational Physics, 336(??):664–681, May 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301316.

Qi:2019:PSD

[QWZ^{+}19a] Hongxin Qi, Yuheng Wang, Jie Zhang, Xianghui Wang, and Jianguo Wang. A partially staggered discontinuous Galerkin method for transient electromagnetics. Journal of Computational Physics, 387(??):30–44, June 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301597.

Qian:2019:CFE

[QWZ19b] Yiran Qian, Zhongming Wang, and Shenggao Zhou. A conservative, free energy dissipating, and positivity preserving finite difference scheme for multi-dimensional nonlocal Fokker–Planck equation. Journal of Computational Physics, 386(??):22–36, June 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301457.

Qiu:2015:TGB

[QYF15] Linhai Qiu, Yue Yu, and Ronald Fedkiw. On thin gaps between rigid bodies two-way coupled to incompressible flow. Journal of Computational Physics, 292(??):1–29, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001746.

Qing:2019:RMM

[QYJ19] Fang Qing, Xijun Yu, and Zupeng Jia. A robust MoF method applicable to severely deformed polygonal mesh. Journal of Computational Physics, 377(??):162–182, January 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306995.

Romick:2017:HOS

[RA17] Christopher M. Romick and Tariq D. Aslam. High-order shock-fitted detonation propagation in high explosives. Journal of Computational Physics, 332(??):210–235, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306416.

Romick:2019:EHO

[RA19] Christopher M. Romick and Tariq D. Aslam. An extension of high-order shock-fitted detonation propagation in explosives. Journal of Computational Physics, 395(??):765–771, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119304188.

Ragusa:2015:DFE

[Rag15] Jean C. Ragusa. Discontinuous finite element solution of the radiation diffusion equation on arbitrary polygonal meshes and locally adapted quadrilateral grids. Journal of Computational Physics, 280(??):195–213, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006494.

Ramis:2017:ODL

[Ram17] Rafael Ramis. One-dimensional Lagrangian implicit hydrodynamic algorithm for Inertial Confinement Fusion applications. Journal of Computational Physics, 330(??):173–191, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305976.

Ramadan:2018:RSH

[Ram18] Omar Ramadan. Revisiting the stability of the HIE–FDTD technique for modeling graphene dispersion. Journal of Computational Physics, 372(??):719–725, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304467.

Rodrigues:2015:SIF

[RAMB15] Diego S. Rodrigues, Roberto F. Ausas, Fernando Mut, and Gustavo C. Buscaglia. A semi-implicit finite element method for viscous lipid membranes. Journal of Computational Physics, 298(??):565–584, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003964.

Ranocha:2018:GSP

[Ran18] Hendrik Ranocha. Generalised summation-by-parts operators and variable coefficients. Journal of Computational Physics, 362(??):20–48, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301037.

Ruiz-Baier:2015:PMF

[RB15] Ricardo Ruiz-Baier. Primal-mixed formulations for reactiondiffusion systems on deforming domains. Journal of Computational Physics, 299(??):320–338, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004635.

Roberts:2018:MID

[RB18] Malcolm Roberts and John C. Bowman. Multithreaded implicitly dealiased convolutions. Journal of Computational Physics, 356(??):98–114, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117308641.

Reynolds:2017:OSR

[RBD17] Matthew J. Reynolds, Gregory Beylkin, and Alireza Doostan. Optimization via separated representations and the canonical tensor decomposition. Journal of Computational Physics, 348 (??):220–230, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117305144.

Rodriguez:2015:FES

[RBGV15] Ana Alonso Rodríguez, Enrico Bertolazzi, Riccardo Ghiloni, and Alberto Valli. Finite element simulation of eddy current problems using magnetic scalar potentials. Journal of Computational Physics, 294(??):503–523, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002211.

Raeli:2018:FDM

[RBI18] Alice Raeli, Michel Bergmann, and Angelo Iollo. A finitedifference method for the variable coefficient Poisson equation on hierarchical Cartesian meshes. Journal of Computational Physics, 355(??):59–77, February 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308343.

Ramaswamy:2015:HPM

[RBJS15] Rajesh Ramaswamy, George Bourantas, Frank Jülicher, and Ivo F. Sbalzarini. A hybrid particle-mesh method for incompressible active polar viscous gels. Journal of Computational Physics, 291(??):334–361, June 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001412.

Rudy:2019:SPE

[RBK19] Samuel H. Rudy, Steven L. Brunton, and J. Nathan Kutz. Smoothing and parameter estimation by soft-adherence to governing equations. Journal of Computational Physics, 398(??):Article 108860, December 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305443.

Ruiz-Baier:2016:MFE

[RBL16] Ricardo Ruiz-Baier and Ivan Lunati. Mixed finite element– discontinuous finite volume element discretization of a general class of multicontinuum models. Journal of Computational Physics, 322(??):666–688, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302790.

Raoult:2019:DVR

[RBY19] Cécile Raoult, Michel Benoit, and Marissa L. Yates. Development and validation of a 3D RBF-spectral model for coastal wave simulation. Journal of Computational Physics, 378(??):278–302, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118307204.

Razaaly:2018:NAU

[RC18] Nassim Razaaly and Pietro Marco Congedo. Novel algorithm using Active Metamodel Learning and Importance Sampling: Application to multiple failure regions of low probability. Journal of Computational Physics, 368(??):92–114, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302808.

Ricketson:2016:ADE

[RCRF16] L. F. Ricketson, A. J. Cerfon, M. Rachh, and J. P. Freidberg. Accurate derivative evaluation for any Grad–Shafranov solver. Journal of Computational Physics, 305(??):744–757, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500755X.

Re:2017:IFA

[RDG17] B. Re, C. Dobrzynski, and A. Guardone. An interpolation-free ALE scheme for unsteady inviscid flows computations with large boundary displacements over three-dimensional adaptive grids. Journal of Computational Physics, 340(??):26–54, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302292.

Roberts:2015:DPG

[RDM15] Nathan V. Roberts, Leszek Demkowicz, and Robert Moser. A discontinuous Petrov–Galerkin methodology for adaptive solutions to the incompressible Navier–Stokes equations. Journal of Computational Physics, 301(??):456–483, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004593.

Rife:2019:NFF

[RdM19] Max E. Rife and Luca di Mare. Numerical flux function for flow through porous media with discontinuous properties. Journal of Computational Physics, 397(??):Article 108833, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305170.

Regazzoni:2019:MLF

[RDQ19] F. Regazzoni, L. Dedè, and A. Quarteroni. Machine learning for fast and reliable solution of time-dependent differential equations. Journal of Computational Physics, 397(??):Article 108852, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305364.

Renac:2019:ESD

[Ren19] Florent Renac. Entropy stable DGSEM for nonlinear hyperbolic systems in nonconservative form with application to twophase flows. Journal of Computational Physics, 382(??):1–26, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911930021X.

Reeger:2018:NQS

[RF18] Jonah A. Reeger and Bengt Fornberg. Numerical quadrature over smooth surfaces with boundaries. Journal of Computational Physics, 355(??):176–190, February 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308483.

Rahbaralam:2015:DWR

[RFGSV15] Maryam Rahbaralam, Daniel Fernàndez-Garcia, and Xavier Sanchez-Vila. Do we really need a large number of particles to simulate bimolecular reactive transport with random walk methods? A kernel density estimation approach. Journal of Computational Physics, 303(??):95–104, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006233.

Radivojevic:2018:MSS

[RFPSSA18] Tijana Radivojević, Mario Fernández-Pendás, Jesús María Sanz-Serna, and Elena Akhmatskaya. Multi-stage splitting integrators for sampling with modified Hamiltonian Monte Carlo methods. Journal of Computational Physics, 373(??):900–916, November 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304844.

Rangarajan:2015:FEM

[RG15] Ramsharan Rangarajan and Huajian Gao. A finite element method to compute three-dimensional equilibrium configurations of fluid membranes: Optimal parameterization, variational formulation and applications. Journal of Computational Physics, 297(??):266–294, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003198.

Rietmann:2017:NLT

[RGPS17] Max Rietmann, Marcus Grote, Daniel Peter, and Olaf Schenk. Newmark local time stepping on high-performance computing architectures. Journal of Computational Physics, 334(??):308– 326, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305988.

Remacle:2016:GAS

[RGW16] J.-F. Remacle, R. Gandham, and T. Warburton. GPU accelerated spectral finite elements on all-hex meshes. Journal of Computational Physics, 324(??):246–257, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630345X.

Ray:2018:ANN

[RH18] Deep Ray and Jan S. Hesthaven. An artificial neural network as a troubled-cell indicator. Journal of Computational Physics, 367(??):166–191, August 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302547.

Ray:2019:DTC

[RH19] Deep Ray and Jan S. Hesthaven. Detecting troubled-cells on two-dimensional unstructured grids using a neural network. Journal of Computational Physics, 397(??):Article 108845, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305297.

Renaud:2018:DGM

[RHS18] Adrien Renaud, Thomas Heuzé, and Laurent Stainier. A Discontinuous Galerkin Material Point Method for the solution of impact problems in solid dynamics. Journal of Computational Physics, 369(??):80–102, September 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302936.

Rossinelli:2015:MIM

[RHvR^{+}15] Diego Rossinelli, Babak Hejazialhosseini, Wim van Rees, Mattia Gazzola, Michael Bergdorf, and Petros Koumoutsakos. MRAG-I2D: Multi-resolution adapted grids for remeshed vortex methods on multicore architectures. Journal of Computational Physics, 288(??):1–18, May 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500039X.

Ricchiuto:2015:ERB

[Ric15] Mario Ricchiuto. An explicit residual based approach for shallow water flows. Journal of Computational Physics, 280 (??):306–344, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006639.

Riesch:2019:APP

[RJ19a] Michael Riesch and Christian Jirauschek. Analyzing the positivity preservation of numerical methods for the Liouville– von Neumann equation. Journal of Computational Physics, 390(??):290–296, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119302396.

Rodriguez:2019:HOA

[RJ19b] Mauro Rodriguez and Eric Johnsen. A high-order accurate five-equations compressible multiphase approach for viscoelastic fluids and solids with relaxation and elasticity. Journal of Computational Physics, 379(??):70–90, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307022.

Roy:2019:BPN

[RJLW19] Thomas Roy, Tom B. Jönsthövel, Christopher Lemon, and Andrew J. Wathen. A block preconditioner for nonisothermal flow in porous media. Journal of Computational Physics, 395(??):636–652, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119304462.

Raissi:2018:HPM

[RK18] Maziar Raissi and George Em Karniadakis. Hidden physics models: Machine learning of nonlinear partial differential equations. Journal of Computational Physics, 357(??):125–141, March 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117309014.

Rudy:2019:DLD

[RKB19] Samuel H. Rudy, J. Nathan Kutz, and Steven L. Brunton. Deep learning of dynamics and signal-noise decomposition with time-stepping constraints. Journal of Computational Physics, 396(??):483–506, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304644.

Runov:2015:EMP

[RKH15] A. M. Runov, S. V. Kasilov, and P. Helander. Energy and momentum preserving Coulomb collision model for kinetic Monte Carlo simulations of plasma steady states in toroidal fusion devices. Journal of Computational Physics, 300(??):605–622, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005185.

Ramos:2018:LSA

[RKL18] Pablo Miguel Ramos, Nikos Ch. Karayiannis, and Manuel Laso. Off-lattice simulation algorithms for athermal chain molecules under extreme confinement. Journal of Computational Physics, 375(??):918–934, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305850.

Razi:2019:FPM

[RKN19] Mani Razi, Robert M. Kirby, and Akil Narayan. Fast predictive multi-fidelity prediction with models of quantized fidelity levels. Journal of Computational Physics, 376(??):992–1008, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306922.

Rachh:2017:FAQ

[RKO17a] Manas Rachh, Andreas Klöckner, and Michael O’Neil. Fast algorithms for Quadrature by Expansion I: Globally valid expansions. Journal of Computational Physics, 345(??):706–731, September 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117303418.

Rosen:2017:HAR

[RKO^{+}17b] A. L. Rosen, M. R. Krumholz, J. S. Oishi, A. T. Lee, and R. I. Klein. Hybrid Adaptive Ray-Moment Method (HARM^{2}): a highly parallel method for radiation hydrodynamics on adaptive grids. Journal of Computational Physics, 330(??):924–942, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305563.

Reitzle:2017:VFM

[RKRGW17] M. Reitzle, C. Kieffer-Roth, H. Garcke, and B. Weigand. A volume-of-fluid method for three-dimensional hexagonal solidification processes. Journal of Computational Physics, 339(??):356–369, June 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301882.

Russo:2017:FAE

[RL17] Serena Russo and Paolo Luchini. A fast algorithm for the estimation of statistical error in DNS (or experimental) time averages. Journal of Computational Physics, 347(??):328–340, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305077.

Rudyak:2018:SAS

[RL18] V. Ya. Rudyak and E. V. Lezhnev. Stochastic algorithm for simulating gas transport coefficients. Journal of Computational Physics, 355(??):95–103, February 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308288.

Reyes:2019:VHO

[RLGT19] Adam Reyes, Dongwook Lee, Carlo Graziani, and Petros Tzeferacos. A variable high-order shock-capturing finite difference method with GP–WENO. Journal of Computational Physics, 381(??):189–217, March 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300099.

Rodriguez-Linan:2019:GBV

[RLH19] Gustavo M. Rodríguez-Liñán and Marco Heinen. Granular beads in a vibrating, quasi two-dimensional cell: the true shape of the effective pair potential. Journal of Computational Physics, 394(??):232–242, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303997.

Rieutord:2016:ACS

[RLP16] Michel Rieutord, Francisco Espinosa Lara, and Bertrand Putigny. An algorithm for computing the 2D structure of fast rotating stars. Journal of Computational Physics, 318(??):277–304, August 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116301395.

Reckinger:2016:CNM

[RLV16] Scott J. Reckinger, Daniel Livescu, and Oleg V. Vasilyev. Comprehensive numerical methodology for direct numerical simulations of compressible Rayleigh–Taylor instability. Journal of Computational Physics, 313(??):181–208, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007354.

Rizzuti:2016:MBS

[RM16] G. Rizzuti and W. A. Mulder. Multigrid-based ‘shifted-Laplacian’ preconditioning for the time-harmonic elastic wave equation. Journal of Computational Physics, 317(??):47–65, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301048.

Rahmouni:2017:TVI

[RMA17] Lyes Rahmouni, Rajendra Mitharwal, and Francesco P. Andriulli. Two volume integral equations for the inhomogeneous and anisotropic forward problem in electroencephalography. Journal of Computational Physics, 348(??):732–743, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305156.

Roberts:2018:VCD

[RMBN18] Scott A. Roberts, Hector Mendoza, Victor E. Brunini, and David R. Noble. A verified conformal decomposition finite element method for implicit, many-material geometries. Journal of Computational Physics, 375(??):352–367, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830545X.

Romano:2015:DMC

[RMC15] Vittorio Romano, Armando Majorana, and Marco Coco. DSMC method consistent with the Pauli exclusion principle and comparison with deterministic solutions for charge transport in graphene. Journal of Computational Physics, 302 (??):267–284, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115005768.

Riahi:2018:PCI

[RMF^{+}18] H. Riahi, M. Meldi, J. Favier, E. Serre, and E. Goncalves. A pressure-corrected Immersed Boundary Method for the numerical simulation of compressible flows. Journal of Computational Physics, 374(??):361–383, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304947.

Rauser:2015:ETN

[RMK15] Florian Rauser, Jochem Marotzke, and Peter Korn. Ensembletype numerical uncertainty information from single model integrations. Journal of Computational Physics, 292(??):30–42, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001138.

Reimann:2018:RRB

[RMLvR18] Franziska Reimann, Peter Michel, Ulf Lehnert, and Ursula van Rienen. Rayleigh–Ritz based expansion method for wakefields in dielectrically lined rectangular waveguides. Journal of Computational Physics, 372(??):299–315, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303826.

Rao:2018:SCI

[RMP18] Kaustubh Rao, Paul Malan, and J. Blair Perot. A stopping criterion for the iterative solution of partial differential equations. Journal of Computational Physics, 352(??):265–284, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306939.

Ringue:2018:OBF

[RN18a] N. Ringue and S. Nadarajah. An optimization-based framework for anisotropic hp -adaptation of high-order discretizations. Journal of Computational Physics, 375(??):589–618, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305990.

Ruggiu:2018:PST

[RN18b] Andrea A. Ruggiu and Jan Nordström. On pseudo-spectral time discretizations in summation-by-parts form. Journal of Computational Physics, 360(??):192–201, May 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300536.

Rakhuba:2019:LRR

[RNO19] Maxim Rakhuba, Alexander Novikov, and Ivan Oseledets. Low-rank Riemannian eigensolver for high-dimensional Hamiltonians. Journal of Computational Physics, 396(??):718–737, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304875.

Rakhuba:2016:GBE

[RO16] M. V. Rakhuba and I. V. Oseledets. Grid-based electronic structure calculations: the tensor decomposition approach. Journal of Computational Physics, 312(??):19–30, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000772.

Rostami:2019:FAL

[RO19] Minghao W. Rostami and Sarah D. Olson. Fast algorithms for large dense matrices with applications to biofluids. Journal of Computational Physics, 394(??):364–384, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303869.

Rodionov:2017:AVG

[Rod17] Alexander V. Rodionov. Artificial viscosity in Godunovtype schemes to cure the carbuncle phenomenon. Journal of Computational Physics, 345(??):308–329, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303996.

Rodionov:2018:AVC

[Rod18] Alexander V. Rodionov. Artificial viscosity to cure the carbuncle phenomenon: The three-dimensional case. Journal of Computational Physics, 361(??):50–55, May 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300779.

Ranocha:2016:SPO

[RÖS16] Hendrik Ranocha, Philipp Öffner, and Thomas Sonar. Summation-by-parts operators for correction procedure via reconstruction. Journal of Computational Physics, 311(??): 299–328, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116000632.

Ranocha:2017:ESS

[RÖS17] Hendrik Ranocha, Philipp Öffner, and Thomas Sonar. Extended skew-symmetric form for summation-by-parts operators and varying Jacobians. Journal of Computational Physics, 342(??):13–28, August 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117303236.

Roy:2015:CFS

[Roy15] Indrajit G. Roy. On computing first and second order derivative spectra. Journal of Computational Physics, 295 (??):307–321, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115002612.

Rubino:2018:ABF

[RPC^{+}18] A. Rubino, M. Pini, P. Colonna, T. Albring, S. Nimmagadda, T. Economon, and J. Alonso. Adjoint-based fluid dynamic design optimization in quasi-periodic unsteady flow problems using a harmonic balance method. Journal of Computational Physics, 372(??):220–235, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304017.

Raissi:2017:ISD

[RPK17a] Maziar Raissi, Paris Perdikaris, and George Em Karniadakis. Inferring solutions of differential equations using noisy multi-fidelity data. Journal of Computational Physics, 335(??):736–746, April 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300761.

Raissi:2017:MLL

[RPK17b] Maziar Raissi, Paris Perdikaris, and George Em Karniadakis. Machine learning of linear differential equations using Gaussian processes. Journal of Computational Physics, 348(??):683–693, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305582.

Raissi:2019:PIN

[RPK19] M. Raissi, P. Perdikaris, and G. E. Karniadakis. Physicsinformed neural networks: a deep learning framework for solving forward and inverse problems involving nonlinear partial differential equations. Journal of Computational Physics, 378(??):686–707, February 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307125.

Royston:2018:PRU

[RPL^{+}18] Michael Royston, Andre Pradhana, Byungjoon Lee, Yat Tin Chow, Wotao Yin, Joseph Teran, and Stanley Osher. Parallel redistancing using the Hopf–Lax formula. Journal of Computational Physics, 365(??):7–17, July 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300457.

Rattia:2018:MLS

[RPNP18] J. M. Nunez Rattia, J. R. Percival, S. J. Neethling, and M. D. Piggott. Modelling local scour near structures with combined mesh movement and mesh optimisation. Journal of Computational Physics, 375(??):1220–1237, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306260.

Rodriguez-Rozas:2016:NCC

[RRD16] Ángel Rodríguez-Rozas and Julien Diaz. Non-conforming curved finite element schemes for time-dependent elasticacoustic coupled problems. Journal of Computational Physics, 305(??):44–62, January 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006920.

Rosenfeld:2019:MFP

[RRD19] Joel A. Rosenfeld, Spencer A. Rosenfeld, and Warren E. Dixon. A mesh-free pseudospectral approach to estimating the fractional Laplacian via radial basis functions. Journal of Computational Physics, 390(??):306–322, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301329.

Reiser:2019:PTL

[RRL19] D. Reiser, J. Romazanov, and Ch. Linsmeier. On the possibility of track length based Monte-Carlo algorithms for stationary drift-diffusion systems with sources and sinks. Journal of Computational Physics, 377(??):219–231, January 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830514X.

Ravu:2016:CAD

[RRM^{+}16] Bharath Ravu, Murray Rudman, Guy Metcalfe, Daniel R. Lester, and Devang V. Khakhar. Creating analytically divergence-free velocity fields from grid-based data. Journal of Computational Physics, 323(??):75–94, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303138.

Rueda-Ramirez:2019:PMS

[RRMF^{+}19] Andrés M. Rueda-Ramírez, Juan Manzanero, Esteban Ferrer, Gonzalo Rubio, and Eusebio Valero. A p-multigrid strategy with anisotropic p-adaptation based on truncation errors for high-order discontinuous Galerkin methods. Journal of Computational Physics, 378(??):209–233, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307277.

Ramani:2019:STSb

[RRS19a] Raaghav Ramani, Jon Reisner, and Steve Shkoller. A spacetime smooth artificial viscosity method with wavelet noise indicator and shock collision scheme, Part 1: the 1-D case. Journal of Computational Physics, 387(??):81–116, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301664.

Ramani:2019:STSa

[RRS19b] Raaghav Ramani, Jon Reisner, and Steve Shkoller. A spacetime smooth artificial viscosity method with wavelet noise indicator and shock collision scheme, Part 2: the 2-D case. Journal of Computational Physics, 387(??):45–80, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301652.

Rao:2015:NSE

[RS15a] Parthib R. Rao and Laura A. Schaefer. Numerical stability of explicit off-lattice Boltzmann schemes: a comparative study. Journal of Computational Physics, 285(??):251–264, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000212.

Rolland:2015:SBA

[RS15b] Joran Rolland and Eric Simonnet. Statistical behaviour of adaptive multilevel splitting algorithms in simple models. Journal of Computational Physics, 283(??):541–558, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008171.

Rao:2016:TPA

[RS16a] Vishwas Rao and Adrian Sandu. A time-parallel approach to strong-constraint four-dimensional variational data assimilation. Journal of Computational Physics, 313(??):583–593, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001042.

Rapaka:2016:IBM

[RS16b] Narsimha R. Rapaka and Sutanu Sarkar. An immersed boundary method for direct and large eddy simulation of stratified flows in complex geometry. Journal of Computational Physics, 322(??):511–534, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302613.

Reeve:2017:ECM

[RS17] Samuel Temple Reeve and Alejandro Strachan. Error correction in multi-fidelity molecular dynamics simulations using functional uncertainty quantification. Journal of Computational Physics, 334(??):207–220, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630701X.

Rapaka:2018:NNS

[RS18] Narsimha Reddy Rapaka and Ravi Samtaney. Non-normal stability of embedded boundary methods through pseudospectra. Journal of Computational Physics, 373(??):975–999, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304376.

Rycroft:2015:EPM

[RSB15] Chris H. Rycroft, Yi Sui, and Eran Bouchbinder. An Eulerian projection method for quasi-static elastoplasticity. Journal of Computational Physics, 300(??):136–166, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004453.

Rodrigues:2016:CME

[RSB16] Caio F. Rodrigues, Jorge L. Suzuki, and Marco L. Bittencourt. Construction of minimum energy high-order Helmholtz bases for structured elements. Journal of Computational Physics, 306(??):269–290, February 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007731.

Rezaei:2019:AFM

[RSBS19] Ali Rezaei, Fahd Siddiqui, Giorgio Bornia, and Mohammed Soliman. Applications of the fast multipole fully coupled poroelastic displacement discontinuity method to hydraulic fracturing problems. Journal of Computational Physics, 399(??):Article 108955, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306606.

Romer:2017:DCF

[RSD17] Ulrich Römer, Sebastian Schöps, and Herbert De Gersem. A defect corrected finite element approach for the accurate evaluation of magnetic fields on unstructured grids. Journal of Computational Physics, 335(??):688–699, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300578.

Rona:2017:OPC

[RSH^{+}17] A. Rona, I. Spisso, E. Hall, M. Bernardini, and S. Pirozzoli. Optimised prefactored compact schemes for linear wave propagation phenomena. Journal of Computational Physics, 328(??):66–85, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305125.

Roohi:2018:GFB

[RSSSE18] Ehsan Roohi, Stefan Stefanov, Ahmad Shoja-Sani, and Hossein Ejraei. A generalized form of the Bernoulli Trial collision scheme in DSMC: Derivation and evaluation. Journal of Computational Physics, 354(??):476–492, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307921.

Rainwater:2016:NAC

[RT16] G. Rainwater and M. Tokman. A new approach to constructing efficient stiffly accurate EPIRK methods. Journal of Computational Physics, 323(??):283–309, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303217.

Rieke:2015:CVT

[RTG15] M. Rieke, T. Trost, and R. Grauer. Coupled Vlasov and two-fluid codes on GPUs. Journal of Computational Physics, 283(??):436–452, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400833X.

Rabinovich:2018:ATR

[RTG18] Daniel Rabinovich, Eli Turkel, and Dan Givoli. An augmented time reversal method for source and scatterer identification. Journal of Computational Physics, 375(??):99–119, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305497.

Rojas:2015:PFL

[RTO15] Roberto Rojas, Tomohiro Takaki, and Munekazu Ohno. A phase-field-lattice Boltzmann method for modeling motion and growth of a dendrite for binary alloy solidification in the presence of melt convection. Journal of Computational Physics, 298(??):29–40, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003794.

Rapun:2017:LCP

[RTV17] M.-L. Rapún, F. Terragni, and J. M. Vega. LUPOD: Collocation in POD via LU decomposition. Journal of Computational Physics, 335(??):1–20, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300153.

Ruan:2018:NGF

[Rua18] Xinran Ruan. A normalized gradient flow method with attractive-repulsive splitting for computing ground states of Bose–Einstein condensates with higher-order interaction. Journal of Computational Physics, 367(??):374–390, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302638.

Reuther:2016:ITP

[RV16] Sebastian Reuther and Axel Voigt. Incompressible two-phase flows with an inextensible Newtonian fluid interface. Journal of Computational Physics, 322(??):850–858, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303187.

Rodriguez:2018:STF

[RVK^{+}18] C. Rodriguez, A. Vidal, P. Koukouvinis, M. Gavaises, and M. A. McHugh. Simulation of transcritical fluid jets using the PC-SAFT EoS. Journal of Computational Physics, 374(??):444–468, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304911.

Rosenthal:2017:DDA

[RVMR17] W. Steven Rosenthal, Shankar Venkataramani, Arthur J. Mariano, and Juan M. Restrepo. Displacement data assimilation.

Journal of Computational Physics, 330(??):594–614, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630523X.

Rahimian:2015:BIM

[RVZB15] Abtin Rahimian, Shravan K. Veerapaneni, Denis Zorin, and George Biros. Boundary integral method for the flow of vesicles with viscosity contrast in three dimensions. Journal of Computational Physics, 298(??):766–786, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004143.

Rauschenberger:2015:DNS

[RW15a] P. Rauschenberger and B. Weigand. Direct numerical simulation of rigid bodies in multiphase flow within an Eulerian framework. Journal of Computational Physics, 291(??):238– 253, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001709.

Rauschenberger:2015:VFM

[RW15b] P. Rauschenberger and B. Weigand. A volume-of-fluid method with interface reconstruction for ice growth in supercooled water. Journal of Computational Physics, 282(??):98–112, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007190.

Ran:2019:EID

[RW19] Xin Ran and Moran Wang. Efficiency improvement of discreteordinates method for interfacial phonon transport by Gauss– Legendre integral for frequency domain. Journal of Computational Physics, 399(??):Article 108920, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306254.

Rasthofer:2018:EAV

[RWG18] U. Rasthofer, W. A. Wall, and V. Gravemeier. An extended algebraic variational multiscale–multigrid–multifractal method (XAVM 4) for large-eddy simulation of turbulent two-phase flow. Journal of Computational Physics, 359(??):1–19, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300238.

Rider:2016:RVA

[RWKW16] William Rider, Walt Witkowski, James R. Kamm, and Tim Wildey. Robust verification analysis. Journal of Computational Physics, 307(??):146–163, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008025.

Ruggiu:2018:NMF

[RWN18] Andrea A. Ruggiu, Per Weinerfelt, and Jan Nordström. A new multigrid formulation for high order finite difference methods on summation-by-parts form. Journal of Computational Physics, 359(??):216–238, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300214.

Ren:2016:MDH

[RXS16] Xiaodong Ren, Kun Xu, and Wei Shyy. A multi-dimensional high-order DG–ALE method based on gas-kinetic theory with application to oscillating bodies. Journal of Computational Physics, 316(??):700–720, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300742.

Ren:2015:MDH

[RXSG15] Xiaodong Ren, Kun Xu, Wei Shyy, and Chunwei Gu. A multi-dimensional high-order discontinuous Galerkin method based on gas kinetic theory for viscous flow computations. Journal of Computational Physics, 292(??):176–193, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001783.

Ran:2018:GBA

[RYZ18] Chunjiang Ran, Haitian Yang, and Guoqing Zhang. A gradient based algorithm to solve inverse plane bimodular problems of identification. Journal of Computational Physics, 355(??):78–94, February 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911730832X.

Rosatti:2015:MTB

[RZ15] Giorgio Rosatti and Daniel Zugliani. Modelling the transition between fixed and mobile bed conditions in twophase free-surface flows: the Composite Riemann Problem and its numerical solution. Journal of Computational Physics, 285(??):226–250, March 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000157.

Reiter:2017:MDB

[RZ17] Paul Reiter and Harald Ziegelwanger. Multi-domain boundary element method for axi-symmetric layered linear acoustic systems. Journal of Computational Physics, 350(??):759–781, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306435.

Rundell:2018:RUS

[RZ18] William Rundell and Zhidong Zhang. Recovering an unknown source in a fractional diffusion problem. Journal of Computational Physics, 368(??):299–314, September 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302791.

Ragnoli:2019:LSS

[RZOZ19] Emanuele Ragnoli, Mykhaylo Zayats, Fearghal O’Donncha, and Sergiy Zhuk. Localised sequential state estimation for advection dominated flows with non-Gaussian uncertainty description. Journal of Computational Physics, 387(??):356–375, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300919.

Ren:2019:FAR

[RZZ19] Kui Ren, Rongting Zhang, and Yimin Zhong. A fast algorithm for radiative transport in isotropic media. Journal of Computational Physics, 399(??):Article 108958, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119306631.

Soghrati:2015:HIE

[SA15] Soheil Soghrati and Hossein Ahmadian. 3D hierarchical interface-enriched finite element method: Implementation and applications. Journal of Computational Physics, 299 (??):45–55, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004337.

Schranner:2016:CII

[SA16] Felix S. Schranner and Nikolaus A. Adams. A conservative interface-interaction model with insoluble surfactant. Journal of Computational Physics, 327(??):653–677, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630479X.

Sitompul:2019:FCL

[SA19] Yos Panagaman Sitompul and Takayuki Aoki. A filtered cumulant lattice Boltzmann method for violent two-phase flows. Journal of Computational Physics, 390(??):93–120, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302529.

Schneider:2017:CNF

[SAEF17] Martin Schneider, Léo Agélas, Guillaume Enchéry, and Bernd Flemisch. Convergence of nonlinear finite volume schemes for heterogeneous anisotropic diffusion on general meshes. Journal of Computational Physics, 351(??):80–107, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306538.

Silberman:2019:NGV

[SAF^{+}19] Zachary J. Silberman, Thomas R. Adams, Joshua A. Faber, Zachariah B. Etienne, and Ian Ruchlin. Numerical generation of vector potentials from specified magnetic fields. Journal of Computational Physics, 379(??):421–437, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307940.

Semkiv:2017:TSM

[SAH17] Mykhailo Semkiv, Patrick D. Anderson, and Markus Hütter. Two-scale model for the effect of physical aging in elastomers filled with hard nanoparticles. Journal of Computational Physics, 350(??):184–206, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306393.

Seric:2018:DNS

[SAK18] Ivana Seric, Shahriar Afkhami, and Lou Kondic. Direct numerical simulation of variable surface tension flows using a volume-of-fluid method. Journal of Computational Physics, 352(??):615–636, January 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307465.

Sun:2017:ADM

[SAOW17] X. Sun, M. P. Ariza, M. Ortiz, and K. G. Wang. Acceleration of diffusive molecular dynamics simulations through mean field approximation and subcycling time integration. Journal of Computational Physics, 350(??):470–492, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306502.

Saye:2017:IMDa

[Say17a] Robert Saye. Implicit mesh discontinuous Galerkin methods and interfacial gauge methods for high-order accurate interface dynamics, with applications to surface tension dynamics, rigid body fluid-structure interaction, and free surface flow: Part I. Journal of Computational Physics, 344(??):647–682, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303649.

Saye:2017:IMDb

[Say17b] Robert Saye. Implicit mesh discontinuous Galerkin methods and interfacial gauge methods for high-order accurate interface dynamics, with applications to surface tension dynamics, rigid body fluid-structure interaction, and free surface flow: Part II. Journal of Computational Physics, 344(??):683–723, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303728.

Schram:2018:SRP

[SB18] R. D. Schram and G. T. Barkema. Simulation of ring polymer melts with GPU acceleration. Journal of Computational Physics, 363(??):128–139, June 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301098.

Seo:2017:MCM

[SBG^{+}17] Jung Hee Seo, Hani Bakhshaee, Guillaume Garreau, Chi Zhu, Andreas Andreou, William R. Thompson, and Rajat Mittal. A method for the computational modeling of the physics of heart murmurs. Journal of Computational Physics, 336(??):546–568, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301110.

Sokolova:2019:MFV

[SBH19] Irina Sokolova, Muhammad Gusti Bastisya, and Hadi Hajibeygi. Multiscale finite volume method for finite-volume-based simulation of poroelasticity. Journal of Computational Physics, 379(??):309–324, February 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307848.

Serino:2019:SAM

[SBHS19] D. A. Serino, J. W. Banks, W. D. Henshaw, and D. W. Schwendeman. A stable added-mass partitioned (AMP) algorithm for elastic solids and incompressible flow. Journal of Computational Physics, 399(??):Article 108923, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S002199911930628X.

Schaerer:2017:EAI

[SBT17] Roman Pascal Schaerer, Pratyuksh Bansal, and Manuel Torrilhon. Efficient algorithms and implementations of entropybased moment closures for rarefied gases. Journal of Computational Physics, 340(??):138–159, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301675.

Sousbie:2016:CPV

[SC16] Thierry Sousbie and Stéphane Colombi. ColDICE: a parallel Vlasov–Poisson solver using moving adaptive simplicial tessellation. Journal of Computational Physics, 321(??):644–697, September 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116301930.

Sevilla:2018:IPF

[SC18a] Ruben Sevilla and Régis Cottereau. Influence of periodically fluctuating material parameters on the stability of explicit high-order spectral element methods. Journal of Computational Physics, 373(??):304–323, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304558.

Stupfel:2018:HOT

[SC18b] Bruno Stupfel and Mathieu Chanaud. High-order transmission conditions in a domain decomposition method for the time-harmonic Maxwell’s equations in inhomogeneous media. Journal of Computational Physics, 372(??):385–405, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304285.

Stanier:2019:FIC

[SCC19] A. Stanier, L. Chacón, and G. Chen. A fully implicit, conservative, non-linear, electromagnetic hybrid particleion/fluid-electron algorithm. Journal of Computational Physics, 376(??):597–616, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306405.

Stadlbauer:2019:BEP

[SCE^{+}19] Benjamin Stadlbauer, Andrea Cossettini, José A. Morales E., Daniel Pasterk, Paolo Scarbolo, Leila Taghizadeh, Clemens Heitzinger, and Luca Selmi. Bayesian estimation of physical and geometrical parameters for nanocapacitor array biosensors. Journal of Computational Physics, 397(??):Article 108874, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305650.

Schneider:2016:KCLa

[Sch16a] Florian Schneider. Kershaw closures for linear transport equations in slab geometry I: Model derivation. Journal of Computational Physics, 322(??):905–919, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300596.

Schneider:2016:KCLb

[Sch16b] Florian Schneider. Kershaw closures for linear transport equations in slab geometry II: High-order realizabilitypreserving discontinuous-Galerkin schemes. Journal of Computational Physics, 322(??):920–935, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302959.

Shin:2018:HIT

[SCJ^{+}18] Seungwon Shin, Jalel Chergui, Damir Juric, Lyes Kahouadji, Omar K. Matar, and Richard V. Craster. A hybrid interface tracking — level set technique for multiphase flow with soluble surfactant. Journal of Computational Physics, 359(??):409–435, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300202.

Sellier:2019:SPN

[SCL19] Jean Michel Sellier, Gaétan Marceau Caron, and Jacob Leygonie. Signed particles and neural networks, towards efficient simulations of quantum systems. Journal of Computational Physics, 387(??):154–162, June 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301536.

Strychalski:2015:PIB

[SCLG15] Wanda Strychalski, Calina A. Copos, Owen L. Lewis, and Robert D. Guy. A poroelastic immersed boundary method with applications to cell biology. Journal of Computational Physics, 282(??):77–97, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006780.

Sanchez:2017:SHH

[SCN^{+}17] M. A. Sánchez, C. Ciuca, N. C. Nguyen, J. Peraire, and B. Cockburn. Symplectic Hamiltonian HDG methods for wave propagation phenomena. Journal of Computational Physics, 350(??):951–973, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306691.

Shi:2016:SSP

[SCQP16] Lingling Shi, Suncica Canić, Annalisa Quaini, and Tsorng-Whay Pan. A study of self-propelled elastic cylindrical microswimmers using modeling and computation. Journal of Computational Physics, 314(??):264–286, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001431.

Simmons:2016:HBE

[SCS16] Daniel Simmons, Kristof Cools, and Phillip Sewell. A hybrid Boundary Element Unstructured Transmission-line (BEUT) method for accurate 2D electromagnetic simulation. Journal of Computational Physics, 324(??):275–288, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116303424.

Sun:2018:DGM

[SCS18] Zheng Sun, José A. Carrillo, and Chi-Wang Shu. A discontinuous Galerkin method for nonlinear parabolic equations and gradient flow problems with interaction potentials. Journal of Computational Physics, 352(??):76–104, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307106.

Chen:2018:ASS

[sCYxL^{+}18] Shu sheng Chen, Chao Yan, Bo xi Lin, Li yuan Liu, and Jian Yu. Affordable shock-stable item for Godunov-type schemes against carbuncle phenomenon. Journal of Computational Physics, 373(??):662–672, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304832.

Sellier:2015:SIF

[SD15] J. M. Sellier and I. Dimov. On the simulation of indistinguishable fermions in the many-body Wigner formalism. Journal of Computational Physics, 280(??):287–294, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006627.

Stewart:2016:EPE

[SD16] Andrew L. Stewart and Paul J. Dellar. An energy and potential enstrophy conserving numerical scheme for the multilayer shallow water equations with complete Coriolis force. Journal of Computational Physics, 313(??):99–120, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500861X.

Shaw:2017:SME

[SD17] Jeremy A. Shaw and Dacian N. Daescu. Sensitivity of the model error parameter specification in weak-constraint fourdimensional variational data assimilation. Journal of Computational Physics, 343(??):115–129, August 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303297.

Sun:2018:IUA

[SD18] Guangrui Sun and Julian A. Domaradzki. Implicit LES using adaptive filtering. Journal of Computational Physics, 359(??):380–408, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118300196.

Silva:2017:NAQ

[SDFA17] J. J. B. Silva, G. C. Duarte-Filho, and F. A. G. Almeida. Numerical approach of the quantum circuit theory. Journal of Computational Physics, 333(??):427–439, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306829.

Smolarkiewicz:2016:FVM

[SDH^{+}16] Piotr K. Smolarkiewicz, Willem Deconinck, Mats Hamrud, Christian Kühnlein, George Mozdzynski, Joanna Szmelter, and Nils P. Wedi. A finite-volume module for simulating global all-scale atmospheric flows. Journal of Computational Physics, 314(??):287–304, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001674.

Sen:2015:ECB

[SDJU15] Oishik Sen, Sean Davis, Gustaaf Jacobs, and H. S. Udaykumar. Evaluation of convergence behavior of metamodeling techniques for bridging scales in multi-scale multimaterial simulation. Journal of Computational Physics, 294(??):585–604, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001989.

Sibra:2017:SRP

[SDM^{+}17] A. Sibra, J. Dupays, A. Murrone, F. Laurent, and M. Massot. Simulation of reactive polydisperse sprays strongly coupled to unsteady flows in solid rocket motors: Efficient strategy using Eulerian multi-fluid methods. Journal of Computational Physics, 339(??):210–246, June 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300967.

Sauer:2017:SFE

[SDMS17] Roger A. Sauer, Thang X. Duong, Kranthi K. Mandadapu, and David J. Steigmann. A stabilized finite element formulation for liquid shells and its application to lipid bilayers. Journal of Computational Physics, 330(??):436–466, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305836.

Song:2016:CFE

[SDW16] Fei Song, Weibing Deng, and Haijun Wu. A combined finite element and oversampling multiscale Petrov–Galerkin method for the multiscale elliptic problems with singularities. Journal of Computational Physics, 305(??):722–743, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007536.

Su:2018:DPP

[SDW18] Shuai Su, Qiannan Dong, and Jiming Wu. A decoupled and positivity-preserving discrete duality finite volume scheme for anisotropic diffusion problems on general polygonal meshes. Journal of Computational Physics, 372(??):773–798, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304303.

Stasyszyn:2015:VPI

[SE15] Federico A. Stasyszyn and Detlef Elstner. A vector potential implementation for smoothed particle magnetohydrodynamics. Journal of Computational Physics, 282(??):148–156, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007633.

Sousedik:2016:SGM

[SE16] Bedrich Sousedík and Howard C. Elman. Stochastic Galerkin methods for the steady-state Navier–Stokes equations. Journal of Computational Physics, 316(??):435–452, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300456.

Sellier:2015:SPF

[Sel15] Jean Michel Sellier. A signed particle formulation of nonrelativistic quantum mechanics. Journal of Computational Physics, 297(??):254–265, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003708.

Stinchcombe:2016:EMS

[SF16] Adam R. Stinchcombe and Daniel B. Forger. An efficient method for simulation of noisy coupled multi-dimensional oscillators. Journal of Computational Physics, 321(??):932– 946, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116301620.

Setukha:2018:MRB

[SF18a] Alexey Setukha and Sergey Fetisov. The method of relocation of boundary condition for the problem of electromagnetic wave scattering by perfectly conducting thin objects. Journal of Computational Physics, 373(??):631–647, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304741.

Shankar:2018:HBS

[SF18b] Varun Shankar and Aaron L. Fogelson. Hyperviscosity-based stabilization for radial basis function–finite difference (RBF– FD) discretizations of advection–diffusion equations. Journal of Computational Physics, 372(??):616–639, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304145.

Sommer:2015:CDC

[SFDE15] A. Sommer, O. Farle, and R. Dyczij-Edlinger. Certified dual-corrected radiation patterns of phased antenna arrays by offline-online order reduction of finite-element models. Journal of Computational Physics, 299(??):22–44, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004210.

Strobl:2016:ECO

[SFP16] Severin Strobl, Arno Formella, and Thorsten Pöschel. Exact calculation of the overlap volume of spheres and mesh elements. Journal of Computational Physics, 311(??):158–172, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000577.

Shashkin:2016:CCS

[SFT16] Vladimir Shashkin, Rostislav Fadeev, and Mikhail Tolstykh. 3D conservative cascade semi-Lagrangian transport scheme using reduced latitude-longitude grid (CCS-RG). Journal of Computational Physics, 305(??):700–721, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007391.

Soize:2016:DDP

[SG16] C. Soize and R. Ghanem. Data-driven probability concentration and sampling on manifold. Journal of Computational Physics, 321(??):242–258, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301899.

Soize:2017:PCR

[SG17] C. Soize and R. Ghanem. Polynomial chaos representation of databases on manifolds. Journal of Computational Physics, 335(??):201–221, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300475.

Scarnati:2018:JIF

[SG18] Theresa Scarnati and Anne Gelb. Joint image formation and two-dimensional autofocusing for synthetic aperture radar data. Journal of Computational Physics, 374(??):803–821, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305242.

Sadr:2019:TLR

[SG19a] Mohsen Sadr and M. Hossein Gorji. Treatment of long-range interactions arising in the Enskog–Vlasov description of dense fluids. Journal of Computational Physics, 378(??):129–142, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911830723X.

Shi:2019:FOA

[SG19b] YuFeng Shi and Yan Guo. A fifth order alternative Compact-WENO finite difference scheme for compressible Euler equations. Journal of Computational Physics, 397(??):Article 108873, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305649.

Spill:2015:HAM

[SGA^{+}15] Fabian Spill, Pilar Guerrero, Tomas Alarcon, Philip K. Maini, and Helen Byrne. Hybrid approaches for multiple-species stochastic reaction-diffusion models. Journal of Computational Physics, 299(??):429–445, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004477.

Sen:2017:EKB

[SGC^{+}17] Oishik Sen, Nicholas J. Gaul, K. K. Choi, Gustaaf Jacobs, and H. S. Udaykumar. Evaluation of kriging based surrogate models constructed from mesoscale computations of shock interaction with particles. Journal of Computational Physics, 336(??):235–260, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300621.

Sen:2018:EMS

[SGC^{+}18a] Oishik Sen, Nicholas J. Gaul, K. K. Choi, Gustaaf Jacobs, and H. S. Udaykumar. Evaluation of multifidelity surrogate modeling techniques to construct closure laws for drag in shock-particle interactions. Journal of Computational Physics, 371(??):434–451, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303486.

Shipton:2018:HOC

[SGC18b] J. Shipton, T. H. Gibson, and C. J. Cotter. Higher-order compatible finite element schemes for the nonlinear rotating shallow water equations on the sphere. Journal of Computational Physics, 375(??):1121–1137, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305503.

Skarysz:2018:IIR

[SGD18] M. Skarysz, A. Garmory, and M. Dianat. An iterative interface reconstruction method for PLIC in general convex grids as part of a Coupled Level Set Volume of Fluid solver. Journal of Computational Physics, 368(??):254–276, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302699.

Stone:2017:ADE

[SGL17] D. Stone, S. Geiger, and G. J. Lord. Asynchronous discrete event schemes for PDEs. Journal of Computational Physics, 342(??):161–176, August 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302978.

Schneck:2018:OEF

[SGL18] William C. Schneck, Elizabeth D. Gregory, and Cara A. C. Leckey. Optimization of elastodynamic finite integration technique on Intel Xeon Phi Knights Landing processors. Journal of Computational Physics, 374(??):550–562, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305114.

Schneiders:2016:ECC

[SGMS16] Lennart Schneiders, Claudia Günther, Matthias Meinke, and Wolfgang Schröder. An efficient conservative cut-cell method for rigid bodies interacting with viscous compressible flows. Journal of Computational Physics, 311(??):62–86, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000346.

Sanchez:2016:RCM

[SGN16] J. Sánchez, F. Garcia, and M. Net. Radial collocation methods for the onset of convection in rotating spheres. Journal of Computational Physics, 308(??):273–288, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008591.

Sanders:2017:CSI

[SGP17a] Toby Sanders, Anne Gelb, and Rodrigo B. Platte. Composite SAR imaging using sequential joint sparsity. Journal of Computational Physics, 338(??):357–370, June 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301869.

Sun:2017:RTN

[SGP17b] Dan Sun, Andrew Garmory, and Gary J. Page. A robust two-node, 13 moment quadrature method of moments for dilute particle flows including wall bouncing. Journal of Computational Physics, 330(??):493–509, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306118.

Soize:2019:EBC

[SGS^{+}19] C. Soize, R. Ghanem, C. Safta, X. Huan, Z. P. Vane, J. Oefelein, G. Lacaze, H. N. Najm, Q. Tang, and X. Chen. Entropy-based closure for probabilistic learning on manifolds.

Journal of Computational Physics, 388(??):518–533, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300105.

Stein:2016:IBS

[SGT16] David B. Stein, Robert D. Guy, and Becca Thomases. Immersed boundary smooth extension: a high-order method for solving PDE on arbitrary smooth domains using Fourier spectral methods. Journal of Computational Physics, 304 (??):252–274, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115006877.

Stein:2017:IBS

[SGT17] David B. Stein, Robert D. Guy, and Becca Thomases. Immersed Boundary Smooth Extension (IBSE): a high-order method for solving incompressible flows in arbitrary smooth domains. Journal of Computational Physics, 335(??):155–178, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300207.

Sirajuddin:2019:TFS

[SH19] David Sirajuddin and William N. G. Hitchon. A truly forward semi-Lagrangian WENO scheme for the Vlasov–Poisson system. Journal of Computational Physics, 392(??):619– 665, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303079.

Schranner:2016:CWC

[SHA16] Felix S. Schranner, Xiangyu Hu, and Nikolaus A. Adams. On the convergence of the weakly compressible sharpinterface method for two-phase flows. Journal of Computational Physics, 324(??):94–114, November 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303321.

Shahbazi:2017:RSO

[Sha17a] Khosro Shahbazi. Robust second-order scheme for multiphase flow computations. Journal of Computational Physics, 339(??):163–178, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302206.

Shankar:2017:ORB

[Sha17b] Varun Shankar. The overlapped radial basis function–finite difference (RBF–FD) method: a generalization of RBF–FD. Journal of Computational Physics, 342(??):211–228, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303169.

Shim:2017:PLB

[Shi17] Jae Wan Shim. Parametric lattice Boltzmann method. Journal of Computational Physics, 338(??):240–251, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301602.

Shirangi:2019:CLF

[Shi19] Mehrdad G. Shirangi. Closed-loop field development with multipoint geostatistics and statistical performance assessment. Journal of Computational Physics, 390(??):249–264, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302360.

Seol:2016:IBM

[SHKL16] Yunchang Seol, Wei-Fan Hu, Yongsam Kim, and Ming-Chih Lai. An immersed boundary method for simulating vesicle dynamics in three dimensions. Journal of Computational Physics, 322(??):125–141, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302601.

Schneider:2015:SAN

[SHLG15] Nicolas Schneider, Zohra Hammouch, Gérard Labrosse, and Serge Gauthier. A spectral anelastic Navier–Stokes solver for a stratified two-miscible-layer system in infinite horizontal channel. Journal of Computational Physics, 299(??):374–403, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004428.

Stotsky:2016:VVD

[SHP^{+}16] Jay A. Stotsky, Jason F. Hammond, Leonid Pavlovsky, Elizabeth J. Stewart, John G. Younger, Michael J. Solomon, and David M. Bortz. Variable viscosity and density biofilm simulations using an immersed boundary method, part II: Experimental validation and the heterogeneous rheology-IBM. Journal of Computational Physics, 317(??):204–222, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300730.

Stinis:2019:ECI

[SHTY19] Panos Stinis, Tobias Hagge, Alexandre M. Tartakovsky, and Enoch Yeung. Enforcing constraints for interpolation and extrapolation in Generative Adversarial Networks. Journal of Computational Physics, 397(??):Article 108844, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305285.

Shu:2016:HOW

[Shu16] Chi-Wang Shu. High order WENO and DG methods for time-dependent convection-dominated PDEs: a brief survey of several recent developments. Journal of Computational Physics, 316(??):598–613, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300766.

Spietz:2017:IBP

[SHW17] Henrik Juul Spietz, Mads Mølholm Hejlesen, and Jens Honoré Walther. Iterative Brinkman penalization for simulation of impulsively started flow past a sphere and a circular disc. Journal of Computational Physics, 336(??):261–274, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300803.

Spietz:2018:RMS

[SHW18] Henrik Juul Spietz, Mads Mølholm Hejlesen, and Jens Honoré Walther. A regularization method for solving the Poisson equation for mixed unbounded-periodic domains. Journal of Computational Physics, 356(??):439–447, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117309038.

Sidilkover:2018:TUV

[Sid18] David Sidilkover. Towards unification of the Vorticity Confinement and Shock Capturing (TVD and ENO/WENO) methods. Journal of Computational Physics, 358(??):235–255, April 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117309270.

Sugiura:2016:EGS

[SiI16] Keisuke Sugiura and Shu ichiro Inutsuka. An extension of Godunov SPH: Application to negative pressure media. Journal of Computational Physics, 308(??):171–197, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008499.

Sugiura:2017:EGS

[SiI17] Keisuke Sugiura and Shu ichiro Inutsuka. An extension of Godunov SPH II: Application to elastic dynamics. Journal of Computational Physics, 333(??):78–103, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306805.

Siryk:2019:NAG

[Sir19] Sergii V. Siryk. A note on the application of the Guermond– Pasquetti mass lumping correction technique for convection– diffusion problems. Journal of Computational Physics, 376 (??):1273–1291, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118306739.

Sun:2016:BVD

[SIX16] Ziyao Sun, Satoshi Inaba, and Feng Xiao. Boundary Variation Diminishing (BVD) reconstruction: a new approach to improve Godunov schemes. Journal of Computational Physics, 322(??):309–325, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302765.

Sun:2015:PML

[SJH^{+}15] X. F. Sun, Z. H. Jiang, X. W. Hu, G. Zhuang, J. F. Jiang, and W. X. Guo. Perfectly matched layer absorbing boundary condition for nonlinear two-fluid plasma equations. Journal of Computational Physics, 286(??):128–142, April 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000376.

Sun:2015:APUa

[SJX15] Wenjun Sun, Song Jiang, and Kun Xu. An asymptotic preserving unified gas kinetic scheme for gray radiative transfer equations. Journal of Computational Physics, 285(??):265–279, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000121.

Sun:2017:MUG

[SJX17] Wenjun Sun, Song Jiang, and Kun Xu. A multidimensional unified gas-kinetic scheme for radiative transfer equations on unstructured mesh. Journal of Computational Physics, 351(??):455–472, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306964.

Sun:2015:APUb

[SJXL15] Wenjun Sun, Song Jiang, Kun Xu, and Shu Li. An asymptotic preserving unified gas kinetic scheme for frequencydependent radiative transfer equations. Journal of Computational Physics, 302(??):222–238, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005823.

Shevchenko:2015:ABC

[SK15a] Igor Shevchenko and Barbara Kaltenbacher. Absorbing boundary conditions for nonlinear acoustics: the Westervelt equation. Journal of Computational Physics, 302(??):200–221, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005811.

Sun:2015:NSH

[SK15b] Yifei Sun and Mrinal Kumar. A numerical solver for high dimensional transient Fokker–Planck equation in modeling polymeric fluids. Journal of Computational Physics, 289(??):149– 168, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000960.

Su:2018:ICC

[SK18] Yunde Su and Seung Hyun Kim. An improved consistent, conservative, non-oscillatory and high order finite difference scheme for variable density low Mach number turbulent flow simulation. Journal of Computational Physics, 372(??):202–219, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303991.

Saini:2019:NCM

[SK19a] N. Saini and C. Kleinstreuer. A new collision model for ellipsoidal particles in shear flow. Journal of Computational Physics, 376(??):1028–1050, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306417.

Song:2019:FMH

[SK19b] Fangying Song and George Em Karniadakis. Fractional magneto-hydrodynamics: Algorithms and applications. Journal of Computational Physics, 378(??):44–62, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307149.

Suchde:2019:FLM

[SK19c] Pratik Suchde and Jörg Kuhnert. A fully Lagrangian meshfree framework for PDEs on evolving surfaces. Journal of Computational Physics, 395(??):38–59, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304395.

Subbareddy:2017:SCB

[SKC17] Pramod K. Subbareddy, Anand Kartha, and Graham V. Candler. Scalar conservation and boundedness in simulations of compressible flow. Journal of Computational Physics, 348(??):827–846, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305727.

Schwarz:2015:TDS

[SKF15] S. Schwarz, T. Kempe, and J. Fröhlich. A temporal discretization scheme to compute the motion of light particles in viscous flows by an immersed boundary method. Journal of Computational Physics, 281(??):591–613, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007219.

Schwarz:2016:IBM

[SKF16] Stephan Schwarz, Tobias Kempe, and Jochen Fröhlich. An immersed boundary method for the simulation of bubbles with varying shape. Journal of Computational Physics, 315(??):124– 149, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000413.

Smolarkiewicz:2017:FVM

[SKG17] Piotr K. Smolarkiewicz, Christian Kühnlein, and Wojciech W. Grabowski. A finite-volume module for cloud-resolving simulations of global atmospheric flows. Journal of Computational Physics, 341(??):208–229, July 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302796.

Shiroto:2017:FVC

[SKO17] Takashi Shiroto, Soshi Kawai, and Naofumi Ohnishi. Finitevolume-concept-based Padé-type filters. Journal of Computational Physics, 349(??):215–219, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305983.

Shiroto:2018:SPO

[SKO18] Takashi Shiroto, Soshi Kawai, and Naofumi Ohnishi. Structurepreserving operators for thermal-nonequilibrium hydrodynamics. Journal of Computational Physics, 364(??):1–17, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301608.

Sadovskyy:2015:SLS

[SKP^{+}15] I. A. Sadovskyy, A. E. Koshelev, C. L. Phillips, D. A. Karpeyev, and A. Glatz. Stable large-scale solver for Ginzburg–Landau equations for superconductors. Journal of Computational Physics, 294(??):639–654, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002284.

Selim:2017:IAR

[SKS17] Mohamed M. Selim, Roy P. Koomullil, and Ahmed S. Shehata. Incremental approach for radial basis functions mesh deformation with greedy algorithm. Journal of Computational Physics, 340(??):556–574, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302322.

Smolarkiewicz:2019:SII

[SKW19] Piotr K. Smolarkiewicz, Christian Kühnlein, and Nils P. Wedi. Semi-implicit integrations of perturbation equations for all-scale atmospheric dynamics. Journal of Computational Physics, 376(??):145–159, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830634X.

Sheu:2015:DRE

[SL15] Tony W. H. Sheu and Le Lin. Dispersion relation equation preserving FDTD method for nonlinear cubic Schrödinger equation. Journal of Computational Physics, 299(??):1–21, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004209.

Shlivinski:2016:GWF

[SL16a] A. Shlivinski and V. Lomakin. Gaussian-windowed frame based method of moments formulation of surface-integral-equation for extended apertures. Journal of Computational Physics, 308(??):289–304, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115008475.

Stupfel:2016:OWD

[SL16b] Bruno Stupfel and Matthieu Lecouvez. One-way domain decomposition method with exact radiation condition and fast GMRES solver for the solution of Maxwell’s equations. Journal of Computational Physics, 322(??):882–904, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302960.

Su:2016:EDP

[SL16c] Hongling Su and Shengtai Li. Energy/dissipation-preserving Birkhoffian multi-symplectic methods for Maxwell’s equations with dissipation terms. Journal of Computational Physics, 311(??):213–240, April 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116000541.

Schroeder:2017:SDF

[SL17] Philipp W. Schroeder and Gert Lube. Stabilised dG–FEM for incompressible natural convection flows with boundary and moving interior layers on non-adapted meshes. Journal of Computational Physics, 335(??):760–779, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300712.

Semplice:2018:AMR

[SL18] Matteo Semplice and Raphaël Loubère. Adaptive-mesh-refinement for hyperbolic systems of conservation laws based on a posteriori stabilized high order polynomial reconstructions. Journal of Computational Physics, 354(??):86–110, February 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307842.

Shen:2019:GEI

[SL19a] Xuefeng Shen and Melvin Leok. Geometric exponential integrators. Journal of Computational Physics, 382(??):27–42, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300221.

Shi:2019:LDG

[SL19b] Hui Shi and Ying Li. Local discontinuous Galerkin methods with implicit-explicit multistep time-marching for solving the nonlinear Cahn–Hilliard equation. Journal of Computational Physics, 394(??):719–731, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303845.

Smyl:2019:LOM

[SL19c] Danny Smyl and Dong Liu. Less is often more: Applied inverse problems using hp-forward models. Journal of Computational Physics, 399(??):Article 108949, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306540.

Slattery:2016:MFD

[Sla16] Stuart R. Slattery. Mesh-free data transfer algorithms for partitioned multiphysics problems: Conservation, accuracy, and parallelism. Journal of Computational Physics, 307 (??):164–188, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115008037.

Sousa:2019:FDM

[SLA^{+}19] F. S. Sousa, C. F. Lages, J. L. Ansoni, A. Castelo, and A. Simao. A finite difference method with meshless interpolation for incompressible flows in non-graded tree-based grids. Journal of Computational Physics, 396(??):848–866, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304966.

Shinde:2016:GFM

[SLB^{+}16] Vilas Shinde, Elisabeth Longatte, Franck Baj, Yannick Hoarau, and Marianna Braza. A Galerkin-free model reduction approach for the Navier–Stokes equations. Journal of Computational Physics, 309(??):148–163, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008700.

Shinde:2019:GFM

[SLB^{+}19] Vilas Shinde, Elisabeth Longatte, Franck Baj, Yannick Hoarau, and Marianna Braza. Galerkin-free model reduction for fluid-structure interaction using proper orthogonal decomposition. Journal of Computational Physics, 396(??):579–595, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304826.

Shao:2018:CFI

[SLC^{+}18] Changxiao Shao, Kun Luo, Min Chai, Haiou Wang, and Jianren Fan. A computational framework for interface-resolved DNS of simultaneous atomization, evaporation and combustion. Journal of Computational Physics, 371(??):751–778, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303899.

Spandan:2018:FML

[SLdTV18] Vamsi Spandan, Detlef Lohse, Marco D. de Tullio, and Roberto Verzicco. A fast moving least squares approximation with adaptive Lagrangian mesh refinement for large scale immersed boundary simulations. Journal of Computational Physics, 375(??):228–239, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305710.

Schneider:2018:EBS

[SLH18] Moritz Schneider, Jens Lang, and Willem Hundsdorfer. Extrapolation-based super-convergent implicit-explicit Peer methods with A-stable implicit part. Journal of Computational Physics, 367(??):121–133, August 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302201.

Shin:2016:FSO

[SLL16] Jaemin Shin, Hyun Geun Lee, and June-Yub Lee. First and second order numerical methods based on a new convex splitting for phase-field crystal equation. Journal of Computational Physics, 327(??):519–542, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304740.

Shin:2017:USM

[SLL17] Jaemin Shin, Hyun Geun Lee, and June-Yub Lee. Unconditionally stable methods for gradient flow using Convex Splitting Runge–Kutta scheme. Journal of Computational Physics, 347(??):367–381, October 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305089.

Santos:2019:ONM

[SLL19] Tatiane P. Santos, Leandro R. F. Lima, and Caio H. Lewenkopf. An order N numerical method to efficiently calculate the transport properties of large systems: an algorithm optimized for sparse linear solvers. Journal of Computational Physics, 394(??):440–455, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303729.

Sorgentone:2015:NHO

[SLN15] Chiara Sorgentone, Cristina La Cognata, and Jan Nordström. A new high order energy and enstrophy conserving Arakawa-like Jacobian differential operator. Journal of Computational Physics, 301(??):167–177, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005550.

Sheikh:2016:ASL

[SLR^{+}16] A. H. Sheikh, D. Lahaye, L. Garcia Ramos, R. Nabben, and C. Vuik. Accelerating the shifted Laplace preconditioner for the Helmholtz equation by multilevel deflation. Journal of Computational Physics, 322(??):473–490, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302509.

Simpson:2018:IBE

[SLVE18] R. N. Simpson, Z. Liu, R. Vázquez, and J. A. Evans. An isogeometric boundary element method for electromagnetic scattering with compatible B-spline discretizations. Journal of Computational Physics, 362(??):264–289, June 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300354.

Sifounakis:2016:CFV

[SLY16] Adamandios Sifounakis, Sangseung Lee, and Donghyun You. A conservative finite volume method for incompressible Navier– Stokes equations on locally refined nested Cartesian grids. Journal of Computational Physics, 326(??):845–861, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304405.

Sun:2017:FSD

[SLZ^{+}17] HongGuang Sun, Xiaoting Liu, Yong Zhang, Guofei Pang, and Rhiannon Garrard. A fast semi-discrete Kansa method to solve the two-dimensional spatiotemporal fractional diffusion equation. Journal of Computational Physics, 345 (??):74–90, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117303819.

Shamasundar:2016:IAM

[SM16] R. Shamasundar and W. A. Mulder. Improving the accuracy of mass-lumped finite-elements in the first-order formulation of the wave equation by defect correction. Journal of Computational Physics, 322(??):689–707, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302856.

Simon:2019:SCN

[SM19a] Sangeeth Simon and J. C. Mandal. A simple cure for numerical shock instability in the HLLC Riemann solver. Journal of Computational Physics, 378(??):477–496, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307484.

Subramanian:2019:EEC

[SM19b] Abhinav Subramanian and Sankaran Mahadevan. Error estimation in coupled multi-physics models. Journal of Computational Physics, 395(??):19–37, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304206.

Siguenza:2016:VIT

[SMA^{+}16] J. Sigüenza, S. Mendez, D. Ambard, F. Dubois, F. Jourdan, R. Mozul, and F. Nicoud. Validation of an immersed thick boundary method for simulating fluid-structure interactions of deformable membranes. Journal of Computational Physics, 322(??):723–746, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302662.

Spellings:2017:GAD

[SMAG17] Matthew Spellings, Ryan L. Marson, Joshua A. Anderson, and Sharon C. Glotzer. GPU accelerated Discrete Element Method (DEM) molecular dynamics for conservative, faceted particle simulations. Journal of Computational Physics, 334(??):460– 467, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300244.

Sabzikar:2015:TFC

[SMC15] Farzad Sabzikar, Mark M. Meerschaert, and Jinghua Chen. Tempered fractional calculus. Journal of Computational Physics, 293(??):14–28, July 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114002873.

Sharan:2018:MSB

[SMD18a] Nek Sharan, Georgios Matheou, and Paul E. Dimotakis. Mixing, scalar boundedness, and numerical dissipation in large-eddy simulations. Journal of Computational Physics, 369(??):148–172, September 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302973.

Slevinsky:2018:SMN

[SMD18b] Richard Mikaël Slevinsky, Hadrien Montanelli, and Qiang Du. A spectral method for nonlocal diffusion operators on the sphere. Journal of Computational Physics, 372(??):893–911, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304029.

Shao:2019:EMS

[SMG19] Q. Shao, S. K. Matthäi, and L. Gross. Efficient modelling of solute transport in heterogeneous media with discrete event simulation. Journal of Computational Physics, 384(??):134– 150, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300671.

Smith:2018:NND

[Smi18] David J. Smith. A nearest-neighbour discretisation of the regularized stokeslet boundary integral equation. Journal of Computational Physics, 358(??):88–102, April 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308847.

Soghrati:2015:BCM

[SMLB15] Soheil Soghrati, Weijie Mai, Bowen Liang, and Rudolph G. Buchheit. A boundary collocation meshfree method for the treatment of Poisson problems with complex morphologies. Journal of Computational Physics, 281(??):225–236, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007128.

Spandan:2017:PIP

[SMOM^{+}17] Vamsi Spandan, Valentina Meschini, Rodolfo Ostilla-Mónico, Detlef Lohse, Giorgio Querzoli, Marco D. de Tullio, and Roberto Verzicco. A parallel interaction potential approach coupled with the immersed boundary method for fully resolved simulations of deformable interfaces and membranes. Journal of Computational Physics, 348(??):567–590, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305442.

Sheldon:2016:HDG

[SMP16] Jason P. Sheldon, Scott T. Miller, and Jonathan S. Pitt. A hybridizable discontinuous Galerkin method for modeling fluid-structure interaction. Journal of Computational Physics, 326(??):91–114, December 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303941.

Serson:2016:VCS

[SMS16] D. Serson, J. R. Meneghini, and S. J. Sherwin. Velocitycorrection schemes for the incompressible Navier–Stokes equations in general coordinate systems. Journal of Computational Physics, 316(??):243–254, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300687.

Song:2018:SBM

[SMSR18] T. Song, A. Main, G. Scovazzi, and M. Ricchiuto. The shifted boundary method for hyperbolic systems: Embedded domain computations of linear waves and shallow water flows. Journal of Computational Physics, 369(??):45–79, September 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302857.

Shah:2016:MRS

[SMT^{+}16] Swej Shah, Olav Møyner, Matei Tene, Knut-Andreas Lie, and Hadi Hajibeygi. The multiscale restriction smoothed basis method for fractured porous media (F-MsRSB). Journal of Computational Physics, 318(??):36–57, August 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301267.

Sato:2015:DML

[SN15] Yohei Sato and Bojan Niceno. A depletable micro-layer model for nucleate pool boiling. Journal of Computational Physics, 300(??):20–52, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500491X.

Svard:2019:CRE

[SN19] Magnus Svärd and Jan Nordström. On the convergence rates of energy-stable finite-difference schemes. Journal of Computational Physics, 397(??):Article 108819, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305030.

Samaras:2015:URL

[SNB^{+}15] Stefanos Samaras, Doina Nicolae, Christine Böckmann, Jeni Vasilescu, Ioannis Binietoglou, Lev Labzovskii, Florica Toanca, and Alexandros Papayannis. Using Raman-lidar-based regularized microphysical retrievals and Aerosol Mass Spectrometer measurements for the characterization of biomass burning aerosols. Journal of Computational Physics, 299(??):156–174, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004441.

Shankar:2018:RLA

[SNK18] Varun Shankar, Akil Narayan, and Robert M. Kirby. RBF-LOI: Augmenting radial basis functions (RBFs) with least orthogonal interpolation (LOI) for solving PDEs on surfaces. Journal of Computational Physics, 373(??):722–735, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304765.

Safdari:2016:NBG

[SNSG16] Masoud Safdari, Ahmad R. Najafi, Nancy R. Sottos, and Philippe H. Geubelle. A NURBS-based generalized finite element scheme for 3D simulation of heterogeneous materials. Journal of Computational Physics, 318(??):373–390, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301310.

Starinshak:2015:SRA

[SO15] D. P. Starinshak and J. M. Owen. A subzone reconstruction algorithm for efficient staggered compatible remapping. Journal of Computational Physics, 296(??):263–292, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003034.

Starinshak:2016:MES

[SO16] D. P. Starinshak and J. M. Owen. A multimaterial extension to subzonal reconstruction. Journal of Computational Physics, 313(??):594–616, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008049.

Slevinsky:2017:FWC

[SO17] Richard Mikael Slevinsky and Sheehan Olver. A fast and well-conditioned spectral method for singular integral equations. Journal of Computational Physics, 332(??):290–315, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306507.

Shiroto:2019:QCS

[SOS19] Takashi Shiroto, Naofumi Ohnishi, and Yasuhiko Sentoku. Quadratic conservative scheme for relativistic Vlasov–Maxwell system. Journal of Computational Physics, 379(??):32–50, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307083.

Sovinec:2016:SNI

[Sov16] C. R. Sovinec. Stabilization of numerical interchange in spectral-element magnetohydrodynamics. Journal of Computational Physics, 319(??):61–78, August 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301383.

Schmitt:2015:RLE

[SP15a] Christoph Schmitt and Heinz Pitsch. Reactive linearized equations of perturbed compressible variables for low-Mach number variable-density flows. Journal of Computational Physics, 281(??):1–27, January 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114006895.

Stevens:2015:RBF

[SP15b] D. Stevens and H. Power. The radial basis function finite collocation approach for capturing sharp fronts in time dependent advection problems. Journal of Computational Physics, 298(??):423–445, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003666.

Shamir:2016:GBS

[SP16a] Ofer Shamir and Nathan Paldor. A Gegenbauer-based shallow water solver for a thick “ocean” over a rotating sphere. Journal of Computational Physics, 304(??):487–505, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500683X.

Sierakowski:2016:RPS

[SP16b] Adam J. Sierakowski and Andrea Prosperetti. Resolvedparticle simulation by the Physalis method: Enhancements and new capabilities. Journal of Computational Physics, 309(??):164–184, March 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115008827.

Sturdevant:2016:FTS

[SP16c] Benjamin J. Sturdevant and Scott E. Parker. Finite time step and spatial grid effects in δf simulation of warm plasmas. Journal of Computational Physics, 305(??):647–663, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007329.

Salmasi:2018:DEC

[SP18] Mahbod Salmasi and Michael Potter. Discrete exterior calculus approach for discretizing Maxwell’s equations on face-centered cubic grids for FDTD. Journal of Computational Physics, 364(??):298–313, July 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301773.

Shawki:2019:PMS

[SP19a] Karim Shawki and George Papadakis. A preconditioned Multiple Shooting Shadowing algorithm for the sensitivity analysis of chaotic systems. Journal of Computational Physics, 398(??):Article 108861, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305455.

Shen:2019:RFC

[SP19b] Hua Shen and Matteo Parsani. A rezoning-free CESE scheme for solving the compressible Euler equations on moving unstructured meshes. Journal of Computational Physics, 397(??):Article 108858, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930542X.

Straub:2016:BAR

[SPB16] Daniel Straub, Iason Papaioannou, and Wolfgang Betz. Bayesian analysis of rare events. Journal of Computational Physics, 314(??):538–556, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001704.

Schmidt:2017:KBL

[SPB17] Michael J. Schmidt, Stephen Pankavich, and David A. Benson. A kernel-based Lagrangian method for imperfectlymixed chemical reactions. Journal of Computational Physics, 336(??):288–307, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117301055.

Sharan:2018:TSO

[SPB18] Nek Sharan, Carlos Pantano, and Daniel J. Bodony. Timestable overset grid method for hyperbolic problems using summation-by-parts operators. Journal of Computational Physics, 361(??):199–230, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300597.

Sturdevant:2016:IPC

[SPCH16] Benjamin J. Sturdevant, Scott E. Parker, Yang Chen, and Benjamin B. Hause. An implicit δf particle-in-cell method with sub-cycling and orbit averaging for Lorentz ions. Journal of Computational Physics, 316(??):519–533, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300821.

Schmidmayer:2017:MNM

[SPD^{+}17] Kevin Schmidmayer, Fabien Petitpas, Eric Daniel, Nicolas Favrie, and Sergey Gavrilyuk. A model and numerical method for compressible flows with capillary effects. Journal of Computational Physics, 334(??):468–496, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300116.

Schmidmayer:2019:AMR

[SPD19] Kevin Schmidmayer, Fabien Petitpas, and Eric Daniel. Adaptive mesh refinement algorithm based on dual trees for cells and faces for multiphase compressible flows. Journal of Computational Physics, 388(??):252–278, July 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301949.

Spence:2015:GAO

[Spe15] Peter J. Spence. The generation of arbitrary order, nonclassical, Gauss-type quadrature for transport applications. Journal of Computational Physics, 296(??):25–57, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002958.

Snedden:2015:NMS

[SPM^{+}15] Ali Snedden, Lara Arielle Phillips, Grant J. Mathews, Jared Coughlin, In-Saeng Suh, and Aparna Bhattacharya. A new multi-scale structure finding algorithm to identify cosmological structure. Journal of Computational Physics, 299(??):92–97, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004490.

Stoll:2016:FSO

[SPM16] Martin Stoll, John W. Pearson, and Philip K. Maini. Fast solvers for optimal control problems from pattern formation. Journal of Computational Physics, 304(??):27–45, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006683.

Stern:2019:NFT

[SPN^{+}19] Eric C. Stern, Savio Poovathingal, Ioannis Nompelis, Thomas E. Schwartzentruber, and Graham V. Candler. Nonequilibrium flow through porous thermal protection materials, Part I: Numerical methods. Journal of Computational Physics, 380(??):408–426, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117306708.

Sandhu:2016:BIN

[SPP^{+}16a] Rimple Sandhu, Dominique Poirel, Chris Pettit, Mohammad Khalil, and Abhijit Sarkar. Bayesian inference of nonlinear unsteady aerodynamics from aeroelastic limit cycle oscillations. Journal of Computational Physics, 316(??):534–557, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001571.

Smith:2016:RLN

[SPP16b] Timothy A. Smith, David J. Petty, and Carlos Pantano. A Roe-like numerical method for weakly hyperbolic systems of equations in conservation and non-conservation form. Journal of Computational Physics, 316(??):117–138, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300389.

Stokes:2015:ENS

[SPRW15] Peter W. Stokes, Bronson Philippa, Wayne Read, and Ronald D. White. Efficient numerical solution of the time fractional diffusion equation by mapping from its Brownian counterpart. Journal of Computational Physics, 282(??):334–344, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007852.

Singh:2018:AJN

[SPW18] Gurpreet Singh, Gergina Pencheva, and Mary F. Wheeler. An approximate Jacobian nonlinear solver for multiphase flow and transport. Journal of Computational Physics, 375(??):337–351, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305746.

Salinas:2018:DCV

[SPX^{+}18] P. Salinas, D. Pavlidis, Z. Xie, H. Osman, C. C. Pain, and M. D. Jackson. A discontinuous control volume finite element method for multi-phase flow in heterogeneous porous media. Journal of Computational Physics, 352(??):602–614, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307313.

Srinivasan:2018:PPH

[SPZ18] Sashank Srinivasan, Jonathan Poggie, and Xiangxiong Zhang. A positivity-preserving high order discontinuous Galerkin scheme for convection-diffusion equations. Journal of Computational Physics, 365(??):120–143, August 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302158.

Serkh:2016:SEP

[SR16] Kirill Serkh and Vladimir Rokhlin. On the solution of elliptic partial differential equations on regions with corners. Journal of Computational Physics, 305(??):150–171, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006889.

Subramaniam:2018:PVI

[SR18] Vivek Subramaniam and Laxminarayan L. Raja. A plasmavacuum interface tracking algorithm for magnetohydrodynamic simulations of coaxial plasma accelerators. Journal of Computational Physics, 365(??):207–225, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302079.

Scheufler:2019:AES

[SR19] Henning Scheufler and Johan Roenby. Accurate and efficient surface reconstruction from volume fraction data on general meshes. Journal of Computational Physics, 383(??):1–23, April 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300269.

Shams:2018:NMT

[SRBB18] Mosayeb Shams, Ali Q. Raeini, Martin J. Blunt, and Branko Bijeljic. A numerical model of two-phase flow at the microscale using the volume-of-fluid method. Journal of Computational Physics, 357(??):159–182, March 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730921X.

Samake:2017:PIL

[SRBÓ17] Abdoulaye Samaké, Pierre Rampal, Sylvain Bouillon, and Einar Ólason. Parallel implementation of a Lagrangian-based model on an adaptive mesh in C++: Application to seaice. Journal of Computational Physics, 350(??):84–96, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306368.

Soligo:2019:CSL

[SRS19] Giovanni Soligo, Alessio Roccon, and Alfredo Soldati. Coalescence of surfactant-laden drops by Phase Field Method. Journal of Computational Physics, 376(??):1292–1311, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118306880.

Santilli:2015:SOM

[SS15a] Edward Santilli and Alberto Scotti. The Stratified Ocean Model with Adaptive Refinement (SOMAR). Journal of Computational Physics, 291(??):60–81, June 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001424.

Shao:2015:CDS

[SS15b] Sihong Shao and Jean Michel Sellier. Comparison of deterministic and stochastic methods for time-dependent Wigner simulations. Journal of Computational Physics, 300(??):167–185, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005197.

Saye:2016:MME

[SS16a] R. I. Saye and J. A. Sethian. Multiscale modelling of evolving foams. Journal of Computational Physics, 315(??):273–301, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300158.

Sengupta:2016:NAB

[SS16b] Tapan K. Sengupta and Aditi Sengupta. A new alternating bi-diagonal compact scheme for non-uniform grids. Journal of Computational Physics, 310(??):1–25, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000152.

Sousa:2016:BCD

[SS16c] E. M. Sousa and U. Shumlak. A blended continuousdiscontinuous finite element method for solving the multifluid plasma model. Journal of Computational Physics, 326 (??):56–75, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304016.

Sauter:2017:CQW

[SS17a] S. A. Sauter and M. Schanz. Convolution quadrature for the wave equation with impedance boundary conditions. Journal of Computational Physics, 334(??):442–459, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300232.

Schobi:2017:UPB

[SS17b] Roland Schöbi and Bruno Sudret. Uncertainty propagation of p-boxes using sparse polynomial chaos expansions. Journal of Computational Physics, 339(??):307–327, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302164.

Strumik:2017:MHM

[SS17c] Marek Strumik and Krzysztof Stasiewicz. Multidimensional Hall magnetohydrodynamics with isotropic or anisotropic thermal pressure: Numerical scheme and its validation using solitary waves. Journal of Computational Physics, 330 (??):846–862, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116305666.

Schlachter:2018:HPS

[SS18a] Louisa Schlachter and Florian Schneider. A hyperbolicitypreserving stochastic Galerkin approximation for uncertain hyperbolic systems of equations. Journal of Computational Physics, 375(??):80–98, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830487X.

Sirignano:2018:DDL

[SS18b] Justin Sirignano and Konstantinos Spiliopoulos. DGM: a deep learning algorithm for solving partial differential equations. Journal of Computational Physics, 375(??):1339–1364, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305527.

Subber:2018:PTI

[SS18c] Waad Subber and Abhijit Sarkar. A parallel time integrator for noisy nonlinear oscillatory systems. Journal of Computational Physics, 362(??):190–207, June 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300299.

Solomenko:2017:LSM

[SSA17] Zlatko Solomenko, Peter D. M. Spelt, and Pascal Alix. A levelset method for large-scale simulations of three-dimensional flows with moving contact lines. Journal of Computational Physics, 348(??):151–170, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305132.

Shadid:2016:SFS

[SSC^{+}16] J. N. Shadid, T. M. Smith, E. C. Cyr, T. M. Wildey, and R. P. Pawlowski. Stabilized FE simulation of prototype thermalhydraulics problems with integrated adjoint-based capabilities. Journal of Computational Physics, 321(??):321–341, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301280.

Safta:2015:HDC

[SSDN15] Cosmin Safta, Khachik Sargsyan, Bert Debusschere, and Habib N. Najm. Hybrid discrete/continuum algorithms for stochastic reaction networks. Journal of Computational Physics, 281(??):177–198, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007086.

Schmitt:2016:DMN

[SSL^{+}16a] Nikolai Schmitt, Claire Scheid, Stéphane Lanteri, Antoine Moreau, and Jonathan Viquerat. A DGTD method for the numerical modeling of the interaction of light with nanometer scale metallic structures taking into account non-local dispersion effects. Journal of Computational Physics, 316(??):396– 415, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300614.

Stosic:2016:GAE

[SSL^{+}16b] Darko Stosić, Dusan Stosić, Teresa Ludermir, Borko Stosić, and Milorad V. Milosević. GPU-advanced 3D electromagnetic simulations of superconductors in the Ginzburg–Landau formalism. Journal of Computational Physics, 322(??):183–198, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116302650.

Sadovskii:2017:MWP

[SSL17] Vladimir M. Sadovskii, Oxana V. Sadovskaya, and Alexander A. Lukyanov. Modeling of wave processes in blocky media with porous and fluid-saturated interlayers. Journal of Computational Physics, 345(??):834–855, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304461.

Smedley-Stevenson:2015:ADL

[SSM15] Richard P. Smedley-Stevenson and Ryan G. McClarren. Asymptotic diffusion limit of cell temperature discretisation schemes for thermal radiation transport. Journal of Computational Physics, 286(??):214–235, April 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007146.

Shi:2017:NBF

[SSM^{+}17] Sheng-Bing Shi, Wei Shao, Jing Ma, Congjun Jin, and Xiao-Hua Wang. Newmark–Beta–FDTD method for superresolution analysis of time reversal waves. Journal of Computational Physics, 345(??):475–483, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304114.

Stefanescu:2015:PDR

[SSN15] R. Stefanescu, A. Sandu, and I. M. Navon. POD/ DEIM reduced-order strategies for efficient four dimensional variational data assimilation. Journal of Computational Physics, 295(??):569–595, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002867.

Sondak:2015:NCF

[SSO^{+}15] D. Sondak, J. N. Shadid, A. A. Oberai, R. P. Pawlowski, E. C. Cyr, and T. M. Smith. A new class of finite element variational multiscale turbulence models for incompressible magnetohydrodynamics. Journal of Computational Physics, 295(??):596–616, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002922.

Sun:2015:EFG

[SST^{+}15] Y. Sun, C. Shu, C. J. Teo, Y. Wang, and L. M. Yang. Explicit formulations of gas-kinetic flux solver for simulation of incompressible and compressible viscous flows. Journal of Computational Physics, 300(??):492–519, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005136.

Schmitt:2018:STD

[SSVL18] Nikolai Schmitt, Claire Scheid, Jonathan Viquerat, and Stéphane Lanteri. Simulation of three-dimensional nanoscale light interaction with spatially dispersive metals using a high order curvilinear DGTD method. Journal of Computational Physics, 373(??):210–229, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830411X.

Smolarkiewicz:2016:SAS

[SSX16] Piotr K. Smolarkiewicz, Joanna Szmelter, and Feng Xiao. Simulation of all-scale atmospheric dynamics on unstructured meshes. Journal of Computational Physics, 322(??):267–287, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630273X.

Seibold:2019:USM

[SSZ19] Benjamin Seibold, David Shirokoff, and Dong Zhou. Unconditional stability for multistep ImEx schemes: Practice. Journal of Computational Physics, 376(??):295–321, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306466.

Szmelter:2019:NOF

[SSZC19] Joanna Szmelter, Piotr K. Smolarkiewicz, Zhao Zhang, and Zhixin Cao. Non-oscillatory forward-in-time integrators for viscous incompressible flows past a sphere. Journal of Computational Physics, 386(??):365–383, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301214.

Singh:2015:TPI

[ST15] Amit Singh and Ellad B. Tadmor. Thermal parameter identification for non-Fourier heat transfer from molecular dynamics. Journal of Computational Physics, 299(??):667–686, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004532.

Sainath:2016:FWA

[ST16] Kamalesh Sainath and Fernando L. Teixeira. Full-wave algorithm to model effects of bedding slopes on the response of subsurface electromagnetic geophysical sensors near unconformities. Journal of Computational Physics, 313(??):328–351, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001005.

Shamshirgar:2017:SEM

[ST17] Davoud Saffar Shamshirgar and Anna-Karin Tornberg. The Spectral Ewald method for singly periodic domains. Journal of Computational Physics, 347(??):341–366, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730503X.

Sarna:2018:ESH

[ST18a] Neeraj Sarna and Manuel Torrilhon. Entropy stable Hermite approximation of the linearised Boltzmann equation for inflow and outflow boundaries. Journal of Computational Physics, 369(??):16–44, September 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302833.

Siegel:2018:LTS

[ST18b] Michael Siegel and Anna-Karin Tornberg. A local target specific quadrature by expansion method for evaluation of layer potentials in 3D. Journal of Computational Physics, 364(??):365–392, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830158X.

Sorgentone:2018:HAB

[ST18c] Chiara Sorgentone and Anna-Karin Tornberg. A highly accurate boundary integral equation method for surfactant-laden drops in 3D. Journal of Computational Physics, 360(??):167– 191, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300433.

Schaefer:2017:SGA

[STEK17] Ido Schaefer, Hillel Tal-Ezer, and Ronnie Kosloff. Semi-global approach for propagation of the time-dependent Schrödinger equation for time-dependent and nonlinear problems. Journal of Computational Physics, 343(??):368–413, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302887. See corrigendum

[STEK22].

Schaefer:2022:CSG

[STEK22] Ido Schaefer, Hillel Tal-Ezer, and Ronnie Kosloff. Corrigendum to “Semi-global approach for propagation of the time-dependent Schrödinger equation for time-dependent and nonlinear problems” [j. comput. phys. 343 (2017) 368–413]. Journal of Computational Physics, 463(??):??, August 15, 2022. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199912200362X. See [STEK17].

Schoder:2019:RIM

[STFK19] S. Schoder, F. Toth, C. Freidhager, and M. Kaltenbacher. Revisiting infinite mapping layer for open domain problems. Journal of Computational Physics, 392(??):354–367, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303201.

Silva:2017:LHO

[STG17] Goncalo Silva, Laurent Talon, and Irina Ginzburg. Lowand high-order accurate boundary conditions: From Stokes to Darcy porous flow modeled with standard and improved Brinkman lattice Boltzmann schemes. Journal of Computational Physics, 335(??):50–83, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300335.

Sarkar:2017:EOP

[STHW17] Kanchan Sarkar, Mehmet Topsakal, N. A. W. Holzwarth, and Renata M. Wentzcovitch. Evolutionary optimization of PAW data-sets for accurate high pressure simulations. Journal of Computational Physics, 347(??):39–55, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304771.

Stiller:2016:RMH

[Sti16] Jörg Stiller. Robust multigrid for high-order discontinuous Galerkin methods: a fast Poisson solver suitable for high-aspect ratio Cartesian grids. Journal of Computational Physics, 327(??):317–336, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304624.

Sato:2016:CDD

[STK^{+}16] Norikazu Sato, Shintaro Takeuchi, Takeo Kajishima, Masahide Inagaki, and Nariaki Horinouchi. A consistent direct discretization scheme on Cartesian grids for convective and conjugate heat transfer. Journal of Computational Physics, 321(??):76–104, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116301796.

Shrestha:2015:NSM

[STKH15] Samir Shrestha, Sudarshan Tiwari, Axel Klar, and Steffen Hardt. Numerical simulation of a moving rigid body in a rarefied gas. Journal of Computational Physics, 292(??):239–252, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001771.

Seol:2019:IBM

[STKL19] Yunchang Seol, Yu-Hau Tseng, Yongsam Kim, and Ming-Chih Lai. An immersed boundary method for simulating Newtonian vesicles in viscoelastic fluid. Journal of Computational Physics, 376(??):1009–1027, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306946.

Stolk:2016:DMS

[Sto16] Christiaan C. Stolk. A dispersion minimizing scheme for the 3-D Helmholtz equation based on ray theory. Journal of Computational Physics, 314(??):618–646, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001753.

Stoltz:2017:SSD

[Sto17] Gabriel Stoltz. Stable schemes for dissipative particle dynamics with conserved energy. Journal of Computational Physics, 340(??):451–469, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302607.

Salimi:2015:NLR

[STR15] M. R. Salimi and M. Taeibi-Rahni. New lifting relations for estimating LBM distribution functions from corresponding macroscopic quantities, based on equilibrium and non-equilibrium moments. Journal of Computational Physics, 302(??):155–175, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005732.

Stricker:2017:NSA

[Str17] L. Stricker. Numerical simulation of artificial microswimmers driven by Marangoni flow. Journal of Computational Physics, 347(??):467–489, October 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305090.

Strain:2018:FFT

[Str18] John Strain. Fast Fourier Transforms of piecewise polynomials. Journal of Computational Physics, 373(??):346–369, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304583.

Stuck:2015:AVF

[Stü15] Arthur Stück. An adjoint view on flux consistency and strong wall boundary conditions to the Navier–Stokes equations. Journal of Computational Physics, 301(??):247–264, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005495.

Stuck:2017:DCS

[Stü17] Arthur Stück. Dual-consistency study for Green–Gauss gradient schemes in an unstructured Navier–Stokes method. Journal of Computational Physics, 350(??):530–549, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306241.

Shervani-Tabar:2018:SCL

[STV18] Navid Shervani-Tabar and Oleg V. Vasilyev. Stabilized conservative level set method. Journal of Computational Physics, 375(??):1033–1044, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830620X.

Sorgentone:2019:BIM

[STV19] Chiara Sorgentone, Anna-Karin Tornberg, and Petia M. Vlahovska. A 3D boundary integral method for the electrohydrodynamics of surfactant-covered drops. Journal of Computational Physics, 389(??):111–127, July 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302244.

Shi:2016:STC

[STW16] Y. Shi, G. H. Tang, and Y. Wang. Simulation of threecomponent fluid flows using the multiphase lattice Boltzmann flux solver. Journal of Computational Physics, 314(??):228– 243, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001637.

Sibatov:2015:DTC

[SU15] R. T. Sibatov and V. V. Uchaikin. Dispersive transport of charge carriers in disordered nanostructured materials. Journal of Computational Physics, 293(??):409–426, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000261.

Subich:2015:RMM

[Sub15] Christopher J. Subich. A robust moving mesh method for spectral collocation solutions of time-dependent partial differential equations. Journal of Computational Physics, 294(??):297–311, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002296.

Subich:2018:HOF

[Sub18] Christopher J. Subich. Higher-order finite volume differential operators with selective upwinding on the icosahedral spherical grid. Journal of Computational Physics, 368 (??):21–46, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118302869.

Sun:2019:DSP

[Sun19a] Shuyu Sun. Darcy-scale phase equilibrium modeling with gravity and capillarity. Journal of Computational Physics, 399(??):Article 108908, December 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306138.

Sun:2019:MSQ

[Sun19b] Zhengjie Sun. Multi-symplectic quasi-interpolation method for Hamiltonian partial differential equations. Journal of Computational Physics, 395(??):125–143, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304346.

Sekulic:2018:VAS

[SUR18] Ivan Sekulic, Eduard Ubeda, and Juan M. Rius. Versatile and accurate schemes of discretization for the electromagnetic scattering analysis of arbitrarily shaped piecewise homogeneous objects. Journal of Computational Physics, 374(??):478–494, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304959.

Suzuki:2018:BFE

[Suz18] Yukihito Suzuki. Bracket formulations and energy- and helicity-preserving numerical methods for incompressible twophase flows. Journal of Computational Physics, 356(??):64–97, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308720.

Subramaniam:2017:TPI

[SV17] Gnana M. Subramaniam and Prakash Vedula. A transformed path integral approach for solution of the Fokker–Planck equation. Journal of Computational Physics, 346(??):49–70, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304473.

Sanderse:2019:CCR

[SV19] B. Sanderse and A. E. P. Veldman. Constraint-consistent Runge–Kutta methods for one-dimensional incompressible multiphase flow. Journal of Computational Physics, 384(??): 170–199, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119300683.

Svard:2015:WSC

[Svä15] Magnus Svärd. Weak solutions and convergent numerical schemes of modified compressible Navier–Stokes equations. Journal of Computational Physics, 288(??):19–51, May 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000753.

Sekaran:2018:ANC

[SVG18] Aarthi Sekaran, Philip Varghese, and David Goldstein. An analysis of numerical convergence in discrete velocity gas dynamics for internal flows. Journal of Computational Physics, 365(??):226–242, July 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301839.

Shi:2015:ABE

[SW15] Lei Shi and Z. J. Wang. Adjoint-based error estimation and mesh adaptation for the correction procedure via reconstruction method. Journal of Computational Physics, 295(??):261–284, August 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115002570.

Shen:2016:CST

[SW16] Hua Shen and Chih-Yung Wen. A characteristic space–time conservation element and solution element method for conservation laws II. Multidimensional extension. Journal of Computational Physics, 305(??):775–792, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007573.

Schmidtmann:2017:HES

[SW17a] Birte Schmidtmann and Andrew R. Winters. Hybrid entropy stable HLL-type Riemann solvers for hyperbolic conservation laws. Journal of Computational Physics, 330(??):566–570, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305320.

Song:2017:MGN

[SW17b] Shufang Song and Lu Wang. Modified GMDH-NN algorithm and its application for global sensitivity analysis. Journal of Computational Physics, 348(??):534–548, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305351.

Shankar:2018:MFS

[SW18a] Varun Shankar and Grady B. Wright. Mesh-free semi-Lagrangian methods for transport on a sphere using radial basis functions. Journal of Computational Physics, 365 (??):170–190, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118302213.

Singh:2018:STD

[SW18b] Gurpreet Singh and Mary F. Wheeler. A space-time domain decomposition approach using enhanced velocity mixed finite element method. Journal of Computational Physics, 374(??):893–911, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305369.

Samulyak:2018:LPM

[SWC18] Roman Samulyak, Xingyu Wang, and Hsin-Chiang Chen. Lagrangian particle method for compressible fluid dynamics. Journal of Computational Physics, 362(??):1–19, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300809.

Sweezy:2018:MCV

[Swe18] Jeremy E. Sweezy. A Monte Carlo volumetric-ray-casting estimator for global fluence tallies on GPUs. Journal of Computational Physics, 372(??):426–445, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118304108.

Schunert:2017:FND

[SWG^{+}17] Sebastian Schunert, Yaqi Wang, Frederick Gleicher, Javier Ortensi, Benjamin Baker, Vincent Laboure, Congjian Wang, Mark DeHart, and Richard Martineau. A flexible nonlinear diffusion acceleration method for the S_{N} transport equations discretized with discontinuous finite elements. Journal of Computational Physics, 338(??):107–136, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301286.

Sonnendrucker:2015:SCV

[SWHK15] Eric Sonnendrücker, Abigail Wacher, Roman Hatzky, and Ralf Kleiber. A split control variate scheme for PIC simulations with collisions. Journal of Computational Physics, 295(??):402–419, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002442.

Souopgui:2016:STA

[SWHV16] Innocent Souopgui, Scott A. Wieland, M. Yousuff Hussaini, and Oleg V. Vasilyev. Space–time adaptive approach to variational data assimilation using wavelets. Journal of Computational Physics, 306(??):253–268, February 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007706.

Shuang:2019:PUG

[SWJG19] Tan Shuang, Sun Wenjun, Wei Junxia, and Ni Guoxi. A parallel unified gas kinetic scheme for three-dimensional multi-group neutron transport. Journal of Computational Physics, 391(??):37–58, August 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302876.

Smith:2018:NIS

[SWK18] Alastair J. Smith, Clive G. Wells, and Markus Kraft. A new iterative scheme for solving the discrete Smoluchowski equation. Journal of Computational Physics, 352(??):373–387, January 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117307052.

Subramaniam:2019:HOW

[SWL19] Akshay Subramaniam, Man Long Wong, and Sanjiva K. Lele. A high-order weighted compact high resolution scheme with boundary closures for compressible turbulent flows with shocks. Journal of Computational Physics, 397(??):Article 108822, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305066.

Su:2019:AEC

[SWLW19] Wei Su, Peng Wang, Haihu Liu, and Lei Wu. Accurate and efficient computation of the Boltzmann equation for Couette flow: Influence of intermolecular potentials on Knudsen layer function and viscous slip coefficient. Journal of Computational Physics, 378(??):573–590, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307332.

Shen:2015:RHO

[SWLZ15] Hua Shen, Chih-Yung Wen, Kaixin Liu, and Deliang Zhang. Robust high-order space–time conservative schemes for solving conservation laws on hybrid meshes. Journal of Computational Physics, 281(??):375–402, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007050.

Shaw:2017:CMM

[SWMD17a] James Shaw, Hilary Weller, John Methven, and Terry Davies. Corrigendum to “Multidimensional method-of-lines transport for atmospheric flows over steep terrain using arbitrary meshes” [j. comput. phys. 344 (2017) 86–107]. Journal of Computational Physics, 349(??):636, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305958. See [SWMD17b].

Shaw:2017:MML

[SWMD17b] James Shaw, Hilary Weller, John Methven, and Terry Davies. Multidimensional method-of-lines transport for atmospheric flows over steep terrain using arbitrary meshes. Journal of Computational Physics, 344(??):86–107, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303406. See corrigendum

[SWMD17a].

Scovazzi:2017:AVN

[SWML17] Guglielmo Scovazzi, Mary F. Wheeler, Andro Mikelić, and Sanghyun Lee. Analytical and variational numerical methods for unstable miscible displacement flows in porous media. Journal of Computational Physics, 335(??):444–496, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300311.

Shen:2017:MPS

[SWPS17] Hua Shen, Chih-Yung Wen, Matteo Parsani, and Chi-Wang Shu. Maximum-principle-satisfying space-time conservation element and solution element scheme applied to compressible multifluids. Journal of Computational Physics, 330(??):668–692, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305344.

Sheng:2016:SMP

[SwS16] Qin Sheng and Hai wei Sun. Stability of a modified Peaceman–Rachford method for the paraxial Helmholtz equation on adaptive grids. Journal of Computational Physics, 325(??):259–271, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303977.

Schlanderer:2017:BDI

[SWS17] Stefan C. Schlanderer, Gabriel D. Weymouth, and Richard D. Sandberg. The boundary data immersion method for compressible flows with application to aeroacoustics. Journal of Computational Physics, 333(??):440–461, March 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116307148.

Schoepplein:2018:AEA

[SWS^{+}18] Matthias Schoepplein, Jack Weatheritt, Richard Sandberg, Mohsen Talei, and Markus Klein. Application of an evolutionary algorithm to LES modelling of turbulent transport in premixed flames. Journal of Computational Physics, 374(??):1166–1179, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305394.

Shin:2018:SFA

[SWX18] Yeonjong Shin, Kailiang Wu, and Dongbin Xiu. Sequential function approximation with noisy data. Journal of Computational Physics, 371(??):363–381, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303516.

Shen:2015:CST

[SWZ15] Hua Shen, Chih-Yung Wen, and De-Liang Zhang. A characteristic space–time conservation element and solution element method for conservation laws. Journal of Computational Physics, 288(??):101–118, May 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000881.

Su:2017:MED

[SWZ17] Xiao-Xing Su, Yue-Sheng Wang, and Chuanzeng Zhang. A matrix-exponential decomposition based time-domain method for calculating the defect states of scalar waves in twodimensional periodic structures. Journal of Computational Physics, 337(??):403–420, May 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301456.

Su:2019:HOH

[SWZW19] Wei Su, Peng Wang, Yonghao Zhang, and Lei Wu. A highorder hybridizable discontinuous Galerkin method with fast convergence to steady-state solutions of the gas kinetic equation. Journal of Computational Physics, 376(??):973–991, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305825.

Song:2015:SDS

[SX15] Fangying Song and Chuanju Xu. Spectral direction splitting methods for two-dimensional space fractional diffusion equations. Journal of Computational Physics, 299(??):196–214, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004568.

Shin:2016:NOS

[SX16] Yeonjong Shin and Dongbin Xiu. On a near optimal sampling strategy for least squares polynomial regression. Journal of Computational Physics, 326(??):931–946, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304533.

Savard:2015:CES

[SXBB15] B. Savard, Y. Xuan, B. Bobbitt, and G. Blanquart. A computationally-efficient, semi-implicit, iterative method for the time-integration of reacting flows with stiff chemistry. Journal of Computational Physics, 295(??):740–769, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002648.

Shen:2018:SAV

[SXY18] Jie Shen, Jie Xu, and Jiang Yang. The scalar auxiliary variable (SAV) approach for gradient flows. Journal of Computational Physics, 353(??):407–416, January 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730774X.

Sheng:2016:NNF

[SY16] Zhiqiang Sheng and Guangwei Yuan. A new nonlinear finite volume scheme preserving positivity for diffusion equations. Journal of Computational Physics, 315(??):182–193, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300134.

Sakamoto:2017:IPE

[SY17] Hiroki Sakamoto and Toshihiro Yamamoto. Improvement and performance evaluation of the perturbation source method for an exact Monte Carlo perturbation calculation in fixed source problems. Journal of Computational Physics, 345(??):245–259, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911730373X.

Shojaei:2018:CSM

[SY18a] Mostafa Faghih Shojaei and Arash Yavari. Compatible-strain mixed finite element methods for incompressible nonlinear elasticity. Journal of Computational Physics, 361(??):247–279, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300755.

Sjogreen:2018:HOE

[SY18b] Björn Sjögreen and H. C. Yee. High order entropy conservative central schemes for wide ranges of compressible gas dynamics and MHD flows. Journal of Computational Physics, 364(??):153–185, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300792.

Sato:2019:TOM

[SYI^{+}19] A. Sato, T. Yamada, K. Izui, S. Nishiwaki, and S. Takata. A topology optimization method in rarefied gas flow problems using the Boltzmann equation. Journal of Computational Physics, 395(??):60–84, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304309.

Simmons:2015:PNS

[SYM15] Alex Simmons, Qianqian Yang, and Timothy Moroney. A preconditioned numerical solver for stiff nonlinear reaction diffusion equations with fractional Laplacians that avoids dense matrices. Journal of Computational Physics, 287 (??):254–268, 2015. CODEN JCTPAH. ISSN 0021-9991.

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000741.

Simmons:2017:FVM

[SYM17] Alex Simmons, Qianqian Yang, and Timothy Moroney. A finite volume method for two-sided fractional diffusion equations on non-uniform meshes. Journal of Computational Physics, 335(??):747–759, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300773.

Sakurai:2019:VPI

[SYOS19] Teluo Sakurai, Katsunori Yoshimatsu, Naoya Okamoto, and Kai Schneider. Volume penalization for inhomogeneous Neumann boundary conditions modeling scalar flux in complicated geometry. Journal of Computational Physics, 390(??):452–469, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302414. See critique

[TNB21] and corrigendum [SYOS21].

Sakurai:2021:CVP

[SYOS21] Teluo Sakurai, Katsunori Yoshimatsu, Naoya Okamoto, and Kai Schneider. Corrigendum to “Volume penalization for inhomogeneous Neumann boundary conditions modeling scalar flux in complicated geometry” [J. Comput. Phys. 390 (2019) 452–469]. Journal of Computational Physics, 443(??):Article 110497, October 15, 2021. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999121003922.

Sjogreen:2014:HOF

[SYV14] Björn Sjögreen, H. C. Yee, and Marcel Vinokur. On high order finite-difference metric discretizations satisfying GCL on moving and deforming grids. Journal of Computational Physics, 265(??):211–220, May 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000904. See corrigendum [SYV17].

Sjogreen:2017:CHO

[SYV17] Björn Sjögreen, H. C. Yee, and Marcel Vinokur. Corrigendum to “On High Order Finite-Difference Metric Discretizations Satisfying GCL on Moving and Deforming Grids” [J. Comput. Phys. 265 (2014) 211–220]. Journal of Computational Physics, 338(??):650, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301997. See [SYV14].

Shen:2015:EES

[SYY15] Jie Shen, Xiaofeng Yang, and Haijun Yu. Efficient energy stable numerical schemes for a phase field moving contact line model. Journal of Computational Physics, 284(??):617–630, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008699.

Shen:2016:RHT

[SYY16] Zhijun Shen, Wei Yan, and Guangwei Yuan. A robust HLLCtype Riemann solver for strong shock. Journal of Computational Physics, 309(??):185–206, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000024.

Semenikhin:2015:AIA

[SZ15a] Igor Semenikhin and Mauro Zanuccoli. Application of the iterative approach to modal methods for the solution of Maxwell’s equations. Journal of Computational Physics, 300(??):438–454, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005057.

Song:2015:SPP

[SZ15b] Lunji Song and Zhimin Zhang. Superconvergence property of an over-penalized discontinuous Galerkin finite element gradient recovery method. Journal of Computational Physics, 299(??):1004–1020, October 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004817.

Song:2017:MOS

[SZ17] Wanjun Song and Hou Zhang. Memory-optimized shift operator alternating direction implicit finite difference time domain method for plasma. Journal of Computational Physics, 349(??):122–136, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305880.

Sun:2018:NMS

[SZCL18] Hui Sun, Shenggao Zhou, Li-Tien Cheng, and Bo Li. Numerical methods for solvent Stokes flow and solute-solvent interfacial dynamics of charged molecules. Journal of Computational Physics, 374(??):533–549, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305072.

Sun:2015:SGC

[SZF15] Weijia Sun, Binzhong Zhou, and Li-Yun Fu. A staggeredgrid convolutional differentiator for elastic wave modelling. Journal of Computational Physics, 301(??):59–76, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005367.

Schoberl:2017:PCG

[SZK17] Markus Schöberl, Nicholas Zabaras, and Phaedon-Stelios Koutsourelakis. Predictive coarse-graining. Journal of Computational Physics, 333(??):49–77, March 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306763.

Samiee:2019:USMa

[SZM19a] Mehdi Samiee, Mohsen Zayernouri, and Mark M. Meerschaert. A unified spectral method for FPDEs with twosided derivatives; Part I: a fast solver. Journal of Computational Physics, 385(??):225–243, May 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300901.

Samiee:2019:USMb

[SZM19b] Mehdi Samiee, Mohsen Zayernouri, and Mark M. Meerschaert. A unified spectral method for FPDEs with two-sided derivatives; part II: Stability, and error analysis. Journal of Computational Physics, 385(??):244–261, May 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305023.

S:2019:AMH

[SZN19] Vevek U S, B. Zang, and T. H. New. Adaptive mapping for high order WENO methods. Journal of Computational Physics, 381(??):162–188, March 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300154. See corrigendum [SZN20].

S:2020:CAM

[SZN20] Vevek U. S, B. Zang, and T. H. New. Corrigendum to “Adaptive mapping for high order WENO methods” [j. comput. phys. 381 (2019) 162–188]. Journal of Computational Physics, 414(??):Article 109492, August 1, 2020. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999120302667. See [SZN19].

Szmelter:2015:UMA

[SZS15] Joanna Szmelter, Zhao Zhang, and Piotr K. Smolarkiewicz. An unstructured-mesh atmospheric model for nonhydrostatic dynamics: Towards optimal mesh resolution. Journal of Computational Physics, 294(??):363–381, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002156.

Shi:2016:SEM

[SZW^{+}16] Linlin Shi, Yuanguo Zhou, Jia-Min Wang, Mingwei Zhuang, Na Liu, and Qing Huo Liu. Spectral element method for elastic and acoustic waves in frequency domain. Journal of Computational Physics, 327(??):19–38, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304570.

Santillana:2016:ENE

[SZY16] Mauricio Santillana, Lin Zhang, and Robert Yantosca. Estimating numerical errors due to operator splitting in global atmospheric chemistry models: Transport and chemistry. Journal of Computational Physics, 305(??):372–386, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007299.

Tugnoli:2017:LAA

[TABR17] Matteo Tugnoli, Antonella Abbà, Luca Bonaventura, and Marco Restelli. A locally p-adaptive approach for Large Eddy Simulation of compressible flows in a DG framework. Journal of Computational Physics, 349(??):33–58, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305788.

Tene:2016:AMM

[TAH16] Matei Tene, Mohammed Saad Al Kobaisi, and Hadi Hajibeygi. Algebraic multiscale method for flow in heterogeneous porous media with embedded discrete fractures (F-AMS). Journal of Computational Physics, 321(??):819–845, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302376.

Terrier:2017:CDM

[TAJ^{+}17] Pierre Terrier, Manuel Athènes, Thomas Jourdan, Gilles Adjanor, and Gabriel Stoltz. Cluster dynamics modelling of materials: a new hybrid deterministic/stochastic coupling approach. Journal of Computational Physics, 350(??):280–295, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305867.

Tao:2016:EHO

[Tao16] Molei Tao. Explicit high-order symplectic integrators for charged particles in general electromagnetic fields. Journal of Computational Physics, 327(??):245–251, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304685.

Trujillo:2017:DLS

[TAR17] Mario F. Trujillo, Lakshman Anumolu, and Doug Ryddner. The distortion of the level set gradient under advection. Journal of Computational Physics, 334(??):81–101, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116307045.

Tofighian:2019:PND

[TASA19] H. Tofighian, E. Amani, and M. Saffar-Avval. Parcelnumber-density control algorithms for the efficient simulation of particle-laden two-phase flows. Journal of Computational Physics, 387(??):569–588, June 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930169X.

Tavakoli:2015:CEA

[Tav15] Rouhollah Tavakoli. Computationally efficient approach for the minimization of volume constrained vector-valued Ginzburg–Landau energy functional. Journal of Computational Physics, 295(??):355–378, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002661.

Tavakoli:2016:UES

[Tav16] Rouhollah Tavakoli. Unconditionally energy stable time stepping scheme for Cahn–Morral equation: Application to multicomponent spinodal decomposition and optimal space tiling. Journal of Computational Physics, 304(??):441–464, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006828.

Tripathy:2018:DUL

[TB18] Rohit K. Tripathy and Ilias Bilionis. Deep UQ: Learning deep neural network surrogate models for high dimensional uncertainty quantification. Journal of Computational Physics, 375(??):565–588, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305655.

Tlupova:2019:RSD

[TB19] Svetlana Tlupova and J. Thomas Beale. Regularized single and double layer integrals in 3D Stokes flow. Journal of Computational Physics, 386(??):568–584, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301482.

Tsoupas:2019:CMF

[TBB^{+}19] Nicholaos Tsoupas, Joseph S. Berg, Stephen Brooks, François Méot, Vadim Ptitsyn, Dejan Trbojevic, and Shinji Machida. Computation of magnetic fields from field components on a plane grid. Journal of Computational Physics, 396(??):653– 668, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0021999119304929.

Tamain:2016:TCE

[TBC^{+}16] P. Tamain, H. Bufferand, G. Ciraolo, C. Colin, D. Galassi, Ph. Ghendrih, F. Schwander, and E. Serre. The TOKAM3X code for edge turbulence fluid simulations of tokamak plasmas in versatile magnetic geometries. Journal of Computational Physics, 321(??):606–623, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301838.

Tripathy:2016:GPB

[TBG16] Rohit Tripathy, Ilias Bilionis, and Marcial Gonzalez. Gaussian processes with built-in dimensionality reduction: Applications to high-dimensional uncertainty propagation. Journal of Computational Physics, 321(??):191–223, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630184X.

Takash:2018:NAB

[TBHG18] Ahmad Al Takash, Marianne Beringhier, Mohammad Hammoud, and Jean-Claude Grandidier. Numerical approach based on the collection of the most significant modes to solve cyclic transient thermal problems involving different time scales. Journal of Computational Physics, 375(??):950–959, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305928.

Trozzo:2015:ABE

[TBLJ15] R. Trozzo, G. Boedec, M. Leonetti, and M. Jaeger. Axisymmetric Boundary Element Method for vesicles in a capillary. Journal of Computational Physics, 289(??):62–82, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000923.

Tveit:2015:ISS

[TBLM15] Svenn Tveit, Shaaban A. Bakr, Martha Lien, and Trond Mannseth. Identification of subsurface structures using electromagnetic data and shape priors. Journal of Computational Physics, 284(??):505–527, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008584.

Tome:2016:FDT

[TBO^{+}16] M. F. Tomé, J. Bertoco, C. M. Oishi, M. S. B. Araujo, D. Cruz, F. T. Pinho, and M. Vynnycky. A finite difference technique for solving a time strain separable K-BKZ constitutive equation for two-dimensional moving free surface flows. Journal of Computational Physics, 311(??):114–141, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000401.

Taitano:2015:CECa

[TC15a] William T. Taitano and Luis Chacón. Charge-and-energy conserving moment-based accelerator for a multi-species Vlasov– Fokker–Planck–Ampère system, part i: Collisionless aspects. Journal of Computational Physics, 284(??):718–736, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008559.

Thuburn:2015:PDM

[TC15b] John Thuburn and Colin J. Cotter. A primal-dual mimetic finite element scheme for the rotating shallow water equations on polygonal spherical meshes. Journal of Computational Physics, 290(??):274–297, June 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001151.

Towne:2015:OWS

[TC15c] Aaron Towne and Tim Colonius. One-way spatial integration of hyperbolic equations. Journal of Computational Physics, 300(??):844–861, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005343.

Turkoz:2015:ASP

[TC15d] Emre Turkoz and Murat Celik. AETHER: a simulation platform for inductively coupled plasma. Journal of Computational Physics, 286(??):87–102, April 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000315.

Tang:2016:ASM

[TCA16] Kunkun Tang, Pietro M. Congedo, and Rémi Abgrall. Adaptive surrogate modeling by ANOVA and sparse polynomial dimensional decomposition for global sensitivity analysis in fluid simulation. Journal of Computational Physics, 314(??):557– 589, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001789.

Tanaka:2018:MRM

[TCB18] Masayuki Tanaka, Rui Cardoso, and Hamid Bahai. Multiresolution MPS method. Journal of Computational Physics, 359(??):106–136, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118300019.

Takahashi:2017:AIF

[TCD17] Toru Takahashi, Pieter Coulier, and Eric Darve. Application of the inverse fast multipole method as a preconditioner in a 3D Helmholtz boundary element method. Journal of Computational Physics, 341(??):406–428, July 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302875.

Toro:2015:NNF

[TCL15] Eleuterio F. Toro, Cristóbal E. Castro, and Bok Jik Lee. A novel numerical flux for the 3D Euler equations with general equation of state. Journal of Computational Physics, 303(??):80–94, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006300.

Taitano:2016:ACM

[TCS16a] W. T. Taitano, L. Chacón, and A. N. Simakov. An adaptive, conservative 0D-2V multispecies Rosenbluth–Fokker–Planck solver for arbitrarily disparate mass and temperature regimes. Journal of Computational Physics, 318(??):391–420, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300729.

Turner:2016:VER

[TCS^{+}16b] John A. Turner, Kevin Clarno, Matt Sieger, Roscoe Bartlett, Benjamin Collins, Roger Pawlowski, Rodney Schmidt, and Randall Summers. The Virtual Environment for Reactor Applications (VERA): Design and architecture. Journal of Computational Physics, 326(??):544–568, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304156.

Taitano:2017:EPD

[TCS17] W. T. Taitano, L. Chacón, and A. N. Simakov. An equilibriumpreserving discretization for the nonlinear Rosenbluth–Fokker– Planck operator in arbitrary multi-dimensional geometry. Journal of Computational Physics, 339(??):453–460, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302279.

Taitano:2015:MME

[TCSM15] W. T. Taitano, L. Chacón, A. N. Simakov, and K. Molvig. A mass, momentum, and energy conserving, fully implicit, scalable algorithm for the multi-dimensional, multi-species Rosenbluth–Fokker–Planck equation. Journal of Computational Physics, 297(??):357–380, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003599.

Tavelli:2016:SST

[TD16a]  Maurizio Tavelli and Michael Dumbser.  A staggered

space–time discontinuous Galerkin method for the threedimensional incompressible Navier–Stokes equations on unstructured tetrahedral meshes. Journal of Computational Physics, 319(??):294–323, August 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911630136X.

Trehan:2016:TPQ

[TD16b] Sumeet Trehan and Louis J. Durlofsky. Trajectory piecewise quadratic reduced-order model for subsurface flow, with application to PDE-constrained optimization. Journal of Computational Physics, 326(??):446–473, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303898.

Tavelli:2017:PBS

[TD17] Maurizio Tavelli and Michael Dumbser. A pressure-based semiimplicit space-time discontinuous Galerkin method on staggered unstructured meshes for the solution of the compressible Navier–Stokes equations at all Mach numbers. Journal of Computational Physics, 341(??):341–376, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302255.

Tavelli:2018:AHO

[TD18] Maurizio Tavelli and Michael Dumbser. Arbitrary high order accurate space-time discontinuous Galerkin finite element schemes on staggered unstructured meshes for linear elasticity. Journal of Computational Physics, 366(??):386–414, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302043.

Tavelli:2019:SDI

[TDC^{+}19] Maurizio Tavelli, Michael Dumbser, Dominic Etienne Charrier, Leonhard Rannabauer, Tobias Weinzierl, and Michael Bader. A simple diffuse interface approach on adaptive Cartesian grids for the linear elastic wave equations with complex topography. Journal of Computational Physics, 386(??):158– 189, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300786.

Tarakanov:2019:RBS

[TE19] Alexander Tarakanov and Ahmed H. Elsheikh. Regressionbased sparse polynomial chaos for uncertainty quantification of subsurface flow models. Journal of Computational Physics, 399(??):Article 108909, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S002199911930614X.

tenEikelder:2017:ACS

[tEDKT17] M. F. P. ten Eikelder, F. Daude, B. Koren, and A. S. Tijsseling. An acoustic-convective splitting-based approach for the Kapila two-phase flow model. Journal of Computational Physics, 331(??):188–208, February 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306234.

Tripathi:2019:MST

[TEP19] Bharat B. Tripathi, David Espíndola, and Gianmarco F. Pinton. Modeling and simulations of two dimensional propagation of shear shock waves in relaxing soft solids. Journal of Computational Physics, 395(??):205–222, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304218.

Terekhov:2018:SHO

[Ter18] Andrew V. Terekhov. The stabilization of high-order multistep schemes for the Laguerre one-way wave equation solver. Journal of Computational Physics, 368(??):115–130, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302924.

Teukolsky:2015:SNM

[Teu15] Saul A. Teukolsky. Short note on the mass matrix for Gauss–Lobatto grid points. Journal of Computational Physics, 283(??):408–413, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114008201.

Teukolsky:2016:FDG

[Teu16] Saul A. Teukolsky. Formulation of discontinuous Galerkin methods for relativistic astrophysics. Journal of Computational Physics, 312(??):333–356, May 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000954.

Takeuchi:2018:IPB

[TFGK18] Shintaro Takeuchi, Hiroki Fukuoka, Jingchen Gu, and Takeo Kajishima. Interaction problem between fluid and membrane by a consistent direct discretisation approach. Journal of Computational Physics, 371(??):1018–1042, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303413.

Tsilifis:2017:RWC

[TG17] Panagiotis Tsilifis and Roger G. Ghanem. Reduced Wiener Chaos representation of random fields via basis adaptation and projection. Journal of Computational Physics, 341(??):102– 120, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302802.

Theillard:2019:SNS

[TGS19] Maxime Theillard, Frédéric Gibou, and David Saintillan. Sharp numerical simulation of incompressible two-phase flows. Journal of Computational Physics, 391(??):91–118, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302578.

Thornber:2018:FEM

[TGY18] Ben Thornber, Michael Groom, and David Youngs. A five-equation model for the simulation of miscible and viscous compressible fluids. Journal of Computational Physics, 372(??):256–280, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304066.

Treister:2016:FMA

[TH16] Eran Treister and Eldad Haber. A fast marching algorithm for the factored eikonal equation. Journal of Computational Physics, 324(??):210–225, November 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303552.

Taneja:2018:FCD

[TH18] Ankur Taneja and Jonathan Higdon. A fully-coupled discontinuous Galerkin spectral element method for twophase flow in petroleum reservoirs. Journal of Computational Physics, 352(??):341–372, January 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307325.

Tsilifis:2019:CSA

[THS^{+}19] Panagiotis Tsilifis, Xun Huan, Cosmin Safta, Khachik Sargsyan, Guilhem Lacaze, Joseph C. Oefelein, Habib N. Najm, and Roger G. Ghanem. Compressive sensing adaptation for polynomial chaos expansions. Journal of Computational Physics, 380(??):29–47, March 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118308064.

Tierens:2016:HOH

[Tie16] W. Tierens. Higher-order hybrid implicit/explicit FDTD time-stepping. Journal of Computational Physics, 327(??): 643–652, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304922.

Tierens:2018:EPS

[Tie18] W. Tierens. Explicit and provably stable spatiotemporal FDTD refinement. Journal of Computational Physics, 375(??):901–917, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306090.

Terashima:2012:ASG

[TK12] Hiroshi Terashima and Mitsuo Koshi. Approach for simulating gas-liquid-like flows under supercritical pressures using a high-order central differencing scheme. Journal of Computational Physics, 231(20):6907–6923, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003361. See corrigendum [TK15b].

Tauriello:2015:CSP

[TK15a] Gerardo Tauriello and Petros Koumoutsakos. A comparative study of penalization and phase field methods for the solution of the diffusion equation in complex geometries. Journal of Computational Physics, 283(??):388–407, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007955.

Terashima:2015:CAS

[TK15b] Hiroshi Terashima and Mitsuo Koshi. Corrigendum to “Approach for simulating gas-liquid-like flows under supercritical pressures using a high-order central differencing scheme” [J. Comput. Phys. 231 (20) (2012) 6907–6923]. Journal of Computational Physics, 283(??):609–610, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007657. See [TK12].

Turinsky:2016:MSC

[TK16] Paul J. Turinsky and Douglas B. Kothe. Modeling and simulation challenges pursued by the Consortium for Advanced Simulation of Light Water Reactors (CASL). Journal of Computational Physics, 313(??):367–376, May 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001133.

Tang:2015:MUI

[TKB^{+}15] Yu-Hang Tang, Shuhei Kudo, Xin Bian, Zhen Li, and George Em Karniadakis. Multiscale Universal Interface: a concurrent framework for coupling heterogeneous solvers. Journal of Computational Physics, 297(??):13–31, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003228.

Taitano:2015:CECb

[TKC15] William T. Taitano, Dana A. Knoll, and Luis Chacón. Chargeand-energy conserving moment-based accelerator for a multispecies Vlasov–Fokker–Planck–Ampère system, part II: Collisional aspects. Journal of Computational Physics, 284(??): 737–757, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006305.

Tschisgale:2017:NII

[TKF17] Silvio Tschisgale, Tobias Kempe, and Jochen Fröhlich. A non-iterative immersed boundary method for spherical particles of arbitrary density ratio. Journal of Computational Physics, 339(??):432–452, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302218.

Trisjono:2016:CHO

[TKP16] Philipp Trisjono, Seongwon Kang, and Heinz Pitsch. On a consistent high-order finite difference scheme with kinetic energy conservation for simulating turbulent reacting flows. Journal of Computational Physics, 327(??):612–628, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304739.

Tomin:2015:LGS

[TL15] Pavel Tomin and Ivan Lunati. Local-global splitting for spatiotemporal-adaptive multiscale methods. Journal of Computational Physics, 280(??):214–231, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006585.

Tu:2017:ULP

[TL17] Qingsong Tu and Shaofan Li. An updated Lagrangian particle hydrodynamics (ULPH) for Newtonian fluids. Journal of Computational Physics, 348(??):493–513, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305399.

Tripathi:2018:ECS

[TLB^{+}18] Bharat B. Tripathi, Adrian Luca, Sambandam Baskar, François Coulouvrat, and Régis Marchiano. Element centered smooth artificial viscosity in discontinuous Galerkin method for propagation of acoustic shock waves on unstructured meshes. Journal of Computational Physics, 365(??):298–319, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302249.

Tavakoli:2015:HOP

[TLH15] Ehsan Tavakoli, Bamdad Lessani, and Reza Hosseini. Highorder pole-treatment in cylindrical coordinates for incompressible flow simulations with finite-difference collocated schemes. Journal of Computational Physics, 296(??):1–24, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002995.

Tillenius:2015:SRF

[TLLF15] Martin Tillenius, Elisabeth Larsson, Erik Lehto, and Natasha Flyer. A scalable RBF–FD method for atmospheric flow. Journal of Computational Physics, 298(??):406–422, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003824.

Tao:2015:HOC

[TLQ15] Zhanjing Tao, Fengyan Li, and Jianxian Qiu. Highorder central Hermite WENO schemes on staggered meshes for hyperbolic conservation laws. Journal of Computational Physics, 281(??):148–176, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007098.

Tao:2016:HOC

[TLQ16] Zhanjing Tao, Fengyan Li, and Jianxian Qiu. High-order central Hermite WENO schemes: Dimension-by-dimension moment-based reconstructions. Journal of Computational Physics, 318(??):222–251, August 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116301322.

Toulorge:2016:OGA

[TLR16] Thomas Toulorge, Jonathan Lambrechts, and Jean-François Remacle. Optimizing the geometrical accuracy of curvilinear meshes. Journal of Computational Physics, 310(??):361–380, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000310.

Thompson:2015:DWC

[TM15a] Richard J. Thompson and Trevor Moeller. A discontinuous wave-in-cell numerical scheme for hyperbolic conservation laws. Journal of Computational Physics, 299(??):404–428, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004465.

Toro:2015:ISA

[TM15b] Eleuterio F. Toro and Gino I. Montecinos. Implicit, semianalytical solution of the generalized Riemann problem for stiff hyperbolic balance laws. Journal of Computational Physics, 303(??):146–172, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006324.

Turinsky:2017:SIC

[TM17] Paul J. Turinsky and William R. Martin. Special issue on the “Consortium for Advanced Simulation of Light Water Reactors Research and Development Progress”. Journal of Computational Physics, 334(??):687–688, April 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730044X.

Teodoro:2019:RDF

[TMdO19] G. Sales Teodoro, J. A. Tenreiro Machado, and E. Capelas de Oliveira. A review of definitions of fractional derivatives and other operators. Journal of Computational Physics, 388(??):195–208, July 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119301913.

Torre:2019:DDP

[TMES19] Emiliano Torre, Stefano Marelli, Paul Embrechts, and Bruno Sudret. Data-driven polynomial chaos expansion for machine learning regression. Journal of Computational Physics, 388(??):601–623, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302220.

Trask:2016:CML

[TMH16] Nathaniel Trask, Martin Maxey, and Xiaozhe Hu. Compact moving least squares: an optimization framework for generating high-order compact meshless discretizations. Journal of Computational Physics, 326(??):596–611, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116304028.

Trask:2018:CHO

[TMH18] Nathaniel Trask, Martin Maxey, and Xiaozhe Hu. A compatible high-order meshless method for the Stokes equations with applications to suspension flows. Journal of Computational Physics, 355(??):310–326, February 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308069.

Timokhin:2019:NMS

[TMS^{+}19] I. V. Timokhin, S. A. Matveev, N. Siddharth, E. E. Tyrtyshnikov, A. P. Smirnov, and N. V. Brilliantov. Newton method for stationary and quasi-stationary problems for Smoluchowski-type equations. Journal of Computational Physics, 382(??):124–137, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300361.

Terekhov:2017:CCN

[TMT17] Kirill M. Terekhov, Bradley T. Mallison, and Hamdi A. Tchelepi. Cell-centered nonlinear finite-volume methods for the heterogeneous anisotropic diffusion problem. Journal of Computational Physics, 330(??):245–267, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305964.

Tang:2018:ALS

[TMWF18] Kunkun Tang, Luca Massa, Jonathan Wang, and Jonathan B. Freund. An adaptive least-squares global sensitivity method and application to a plasma-coupled combustion prediction with parametric correlation. Journal of Computational Physics, 361(??):167–198, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300524.

Thirumalaisamy:2021:CVP

[TNB21] Ramakrishnan Thirumalaisamy, Nishant Nangia, and Amneet Pal Singh Bhalla. Critique on “Volume penalization for inhomogeneous Neumann boundary conditions modeling scalar flux in complicated geometry”. Journal of Computational Physics, 433(??):Article 110163, May 15, 2021. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999121000553. See [SYOS19].

Thirumalaisamy:2018:TIC

[TND18] Ramakrishnan Thirumalaisamy, Ganesh Natarajan, and Amaresh Dalal. Towards an improved conservative approach for simulating electrohydrodynamic two-phase flows using volume-of-fluid. Journal of Computational Physics, 367 (??):391–398, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118302407.

Townsend:2015:ASP

[TO15] Alex Townsend and Sheehan Olver. The automatic solution of partial differential equations using a global spectral method. Journal of Computational Physics, 299(??):106–123, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004295.

Turnquist:2019:MIU

[TO19] Brian Turnquist and Mark Owkes. multiUQ: an intrusive uncertainty quantification tool for gas-liquid multiphase flows. Journal of Computational Physics, 399(??):Article 108951, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119306564.

Tofighi:2015:ISP

[TOR^{+}15] N. Tofighi, M. Ozbulut, A. Rahmat, J. J. Feng, and M. Yildiz. An incompressible smoothed particle hydrodynamics method for the motion of rigid bodies in fluids. Journal of Computational Physics, 297(??):207–220, September 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003435.

Toutant:2018:NSU

[Tou18] Adrien Toutant. Numerical simulations of unsteady viscous incompressible flows using general pressure equation. Journal of Computational Physics, 374(??):822–842, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305230.

Towers:2018:STM

[Tow18] John D. Towers. A source term method for Poisson problems on irregular domains. Journal of Computational Physics, 361(??):424–441, May 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300482.

Tartakovsky:2016:PFS

[TP16a] Alexandre M. Tartakovsky and Alexander Panchenko. Pairwise Force Smoothed Particle Hydrodynamics model for multiphase flow: Surface tension and contact line dynamics. Journal of Computational Physics, 305(??):1119–1146, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005641.

Topic:2016:SDB

[TP16b] Nikola Topic and Thorsten Pöschel. Steepest descent ballistic deposition of complex shaped particles. Journal of Computational Physics, 308(??):421–437, March 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008712.

Thirard:2017:WSM

[TP17] Christophe Thirard and Jean-Marc Parot. On a way to save memory when solving time domain boundary integral equations for acoustic and vibroacoustic applications. Journal of Computational Physics, 348(??):744–753, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730548X.

Thaler:2019:SIT

[TPA19] Stephan Thaler, Ludger Paehler, and Nikolaus A. Adams. Sparse identification of truncation errors. Journal of Computational Physics, 397(??):Article 108851, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305352.

Tricco:2016:CHD

[TPB16] Terrence S. Tricco, Daniel J. Price, and Matthew R. Bate. Constrained hyperbolic divergence cleaning in smoothed particle magnetohydrodynamics with variable cleaning speeds. Journal of Computational Physics, 322(??):326–344, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302789.

Taverniers:2016:CTC

[TPT16] Søren Taverniers, Alexander Y. Pigarov, and Daniel M. Tartakovsky. Conservative tightly-coupled simulations of stochastic multiscale systems. Journal of Computational Physics, 313(??):400–414, May 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116001182.

Tranchida:2018:MPS

[TPTT18] J. Tranchida, S. J. Plimpton, P. Thibaudeau, and A. P. Thompson. Massively parallel symplectic algorithm for coupled magnetic spin dynamics and molecular dynamics. Journal of Computational Physics, 372(??):406–425, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304200.

Trstanova:2017:ESA

[TR17] Zofia Trstanova and Stephane Redon. Estimating the speedup of adaptively restrained Langevin dynamics. Journal of Computational Physics, 336(??):412–428, May 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117301031.

Tyson:2019:HOE

[TR19] William C. Tyson and Christopher J. Roy. A higher-order error estimation framework for finite-volume CFD. Journal of Computational Physics, 394(??):632–657, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119304243.

Treyssede:2016:SEC

[Tre16] F. Treyssède. Spectral element computation of high-frequency leaky modes in three-dimensional solid waveguides. Journal of Computational Physics, 314(??):341–354, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001819.

Tobon:2015:NED

[TRL15] Luis E. Tobón, Qiang Ren, and Qing Huo Liu. A new efficient 3D Discontinuous Galerkin Time Domain (DGTD) method for large and multiscale electromagnetic simulations. Journal of Computational Physics, 283(??):374–387, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400816X.

Trenchant:2018:FDS

[TRLK18] Vincent Trenchant, Hector Ramirez, Yann Le Gorrec, and Paul Kotyczka. Finite differences on staggered grids preserving the port-Hamiltonian structure with application to an acoustic duct. Journal of Computational Physics, 373(??):673–697, November 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304297.

Tredak:2016:EIM

[TRM16] Przemyslaw Tredak, Witold R. Rudnicki, and Jacek A. Majewski. Efficient implementation of the many-body Reactive Bond Order (REBO) potential on GPU. Journal of Computational Physics, 321(??):556–570, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302066.

Torrilhon:2017:HBS

[TS17] Manuel Torrilhon and Neeraj Sarna. Hierarchical Boltzmann simulations and model error estimation. Journal of Computational Physics, 342(??):66–84, August 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303200.

Thiagarajan:2018:SAQ

[TS18] Vaidyanathan Thiagarajan and Vadim Shapiro. Shape aware quadratures. Journal of Computational Physics, 374(??):1239–1260, December 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118303243.

Theillard:2019:CMF

[TS19] Maxime Theillard and David Saintillan. Computational meanfield modeling of confined active fluids. Journal of Computational Physics, 397(??):Article 108841, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930525X.

Tsalamengas:2015:QRW

[Tsa15] John L. Tsalamengas. Quadrature rules for weakly singular, strongly singular, and hypersingular integrals in boundary integral equation methods. Journal of Computational Physics, 303(??):498–513, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006622.

Tsalamengas:2016:GJQ

[Tsa16] John L. Tsalamengas. Gauss–Jacobi quadratures for weakly, strongly, hyper- and nearly-singular integrals in boundary integral equation methods for domains with sharp edges and corners. Journal of Computational Physics, 325(??):338–357, November 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303849.

Trahan:2018:FAA

[TSB^{+}18] C. J. Trahan, G. Savant, R. C. Berger, M. Farthing, T. O. McAlpin, L. Pettey, G. K. Choudhary, and C. N. Dawson. Formulation and application of the adaptive hydraulics three-dimensional shallow water and transport models. Journal of Computational Physics, 374(??):47–90, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302882.

Tang:2017:DGM

[TSC17] Wensheng Tang, Yajuan Sun, and Wenjun Cai. Discontinuous Galerkin methods for Hamiltonian ODEs and PDEs. Journal of Computational Physics, 330(??):340–364, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630609X.

Tramm:2017:RRM

[TSFS17] John R. Tramm, Kord S. Smith, Benoit Forget, and Andrew R. Siegel. The Random Ray Method for neutral particle transport. Journal of Computational Physics, 342(??):229–252, August 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303170.

Tayebi:2017:MMS

[TSH17] A. Tayebi, Y. Shekari, and M. H. Heydari. A meshless method for solving two-dimensional variable-order time fractional advection-diffusion equation. Journal of Computational Physics, 340(??):655–669, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302620.

Thiele:2016:BCA

[TSN16] Illia Thiele, Stefan Skupin, and Rachel Nuter. Boundary conditions for arbitrarily shaped and tightly focused laser pulses in electromagnetic codes. Journal of Computational Physics, 321(??):1110–1119, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302297.

Tsoutsanis:2018:EBL

[Tso18] Panagiotis Tsoutsanis. Extended bounds limiter for highorder finite-volume schemes on unstructured meshes. Journal of Computational Physics, 362(??):69–94, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300858.

Turrell:2015:SCI

[TSR15] A. E. Turrell, M. Sherlock, and S. J. Rose. Self-consistent inclusion of classical large-angle Coulomb collisions in plasma Monte Carlo simulations. Journal of Computational Physics, 299(??):144–155, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004325.

Troshin:2016:POD

[TSST16] Victor Troshin, Avi Seifert, David Sidilkover, and Gilead Tadmor. Proper orthogonal decomposition of flow-field in non-stationary geometry. Journal of Computational Physics, 311(??):329–337, April 15, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116000607.

Thompson:2015:SNA

[TST^{+}15] A. P. Thompson, L. P. Swiler, C. R. Trott, S. M. Foiles, and G. J. Tucker. Spectral neighbor analysis method for automated generation of quantum-accurate interatomic potentials. Journal of Computational Physics, 285(??):316–330, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008353.

Tipireddy:2017:BAD

[TST17] R. Tipireddy, P. Stinis, and A. M. Tartakovsky. Basis adaptation and domain decomposition for steady-state partial differential equations with random coefficients. Journal of Computational Physics, 351(??):203–215, December 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117306484.

Tokareva:2016:FSM

[TT16] S. A. Tokareva and E. F. Toro. A flux splitting method for the Baer–Nunziato equations of compressible two-phase flow. Journal of Computational Physics, 323(??):45–74, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630314X.

Taverniers:2017:IPU

[TT17a] Søren Taverniers and Daniel M. Tartakovsky. Impact of parametric uncertainty on estimation of the energy deposition into an irradiated brain tumor. Journal of Computational Physics, 348(??):139–150, November 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305107.

Taverniers:2017:TCD

[TT17b] Søren Taverniers and Daniel M. Tartakovsky. A tightlycoupled domain-decomposition approach for highly nonlinear stochastic multiphysics systems. Journal of Computational Physics, 330(??):884–901, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305605.

Trofimov:2019:EAE

[TT19] Vyacheslav A. Trofimov and Evgeny M. Trykin. Enhancement of ABCs efficiency at computer simulation of optical pulse interaction with inhomogeneous nonlinear medium. Journal of Computational Physics, 399(??):Article 108947, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119306527.

Togo:2016:SCT

[TTN^{+}16] Satoshi Togo, Tomonori Takizuka, Makoto Nakamura, Kazuo Hoshino, Kenzo Ibano, Tee Long Lang, and Yuichi Ogawa. Self-consistent treatment of the sheath boundary conditions by introducing anisotropic ion temperatures and virtual divertor model. Journal of Computational Physics, 310(??):109–126, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000127.

Taverniers:2019:TWC

[TUJ19] Søren Taverniers, H. S. Udaykumar, and Gustaaf B. Jacobs. Two-way coupled Cloud-In-Cell modeling of non-isothermal particle-laden flows: a Subgrid Particle-Averaged Reynolds Stress-Equivalent (SPARSE) formulation. Journal of Computational Physics, 390(??):595–618, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300166.

Terekhov:2019:FVM

[TV19] Kirill M. Terekhov and Yuri V. Vassilevski. Finite volume method for coupled subsurface flow problems, I: Darcy problem. Journal of Computational Physics, 395(??):298–306, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304164.

Todarello:2016:FVG

[TVB^{+}16] Giovanni Todarello, Floris Vonck, Sébastien Bourasseau, Jacques Peter, and Jean-Antoine Désidéri. Finite-volume goal-oriented mesh adaptation for aerodynamics using functional derivative with respect to nodal coordinates. Journal of Computational Physics, 313(??):799–819, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001352.

Tang:2017:APM

[TW17] Min Tang and Yihong Wang. An asymptotic preserving method for strongly anisotropic diffusion equations based on field line integration. Journal of Computational Physics, 330(??):735–748, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305708.

Tang:2019:ASP

[TWF19] Kunkun Tang, Jonathan M. Wang, and Jonathan B. Freund. Adaptive sparse polynomial dimensional decomposition for derivative-based sensitivity. Journal of Computational Physics, 391(??):303–321, August 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302888.

Tene:2015:AAM

[TWH15] Matei Tene, Yixuan Wang, and Hadi Hajibeygi. Adaptive algebraic multiscale solver for compressible flow in heterogeneous porous media. Journal of Computational Physics, 300(??):679–694, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005264.

Till:2018:ALM

[TWM18] Andrew T. Till, James S. Warsa, and Jim E. Morel. Application of linear multifrequency-grey acceleration to preconditioned Krylov iterations for thermal radiation transport. Journal of Computational Physics, 372(??):931–955, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303954.

Ta:2015:IFM

[TWN15] Catherine Ta, Dongyong Wang, and Qing Nie. An integration factor method for stochastic and stiff reaction-diffusion systems. Journal of Computational Physics, 295(??):505–522, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002843.

Tanaka:2019:MPM

[TWN19] Shusei Tanaka, Tomoaki Watanabe, and Koji Nagata. Multiparticle model of coarse-grained scalar dissipation rate with volumetric tensor in turbulence. Journal of Computational Physics, 389(??):128–146, July 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302177.

Tian:2015:LDG

[TXKvdV15] Lulu Tian, Yan Xu, J. G. M. Kuerten, and J. J. W. van der Vegt. A local discontinuous Galerkin method for the (non)-isothermal Navier–Stokes–Korteweg equations. Journal of Computational Physics, 295(??):685–714, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002818.

Tian:2016:ALD

[TXKvdV16] Lulu Tian, Yan Xu, J. G. M. Kuerten, and J. J. W. van der Vegt. An h-adaptive local discontinuous Galerkin method for the Navier–Stokes–Korteweg equations. Journal of Computational Physics, 319(??):242–265, August 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301644.

Tang:2017:HAD

[TY17] Shaoqiang Tang and Yuping Ying. Homogenizing atomic dynamics by fractional differential equations. Journal of Computational Physics, 346(??):539–551, October 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304928.

Tao:2016:SSD

[TYD16] Yutao Tao, Andrew Yeckel, and Jeffrey J. Derby. Steady-state and dynamic models for particle engulfment during solidification. Journal of Computational Physics, 315(??):238–263, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300092.

Tyson:2019:RET

[TYROG19] William C. Tyson, Gary K. Yan, Christopher J. Roy, and Carl F. Ollivier-Gooch. Relinearization of the error transport equations for arbitrarily high-order error estimates. Journal of Computational Physics, 397(??):Article 108867, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119305583.

Tang:2016:NEM

[TZ16] Zhili Tang and Lianhe Zhang. Nash equilibrium and multi criterion aerodynamic optimization. Journal of Computational Physics, 314(??):107–126, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001455.

Tao:2018:CIB

[TZGW18] Shi Tao, Haolong Zhang, Zhaoli Guo, and Lian-Ping Wang. A combined immersed boundary and discrete unified gas kinetic scheme for particle-fluid flows. Journal of Computational Physics, 375(??):498–518, December 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305783.

Teichert:2017:IRG

[TZSS17] Fabian Teichert, Andreas Zienert, Jörg Schuster, and Michael Schreiber. Improved recursive Green’s function formalism for quasi one-dimensional systems with realistic defects. Journal of Computational Physics, 334(??):607–619, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300347.

Unnikrishnan:2016:HFM

[UG16] S. Unnikrishnan and Datta V. Gaitonde. A high-fidelity method to analyze perturbation evolution in turbulent flows. Journal of Computational Physics, 310(??):45–62, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000188.

Um:2019:CBN

[UHKT19] Kimoon Um, Eric J. Hall, Markos A. Katsoulakis, and Daniel M. Tartakovsky. Causality and Bayesian network PDEs for multiscale representations of porous media. Journal of Computational Physics, 394(??):658–678, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119304140.

Kim:2019:HOF

[uKHGK19] Hyeon uk Kim, Heejae Han, Dalhyeon Gwon, and Myungjoo Kang. High order face-offsetting method for interface tracking problem using WENO schemes. Journal of Computational Physics, 376(??):863–893, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306223.

Ueckermann:2016:HDG

[UL16] M. P. Ueckermann and P. F. J. Lermusiaux. Hybridizable discontinuous Galerkin projection methods for Navier– Stokes and Boussinesq equations. Journal of Computational Physics, 306(??):390–421, February 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007688.

Ullrich:2018:IIH

[URGT18] Paul A. Ullrich, Daniel R. Reynolds, Jorge E. Guerra, and Mark A. Taylor. Impact and importance of hyperdiffusion on the spectral element method: a linear dispersion analysis. Journal of Computational Physics, 375(??):427–446, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304133.

Ullmann:2016:PGR

[URL16] Sebastian Ullmann, Marko Rotkvic, and Jens Lang. POD– Galerkin reduced-order modeling with adaptive finite element snapshots. Journal of Computational Physics, 325(??): 244–258, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116303679.

Uber:2017:ANS

[UWH17] Richard Uber, Aihua Wood, and Michael Havrilla. Analysis and numerical solution of transient electromagnetic scattering from two cavities. Journal of Computational Physics, 343(??):217–234, August 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117303224.

Upperman:2019:ESA

[UY19] Johnathon Upperman and Nail K. Yamaleev. Entropy stable artificial dissipation based on Brenner regularization of the Navier–Stokes equations. Journal of Computational Physics, 393(??):74–91, September 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303316.

Vabishchevich:2015:NSE

[Vab15] Petr N. Vabishchevich. Numerically solving an equation for fractional powers of elliptic operators. Journal of Computational Physics, 282(??):289–302, February 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007840.

Vabishchevich:2018:TLS

[Vab18] Petr N. Vabishchevich. Two-level schemes for the advection equation. Journal of Computational Physics, 363(??):158–177, June 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911830127X.

Villamizar:2017:HOL

[VAD17] Vianey Villamizar, Sebastian Acosta, and Blake Dastrup. High order local absorbing boundary conditions for acoustic waves in terms of farfield expansions. Journal of Computational Physics, 333(??):331–351, March 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116307124.

Vaibhav:2015:TBC

[Vai15] V. Vaibhav. Transparent boundary condition for numerical modeling of intense laser-molecule interaction. Journal of Computational Physics, 283(??):478–494, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114008122.

Villegas:2016:GFL

[VALT16] Lucia Rueda Villegas, Romain Alis, Mathieu Lepilliez, and Sébastien Tanguy. A Ghost Fluid/Level Set Method for boiling flows and liquid evaporation: Application to the Leidenfrost effect. Journal of Computational Physics, 316(??):789– 813, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300778.

Vishnampet:2015:PDA

[VBF15] Ramanathan Vishnampet, Daniel J. Bodony, and Jonathan B. Freund. A practical discrete-adjoint method for high-fidelity compressible turbulence simulations. Journal of Computational Physics, 285(??):173–192, March 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000133.

Villa:2015:SDE

[VBG^{+}15] Andrea Villa, Luca Barbieri, Marco Gondola, Andres R. Leon-Garzon, and Roberto Malgesini. Stability of the discretization of the electron avalanche phenomenon. Journal of Computational Physics, 296(??):369–381, September 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115003411.

Veeraragavan:2016:UMM

[VBG16] A. Veeraragavan, J. Beri, and R. J. Gollan. Use of the method of manufactured solutions for the verification of conjugate heat transfer solvers. Journal of Computational Physics, 307(??):308–320, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008190.

Villa:2017:EAC

[VBG^{+}17a] Andrea Villa, Luca Barbieri, Marco Gondola, Andres R. Leon-Garzon, and Roberto Malgesini. An efficient algorithm for corona simulation with complex chemical models. Journal of Computational Physics, 337(??):233–251, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301328.

Villa:2017:PBP

[VBG^{+}17b] Andrea Villa, Luca Barbieri, Marco Gondola, Andres R. Leon-Garzon, and Roberto Malgesini. A PDE-based partial discharge simulator. Journal of Computational Physics, 345(??):687–705, September 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911730428X.

Vasil:2016:TCP

[VBL^{+}16] Geoffrey M. Vasil, Keaton J. Burns, Daniel Lecoanet, Sheehan Olver, Benjamin P. Brown, and Jeffrey S. Oishi. Tensor calculus in polar coordinates using Jacobi polynomials. Journal of Computational Physics, 325(??):53–73, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630362X.

Vasilyeva:2019:CEM

[VCEK19] Maria Vasilyeva, Eric T. Chung, Yalchin Efendiev, and Jihoon Kim. Constrained energy minimization based upscaling for coupled flow and mechanics. Journal of Computational Physics, 376(??):660–674, January 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306569.

Vidal-Codina:2015:MVR

[VCNGP15] F. Vidal-Codina, N. C. Nguyen, M. B. Giles, and J. Peraire. A model and variance reduction method for computing statistical outputs of stochastic elliptic partial differential equations. Journal of Computational Physics, 297(??):700–720, September 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003757.

Vidal-Codina:2018:HDG

[VCNOP18] F. Vidal-Codina, N. C. Nguyen, S.-H. Oh, and J. Peraire. A hybridizable discontinuous Galerkin method for computing nonlocal electromagnetic effects in three-dimensional metallic nanostructures. Journal of Computational Physics, 355 (??):548–565, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911730863X.

Vidal-Codina:2018:CPS

[VCNP18] F. Vidal-Codina, N. C. Nguyen, and J. Peraire. Computing parametrized solutions for plasmonic nanogap structures. Journal of Computational Physics, 365(??):89–106, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302237.

Vo:2016:RKP

[VD16] Hai X. Vo and Louis J. Durlofsky. Regularized kernel PCA for the efficient parameterization of complex geological models. Journal of Computational Physics, 322(??):859–881, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911630290X.

vandenBos:2017:NIU

[vdBKD17] L. M. M. van den Bos, B. Koren, and R. P. Dwight. Nonintrusive uncertainty quantification using reduced cubature rules. Journal of Computational Physics, 332(??):418–445, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116306647.

vanderKaap:2016:MPK

[vdKK16] N. J. van der Kaap and L. J. A. Koster. Massively parallel kinetic Monte Carlo simulations of charge carrier transport in organic semiconductors. Journal of Computational Physics, 307(??):321–332, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008153.

vanderLinden:2016:PSL

[vdLJLV16] J. H. van der Linden, T. B. Jönsthövel, A. A. Lukyanov, and C. Vuik. The parallel subdomain-levelset deflation method in reservoir simulation. Journal of Computational Physics, 304(??):340–358, January 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006804.

Vidovic:2015:PLT

[VDPP15] D. Vidović, M. Dotlić, M. Pusić, and B. Pokorni. Piecewise linear transformation in diffusive flux discretization. Journal of Computational Physics, 282(??):227–237, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007864.

Veerapaneni:2016:IEM

[Vee16] Shravan Veerapaneni. Integral equation methods for vesicle electrohydrodynamics in three dimensions. Journal of Computational Physics, 326(??):278–289, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304090.

vanEs:2016:FVS

[vEKdB16] Bram van Es, Barry Koren, and Hugo J. de Blank. Finitevolume scheme for anisotropic diffusion. Journal of Computational Physics, 306(??):422–442, February 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007810.

Veldman:2019:GCK

[Vel19] Arthur E. P. Veldman. A general condition for kinetic-energy preserving discretization of flow transport equations. Journal of Computational Physics, 398(??):Article 108894, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119305923.

Vermeire:2019:PER

[Ver19] Brian C. Vermeire. Paired explicit Runge–Kutta schemes for stiff systems of equations. Journal of Computational Physics, 393(??):465–483, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303390.

Vico:2016:FCF

[VGF16] Felipe Vico, Leslie Greengard, and Miguel Ferrando. Fast convolution with free-space Green’s functions. Journal of Computational Physics, 323(??):191–203, October 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303230.

Velikovich:2018:GNS

[VGZ18] A. L. Velikovich, J. L. Giuliani, and S. T. Zalesak. Generalized Noh self-similar solutions of the compressible Euler equations for hydrocode verification. Journal of Computational Physics, 374(??):843–862, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305175.

Vilar:2019:PCH

[Vil19] François Vilar. A posteriori correction of high-order discontinuous Galerkin scheme through subcell finite volume formulation and flux reconstruction. Journal of Computational Physics, 387(??):245–279, June 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307174.

Virta:2015:IWA

[VK15] Kristoffer Virta and Gunilla Kreiss. Interface waves in almost incompressible elastic materials. Journal of Computational Physics, 303(??):313–330, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006592.

Vreman:2016:TOM

[VK16] A. W. Vreman and J. G. M. Kuerten. A third-order multistep time discretization for a Chebyshev tau spectral method. Journal of Computational Physics, 304(??):162–169, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115006865.

Vinas:2018:FBV

[VK18] Adolfo. F. Viñas and Alexander J. Klimas. Flux-balance Vlasov simulation with filamentation filtration. Journal of Computational Physics, 375(??):983–1004, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118306089.

Veske:2018:DCF

[VKE^{+}18] Mihkel Veske, Andreas Kyritsakis, Kristjan Eimre, Vahur Zadin, Alvo Aabloo, and Flyura Djurabekova. Dynamic coupling of a finite element solver to large-scale atomistic simulations. Journal of Computational Physics, 367(??):279–294, August 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302560.

Violeau:2015:OTS

[VL15] Damien Violeau and Agnès Leroy. Optimal time step for incompressible SPH. Journal of Computational Physics, 288 (??):119–130, May 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115000856.

VanLangenhove:2018:GOE

[VLAB18] J. Van Langenhove, D. Lucor, F. Alauzet, and A. Belme. Goal-oriented error control of stochastic system approximations using metric-based anisotropic adaptations. Journal of Computational Physics, 374(??):384–412, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305059.

Voronin:2018:STD

[VLN^{+}18] Kirill Voronin, Chak Shing Lee, Martin Neumüller, Paulina Sepulveda, and Panayot S. Vassilevski. Space-time discretizations using constrained first-order system least squares (CFOSLS). Journal of Computational Physics, 373(??):863– 876, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118304856.

Vergara:2016:CNM

[VLP^{+}16] Christian Vergara, Matthias Lange, Simone Palamara, Toni Lassila, Alejandro F. Frangi, and Alfio Quarteroni. A coupled 3D–1D numerical monodomain solver for cardiac electrical activation in the myocardium with detailed Purkinje network. Journal of Computational Physics, 308(??):218–238, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008359.

Vidal:2016:PDS

[VLTPS16] A. Báez Vidal, O. Lehmkuhl, F. X. Trias, and C. D. Pérez-Segarra. On the properties of discrete spatial filters for CFD. Journal of Computational Physics, 326(??):474–498, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304144.

vanLith:2017:EWD

[vLtTBI17] Bart S. van Lith, Jan H. M. ten Thije Boonkkamp, and Wilbert L. IJzerman. Embedded WENO: a design strategy to improve existing WENO schemes. Journal of Computational Physics, 330(??):529–549, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911630612X.

Vermeire:2019:ORK

[VLV19] B. C. Vermeire, N. A. Loppi, and P. E. Vincent. Optimal Runge–Kutta schemes for pseudo time-stepping with high-order unstructured methods. Journal of Computational Physics, 383(??):55–71, April 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911930018X.

Valizadeh:2015:SSW

[VM15] Alireza Valizadeh and Joseph J. Monaghan. A study of solid wall models for weakly compressible SPH. Journal of Computational Physics, 300(??):5–19, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004787.

Vymazal:2019:WDB

[VMC^{+}19] Martin Vymazal, David Moxey, Chris D. Cantwell, Spencer J. Sherwin, and Robert M. Kirby. On weak Dirichlet boundary conditions for elliptic problems in the continuous Galerkin method. Journal of Computational Physics, 394(??):732–744, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303523.

Vesselinov:2019:UML

[VMK^{+}19] V. V. Vesselinov, M. K. Mudunuru, S. Karra, D. O’Malley, and B. S. Alexandrov. Unsupervised machine learning based on non-negative tensor factorization for analyzing reactivemixing. Journal of Computational Physics, 395(??):85–104, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119303833.

Velasco:2019:LBM

[VMM19] A. M. Velasco, J. D. Muñoz, and M. Mendoza. Lattice Boltzmann model for the simulation of the wave equation in curvilinear coordinates. Journal of Computational Physics, 376(??):76–97, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306338.

Beeumen:2018:CRM

[VMN^{+}18] Roel Van Beeumen, Osni Marques, Esmond G. Ng, Chao Yang, Zhaojun Bai, Lixin Ge, Oleksiy Kononenko, Zenghai Li, Cho-Kuen Ng, and Liling Xiao. Computing resonant modes of accelerator cavities by solving nonlinear eigenvalue problems via rational approximation. Journal of Computational Physics, 374(??):1031–1043, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305400.

Vermeire:2015:AIS

[VN15] Brian C. Vermeire and Siva Nadarajah. Adaptive IMEX schemes for high-order unstructured methods. Journal of Computational Physics, 280(??):261–286, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006524.

Vides:2015:STD

[VNA15] Jeaniffer Vides, Boniface Nkonga, and Edouard Audit. A simple two-dimensional extension of the HLL Riemann solver for hyperbolic systems of conservation laws. Journal of Computational Physics, 280(??):643–675, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006950.

Vogl:2017:CAC

[Vog17] Christopher J. Vogl. The Curvature-Augmented Closest Point method with vesicle inextensibility application. Journal of Computational Physics, 345(??):818–833, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304497.

vanOers:2017:HDG

[vOMB17] Alexander M. van Oers, Leo R. M. Maas, and Onno Bokhove. Hamiltonian discontinuous Galerkin FEM for linear, stratified (in)compressible Euler equations: internal gravity waves. Journal of Computational Physics, 330(??):770–793, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305307.

Voskov:2017:OBL

[Vos17] Denis V. Voskov. Operator-based linearization approach for modeling of multiphase multi-component flow in porous media. Journal of Computational Physics, 337(??):275–288, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301444.

Vanharen:2015:TNA

[VPM15] Julien Vanharen, Guillaume Puigt, and Marc Montagnac. Theoretical and numerical analysis of nonconforming grid interface for unsteady flows. Journal of Computational Physics, 285(??):111–132, March 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115000170.

Vanharen:2017:RSA

[VPV^{+}17] Julien Vanharen, Guillaume Puigt, Xavier Vasseur, Jean-François Boussuge, and Pierre Sagaut. Revisiting the spectral analysis for high-order spectral discontinuous methods. Journal of Computational Physics, 337(??):379–402, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301468.

Vreman:2017:SOG

[Vre17] A. W. Vreman. A staggered overset grid method for resolved simulation of incompressible flow around moving spheres. Journal of Computational Physics, 333(??):269–296, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306817. See corrigendum

[Vre21].

Vreman:2021:CSO

[Vre21] A. W. Vreman. Corrigendum to “A staggered overset grid method for resolved simulation of incompressible flow around moving spheres” [j. comput. phys. 333 (2017) 269–296]. Journal of Computational Physics, 435(??):Article 110302, June 15, 2021. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999121001972. See [Vre17].

Ververis:2017:CIP

[VS17] Antonios Ververis and Markus Schmuck. Computational investigation of porous media phase field formulations: Microscopic, effective macroscopic, and Langevin equations. Journal of Computational Physics, 344(??):485–498, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303534.

Vogman:2018:CFO

[VSC18] G. V. Vogman, U. Shumlak, and P. Colella. Conservative fourth-order finite-volume Vlasov–Poisson solver for axisymmetric plasmas in cylindrical (r, v_{r}, v_{θ}) phase space coordinates. Journal of Computational Physics, 373(??):877–899, November 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830490X.

Vamaraju:2018:EGF

[VSDW18] Janaki Vamaraju, Mrinal K. Sen, Jonas De Basabe, and Mary Wheeler. Enriched Galerkin finite element approximation for elastic wave propagation in fractured media. Journal of Computational Physics, 372(??):726–747, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304273.

Vilar:2016:PPCa

[VSM16a] François Vilar, Chi-Wang Shu, and Pierre-Henri Maire. Positivity-preserving cell-centered Lagrangian schemes for multi-material compressible flows: From first-order to highorders. Part I: the one-dimensional case. Journal of Computational Physics, 312(??):385–415, May 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000917.

Vilar:2016:PPCb

[VSM16b] François Vilar, Chi-Wang Shu, and Pierre-Henri Maire. Positivity-preserving cell-centered Lagrangian schemes for multi-material compressible flows: From first-order to highorders. Part II: the two-dimensional case. Journal of Computational Physics, 312(??):416–442, May 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000802.

VandePut:2017:ESW

[VSM17] Maarten L. Van de Put, Bart Sorée, and Wim Magnus. Efficient solution of the Wigner–Liouville equation using a spectral decomposition of the force field. Journal of Computational Physics, 350(??):314–325, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730640X.

Vanzo:2016:PTS

[VST16] Davide Vanzo, Annunziato Siviglia, and Eleuterio F. Toro. Pollutant transport by shallow water equations on unstructured meshes: Hyperbolization of the model and numerical solution via a novel flux splitting scheme. Journal of Computational Physics, 321(??):1–20, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301607.

Vermeire:2016:PES

[VV16] B. C. Vermeire and P. E. Vincent. On the properties of energy stable flux reconstruction schemes for implicit large eddy simulation. Journal of Computational Physics, 327(??):368–388, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304557.

Viguerie:2019:DBS

[VV19] Alex Viguerie and Alessandro Veneziani. Deconvolution-based stabilization of the incompressible Navier–Stokes equations. Journal of Computational Physics, 391(??):226–242, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830768X.

Vachal:2016:PSS

[VW16] Pavel Váchal and Burton Wendroff. On preservation of symmetry in r–z staggered Lagrangian schemes. Journal of Computational Physics, 307(??):496–507, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008268.

Vachal:2018:VCE

[VW18] Pavel Váchal and Burton Wendroff. Volume change and energy exchange: How they affect symmetry in the Noh problem. Journal of Computational Physics, 364(??):416–419, July 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301815.

Vermeire:2017:UGA

[VWV17] B. C. Vermeire, F. D. Witherden, and P. E. Vincent. On the utility of GPU accelerated high-order methods for unsteady flow simulations: a comparison with industry-standard tools. Journal of Computational Physics, 334(??):497–521, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116307136.

Vecharynski:2015:PPC

[VYP15] Eugene Vecharynski, Chao Yang, and John E. Pask. A projected preconditioned conjugate gradient algorithm for computing many extreme eigenpairs of a Hermitian matrix. Journal of Computational Physics, 290(??):73–89, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500100X.

Wu:2018:HAF

[WA18] Zedong Wu and Tariq Alkhalifah. A highly accurate finitedifference method with minimum dispersion error for solving the Helmholtz equation. Journal of Computational Physics, 366(??):350–361, July 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302134.

Waclawczyk:2015:CSR

[Wac15] Tomasz Waclawczyk. A consistent solution of the reinitialization equation in the conservative level-set method. Journal of Computational Physics, 299(??):487–525, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004271.

Wieland:2019:ENF

[WAF^{+}19] Manuel Wieland, Walter Arne, Robert Feßler, Nicole Marheineke, and Raimund Wegener. An efficient numerical framework for fiber spinning scenarios with evaporation effects in airflows. Journal of Computational Physics, 384(??):326–348, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300853.

Walker:2016:SOS

[Wal16] Shawn W. Walker. Shape optimization of self-avoiding curves. Journal of Computational Physics, 311(??):275–298, April 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000656.

Wu:2019:PLR

[WAZ19] Zedong Wu, Tariq Alkhalifah, and Zhendong Zhang. A partial-low-rank method for solving acoustic wave equation. Journal of Computational Physics, 385(??):1–12, May 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301123.

Wang:2016:SEA

[WB16] Mu Wang and John F. Brady. Spectral Ewald acceleration of Stokesian dynamics for polydisperse suspensions. Journal of Computational Physics, 306(??):443–477, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007822.

Weiner:2017:ASS

[WB17] Andre Weiner and Dieter Bothe. Advanced subgrid-scale modeling for convection-dominated species transport at fluid interfaces with application to mass transfer from rising bubbles. Journal of Computational Physics, 347(??):261–289, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304941.

Weller:2016:MAS

[WBBC16] Hilary Weller, Philip Browne, Chris Budd, and Mike Cullen. Mesh adaptation on the sphere using optimal transport and the numerical solution of a Monge–Ampère type equation. Journal of Computational Physics, 308(??):102–123, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008372.

Weitz:2016:MCE

[WBC^{+}16] Sebastian Weitz, Stéphane Blanco, Julien Charon, Jérémi Dauchet, Mouna El Hafi, Vincent Eymet, Olivier Farges, Richard Fournier, and Jacques Gautrais. Monte Carlo efficiency improvement by multiple sampling of conditioned integration variables. Journal of Computational Physics, 326(??):30–34, December 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911630393X.

Whalen:2015:ETD

[WBM15a] P. Whalen, M. Brio, and J. V. Moloney. Exponential timedifferencing with embedded Runge–Kutta adaptive step control. Journal of Computational Physics, 280(??):579–601, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006743.

Wirasaet:2015:ABL

[WBM^{+}15b] D. Wirasaet, S. R. Brus, C. E. Michoski, E. J. Kubatko, J. J. Westerink, and C. Dawson. Artificial boundary layers in discontinuous Galerkin solutions to shallow water equations in channels. Journal of Computational Physics, 299(??):597–612, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500460X.

Wen:2018:RMP

[WC18] Baole Wen and Gregory P. Chini. Reduced modeling of porous media convection in a minimal flow unit at large Rayleigh number. Journal of Computational Physics, 371(??):551–563, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303784.

Wang:2019:ABC

[WC19] Yanli Wang and Zhenning Cai. Approximation of the Boltzmann collision operator based on Hermite spectral method. Journal of Computational Physics, 397(??):Article 108815, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119304991.

Wolf:2016:PCM

[WCCB16] Eric M. Wolf, Matthew Causley, Andrew Christlieb, and Matthew Bettencourt. A particle-in-cell method for the simulation of plasmas based on an unconditionally stable field solver. Journal of Computational Physics, 326(??):342–372, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303461.

Wang:2017:IBM

[WCH^{+}17] Li Wang, Gaetano M. D. Currao, Feng Han, Andrew J. Neely, John Young, and Fang-Bao Tian. An immersed boundary method for fluid-structure interaction with compressible multiphase flows. Journal of Computational Physics, 346(??):131–151, October 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304539.

Wilkening:2015:ASN

[WCL15] Jon Wilkening, Antoine J. Cerfon, and Matt Landreman. Accurate spectral numerical schemes for kinetic equations with energy diffusion. Journal of Computational Physics, 294(??):58–77, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001941.

Wang:2015:SII

[WCN15] Dongyong Wang, Weitao Chen, and Qing Nie. Semi-implicit integration factor methods on sparse grids for high-dimensional systems. Journal of Computational Physics, 292(??):43–55, July 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001801.

Wise:2018:BBM

[WCT18] Elliott S. Wise, Ben T. Cox, and Bradley E. Treeby. Bandwidth-based mesh adaptation in multiple dimensions. Journal of Computational Physics, 371(??):651–662, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303875.

Wrobel:2016:RIS

[WCVF16] Jacek K. Wróbel, Ricardo Cortez, Douglas Varela, and Lisa Fauci. Regularized image system for Stokes flow outside a solid sphere. Journal of Computational Physics, 317(??):165–184, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300894.

Wang:2019:VLDb

[WCWY19] Yanxing Wang, Xiaodong Chen, Xingjian Wang, and Vigor Yang. Vaporization of liquid droplet with large deformation and high mass transfer rate, II: Variabledensity, variable-property case. Journal of Computational Physics, 394(??):1–17, October 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303055.

Wackers:2017:CAG

[WDG^{+}17] Jeroen Wackers, Ganbo Deng, Emmanuel Guilmineau, Alban Leroyer, Patrick Queutey, Michel Visonneau, Alexandro Palmieri, and Alfredo Liverani. Can adaptive grid refinement produce grid-independent solutions for incompressible flows? Journal of Computational Physics, 344(??):364–380, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303650.

Winters:2017:UDE

[WDGW17] Andrew R. Winters, Dominik Derigs, Gregor J. Gassner, and Stefanie Walch. A uniquely defined entropy stable matrix dissipation operator for high Mach number ideal MHD and compressible Euler simulations. Journal of Computational Physics, 332(??):274–289, March 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306477.

Wang:2015:PAM

[WDS15] Cheng Wang, XinZhuang Dong, and Chi-Wang Shu. Parallel adaptive mesh refinement method based on WENO finite difference scheme for the simulation of multi-dimensional detonation. Journal of Computational Physics, 298(??):161–175, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115003800.

Wang:2019:SPF

[WDT^{+}19] Zhicheng Wang, Suchuan Dong, Michael S. Triantafyllou, Yiannis Constantinides, and George Em Karniadakis. A stabilized phase-field method for two-phase flow at high Reynolds number and large density/viscosity ratio. Journal of Computational Physics, 397(??):Article 108832, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305169.

Wang:2015:SCD

[WE15] Chengjie Wang and Jeff D. Eldredge. Strongly coupled dynamics of fluids and rigid-body systems with the immersed boundary projection method. Journal of Computational Physics, 295(??):87–113, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002454.

Webb:2014:SIM

[Web14] Stephen D. Webb. Symplectic integration of magnetic systems. Journal of Computational Physics, 270(??):570–576, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002368. See com-

ments [ZJS15, EBQ15].

Wendt:2015:PCS

[WED15] Gunnar Wendt, Patrick Erbts, and Alexander Düster. Partitioned coupling strategies for multi-physically coupled radiative heat transfer problems. Journal of Computational Physics, 300(??):327–351, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005161.

Wright:2017:SCF

[WF17] Grady B. Wright and Bengt Fornberg. Stable computations with flat radial basis functions using vectorvalued rational approximations. Journal of Computational Physics, 331(??):137–156, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306222.

Winters:2015:CTE

[WG15] Andrew R. Winters and Gregor J. Gassner. A comparison of two entropy stable discontinuous Galerkin spectral element approximations for the shallow water equations with non-constant topography. Journal of Computational Physics, 301(??):357–376, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005616.

Waindim:2016:BFB

[WG16a] M. Waindim and D. V. Gaitonde. A body-force based method to generate supersonic equilibrium turbulent boundary layer profiles. Journal of Computational Physics, 304(??):1–26, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500666X.

Winters:2016:AEC

[WG16b] Andrew R. Winters and Gregor J. Gassner. Affordable, entropy conserving and entropy stable flux functions for the ideal MHD equations. Journal of Computational Physics, 304(??):72–108, January 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006737.

Wenzel:2019:PMP

[WG19] E. A. Wenzel and S. C. Garrick. A point-mass particle method for the simulation of immiscible multiphase flows on an Eulerian grid. Journal of Computational Physics, 397(??):Article 108835, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305194.

Watwood:2019:ECG

[WGJS19] Matthew Watwood, Ian Grooms, Keith A. Julien, and K. Shafer Smith. Energy-conserving Galerkin approximations for quasigeostrophic dynamics. Journal of Computational Physics, 388(??):23–40, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302128.

Wissocq:2017:RCB

[WGME17] Gauthier Wissocq, Nicolas Gourdain, Orestis Malaspinas, and Alexandre Eyssartier. Regularized characteristic boundary conditions for the Lattice-Boltzmann methods at high Reynolds number flows. Journal of Computational Physics, 331(??):1–18, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306295.

Wang:2015:ECD

[WH15] Pengde Wang and Chengming Huang. An energy conservative difference scheme for the nonlinear fractional Schrödinger equations. Journal of Computational Physics, 293(??):238–251, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002241.

Wang:2016:IMD

[WH16a] Pengde Wang and Chengming Huang. An implicit midpoint difference scheme for the fractional Ginzburg–Landau equation. Journal of Computational Physics, 312(??):31–49, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116000723.

Wu:2016:BDB

[WH16b] Zhizhang Wu and Zhongyi Huang. A Bloch decompositionbased stochastic Galerkin method for quantum dynamics with a random external potential. Journal of Computational Physics, 317(??):257–275, July 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301061.

Wang:2017:CSP

[WHCN17] Lijin Wang, Xiaoying Han, Yanzhao Cao, and Habib N. Najm. Computational singular perturbation analysis of stochastic chemical systems with stiffness. Journal of Computational Physics, 335(??):404–425, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300566.

Wiesenberger:2017:SIM

[WHE17] M. Wiesenberger, M. Held, and L. Einkemmer. Streamline integration as a method for two-dimensional elliptic grid generation. Journal of Computational Physics, 340(??):435–450, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302577.

Wiesenberger:2018:SIM

[WHEK18] M. Wiesenberger, M. Held, L. Einkemmer, and A. Kendl. Streamline integration as a method for structured grid generation in X-point geometry. Journal of Computational Physics, 373(??):370–384, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118304625.

Wang:2017:NDF

[WHL17] Qi Wang, Yanren Hou, and Jingzhi Li. Numerical design of FSHL-based approximate cloaks with arbitrary shapes. Journal of Computational Physics, 333(??):142–159, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306994.

Wang:2019:NIR

[WHR19] Qian Wang, Jan S. Hesthaven, and Deep Ray. Non-intrusive reduced order modeling of unsteady flows using artificial neural networks with application to a combustion problem. Journal of Computational Physics, 384(??):289–307, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300828.

Wang:2019:OOB

[WHRL19] Yuepeng Wang, Kun Hu, Lanlan Ren, and Guang Lin. Optimal observations-based retrieval of topography in 2D shallow water equations using PC-EnKF. Journal of Computational Physics, 382(??):43–60, April 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300208.

Wong:2018:SIN

[WHT18] Zhi Yang Wong, Roland N. Horne, and Hamdi A. Tchelepi. Sequential implicit nonlinear solver for geothermal simulation. Journal of Computational Physics, 368(??):236–253, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302687.

Wang:2017:PST

[WHY17] Shuai Wang, Xudeng Hang, and Guangwei Yuan. A pyramid scheme for three-dimensional diffusion equations on polyhedral meshes. Journal of Computational Physics, 350(??):590–606, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306411.

Wang:2018:PPP

[WHY18] Shuai Wang, Xudeng Hang, and Guangwei Yuan. A positivitypreserving pyramid scheme for anisotropic diffusion problems on general hexahedral meshes with nonplanar cell faces. Journal of Computational Physics, 371(??):152–167, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118303267.

Wu:2018:MMF

[WHZ18] Shengyang Wu, Xianliang Hu, and Shengfeng Zhu. A multi-mesh finite element method for phase-field based photonic band structure optimization. Journal of Computational Physics, 357(??):324–337, March 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117309257.

Wick:2016:CFS

[Wic16] Thomas Wick. Coupling fluid-structure interaction with phase-field fracture. Journal of Computational Physics, 327 (??):67–96, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116304387.

Williams:2018:SGP

[Wil18] R. J. R. Williams. Sub-grid properties and artificial viscous stresses in staggered-mesh schemes. Journal of Computational Physics, 374(??):413–443, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830473X.

Williams:2019:FCC

[Wil19] R. J. R. Williams. Fully-conservative contact-capturing schemes for multi-material advection. Journal of Computational Physics, 398(??):Article 108809, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304930.

Wang:2016:ESE

[WJD16] Xiaoqiang Wang, Lili Ju, and Qiang Du. Efficient and stable exponential time differencing Runge–Kutta methods for phase field elastic bending energy models. Journal of Computational Physics, 316(??):21–38, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300365.

Wala:2018:FAE

[WK18] Matt Wala and Andreas Klöckner. A fast algorithm with error bounds for Quadrature by Expansion. Journal of Computational Physics, 374(??):135–162, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302985.

Wala:2019:FAQ

[WK19] Matt Wala and Andreas Klöckner. A fast algorithm for Quadrature by Expansion in three dimensions. Journal of Computational Physics, 388(??):655–689, July 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302074.

Wong:2019:SIN

[WKHT19] Zhi Yang Wong, Felix Kwok, Roland N. Horne, and Hamdi A. Tchelepi. Sequential-implicit Newton method for multiphysics simulation. Journal of Computational Physics, 391(??):155–178, August 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302566.

Welch:2017:GPC

[WKOE17] J. A. Welch, J. Kópházi, A. R. Owens, and M. D. Eaton. A geometry preserving, conservative, mesh-to-mesh isogeometric interpolation algorithm for spatial adaptivity of the multigroup, second-order even-parity form of the neutron transport equation. Journal of Computational Physics, 347(??):129–146, October 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304606.

Waruszewski:2018:MTO

[WKPS18] Maciej Waruszewski, Christian Kühnlein, Hanna Pawlowska, and Piotr K. Smolarkiewicz. MPDATA: Third-order accuracy for variable flows. Journal of Computational Physics, 359(??):361–379, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118300159.

Wu:2015:SFC

[WKSS15] Yuanqing Wu, Christoph Kowitz, Shuyu Sun, and Amgad Salama. Speeding up the flash calculations in two-phase compositional flow simulations — the application of sparse grids. Journal of Computational Physics, 285(??):88–99, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000169.

Wu:2016:SAM

[WL16] Keyi Wu and Jinglai Li. A surrogate accelerated multicanonical Monte Carlo method for uncertainty quantification. Journal of Computational Physics, 321(??):1098–1109, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116302455.

Wong:2017:HOL

[WL17] Man Long Wong and Sanjiva K. Lele. High-order localized dissipation weighted compact nonlinear scheme for shock- and interface-capturing in compressible flows. Journal of Computational Physics, 339(??):179–209, June 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911730195X.

Wang:2018:NSI

[WL18] Haibing Wang and Yi Li. Numerical solution of an inverse boundary value problem for the heat equation with unknown inclusions. Journal of Computational Physics, 369 (??):1–15, September 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118303000.

Wang:2015:SPB

[WLC15] Kun Wang, Hui Liu, and Zhangxin Chen. A scalable parallel black oil simulator on distributed memory parallel computers. Journal of Computational Physics, 301(??):19–34, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005355.

Ward:2017:DGM

[WLE17] N. F. Dudley Ward, T. Lähivaara, and S. Eveson. A discontinuous Galerkin method for poroelastic wave propagation: the two-dimensional case. Journal of Computational Physics, 350(??):690–727, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306587.

Wang:2018:IFO

[WLGD18] Bao-Shan Wang, Peng Li, Zhen Gao, and Wai Sun Don. An improved fifth order alternative WENO-Z finite difference scheme for hyperbolic conservation laws. Journal of Computational Physics, 374(??):469–477, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305151.

Wu:2016:ETM

[WLK^{+}16] Lingfei Wu, Jesse Laeuchli, Vassilis Kalantzis, Andreas Stathopoulos, and Efstratios Gallopoulos. Estimating the trace of the matrix inverse by interpolating from the diagonal of an approximate inverse. Journal of Computational Physics, 326(??):828–844, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304120.

Wang:2016:GPS

[WLL16] Hongqiao Wang, Guang Lin, and Jinglai Li. Gaussian process surrogates for failure detection: a Bayesian experimental design approach. Journal of Computational Physics, 313(??):247– 259, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911600125X.

Wang:2015:CHO

[WLM15] Junfeng Wang, Chunlei Liang, and Mark S. Miesch. A compressible high-order unstructured spectral difference code for stratified convection in rotating spherical shells. Journal of Computational Physics, 290(??):90–111, June 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115001175.

Wang:2018:NCC

[WLW^{+}18] Zhikai Wang, Jingye Li, Benfeng Wang, Yiran Xu, and Xiaohong Chen. A new central compact finite difference scheme with high spectral resolution for acoustic wave equation. Journal of Computational Physics, 365(??):191–206, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301906.

Wang:2017:EIT

[WLWW17] Dong Wang, Haohan Li, Xiaoyu Wei, and Xiao-Ping Wang. An efficient iterative thresholding method for image segmentation. Journal of Computational Physics, 350(??):657–667, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305910.

Winters:2018:CSP

[WMM^{+}18] Andrew R. Winters, Rodrigo C. Moura, Gianmarco Mengaldo, Gregor J. Gassner, Stefanie Walch, Joaquim Peiro, and Spencer J. Sherwin. A comparative study on polynomial dealiasing and split form discontinuous Galerkin schemes for under-resolved turbulence computations. Journal of Computational Physics, 372(??):1–21, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303942.

Wen:2018:ECS

[WMS18] C. Y. Wen, H. Saldivar Massimi, and H. Shen. Extension of CE/SE method to non-equilibrium dissociating flows. Journal of Computational Physics, 356(??):240–260, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911730880X.

Wang:2016:HMS

[WMY16] Jinghua Wang, Q. W. Ma, and S. Yan. A hybrid model for simulating rogue waves in random seas on a large temporal and spatial scale. Journal of Computational Physics, 313(??):279– 309, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116001145.

Wang:2018:FNN

[WMY18] Jinghua Wang, Qingwei Ma, and Shiqiang Yan. A fully nonlinear numerical method for modeling wave-current interactions. Journal of Computational Physics, 369(??):173–190, September 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118302900.

Wasserman:2016:PPI

[WMYG16] M. Wasserman, Y. Mor-Yossef, and J. B. Greenberg. A positivity-preserving, implicit defect-correction multigrid method for turbulent combustion. Journal of Computational Physics, 316(??):303–337, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116300377.

Watanabe:2017:GER

[WN17] Tomoaki Watanabe and Koji Nagata. Gradients estimation from random points with volumetric tensor in turbulence. Journal of Computational Physics, 350(??):518–529, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306381.

Wahlsten:2018:RBC

[WN18] Markus Wahlsten and Jan Nordström. Robust boundary conditions for stochastic incompletely parabolic systems of equations. Journal of Computational Physics, 371(??):192–213, October 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118303437.

Weston:2019:PNK

[WNDB19] Brian Weston, Robert Nourgaliev, Jean-Pierre Delplanque, and Andrew T. Barker. Preconditioning a Newton–Krylov solver for all-speed melt pool flow physics. Journal of Computational Physics, 397(??):Article 108847, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119305315.

Weeber:2019:ACD

[WNW^{+}19] Rudolf Weeber, Franziska Nestler, Florian Weik, Michael Pippig, Daniel Potts, and Christian Holm. Accelerating the calculation of dipolar interactions in particle based simulations with open boundary conditions by means of the P^{2} NFFT method. Journal of Computational Physics, 391(??):243–258, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301020.

Watvisave:2015:HMD

[WPB15] D. S. Watvisave, B. P. Puranik, and U. V. Bhandarkar. A hybrid MD-DSMC coupling method to investigate flow characteristics of micro-devices. Journal of Computational Physics, 302(??):603–617, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115006051.

Wang:2015:DGR

[WQZ15] Yibin Wang, Ning Qin, and Ning Zhao. Delaunay graph and radial basis function for fast quality mesh deformation. Journal of Computational Physics, 294(??):149–172, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002016.

Wang:2015:ARF

[WR15] Qiuju Wang and Yu-Xin Ren. An accurate and robust finite volume scheme based on the spline interpolation for solving the Euler and Navier–Stokes equations on non-uniform curvilinear grids. Journal of Computational Physics, 284(??):648–667, March 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115000066.

Winges:2016:HOB

[WR16] Johan Winges and Thomas Rylander. Higher-order bricktetrahedron hybrid method for Maxwell’s equations in time domain. Journal of Computational Physics, 321(??):698– 707, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116302248.

Wang:2016:CHOa

[WRL16a] Qian Wang, Yu-Xin Ren, and Wanai Li. Compact high order finite volume method on unstructured grids I: Basic formulations and one-dimensional schemes. Journal of Computational Physics, 314(??):863–882, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116000784.

Wang:2016:CHOb

[WRL16b] Qian Wang, Yu-Xin Ren, and Wanai Li. Compact high order finite volume method on unstructured grids II: Extension to two-dimensional Euler equations. Journal of Computational Physics, 314(??):883–908, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911600200X.

Wu:2018:CMS

[WRL18] Fuyuan Wu, Rafael Ramis, and Zhenghong Li. A conservative MHD scheme on unstructured Lagrangian grids for Z-pinch hydrodynamic simulations. Journal of Computational Physics, 357(??):206–229, March 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308987.

Winovich:2019:CUC

[WRL19] Nick Winovich, Karthik Ramani, and Guang Lin. ConvPDE-UQ: Convolutional neural networks with quantified uncertainty for heterogeneous elliptic partial differential equations on varied domains. Journal of Computational Physics, 394 (??):263–279, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303572.

Wang:2017:CHO

[WRPL17] Qian Wang, Yu-Xin Ren, Jianhua Pan, and Wanai Li. Compact high order finite volume method on unstructured grids III: Variational reconstruction. Journal of Computational Physics, 337(??):1–26, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301249.

Wiens:2015:EPI

[WS15a] Jeffrey K. Wiens and John M. Stockie. An efficient parallel immersed boundary algorithm using a pseudo-compressible fluid solver. Journal of Computational Physics, 281(??):917–941, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007463.

Woods:2015:VFD

[WS15b] C. Nathan Woods and Ryan P. Starkey. Verification of fluiddynamic codes in the presence of shocks and other discontinuities. Journal of Computational Physics, 294(??):312–328, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002168.

Weatheritt:2016:NEA

[WS16] Jack Weatheritt and Richard Sandberg. A novel evolutionary algorithm applied to algebraic modifications of the RANS stress-strain relationship. Journal of Computational Physics, 325(??):22–37, November 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303643.

Wissocq:2019:ESA

[WSB19] Gauthier Wissocq, Pierre Sagaut, and Jean-François Boussuge. An extended spectral analysis of the lattice Boltzmann method: modal interactions and stability issues. Journal of Computational Physics, 380(??):311–333, March 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118308118.

Wu:2017:USG

[WSF17] Jilian Wu, Jie Shen, and Xinlong Feng. Unconditionally stable gauge–Uzawa finite element schemes for incompressible natural convection problems with variable density. Journal of Computational Physics, 348(??):776–789, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305533.

Weyens:2017:PNC

[WSH^{+}17] T. Weyens, R. Sánchez, G. Huijsmans, A. Loarte, and L. García. PB3D: a new code for edge 3-D ideal linear peeling-ballooning stability. Journal of Computational Physics, 330(??):997–1009, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305629.

Woods:2019:NSE

[WSH19] M. Woods, W. Sailor, and M. Holmes. Numerical solution of the electron transport equation in the upper atmosphere. Journal of Computational Physics, 376(??):129–144, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306302.

Wang:2015:MLB

[WSHT15] Y. Wang, C. Shu, H. B. Huang, and C. J. Teo. Multiphase lattice Boltzmann flux solver for incompressible multiphase flows with large density ratio. Journal of Computational Physics, 280(??):404–423, January 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006718.

Wang:2016:ACA

[WSJY16] Xingyu Wang, Roman Samulyak, Xiangmin Jiao, and Kwangmin Yu. AP-Cloud: Adaptive Particle-in-Cloud method for optimal solutions to Vlasov–Poisson equation. Journal of Computational Physics, 316(??):682–699, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300833.

Wu:2019:DRB

[WSK19] Xu Wu, Koroush Shirvan, and Tomasz Kozlowski. Demonstration of the relationship between sensitivity and identifiability for inverse uncertainty quantification. Journal of Computational Physics, 396(??):12–30, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304401.

Watanabe:2015:LPM

[WSN^{+}15] Tomoaki Watanabe, Yasuhiko Sakai, Kouji Nagata, Yasumasa Ito, and Toshiyuki Hayase. LES–Lagrangian particle method for turbulent reactive flows based on the approximate deconvolution model and mixing model. Journal of Computational Physics, 294(??):127–148, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500193X.

Wang:2018:DSG

[WSN^{+}18] Guoyin Wang, Guglielmo Scovazzi, Léo Nouveau, Christopher E. Kees, Simone Rossi, Oriol Colomés, and Alex Main. Dual-scale Galerkin methods for Darcy flow. Journal of Computational Physics, 354(??):111–134, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117308148.

Wei:2016:EHO

[WSOW16] Xiao-Kun Wei, Wei Shao, Haiyan Ou, and Bing-Zhong Wang. An efficient higher-order PML in WLP-FDTD method for time reversed wave simulation. Journal of Computational Physics, 321(??):1206–1216, September 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302571.

Wang:2017:FRS

[WSP17] Yayun Wang, Adam J. Sierakowski, and Andrea Prosperetti. Fully-resolved simulation of particulate flows with particles-fluid heat transfer. Journal of Computational Physics, 350(??):638–656, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305521.

Winkel:2015:HOB

[WSR15] Mathias Winkel, Robert Speck, and Daniel Ruprecht. A high-order Boris integrator. Journal of Computational Physics, 295(??):456–474, August 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115002685.

Wang:2015:MCD

[WSS^{+}15] Y. Wang, C. Shu, J. Y. Shao, J. Wu, and X. D. Niu. A mass-conserved diffuse interface method and its application for incompressible multiphase flows with large density ratio. Journal of Computational Physics, 290(??):336–351, June 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001382.

Wingen:2015:RSX

[WSU^{+}15] A. Wingen, M. W. Shafer, E. A. Unterberg, J. C. Hill, and D. L. Hillis. Regularization of soft-X-ray imaging in the DIII-D tokamak. Journal of Computational Physics, 289(??):83–95, May 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001102.

Wang:2015:IML

[WSY15] Y. Wang, C. Shu, and L. M. Yang. An improved multiphase lattice Boltzmann flux solver for three-dimensional flows with large density ratio and high Reynolds number. Journal of Computational Physics, 302(??):41–58, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115005781.

Wang:2016:BCE

[WSY16] Y. Wang, C. Shu, and L. M. Yang. Boundary conditionenforced immersed boundary-lattice Boltzmann flux solver for thermal flows with Neumann boundary conditions. Journal of Computational Physics, 306(??):237–252, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500786X.

Wu:2015:HOA

[WT15]  Kailiang Wu and Huazhong Tang.  High-order accu-

rate physical-constraints-preserving finite difference WENO schemes for special relativistic hydrodynamics. Journal of Computational Physics, 298(??):539–564, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003988.

Wu:2016:FIA

[WT16] Yu Mao Wu and Si Jia Teng. Frequency-independent approach to calculate physical optics radiations with the quadratic concave phase variations. Journal of Computational Physics, 324(??):44–61, November 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303242.

Westerkamp:2019:FEM

[WT19] A. Westerkamp and M. Torrilhon. Finite element methods for the linear regularized 13-moment equations describing slow rarefied gas flows. Journal of Computational Physics, 389(??):1– 21, July 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302050.

Wang:2016:SGD

[WTGC16] Zixuan Wang, Qi Tang, Wei Guo, and Yingda Cheng. Sparse grid discontinuous Galerkin methods for high-dimensional elliptic equations. Journal of Computational Physics, 314(??): 244–263, June 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911600156X.

Williams:2017:EPW

[WTL17] James Williams, L. Bruno Tremblay, and Jean-François Lemieux. The effects of plastic waves on the numerical convergence of the viscous-plastic and elastic-viscous-plastic sea-ice models. Journal of Computational Physics, 340(??):519–533, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302498.

Wang:2019:SRC

[WTL19] Zekun Wang, Yujun Teng, and Moubin Liu. A semi-resolved CFD–DEM approach for particulate flows with kernel based approximation and Hilbert curve based searching strategy. Journal of Computational Physics, 384(??):151–169, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300427.

White:2017:IDM

[WTS^{+}17] T. G. White, A. Tikku, M. F. Alves Silva, G. Gregori, A. Higginbotham, and D. E. Eakins. Identifying deformation mechanisms in molecular dynamics simulations of laser shocked matter. Journal of Computational Physics, 350(??):16–24, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306174.

Wu:2017:SGM

[WTX17] Kailiang Wu, Huazhong Tang, and Dongbin Xiu. A stochastic Galerkin method for first-order quasilinear hyperbolic systems with uncertainty. Journal of Computational Physics, 345(??):224–244, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304023.

Wu:2016:SOP

[Wu16] Shu-Lin Wu. A second-order parareal algorithm for fractional PDEs. Journal of Computational Physics, 307(??):280–290, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008220.

Wu:2019:FSM

[Wu19] Lei Wu. A fast spectral method for the Uehling–Uhlenbeck equation for quantum gas mixtures: Homogeneous relaxation and transport coefficients. Journal of Computational Physics, 399(??):Article 108924, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119306291.

Wang:2019:HSM

[WVB19] Shizhao Wang, Marcos Vanella, and Elias Balaras. A hydrodynamic stress model for simulating turbulence/particle interactions with immersed boundary methods. Journal of Computational Physics, 382(??):240–263, April 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300270.

Wang:2015:AHO

[WW15] Yue Wang and Shuanghu Wang. Arbitrary high order discontinuous Galerkin schemes based on the GRP method for compressible Euler equations. Journal of Computational Physics, 298(??):113–124, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002855.

Wang:2016:OOP

[WW16] Bao Wang and Guo-Wei Wei. Object-oriented persistent homology. Journal of Computational Physics, 305(??):276–299, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007123.

Wang:2017:FCM

[WW17] Che Wang and Hong Wang. A fast collocation method for a variable-coefficient nonlocal diffusion model. Journal of Computational Physics, 330(??):114–126, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305824.

Wang:2018:FFE

[WW18] Bin Wang and Xinyuan Wu. Functionally-fitted energypreserving integrators for Poisson systems. Journal of Computational Physics, 364(??):137–152, July 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301670.

Wang:2019:VPE

[WW19] Bin Wang and Xinyuan Wu. Volume-preserving exponential integrators and their applications. Journal of Computational Physics, 396(??):867–887, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S002199911930511X.

Wintermeyer:2017:ESN

[WWGK17] Niklas Wintermeyer, Andrew R. Winters, Gregor J. Gassner, and David A. Kopriva. An entropy stable nodal discontinuous Galerkin method for the two dimensional shallow water equations on unstructured curvilinear meshes with discontinuous bathymetry. Journal of Computational Physics, 340(??):200– 242, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302310.

Wintermeyer:2018:ESD

[WWGW18] Niklas Wintermeyer, Andrew R. Winters, Gregor J. Gassner, and Timothy Warburton. An entropy stable discontinuous Galerkin method for the shallow water equations on curvilinear meshes with wet/dry fronts accelerated by GPUs. Journal of Computational Physics, 375(??):447–480, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305692.

Waluga:2016:MCC

[WWR16] Christian Waluga, Barbara Wohlmuth, and Ulrich Rüde. Mass-corrections for the conservative coupling of flow and transport on collocated meshes. Journal of Computational Physics, 305(??):319–332, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007214.

Weinmuller:2017:PAS

[WWRS17] Markus Weinmüller, Michael Weinmüller, Jonathan Rohland, and Armin Scrinzi. Perfect absorption in Schrödingerlike problems using non-equidistant complex grids. Journal of Computational Physics, 333(??):199–211, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306830.

Wang:2019:ITD

[WWX19] Dong Wang, Xiao-Ping Wang, and Xianmin Xu. An improved threshold dynamics method for wetting dynamics. Journal of Computational Physics, 392(??):291–310, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302839.

Wang:2019:DAF

[WWZ19] Mengze Wang, Qi Wang, and Tamer A. Zaki. Discrete adjoint of fractional-step incompressible Navier–Stokes solver in curvilinear coordinates and application to data assimilation. Journal of Computational Physics, 396(??):427–450, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304735.

Wu:2017:MAC

[WX17] Shuonan Wu and Jinchao Xu. Multiphase Allen–Cahn and Cahn–Hilliard models and their discretizations with the effect of pairwise surface tensions. Journal of Computational Physics, 343(??):10–32, August 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303182.

Wu:2018:SFA

[WX18] Kailiang Wu and Dongbin Xiu. Sequential function approximation on arbitrarily distributed point sets. Journal of Computational Physics, 354(??):370–386, February 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307738.

Wu:2019:NAA

[WX19] Kailiang Wu and Dongbin Xiu. Numerical aspects for approximating governing equations using data. Journal of Computational Physics, 384(??):200–221, May 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119300816.

Wei:2019:FEM

[WXSJ19] Huayi Wei, Ming Xu, Wei Si, and Kai Jiang. A finite element method of the self-consistent field theory on general curved surfaces. Journal of Computational Physics, 387(??):230–244, June 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301640.

Wang:2015:SOM

[WXW15] Bao Wang, Kelin Xia, and Guo-Wei Wei. Second order method for solving 3D elasticity equations with complex interfaces. Journal of Computational Physics, 294(??):405–438, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002144.

Wang:2016:APS

[WY16] Li Wang and Bokai Yan. An asymptotic-preserving scheme for linear kinetic equation with fractional diffusion limit. Journal of Computational Physics, 312(??):157–174, May 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911600098X.

Wan:2017:DSC

[WY17] Xiaoliang Wan and Haijun Yu. A dynamic-solver-consistent minimum action method: With an application to 2D Navier– Stokes equations. Journal of Computational Physics, 331 (??):209–226, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116306052.

Wang:2019:VLDa

[WY19] Yanxing Wang and Vigor Yang. Vaporization of liquid droplet with large deformation and high mass transfer rate, I: Constant-density, constant-property case. Journal of Computational Physics, 392(??):56–70, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301962.

Wu:2017:EMP

[WYA^{+}17a] Shaohua Wu, Edward K. Y. Yapp, Jethro Akroyd, Sebastian Mosbach, Rong Xu, Wenming Yang, and Markus Kraft. Extension of moment projection method to the fragmentation process. Journal of Computational Physics, 335(??):516–534, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730061X.

Wu:2017:MPM

[WYA^{+}17b] Shaohua Wu, Edward K. Y. Yapp, Jethro Akroyd, Sebastian Mosbach, Rong Xu, Wenming Yang, and Markus Kraft. A moment projection method for population balance dynamics with a shrinkage term. Journal of Computational Physics, 330(??):960–980, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305289.

Wang:2017:UGK

[WYLX17] Zhao Wang, Hong Yan, Qibing Li, and Kun Xu. Unified gas-kinetic scheme for diatomic molecular flow with translational, rotational, and vibrational modes. Journal of Computational Physics, 350(??):237–259, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306265.

Wang:2018:DMP

[WYZZ18] Junping Wang, Xiu Ye, Qilong Zhai, and Ran Zhang. Discrete maximum principle for the P_{1}–P_{0} weak Galerkin finite element approximations. Journal of Computational Physics, 362(??):114–130, June 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118300895.

Wang:2015:HAP

[WZ15] Hong Wang and Xuhao Zhang. A high-accuracy preserving spectral Galerkin method for the Dirichlet boundary-value problem of variable-coefficient conservative fractional diffusion equations. Journal of Computational Physics, 281(??):67–81, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007001.

Wu:2017:FPI

[WZ17] Shu-Lin Wu and Tao Zhou. Fast parareal iterations for fractional diffusion equations. Journal of Computational Physics, 329(??):210–226, January 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305502.

Wang:2018:MPM

[WZ18a] Jianqiang Wang and Xiaobing Zhang. Modified particle method with integral Navier–Stokes formulation for incompressible flows. Journal of Computational Physics, 365(??): 1–13, August 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118302092.

Wu:2018:PAL

[WZ18b] Shu-Lin Wu and Tao Zhou. Parareal algorithms with local time-integrators for time fractional differential equations. Journal of Computational Physics, 358(??):135–149, April 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117309233.

Wu:2017:FIS

[WZL^{+}17] Lei Wu, Jun Zhang, Haihu Liu, Yonghao Zhang, and Jason M. Reese. A fast iterative scheme for the linearized Boltzmann equation. Journal of Computational Physics, 338(??):431–451, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301894.

Wang:2019:PGF

[WZLS19] Liqun Wang, Hui Zheng, Xin Lu, and Liwei Shi. A Petrov–Galerkin finite element interface method for interface problems with Bloch-periodic boundary conditions and its application in phononic crystals. Journal of Computational Physics, 393(??):117–138, September 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303043.

Wu:2015:FSS

[WZR15] Lei Wu, Yonghao Zhang, and Jason M. Reese. Fast spectral solution of the generalized Enskog equation for dense gases. Journal of Computational Physics, 303(??):66–79, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006270.

Wu:2015:FSM

[WZRZ15] Lei Wu, Jun Zhang, Jason M. Reese, and Yonghao Zhang. A fast spectral method for the Boltzmann equation for monatomic gas mixtures. Journal of Computational Physics, 298(??):602–621, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004167.

Xiang:2018:MBF

[XB18] Dao P. Xiang and Matthys M. Botha. MLFMM-based, fast multiple-reflection physical optics for large-scale electromagnetic scattering analysis. Journal of Computational Physics, 368(??):69–91, September 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118302870.

Xiao:2017:WBU

[XCX17] Tianbai Xiao, Qingdong Cai, and Kun Xu. A wellbalanced unified gas-kinetic scheme for multiscale flow transport under gravitational field. Journal of Computational Physics, 332(??):475–491, March 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306751.

Xie:2019:HOM

[XDLX19] Bin Xie, Xi Deng, ShiJun Liao, and Feng Xiao. High-order multi-moment finite volume method with smoothness adaptive fitting reconstruction for compressible viscous flow. Journal of Computational Physics, 394(??):559–593, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119304036.

Xie:2017:HPD

[XDSX17] Bin Xie, Xi Deng, Ziyao Sun, and Feng Xiao. A hybrid pressure-density-based Mach uniform algorithm for 2D Euler equations on unstructured grids by using multimoment finite volume method. Journal of Computational Physics, 335(??):637–663, April 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300591.

Xiao:2017:FCP

[XDvW17] Cheng-Nian Xiao, Fabian Denner, and Berend G. M. van Wachem. Fully-coupled pressure-based finite-volume framework for the simulation of fluid flows at all speeds in complex geometries. Journal of Computational Physics, 346(??):91–130, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117304540.

Xu:2019:HRB

[XGZ19] Minqiang Xu, Hailong Guo, and Qingsong Zou. Hessian recovery based finite element methods for the Cahn–Hilliard equation. Journal of Computational Physics, 386(??):524–540, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119301159.

Xu:2015:PMT

[XHC15] Qinwu Xu, Jan S. Hesthaven, and Feng Chen. A parareal method for time-fractional differential equations. Journal of Computational Physics, 293(??):173–183, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007967.

Xia:2015:FDS

[Xia15] Yinhua Xia. A fully discrete stable discontinuous Galerkin method for the thin film epitaxy problem without slope selection. Journal of Computational Physics, 280(??):248–260, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006615.

Xie:2015:AMR

[Xie15] Wenjun Xie. An axisymmetric multiple-relaxation-time lattice Boltzmann scheme. Journal of Computational Physics, 281(??):55–66, January 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114007013.

Xie:2016:NMP

[XJ16] Dexuan Xie and Yi Jiang. A nonlocal modified Poisson– Boltzmann equation and finite element solver for computing electrostatics of biomolecules. Journal of Computational Physics, 322(??):1–20, October 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116302534.

Xi:2018:MTE

[XJG18] Yingxia Xi, Xia Ji, and Hongrui Geng. A C^{0}IP method of transmission eigenvalues for elastic waves. Journal of Computational Physics, 374(??):237–248, December 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118305163.

Xiong:2015:HOA

[XJLQ15] Tao Xiong, Juhi Jang, Fengyan Li, and Jing-Mei Qiu. High order asymptotic preserving nodal discontinuous Galerkin IMEX schemes for the BGK equation. Journal of Computational Physics, 284(??):70–94, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114008389.

Xu:2016:NCC

[XL16] Zhiliang Xu and Yingjie Liu. New central and central discontinuous Galerkin schemes on overlapping cells of unstructured grids for solving ideal magnetohydrodynamic equations with globally divergence-free magnetic field. Journal of Computational Physics, 327(??):203–224, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630465X.

Xie:2017:EMM

[XL17a] Liang Xie and Hong Liu. Efficient mesh motion using radial basis functions with volume grid points reduction algorithm.

Journal of Computational Physics, 348(??):401–415, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305508.

Xu:2017:DDO

[XL17b] Xiankun Xu and Peiwen Li. Distance descending ordering method: an O(n) algorithm for inverting the mass matrix in simulation of macromolecules with long branches. Journal of Computational Physics, 349(??):253–264, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117305776.

Xie:2017:NIG

[XLL^{+}17] Wenjia Xie, Wei Li, Hua Li, Zhengyu Tian, and Sha Pan. On numerical instabilities of Godunov-type schemes for strong shocks. Journal of Computational Physics, 350(??):607–637, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117306447.

Xia:2015:EMF

[XLY15] Jianlin Xia, Zhilin Li, and Xin Ye. Effective matrix-free preconditioning for the augmented immersed interface method. Journal of Computational Physics, 303(??):295–312, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006580.

Xuan:2018:PVE

[XM18] Li-Jun Xuan and Joseph Majdalani. A point-value enhanced finite volume method based on approximate delta functions. Journal of Computational Physics, 355(??):37–58, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308264.

Xing:2017:PNM

[XML17] F. Xing, R. Masson, and S. Lopez. Parallel numerical modeling of hybrid-dimensional compositional non-isothermal Darcy flows in fractured porous media. Journal of Computational Physics, 345(??):637–664, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304266.

Xie:2019:DGM

[XOX19] Jiangming Xie, M. Yvonne Ou, and Liwei Xu. A discontinuous Galerkin method for wave propagation in orthotropic poroelastic media with memory terms. Journal of Computational Physics, 397(??):Article 108865, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930556X.

Xu:2015:CJC

[XP15] Sheng Xu and Glen D. Pearson, Jr. Computing jump conditions for the immersed interface method using triangular meshes. Journal of Computational Physics, 302(??):59–67, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500546X.

Xiong:2017:HUH

[XQ17] Tao Xiong and Jing-Mei Qiu. A hierarchical uniformly high order DG–IMEX scheme for the 1D BGK equation. Journal of Computational Physics, 336(??):164–191, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117300487.

Xu:2017:MEL

[XR17] Yuanwei Xu and P. Mark Rodger. MBAR-enhanced lattice Monte Carlo simulation of the effect of helices on membrane protein aggregation. Journal of Computational Physics, 333(??):128–141, March 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116306696.

Xu:2015:SDS

[XRMM15] Shenren Xu, David Radford, Marcus Meyer, and Jens-Dominik Müller. Stabilisation of discrete steady adjoint solvers. Journal of Computational Physics, 299(??):175–195, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004349.

Xing:2015:CPE

[XS15] Tao Xing and Frederick Stern. Comment on “A procedure for the estimation of the numerical uncertainty of CFD calculations based on grid refinement studies” (L. Eça and M. Hoekstra, Journal of Computational Physics 262 (2014) 104–130). Journal of Computational Physics, 301(??):484–486, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005756. See [EH14]

and reply [EH15].

Xuan:2019:CSS

[XS19] Anqing Xuan and Lian Shen. A conservative scheme for simulation of free-surface turbulent and wave flows. Journal of Computational Physics, 378(??):18–43, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307137.

Xu:2018:LSM

[XSL18] Jian-Jun Xu, Weidong Shi, and Ming-Chih Lai. A level-set method for two-phase flows with soluble surfactant. Journal of Computational Physics, 353(??):336–355, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307726.

Xing:2016:MLE

[XTS^{+}16] W. W. Xing, V. Triantafyllidis, A. A. Shah, P. B. Nair, and N. Zabaras. Manifold learning for the emulation of spatial fields from computational models. Journal of Computational Physics, 326(??):666–690, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303722.

Xu:2018:NGA

[XTYL18] Lincheng Xu, Fang-Bao Tian, John Young, and Joseph C. S. Lai. A novel geometry-adaptive Cartesian grid based immersed boundary-lattice Boltzmann method for fluid-structure interactions at moderate and high Reynolds numbers. Journal of Computational Physics, 375(??):22–56, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305473.

Xiao:2015:RVM

[XWB15] Jianping Xiao, Lei Wang, and John P. Boyd. RBF-vortex methods for the barotropic vorticity equation on a sphere. Journal of Computational Physics, 285(??):208–225, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115000078.

Xia:2016:AHF

[XWL^{+}16] Yidong Xia, Chuanjin Wang, Hong Luo, Mark Christon, and Jozsef Bakosi. Assessment of a hybrid finite element and finite volume code for turbulent incompressible flows. Journal of Computational Physics, 307(??):653–669, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008414.

Xiao:2016:QRM

[XWW^{+}16] H. Xiao, J.-L. Wu, J.-X. Wang, R. Sun, and C. J. Roy. Quantifying and reducing model-form uncertainties in Reynoldsaveraged Navier–Stokes simulations: a data-driven, physicsinformed Bayesian approach. Journal of Computational Physics, 324(??):115–136, November 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116303394.

Xu:2017:ETD

[XWW17] Xianmin Xu, Dong Wang, and Xiao-Ping Wang. An efficient threshold dynamics method for wetting on rough surfaces. Journal of Computational Physics, 330(??):510–528, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305939.

Xu:2018:PHD

[XWZ^{+}18] Xiao Xu, Xuede Wang, Mu Zhang, Jun Zhang, and Junjie Tan. A parallelized hybrid N-SDS/MC-IP approach based on adaptive structured/unstructured overlapping grids for hypersonic transitional flows. Journal of Computational Physics, 371(??):409–433, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303218.

Xie:2016:MMC

[XX16] Bin Xie and Feng Xiao. A multi-moment constrained finite volume method on arbitrary unstructured grids for incompressible flows. Journal of Computational Physics, 327(??):747–778, December 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304752.

Xie:2017:TEA

[XX17] Bin Xie and Feng Xiao. Toward efficient and accurate interface capturing on arbitrary hybrid unstructured grids: the THINC method with quadratic surface representation and Gaussian quadrature. Journal of Computational Physics, 349(??):415–440, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305995.

Xie:2018:SRS

[XXR18] Qing Xie, Zhixiang Xiao, and Zhuyin Ren. A spectral radius scaling semi-implicit iterative time stepping method for reactive flow simulations with detailed chemistry. Journal of Computational Physics, 368(??):47–68, September 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118302675.

Xiong:2017:BCM

[XY17] Shiying Xiong and Yue Yang. The boundary-constraint method for constructing vortex-surface fields. Journal of Computational Physics, 339(??):31–45, June 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302000.

Xing:2018:HON

[XY18] Yanyuan Xing and Yubin Yan. A higher order numerical method for time fractional partial differential equations with nonsmooth data. Journal of Computational Physics, 357(??):305–323, March 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117309294.

Xiao:2017:NIR

[XYF^{+}17] D. Xiao, P. Yang, F. Fang, J. Xiang, C. C. Pain, I. M. Navon, and M. Chen. A non-intrusive reduced-order model for compressible fluid and fractured solid coupling and its application to blasting. Journal of Computational Physics, 330(??):221–244, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305769.

Xu:2019:HOB

[XYG19] Ziyao Xu, Yang Yang, and Hui Guo. High-order boundpreserving discontinuous Galerkin methods for wormhole propagation on triangular meshes. Journal of Computational Physics, 390(??):323–341, August 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302293.

Xiao:2016:FSI

[XYPT16] Longfei Xiao, Jianmin Yang, Tao Peng, and Longbin Tao. A free surface interpolation approach for rapid simulation of short waves in meshless numerical wave tank based on the radial basis function. Journal of Computational Physics, 307(??):203–224, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008189.

Xie:2015:MFE

[XZ15] Hehu Xie and Tao Zhou. A multilevel finite element method for Fredholm integral eigenvalue problems. Journal of Computational Physics, 303(??):173–184, December 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006427.

Xu:2019:PIS

[XZJK19] Yong Xu, Wanrong Zan, Wantao Jia, and Jürgen Kurths. Path integral solutions of the governing equation of SDEs excited by Lévy white noise. Journal of Computational Physics, 394(??):41–55, October 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119303547.

Xue:2018:TFS

[XZT18] Tao Xue, Xiaobing Zhang, and Kumar K. Tamma. A twofield state-based peridynamic theory for thermal contact problems. Journal of Computational Physics, 374(??):1180–1195, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305370.

Xu:2015:APB

[XZZ15] Zuwei Xu, Haibo Zhao, and Chuguang Zheng. Accelerating population balance-Monte Carlo simulation for coagulation dynamics from the Markov jump model, stochastic algorithm and GPU parallel computing. Journal of Computational Physics, 281(??):844–863, January 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007438.

Yaman:2019:RGI

[Yam19] Olha Ivanyshyn Yaman. Reconstruction of generalized impedance functions for 3D acoustic scattering. Journal of Computational Physics, 392(??):444–455, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303134.

Yan:2016:HMD

[Yan16a] Bokai Yan. A hybrid method with deviational particles for spatial inhomogeneous plasma. Journal of Computational Physics, 309(??):18–36, March 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008694.

Yang:2016:LFS

[Yan16b] Xiaofeng Yang. Linear, first and second-order, unconditionally energy stable numerical schemes for the phase field model of homopolymer blends. Journal of Computational Physics, 327(??):294–316, December 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304508.

Yano:2017:FAC

[Yan17] Ryosuke Yano. Fast and accurate calculation of dilute quantum gas using Uehling–Uhlenbeck model equation. Journal of Computational Physics, 330(??):1010–1021, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305794.

Yang:2019:UFO

[Yan19] Haizhao Yang. A unified framework for oscillatory integral transforms: When to use NUFFT or butterfly factorization? Journal of Computational Physics, 388(??):103–122, July 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119301615.

Yasuda:2017:MCS

[Yas17] Shugo Yasuda. Monte Carlo simulation for kinetic chemotaxis model: an application to the traveling population wave. Journal of Computational Physics, 330(??):1022–1042, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305745.

Yousefzadeh:2017:PBH

[YB17] Mehrdad Yousefzadeh and Ilenia Battiato. Physics-based hybrid method for multiscale transport in porous media. Journal of Computational Physics, 344(??):320–338, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303340.

Yang:2019:PIC

[YBSTT19] Xiu Yang, David Barajas-Solano, Guzel Tartakovsky, and Alexandre M. Tartakovsky. Physics-informed CoKriging: a Gaussian-process-regression-based multifidelity method for data-model convergence. Journal of Computational Physics, 395(??):410–431, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119304498.

Yan:2015:MCM

[YC15] Bokai Yan and Russel E. Caflisch. A Monte Carlo method with negative particles for Coulomb collisions. Journal of Computational Physics, 298(??):711–740, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004180.

Younsi:2016:AFC

[YC16] Amina Younsi and Alain Cartalade. On anisotropy function in crystal growth simulations using Lattice Boltzmann equation. Journal of Computational Physics, 325(??):1–21, November 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116303631.

Yamaleev:2017:FFO

[YC17] Nail K. Yamaleev and Mark H. Carpenter. A family of fourth-order entropy stable nonoscillatory spectral collocation schemes for the 1-D Navier–Stokes equations. Journal of Computational Physics, 331(??):90–107, February 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306313.

Yan:2015:ASM

[YCBC15] Bokai Yan, Russel E. Caflisch, Farzin Barekat, and Jean-Luc Cambier. Analysis and simulation for a model of electron impact excitation/deexcitation and ionization/recombination. Journal of Computational Physics, 299(??):747–786, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004726.

Yang:2015:HOC

[YCPD15] H. Q. Yang, Z. J. Chen, Andrzej Przekwas, and Jonathan Dudley. A high-order CFD method using successive differentiation. Journal of Computational Physics, 281(??):690–707, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114007281.

Yuan:2017:FEB

[YCS^{+}17] H. Z. Yuan, Z. Chen, C. Shu, Y. Wang, X. D. Niu, and S. Shu. A free energy-based surface tension force model for simulation of multiphase flows by level-set method. Journal of Computational Physics, 345(??):404–426, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303959.

Yang:2018:MFI

[YD18] Zhiguo Yang and Suchuan Dong. Multiphase flows of N immiscible incompressible fluids: An outflow/open boundary condition and algorithm. Journal of Computational Physics, 365(??):33–70, August 1, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911830216X.

Yang:2019:UES

[YD19] Zhiguo Yang and Suchuan Dong. An unconditionally energystable scheme based on an implicit auxiliary energy variable for incompressible two-phase flows with different densities involving only precomputable coefficient matrices. Journal of Computational Physics, 393(??):229–257, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303493.

Yazdani:2016:FCD

[YDCK16] Alireza Yazdani, Mingge Deng, Bruce Caswell, and George Em Karniadakis. Flow in complex domains simulated by Dissipative Particle Dynamics driven by geometry-specific bodyforces. Journal of Computational Physics, 305(??):906–920, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007342.

Yamaleev:2019:ESS

[YDLC19] Nail K. Yamaleev, David C. Del Rey Fernández, Jialin Lou, and Mark H. Carpenter. Entropy stable spectral collocation schemes for the 3-D Navier–Stokes equations on dynamic unstructured grids. Journal of Computational Physics, 399(??):Article 108897, December 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305959.

Yang:2019:NRA

[YDN19] Chang Yang, Fabrice Deluzet, and Jacek Narski. On the numerical resolution of anisotropic equations with high order differential operators arising in plasma physics. Journal of Computational Physics, 386(??):502–523, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301081.

Yeo:2019:DDR

[Yeo19] Kyongmin Yeo. Data-driven reconstruction of nonlinear dynamics from sparse observation. Journal of Computational Physics, 395(??):671–689, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304474.

Yang:2019:TGP

[YFC19] Yanfang Yang, Shubin Fu, and Eric T. Chung. A two-grid preconditioner with an adaptive coarse space for flow simulations in highly heterogeneous media. Journal of Computational Physics, 391(??):1–13, August 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302219.

Yang:2017:HOC

[YFJ17] Yun Yang, Xue-Shang Feng, and Chao-Wei Jiang. A highorder CESE scheme with a new divergence-free method for MHD numerical simulation. Journal of Computational Physics, 349(??):561–581, November 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305909.

Yang:2018:UCS

[YFJ18] Yun Yang, Xue-Shang Feng, and Chao-Wei Jiang. An upwind CESE scheme for 2D and 3D MHD numerical simulation in general curvilinear coordinates. Journal of Computational Physics, 371(??):850–869, October 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118303061.

Yagub:2015:LBM

[YFKS15] A. Yagub, H. Farhat, S. Kondaraju, and T. Singh. A lattice Boltzmann model for substrates with regularly structured surface roughness. Journal of Computational Physics, 301(??):402–414, November 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005689.

Yaghmaie:2018:CMF

[YG18] Reza Yaghmaie and Somnath Ghosh. Computational modeling of finite deformation piezoelectric material behavior coupling transient electrical and mechanical fields. Journal of Computational Physics, 373(??):148–170, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304480.

Yang:2019:SDP

[YG19] Jun Yang and Nikolaos A. Gatsonis. A Smooth Dissipative Particle Dynamics method for nonisothermal liquid and gas flows in bounded domains. Journal of Computational Physics, 394(??):679–699, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302864.

Yang:2017:BVB

[YGEM17] Keren Yang, Nilabja Guha, Yalchin Efendiev, and Bani K. Mallick. Bayesian and variational Bayesian approaches for flows in heterogeneous random media. Journal of Computational Physics, 345(??):275–293, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303054.

Yang:2018:FSC

[YGJ18] Huanhuan Yang, Max Gunzburger, and Lili Ju. Fast spherical centroidal Voronoi mesh generation: A Lloydpreconditioned LBFGS method in parallel. Journal of Computational Physics, 367(??):235–252, August 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302596.

Yang:2017:LFS

[YH17] Xiaofeng Yang and Daozhi Han. Linearly first- and second-order, unconditionally energy stable schemes for the phase field crystal model. Journal of Computational Physics, 330(??):1116–1134, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116305186.

Yang:2017:TMM

[YHKPF17] Jie Yang, Heng Hu, Yao Koutsawa, and Michel Potier-Ferry. Taylor meshless method for solving non-linear partial differential equations. Journal of Computational Physics, 348(??):385–400, November 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305429.

Yang:2015:MMF

[YHQ15] Xiaobo Yang, Weizhang Huang, and Jianxian Qiu. A moving mesh finite difference method for equilibrium radiation diffusion equations. Journal of Computational Physics, 298(??):661–677, October 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115004003.

Yi:2018:TIU

[Yi18] Tae-Hyeong Yi. Time integration of unsteady nonhydrostatic equations with dual time stepping and multigrid methods. Journal of Computational Physics, 374(??):873–892, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305254.

Yan:2017:CPD

[YJ17] Su Yan and Jian-Ming Jin. A continuity-preserving and divergence-cleaning algorithm based on purely and damped hyperbolic Maxwell equations in inhomogeneous media. Journal of Computational Physics, 334(??):392–418, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300220.

Yarmohammadi:2018:SIM

[YJB18] M. Yarmohammadi, S. Javadi, and E. Babolian. Spectral iterative method and convergence analysis for solving nonlinear fractional differential equation. Journal of Computational Physics, 359(??):436–450, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300305.

Yu:2019:CSP

[YJM19] Hans Yu, Matthew P. Juniper, and Luca Magri. Combined state and parameter estimation in level-set methods. Journal of Computational Physics, 399(??):Article 108950, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119306552.

Yan:2015:SRS

[YK15] Yan Yan and David E. Keyes. Smooth and robust solutions for Dirichlet boundary control of fluid-solid conjugate heat transfer problems. Journal of Computational Physics, 281(??):759–786, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114007372.

You:2018:HOM

[YK18] Hojun You and Chongam Kim. High-order multi-dimensional limiting strategy with subcell resolution I. Two-dimensional mixed meshes. Journal of Computational Physics, 375(??): 1005–1032, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118306053.

You:2019:DRM

[YK19] Hojun You and Chongam Kim. Direct reconstruction method for discontinuous Galerkin methods on higher-order mixedcurved meshes I. Volume integration. Journal of Computational Physics, 395(??):223–246, October 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S002199911930422X.

Yildirim:2019:JFA

[YKMM19] Anil Yildirim, Gaetan K. W. Kenway, Charles A. Mader, and Joaquim R. R. A. Martins. A Jacobian-free approximate Newton–Krylov startup strategy for RANS simulations. Journal of Computational Physics, 397(??):Article 108741, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119304255.

Yao:2016:NMA

[YL16] Weigang Yao and Meng-Sing Liou. A nonlinear modeling approach using weighted piecewise series and its applications to predict unsteady flows. Journal of Computational Physics, 318(??):58–84, August 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301073.

Yuan:2017:RIM

[YL17] Lijun Yuan and Ya Yan Lu. Robust iterative method for nonlinear Helmholtz equation. Journal of Computational Physics, 343(??):1–9, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911730325X.

Yu:2019:TOM

[YL19] Hui Yu and Hailiang Liu. Third order maximum-principle-satisfying DG schemes for convection-diffusion problems with anisotropic diffusivity. Journal of Computational Physics, 391(??):14–36, August 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119302748.

Ye:2015:CDS

[YLA15] H. Ye, F. Liu, and V. Anh. Compact difference scheme for distributed-order time-fractional diffusion-wave equation on bounded domains. Journal of Computational Physics, 298(??):652–660, October 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115004222.

Yang:2016:ESH

[YLBL16] Xiu Yang, Huan Lei, Nathan A. Baker, and Guang Lin. Enhancing sparsity of Hermite polynomial expansions by iterative rotations. Journal of Computational Physics, 307(??):94–109, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007780.

Yang:2019:FSO

[YLD19] Zhiguo Yang, Lianlei Lin, and Suchuan Dong. A family of second-order energy-stable schemes for Cahn–Hilliard type equations. Journal of Computational Physics, 383(??):24–54, April 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300373.

Yong:2019:MFF

[YLH^{+}19] Peng Yong, Wenyuan Liao, Jianping Huang, Zhenchun Li, and Yaoting Lin. Misfit function for full waveform inversion based on the Wasserstein metric with dynamic formulation. Journal of Computational Physics, 399(??):Article 108911, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119306163.

Yin:2019:FAB

[YLLH19] Baoli Yin, Yang Liu, Hong Li, and Siriguleng He. Fast algorithm based on TT-M FE system for space fractional Allen– Cahn equations with smooth and non-smooth solutions. Journal of Computational Physics, 379(??):351–372, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307927.

Yan:2019:ULP

[YLZ^{+}19] Jiale Yan, Shaofan Li, A-Man Zhang, Xingyu Kan, and Peng-Nan Sun. Updated Lagrangian Particle Hydrodynamics (ULPH) modeling and simulation of multiphase flows. Journal of Computational Physics, 393(??):406–437, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303481.

Yoon:2015:AFD

[YM15] Gangjoon Yoon and Chohong Min. Analyses on the finite difference method by Gibou et al. for Poisson equation. Journal of Computational Physics, 280(??):184–194, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006457.

Yang:2017:ERJ

[YM17a] Xiang Yang and Rajat Mittal. Efficient relaxed-Jacobi smoothers for multigrid on parallel computers. Journal of Computational Physics, 332(??):135–142, March 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306519.

Yao:2017:NMO

[YM17b] Lingxing Yao and Yoichiro Mori. A numerical method for osmotic water flow and solute diffusion with deformable membrane boundaries in two spatial dimension. Journal of Computational Physics, 350(??):728–746, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306563.

Yoon:2017:CER

[YM17c] Gangjoon Yoon and Chohong Min. Comparison of eigenvalue ratios in artificial boundary perturbation and Jacobi preconditioning for solving Poisson equation. Journal of Computational Physics, 349(??):1–10, November 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117305843.

Yeo:2019:DLA

[YM19] Kyongmin Yeo and Igor Melnyk. Deep learning algorithm for data-driven simulation of noisy dynamical system. Journal of Computational Physics, 376(??):1212–1231, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118306910.

Yang:2017:SDA

[YNW17] Lun Yang, Akil Narayan, and Peng Wang. Sequential data assimilation with multiple nonlinear models and applications to subsurface flow. Journal of Computational Physics, 346(??):356–368, October 1, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117304710.

Yi:2017:VDF

[YP17] Tae-Hyeong Yi and Ja-Rin Park. Vertical discretization with finite elements for a global hydrostatic model on the cubed sphere. Journal of Computational Physics, 338(??):339–356, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301705.

Yang:2019:AUQ

[YP19] Yibo Yang and Paris Perdikaris. Adversarial uncertainty quantification in physics-informed neural networks. Journal of Computational Physics, 394(??):136–152, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303584.

Yang:2019:CQN

[YPC19] Shuo Yang, Samuel F. Potter, and Maria K. Cameron. Computing the quasipotential for nongradient SDEs in 3D. Journal of Computational Physics, 379(??):325–350, February 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118307939.

Yu:2016:FMV

[YPK16] Yue Yu, Paris Perdikaris, and George Em Karniadakis. Fractional modeling of viscoelasticity in 3D cerebral arteries and aneurysms. Journal of Computational Physics, 323(??):219– 242, October 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116302637.

Yang:2019:DAM

[YQNW19] Lun Yang, Yi Qin, Akil Narayan, and Peng Wang. Data assimilation for models with parametric uncertainty. Journal of Computational Physics, 396(??):785–798, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305054.

Yao:2015:NIT

[YR15] Wenqi Yao and Weiqing Ren. Noise-induced transition in barotropic flow over topography and application to kuroshio. Journal of Computational Physics, 300(??):352–364, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005124.

Yang:2015:NID

[YS15]  Jianming Yang and Frederick Stern.  A non-iterative

direct forcing immersed boundary method for stronglycoupled fluid-solid interactions. Journal of Computational Physics, 295(??):779–804, August 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115002971.

Yang:2017:HSM

[YS17] Jianming Yang and Frederick Stern. A highly scalable massively parallel fast marching method for the eikonal equation. Journal of Computational Physics, 332(??):333–362, March 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306659.

Yan:2018:FIP

[YS18a] Wen Yan and Michael Shelley. Flexibly imposing periodicity in kernel independent FMM: a multipole-to-local operator approach. Journal of Computational Physics, 355(??):214–232, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308501.

Yan:2018:UIS

[YS18b] Wen Yan and Michael Shelley. Universal image systems for non-periodic and periodic Stokes flows above a no-slip wall. Journal of Computational Physics, 375(??):263–270, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305722.

Yamazaki:2017:VSM

[YSC^{+}17] Hiroe Yamazaki, Jemma Shipton, Michael J. P. Cullen, Lawrence Mitchell, and Colin J. Cotter. Vertical slice modelling of nonlinear Eady waves using a compatible finite element method. Journal of Computational Physics, 343 (??):130–149, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302772.

Yang:2019:FIC

[YSLY19] Haijian Yang, Shuyu Sun, Yiteng Li, and Chao Yang. A fully implicit constraint-preserving simulator for the black oil model of petroleum reservoirs. Journal of Computational Physics, 396(??):347–363, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119303821.

Yang:2015:TDE

[YSW15] L. M. Yang, C. Shu, and J. Wu. A three-dimensional explicit sphere function-based gas-kinetic flux solver for simulation of inviscid compressible flows. Journal of Computational Physics, 295(??):322–339, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002193.

Yang:2016:DDG

[YSWS16] L. M. Yang, C. Shu, Y. Wang, and Y. Sun. Development of discrete gas kinetic scheme for simulation of 3D viscous incompressible and compressible flows. Journal of Computational Physics, 319(??):129–144, August 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301553.

Yang:2016:NSF

[YSWW16] L. M. Yang, C. Shu, J. Wu, and Y. Wang. Numerical simulation of flows from free molecular regime to continuum regime by a DVM with streaming and collision processes. Journal of Computational Physics, 306(??):291–310, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007834.

Yang:2017:NPS

[YSY17] Haijian Yang, Shuyu Sun, and Chao Yang. Nonlinearly preconditioned semismooth Newton methods for variational inequality solution of two-phase flow in porous media. Journal of Computational Physics, 332(??):1–20, March 1, 2017.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116306283.

Yang:2019:ITD

[YSYW19] L. M. Yang, C. Shu, W. M. Yang, and J. Wu. An improved three-dimensional implicit discrete velocity method on unstructured meshes for all Knudsen number flows. Journal of Computational Physics, 396(??):738–760, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304863.

Yang:2017:MMM

[YT17] Qingcheng Yang and Albert C. To. Multiresolution molecular mechanics: Surface effects in nanoscale materials. Journal of Computational Physics, 336(??):212–234, May 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117300748.

Yu:2019:HOC

[YT19] P. X. Yu and Z. F. Tian. A high-order compact scheme for the pure streamfunction (vector potential) formulation of the 3D steady incompressible Navier–Stokes equations. Journal of Computational Physics, 382(??):65–85, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300087.

Young:2015:NVP

[YTW15] D. L. Young, C. H. Tsai, and C. S. Wu. A novel vector potential formulation of 3D Navier–Stokes equations with through-flow boundaries by a local meshless method. Journal of Computational Physics, 300(??):219–240, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004854.

Yu:2015:ODR

[YWHP15] C. H. Yu, D. Wang, Z. He, and T. Pähtz. An optimized dispersion-relation-preserving combined compact difference scheme to solve advection equations. Journal of Computational Physics, 300(??):92–115, November 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005045.

Yang:2016:HSC

[YWS^{+}16] Yan Yang, Minping Wan, Yipeng Shi, Kun Yang, and Shiyi Chen. A hybrid scheme for compressible magnetohydrodynamic turbulence. Journal of Computational Physics, 306(??):73–91, February 1, 2016. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115007652.

Ying:2015:NFE

[YX15] Jinyong Ying and Dexuan Xie. A new finite element and finite difference hybrid method for computing electrostatics of ionic solvated biomolecule. Journal of Computational Physics, 298(??):636–651, October 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115004027.

Yu:2016:ELB

[YXD^{+}16] Peicheng Yu, Xinlu Xu, Asher Davidson, Adam Tableman, Thamine Dalichaouch, Fei Li, Michael D. Meyers, Weiming An, Frank S. Tsung, Viktor K. Decyk, Frederico Fiuza, Jorge Vieira, Ricardo A. Fonseca, Wei Lu, Luis O. Silva, and Warren B. Mori. Enabling Lorentz boosted frame particle-in-cell simulations of laser wakefield acceleration in quasi-3D geometry. Journal of Computational Physics, 316(??):747–759, July 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116300468.

Yang:2016:MFS

[YXF^{+}16] P. Yang, J. Xiang, F. Fang, D. Pavlidis, J.-P. Latham, and C. C. Pain. Modelling of fluid-structure interaction with multiphase viscous flows using an immersed-body method. Journal of Computational Physics, 321(??):571–592, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301802.

Yacoubi:2019:NMC

[YXW19] Acmae El. Yacoubi, Sheng Xu, and Z. Jane Wang. A new method for computing particle collisions in Navier– Stokes flows. Journal of Computational Physics, 399 (??):Article 108919, December 15, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119306242.

Ye:2016:GCC

[YXX^{+}16] Lei Ye, Yingfeng Xu, Xiaotao Xiao, Zongliang Dai, and Shaojie Wang. A gyrokinetic continuum code based on the numerical Lie transform (NLT) method. Journal of Computational Physics, 316(??):180–192, July 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300699.

Yang:2016:SNE

[YY16] Hyoseon Yang and Jungho Yoon. A short note on the error estimates of Yuan–Shu discontinuous Galerkin method based on non-polynomial approximation spaces. Journal of Computational Physics, 320(??):33–39, September 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116301772.

Yu:2017:NAP

[YY17] Haijun Yu and Xiaofeng Yang. Numerical approximations for a phase-field moving contact line model with variable densities and viscosities. Journal of Computational Physics, 334(??):665–686, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117300426.

Yu:2019:ANI

[YYJ^{+}19] Jian Yu, Chao Yan, Zhenhua Jiang, Wu Yuan, and Shusheng Chen. Adaptive non-intrusive reduced order modeling for compressible flows. Journal of Computational Physics, 397(??):Article 108855, ???? 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911930539X.

Yan:2016:OSG

[YYL16] Hongyong Yan, Lei Yang, and Xiang-Yang Li. Optimal staggered-grid finite-difference schemes by combining Taylor-series expansion and sampling approximation for wave equation modeling. Journal of Computational Physics, 326(??):913–930, December 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116304338.

Yang:2018:DSR

[YYL18] Fenglian Yang, Liang Yan, and Leevan Ling. Doubly stochastic radial basis function methods. Journal of Computational Physics, 363(??):87–97, June 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301256.

Yang:2017:FEM

[YYN^{+}17] Z. Yang, Z. Yuan, Y. Nie, J. Wang, X. Zhu, and F. Liu. Finite element method for nonlinear Riesz space fractional diffusion equations on irregular domains. Journal of Computational Physics, 330(??):863–883, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116305617.

Yaji:2016:TOT

[YYY^{+}16] Kentaro Yaji, Takayuki Yamada, Masato Yoshino, Toshiro Matsumoto, Kazuhiro Izui, and Shinji Nishiwaki. Topology optimization in thermal-fluid flow using the lattice Boltzmann method. Journal of Computational Physics, 307(??):355–377, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008244.

Yan:2019:AMF

[YZ19] Liang Yan and Tao Zhou. Adaptive multi-fidelity polynomial chaos approach to Bayesian inference in inverse problems. Journal of Computational Physics, 381(??):110–128, March 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300063.

Yang:2018:BST

[YZT^{+}18] Jianbin Yang, Guanhua Zhu, Dudu Tong, Lanyuan Lu, and Zuowei Shen. B-spline tight frame based force matching method. Journal of Computational Physics, 362(??):208–219, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118301062.

Yang:2017:NAM

[YZW17] Xiaofeng Yang, Jia Zhao, and Qi Wang. Numerical approximations for the molecular beam epitaxial growth model based on the invariant energy quadratization method. Journal of Computational Physics, 333(??):104–127, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116306799.

Yang:2018:IMF

[YZW^{+}18] Lin Yang, Feng Zhang, Cai-Zhuang Wang, Kai-Ming Ho, and Alex Travesset. Implementation of metal-friendly EAM/FStype semi-empirical potentials in HOOMD-blue: A GPUaccelerated molecular dynamics software. Journal of Computational Physics, 359(??):352–360, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300251.

Yuan:2015:IBM

[YZZ15] Ruifeng Yuan, Chengwen Zhong, and He Zhang. An immersed-boundary method based on the gas kinetic BGK scheme for incompressible viscous flow. Journal of Computational Physics, 296(??):184–208, September 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115003095.

Yan:2019:MCM

[YZZ19] Zhi-Zhong Yan, Cheng-Feng Zheng, and Chuanzeng Zhang. The Monte Carlo Markov chain method for solving the modified anomalous fractional sub-diffusion equation. Journal of Computational Physics, 394(??):477–490, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911930419X.

Zerroukat:2015:MBS

[ZA15a] M. Zerroukat and T. Allen. A moist Boussinesq shallow water equations set for testing atmospheric models. Journal of Computational Physics, 290(??):55–72, June 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500073X.

Zerroukat:2015:MCT

[ZA15b] M. Zerroukat and T. Allen. On the monotonic and conservative transport on overset/Yin–Yang grids. Journal of Computational Physics, 302(??):285–299, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005860.

Zadeh:2011:MCS

[Zad11] Kouroush Sadegh Zadeh. A mass-conservative switching algorithm for modeling fluid flow in variably saturated porous media. Journal of Computational Physics, 230(3):664–679, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005565. See com-

ments [HSK^{+}15].

Zabelok:2015:AKF

[ZAK15] Sergey Zabelok, Robert Arslanbekov, and Vladimir Kolobov. Adaptive kinetic-fluid solvers for heterogeneous computing architectures. Journal of Computational Physics, 303(??): 455–469, December 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115006658.

Zauner:2016:AFF

[Zau16] Thomas Zauner. Application of a force field algorithm for creating strongly correlated multiscale sphere packings. Journal of Computational Physics, 313(??):662–673, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001029.

Zhang:2015:NSP

[ZB15] P. Zhang and A. Benard. Numerical simulation of particle motion using a combined MacCormack and immersed boundary method. Journal of Computational Physics, 294(??):524–546, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115001680.

Zago:2018:SIS

[ZBH^{+}18] Vito Zago, Giuseppe Bilotta, Alexis Hérault, Robert A. Dalrymple, Luigi Fortuna, Annalisa Cappello, Gaetana Ganci, and Ciro Del Negro. Semi-implicit 3D SPH on GPU for lava flows. Journal of Computational Physics, 375(??):854–870, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911830593X.

Zheng:2018:FVA

[ZBZ^{+}18] Chang-Jun Zheng, Chuan-Xing Bi, Chuanzeng Zhang, Hai-Feng Gao, and Hai-Bo Chen. Free vibration analysis of elastic structures submerged in an infinite or semi-infinite fluid domain by means of a coupled FE–BE solver. Journal of Computational Physics, 359(??):183–198, April 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118300287.

Zhao:2017:GFE

[ZBZT17] Yue Zhao, Weiping Bu, Xuan Zhao, and Yifa Tang. Galerkin finite element method for two-dimensional space and time fractional Bloch–Torrey equation. Journal of Computational Physics, 350(??):117–135, December 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117306320.

Zhang:2015:VIR

[ZC15] Jianping Zhang and Ke Chen. Variational image registration by a total fractional-order variation model. Journal of Computational Physics, 293(??):442–461, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000911.

Zhang:2018:AMS

[ZC18] Zhengfang Zhang and Weifeng Chen. An approach for maximizing the smallest eigenfrequency of structure vibration based on piecewise constant level set method. Journal of Computational Physics, 361(??):377–390, May 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300664.

Zhang:2019:ATO

[ZC19a] Zhengfang Zhang and Weifeng Chen. An approach for topology optimization of damping layer under harmonic excitations based on piecewise constant level set method. Journal of Computational Physics, 390(??):470–489, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302554.

Zheng:2019:MSP

[ZC19b] B. X. Zheng and Z. Chen. A multiphase smoothed particle hydrodynamics model with lower numerical diffusion. Journal of Computational Physics, 382(??):177–201, April 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911930035X.

Zakari:2015:AUF

[ZCHS15] M. Zakari, H. Caquineau, P. Hotmar, and P. Ségur. An axisymmetric unstructured finite volume method applied to the numerical modeling of an atmospheric pressure gas discharge. Journal of Computational Physics, 281(??):473–492, January 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400713X.

Zheng:2017:NAT

[ZCL17] Youqi Zheng, Sooyoung Choi, and Deokjung Lee. A new approach to three-dimensional neutron transport solution based on the method of characteristics and linear axial approximation. Journal of Computational Physics, 350(??):25–44, December 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117305971.

Zhang:2019:DDG

[ZCL19] Fan Zhang, Jian Cheng, and Tiegang Liu. A direct discontinuous Galerkin method for the incompressible Navier– Stokes equations on arbitrary grids. Journal of Computational Physics, 380(??):269–294, March 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118307770.

Zhang:2019:HOP

[ZCQ19] Min Zhang, Juan Cheng, and Jianxian Qiu. High order positivity-preserving discontinuous Galerkin schemes for radiative transfer equations on triangular meshes. Journal of Computational Physics, 397(??):Article 108811, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119304954. See corrigendum [ZCQ20].

Zhang:2020:CHO

[ZCQ20] Min Zhang, Juan Cheng, and Jianxian Qiu. Corrigendum to “High order positivity-preserving discontinuous Galerkin schemes for radiative transfer equations on triangular meshes” [j. comput. phys. 397 (2019) 108811]. Journal of Computational Physics, 406(??):Article 109250, April 1, 2020. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999120300243. See [ZCQ19].

Zhang:2019:KTB

[ZCSZ19] L. Q. Zhang, Z. Chen, C. Shu, and M. Q. Zhang. A kinetic theory-based axisymmetric lattice Boltzmann flux solver for isothermal and thermal swirling flows. Journal of Computational Physics, 392(??):141–160, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302943.

Zhang:2019:LDA

[ZCY^{+}19] Yu Zhang, Anirban Chandra, Fan Yang, Ehsan Shams, Onkar Sahni, Mark Shephard, and Assad A. Oberai. A locally discontinuous ALE finite element formulation for compressible phase change problems. Journal of Computational Physics, 393(??):438–464, September 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302852.

Zhou:2019:PWM

[ZCZ19] Yuzhi Zhou, Huajie Chen, and Aihui Zhou. Plane wave methods for quantum eigenvalue problems of incommensurate systems. Journal of Computational Physics, 384(??):99–113, May 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300774.

Zhang:2015:TFQ

[ZD15a] Jianhua Zhang and Hua Dai. A transpose-free quasiminimal residual variant of the CORS method for solving non-Hermitian linear systems. Journal of Computational Physics, 291(??):20–33, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115001497.

Zhang:2015:SIM

[ZD15b] Mingyu Zhang and Xiao-Long Deng. A sharp interface method for SPH. Journal of Computational Physics, 302(??):469–484, December 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006087.

Zhang:2017:SWS

[ZD17] Duan Z. Zhang and Tilak R. Dhakal. Shock waves simulated using the dual domain material point method combined with molecular dynamics. Journal of Computational Physics, 334(??):240–254, April 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911730013X.

Zhang:2016:DIS

[ZDGW16] Chun-Yu Zhang, Hang Ding, Peng Gao, and Yan-Ling Wu. Diffuse interface simulation of ternary fluids in contact with solid. Journal of Computational Physics, 309(??):37–51, March 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115008797.

Zhao:2015:CSD

[ZED15] B. B. Zhao, R. C. Ertekin, and W. Y. Duan. A comparative study of diffraction of shallow-water waves by high-level IGN and GN equations. Journal of Computational Physics, 283(??):129–147, February 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007785.

Zhang:2018:MFI

[ZF18] Lei Zhang and Lixin Feng. A multi-frequency iterative imaging method for discontinuous inverse medium problem. Journal of Computational Physics, 362(??):290–304, June 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118301086.

Zhang:2019:CSI

[ZFL^{+}19] Changjuan Zhang, Abbas Fakhari, Jie Li, Li-Shi Luo, and Tiezheng Qian. A comparative study of interfaceconforming ALE–FE scheme and diffuse interface AMR– LB scheme for interfacial dynamics. Journal of Computational Physics, 395(??):602–619, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119304565.

Zentner:2016:BDI

[ZFPB16] I. Zentner, G. Ferré, F. Poirion, and M. Benoit. A biorthogonal decomposition for the identification and simulation of non-stationary and non-Gaussian random fields. Journal of Computational Physics, 314(??):1–13, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911600139X.

Zhang:2015:CFT

[ZFZL15] Yuxian Zhang, Naixing Feng, Henry Hongxing Zheng, and Qing Huo Liu. A corner-free truncation strategy for FDTD method in target scattering. Journal of Computational Physics, 302(??):567–572, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005793.

Zhang:2017:CSV

[ZG17] Chenglong Zhang and Irene M. Gamba. A conservative scheme for Vlasov Poisson Landau modeling collisional plasmas. Journal of Computational Physics, 340(??):470–497, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117302474.

Zhang:2018:FDF

[ZG18a] Tie Zhang and Qingxin Guo. The finite difference/finite volume method for solving the fractional diffusion equation. Journal of Computational Physics, 375(??):120–134, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305564.

Zhang:2018:FDS

[ZG18b] Yabin Zhang and Adrianna Gillman. A fast direct solver for boundary value problems on locally perturbed geometries. Journal of Computational Physics, 356(??):356–371, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117308975.

Zoni:2019:SHE

[ZG19] Edoardo Zoni and Yaman Güçlü. Solving hyperbolic-elliptic problems on singular mapped disk-like domains with the method of characteristics and spline finite elements. Journal of Computational Physics, 398(??):Article 108889, December 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S002199911930587X.

Zheng:2016:AEA

[ZGD^{+}16] Chang-Jun Zheng, Hai-Feng Gao, Lei Du, Hai-Bo Chen, and Chuanzeng Zhang. An accurate and efficient acoustic eigensolver based on a fast multipole BEM and a contour integral method. Journal of Computational Physics, 305(??):677–699, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007251.

Zhang:2016:QRF

[ZGJ16] Xiaoping Zhang, Max Gunzburger, and Lili Ju. Quadrature rules for finite element approximations of 1D nonlocal problems. Journal of Computational Physics, 310(??):213–236, April 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116000176.

Zhen:2015:AEC

[ZH15] Yicun Zhen and John Harlim. Adaptive error covariances estimation methods for ensemble Kalman filters. Journal of Computational Physics, 294(??):619–638, August 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115002223.

Zhu:2019:FSP

[ZH19] Yujie Zhu and Xiangyu Hu. Free-stream preserving linearupwind and WENO schemes on curvilinear grids. Journal of Computational Physics, 399(??):Article 108907, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0021999119306126.

Zhang:2016:CBT

[Zha16] Xiangxiong Zhang. A curved boundary treatment for discontinuous Galerkin schemes solving time dependent problems. Journal of Computational Physics, 308(??):153–170, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115008554.

Zhang:2017:WCS

[ZHA17a] C. Zhang, X. Y. Hu, and N. A. Adams. A weakly compressible SPH method based on a low-dissipation Riemann solver. Journal of Computational Physics, 335(??):605–620, April 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117300438.

Zhang:2017:GTV

[ZHA17b] Chi Zhang, Xiangyu Y. Hu, and Nikolaus A. Adams. A generalized transport-velocity formulation for smoothed particle hydrodynamics. Journal of Computational Physics, 337(??):216– 232, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117301092.

Zhang:2017:PPH

[Zha17c] Xiangxiong Zhang. On positivity-preserving high order discontinuous Galerkin schemes for compressible Navier–Stokes equations. Journal of Computational Physics, 328(??):301–343, January 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116304958.

Zhang:2018:MMF

[ZHLZ18] Fei Zhang, Weizhang Huang, Xianping Li, and Shicheng Zhang. Moving mesh finite element simulation for phase-field modeling of brittle fracture and convergence of Newton’s iteration. Journal of Computational Physics, 356(??):127–149, March 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308719.

Zhang:2018:NSPa

[ZHS18] Yingnan Zhang, Xingbiao Hu, and Jianqing Sun. A numerical study of the 3-periodic wave solutions to KdV-type equations. Journal of Computational Physics, 355(??):566–581, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308653.

Zhu:2018:LSM

[ZHW18] Shengfeng Zhu, Xianliang Hu, and Qingbiao Wu. A level set method for shape optimization in semilinear elliptic problems. Journal of Computational Physics, 355(??):104–120, February 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117308355.

Zhao:2018:IFP

[ZHWQ18] Meng Zhao, Shuai He, Hong Wang, and Guan Qin. An integrated fractional partial differential equation and molecular dynamics model of anomalously diffusive transport in heterogeneous nano-pore structures. Journal of Computational Physics, 373(??):1000–1012, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300068.

Zillich:2015:CPD

[Zil15] Robert E. Zillich. Combination of the pair density approximation and the Takahashi–Imada approximation for path integral Monte Carlo simulations. Journal of Computational Physics, 301(??):111–118, November 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005471.

Zhang:2015:IFE

[ZILZ15] Qian Zhang, Kazufumi Ito, Zhilin Li, and Zhiyue Zhang. Immersed finite elements for optimal control problems of elliptic PDEs with interfaces. Journal of Computational Physics, 298(??):305–319, October 1, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115003903.

Zhang:2018:ABF

[ZJ18] Yaoxin Zhang and Yafei Jia. 2D automatic body-fitted structured mesh generation using advancing extraction method. Journal of Computational Physics, 353(??):316–335, January 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117307714.

Zhao:2016:PIM

[ZJL16] Zhi Zhao, Xiao-Qing Jin, and Matthew M. Lin. Preconditioned iterative methods for space–time fractional advection–diffusion equations. Journal of Computational Physics, 319(??):266–279, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301589.

Zhang:2015:RTS

[ZJLC15] Di Zhang, Chunbo Jiang, Dongfang Liang, and Liang Cheng. A review on TVD schemes and a refined fluxlimiter for steady-state calculations. Journal of Computational Physics, 302(??):114–154, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115005707.

Zhang:2015:CSI

[ZJS15] Shuangxi Zhang, Yuesong Jia, and Qizhi Sun. Comment on “Symplectic integration of magnetic systems” by Stephen D. Webb [J. Comput. Phys. 270 (2014) 570–576]. Journal of Computational Physics, 282(??):43–46, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007505. See [Web14].

Zayernouri:2015:FSC

[ZK15] Mohsen Zayernouri and George Em Karniadakis. Fractional spectral collocation methods for linear and nonlinear variable order FPDEs. Journal of Computational Physics, 293(??):312– 338, July 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114008067.

Zivcakova:2018:DGM

[ZK18] A. Zivcáková and V. Kucera. Discontinuous Galerkin method for a nonlocal hydrodynamic model of flocking dynamics. Journal of Computational Physics, 372(??):500–523, November 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118304339.

Zhang:2019:CPB

[ZKG19] Lei Zhang, Anela Kumbaro, and Jean-Michel Ghidaglia. A conservative pressure based solver with collocated variables on unstructured grids for two-fluid flows with phase change. Journal of Computational Physics, 390(??):265–289, August 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119302402.

Zhang:2015:FSM

[ZKS^{+}15] Bohai Zhang, Bledar A. Konomi, Huiyan Sang, Georgios Karagiannis, and Guang Lin. Full scale multi-output Gaussian process emulator with nonseparable auto-covariance functions. Journal of Computational Physics, 300(??):623–642, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005239.

Zhang:2015:ITD

[ZL15a] A. M. Zhang and Y. L. Liu. Improved three-dimensional bubble dynamics model based on boundary element method. Journal of Computational Physics, 294(??):208–223, August 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115002041.

Zhang:2015:SEH

[ZL15b] Bin Zhang and Chunlei Liang. A simple, efficient, and highorder accurate curved sliding-mesh interface approach to spectral difference method on coupled rotating and stationary domains. Journal of Computational Physics, 295(??):147–160, August 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115002466.

Zhou:2015:NTD

[ZL15c] Xuan Zhou and Shuixiang Li. A novel three-dimensional mesh deformation method based on sphere relaxation. Journal of Computational Physics, 298(??):320–336, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911500385X.

Zhao:2018:ALC

[ZLC^{+}18] Lifei Zhao, Zhen Li, Bruce Caswell, Jie Ouyang, and George Em Karniadakis. Active learning of constitutive relation from mesoscopic dynamics for macroscopic modeling of non-Newtonian flows. Journal of Computational Physics, 363(??):116–127, June 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118301220.

Zhang:2018:TOG

[ZLFW18] Chao Zhang, Qibing Li, Song Fu, and Z. J. Wang. A thirdorder gas-kinetic CPR method for the Euler and Navier– Stokes equations on triangular meshes. Journal of Computational Physics, 363(??):329–353, June 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118301232.

Zhang:2019:QTU

[ZLGK19] Dongkun Zhang, Lu Lu, Ling Guo, and George Em Karniadakis. Quantifying total uncertainty in physics-informed neural networks for solving forward and inverse stochastic problems. Journal of Computational Physics, 397(??):Article 108850, ???? 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119305340.

Zhou:2018:DMR

[ZLGS18] Di Zhou, Zhiliang Lu, Tongqing Guo, and Ennan Shen. Development of a moving reference frame-based gas-kinetic BGK scheme for viscous flows around arbitrarily moving bodies. Journal of Computational Physics, 373(??):698–721, November 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118304789.

Zhang:2017:ALB

[ZLH^{+}17] Gaigong Zhang, Lin Lin, Wei Hu, Chao Yang, and John E. Pask. Adaptive local basis set for Kohn–Sham density functional theory in a discontinuous Galerkin framework II: Force, vibration, and molecular dynamics calculations. Journal of Computational Physics, 335(??):426–443, April 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116307203.

Zhang:2016:APM

[ZLJ16] Bin Zhang, Hong Liu, and Shi Jin. An asymptotic preserving Monte Carlo method for the multispecies Boltzmann equation. Journal of Computational Physics, 305(??):575–588, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007408.

Zhang:2016:EMP

[ZLL16a] Peng Zhang, Hyunsuk Lee, and Deokjung Lee. Extension of modified power method to two-dimensional problems. Journal of Computational Physics, 320(??):17–32, September 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301619.

Zhang:2016:GSS

[ZLL16b] Peng Zhang, Hyunsuk Lee, and Deokjung Lee. A general solution strategy of modified power method for higher mode solutions. Journal of Computational Physics, 305(??):387–402, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007196.

Zhang:2017:ENP

[ZLL17a] Peng Zhang, Hyunsuk Lee, and Deokjung Lee. Extension of the noise propagation matrix method for higher mode solutions. Journal of Computational Physics, 344(??):440–450, September 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999117303765.

Zheng:2017:NST

[ZLL^{+}17b] Minling Zheng, Fawang Liu, Qingxia Liu, Kevin Burrage, and Matthew J. Simpson. Numerical solution of the time fractional reaction-diffusion equation with a moving boundary. Journal of Computational Physics, 338(??):493–510, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301936.

Zhu:2017:MFS

[ZLX17] Xueyu Zhu, Erin M. Linebarger, and Dongbin Xiu. Multifidelity stochastic collocation method for computation of statistical moments. Journal of Computational Physics, 341 (??):386–396, July 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999117302930.

Zhang:2015:FDM

[ZLY15] Longhui Zhang, Kai Liu, and Changfu You. Fictitious domain method for fully resolved reacting gas-solid flow simulation. Journal of Computational Physics, 299(??):215–228, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115004556.

Zayernouri:2016:FAB

[ZM16a] Mohsen Zayernouri and Anastasios Matzavinos. Fractional Adams–Bashforth/Moulton methods: an application to the fractional Keller–Segel chemotaxis system. Journal of Computational Physics, 317(??):1–14, July 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999116300870.

Zheng:2016:MCB

[ZM16b] Weixiong Zheng and Ryan G. McClarren. Moment closures based on minimizing the residual of the P_{N} angular expansion in radiation transport. Journal of Computational Physics, 314(??):682–699, June 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116001820.

Zhang:2018:GAI

[ZMCC18] Jia-Le Zhang, Zhi-Hua Ma, Hong-Quan Chen, and Cheng Cao. A GPU-accelerated implicit meshless method for compressible flows. Journal of Computational Physics, 360(??):39–56, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118300470.

Zandi:2015:SAA

[ZMF15] S. Zandi, A. Mohammadi, and J. M. Floryan. Spectrallyaccurate algorithm for the analysis of flows in two-dimensional vibrating channels. Journal of Computational Physics, 301(??):425–455, November 15, 2015. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115005525.

Zhao:2019:DTI

[ZMZC19] H. Y. Zhao, P. J. Ming, W. P. Zhang, and J. K. Chen. A direct time-integral THINC scheme for sharp interfaces. Journal of Computational Physics, 393(??):139–161, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303365.

Zwanenburg:2016:EBE

[ZN16] Philip Zwanenburg and Siva Nadarajah. Equivalence between the Energy Stable Flux Reconstruction and Filtered Discontinuous Galerkin Schemes. Journal of Computational Physics, 306(??):343–369, February 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115007767.

Zhang:2018:DNS

[ZN18] Jie Zhang and Ming-Jiu Ni. Direct numerical simulations of incompressible multiphase magnetohydrodynamics with phase change. Journal of Computational Physics, 375(??):717–746, December 15, 2018. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118305941.

Zepeda-Nunez:2016:MPT

[ZND16] Leonardo Zepeda-Núñez and Laurent Demanet. The method of polarized traces for the 2D Helmholtz equation. Journal of Computational Physics, 308(??):347–388, March 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007809.

Zamolo:2019:NMT

[ZNS19] Riccardo Zamolo, Enrico Nobile, and Bozidar Sarler. Novel multilevel techniques for convergence acceleration in the solution of systems of equations arising from RBF–FD meshless discretizations. Journal of Computational Physics, 392 (??):311–334, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999119303171.

Zhou:2015:WDL

[ZNX15] Tao Zhou, Akil Narayan, and Dongbin Xiu. Weighted discrete least-squares polynomial approximation using randomized quadratures. Journal of Computational Physics, 298 (??):787–800, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999115004404.

Zangeneh:2019:SAI

[ZOG19] Reza Zangeneh and Carl F. Ollivier-Gooch. Stability analysis and improvement of the solution reconstruction for cellcentered finite volume methods on unstructured meshes. Journal of Computational Physics, 393(??):375–405, September 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119303274.

Zohdi:2017:CME

[Zoh17] T. I. Zohdi. Computational modeling of electrically-driven deposition of ionized polydisperse particulate powder mixtures in advanced manufacturing processes. Journal of Computational Physics, 340(??):309–329, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302395.

Zhang:2019:MOD

[ZOY^{+}19] Qingfu Zhang, Houman Owhadi, Jun Yao, Florian Schäfer, Zhaoqin Huang, and Yang Li. Multiresolution operator decomposition for flow simulation in fractured porous media.

Journal of Computational Physics, 391(??):381–396, August 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119300130.

Zahr:2016:AMH

[ZP16] M. J. Zahr and P.-O. Persson. An adjoint method for a high-order discretization of deforming domain conservation laws for optimization of flow problems. Journal of Computational Physics, 326(??):516–543, December 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116304247.

Zimon:2016:ENR

[ZPE^{+}16] M. J. Zimoń, R. Prosser, D. R. Emerson, M. K. Borg, D. J. Bray, L. Grinberg, and J. M. Reese. An evaluation of noise reduction algorithms for particle-based fluid simulations in multi-scale applications. Journal of Computational Physics, 325(??):380–394, November 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116303709.

Zhao:2018:WEN

[ZPW18] Fengxiang Zhao, Liang Pan, and Shuanghu Wang. Weighted essentially non-oscillatory scheme on unstructured quadrilateral and triangular meshes for hyperbolic conservation laws. Journal of Computational Physics, 374(??):605–624, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305308.

Zheng:2016:DSH

[ZQ16a] Feng Zheng and Jianxian Qiu. Directly solving the Hamilton– Jacobi equations by Hermite WENO schemes. Journal of Computational Physics, 307(??):423–445, February 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008293.

Zhu:2016:NFO

[ZQ16b] Jun Zhu and Jianxian Qiu. A new fifth order finite difference WENO scheme for solving hyperbolic conservation laws. Journal of Computational Physics, 318(??):110–121, August 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116301371.

Zhu:2017:NTO

[ZQ17] Jun Zhu and Jianxian Qiu. A new third order finite volume weighted essentially non-oscillatory scheme on tetrahedral meshes. Journal of Computational Physics, 349(??):220–232, November 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305922.

Zhou:2015:LBS

[ZQCT15] L. Zhou, Z. G. Qu, L. Chen, and W. Q. Tao. Lattice Boltzmann simulation of gas-solid adsorption processes at pore scale level. Journal of Computational Physics, 300(??):800–813, November 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005331.

Ziegelwanger:2017:PMM

[ZR17] Harald Ziegelwanger and Paul Reiter. The PAC-MAN model: Benchmark case for linear acoustics in computational physics. Journal of Computational Physics, 346(??):152–171, October 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304631.

Zimon:2016:NCN

[ZRE16] M. J. Zimoń, J. M. Reese, and D. R. Emerson. A novel coupling of noise reduction algorithms for particle flow simulations. Journal of Computational Physics, 321(??):169–190, September 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116301942.

Zhong:2018:IBI

[ZRT18] Yimin Zhong, Kui Ren, and Richard Tsai. An implicit boundary integral method for computing electric potential of macromolecules in solvent. Journal of Computational Physics, 359(??):199–215, April 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118300317.

Zhang:2019:CHO

[ZRW19] Yu-Si Zhang, Yu-Xin Ren, and Qian Wang. Compact high order finite volume method on unstructured grids IV: Explicit multi-step reconstruction schemes on compact stencil. Journal of Computational Physics, 396(??):161–192, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304620.

Zhang:2015:EFE

[ZS15] Ruming Zhang and Jiguang Sun. Efficient finite element method for grating profile reconstruction. Journal of Computational Physics, 302(??):405–419, December 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006099.

Zeng:2016:VMF

[ZS16] X. Zeng and G. Scovazzi. A variational multiscale finite element method for monolithic ALE computations of shock hydrodynamics using nodal elements. Journal of Computational Physics, 315(??):577–608, June 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116300122.

Zhu:2017:NSC

[ZS17] Jun Zhu and Chi-Wang Shu. Numerical study on the convergence to steady state solutions of a new class of high order WENO schemes. Journal of Computational Physics, 349(??):80–96, November 15, 2017. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117305831.

Zhu:2018:NTM

[ZS18] Jun Zhu and Chi-Wang Shu. A new type of multi-resolution WENO schemes with increasingly higher order of accuracy. Journal of Computational Physics, 375(??):659–683, December 15, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305977.

Zhu:2019:NTM

[ZS19a] Jun Zhu and Chi-Wang Shu. A new type of multiresolution WENO schemes with increasingly higher order of accuracy on triangular meshes. Journal of Computational Physics, 392(??):19–33, September 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302736.

Zhuang:2019:ESA

[ZS19b] Qingqu Zhuang and Jie Shen. Efficient SAV approach for imaginary time gradient flows with applications to one- and multi-component Bose–Einstein condensates. Journal of Computational Physics, 396(??):72–88, November 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119304516.

Zhang:2019:CAR

[ZSL^{+}19] Binzheng Zhang, Kareem A. Sorathia, John G. Lyon, Viacheslav G. Merkin, and Michael Wiltberger. Conservative averaging-reconstruction techniques (ring average) for 3-D finite-volume MHD solvers with axis singularity. Journal of Computational Physics, 376(??):276–294, January 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999118305436.

Zhu:2019:GPS

[ZSM19] Chi Zhu, Jung-Hee Seo, and Rajat Mittal. A graphpartitioned sharp-interface immersed boundary solver for efficient solution of internal flows. Journal of Computational Physics, 386(??):37–46, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999119300968.

Zhao:2019:GFE

[ZSMP19] Guoyan Zhao, Mingbo Sun, Antonio Memmolo, and Sergio Pirozzoli. A general framework for the evaluation of shock-capturing schemes. Journal of Computational Physics, 376(??):924–936, January 1, 2019. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999118306703.

Zhiliakov:2019:HOA

[ZSO^{+}19] Alexander Zhiliakov, Daniil Svyatskiy, Maxim Olshanskii, Evgeny Kikinzon, and Mikhail Shashkov. A higher order approximate static condensation method for multimaterial diffusion problems. Journal of Computational Physics, 395(??):333–350, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119304528.

Zhang:2015:FNS

[ZSP15] Lu Zhang, Hai-Wei Sun, and Hong-Kui Pang. Fast numerical solution for fractional diffusion equations by exponential quadrature rule. Journal of Computational Physics, 299(??):130–143, October 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999115004416.

Zheng:2017:FDH

[ZSQ17] Feng Zheng, Chi-Wang Shu, and Jianxian Qiu. Finite difference Hermite WENO schemes for the Hamilton–Jacobi equations. Journal of Computational Physics, 337(??):27–41, May 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301262.

Zhang:2017:VCP

[ZSW17] Xiaoping Zhang, Shuai Su, and Jiming Wu. A vertex-centered and positivity-preserving scheme for anisotropic diffusion problems on arbitrary polygonal grids. Journal of Computational Physics, 344(??):419–436, September 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117303583.

Zhao:2017:VBC

[ZSX17] Ying Zhao, Dominik Schillinger, and Bai-Xiang Xu. Variational boundary conditions based on the Nitsche method for fitted and unfitted isogeometric discretizations of the mechanically coupled Cahn–Hilliard equation. Journal of Computational Physics, 340(??):177–199, July 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117302358.

Zhao:2017:RKD

[ZT17] Jian Zhao and Huazhong Tang. Runge–Kutta discontinuous Galerkin methods for the special relativistic magnetohydrodynamics. Journal of Computational Physics, 343(??):33–72, August 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911730298X.

Zeng:2019:DLS

[ZTBW19] Fanhai Zeng, Ian Turner, Kevin Burrage, and Stephen J. Wright. A discrete least squares collocation method for twodimensional nonlinear time-dependent partial differential equations. Journal of Computational Physics, 394(??):177–199, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303882.

Zhang:2016:CBR

[ZV16] Y. F. Zhang and R. Vicquelin. Controlling bulk Reynolds number and bulk temperature in channel flow simulations. Journal of Computational Physics, 305(??):208–216, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115007287.

Zhu:2018:FAM

[ZV18] Yuanran Zhu and Daniele Venturi. Faber approximation of the Mori–Zwanzig equation. Journal of Computational Physics, 372(??):694–718, November 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911830425X.

Zhao:2015:FNC

[ZVO15] Jing Zhao, Edwin A. H. Vollebregt, and Cornelis W. Oosterlee. A fast nonlinear conjugate gradient based method for 3D concentrated frictional contact problems. Journal of Computational Physics, 288(??):86–100, May 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000868.

Zhang:2015:SNC

[ZW15] Bin Zhang and Jian-Hang Wang. A short note on the counter-intuitive spurious behaviors in stiff reacting flow. Journal of Computational Physics, 291(??):52–59, June 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115001643.

Zhang:2016:PFM

[ZW16] Qian Zhang and Xiao-Ping Wang. Phase field modeling and simulation of three-phase flow on solid surfaces. Journal of Computational Physics, 319(??):79–107, August 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911630153X.

Zhao:2019:SOF

[ZW19] Xueping Zhao and Qi Wang. A second order fully-discrete linear energy stable scheme for a binary compressible viscous fluid model. Journal of Computational Physics, 395(??):382–409, October 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119304383.

Zhu:2018:SDI

[ZWB^{+}18] Wei Zhu, Bao Wang, Richard Barnard, Cory D. Hauck, Frank Jenko, and Stanley Osher. Scientific data interpolation with low dimensional manifold model. Journal of Computational Physics, 352(??):213–245, January 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117307088.

Zhu:2017:PEG

[ZWG17] Lianhua Zhu, Peng Wang, and Zhaoli Guo. Performance evaluation of the general characteristics based off-lattice Boltzmann scheme and DUGKS for low speed continuum flows. Journal of Computational Physics, 333(??):227–246, March 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116307057.

Zhang:2019:ACS

[ZWL^{+}19] Zijian Zhang, Chihyung Wen, Yunfeng Liu, Deliang Zhang, and Zonglin Jiang. Application of CE/SE method to gas-particle two-phase detonations under an Eulerian–Lagrangian framework. Journal of Computational Physics, 394(??):18–40, October 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119303560.

Zimmerling:2016:LMO

[ZWUR16] Jörn Zimmerling, Lei Wei, Paul Urbach, and Rob Remis. A Lanczos model-order reduction technique to efficiently simulate electromagnetic wave propagation in dispersive media. Journal of Computational Physics, 315(??):348–362, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116300225.

Zhou:2018:SON

[ZWYW18] Shenggao Zhou, Yu Wang, Xingye Yue, and Cheng Wang. A second order numerical scheme for the annealing of metalintermetallic laminate composite: a ternary reaction system. Journal of Computational Physics, 374(??):1044–1060, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118305011.

Zhang:2019:DGM

[ZX19] Qian Zhang and Yinhua Xia. Discontinuous Galerkin methods for short pulse type equations via hodograph transformations.

Journal of Computational Physics, 399(??):Article 108928, December 15, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0021999119306333.

Zou:2017:SFT

[ZXDL17] Dongyang Zou, Chunguang Xu, Haibo Dong, and Jun Liu. A shock-fitting technique for cell-centered finite volume methods on unstructured dynamic meshes. Journal of Computational Physics, 345(??):866–882, September 15, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999117304400.

Zhou:2017:SFF

[ZXL17] Guangzhao Zhou, Kun Xu, and Feng Liu. Simplification of the flux function for a high-order gas-kinetic evolution model. Journal of Computational Physics, 339(??):146–162, June 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117302188.

Zhang:2019:WCS

[ZXW^{+}19] C. Zhang, G. M. Xiang, B. Wang, X. Y. Hu, and N. A. Adams. A weakly compressible SPH method with WENO reconstruction. Journal of Computational Physics, 392(??):1–18, September 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119302840.

Zhao:2017:SNS

[ZY17] Weifeng Zhao and Wen-An Yong. Single-node second-order boundary schemes for the lattice Boltzmann method. Journal of Computational Physics, 329(??):1–15, January 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116305575.

Zhang:2019:HOI

[ZY19] Jiaqi Zhang and Pengtao Yue. A high-order and interfacepreserving discontinuous Galerkin method for level-set reinitialization. Journal of Computational Physics, 378(??):634– 664, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999118307733.

Zheng:2015:PDD

[ZYCK15] Xiang Zheng, Chao Yang, Xiao-Chuan Cai, and David Keyes. A parallel domain decomposition-based implicit method for the Cahn–Hilliard–Cook phase-field equation in 3D. Journal of Computational Physics, 285(??):55–70, March 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000200.

Zhao:2019:ALE

[ZYD^{+}19a] Xiaolong Zhao, Xijun Yu, Maochang Duan, Fang Qing, and Shijun Zou. An arbitrary Lagrangian–Eulerian RKDG method for compressible Euler equations on unstructured meshes: Single-material flow. Journal of Computational Physics, 396(??):451–469, November 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL https://www.sciencedirect.com/science/article/
pii/S0021999119305005.

Zou:2019:RKD

[ZYD19b] Shijun Zou, Xijun Yu, and Zihuan Dai. A Runge–Kutta discontinuous Galerkin method for Lagrangian ideal magnetohydrodynamics equations in two-dimensions. Journal of Computational Physics, 386(??):384–404, June 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119301366.

Zhang:2018:BDC

[ZYK18] Dongkun Zhang, Liu Yang, and George Em Karniadakis. Bidirectional coupling between a PDE-domain and an adjacent data-domain equipped with multi-fidelity sensors. Journal of Computational Physics, 374(??):121–134, December 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830500X.

Zhao:2016:DES

[ZYSW16] Jia Zhao, Xiaofeng Yang, Jie Shen, and Qi Wang. A decoupled energy stable scheme for a hydrodynamic phase-field model of mixtures of nematic liquid crystals and viscous fluids. Journal of Computational Physics, 305(??):539–556, January 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115006439.

Zhang:2016:LBS

[ZYW16] Jianying Zhang, Guangwu Yan, and Moran Wang. Lattice Boltzmann simulations for the vortex tori pattern in the three-dimensional cubic-quintic complex Ginzburg–Landau equation. Journal of Computational Physics, 306(??):311–319, February 1, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115007792.

Zhang:2017:RSO

[ZZ17a] Bo Zhang and Haiwen Zhang. Recovering scattering obstacles by multi-frequency phaseless far-field data. Journal of Computational Physics, 345(??):58–73, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117303972.

Zhang:2017:NIT

[ZZ17b] Hong Zhang and Paul Andries Zegeling. Numerical investigations of two-phase flow with dynamic capillary pressure in porous media via a moving mesh method. Journal of Computational Physics, 345(??):510–527, September 15, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117304242.

Zhu:2018:BDC

[ZZ18] Yinhao Zhu and Nicholas Zabaras. Bayesian deep convolutional encoder-decoder networks for surrogate modeling and uncertainty quantification. Journal of Computational Physics, 366(??):415–447, August 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302341.

Zhao:2019:FPI

[ZZ19] Meiling Zhao and Na Zhu. A fast preconditioned iterative method for the electromagnetic scattering by multiple cavities with high wave numbers. Journal of Computational Physics, 398(??):Article 108826, December 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0021999119305108.

Zhang:2015:MTS

[ZZDB15] Peng Zhang, Na Zhang, Yuefan Deng, and Danny Bluestein. A multiple time stepping algorithm for efficient multiscale modeling of platelets flowing in blood plasma. Journal of Computational Physics, 284(??):668–686, March 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000054.

Zhong:2016:CCF

[ZZH16] Siyang Zhong, Xin Zhang, and Xun Huang. A controllable canonical form implementation of time domain impedance boundary conditions for broadband aeroacoustic computation. Journal of Computational Physics, 313(??):713–725, May 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999116001467.

Zeng:2016:FDS

[ZZK16] Fanhai Zeng, Zhongqiang Zhang, and George Em Karniadakis. Fast difference schemes for solving high-dimensional time-fractional subdiffusion equations. Journal of Computational Physics, 307(??):15–33, February 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115008062.

Zheng:2015:NID

[ZZKF15] Wen Zheng, Bo Zhu, Byungmoon Kim, and Ronald Fedkiw. A new incompressibility discretization for a hybrid particle MAC grid representation with surface tension. Journal of Computational Physics, 280(??):96–142, January 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006421.

Zhu:2019:PCD

[ZZKP19] Yinhao Zhu, Nicholas Zabaras, Phaedon-Stelios Koutsourelakis, and Paris Perdikaris. Physics-constrained deep learning for high-dimensional surrogate modeling and uncertainty quantification without labeled data. Journal of Computational Physics, 394(??):56–81, October 1, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119303559.

Zhang:2019:LBM

[ZZL19] Mengxin Zhang, Weifeng Zhao, and Ping Lin. Lattice Boltzmann method for general convection-diffusion equations: MRT model and boundary schemes. Journal of Computational Physics, 389(??):147–163, July 15, 2019. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999119302281.

Zhang:2018:NSPc

[ZZPH18a] Ya Zhang, Yonghao Zhang, Guang Pan, and Sina Haeri. Numerical study of the particle sedimentation in a viscous fluid using a coupled DEM–IB–CLBM approach. Journal of Computational Physics, 368(??):1–20, September 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302821.

Zhang:2018:NSPb

[ZZPH18b] Ya Zhang, Yonghao Zhang, Guang Pan, and Sina Haeri. Numerical study of the particle sedimentation in a viscous fluid using a coupled DEM-IB-CLBM approach. Journal of Computational Physics, 368(??):1–20, September 1, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999118302821.

Zhang:2017:IMP

[ZZS^{+}17] Fan Zhang, Xiong Zhang, Kam Yim Sze, Yanping Lian, and Yan Liu. Incompressible material point method for free surface flow. Journal of Computational Physics, 330(??):92–110, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999116305721.

Zhao:2015:SOA

[ZzSK15] Xuan Zhao, Zhi zhong Sun, and George Em Karniadakis. Second-order approximations for variable order fractional derivatives: Algorithms and applications. Journal of Computational Physics, 293(??):184–200, July 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005610.

Zhu:2016:SSM

[ZZT^{+}16] Beibei Zhu, Ruili Zhang, Yifa Tang, Xiongbiao Tu, and Yue Zhao. Splitting K-symplectic methods for non-canonical separable Hamiltonian problems. Journal of Computational Physics, 322(??):387–399, October 1, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999116302698.

Zheng:2016:MLR

[ZZW^{+}16] Hui Zheng, Chuanzeng Zhang, Yuesheng Wang, Jan Sladek, and Vladimir Sladek. A meshfree local RBF collocation method for anti-plane transverse elastic wave propagation analysis in 2D phononic crystals. Journal of Computational Physics, 305(??):997–1014, January 15, 2016. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115006841.

Zhu:2016:IUG

[ZZX16] Yajun Zhu, Chengwen Zhong, and Kun Xu. Implicit unified gas-kinetic scheme for steady state solutions in all flow regimes. Journal of Computational Physics, 315(??):16–38, June 15, 2016. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911600190X.

Zhu:2019:IUG

[ZZX19] Yajun Zhu, Chengwen Zhong, and Kun Xu. An implicit unified gas-kinetic scheme for unsteady flow in all Knudsen regimes. Journal of Computational Physics, 386(??):190–217, June 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999119300841.

Zhang:2019:NVE

[ZZYC19] Bei Zhang, Jikun Zhao, Yongqin Yang, and Shaochun Chen. The nonconforming virtual element method for elasticity problems. Journal of Computational Physics, 378(??):394–410, February 1, 2019. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999118307228.

Zhang:2017:NHO

[ZZZ17] Wensheng Zhang, Yuan Zhuang, and Lina Zhang. A new high-order finite volume method for 3D elastic wave simulation on unstructured meshes. Journal of Computational Physics, 340(??):534–555, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117302516.